A Complete Bibliography of Publications in Limnology and Oceanography: 2010–2019

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254

 E-mail: beebe@math.utah.edu, beebe@acm.org, beebe@computer.org (Internet)
            WWW URL: https://www.math.utah.edu/~beebe/

23 April 2025 Version 1.09

Title word cross-reference

3 [CLB19, WP14]. ^{+} [KRR16]. ^{0} [Kus14]. ^{13} [Bre14, CHL^{+}17, CKCEP10, DPM18, HBZ12, KPJ12, MMXC15, MMPSB14, OEMB10, OEM12, ORGE16, WLG^{+}16, WCCP14, dKYH^{+}12]. ^{14} [KPJ12, MMXC15, WLG^{+}16]. ^{15} [BCF^{+}17, BLJ13, CHL^{+}17, DLP13, DLBF17, EOM16, JSH12, JTH^{+}13, OHKC^{+}12, WE19, ZLLM10]. ^{16} [SSS^{+}16]. ^{18} [FYVU17, HH14, VHR^{+}11]. ^{223} [HGvB^{+}13]. ^{224} [HGvB^{+}13, SBNC^{+}19]. ^{226} [HGvB^{+}13]. ^{228} [HGvB^{+}13, SBNC^{+}19]. ^{56} [WGRS^{+}17]. ^{st} [MTU18]. ^{th} [BLS^{+}16, SSM^{+}19]. _{1} [FLLH18, KSWFG13]. _{12} [BMM^{+}13, BWP^{+}10, KMP^{+}11, KSWFG13]. _{2} [APB^{+}17, AMB^{+}11, BR17, BDP^{+}19, BOT^{+}15, BPL^{+}19b, CCV^{+}18, CHHT18, CvHB^{+}18, CGP^{+}19, CESC13, CWHP14, DSM^{+}18, EO13, FVSL19, HST^{+}14, HLSW^{+}15, HRPW15, HXS^{+}10, HH14, HCAF18, HCC^{+}13, HCL^{+}18, KRR16, LWrDM^{+}12, LWWE^{+}18, Man10, MCGF^{+}11, MSR16, MQJG13, MRH^{+}15, NTK^{+}18, NWT^{+}19, OLC18, PMY19a, PSG^{+}16, QFH18, RR12, RMH^{+}17, RHMSE15, SSU^{+}16, SYW18, SHF^{+}11, TJJ^{+}15, VSdG17, VTH^{+}18, WYL16, WFK^{+}16, YH17, ZCZ^{+}18, ZCY^{+}15]. _{3} [KRR16]. _{4} [CKB^{+}16, NSG^{+}16, PMY19a, RMH^{+}17]. _{CO2} [SKJD^{+}14]. _{O2} [HH14]. ≈ [PE16a]. β [YLJ11]. : [BGB^{+}14, FCC11, HSB^{+}13, KK13, MQJG13, PFH^{+}17, SGCI14, THA17, UA10, WC17, YJO^{+}19]. Δ [KPJ12, BCF^{+}17, BLJ13, CHL^{+}17, DLP13, DLBF17, DPM18, FYVU17, HH14, HBZ12, JSH12, MMXC15, VHR^{+}11, WE19, WGRS^{+}17]. S_{275−−295} [FB12]. ζ [YLJ11].

-carotene [YLJ11]. -driven [KRR16]. -induced [MSR16]. -sink [CCV^{+}18].

1-m [vH19]. 10.1002/lno.10504 [Ano21a].

2 [KPSW10]. 2007 [PCY^{+}10]. 2010 [SWD^{+}14]. 2011 [CWHP14, KJKS18]. 227 [PSH^{+}11]. 28-year [EMH12].

30-yr [Scu16].

48-meter [SLS^{+}11].

5 [MLK11]. 5-bisphosphate [nVOH12]. 5-year [SGRB10].

62 [Ano21a]. 64 [Ano21b].

abatement [AEH19]. aberrant [FDP^{+}18]. Abiotic [ASR^{+}17, DDF^{+}10, RZW11, uGH^{+}11]. above [MMC^{+}10, WMM18]. Absence [CKP^{+}15, KCB^{+}17]. absorption [DVC^{+}17, EM13, EBMR12, LWB^{+}17, PE13, RDT^{+}14, TW10b, XSAHV13, ZD18]. Abundance [PS17, RCJ15, SDMK10, AAIA14a, AAIA14b, AdBVA10, BBCM^{+}13, BSRP^{+}12, DDH^{+}19, DBFL11, GOD^{+}18, HPS^{+}10a, LG16, MRSS12, MPSA17, MCYR17, MBBG^{+}12, PRL18, Piw19, SAPI14, SJ11, SDCF16, SSFR19, WMP^{+}19, YYMN13]. abundances [RG13]. abundant [LZR^{+}17, LCW^{+}17b, MH16]. Abyssal [vOSH12, DRP^{+}17, SSH^{+}14, SSS^{+}19, WRS13]. acantharian [MAC^{+}10]. Acartia [DHK11, JLG10, JLG11, TW10b]. accelerated [HBD^{+}11, ZNVF16]. accelerates [VSD10]. access [LFC17]. acclimation [BG10b, GPL11, KRR16, XNK18]. acclimatory [SGME11]. accomplishments [APS^{+}19]. accretion [AC17, LSD18]. accumulates [YLJ11]. accumulation [BPRG^{+}18, BSC^{+}15, CF13b, GKS12, KNL10, MMHT10, MFL11, STCS10, SM11a, SGVR16, SFLB16, VF10, WFR10]. accurate [CSGW18]. acid [BB10, BISZ17, CWF11, CHL^{+}17, DBC^{+}13, GLS^{+}13, HHW^{+}19, HSTK15, JTV^{+}16, KNA^{+}14, KBVW12, LWWE^{+}18, MVL^{+}10, MMXC15, MGW^{+}13, NBSMN19, SKLG10, SW11, SHF^{+}11, WTC^{+}17, dKYH^{+}12]. acid-based [BISZ17]. acid-producing [HHW^{+}19]. acidic [CCC10, RSJ^{+}18, RHMSE15, SHD^{+}11]. acidification [BRS11, BR17, BHW^{+}12, BG10a, BWD^{+}11, BWD^{+}12, CESC14, CHPH13, CSME13, Edm11, FCC11, GWB^{+}14, HVJ^{+}19, HRG^{+}15, KBHT19, KRR16, LABJ18, Man10, MLGZ16, MSR16, NBDM16, RSTS^{+}18, SLC18, SPTS15, TIN^{+}14, TBSR13, VFS^{+}15, WCS^{+}18, WGH^{+}16, WCI^{+}14, XFH14, ZBSR15, ZHG15]. acidified [WBB^{+}17]. acidity [KKP^{+}19, NWT^{+}19]. acids [CPPdAR^{+}13, GBB19b, HOD^{+}17, IWF19, KMF10, LWrDM^{+}12, MTEM15, MKK15, NCT^{+}15, SRCL^{+}13, TEGL11, TAV^{+}10, UVGS10]. Acoustic [PGB^{+}19, BM16, SWL11]. acoustics [MPM^{+}15]. acquisition [Ano19c, BSCC15, GBB^{+}18, KP13, TBSR13]. across [ARML10, BRM^{+}19, BYD19, BTC^{+}19, BGB^{+}14, CMK^{+}10, DJS18, FZL^{+}14, FWO^{+}18, FOT^{+}15, GLI^{+}15, GBC^{+}17, GRSD^{+}14, HRMD19, JKKM13, JWGH19, KEH^{+}14, LLH^{+}15, LPLH18, LG16, LHS19, MLS^{+}18, MMWR17, MHPW18, NTK^{+}18, NHS^{+}12, PMY^{+}19b, PSS^{+}14, PFH^{+}17, PS17, RAB^{+}17, SES18, SSG^{+}17, SLK^{+}14, SAPI14, SWP11, SvKP^{+}18, SSM^{+}19, WMT^{+}12, WVL^{+}18, WMP^{+}19, XDK^{+}17, YP18, ZHN^{+}10]. acrylate [TKK^{+}17]. act [HCD19]. action [DMSHC16]. activation [WLO^{+}19]. Active [BSSW11, HPL11, LSK11, JKKM13, PK14, SWM^{+}10, WVV^{+}11, WMC^{+}18]. activities [AFSM17, FCD12, Ker17, VPWW10]. activity [BB11, BBK^{+}15, DM17, DDK10, DBV^{+}11, GFPSG13, GBC^{+}17, HGD14, HBB^{+}11, HPS10b, LDY^{+}16, LC11, MLS^{+}18, MGS12, PSNE15, PSZ^{+}13, RGGL^{+}12, RGLM^{+}12, SPP10, SDCF16, SBH^{+}11, SVG^{+}18, SDMK10, UMHH^{+}14, WYW^{+}10, ZMS^{+}18, ZXL^{+}19]. aculeatus [KKHP14]. acuminata [HLG15]. acute [WGH^{+}16]. acutus [PT11]. adaptation [JLC^{+}15, NCT^{+}15, PGRR^{+}19, SASB^{+}15, Sch19]. adaptations [MNW^{+}19, MCT^{+}14, SWP11]. adapting [SSP^{+}18]. Adaptive [KSTA18a, SMH^{+}11, MAV^{+}13, Rie15]. Adding [MHRH11]. addition [EED10, GMS^{+}18, JWS15, WCCP14]. additions [JTH^{+}13, KRB^{+}18, OHKC^{+}12, SCF^{+}15]. Additive [MMBP18]. adds [WGM16]. adenosine [MLK11]. adenosine-5’-triphosphate [MLK11]. adequately [SP11]. adhesion [CGB^{+}18]. adjacent [ZYZ19]. adjust [PT11]. adjustment [AFG^{+}16]. Adriatic [IGP^{+}12]. adrift [vHOM^{+}19]. adsorption [LK15, RSM13]. adult [HBCK10, LDCT11, SGCI14, TDF^{+}17]. Advancing [PE16b, MWR17]. advection [MFM^{+}12, RPI^{+}12, WMC^{+}18]. Advective [SKK^{+}13]. aeolian [CCV^{+}18]. aeolian-dust [CCV^{+}18]. aerial [KDGL19]. Aerobic [OMB^{+}16, BNW^{+}14b, CG17, TLR^{+}13]. aerosol [LYH17]. aeruginosa [BVSM15, FDBW16, HL13, WKK^{+}11]. affect [GPCJ16, HSTK15, KHH19, LdlSB^{+}12, LBHS13, PMP^{+}12, RLC^{+}11]. affected [BK13, DPG^{+}12, MVNG11, SNO^{+}16]. affecting [OrIA10, PHJ12, PMA18]. affects [DRE^{+}10, HAL17, LUM15, NA17, TIN^{+}14, VBC^{+}12]. Africa [GNHGM13, CMK^{+}10, JAZ^{+}10, MRSE14, MRC^{+}16]. African [RDB^{+}18]. after [BAA^{+}13, KHPIP^{+}14, SCF^{+}15, SVG^{+}18, SYW18]. Against [OPA^{+}14, KGRV18, KJKS18, KMF10, SBFC18]. age [MMXC15]. aged [XZGW17]. agent [RBRH10]. aggregate [TGC^{+}10, vdJFS^{+}18]. aggregates [BVvB^{+}19, LKK13, MTSG18]. Aggregation [PDFS14, vdJFS^{+}18, DMN15, GMJW13, MPM^{+}15, SLG^{+}14]. Agricultural [GTR^{+}13, CBK18, DVSV13, KGvdH16]. agriculturally [KGRV18]. agriculture [BLS^{+}16]. agro [FPG11]. agro-urban [FPG11]. aids [Sch19]. Air [RHSD^{+}10, HEBS10, Kus14, PHL^{+}18, RWB^{+}19, TPM^{+}14, ZCZ^{+}18]. air-sea [HEBS10, ZCZ^{+}18]. Air-water [RHSD^{+}10]. akashiwo [HMD11]. al [Ano21b, LGR^{+}12, MLS^{+}14, Ano21a]. al. [ACC^{+}19, CJC^{+}12]. Alanine [DLBF17]. alarm [SBDS^{+}15]. alarm-cues [SBDS^{+}15]. Alaska [AJG13, BLLB12, HXS^{+}10, KMP^{+}11, VHV10]. Alaskan [GSBR11, LKS^{+}16, MGJH18]. Albula [HCD19]. alewife [BRT^{+}10]. alewife-mysid [BRT^{+}10]. Alexandrium [BVSR^{+}15, BRF^{+}17, CBS^{+}17, HLSW^{+}15, MTH^{+}11, SFWP12]. alga [GHSR^{+}16, HCL^{+}18, KSWFG13, SMLC^{+}18, VFS^{+}15, YLJ11]. algae [CHPH13, EM13, GK15, HZC^{+}13, KG18, KGL^{+}16, LCS^{+}19, PBA^{+}15, RKLH11, RBD18, WZR19]. algae-produced [KGL^{+}16]. Algal [JTV^{+}16, OSHS19, AMNU16, BH13, CFB14, DMS^{+}18, DBMP^{+}11, HMD11, KG18, KIH^{+}15, OCB^{+}18, RKBA14, RPH^{+}10, SBM^{+}15, SLA^{+}18, SMN^{+}15, SS12b, SS12c, TDS^{+}10, THA17, TWWY18, VHrI10, WZBW^{+}11, ZD18]. aliases [SSC^{+}17]. alien [PSS^{+}14]. alignment [CFD^{+}19]. Alkaline [DDK10, GFPSG13, DM17, DBV^{+}11, LDY^{+}16, MLS^{+}18, RSJ^{+}18]. Alkalinity [MMG16, FEW^{+}14, KLEH16, MSR16, VPG^{+}19, YKT^{+}15]. allelochemical [WRO^{+}11]. allelochemicals [FPSL18]. allelopathically [KG18]. Alleviation [HD19]. allocation [BJF18, CLWD13, MZB^{+}15]. allochthonous [GMS^{+}18, GBP^{+}12, HBR^{+}14, HDDH^{+}17, RMF11, TTV^{+}13]. Allometric [ETKL12, ETKL15]. allows [LFH^{+}12]. ALOHA [DBH^{+}16, BDK^{+}17, GWB^{+}14, MGK15, MG17]. alone [MCLT15]. along [AC17, BSM17, BWS^{+}14, CJS^{+}17, CAS^{+}17, FEC^{+}16, HS10, HHS^{+}18, HSP^{+}16, HMFF12, JCF^{+}10, KH16, LZR^{+}17, LCBC16, LV16, MvdPK^{+}15, MFL11, MMD18, MCGF^{+}11, RBM14, RLL^{+}10, SPGRP^{+}17, SBH^{+}11, TSSH19, VdRA^{+}19, WWC^{+}13, WC17, WWS11, WSUC^{+}18, ZZW16, ZTW^{+}11, dGCB^{+}11]. Alpine [BNW^{+}14b, FUS^{+}16, BNW^{+}14a, FBFR13, GDCM13, GPH^{+}13, GSB^{+}17, ITO^{+}17, KWRS13, MMGP^{+}12, PJUR15, SFMF15, SS12a, WBZ^{+}13]. alter [BVvB^{+}19, GSPM13, WBB^{+}17]. alteration [SSC^{+}10]. alterations [FAF^{+}12]. altered [FBFR13, NWT^{+}19, WHH^{+}11]. alternates [BSN^{+}14]. alternative [ZHD^{+}16]. alterniflora [TMH^{+}18]. alters [BH13, GTR^{+}13, GSB^{+}17, NPT11, Rie15, SCP^{+}16]. altitude [ZZY^{+}10]. altitudinal [LV16]. aluminum [MdBKL13]. Amazon [AHD^{+}18, BMF^{+}16, CSS^{+}16, ERA^{+}12, LWWC^{+}16, WCC^{+}17, ZKMT^{+}13]. Amazonia [CKCEP10]. Amazonian [APB^{+}17]. ambient [HWZ13, HMHI13]. ambush [XNK18]. ameliorates [Edm11]. America [BHC13, BGW^{+}15, PSS^{+}14]. American [HLG15, MBBG^{+}12, RLPL14, RPL16]. amino [CHL^{+}17, GBB19b, HOD^{+}17, KMF10, LWrDM^{+}12, LWWE^{+}18, MTEM15, MGW^{+}13, NCT^{+}15, SKLG10, SRCL^{+}13, TAV^{+}10, UVGS10, WTC^{+}17]. Ammonia [UMHH^{+}14, AMMH^{+}13, Ano10, BPA12, CMB10, HQB^{+}18, MACM11, MBP^{+}17, NFW13, PWS^{+}11, SDCF16, VFME18]. ammonia-oxidizer [NFW13]. ammonia-oxidizing [AMMH^{+}13, BPA12, MACM11, PWS^{+}11, SDCF16, VFME18]. Ammonium [MFK^{+}13, PPPA14, DBFL11, DSS^{+}11, GWD^{+}16, JAZ^{+}10, KCB^{+}17, MAS^{+}16, RRB^{+}16, RDB^{+}18, STCS10, SMR^{+}17, SSGL19, WBZ^{+}13]. ammonium-oxidizing [JAZ^{+}10]. among [AJG13, BWD^{+}11, BWD^{+}12, CHL^{+}17, CMS^{+}18, FWS^{+}14, HS11, KP13, MRB11, NG13, RCIB14, SWP11]. amounts [CHV^{+}17, YLJ11]. amplifies [RBG^{+}10]. Amundsen [PKB^{+}17, SSFF12, STC^{+}11]. Amyot [SH10a]. anadromous [CBP10, TWP13]. Anaerobic [BK11, DMMV15, àNTS13, TSB^{+}19, WBZ^{+}13, JAZ^{+}10, RDB^{+}18, RETS16, SSS^{+}16, SAP^{+}11, TSDF^{+}16]. analyses [CHL^{+}17, KGL^{+}16, KWB^{+}16, MGHS18, MH16, SNTK15]. Analysis [ABS^{+}19, GK14, AWG^{+}12, BSCG17, BPPF12, BRS18, BSC^{+}15, CS12, CVS^{+}10, Edm15, FBV11, FPP^{+}19, GCH^{+}18, GAK^{+}19, HOD^{+}17, KHCH14, KFP^{+}18, LGW^{+}19, MBB^{+}18, MJJMM17, PE13, RKLH11, RMNZ12, SCQ^{+}17, UIY^{+}11, WVL^{+}18, YLJ11]. analytical [RGM^{+}11]. Anammox [DTFR12, BGR14, DT16, DSS^{+}11, JAZ^{+}10, NTK^{+}18, WBZ^{+}13]. Ancient [GBS17, FHR^{+}15, HMO^{+}18, KPW^{+}11, SHL^{+}18]. anemone [HRPW15]. anemophilous [MBK^{+}11]. Angola [NLO^{+}12, WMBR13]. angustifolia [KOFN11]. animals [Hir12, MHT13, OCR10]. Annecy [PDER10]. Annual [RGGL^{+}12, CB19, GGPM^{+}10, GMD11, MGJH18, NSG^{+}16, RPL16, Scu16]. annularis [Edm15]. anomaly [PNR19]. anophagefferens [KG18, KSWFG13]. anoxia [AMQ^{+}11, BKA^{+}14, GMBL16, MGW^{+}13]. anoxic [BNW^{+}14a, BNW^{+}14b, BKD^{+}16, CHHT18, GMS^{+}18, KPJ12, NCT^{+}14, OMB^{+}16]. anoxygenic [MRC^{+}16]. antagonistic [ŠSP17]. Antarctic [HVJ^{+}19, MMD18, RVvdP^{+}17, TSSH19, ADS^{+}17, BPV^{+}19, DTKMK15, GAK^{+}19, HGD14, MKLKP16, PHB^{+}10, PMA18, RNK^{+}16, RHDTS^{+}11, STCS10, SAS^{+}11, SAPI14, SJ11, SPO^{+}18, SDMK10, TT14, TBSR13, VCM13, VML^{+}19, XFH14]. Antarctica [SWD^{+}14, BIM^{+}16, DKK^{+}14, FYC^{+}18, MWR17, PKB^{+}17, SSS^{+}16, SMA15, VMAS^{+}16, ZCZ^{+}18, KBHT19, MEM^{+}17]. antecedent [KHH19]. Antenucci [PHJ12]. Anthropogenic [BBJ^{+}19, HSC^{+}14, MH16, RKG^{+}11, LYH17, MMGO^{+}17a, MMGO^{+}17b, SLK^{+}10, TZD^{+}15]. antioxidant [DJD^{+}14, HKS^{+}15]. antipredation [GMJW13]. antipredatory [KS13]. Apopka [SLK^{+}10]. Appalachian [REDW10, VB17]. Apparent [KBT16, EMB12, GBK^{+}18, SMLC^{+}18]. appearance [OLF^{+}11, SGH12]. appendicularian [LTPK^{+}18, LBR^{+}13, LEG^{+}10, LSK11, NTI^{+}15, WBB^{+}17]. appendicularian-ciliate [LEG^{+}10]. appendicularians [CGB^{+}18, LRS^{+}10]. Application [GBMG12, JSB^{+}14, SRAB10, CJHR19]. applied [BRR^{+}13, GBL13]. approach [BBS^{+}18, BRT^{+}10, DHW11, DWDH10, GBL13, GRE^{+}16, HSLH^{+}14, HMFB16, HESU13, MW15, RKBA14, RBY^{+}17, RBI^{+}10, SAH^{+}19, Spi15, TRD^{+}14, YWL^{+}17]. approaches [KWF^{+}17, MHT13]. aq [WYL16]. Aqaba [WSB^{+}13]. aquaculture [DCRC16, MAS^{+}16]. Aquatic [HCK10, SDS^{+}16, AP12, AGCA16, AGMR14, GMMV19, GBS17, GN16, HSCM19, HS10, HBR13, HSBA10, JLC^{+}15, KS16, KB15, KWF^{+}17, LN11, MKW^{+}19, NBG17, PvEF12, PWWF18, RDT^{+}14, SSU^{+}16, SM11b, TBHM^{+}13, TCFP19, WLO^{+}19, WC17, XDK^{+}17]. aquaticus [FA10]. aquifer [MBH^{+}15, YMB^{+}18]. Arabian [CRJ^{+}14, MVG^{+}15, TRD^{+}14]. Aragonite [LGC13a, LGC13b, NLHAA^{+}17]. archaea [BPA12]. archaeal [CKCEP10, GGPM^{+}10, UMHH^{+}14]. archaeon [AMMH^{+}13]. archipelago [PMA18, TBK15]. archive [VKC18]. archives [FMGR^{+}11]. arctia [RF13]. Arctic [AMQ^{+}11, BAG^{+}17, PvDM^{+}13, SSFF12, ÅCA^{+}18, ABD^{+}17, AMNU16, BSR^{+}17, BPW^{+}19, BAY^{+}14, CMS17, DKG15, DHG^{+}17, DHZ^{+}19, DL11, EM13, FNSS15, FDS^{+}18, GLKK10, GKT^{+}15, GVS^{+}10, HNHS^{+}15, HKS^{+}15, JBB^{+}16, KHH19, KPW^{+}11, KHCH14, KGL^{+}16, LKT17, LEN^{+}15, LKS^{+}16, LAC^{+}19, LGC16, MCCA18, MKBSK19, MSK^{+}17, MW15, NRS16, ORGE16, OSHS19, PPT12, PH13, PML^{+}19, PSNE15, PMRRA19, RSN16, RS19, SNO^{+}16, SKV^{+}19, SHT^{+}17, SFI^{+}18, SBC^{+}17, TLG^{+}11, VLJ^{+}10, WCB^{+}10, WKAM^{+}19, WHR18]. Arctica [MWC^{+}16]. area [BBT^{+}10, CKB^{+}16, FCD12, GRE^{+}16, GBMG12, HLGA17, RPG13]. areas [HCAF18, MPM^{+}15, QS19]. arenaria [BMDC10]. Arenicola [VPWW10]. Argentina [HPM^{+}10, VBGG^{+}13]. Argyle [WMI^{+}17]. arid [BMBI12, Ker17]. armed [BL18]. aromatic [GPS15, ZZW16]. arriving [BBM11]. Arsenic [WZC13, BHB^{+}19, LCW17a]. Artemia [BPL^{+}19a, JW14]. Artificial [GSB^{+}17]. ascidians [JYS18]. Ascomycete [RJFMG17]. Asellus [FA10]. asexual [HRPW15]. ash [MLL^{+}14]. Asia [MKG^{+}15, TLB^{+}16]. ASLO [Ano19a]. aspergillosis [RBRH10]. Aspergillus [RBRH10]. aspirations [YKBJL12]. assemblage [BDC^{+}14, CPOMA15, JC14, MLL^{+}14, MS13, PMP^{+}12, RRGGCA19, WS13]. assemblage-based [BDC^{+}14]. assemblages [AMNU16, BSMC12, CEB^{+}17, CHPH13, GGPM^{+}10, IHSS^{+}19, LCS^{+}19, MF19, MBTK18, NRS16, RS19, SLA^{+}18, SRM^{+}18, TCG^{+}17, VSP^{+}11, ZCL^{+}19]. assembly [LCW^{+}17b]. assessed [CCK^{+}12, WP14]. Assessing [AC15, FZL^{+}14, JSH12, JHW^{+}19, NBDM16, YJO^{+}19, BJDMH10]. Assessment [EM13, FPGR^{+}13, HPS10b, LALGM18, MTU18, SAS^{+}11]. assessments [BSB^{+}18, BGB^{+}14, KBA^{+}14]. Assimilation [SHL^{+}18, BCRC16, CFB14, CMM^{+}11, FHR^{+}15, GWD^{+}16, KHP18, TEGL11, XLS^{+}19]. assimilatory [RvSM17, TG17]. assist [OBM^{+}11]. associated [AFSM17, BR17, BS18a, BIM^{+}16, BBK^{+}15, BSRP^{+}12, CMS^{+}18, CBF10, DVDB16, FCD12, HNL^{+}13, JB19, KZR^{+}16, LPO^{+}11, LSH^{+}17, Lee18, MVT^{+}17, ORC^{+}17, RWM^{+}19, RBRH10, Sch19, TGC^{+}10, USB^{+}10, VBGG^{+}13, ZOB^{+}15]. Association [SSL^{+}12, SKLG10]. Associations [LEK^{+}18, PGRR^{+}19]. astaxanthin [NZH^{+}11]. astreoides [MPSA17, TEGL11]. asymmetry [JGR^{+}14]. Atchafalaya [SFB12].

Athabasca [RKWH18]. Atlantic [Ano17l, DTL^{+}19, HLJ12, JWGH19, MvdPK^{+}15, NLO^{+}12, PFvO^{+}18, RWM^{+}14, RKMN^{+}13, WMBR13, ABB^{+}14, ASSG12, ÁSNCÁ^{+}13, ASA^{+}18, BFW^{+}13, BLW15, BCRC16, BSB^{+}10, CCV^{+}18, CTA^{+}19, CR16, CSS^{+}16, CWHP14, CSC^{+}11, DKSA19, DVDB16, FPP^{+}19, GMGM^{+}13, HWZ13, HGM10, KCL^{+}14, KKHP14, KMH^{+}17, MRKR^{+}14, MMD15, MCGF^{+}11, PPT12, SLG10, SBBNM14, WWC^{+}13, WCC^{+}17, WB19, WM17, ZXZ17b]. atmosphere [ZYZ19]. atmospheric [ACW^{+}18, HCL^{+}18, KK13, KHVS11, LGC16, MKG^{+}15, WCJ16]. atoll [GJR^{+}19, KCH^{+}12, RMK^{+}16]. attached [TCG^{+}17]. attenuate [GGL^{+}18]. Attenuation [TBK15, BA14, GAH11, LPO^{+}11, NLM^{+}12, OR16, RNT^{+}19, SSG^{+}17, SVMT15, UVGS10, WZTK15, ZWA^{+}14]. attraction [HJMD13]. August [PCY^{+}10, CWHP14]. Aurelia [RG13]. Aureococcus [KG18, KSWFG13]. Aureoumbra [KG18]. austral [KYRMD18]. Australia [AAIA14a, AAIA14b, CSGW18, LCBC16, RMJ^{+}18, REE^{+}12, UA10, VHrI10, WMI^{+}17]. Australian [MS13, CHL10, EMS16, HBM^{+}15, HVD^{+}18, PBL^{+}18, PD11, RHMSE15]. australis [HCK11]. Autochthonous [HDDH^{+}17, GMS^{+}18, RMF11, TTV^{+}13, WCCP14]. automated [FPP^{+}19, HGD14]. autonomous [GPH^{+}13, RGM^{+}11, SBM^{+}15, SPO^{+}18]. autotrophic [DTKMK15, GFT^{+}14, PGP^{+}14, SKJD^{+}14, TEGL11]. autotrophs [NCC14]. autotrophy [FPPA^{+}11]. autumn [JZZY18, KYRMD18]. availability [BVvB^{+}19, BMB^{+}18, CJWS15, ETI^{+}16, HBZ12, HVD^{+}18, IHSS^{+}19, KGRV18, KMF10, KvdPVB13, LTPA17, LRG16, MKB^{+}19, PCF14, PKB^{+}17, QFH18, SKV^{+}19, SBvH^{+}15, Spi15, SHF^{+}11, TIN^{+}14, TSC^{+}19, TFLS14, UMHH^{+}14, XFH14]. avara [SWM^{+}10]. Average [GPCJ16, KTRK11, Kir13, MB10]. Avoidance [HMD11, HJMD13, HL13, TWWY18]. Avoiding [BSB^{+}18]. away [WKSR13]. axenic [SLC^{+}16].

B [BMM^{+}13, BWP^{+}10, FLLH18, KMP^{+}11, KSWFG13, VMF^{+}11]. B1 [PBA^{+}15]. Bachmann [MLS^{+}14]. Bacillariophyta [RASD10]. back [NBG17, TDM^{+}13]. back-barrier [NBG17]. back-reef [TDM^{+}13]. background [SBR^{+}13]. backscattering [ASK^{+}11, BA14, NLM^{+}12, RSN16]. bacteria [ATP^{+}15, AGCA16, BS18a, BPA12, BSB^{+}10, DMSHC16, FDS^{+}14, FYT^{+}12, GVS^{+}10, HAC^{+}11, JAZ^{+}10, KWM^{+}19, KHG^{+}13, LCW^{+}17b, LFGK10, MLK11, MVT^{+}17, MTW12, MDE11, PBA^{+}15, PSZ^{+}13, SPP10, SRCL^{+}13, Sch19, SSS^{+}19, TMK^{+}13, USB^{+}10, WBZ^{+}13, YLH^{+}16]. Bacterial [ASSG12, BNW^{+}14a, MGS12, PCM^{+}16, ŠGH^{+}18, Ano10, BNW^{+}14b, BC10, CMB10, CFF^{+}17, FBV11, HT17a, LRM17, LFGK10, LTX^{+}17, MMPSB14, MW15, OCLW11, PSNE15, PD11, RSTS^{+}18, RGGL^{+}12, RGLM^{+}12, SDMK10, TGC^{+}10, TCG^{+}17, TBAS14, TST^{+}19, UMHH^{+}14, WCJ^{+}15, WZBW^{+}11, dGCB^{+}11]. bacterioneuston [HPS10b]. bacterioplankton [BB11, GKT^{+}15, HGG^{+}17, LZR^{+}17, SNM11, SPFP11, VF10, dKNL^{+}15]. bacterivores [ŠGN^{+}19]. bacterivorous [ŠGH^{+}18]. bacterivory [WSUC^{+}18]. Baikal [PRS^{+}18, KIH^{+}15, KZR^{+}16, KZR^{+}19, OWS^{+}17]. balance [AdGAD14, CR10, DdG10, LRM^{+}19, NPT11, PS13, RBY^{+}17, RAKE05, RWC16, SJB^{+}19, SM10, SSW19, WGC^{+}13, WLR17, YWL^{+}17]. balanced [BMN16]. balances [LKF^{+}18]. Balancing [UFW^{+}18]. Balaton [GTPB^{+}11]. ballast [BBM11, MAC^{+}10]. ballasting [vdJFS^{+}18]. Baltic [RCJ15, ACC^{+}17, BBT^{+}10, BKD^{+}16, DDH^{+}19, EO13, HJB^{+}12, HPS^{+}10a, JTH^{+}11, KBH^{+}19, KKHP14, MDSG18, MSR16, NZH^{+}11, Piw19, RF13, SLE10, SWM^{+}18, SFLQ^{+}19, SBH^{+}11]. Baltimore [RWM^{+}19]. banding [TLB^{+}16]. Bank [MBBG^{+}12]. Barbara [BSG14]. bardawil [YLJ11]. bare [EMO^{+}11]. Barents [LFB^{+}10]. barnacle [LAM12, PRL18]. baroclinic [ILPL13]. barrel [MBLP11, MJH^{+}16, WMP^{+}19]. barretti [LKF^{+}18]. Barrier [BWS10, CUW11, LÁSDC18, MLCD13, RGG^{+}10, UA10, GSZL13, NBG17, TvBR^{+}19]. barriers [NG13]. basal [GFDC11]. base [JTV^{+}16]. based [ALdML^{+}14, BMN16, BPW^{+}19, BISZ17, BLG^{+}15, BDC^{+}14, CSGW18, CMW^{+}19, DRE^{+}10, DB11, FFA13, FYVU17, FPP^{+}19, GM12, HOD^{+}17, JGR^{+}14, KWF^{+}17, LHSG15, LMR14, MMN^{+}10, SOM17, SHM^{+}19, SPMW11, SGRB10, TBLG14, VMC^{+}13, WS18, ZWA^{+}14, ZD18, ZKMT^{+}13]. baseline [BJDMH10, DLP13, SMG12]. basic [HESU13]. Basin [HMV12, LBNT11, NTK^{+}18, NLO^{+}12, RKL14, WMBR13, ABD^{+}17, BMF^{+}16, DL11, JABZ19, Ker17, LEK^{+}18, MAD^{+}15, SSGB^{+}17, SI10, SRAB10, SRA10, VPMrI12, WZG^{+}14, ERA^{+}12, MAC^{+}10, PCY^{+}10, WB19]. Basin-scale [NLO^{+}12, RKL14, SI10, VPMrI12]. basins [CGT16, GBC^{+}17, WFK^{+}16]. basis [JC14, LCCF10, VdRA^{+}19, ZF17]. batch [BRR^{+}13]. bathyal [PCF14]. bathymetric [NSO19]. bathymetry [BSRP^{+}12, VPMrI12]. bathypelagic [GCH^{+}18, MVT^{+}17, YYMN13]. Bay [CKB^{+}16, FYC^{+}18, FGMN17, GGPM^{+}10, GMBL16, HNSM12, HONR11, MDE11, PCD^{+}19, RVvdP^{+}17, TNI19, MF19, CJS^{+}17, DTPP12, DDF^{+}10, GGL^{+}18, GGTC^{+}18, GK14, HTLM18, KHK^{+}19, MGSM10, NHP17, Scu16, SGA^{+}17, SHK13, TK12, TKB18, ZSM14]. Bayesian [CAQS16, CS12, HSBA10]. be [CR11, CBF11, DKSA19, HLFM^{+}10, KPV^{+}11]. beach [GWN^{+}12, MBH^{+}15, SWE^{+}18, TvBR^{+}19]. beam [GAH11, NLM^{+}12]. bearing [JLRK12, UA10]. Beaufort [PvDM^{+}13, ABD^{+}17, LGC16, SSFF12, SLA^{+}18, STC^{+}11]. because [Lat14]. Becker [Bre10]. becomes [HATF17]. bed [GK10, GK14, TMH^{+}10]. beds [SWCL12]. been [BHC13]. before [GBS17, KHPIP^{+}14, SS16]. behavior [BRF^{+}17, CBP12, FDBW16, GPL11, vSGAK17, HV16, HPS^{+}10a, JSFC18, KSTA18a, KSY11, LWE^{+}19, MCT^{+}14, MFM^{+}12, NA17, SDS^{+}11, WCB^{+}10, WMC^{+}18, WD15]. behavioral [BRT^{+}10, CPOMA15, LDCT11]. behaviors [KYRMD18, PGB^{+}19]. below [OMSC13]. Belt [BDB^{+}14]. beluga [BCF^{+}17]. beneath [JTG^{+}11, SNG^{+}14, VMAS^{+}16]. benefits [HCAF18, MBHG11]. Benguela [NLO^{+}12]. Benthic [BVvB^{+}19, BSY^{+}16, CCW^{+}19, DFWPK16, DKG15, GLF18, GAK^{+}19, KYC^{+}15, MBB^{+}18, MDF^{+}14, MGL^{+}16, NHS^{+}12, Spi15, SNG^{+}14, WS13, AWK^{+}17, AGMR14, BDP^{+}19, BSM17, BNW^{+}14a, BHV^{+}17, BRF^{+}17, BBB^{+}14, CFAE^{+}15, DSS^{+}11, DvOR^{+}16, DRP^{+}17, EMO^{+}11, GLS^{+}13, GJWS14, GJWS16, GFDC11, GSB^{+}17, GVS^{+}10, GN16, HSLH^{+}14, HA16, IH11, KBM^{+}14, LGV13, MSSH12, MTSG18, MWS10, MBLD15, MSK^{+}17, MPvBS^{+}18, MDS^{+}10, NCC14, NCT^{+}15, NB17, PMA18, RPI^{+}12, RSG11, RPB17, RBD18, SLK^{+}10, SHKU11, SSH^{+}14, SCP^{+}16, UA10, WLS^{+}11, WZG^{+}14, WXMS10, WKAM^{+}19, ZCL^{+}19, vHOM^{+}19, vOSH12]. Benthic-pelagic [BSY^{+}16, WS13]. bentho [SAS^{+}11]. bentho-pelagic [SAS^{+}11]. benthopelagic [PCF14]. Bering [Tho19]. Bermuda [ZXN^{+}12]. Berry [CMW^{+}19]. beryllium [CSJ^{+}14]. best [KPV^{+}11]. beta [HT17b]. Between [ZKL^{+}14, ALL^{+}10a, AHH^{+}16, AFG^{+}16, AGCA16, BRS11, BMW10, BSN^{+}14, BC19, BISZ17, BVSM15, BDU^{+}19, BSSR10, BCF^{+}17, CL10, CL11, CL17, DBSP^{+}16, DTKMK15, ETI^{+}16, FT11, GLS^{+}13, GKT^{+}15, HAC^{+}11, HMD11, KM10, KHK^{+}19, LEK^{+}18, LLH^{+}15, LKLH10, LFGK10, MCWB10, MHA^{+}18, MBBG^{+}12, PPT12, PS13, QFH18, RMF11, RSJ^{+}18, RRD14, RCIB14, RPG13, RWC16, RKTLM18, SKLG10, SSH^{+}16, SBK18, SSFR19, ŠSP17, TCG^{+}17, TDM^{+}13, VPC10, WZG^{+}14, WC17, XFH14, YP18]. Between- [ZKL^{+}14]. Beyond [KTRK11]. bias [BSB^{+}18, Lat14]. biased [BD15]. Biases [SCL^{+}19, Lan14]. bicarbonate [CF13a]. biennial [DdD^{+}10]. Bight [DTL^{+}19, HSC^{+}14, SNvD^{+}10, SBBNM14]. binding [BBB^{+}14, SH10b]. Bio [DCRC16, ASK^{+}11]. Bio-optical [DCRC16, ASK^{+}11]. bioaccumulation [JW14, TW11]. bioadvection [VPWW10]. bioassay [GBL13]. bioavailability [GdG11, JSK^{+}15, LÁSDC18, PCO^{+}15, RM14, WCJ^{+}15]. Bioavailable [JBLJ12, NSV^{+}14]. biochemical [Ano21c, MPvBS^{+}18, VdRA^{+}19, WRWPG19]. biochemistry [PWF18]. Biodegradation [DBA16]. biodiversity [IBPG17, MTU18]. bioerosion [LSD18]. Biofilm [MACM11, BMBI12, MBP^{+}17, Sch19, TBAS14]. biofilms [BLMS17, MBP^{+}17]. biogenic [FTC10, HSC^{+}11, JZZY18, KBL^{+}10, KNL10, LYL^{+}17, SKLG10, WLR17]. Biogeochemical [BSC^{+}15, CT18b, MTM^{+}16, SSS^{+}16, THH^{+}13, CA08, DHG^{+}17, FPG11, MBH^{+}15, MT11, MBC^{+}18, MAFCD^{+}18, NO17, RGB^{+}19, SH10a, Spi15, SSC^{+}17, SCP^{+}16, TIF^{+}15, TGG^{+}11, UFW^{+}18, XDK^{+}17]. biogeochemically [RDB^{+}16, SGS18]. Biogeochemistry [MVG^{+}15, VMAS^{+}16, ZXL^{+}19, BIS^{+}10, BLWV10, CHL10, GCSO14, HHM^{+}18, KCM^{+}10, MHPW18, RDP^{+}17, SPP^{+}16, TBSL17, WZC13]. Biogeographic [CLJ^{+}19]. Biogeography [PVA^{+}19, LCW^{+}17b, SS17, TSSH19, WDMF13]. biogeomorphological [BBR^{+}14]. biokinetic [TW10a]. biolability [MGJH18]. Biological [CK12, CK13, GRT^{+}14, HGG^{+}17, MQJG13, PML^{+}19, SFLB16, ÅCA^{+}18, BDB^{+}14, BSM17, BVSM15, BBCM^{+}13, CFD15, CGB^{+}18, CMK^{+}10, FSBT16, HDK^{+}12, HKP^{+}16, HLFM^{+}10, HCC^{+}13, HZC^{+}13, KEH^{+}14, KKH11, LBS17, LALM16, LC11, MMC^{+}10, MTH^{+}11, MGL^{+}13, OLF^{+}11, QWRJ10, RLC^{+}11, RNG^{+}13, SBBNM14, TKK^{+}17, WCP^{+}15, ZCY^{+}15]. Biological- [GRT^{+}14]. biologically [HD19]. bioluminescence [VdRA^{+}19]. biomagnification [JSB^{+}14]. biomarker [BBS^{+}18, BCF^{+}17, WCV^{+}12]. biomarkers [BAY^{+}14, GLS^{+}13, JTV^{+}16, dKYH^{+}12]. Biomass [SGJB14, BJF18, BPGE13, BBSK18, FDH^{+}14, GSB^{+}17, KKS10, LdJMS^{+}13, LHSBP18, MRE18, PRS^{+}18, PHG13, PWF18, RVvdP^{+}17, SBT^{+}19, SBK18, SPGRP^{+}17, SPG^{+}11, YWY^{+}15]. biomasses [YP18]. biomechanical [LdlSB^{+}12]. biometry [CNL^{+}15]. biomixing [NL14]. biomolecules [CSJ^{+}14]. biophysical [RAV^{+}17, SSN12]. biopolymers [SH10b]. bioreactor [VPG^{+}19]. bioreactors [DMMV15]. biosynthesis [GvBBB17]. biota [JPH^{+}18]. Biotic [RZW11, DDF^{+}10, HCF^{+}10, WFB^{+}11, uGH^{+}11]. bioturbated [MBB^{+}18]. bioturbation [RF13]. Bioturbator [TTTM^{+}19]. Bioturbator-stimulated [TTTM^{+}19]. biphenyls [CMW^{+}19]. birdfoot [TT12]. birds [PHDH14]. birth [BD15]. Biscay [DTPP12, GGTC^{+}18]. bismuth [FTC10]. bisphosphate [nVOH12]. bivalve [HSR15, WMC^{+}18]. Biwa [THH^{+}13]. black [DPG^{+}12, FYC^{+}18, BRS18, MGHS18]. blade [RN14, ZWA^{+}14]. blades [HRN11, RN14]. Blanes [GGPM^{+}10]. Bleaching [MBLP11, Ano21c, BWS10, FZL^{+}14, GBR14, HBD^{+}11, KHPIP^{+}14, PST^{+}13, SHKU11, SIW^{+}11, WRWPG19]. Bled [MMGP^{+}12]. Bloom [BRF^{+}17, BDB^{+}14, BVSR^{+}15, CR16, DVC^{+}17, GLMG15, GGTC^{+}18, HST^{+}14, HMD11, HZC^{+}13, HKS^{+}15, IHSS^{+}19, JHLK^{+}19, JTG^{+}11, KSWFG13, LFH^{+}12, LBR^{+}13, MTH^{+}11, NAH^{+}11, OCLW11, PKB^{+}17, PCM^{+}16, PCY^{+}10, RKBA14, RKMN^{+}13, SWD^{+}14, SLG^{+}14, SSH^{+}14, SFLQ^{+}19, ŠSP17, SHF^{+}11, TIF^{+}15, WCJ^{+}17, WCJ^{+}15, WSTD10, ZXM^{+}11]. bloom-derived [WCJ^{+}17]. blooms [BSY^{+}16, CBS^{+}17, GCSO14, HLH13, HZC^{+}13, JTV^{+}16, KG18, KIH^{+}15, KBVW12, MQP^{+}16, MGL^{+}16, OFGF12, OSB^{+}15, PWS^{+}11, QHVM18, SWZ^{+}15, SBM^{+}15, SK19, SMN^{+}15, SS12b, SS12c, TF11, VHrI10]. blow [NA17]. blowout [SSB^{+}16]. blue [BBS^{+}18, HBM^{+}15, LPLH18, Les16, NSO19, OWM^{+}18, VdSLC^{+}16, HCAF18]. bodies [GGC^{+}14]. Body [DOD10, DRE^{+}10, HLGA17, Kiø13, OR16, PWF18]. body-mass [HLGA17]. Boersma [Bre10]. bog [CCC10]. Bohai [SW14, SCQ^{+}17]. border [HPS10b]. bordering [FDS^{+}18]. boreal [AAC^{+}19, BLMS17, CA08, CKD^{+}16, CGT16, GBP^{+}12, HHE^{+}19, HGdG^{+}19, JBLJ12, JTV^{+}16, KHTO13, LKF^{+}18, OBT^{+}11, PSB^{+}16, SS16, SPSG14, SBK18, SH10a, SBB^{+}18, SPG^{+}11, SSM^{+}19]. bores [GJR^{+}19]. borne [KZB^{+}10, LKLH10, SS12b, SS12c]. Bosmina [KM10, FSST11]. both [HDK^{+}12, RWM^{+}19, RVvdP^{+}17, TMK^{+}13, WDL^{+}17]. bottle [SSC^{+}17]. Bottom [LJL^{+}18, WD15, BH13, DHH15, EM13, GMBL16, GdVT^{+}11, KT13, LBR^{+}13, MGW^{+}13, PDER10, RSG11, SPP^{+}16, SHK13, SWL11, WCJ^{+}17]. bottom-layer [KT13]. Bottom-up [LJL^{+}18, WD15, BH13, LBR^{+}13]. bottom-up- [PDER10]. bottom-water [RSG11]. bound [LFC17]. boundary [BBB^{+}14, CT18a, CHPH13, DHH15, HEBS10, HCH^{+}19, RLC^{+}11, SMLC^{+}18, SWL11, WKSR13]. boundary-layer [SWL11]. brackish [FYT^{+}12, PSZ^{+}13]. Brady [HAL17]. branched [BAY^{+}14, ZKMT^{+}13]. Bransfield [MVT^{+}17]. Brazil [AHD^{+}18, CKB^{+}16, FMP^{+}13, NEH^{+}19, PMP^{+}17]. breadth [PWWF18]. break [OPZ13, SS16]. breaking [LHS19]. breakup [LHS19]. Breton [CJC^{+}12]. brevetoxin [KPSW10]. brevetoxins [HST^{+}14]. brevis [HST^{+}14]. brightest [BDB^{+}14]. Brine [BPL^{+}19a, JW14]. Bringing [SIH^{+}17]. British [CHHT18, VHM^{+}10]. Brittany [KCL^{+}14]. broad [AAO^{+}19, BBT^{+}10, MBAS^{+}17]. broad-scale [BBT^{+}10]. broadcast [BMC^{+}16]. Bromoform [LM12]. bromoperoxidase [JBPM15, LM12]. brown [CBP10, KG18, KSWFG13, OBT^{+}11]. brown-water [OBT^{+}11]. brownification [BKA^{+}14]. browning [NBSMN19, WSUC^{+}18]. Bryozoa [SMF10]. bubble [LVDM19, VHM^{+}10]. bubbles [RWB^{+}19]. budget [AACS11, ACA^{+}11, CKD^{+}16, CAS^{+}17, CWRX19, LKF^{+}18, MGL^{+}16, MAD^{+}15, SOM17]. budgeting [SSB^{+}16]. budgets [AMNU16, EMS16, HBR^{+}14, MRBR10, TDM^{+}13, VW17]. buffered [MMG16]. bugensis [KKS10]. build [SS16]. build-up [SS16]. building [BBR^{+}14, CRS^{+}17, ELJ^{+}16, GFPSG13, JLRK12, PGRR^{+}19, WGH^{+}16, YLH^{+}16]. bulk [CHL^{+}17, KWB^{+}16, RSN16, RGLM^{+}12]. Buoyancy [STCS10, PT11, WMI^{+}17]. buoyant [LN11, SVS^{+}19]. burial [AC17, EMS16, KBH^{+}19, MMGO^{+}17a, RRAS17, SML^{+}19]. burning [Dem19, YWY^{+}15]. Burragorang [VHrI10]. Bythotrephes [BSBK13, BBB^{+}17, BBS12, WLV17, WLV18].

C [Bre14, CHL^{+}17, CKCEP10, DPM18, HBZ12, KPJ12, MMGO^{+}17b, MMXC15, MMPSB14, OEMB10, OEM12, ORGE16, PFH^{+}17, UA10, WLG^{+}16, WCCP14, YJO^{+}19, dKYH^{+}12]. C-labeling [OEMB10, OEM12, ORGE16]. C. [JMN15, PPT12]. Ca [HATF17]. Cable [KWM^{+}19]. Cadmium [BC19, BAG^{+}14, LFC17, MBC^{+}18, TW11, WMBR13, XSAM12]. cadmium-phosphate [WMBR13]. caespitosa [FPPA^{+}11, FPGR^{+}13]. calanoid [BD15, DHK11, GPL11, KNA^{+}14, MTU18, PJ16, TW10b, WFR10]. Calanoida [HAL17]. Calanoides [PT11]. Calanus [CBP12, FGMN17, FNSS15, HTL^{+}18, HKS^{+}15, JWGH19, JMNG^{+}13, JMN15, MHA^{+}18, MMJ^{+}12, PPT12, TGG^{+}11, VGJ17]. Calcareous [VFS^{+}15, TTV^{+}13]. calcification [BSCC15, DSM^{+}18, FRA^{+}17, MBC^{+}16, MLGZ16, SBdB10, SLC18, SHD^{+}11, THFG16, WCS^{+}18]. calcifiers [CESC13, CESC14, SHD^{+}11]. calcifying [SBdB10]. Calcite [BDB^{+}14, PE17]. Calcium [KWGN^{+}10, SPS19, AA18, MMG16, PWF16, RPI^{+}12, SPTS15, TW10a]. calculations [Kus14]. California [HSC^{+}14, KH16, SNvD^{+}10, BSG14, BPA12, BPPF12, BTC^{+}19, BWS^{+}14, BBB^{+}14, Car10, CDA16, DLP13, FCD12, HMV12, HONR11, LBNT11, MMC^{+}10, MBC^{+}18, MQJG13, NSO19, OFGF12, PMLC^{+}10, PMPD13, SMM11, SBM^{+}15, SLBNG11, SSGM18, TSC^{+}19, TGGZS^{+}10, WM12, WAB^{+}17]. camouflage [JGR^{+}14]. Can [AdBVA10, CBF11, HBM^{+}15, LFC17, CR11, FPD^{+}10, GHS14, KPV^{+}11, SGH12, SOM17, Sha10]. Canada [RKWH18, AMQ^{+}11, BSCG17, BPW^{+}19, BLS^{+}16, FLM^{+}19, Ker17, MPM^{+}15, MGSM10, RPMK17, RPH^{+}10, ZHN^{+}10, vdHHC^{+}19]. Canadian [AA18, BAG^{+}17, BBM11, DMMV15, TLG^{+}11]. Canary [BAA^{+}13, BSB^{+}10]. cannibalism [FGMN17]. cannot [MLS^{+}14]. canonical [WLO^{+}19]. canopies [AGLM17, AGML18, HE10, WZTK15]. canopy [ARB^{+}19, GK10, SVLS^{+}16, SCPE15]. canopy-forming [ARB^{+}19, GK10]. Canyon [HYK^{+}15, SPFP11]. canyon-shaped [SPFP11]. canyons [KCL^{+}14]. capabilities [TSK13, VIS^{+}13]. capacities [CMS^{+}18]. capacity [CCV^{+}18, MMGO^{+}17b, MHH^{+}17]. Cape [MBBG^{+}12]. capricorni [EMO^{+}11]. capsules [CSGW18]. capture [SGCC16, TSK13]. carbohydrate [AMNU16]. carbohydrates [OEMB10]. Carbon [BTH^{+}16, CAS^{+}17, CSME13, DvOR^{+}16, GCH^{+}12, HV16, HAC^{+}11, HLFM^{+}10, HCAF18, JM16, LCM^{+}12, OBM^{+}11, OBT^{+}11, SML^{+}19, WYL16, vOSH12, AWG^{+}12, AACS11, ACA^{+}11, AC15, AC17, ARB^{+}19, BRR^{+}13, BSCG17, BBS^{+}18, BHW^{+}12, BBLN11, BMBI12, BDS11, BHD^{+}17, BMD17, BOT^{+}15, BAY^{+}14, BCF^{+}17, CF13a, CEPPR14, CPPdAR^{+}13, CKP^{+}15, CRJ^{+}14, CFB14, CTG15, CCW^{+}19, CKD^{+}16, DFWPK16, DIC^{+}18, DTPP12, DBA16, DNH^{+}18, Dem19, DPG^{+}12, DKSA19, DRP^{+}17, DBC^{+}13, DVDB16, EBMR12, FYC^{+}18, FHR^{+}15, FCRW^{+}16, FB12, FLP^{+}10, FDS^{+}14, GFH13, GJWS14, GJWS16, GWB^{+}14, GBP^{+}12, GdG11, GBS17, GHSR^{+}16, HHE^{+}19, HGG^{+}17, HBR^{+}14, HHM^{+}18, HLJ12, HNHS^{+}15, HBM^{+}15, HEH^{+}17, HSTK15, HMH^{+}16, HRPW15, HGT^{+}18, HDDH^{+}17, HEBS10, HVD^{+}18, JHD^{+}11, JMM14, Joh10, JP10, JTG^{+}11, KYC^{+}15, KHTO13, KCL^{+}14, KZB^{+}10, KKH11, KLEH16, KBT16, KGL^{+}16, KPJ12]. carbon [KRB^{+}18, KOFN11, LRM17, LH17, LdJMS^{+}13, LKF^{+}18, LCH^{+}14, LZC^{+}14, MSGS^{+}13, MMC^{+}10, MMGO^{+}17a, MZH15, MBLD15, MGS12, MHH^{+}17, MMPSB14, MPvBS^{+}18, MCYR17, MGSM10, OCR10, OEM12, ORGE16, OCLW11, OWM^{+}18, OVRJ13, PCD^{+}19, PBL^{+}18, PLS^{+}16, PHG13, PHLSSS19, RRAS17, RMF11, RR13, RM14, RCH^{+}15, RASV^{+}17, RHSD^{+}10, RHDTS^{+}11, RKTLM18, SSFF12, SVLS^{+}16, SCR^{+}12, SBvH^{+}15, SHT^{+}17, SLP^{+}14, SLA^{+}15, STC^{+}11, SLH^{+}15, SMG12, SFLQ^{+}19, SPR^{+}15, SBH^{+}11, SL10b, SLBNG11, SBKO18, SSGM18, SHL^{+}18, SH11, SCP^{+}16, SSS^{+}19, TYX^{+}19, TFH17, TJJ^{+}15, TDM^{+}13, TTTM^{+}19, TW10b, TBSR13, TMH^{+}10, UFW^{+}18, UIY^{+}11, VW17, VFS^{+}15, WLS^{+}11, WWC^{+}13, WLG^{+}16, WKG^{+}16, WDCH18, WWC^{+}18, WMC^{+}15, WCJ^{+}15, WC17, WGH^{+}10, WBB^{+}17, WDL^{+}17, WLHW13, XZGW17, ZXN^{+}12, ZHN^{+}10, ZYZ19, ZMWM11, ZHD^{+}16, ZCK^{+}16, dKYH^{+}12, dKNL^{+}15, vdJFS^{+}18, vEG10]. Carbon-to-chlorophyll [JM16]. Carbonate [BG10a, JCF^{+}10, CSME13, GDD^{+}16, HCAF18, KWGN^{+}10, MBC^{+}16, MMG16, NEH^{+}19, PLS^{+}16, RPI^{+}12, SPS19, SPTS15, WYL16, YH17]. carbonate-buffered [MMG16]. carbonation [BRS11]. carbonyl [ZYZ19]. carboxylase [nVOH12]. carboxylase/oxygenase [nVOH12]. carcasses [DJS18, EHT10, GGL^{+}15, KGT12]. Carcinus [GGC^{+}14, MCT^{+}14]. Caribbean [ASR^{+}17, BJDMH10, CMMKH12, Edm15, HGT^{+}18, LABJ18, MDS^{+}10]. carotene [YLJ11]. Carotenoid [SGVR16, SGME11]. Caryophyllia [CRB^{+}17]. Caryophylliidae [CRB^{+}17]. cascade [WLV17]. cascades [FPSL18, PLE^{+}17]. cascading [WHL^{+}11]. case [BAY^{+}14, CSJ^{+}14, IGP^{+}12, LDY^{+}16, PHL^{+}18]. Cast [vHOM^{+}19]. catchment [BBLN11, BSM17, BHM^{+}17, KKP^{+}19, RAB^{+}17, TTV^{+}13]. catenella [BRF^{+}17]. Caught [AAO^{+}19]. Caulerpa [EMO^{+}11, OBM^{+}11, RSTS^{+}18]. cause [Les16, SKV^{+}19, SHD^{+}11]. caused [BLS^{+}16, HZC^{+}13, LC11]. causes [FEW^{+}14, HCH^{+}19]. causing [SMN^{+}15]. cavernicolous [MGT15]. cavity [VMAS^{+}16]. Cayuga [EP14, PE16b]. CDOM [CDA16, DVC^{+}17, WSM^{+}19]. Cell [FAF^{+}12, BFW^{+}13, CL10, CBS^{+}17, CLWD13, DSM^{+}18, GC16, HPS10b, MDE11, NTA14, RGLM^{+}12, SDMK10, SBFB17, TNMV^{+}10]. cells [BCRC16, Clo18, KS13]. Cellular [FDBW16, BRR^{+}13, DBC^{+}13, HST^{+}14, KBHT19, SMH^{+}11]. centenary [GPA^{+}14]. Center [SPB^{+}14]. central [ÁSNCÁ^{+}13, ERA^{+}12, GFT^{+}14, HWZ13, KGL^{+}16, MVL^{+}10, MGW^{+}13, NO17, PCY^{+}10, SWM^{+}18, YYMN13, GTPB^{+}11]. centric [QFH18]. Century [MTU18, BLS^{+}16, Edm15, PDER10, RPH^{+}10, SSM^{+}19]. CH [CKB^{+}16, NSG^{+}16, PMY19a, RMH^{+}17]. chain [BTJ^{+}12, FLP^{+}10, YKBJL12]. chain-forming [YKBJL12]. Challa [WKJS^{+}14]. challenged [JSFC18]. challenges [APS^{+}19, GM12, SOO^{+}17]. chamber [VPC10]. chambers [GJWS14, GJWS16]. change [BSB^{+}18, BBQ^{+}10, BLS^{+}16, DDF^{+}10, FDB^{+}15, GSBR11, HMO^{+}18, Hir12, JBB^{+}16, KTK^{+}13, LSH^{+}17, Les16, LHS19, NUH^{+}12, PHDH14, RBG^{+}10, RG13, SFS^{+}16, VHrI10, VBGG^{+}13, WRO^{+}11, WGM16, WBZ^{+}14, WRH^{+}17, WHR18, ZEXH15]. changed [BHC13]. Changes [DMSHC16, JSK^{+}15, KK13, LMR14, MU17, RWM^{+}19, TMH^{+}18, YP18, BGW^{+}15, BAG^{+}17, BBK^{+}15, BCF^{+}17, BSH16, DCCB17, DML17, DHZ^{+}19, FWWF18, GMD11, GdVT^{+}11, HPS^{+}10a, HML^{+}14, KMC^{+}15, LG16, MTH^{+}11, MKK15, MPvBS^{+}18, MMJ^{+}12, PvEF12, PCO^{+}15, PMP^{+}12, PCM^{+}16, PDER10, PSNE15, QHVM18, RM14, RSE^{+}17, RGLM^{+}12, RPL16, SMLC^{+}18, SGA^{+}17, SW11, SSM^{+}19, TWP13, VKC18, WP14, ZWL^{+}14]. changing [FOT^{+}15, GDD^{+}16, JMNG^{+}13, PHL^{+}18, Spi15, SSM^{+}19, SHF^{+}11, VPG^{+}19]. Changjiang [GLI^{+}15, WLG^{+}16, WCJ^{+}17, ZYZ19]. Channel [BSG14, CVS^{+}10, GNHGM13, JWS15, KH16]. channeled [FRP^{+}14]. Characteristics [ZZY^{+}10, AJ15, CT18b, FBV11, FPG11, FDBW16, GSBR11, JZZY18, LC12, ŠGH^{+}18, SHL^{+}18, WYW^{+}10]. Characterization [DWDH10, DBV^{+}11, LTH^{+}12, SS17, TLH^{+}11, NRS16, RS19, SCPE15, WM12, WDX^{+}11]. Characterizations [PE17]. Characterizing [BBS^{+}18, JD16, WSTD10]. chase [EOM16, MMPSB14]. Chelator [MTW12]. Chelator-induced [MTW12]. Chemical [DHH15, FMP^{+}13, HJMD13, LCS^{+}19, RAKE05, SBC^{+}17, BDB^{+}14, BMPF19, BSBK13, CMK^{+}10, DL11, LG16, LKK13, MVL^{+}10, MTH^{+}11, SBDS^{+}15, TWWY18, CR10]. chemically [SPO^{+}18]. chemistry [BG10a, GDD^{+}16, GM12, LSHK11, MBC^{+}16, MHRH11, MMH^{+}18, PCD^{+}19, SBdB10, TMH^{+}18]. chemoautotroph [MWS10]. chemoautotrophic [MGT15]. Chemoautotrophy [MRC^{+}16]. chemocline [BNW^{+}14b]. chemostat [FDBW16]. chemostats [NCC14]. chemosynthetic [LFB^{+}10]. chemotypes [ALdML^{+}14]. Chesapeake [DDF^{+}10, GMBL16, GGL^{+}18, GK14, Scu16, SHK13, TK12, TKB18, ZSM14, ZHG15]. Chidami [SH10a]. Chile [GFT^{+}14, dlFN10]. Chiloscyllium [WLS^{+}11]. China [CFD15, DWDH10, GLI^{+}15, GBD^{+}10, HCW^{+}10, HCLS11, JZZY18, SW14, WDX^{+}11, XDC^{+}19, ZCY^{+}15, ZYZ19, ZXL^{+}19, CCK^{+}12, JHW^{+}19, LCW^{+}17b, LYL^{+}17, MQP^{+}16, TGC^{+}10, WXF^{+}15, WLG^{+}16, XXZ^{+}19, XPQ^{+}10, ZZY^{+}10, ZWL^{+}14, ZZW16, dKYH^{+}12, dKNL^{+}15]. Chinese [MNW^{+}19, PCPZ18, WWC^{+}18]. chironomid [CSGW18, SWP11, VHR^{+}11]. Chironomidae [REDW10]. chironomids [RHV^{+}13]. Chironomini [HNHS^{+}15]. Chironomus [SPPS10]. chitinase [BB11]. Chl [LLB17]. Chl-a [LLB17]. chloride [CR10, RAKE05]. Chlorophyll [ESMS13, BRR^{+}13, FWS^{+}14, FWO^{+}18, FAF^{+}12, HGD14, IH18, JM16, LCM^{+}17, LBC^{+}18, Lee18, LBS17, MSSH12, MRH^{+}15, XDC^{+}19]. chlorophyll-a [FWO^{+}18, Lee18]. Chlorophyll-normalized [ESMS13]. chlorophyll/biological [LBS17]. chlorophyte [GBL13]. cholesterol [SW11]. cholesterol-induced [SW11]. Chondrosia [SWM^{+}10]. chromophoric [CSÁS^{+}10, CDA16, DVC^{+}17, FB12, RCSÁS^{+}10, XSAHV13, ZZY^{+}10]. chronically [PKWS19]. Chrysaora [RG13]. Chubut [HPM^{+}10]. Chukchi [MBLD15, PvDM^{+}13, SLA^{+}18, SFI^{+}18]. ciliate [FPP^{+}19, JJ17, LEG^{+}10, ŠGN^{+}19]. ciliated [WOC^{+}18]. Ciliates [ZPK^{+}12, JB19, SBFB17]. Circle [PMRRA19]. circuit [PD11]. circulation [MGSM10, NI10, NHP17, RKL14, RPL16, SPSG14, SMA15, TvBR^{+}19, VAH11]. Circumpolar [TT14]. citrate [SMLC^{+}18]. Clade [DNH^{+}18, RWM^{+}14]. clades [AGCA16]. cladocera [PZHD18, FSST11, MXWC11]. cladoceran [BBB^{+}17, KM10, VHR^{+}11]. cladocerans [TW10a, TW11]. Cladocora [FPPA^{+}11, FPGR^{+}13]. Cladosporium [RJFMG17]. clam [BMDC10, MAS^{+}16, SDS^{+}11]. clams [BRM^{+}19]. clarity [BGW^{+}15, GDCM13, LH19]. classification [AC15, NRS16]. classifying [SAH^{+}19]. clay [DMN15]. cleaning [CGB^{+}18]. clear [CKCEP10, HS11, MSSH12, OBT^{+}11, TAE^{+}18]. clear-water [CKCEP10, OBT^{+}11]. clearance [AvSGK18]. clearly [PSH^{+}11]. Climate [EKS^{+}18, Les16, LHSBP18, RSE^{+}17, SFS^{+}16, SGG^{+}11, BJ15, BBQ^{+}10, BLWV10, CJHR19, DHZ^{+}19, FWO^{+}18, FVSL19, GMGM^{+}13, GSBR11, HW16, JBB^{+}16, KTK^{+}13, LHS19, MMB17, MWC^{+}16, NWT^{+}19, PMLC^{+}10, PHL^{+}18, RBG^{+}10, RCIB14, RG13, RAV^{+}17, SLE10, SRAB10, SRA10, VHrI10, VBGG^{+}13, WGM16, WBZ^{+}14, WRH^{+}17, WHR18, ZHN^{+}10, ZEXH15, vEG10]. Climate-index [SGG^{+}11]. Climate-induced [EKS^{+}18, DHZ^{+}19]. Climatic [WDH^{+}17, MHPW18, WP14, WOC^{+}18, ZWL^{+}14]. clocks [HTL^{+}18]. Clonal [MNW^{+}19, LGW^{+}19]. closed [SRAB10, SRA10]. closed-basin [SRAB10, SRA10]. Closing [LRM^{+}19]. closterium [BC19]. Clupea [DDH^{+}19, KKHP14, KMH^{+}17]. cnidarian [DBMP^{+}11]. Co [CLFW17, CR16, GWSEA10, MdBKL13, MCYR17, APB^{+}17, BR17, BOT^{+}15, CCV^{+}18, CGP^{+}19, CESC13, CWHP14, DSM^{+}18, EO13, HST^{+}14, HLSW^{+}15, HXS^{+}10, HCAF18, HCC^{+}13, HCL^{+}18, KRR16, Man10, MSR16, NWT^{+}19, OLC18, PMY19a, PSG^{+}16, QFH18, RR12, RMH^{+}17, RHMSE15, SSU^{+}16, SYW18, TJJ^{+}15, VSdG17, VTH^{+}18, WYL16, YH17]. co-exist [CR16]. Co-existence [CLFW17, MCYR17]. co-limitation [GWSEA10, MdBKL13]. coal [VB17]. coarse [CHW14, WLL^{+}11]. coarse-grained [CHW14]. coast [BWS^{+}14, GHSR^{+}16, KSG^{+}10, MQJG13, RPMK17, VdRA^{+}19, CWHP14, JCF^{+}10, LLH^{+}15, MWBM19]. Coastal [GYP^{+}18, JAS^{+}15, WMT^{+}12, AGLM17, AGML18, AWG^{+}12, AWK^{+}17, AAIA14a, AAIA14b, ACA^{+}11, AJC15, ABD^{+}17, ADS^{+}17, ACC^{+}17, ARB^{+}19, BSR^{+}17, BSCG17, BPPF12, BAG^{+}17, BG10a, BPV^{+}19, BSB^{+}18, BBJ^{+}19, BBQ^{+}10, BHM^{+}17, CCV^{+}18, CPG^{+}10, CMM^{+}11, CDA16, CFVU11, CKB^{+}16, CSD10, CWHP14, DMS^{+}18, DL11, DBC^{+}13, EBMR12, FDS^{+}14, GFT^{+}14, GGPM^{+}10, GGC^{+}14, GLMG15, GWSEA10, GvBBB17, GBB19b, GK10, HDK^{+}12, HVJ^{+}19, HSC^{+}14, HCAF18, HVD^{+}18, JM16, JHW^{+}19, JPH^{+}18, KBH^{+}19, KYR^{+}12, KYG^{+}12, KPSW10, KTK^{+}13, KSG^{+}10, KMH^{+}17, LRY12, LK15, LSH^{+}17, LCH^{+}14, LCZ^{+}19, LÁSDC18, LRM^{+}19, MGGS18, MSSH12, MJJMM17, MCC^{+}10, MBO^{+}16, MBBG^{+}12, MS13, NEH^{+}19, NPT11, OPA^{+}14, PSG^{+}16, PWS^{+}11, Piw19, QWRJ10, RSG11, RSM13, RPG13, RGB^{+}19, RETS16, SBT^{+}19, SLE10, SKGT17, SWD^{+}14, SSH^{+}16, SEYJ11, SSP^{+}18, SBC^{+}17, SRM^{+}18, SMW^{+}18, SOH^{+}18]. coastal [SH10b, SDMK10, TNI19, TDF^{+}17, TWP13, TZD^{+}15, VFME18, VLMTEW11, VPG^{+}19, WWC^{+}13, WKG^{+}16, WSM^{+}19, WGJ^{+}19, WM17, WDL^{+}17, YMB^{+}18, ZCY^{+}15]. coasts [BBM11, HW16, MFL11, WWC^{+}13]. cobalt [HS18, MBC^{+}18, NLO^{+}12]. Coccolithophore [MBC^{+}16, PCY^{+}10, BRS11, BDB^{+}14, BPL^{+}19b, FRA^{+}17, FCC11, KRR16, RR12, THFG16, ZKL^{+}14, ZBSR15]. coccolithophores [DTPP12, GYP^{+}18]. coccoliths [BSCC15, SBFC18]. Cochlodinium [JLG10, JLG11]. cocultures [BSMC12]. cod [JTH^{+}11]. codeterminants [PH13]. coefficient [FB12, GAH11, Kus14, ZD18]. coefficients [BA14, RDT^{+}14]. coexist [WBZ^{+}13]. Coexisting [LLL10]. Coherent [SSM^{+}19, dGCB^{+}11]. cohesive [SPP^{+}16]. cold [ÅCA^{+}18, CCW^{+}19, CBF10, GDD^{+}16, KCL^{+}14, LGC13a, LGC13b, MKB^{+}19, Tho19, WRB^{+}19]. cold-core [WRB^{+}19]. cold-pool [Tho19]. cold-seep [CBF10]. cold-water [GDD^{+}16, KCL^{+}14, LGC13a, LGC13b, MKB^{+}19]. coli [GWN^{+}12]. Colimitation [GFH13, ARML10, BG10b, OWS^{+}17]. collaborative [KWF^{+}17]. Collecting [Ano19c, GBB^{+}18]. collection [SSC^{+}17]. colloidal [SH10b]. colonial [SJM11, WKK^{+}11]. colonies [HBD^{+}16, SIW^{+}11]. Colonization [BBB^{+}17, BDC^{+}14, MKBSK19, RBM14]. Colonizing [MSS^{+}18]. colony [HNZ^{+}16]. color [FSCB11, SBK18]. Colorado [HYK^{+}15]. Colored [NWT^{+}19, TZD^{+}15, TAV^{+}10, UVGS10]. Columbia [CHHT18, CFF^{+}17, DBRB^{+}15, GPS15, PHPH^{+}16, SRAB10, VHM^{+}10]. column [AFG^{+}16, BNW^{+}14b, DWDH10, ERA^{+}12, FDS^{+}18, HCK10, HHM^{+}18, HJB^{+}12, HD19, Kir13, MRC^{+}16, SSB^{+}16, UA10, VLDM19, WDX^{+}11, WBZ^{+}13, ZOB^{+}15]. columns [AdBVA10]. combination [HAL17]. Combined [TJJ^{+}15, BRT^{+}10, HSLH^{+}14, KGC^{+}12, LABJ18]. Combining [KFJ13, WS18]. Comment [APF^{+}18, CL11, KBA^{+}14, Lat14, MLS^{+}14, PJ16, PSH^{+}11, PHJ12, SLU11, SM11b, SDH^{+}14, ACC^{+}19, KVA18, KGC^{+}16, Lan14, SH10a]. comments [BHC14]. Common [FMGR^{+}11, ASR^{+}17, BH16, HS10, HTLM18, KNA^{+}14]. communities [ASA^{+}18, BFW^{+}13, BLW15, BDS11, BMM^{+}13, BPPF12, CBF10, CPF16, DTKMK15, DdG10, DBC^{+}13, EMO^{+}11, FPP^{+}19, FT11, FBFR13, FEC^{+}16, GBC^{+}17, GAK^{+}19, HRMD19, HS11, HVD^{+}18, KCH^{+}12, KTK^{+}13, KGvdH16, KvdPB18, LUM15, MTM^{+}16, MWS10, MVNG11, MU17, MSM^{+}17, ORC^{+}17, PCPZ18, RLC^{+}11, Rie15, RPB17, RSTS^{+}18, RPH^{+}10, SGJB14, SLA^{+}18, SPHVA19, SMR^{+}17, TGC^{+}10, TBAS14, VFME18, VP15a, VLJ^{+}10, VML^{+}19, WVV^{+}11, WKAM^{+}19, ZEXH15, ZOB^{+}15]. Community [CSC^{+}11, KHPIP^{+}14, SPB^{+}14, ABB^{+}14, AdGAD14, ANP^{+}14, AJC15, ABD^{+}17, BA14, BSG14, BAG^{+}17, BRS18, BMD17, BHB^{+}12, CB12, CCK^{+}12, CFVU11, CKCEP10, CVS^{+}10, DKG15, DBRB^{+}15, DBFL11, DDF^{+}10, DMB^{+}12, ETKL16, FMM^{+}14, GTPB^{+}11, GEC^{+}17, GWD^{+}16, GWB^{+}14, GSB^{+}17, GCH^{+}12, GN16, HSLH^{+}14, HMFB16, HHHT19, HVJ^{+}19, HLJ12, HHS^{+}18, HPS10b, HEBS10, HCH^{+}19, KEH^{+}14, KT13, KHH19, KTS^{+}14, KMP^{+}11, KMH^{+}17, KPV^{+}11, LSH^{+}17, LCW^{+}17b, LFGK10, LTX^{+}17, LDT^{+}11, MH16, MPSA17, ML19, Meh10, MVT^{+}17, MGL^{+}13, MvdPK^{+}15, MRE18, MGT15, NFW13, OALD10, PCM^{+}16, RBCS16, SNM11, SPP^{+}16, SFI^{+}18, ŠGH^{+}18, SSH^{+}14, SKKV11, SPG^{+}11, SVG^{+}18, SSC^{+}17, TCG^{+}17, TLR^{+}13, VKC18, VMCM^{+}17, WVGB10, WCJ^{+}15, WXMS10, WZBW^{+}11]. community-wide [Meh10]. Comparative [ACD10, DBFL11, MTK^{+}17, SH10a, XFH14, BPPF12, RF13, UIY^{+}11]. compared [EMO^{+}11]. Comparison [LSHK11, WCB^{+}10]. comparisons [TDS^{+}10]. competing [APF^{+}18, KG18, KVA18]. Competition [BVSM15]. Competitive [HS18]. complementary [AHJS15, APF^{+}18, KVA18]. Complex [FT11, CHPH13, FDB^{+}15, LTH^{+}12, MNW^{+}19, PMP^{+}12, PTS12, RNT^{+}19, VMMS^{+}13]. complexed [SMH^{+}11]. complexes [JBT11, XSAM12]. complexity [PH13, SVLS^{+}16]. component [HLFM^{+}10, KFP^{+}18, PDP^{+}10, PE16b]. components [LBR^{+}12]. composed [GN16]. Composition [CBP12, OEMB10, SLA^{+}18, ALL^{+}10a, AAIA14a, AAIA14b, Ano21c, ABD^{+}17, BHB^{+}12, BSMC12, CWF11, CKCEP10, DBFL11, DMB^{+}12, FUS^{+}16, GWD^{+}16, GSB^{+}17, HVJ^{+}19, HSTK15, HCW^{+}10, HCLS11, HMFF10, HMFF12, JSK^{+}15, Kiø13, KPV^{+}11, LLB17, LVDM19, LGV13, LYH17, LFGK10, LBNT11, MVL^{+}10, MPONC^{+}17, MTM^{+}16, MMXC15, MPSA17, ML19, MVT^{+}17, MGJH18, NLM^{+}12, NFW13, NCT^{+}15, OWFS11, PCO^{+}15, RKG^{+}11, RSTP12, RSN16, RVvdP^{+}17, SKLG10, SBvH^{+}15, SFB12, SKKV11, SPG^{+}11, SVG^{+}18, SYW18, SSC^{+}17, TCG^{+}17, TEZ^{+}18, WM12, WRWPG19, WXMS10, WTC^{+}17, WJHS18, WSB^{+}13, YJO^{+}19, ZZAC13]. Compositional [SLC^{+}16, BWBB15]. compositions [CFD15, CPHD15, KFP^{+}18, PMA18]. compound [HOD^{+}17]. compound-specific [HOD^{+}17]. compounds [DTL^{+}19, DJD^{+}14, GRPB^{+}17, JZZY18, TWWY18]. comprehensive [RASD10, WFB^{+}11]. computation [KLEH16]. Concentration [CHPH13, MWBM19, BLG^{+}15, GC16, GJWS14, GJWS16, HWZ13, HSTK15, Lee18, PSG^{+}16, RNK^{+}16, RSN16, SZH^{+}10, SMG12, ZF17, ZTS13, ZMS^{+}18]. Concentrations [HKS^{+}15, TKK^{+}17, ADS^{+}17, AAC^{+}19, BBJ^{+}19, BRS^{+}13, CKB^{+}16, FNSS15, GMBL16, GNHGM13, HKU^{+}10, KHK^{+}19, ML19, MKG^{+}15, PHG13, RR13, RMNZ12, SES18, SLP^{+}14, TAV^{+}10]. concept [GMJW13]. concerted [BVSR^{+}15]. conchilega [BBR^{+}14]. concordance [FSST11]. condition [BRNS18, LBR^{+}13]. conditioned [SGME11]. conditions [ANP^{+}14, ASR^{+}17, BHB^{+}12, BSBK13, DBA16, FVSL19, GPCJ16, GAH11, GWD^{+}16, HRPW15, IGP^{+}12, KIH^{+}15, KHH19, LG10, MKB^{+}19, MU17, NCT^{+}14, NLHAA^{+}17, PCY^{+}10, RLC^{+}11, SHD^{+}11, VTH^{+}18, VFS^{+}15, WDH^{+}17, WBB^{+}17, ZSM14, dCGS19]. conducted [UFW^{+}18]. conduit [PMY^{+}19b]. conduit- [PMY^{+}19b]. conduits [NZH^{+}11]. configuration [JLR^{+}17]. confirmed [ZXM^{+}11]. Congo [HSC^{+}11, SSC^{+}10, WMBR13]. Congruent [PHDH14]. conjunction [NCT^{+}14]. connection [GGC^{+}14]. connectivity [AWG^{+}12, BCDR^{+}19, Car10, KPP^{+}18, NG13, OMSC13, RNG^{+}13, SBB^{+}18, SS19, WMT^{+}12]. Consequences [AdGAD14, GPL11, MCWB10, BL13, HL13, VMC^{+}13]. conservation [MTU18]. consideration [SM11b]. Consistency [WTC^{+}17]. Constance [BSSW11, WBS^{+}10, WP14]. constants [CLLH14]. Constrained [GBP^{+}12]. constraints [AAO^{+}19, HJT^{+}13a, HJT^{+}13b, MHH^{+}17, SMC^{+}10]. constructed [EED10]. construction [KTH^{+}19]. consumer [BH13, CJWS15, LGV13]. consumer-resource [BH13]. consumers [BLJ13, CWF11, DRE^{+}10, KBA^{+}12, KBA^{+}14, MDF^{+}14, WSUC^{+}18, WKAM^{+}19]. Consumption [HGT^{+}18, BPB^{+}17, CBP10, FWFB10, HDK^{+}12, KBE^{+}17, LALM16, LALGM18, MMN^{+}10, SRCL^{+}13, UFW^{+}18]. consumptive [MHA^{+}18]. contaminated [BHB^{+}19]. contamination [GWN^{+}12]. contemporary [GBS17]. content [BWS^{+}14, CBF11, CFB14, FLLH18, JWGH19, LLB17, ORC^{+}17, TW10a]. contents [GAM^{+}19]. context [MHRH11, RG19]. Continental [CBP10, BRR^{+}13, BK11, GFDC11, HDK^{+}12, HGM10, JAZ^{+}10, JBT11, KSFT13, LPLH18, LDT^{+}11, MRBR10, MBAS^{+}17, NTK^{+}18, WS13, WDL^{+}17]. continuous [GBL13, MSS^{+}18, MFK^{+}13, SGRB10]. continuum [BSM17, CAS^{+}17, FEC^{+}16, GKT^{+}15, WC17, WWS11, XDK^{+}17]. contrasted [BPW^{+}19]. Contrasting [BHB^{+}19, BAG^{+}17, BHV^{+}17, JMJ^{+}19, LEN^{+}15, LZR^{+}17, RSJ^{+}18, SNvD^{+}10, TW11, dBWL^{+}13, ASK^{+}11, BBK^{+}15, BBQ^{+}10, CUW11, LABJ18, LALM16, MKW^{+}19, OSC14, RAB^{+}17, SSJR^{+}10, TBHM^{+}13, WZR19]. contrasts [XDK^{+}17]. Contrib [Ano17g, Ano17h, Ano17i, Ano17j, Ano17k, Ano18h, Ano18i, Ano18j, Ano18k, Ano19j, Ano19k, Ano19l, Ano19m, CESC19, GBB^{+}19c, KSTA18b, LF17a, SHT^{+}18, ZXZ17a]. contribute [HBM^{+}15, OCLW11, PHG13, dBWL^{+}13]. contributes [WA14]. Contribution [KZB^{+}10, SL10b, UVGS10, XSAHV13, ZCZ^{+}18, BCF^{+}17, EMS16, HLG15, QS19, SWM^{+}18, Scu16, TDF^{+}17, VSdG17]. Contributions [KOFN11, NXL^{+}18, TTV^{+}13, BAY^{+}14, DNH^{+}18, HDDH^{+}17, KPW^{+}11, MBTK18, MGW^{+}13, OWM^{+}18, RMF11, VBBR17, WCJ16, WCP^{+}15, XDC^{+}19]. Control [MAV^{+}13, SPPS10, AdBVA10, BH13, BHB^{+}19, BSFH10, CGB^{+}18, DDF^{+}10, HYK^{+}15, HMV12, JZZY18, LDT^{+}11, LDL^{+}19, MTM^{+}16, Meh10, PvDM^{+}13, SM11b, THH^{+}13, UA10, WCM19, XPQ^{+}10]. controlled [BPRG^{+}18, LZK18, LBR^{+}13]. controlling [ASH^{+}14, ERA^{+}12, GLMG15, KBH^{+}19, PSH^{+}11, YHS^{+}17]. Controls [CRCGG^{+}17, HC12, WLG^{+}16, BLW15, BGR14, CFF^{+}17, EKS^{+}18, FUS^{+}16, FRA^{+}17, FLP^{+}10, GJWS14, GJWS16, HHHT19, HBCK10, KEH^{+}14, KCB^{+}17, KBM^{+}14, LH17, MFMC^{+}10, NAH^{+}11, PKB^{+}17, RCH^{+}15, RNG^{+}13, RETS16, SLK^{+}14, SK19, SBS^{+}13, SMG12, TSSH19, VCM13]. Convection [SCR^{+}12, Aus19, SBe10, TF11, vH19]. Convection-driven [SCR^{+}12]. convective [VAH11]. cooling [VAH11]. Copepod [GGL^{+}15, TGG^{+}11, TWWY18, AACS11, BPPF12, BD15, CWF11, CFB14, DHK11, EHT10, FNSS15, GPL11, GOD^{+}18, HHW^{+}19, HBBM19, JLG10, JLG11, KSY11, LRY12, NG13, PLE^{+}17, RBI^{+}10, SGCI14, SMA13, SM11a, SNTK15, SPR^{+}15, TW10b, WB19, WD15, ZTS13]. copepod-mediated [PLE^{+}17]. Copepoda [HAL17, PPT12]. copepodamide [GBB^{+}19a]. Copepods [LKK13, AvSGK18, BAB^{+}16, CBF11, FOT^{+}15, GK15, vSGAK17, HBCK10, KGC^{+}16, KJKS18, MTU18, MSAM18, NTI^{+}15, PJ16, SSFF12, STCS10, SGVR16, SMC^{+}10, TNI19, TIS^{+}13, TAE^{+}18, TSK13, VFME18, VIS^{+}13, WFR10, XNK18]. Coping [SSP^{+}18]. copious [CHV^{+}17]. Copper [AMMH^{+}13, WA14, Alo17, HNZ^{+}16, JKKM13, LBHS13, MBC^{+}18, MTW12, RLSC^{+}13, STB^{+}16, SMW^{+}18]. Copyright [Ano19d, Ano19e, Ano19f, Ano19g, Ano19h, Ano19i]. coral [CHH^{+}17, CNL^{+}15, CRB^{+}17, CGP^{+}19, CPPdAR^{+}13, CESC13, CESC14, DJD^{+}14, DSM^{+}18, Edm11, Edm15, ETI^{+}16, FZL^{+}14, FPPA^{+}11, FPGR^{+}13, FDH^{+}14, GDD^{+}16, GJR^{+}19, GBR14, HBD^{+}11, HRG^{+}15, IPGP10, KHPIP^{+}14, KCL^{+}14, KTH^{+}19, LABJ18, LSD18, LGC13a, LGC13b, MKB^{+}19, Man10, MBHG11, MPSA17, OBL^{+}19, PCD^{+}19, PJFJ^{+}15, RDC^{+}19, RMK^{+}16, SLC18, SPTS15, SIW^{+}11, SHD^{+}11, WGDAA19, WHD10, WTC^{+}17, WFL^{+}12, WLHW13, YKT^{+}15, YLH^{+}16]. coralline [CHPH13, GHSR^{+}16]. Corals [PST^{+}13, Ano21c, BJDMH10, CRS^{+}17, ELJ^{+}16, GFPSG13, JLRK12, Les19, LCBC16, NFRU11, PGRR^{+}19, RBRH10, TIN^{+}14, TFH17, TLB^{+}16, TEGL11, WRWPG19, WPL^{+}14]. core [GKS12, WRB^{+}19]. Coregulation [XXZ^{+}19]. coring [HMFB16]. correlates [RBI^{+}10]. correlating [CJC^{+}12, LGR^{+}12]. correlation [AGMR14, BHW^{+}12, BLH^{+}13, HBM11, RBM14]. correlations [BISZ17]. Corrigendum [Ano21a, Ano21b, Ano21c]. Coscinodiscus [QFH18]. cosmopolitan [SGCC16]. Costa [ANP^{+}14, GRSD^{+}14]. Could [HLFM^{+}10]. count [MTT17]. Counter [KS16, HBD^{+}11]. Counter-measures [KS16]. Counteracting [ZEXH15]. counteracts [HBB^{+}11]. Coupled [BCF^{+}17, IH18, MHL^{+}16, WCM19, XLS^{+}19, FSBT16, OVRJ13, RSG11, RRB^{+}16, SDS^{+}11, VCPC^{+}16, VPG^{+}19]. couples [CMK^{+}10]. coupling [BCVAn10, BSY^{+}16, DMSHC16, FMGR^{+}11, HC10, KVMA17, MBLD15, RKTLM18, SAS^{+}11, SHT^{+}17, SLA^{+}18, WS18, WS13, ZCY^{+}15]. course [WCJ^{+}15]. covariance [KB15, PSB^{+}16, YKT^{+}15]. cover [HKS^{+}15, KFJ13, LLH^{+}15, NHP17, RMK^{+}16, RVvdP^{+}17, SKV^{+}19, SAPI14, VLWV14, WCB^{+}10, YJO^{+}19]. covered [CDW^{+}16, CMS17, DTKMK15, FLPL13, HGD14, JLR^{+}17, MKLKP16, RKL14, SSS^{+}16, SPO^{+}18, SMA15, WCB^{+}10]. Cr [WZR19]. crab [LDCT11, MCT^{+}14]. crabs [KLM^{+}17]. Crassostrea [BHW^{+}12, BMC^{+}16]. craving [MAFCD^{+}18]. created [BSA^{+}16]. creation [CZB^{+}18]. Creek [CMW^{+}19, AC17, DTM18, MSGS^{+}13, SML^{+}19]. creeks [ALG^{+}13, DCCB17]. Crenarchaea [FMGR^{+}11]. Crenarchaeol [PWS^{+}11, ZKMT^{+}13]. Crepidula [NBDM16]. criterion [TF11]. Critical [LH17, KSY11, ZF17]. critically [HATF17]. Crocosphaera [BWP^{+}10, GFH13, MFK^{+}13]. Cross [LGV13, TDS^{+}10, WBG^{+}16, CAQS16, HCC^{+}13, MH16, MAF19]. Cross-ecosystem [LGV13, MH16]. cross-scale [CAQS16]. Cross-shelf [WBG^{+}16, HCC^{+}13, MAF19]. Cross-species [TDS^{+}10]. crucial [ZPK^{+}12]. Crustacea [MXWC11]. Crustacean [CHV^{+}17, AA18, BPL^{+}19a, CCK^{+}12, ITO^{+}17, ŚF19, ŠSP17, WVV^{+}11]. crustaceans [WVGB10]. Cryptic [PFJ10, PTS12, XBR^{+}18]. Crystal [CF10]. ctenophore [CMG^{+}15, JTH^{+}11, JCS^{+}18]. cues [BMPF19, BSH16, HJMD13, LWE^{+}19, SBDS^{+}15]. Cultural [AMQ^{+}11, SDH^{+}14, BHC13, BHC14, MLS^{+}14]. culture [BRR^{+}13]. cultured [BS18a, SES18]. cultures [BMM^{+}13, BWP^{+}10, DSM^{+}18, ESMS13, FYT^{+}12, GBL13, HLSW^{+}15, MFK^{+}13, SLC^{+}16, WCV^{+}12]. Cumberland [BAY^{+}14, BCF^{+}17]. Current [BTC^{+}19, TSC^{+}19, BLH^{+}13, DIC^{+}18, GK15, KGC^{+}16, NSO19, WZTK15, XNK18, BPPF12, BSB^{+}10, BBB^{+}14, DLP13, MBC^{+}18, MS13, PMPD13, SLBNG11, SSGM18, TT14, uGH^{+}11]. current- [WZTK15]. Currents [Ano17l, RMDK10, HSR15, RDZ^{+}13, ZXZ17b]. curtain [VHM^{+}10]. curve [KTL17]. CUSUM [LGR^{+}12, CJC^{+}12].

CUSUM-transformed [LGR^{+}12, CJC^{+}12]. Cyanate [WM17, WMM18, KP13]. Cyanobacteria [dKYH^{+}12, ABB^{+}14, BH16, BVSM15, BWP^{+}10, CJ17, FLLH18, GCSO14, GPH^{+}13, GWSEA10, JLC^{+}15, KWGN^{+}10, LEK^{+}18, MGL^{+}16, MLK11, MDSG18, QHVM18, RCIB14, SWZ^{+}15, ŠSP17, VTH^{+}18, VSP^{+}11]. Cyanobacterial [WCV^{+}12, BPGE13, JHLK^{+}19, KP13, MQP^{+}16, OSB^{+}15, SK19, WSTD10, ZLLM10]. cyanobacterium [FWvD^{+}18, HS18, MZB^{+}15, SJM11]. cyanophages [ŠSP17]. cycle [BMD17, CSME13, DdD^{+}10, FMM^{+}14, GYP^{+}18, HTL^{+}18, JKKM13, KMF10, LCH^{+}14, NPT11, RR12, SFI^{+}18, SBKO18, WRH^{+}18]. cycles [GAH11, HPM^{+}10, HLFM^{+}10, MAFCD^{+}18, NXL^{+}18, OMSC13, PCF14, TIF^{+}15, TKB18, VBGG^{+}13, WMC^{+}15]. cyclic [BvBB^{+}16]. cycling [BHB^{+}19, BHV^{+}17, BGM^{+}13, CSME13, DBV^{+}11, DKSA19, DvOR^{+}16, EED10, FPG11, GLKK10, GRE^{+}16, HGG^{+}17, HCW^{+}10, HCLS11, HGM10, KGRV18, KCL^{+}14, KBL^{+}10, KOFN11, LH17, LK14, MTSG18, MBH^{+}15, MBC^{+}18, MGS12, NUH^{+}12, OCLW11, OVRJ13, PBA^{+}15, PPPA14, PGP^{+}14, SLBNG11, SH11, SCP^{+}16, SSS^{+}19, TK12, WLG^{+}16, WBB^{+}17, WZBW^{+}11, XBR^{+}18, ZXN^{+}12, ZMWM11]. cyclone [LWE^{+}11]. cyclones [AJC15]. cyclonic [FLPL13]. cyclopoid [AACS11, SGCI14, ZTS13]. Cyclotella [RSE^{+}17]. Cylindrotheca [BC19]. Cymodocea [IOB^{+}11]. cyst [BRF^{+}17]. cysts [BPL^{+}19a, MAC^{+}10]. cytometer [FPP^{+}19]. cytometry [AJC15]. Czech [KKP^{+}19].

D [CLB19, WP14]. D. [PMP^{+}12, SZH^{+}10]. D1 [HBD^{+}11]. Dactyliosolen [TJJ^{+}15]. daily [AEH19, FPGR^{+}13, SBR^{+}13, TGGZS^{+}10]. dam [FDB^{+}15]. damicornis [WHD10]. damming [RPH^{+}10]. damped [ILPL13, SI10]. Dampened [PLE^{+}17]. Danish [àNTS13]. Danube [MD15, SDS^{+}16]. Daphnia [BMPF19, HMV^{+}18, HL13, HSR^{+}10, HNL^{+}13, IWF19, ITO^{+}17, JSFC18, LLL10, LWE^{+}19, LJ18, LGW^{+}19, MNW^{+}19, MCWB10, NBSMN19, PMP^{+}12, PDP^{+}10, PTS12, PWF16, PWF18, RKG^{+}11, RG19, RWF^{+}12, SBvH^{+}15, SZH^{+}10, ŚF19, SOM^{+}15, SW11, TYX^{+}19]. Daphnia-associated [HNL^{+}13]. Dark [VHV10, SDCF16, VvO11]. darkening [SLK^{+}10]. Darkness [OPZ13, CGL^{+}16, SSC^{+}10]. data [BPGE13, BM16, Bre10, BCRW15, CJHR19, FSST11, GK14, INF12, LLW^{+}18, LC12, SMA13, SPO^{+}18, VdSLC^{+}16, WCM19, WSTD10]. database [Lan14]. databases [Lat14]. daughters [LOS12]. davisae [AACS11, SGCI14, VIS^{+}13, ZTS13]. Day [BPL^{+}19b, SSH^{+}14, RKBA14]. daylight [AAO^{+}19]. daytime [HH14]. Dead [EHT10, KOFN11, SSFF12, ASL16, ABS^{+}19]. death [FAF^{+}12]. debris [WLL^{+}11]. Decadal [MMHT10, AAIA14a, AAIA14b, GMGM^{+}13, HPS^{+}10a, LH19, WTC^{+}17]. decadal-scale [GMGM^{+}13]. decade [CB19]. decades [MTU18]. decipher [NTM^{+}10]. Deciphering [JHLK^{+}19, SKK^{+}15]. decline [AAIA14a, AAIA14b, AA18, CBS^{+}17, MTK^{+}17, SPTS15]. declines [SM11b]. declining [BWS^{+}14]. Decomposition [CPG^{+}10, CA08, HEH^{+}17, SH10a, SCP^{+}16]. decoupled [WKAM^{+}19]. decoupling [DT16, ZSZ12]. decrease [CCV^{+}18, GTPB^{+}11, KRR16, LdlSB^{+}12]. Decreased [SK19, SHD^{+}11]. decreases [CGP^{+}19, GSB^{+}17]. Deep [BHD^{+}17, KCL^{+}14, SWE^{+}18, SWZ^{+}15, ÅCA^{+}18, ÁSNCÁ^{+}13, Aus19, BNW^{+}14b, BCRW15, CPOMA15, DKG15, DvOR^{+}16, FYC^{+}18, GVS^{+}10, HCLS11, JMM14, JTG^{+}11, KYC^{+}15, KCM^{+}10, LCM^{+}17, LBC^{+}18, LBS17, LKF^{+}18, LCM^{+}12, MSSH12, MAC^{+}10, MKBSK19, nVOH12, PMRRA19, SPFP11, SOM17, SSB^{+}18, SLBH^{+}19, SVG^{+}18, VLDM19, VMI13, WP14, WCCP14, WSB^{+}13, ZWL^{+}14, ZXL^{+}19, FCRW^{+}16]. deep-sea [GVS^{+}10, MKBSK19, SLBH^{+}19, SVG^{+}18, ZXL^{+}19]. deep-water [JMM14, KYC^{+}15, LKF^{+}18, WP14]. deepening [SGJB14]. deeper [Ano21a, GSG^{+}17]. deeply [AdBVA10, LK14, LZK18]. defence [LG10]. defense [GMJW13, HCL^{+}18, KS13, RKLH11]. defenses [Rie15, RG19, SBFC18]. defensive [GBB^{+}19a]. deficiency [GHS14]. deficient [CRJ^{+}14, ORC^{+}17, WFK^{+}16, WMM18]. define [WDMF13]. defining [ITO^{+}17, JSB^{+}14]. Degradation [BPV^{+}19, NTI^{+}15, RZW11, BVSM15, CKP^{+}15, CCC10, HKP^{+}16, JP10, KHCH14, MKW^{+}19, MBLD15, MW15, NTM^{+}10, PSNE15, WE19, YWY^{+}15]. degrade [FPD^{+}10]. degraded [PJFJ^{+}15]. degrading [TMK^{+}13]. Degree [MFMC^{+}10, RKBA14, SZH^{+}10]. degree-day [RKBA14]. degrees [LCBC16]. Delaware [FYVU17, Sha10]. delayed [BH16]. delicatissima [LBHS13]. delivery [WGH^{+}10]. Delta [RKWH18, LYL^{+}17, TT12, TLG^{+}11]. Delving [Ano21a, GSG^{+}17]. demand [HDP15, SWP11]. demands [vOSH12]. demersal [FCD12]. demographic [Edm15]. demography [KKB^{+}18]. demonstrate [TLG^{+}11]. demonstrating [ZS18]. Denitrification [BSN^{+}14, BKD^{+}16, GCR^{+}10, RDB^{+}18, BGR14, DT16, DTFR12, DSS^{+}11, EMO^{+}11, EMS16, GFT^{+}14, GGL^{+}15, GSPM13, GKS12, KGC^{+}12, KJG10, MHL^{+}16, MGL^{+}16, MAS^{+}16, NTK^{+}18, RWC16]. denitrifiers [WBZ^{+}13]. denitrifying [MTW12]. Denmark [JP10]. dense [ANP^{+}14]. densities [MG14]. Density [JLG10, CNL^{+}15, KSY11, MTT17, Meh10, NLM^{+}12, PHDH14, SVG^{+}18, TLB^{+}16]. Density-dependent [JLG10]. depend [PSG^{+}16]. Dependence [LK15, ASW^{+}19, BPB^{+}17, BSCC15, BYD19, BISZ17, CBFK19, DOD10, Fie13, FMGR^{+}11, GRGL^{+}13, PH15, RKBA14, RM14, Tad10]. dependency [KWGS18, RRAS17]. dependent [CEES14, CCK^{+}12, CdC^{+}11, JLG10, JBPM15, KHTO13, RCIB14, RETS16, SSPK^{+}12, TBSL17, WBZ^{+}13, WSTG18]. depending [SHL^{+}18]. depleted [Bre14, DKSA19, FDBW16, NCT^{+}15]. depletion [KKB^{+}18, LCCF10, PHG13, SSB^{+}18, SYW18, WBG^{+}16]. Deposition [EMH12, AWG^{+}12, ACW^{+}18, BAA^{+}13, KK13, KHVS11, LYH17, MKG^{+}15, RQC^{+}15, vdJFS^{+}18]. depression [Les16]. deprivation [IGP^{+}12]. Depth [GEC^{+}17, LPO^{+}11, NFW13, SMM11, AAO^{+}19, CR11, DJS18, HONR11, Kir13, LSDW18, MMGO^{+}17b, MSK^{+}17, OMSC13, PS17, SVLS^{+}16, SAPI14, SLS^{+}11, WGJ^{+}19, WCJ16, WTC^{+}17, YWL^{+}17, vEG10]. Depth-integrated [SMM11]. depth-tiered [OMSC13]. depths [GLS^{+}13]. derived [BA14, BMPF19, BBTK^{+}16, CPG^{+}10, GAM^{+}19, HNHS^{+}15, LRG16, MSGS^{+}13, OCB^{+}18, QHVM18, RPMK17, SBC^{+}17, TSDF^{+}16, WCJ^{+}17, WGC^{+}13, ZD18]. desalination [SJ11]. desaturase [YLJ11]. describe [MLS^{+}14, SP11]. describing [HS10]. deserves [SM11b]. design [GBMG12]. desorption [RSM13]. destabilizes [LS15]. details [vH19]. detect [GK15]. Detecting [ZD18]. detection [BOT^{+}15, KYG^{+}12, PGB^{+}19]. Determinants [FCC11, DDH^{+}19, UIY^{+}11]. determination [GLKK10]. determine [BRNS18, CG17, CLWD13, HMV^{+}18, RASV^{+}17]. determined [AWG^{+}12, CR11, KB15, Kus14, SW11]. determines [AvSGK18, DPLG^{+}19, GBD^{+}10, SPHVA19]. determining [LFGK10, WJHS18]. detoxification [HHW^{+}19]. detrital [WLS^{+}11]. detritivorous [SWP11]. detritus [BPV^{+}19, DRE^{+}10, FGBS^{+}18, KBL^{+}10, OCR10, RPMK17, SGRB10, ZLLM10, dBWL^{+}13, vOSH12]. detritus-based [DRE^{+}10, SGRB10]. Developing [WJHS18]. Development [SBNC^{+}19, SSPK^{+}12, DL11, JMNG^{+}13, JMN15, KBE^{+}17, MMJ^{+}12, RKMN^{+}13, SKGT17, TIF^{+}15, WXF^{+}15]. developmental [AACS11, RG19]. diagnosed [FWWF18]. dialogue [KM10]. diapause [CBP12, HTL^{+}18, STCS10, ŚF19, TNI19]. diapausing [All10b, BPL^{+}19a, PT11]. Diaphanosoma [PZHD18]. diapycnal [ASL16]. Diatom [FPSL18, MD10, WCJ^{+}17, WHH^{+}11, AJ15, BFW^{+}13, BC19, BTJ^{+}12, BMM^{+}13, BHV^{+}17, BBTK^{+}16, CR16, CEB^{+}17, DdD^{+}10, DMB^{+}12, DBC^{+}13, HV19, HHW^{+}19, HBB^{+}11, KBL^{+}10, KBVW12, LBHS13, MDB19, MTSG18, MBTK18, MEM^{+}17, NTM^{+}10, PRS^{+}18, QFH18, RASD10, RLSC^{+}13, Sch19, SLH^{+}15, SLG^{+}14, SHF^{+}11, SH11, TBLG14, VvO11, WHR18, YKBJL12, ZD18]. diatom-aggregates [MTSG18]. Diatom-produced [FPSL18]. Diatoms [LOS12, CSJ^{+}14, CMS^{+}18, MRBR10, MEM^{+}17, MPvBS^{+}18, NTA14, PHB^{+}10, RASD10, SHKU11, SS17, SYW18, TJJ^{+}15, TNK^{+}14, WCI^{+}14]. Diazotroph [BBTK^{+}16, SFI^{+}18, GFH13]. diazotrophic [ABB^{+}14, MDSG18, SWZ^{+}15]. diazotrophs [BAA^{+}13, CLJ^{+}19, MBBG^{+}12, SKK^{+}13]. DIC [MZH15]. die-off [SKGT17]. dieback [KKP^{+}19]. Diel [DHG^{+}17, EHW^{+}15, FOT^{+}15, GAH11, OMSC13, OR16, RGLM^{+}12, SC10, WMC^{+}15, ZXN^{+}12, BA14, BM16, CSME13, GLF17, HSR^{+}10, HC10, HPS^{+}10a, HH14, MWSB18, PGB^{+}19, RRCH^{+}19, RK13, SNM11, SWD11, WFB^{+}11, vBBM^{+}19]. Diet [BLJ13, BCF^{+}17, HSTK15, KNA^{+}14, MTEM15]. Diet-tissue [BLJ13]. dietary [LJ18, SMC^{+}10]. diets [DPLG^{+}19, LRS^{+}10, RMF11]. differ [SHF^{+}12]. difference [CL17]. Differences [DRP^{+}17, EMO^{+}11, GKT^{+}15, LSHK11, RRCH^{+}19, SWP11, GLS^{+}13, LLL10, LGV13, MHA^{+}18, PBA^{+}15, PRS^{+}18, SYdTP^{+}11, SI10, TW10a, WC17, WHR18]. different [ALdML^{+}14, ASSG12, BBS^{+}18, FWFB10, FEC^{+}16, GMD11, GN16, HRPW15, LFB^{+}10, LG10, MF19, MG14, MU17, MCT^{+}14, MZB^{+}15, RF13, SPFP11, SNG^{+}14, TPM^{+}14, XDC^{+}19]. Differential [BWD^{+}11, BWD^{+}12, HRPW15, SMR^{+}17, SSGL19, TNK^{+}14, VAH11, WFK^{+}16, BJF18, ML19, MP17]. differentially [DRE^{+}10]. differentiation [TB18, WOC^{+}18]. differently [LBHS13]. differing [HSTK15, ZZY^{+}10]. differs [HS11]. diffuse [PMY^{+}19b]. diffusion [ASL16, DBSP^{+}16, JJ17, Kus14, MCGF^{+}11, SSW19]. Diffusive [BMF^{+}16, CKB^{+}16, SBe10, TBK15]. digitata [HONR11]. Dilution [CPF16, CLLH14, Lan14, Lat14, LFL17, SMMF19, SNM^{+}15]. dim [SLS^{+}11]. dimensional [HSLH^{+}14, HE10, LWE^{+}19, MMFBB18, OBI12]. Dimethyl [ARW^{+}10, ZYZ19]. Dimethylated [GRPB^{+}17, DJD^{+}14, JZZY18]. dimethylsulfide [GRGL^{+}13]. dimethylsulfoniopropionate [ASA^{+}18, FAF^{+}12, RLL^{+}10]. dimethylsulfoxide [TKK^{+}17]. dinitrogen [CJW^{+}19, DHW11, GKS12, MFK^{+}13, MBBG^{+}12]. dinoflagellate [BFW^{+}13, BVSR^{+}15, BRF^{+}17, HST^{+}14, HLG15, HLSW^{+}15, JLG10, JLG11, LKLH10, NAH^{+}11, SBDS^{+}15, VdRA^{+}19]. dinoflagellates [Les19, MAFCD^{+}18, PK14, SFWP12, XNK18]. Dinophyceae [BWD^{+}11, BWD^{+}12, BVSR^{+}15, KMF10]. Dinophysis [HLG15]. dioica [LTPK^{+}18, LBR^{+}13, LSK11]. dioxide [BHW^{+}12, CF13a, CCW^{+}19, GFH13, GWB^{+}14, Joh10, KHTO13, SLH^{+}15, SMG12, UIY^{+}11, VFS^{+}15]. Diptera [REDW10]. Direct [BPB^{+}17, GLKK10, HC10, HCK11, NL14, SBA^{+}11, JTH^{+}11]. discarded [NTI^{+}15]. Discharge [DPG^{+}12, BBLN11, DB11, GBD^{+}10, KDGL19, KKH11, KSG^{+}10, LKS^{+}16, LSH^{+}17, LCH^{+}14, LSD18, MT11, OBL^{+}19, PVLMT^{+}16, RDP^{+}17, RGM15, WGH^{+}10]. discharges [PLS^{+}16, VLMTEW11]. discontinuous [KMC^{+}15]. discrepancy [BMW10]. discrimination [KLM^{+}17, MTEM15, MZH15]. discriminatory [AC15]. disease [JSFC18, PHCD14, SLG10]. Disentangling [SLHA19, SKKV11, VABMS^{+}12, ZKMT^{+}13]. disequilibria [SBNC^{+}19]. Disko [FGMN17, HNSM12]. dislodgement [dBWL^{+}13]. Dispersal [RLPL14, GBMG12, KTK^{+}13, LDCT11, MVT^{+}17, MCT^{+}14, NG13, OIS10, PHJ12, PBV16, RNG^{+}13, RCJ15, RMLVK12, RPL16, WJHS18, uGH^{+}11, vHOM^{+}19]. Dispersion [SGA10, CTH15, HSR15, MCC^{+}10, OrIA10, PHJ12, PH15]. displacement [GPH^{+}13]. Disrupted [LH17]. dissects [LGW^{+}19]. Dissimilatory [DSS^{+}11, RRB^{+}16, LTH^{+}12, MAS^{+}16, RvSM17, RDB^{+}18, TG17]. Dissipation [SLPM15, NBG17, SWL11]. Dissolution [WLR17, KWM^{+}19, KNL10, SEYJ11]. Dissolved [AHJS15, BAG^{+}14, BHD^{+}17, CFD15, GJWS14, GJWS16, HTLM18, HBM11, LGC16, NO17, OALD10, OWFS11, OVRJ13, SMG12, AFG^{+}16, ÁSNCÁ^{+}13, BSCG17, BBLN11, BVSM15, BDU^{+}19, BBB^{+}14, BLWV10, CRCGG^{+}17, CSÁS^{+}10, CÁSO^{+}16, CKP^{+}15, CPG^{+}10, CDA16, CHV^{+}17, CK12, CK13, CF10, DFWPK16, DVC^{+}17, DBA16, Dem19, DPG^{+}12, EO13, EKS^{+}18, FYC^{+}18, FUS^{+}16, FHS10, FPG11, FB12, FEC^{+}16, GdG11, HKP^{+}16, HEB^{+}19, HSTK15, HMH^{+}16, HGT^{+}18, HSC^{+}11, IH18, JTH^{+}13, JSK^{+}15, KWRS13, KBT16, KHK^{+}19, LCW17a, LF19, LG16, LZC^{+}14, LTX^{+}17, LBR^{+}12, MWBM19, MSGS^{+}13, MGHS18, MPONC^{+}17, MA18, MBAS^{+}17, MdBKL13, MCC^{+}10, MSD^{+}14, MGSM10, MGJH18, PRS^{+}18, PCO^{+}15, PML^{+}19, RR13, RM14, RCH^{+}15, RvSM17, RCSÁS^{+}10, RHDTS^{+}11, SLC^{+}16, SHSK14, SCF^{+}15, SCR^{+}12, SRCL^{+}13, SLP^{+}14, SLA^{+}15, SFB12, SHK13, SFLB16, SBC^{+}17, SSC^{+}10, SYW18, TLG^{+}11]. dissolved [THH^{+}13, TAV^{+}10, TTV^{+}13, TKK^{+}17, TZD^{+}15, UVGS10, WM12, WDX^{+}11, WLG^{+}16, WMC^{+}15, WCJ^{+}15, WSM^{+}19, WYW^{+}10, WDL^{+}17, WSTG18, WSB^{+}13, XSAHV13, YHS^{+}17, YJO^{+}19, YMB^{+}18, ZZY^{+}10, ZHN^{+}10, ZZAC13, ZCK^{+}16, dCGS19, vEG10]. Distance [FYVU17, BMB^{+}18]. Distance-based [FYVU17]. Distinct [BBB^{+}14, HRMD19, OCLW11, ANP^{+}14, CR16, CLFW17, EMS16, FYVU17, GRSD^{+}14, RDB^{+}16, SGS18, TLG^{+}11]. Distinctions [LRS^{+}10]. Distinguishing [BRS11, LBR^{+}12]. Distribution [BAG^{+}14, KP13, LCM^{+}17, MMH^{+}18, WBS^{+}10, ZTW^{+}11, ASH^{+}14, BHS^{+}16, BHB^{+}12, BSRP^{+}12, BSSW11, CR11, CUW11, DML17, DB11, DDH^{+}19, DTKMK15, DKK^{+}14, EHW^{+}15, FDS^{+}18, GPCJ16, GBD^{+}10, HCD19, HPM^{+}10, HONR11, IPGP10, JZZY18, KHP18, LMR14, MRSS12, MSAM18, NFW13, NO17, PZHD18, RS16, RPMK17, RSN16, RWF^{+}12, SPP10, SPFP11, SP11, SLA^{+}18, SS17, Tho19, UMHH^{+}14, VPMrI12, WMC^{+}18, WM17, WMM18, YHS^{+}17, vdHHC^{+}19]. distributions [AAO^{+}19, BTC^{+}19, BLLB12, CG17, KYRMD18, KT13, LCW17a, LRS^{+}10, RWM^{+}14, SSG^{+}17, SYdTP^{+}11, SHK13, SH10b, WFK^{+}16, ZYZ19]. districts [AJG13]. Disturbance [KHH19, PH13, AP12, CZB^{+}18, KJKS18, SVG^{+}18, WVV^{+}11]. disturbances [PRL18]. disturbed [KYG^{+}12]. disulfide [ZYZ19]. Diurnal [QHVM18, SMN^{+}15, FWWF18, LCW17a, SSU^{+}16, VBGG^{+}13]. dive [WHAM15]. divergence [MNW^{+}19]. Divergent [RG19, LH19]. Diverse [KTRK11, HLGA17, LF16, LF17b, MH16]. diversification [MXWC11]. Diversity [DTKMK15, GWB^{+}14, GBC^{+}17, MVT^{+}17, RDB^{+}16, WCV^{+}12, BDS11, BL13, CVS^{+}10, FMP^{+}13, GYP^{+}18, HT17b, JHLK^{+}19, JB19, LZR^{+}17, MNW^{+}19, MVL^{+}10, OPA^{+}14, PTS12, RBI^{+}10, SJ11, SS12a, SS17, TBLG14, VMAS^{+}16, VB17, WKAM^{+}19, WHR18, XZC^{+}16, ZCL^{+}19, ZXL^{+}19]. DMS [ADS^{+}17, GRPB^{+}17]. DMSO [ADS^{+}17, BRS^{+}13]. DMSP [ADS^{+}17, BRS^{+}13, FAF^{+}12, ML19, RLL^{+}10]. DNA [HNL^{+}13]. DNRA [BKD^{+}16, KWM^{+}19, KJG10, MAS^{+}16]. Do [CEPPR14, GPCJ16, SBFC18, BB10, CESC13, PMP^{+}17, WRH^{+}17]. DOC [DVC^{+}17]. documentation [WKJS^{+}14]. Does [CLHL12, HBBM19, HBCK10, PD11, SP11, SS19]. doi [Ano21a]. doliolid [TIS^{+}13, TIF^{+}15]. doliolids [TIS^{+}13]. DOM [AGCA16, HKP^{+}16, YJO^{+}19]. domain [SLHA19]. dome [ANP^{+}14, GRSD^{+}14]. Dominance [ABB^{+}14, HCK10, GNWDL19, MM11, RRD14, SLK^{+}10, VSP^{+}11]. dominant [CTA^{+}19, GHSR^{+}16, MMB17, XFH14, vdHHC^{+}19]. dominate [Clo18]. dominated [ARML10, GLF17, GLF18, JMJ^{+}19, KSG^{+}10, PMY^{+}19b, PHPH^{+}16, PLE^{+}17, RMH^{+}17, WDCH18]. dominates [DSS^{+}11]. dominating [MPAS17]. domoic [HHW^{+}19, SHF^{+}11]. dormant [BBM11, WVV^{+}11]. dose [RM14]. Double [SBe10, ASL16, SSW19]. Double-diffusive [SBe10]. down [Meh10, PDER10]. downstream [AP12, DPSW16, LS15, RHMSE15, SSN12, WC17]. drag [RN14]. drainage [MAD^{+}15]. drawdown [MD15]. Dreissena [KKS10]. Dreissenid [KKS10, KKB^{+}18]. drift [BSRP^{+}12, FDP^{+}18]. drinking [DFWPK16]. drive [ASR^{+}17, BNW^{+}14a, BvBB^{+}16, JLR^{+}17, Ker17, MBH^{+}15, SCAB^{+}16, TT12]. driven [Aus19, BH13, BSSR10, CFVU11, DML17, FEW^{+}14, GJR^{+}19, GBB19b, GBR14, HCD19, HSR15, HCC^{+}13, HMFF10, KRR16, LS15, MAF19, MXWC11, PS13, PDER10, RBCS16, RRCH^{+}19, SMLC^{+}18, SCR^{+}12, SMA13, TBAS14, VSD10, WKSR13, WCJ^{+}15, WZTK15, WMI^{+}17]. driver [KBVW12]. Drivers [CÁSO^{+}16, SBR^{+}13, SSJR^{+}10, AMNU16, BWS^{+}14, CLJ^{+}19, DPM18, FZL^{+}14, FSCB11, GAK^{+}19, HGdG^{+}19, HT17b, HATF17, KH16, LBC^{+}18, LHSBP18, LSHK11, LCBC16, PGP^{+}14, PHCD14, RPB17, SBM16, VP15a, ZNVF16, vBBM^{+}19]. drives [APF^{+}18, BSY^{+}16, KHTO13, KVA18, LH19, RDP^{+}17, SWE^{+}18, VAH11]. Driving [SCQ^{+}17, ADS^{+}17, SGG^{+}11]. drought [BGM^{+}13, BHM^{+}17, HEB^{+}19, HAA^{+}19, PMP^{+}17, SCAB^{+}16, YH17]. droughts [WHL^{+}11]. Dry [DTKMK15, DKK^{+}14, dCGS19]. Dryas [Ano17l, ZXZ17b]. drying [RKWH18]. dual [GRE^{+}16]. due [AFG^{+}16, HBB^{+}11, NI10, OIS10, PHJ12, RKWH18, SKV^{+}19, SNM^{+}15, WLV18]. Dunaliella [GBL13, YLJ11]. Durability [TCFP19]. during [Ano10, Ano17l, Ano21c, BBLN11, BPRG^{+}18, BMD17, BMDC10, BBCM^{+}13, BHM^{+}17, BC10, CMB10, CFD^{+}11, CR16, CCC10, CWHP14, DT16, DVC^{+}17, DKG15, FDB^{+}15, GMBL16, GBD^{+}10, HNSM12, HBD^{+}11, HAA^{+}19, HCC^{+}13, HZC^{+}13, IHSS^{+}19, JZZY18, JWS15, JSK^{+}15, KYRMD18, KG18, KHPIP^{+}14, LWWE^{+}18, MC16, MFM^{+}12, NXL^{+}18, NTM^{+}10, OCLW11, OBT^{+}11, OBI12, PMP^{+}17, PCM^{+}16, PCY^{+}10, RMNZ12, RHMSE15, SSFF12, SLG^{+}14, SLPM15, SFLQ^{+}19, VMCM^{+}17, WRWPG19, XDC^{+}19, ZCY^{+}15, ZXZ17b]. dust [BAA^{+}13, BBK^{+}15, CCV^{+}18, MFMC^{+}10, vdJFS^{+}18]. Dutch [LRM17]. dwelling [RPB17, SHD^{+}11]. Dynamic [LG16, MBH^{+}15, BRR^{+}13, BLM^{+}10, HRMD19, MCH12, SWZ^{+}15, SBF18]. dynamically [RMDK10, RDZ^{+}13]. Dynamics [BDP^{+}19, HBD^{+}16, MMGO^{+}17a, RGG^{+}10, TGC^{+}10, WGC^{+}13, WDL^{+}17, ALdML^{+}14, APP12, ABS^{+}19, BH13, BBLN11, BRNS18, BFD^{+}11, BCRC16, BHV^{+}17, BvBB^{+}16, BBB^{+}17, BCRW15, BHM^{+}17, CHHT18, CRB^{+}17, CRCGG^{+}17, CFVU11, CHL10, CMK^{+}10, CdC^{+}11, CAS^{+}17, CWRX19, CSC^{+}11, DHH15, DPSW16, DMMV15, DBRB^{+}15, DBH^{+}16, DVSV13, FHS10, GFT^{+}14, GLI^{+}15, GRT^{+}14, GBT^{+}17, GBB19b, GCH^{+}12, HDK^{+}12, HPCD13, HAC^{+}11, HC10, HPM^{+}10, HHS^{+}18, HNL^{+}13, HSBA10, IH18, JAD^{+}13, JHD^{+}11, JLR^{+}17, KHPIP^{+}14, KSWFG13, KNA^{+}14, KCB^{+}17, LVDM19, LCM^{+}12, LBB18, LBR^{+}12, MCH12, Man10, MGHS18, MMXC15, MWR17, MMWR17, MMD15, MBO^{+}16, MGT15, OBL^{+}19, OBI12, PCF14, PSZ^{+}13, Piw19, PCY^{+}10, PMRRA19, QS19, QWRJ10, RKBA14, RPI^{+}12, RQC^{+}15, RSG11, RRGGCA19, RPG13, RLL^{+}10, RKMN^{+}13, SFS^{+}16, SCAB^{+}16, SNvD^{+}10, SFMF15, SSGB^{+}17, SKJD^{+}14, Sha10, SDS^{+}16]. dynamics [TCG^{+}17, THH^{+}13, TBAS14, TMO^{+}18, TZD^{+}15, VSdG17, VvO11, WDCH18, WRO^{+}11, WFK^{+}16, WFR10, WLHW13, ZWA^{+}14, dCGS19, vdJFS^{+}18]. Dysida [SWM^{+}10]. dystrophic [WMC^{+}15].

Early [JMNG^{+}13, JMN15, MLL^{+}14, MMD15, AACS11, BJ15, BMDC10, HZC^{+}13, LAM12, MTU18, PCJK13, SLA^{+}18, WXF^{+}15]. Earthquake [KJKS18]. East [CMK^{+}10, GNHGM13, MRSE14, MRC^{+}16, NTK^{+}18, RDB^{+}16, CFD15, GLI^{+}15, JZZY18, JCF^{+}10, KK13, MKG^{+}15, MS13, PHB^{+}10, RDB^{+}18, SW14, ZYZ19]. eastern [BSCG17, BPA12, BPW^{+}19, CRJ^{+}14, CJW^{+}19, DTFR12, DvOR^{+}16, HOD^{+}17, HSP^{+}16, JKKM13, KBL^{+}10, Man10, MPM^{+}15, RBG^{+}10, SSS^{+}19, Tho19, VGM14, WBG^{+}16, WHAM15, ZHN^{+}10, DSLLL19, JWGH19, SPB^{+}14, WMM18]. eating [KLM^{+}17, MWSB18]. Ebullition [CHW14, DBSP^{+}16, SOM17]. Ebullition-enhanced [CHW14]. echo [RK13]. Eco [SYdTP^{+}11]. Eco-evolutionary [SYdTP^{+}11]. ecogeochemistry [MHT13]. Ecological [BBK^{+}15, BVP^{+}15, ELJ^{+}16, HT17b, HESU13, MAB^{+}17, MKBSK19, RG19, XDK^{+}17, ALdML^{+}14, APS^{+}19, ABD^{+}17, CJC^{+}12, DSLLL19, HMO^{+}18, KPP^{+}18, LGR^{+}12, PHCD14, PJUR15, PJFJ^{+}15, SMA13]. ecologically [PSD^{+}17]. Ecology [SLBH^{+}19, BRS18, KWF^{+}17, KTL17, MH16, RRCH^{+}19, RGM^{+}11, WLO^{+}19, ZTW^{+}11]. Ecophysiological [CG17, CPOMA15]. Ecophysiology [PGRR^{+}19, PBA^{+}15, PSD^{+}17, PWF16]. ecoregion [RQC^{+}15]. ecosphere [WGDAA19]. Ecosystem [AP12, CJS^{+}17, DLP13, DTM18, RBM14, SBR^{+}13, SCBR12, SLBNG11, SSGM18, SGRB10, ARML10, AMQ^{+}11, BRR^{+}13, BAY^{+}14, BGM^{+}13, BLS^{+}16, CHL^{+}17, CPHD15, CFVU11, FDH^{+}14, GFT^{+}14, GLMG15, GNHGM13, HEB^{+}19, HBR13, HSBA10, HH14, HBM11, KGL^{+}16, KRB^{+}18, LGV13, LALGM18, MH16, MMB17, OHKC^{+}12, OPA^{+}14, PHDH14, QWRJ10, RHDTS^{+}11, SMM11, SHSK14, SFS^{+}16, SWCL12, SLC18, SSP^{+}18, SNG^{+}14, VBC^{+}12, VCPC^{+}16, VMCM^{+}17, VZJ^{+}17, WRB^{+}19, WGRS^{+}17, WTN^{+}15]. ecosystem-level [VCPC^{+}16]. ecosystems [BBT^{+}10, BDU^{+}19, CJWS15, FHS10, GMGM^{+}13, GdG11, KCL^{+}14, LRM^{+}19, MGGS18, MJJMM17, MBBW11, NNE12, PGP^{+}14, SMF10, SGS18, TCFP19, TBF^{+}13, WYW^{+}10, SM11b]. ecotype [SSG^{+}17]. ecotypes [CLFW17]. eddies [BSB^{+}10, CHS^{+}18, KZR^{+}19, Lee18, TNMV^{+}10, WRB^{+}19]. Eddy [BLH^{+}13, AGMR14, BLLB12, HBM11, KB15, KNL10, LBS17, MCGF^{+}11, MS13, PSB^{+}16, RBM14, WBG^{+}16, XDC^{+}19, YKT^{+}15]. edges [AFG^{+}16]. edifice [CSC^{+}11]. Editorial [How15a, How15b, How19, Xen19]. eDNA [RASV^{+}17]. edulis [Les16]. Edward [VML^{+}19]. eel [RLPL14, RPL16]. Eelgrass [MZH15, DDF^{+}10, HHHT19, HBM11, PHLSSS19, RBM14, ZHG15, vdHHC^{+}19]. Effect [BSB^{+}10, CHL10, HSC^{+}11, KSWFG13, LCH^{+}14, MTSG18, MMGO^{+}17b, MBE^{+}13, MHPW18, RQC^{+}15, RAB^{+}17, WP14, WVGB10, BC19, BSSR10, BOT^{+}15, BMB^{+}18, CKP^{+}15, CRJ^{+}14, CLHL12, CBP10, CMG^{+}15, CJ17, DKK^{+}14, DvOR^{+}16, GPA^{+}14, HA16, HXS^{+}10, KLEH16, KMP^{+}11, KJG10, KSY11, MHRH11, MT11, MD15, MMJ^{+}12, RLB^{+}10, TIF^{+}15, Tho19, VFME18, XSAM12, XLS^{+}19, ZHN^{+}10, ZBSR15, vdJFS^{+}18]. effective [FLM^{+}19, SHM^{+}19, WGH^{+}10]. Effects [BMBI12, BMPF19, BSBK13, BSSW11, CZB^{+}18, CSD10, DMN15, DJD^{+}14, FBV11, FDS^{+}14, GC16, GSBR11, HSLH^{+}14, HKP^{+}16, HCC^{+}13, JBB^{+}16, JCS^{+}18, KT13, KGT12, KFJ13, KHG^{+}13, KKP^{+}19, KNA^{+}14, LBB18, LSD18, MGL^{+}16, MKK15, MEM^{+}17, MW15, NBSMN19, PSG^{+}16, PWF16, QWRJ10, RR12, RN14, REDW10, SLE10, SJB^{+}19, SD10, SHF^{+}11, THFG16, VMF^{+}11, VB17, YMB^{+}18, ZTS13, ZMS^{+}18, vEG10, AJ15, BRS11, BHW^{+}12, BH16, BRT^{+}10, BWD^{+}11, BWD^{+}12, BBQ^{+}10, BCM^{+}17, CFAE^{+}15, CL10, CCK^{+}12, CHPH13, CRS^{+}17, DLP13, DHH15, DBFL11, Edm11, FVSL19, FBL15, GRGL^{+}13, GK10, HMV^{+}18, HST^{+}14, HLSW^{+}15, HCS11, JHLK^{+}19, JLRK12, KCM^{+}10, KTK^{+}13, KBL^{+}10, LAM12, LJ18, MVL^{+}10, MLGZ16, MDS^{+}10, MGSM10, OR16, PvEF12, PE17, PSNE15, PSD^{+}17, PWF18, RKWH18, RSTS^{+}18, SWCL12, SNK12, SLH^{+}15, SMR^{+}17]. effects [SLG^{+}14, SBA^{+}11, SSM^{+}19, SH11, Tad10, TJJ^{+}15, TG17, TRA19, VABMS^{+}12, WCS^{+}18, WGH^{+}16, WVV^{+}11, WB19, WGM16, WHL^{+}11, WRH^{+}17, YWY^{+}15, ZEXH15, ZHG15]. Efficiency [ACC^{+}17, CGP^{+}19, HNZ^{+}16, KBVW12, LWWC^{+}16, MA18, MJH^{+}16, MGS12, RRAS17, RM14, TW10b, dGCB^{+}11]. Efficient [JYS18]. efficiently [LTPK^{+}18]. effluent [KCB^{+}17]. efflux [HNHS^{+}15, HEH^{+}17, OLC18]. effluxes [KHTO13]. egg [DHK11, FGMN17, SM11a, ZTS13]. eggs [All10b, BSBK13, BBM11, JTH^{+}11, KKHP14, KMH^{+}17, RCV^{+}14, SVS^{+}19]. eicosapentaenoic [BB10, SW11]. EIFEX [CFD^{+}11]. Eiffel [CSC^{+}11]. eight [CESC13]. Elat [WSB^{+}13]. electron [BMB^{+}18, HVD^{+}18, RKTLM18, SHT^{+}17]. element [CJ17, SH10b]. Elemental [WM12, FWWF18, HBBM19, Kus14, LLB17, LF19, MTM^{+}16, MEM^{+}17, SD10, SYW18, WJHS18]. elementary [HESU13]. elements [GMMV19, MMH^{+}18, SH10b, TNK^{+}14]. Elevated [DM17, HCL^{+}18, ORC^{+}17, BHW^{+}12, BPL^{+}19b, HLSW^{+}15, HBB^{+}11, HRPW15, QFH18]. Elevating [CMG^{+}15]. Elevation [CEES14, CJHR19, LZR^{+}17, SMM11, SNM11]. Elevation-dependent [CEES14]. Elodea [ZLLM10]. elucidated [WGCC14]. elucidating [BSCC15]. Elwha [FDB^{+}15]. embayed [GWN^{+}12]. embayment [CKB^{+}16]. embayments [WMI^{+}17]. embryos [ZS18]. Emerald [SMM11]. emerged [SYW18]. emergence [MBBW11]. emersion [BMD17]. Emiliania [ARW^{+}10, BRS11, BSCC15, FRA^{+}17, Fie13, FCC11, FAF^{+}12, KS13, LCCF10, MMWR17, MLGZ16, RR12, SES18, SBFC18, WA14, WRH^{+}18, ZKL^{+}14]. emission [CCW^{+}19, NSG^{+}16, SPPS10, TMH^{+}18]. emissions [BMN16, CWHP14, HW16, JBB^{+}16, JMJ^{+}19, KBJ^{+}18, LVM^{+}10, LDL^{+}19, MLD^{+}16, OMB^{+}16, TSDF^{+}16, VSdG17, XXZ^{+}19, vBBM^{+}19]. Emphasis [CGT16, GWD^{+}16]. empirical [Meh10, SBT^{+}19, SL10a, VTH^{+}18]. enclosed [GEC^{+}17]. enclosure [CCK^{+}12]. Endemicity [WOC^{+}18]. endobionts [NCT^{+}14]. endogenous [HTL^{+}18]. endosymbionts [TIN^{+}14]. endosymbiotic [Les19, SHKU11]. Enduring [MBHG11]. energetic [BAB^{+}16, RNG^{+}13]. energies [WLO^{+}19]. energy [Aus13, BSY^{+}16, CT18a, CG17, Kir13, LV16, PHDH14, SLH^{+}15, WC17, WWS11, YWL^{+}17]. engineering [TT12]. engineers [SSP^{+}18]. England [BGR14, TWP13]. enhance [CLHL12, RWB^{+}19, SGH12]. Enhanced [Sch19, CHW14, GTPB^{+}11, MBHG11, MBH^{+}15, SEYJ11, AdGAD14]. Enhancement [BAA^{+}13, HAA^{+}19, GWB^{+}14]. enhances [DIC^{+}18, HCL^{+}18, MJH^{+}16, MM11, NNE12, OPA^{+}14, PHLSSS19, WCI^{+}14]. enriched [GWD^{+}16, UCOG16]. Enriching [GMMV19]. enrichment [ATP^{+}15, BBT^{+}10, BHD^{+}17, CF13b, DRE^{+}10, KWRS13, KBJ^{+}18, NB17, OWS^{+}17, OCR10, PHG13, SGRB10, SSGL19, VABMS^{+}12, WGM16, ZCL^{+}19]. enrichments [LBR^{+}12]. ENSO [XDC^{+}19]. entered [FCRW^{+}16]. enters [MPK^{+}13]. Entrainment [MS13, KFJ13, MFL11]. entrance [MMC^{+}10]. entry [RBG^{+}10]. environment [AMMH^{+}13, BLG^{+}15, CBP12, DM17, DBMP^{+}11, DMB^{+}12, EO13, JD16, LFGK10, LYL^{+}17, NBSMN19, RDC^{+}19, TCFP19, TDS^{+}10]. Environmental [BISZ17, BSFH10, BGR14, BCM^{+}17, CLWD13, DLP13, DDH^{+}19, FRA^{+}17, HGdG^{+}19, HJT^{+}13a, LCBC16, TSSH19, All10b, BL13, BMC^{+}16, BHB^{+}12, DJD^{+}14, ETKL15, GM12, GRSD^{+}14, HS10, JZZY18, KIH^{+}15, KFP^{+}18, LJL^{+}18, MMGO^{+}17b, MZH15, MMBP18, PSS^{+}14, PSNE15, RBI^{+}10, Sch19, SBM^{+}15, TNI19, TGC^{+}10, WJHS18, WCV^{+}12, ZTW^{+}11, HJT^{+}13b]. environments [CMMKH12, GPCJ16, KYG^{+}12, KLM^{+}17, MCC^{+}10, MCT^{+}14, NCT^{+}15, PST^{+}13, SPS19, SNG^{+}14]. Enzyme [TG17, FCD12]. EPA [Bre10]. ephemeral [DBRB^{+}15]. ephippia [SBvH^{+}15]. epibenthic [CG17]. epibiont [FA10]. epilimnetic [SNO^{+}16]. epilimnion [SZH^{+}10, vEG10]. epilithic [MU17]. epipelagic [CÁSO^{+}16, CPHD15]. Episodes [CF10]. Episodic [JABZ19, OFGF12]. epizootics [FSBT16]. Eppley [KTL17]. EPS [TMK^{+}13]. EPS-degrading [TMK^{+}13]. Equal [IWF19]. equally [GPCJ16]. Equatorial [Ano17l, NG13, ZXZ17b, CFAE^{+}15, KBL^{+}10, MVL^{+}10, RS16, RZW11, SSG^{+}17, SL10b]. equilibrium [HBR^{+}14]. Erie [JABZ19, JHLK^{+}19, LEK^{+}18, MWBM19, NXL^{+}18, PE13, PFH^{+}17, TSDF^{+}16, VBBR15, WSTD10]. erodibility [JPH^{+}18]. erosion [BBR^{+}14, KTH^{+}19, dBWL^{+}13]. Erratum [AAIA14b, Ano10, Ano15, Ano18a, Ano18b, Ano19b, Ano19c, BWD^{+}12, CR10, CK13, DdG10, GJWS16, HJT^{+}13b, LGC13b, RCSÁS^{+}10, SS12b, ZXN^{+}12]. eruption [MBE^{+}13]. erythraeum [BRS^{+}13]. Erythrobacter [FYT^{+}12]. escape [PST^{+}13]. Escherichia [GWN^{+}12]. Esox [MF19]. Essential [VdSLC^{+}16]. established [OLC18]. establishment [CZB^{+}18, VP15b, ZEXH15]. ester [OALD10]. esters [CBP12, PT11]. estimate [ACD10, HGvB^{+}13, RPMK17, SSB^{+}18, SH10a, SW11]. estimated [HE10, INF12, MQJG13, YKT^{+}15]. Estimates [BMN16, ADCH18, BGB^{+}14, DVC^{+}17, KTH^{+}19, MRSS12, NSG^{+}16, SMM11, VBC^{+}12]. Estimating [CLLH14, CJHR19, HGD14, MHT13, SBT^{+}19, SL10a]. estimation [SC10]. estuaries [CF14, DSS^{+}11, EMS16, HHHT19, HLH13, LDY^{+}16, LLH^{+}15, LS14, MHPW18, PMY^{+}19b, PMY19a, RMH^{+}17, SL10a, WLG^{+}16, WE19, WJHS18, WTN^{+}15]. Estuarine [BGM^{+}13, Sha10, BLG^{+}15, BGR14, CJS^{+}17, Clo18, ES13, FC11, GPCJ16, HMH^{+}16, JBT11, KGM14, KPP^{+}18, MMXC15, MT11, MD15, RKBA14, SML^{+}19, SSL^{+}12, SPGRP^{+}17, VIS^{+}13, WHL^{+}11, WJHS18, vdHHC^{+}19]. estuarine-scale [KPP^{+}18]. estuary [ADCH18, AC15, BWBB15, BBJ^{+}19, BGM^{+}13, CMW^{+}19, Clo19, CAS^{+}17, CFF^{+}17, EHT10, FPG11, FDL17, FYVU17, FDB^{+}15, GLI^{+}15, GMBL16, GPS15, HPM^{+}10, HMFF10, KT13, LC11, MAB^{+}17, MDE11, NBG17, PCPZ18, PHPH^{+}16, REE^{+}12, RRB^{+}16, RGB^{+}19, RNT^{+}19, RHMSE15, SCR^{+}12, SLK^{+}14, Spi15, TCFP19, VW17, WLS^{+}11, WDCH18, WGC^{+}13, WGCC14, YH17, BPW^{+}19, CWRX19, GOD^{+}18, HT17a, LZC^{+}14, Sha10, UMHH^{+}14, WCJ^{+}17, ZYZ19]. Ethiopian [HMV^{+}18]. ethyl [HKU^{+}10]. Etive [HGvB^{+}13]. Eubosmina [FSST11]. Eucampia [MEM^{+}17]. Eudiaptomus [GPL11, KNA^{+}14]. eukaryotes [DKSA19]. eukaryotic [BMM^{+}13, CFVU11, SKKV11, TFLS14]. euphausiid [DOD10, NSO19]. euphotic [KBL^{+}10, LKT17, MGK15, XLS^{+}19]. Europe [GTPB^{+}11, SvKP^{+}18]. European [FSST11, RLB^{+}10, SJB^{+}19]. Eurytemora [LRY12]. eutrophic [DHW11, DHH15, GMBL16, JP10, MMGP^{+}12, MMN^{+}10, MGL^{+}16, MGW^{+}13, NWT^{+}19, PD11, QWRJ10, SWD11, SSYT14, SSGL19, TGC^{+}10, TCG^{+}17, TTV^{+}13, TST^{+}19, VLDM19, WCM19, XPQ^{+}10, ZPK^{+}12, dKYH^{+}12, dKNL^{+}15]. Eutrophication [LH19, AMQ^{+}11, BHC13, BHC14, BG10a, ES13, JHW^{+}19, KGRV18, MJJMM17, MLS^{+}14, PRS^{+}18, PSH^{+}11, PSD^{+}17, RQC^{+}15, RKG^{+}11, RCIB14, SLE10, SWCL12, SM10, SDH^{+}14, Tad10, YP18]. eutrophied [TBLG14]. evaluate [SSYT14]. evaluated [BTH^{+}16]. Evaluating [BSG14, BSSR10]. Evaluation [DB11, HZC^{+}13, MCH12, SSPK^{+}12, SC10]. evaluations [PE17]. evaporation [BWBB15, BGB^{+}14, VLWV14]. evaporative [SBB^{+}18]. evasion [RHMSE15, SSU^{+}16]. Even [WGM16]. evenness [LTPA17]. event [BAA^{+}13, CHL10, KHPIP^{+}14, MMB17, PMP^{+}17, PST^{+}13]. events [BBLN11, CH11, LWS^{+}17, MFM^{+}12, NXL^{+}18, OBT^{+}11, PRS^{+}18, PKWS19, REE^{+}12, VBBR17]. Everglades [HCK10, SGA^{+}17]. Evidence [BHB^{+}19, BMDC10, DVSV13, FHR^{+}15, FWO^{+}18, GFT^{+}14, GSZL13, HHHT19, KBA^{+}12, MCC^{+}10, MVT^{+}17, NMST18, PK14, SNTK15, Tad10, TGG^{+}11, UFW^{+}18, WSB^{+}13, ZHN^{+}10, Alo17, BWBB15, BD15, FPD^{+}10, GdVT^{+}11, GWB^{+}14, HEB^{+}19, JAZ^{+}10, JTV^{+}16, LL11, LEG^{+}10, Meh10, RRD14, SSS^{+}16, SAP^{+}11, THA17, UCOG16]. evolution [GSBR11, LYL^{+}17, NTA14, NAH^{+}11, PDP^{+}10, RSG11, SI10, SOM^{+}15]. Evolutionary [HST^{+}14, HL13, JLG11, RG19, SYdTP^{+}11, SBDS^{+}15]. Exaiptasia [HRPW15]. Examining [MGL^{+}13, PRS^{+}18, KCB^{+}17]. Examples [CJS^{+}17]. Exceptional [GDCM13, SWD^{+}14]. exceptionally [RLB^{+}10]. Exchange [ALL^{+}10a, AFG^{+}16, Ano10, BC10, CF14, CMB10, FEW^{+}14, GLF18, Kus14, MRSE14, NHP17, RHSD^{+}10, SCR^{+}12, TvBR^{+}19, VPC10, WMI^{+}17]. excited [MA18]. excretion [BWP^{+}10, HLGA17, VFME18]. exhibit [BRM^{+}19, CESC13]. exhibited [WHD10]. exist [CR16]. existence [CLFW17, MCYR17]. exopolymer [AAC^{+}19]. exotic [CBP10, PCPZ18]. exotics [KS16]. experiences [APS^{+}19]. experiencing [PK14]. experiment [BBT^{+}10, CFD^{+}11, CCK^{+}12, DIC^{+}18, EED10, GVS^{+}10, KCB^{+}17, MKBSK19, PGP^{+}14, PBV16, SJB^{+}19, Spi15, SVG^{+}18, ZS18, ZLLM10]. Experimental [Alo17, FPD^{+}10, HEB^{+}19, KRB^{+}18, LL11, LEG^{+}10, THA17, MCH12, MBC^{+}18, MU17, MW15, RKLH11, RRD14, SGJB14, VMF^{+}11, ZCK^{+}16]. experimentally [Kus14]. experiments [CLLH14, GKS12, JWS15, KGC^{+}12, KFJ13, Lan14, Lat14, LFL17, OWS^{+}17, PCW19, SMMF19, SNM^{+}15, ZWA^{+}14]. explain [JPH^{+}18, KRR16, LFH^{+}12, PZHD18, WTN^{+}15, ZSM14]. explained [TW10a]. explains [BPW^{+}19, MCLT15]. exploitation [GMJW13]. Exploring [ES13, GN16, SMMF19]. exponents [MPM^{+}15]. Export [MBAS^{+}17, RGB^{+}19, ARB^{+}19, BHD^{+}17, BDK^{+}17, DNH^{+}18, DBC^{+}13, FEW^{+}14, FUS^{+}16, HV16, KHVS11, KNL10, LPO^{+}11, LdJMS^{+}13, MGSM10, SSFF12, SFB12, SSH^{+}14, SL10b, SLBNG11, SSGM18, WM12, WA14, WCC^{+}17, XZGW17, vdJFS^{+}18]. exports [MSGS^{+}13]. exposed [BLH^{+}13, CvHB^{+}18, MF19, MBBW11, SIW^{+}11]. exposure [BH16, CCV^{+}18, IOB^{+}11, MMGO^{+}17b, SPG^{+}13, VSD10]. expression [CMS^{+}18, HTL^{+}18, HBB^{+}11, KP13, PDP^{+}10, TAE^{+}18]. extant [RPB17]. extend [FFA13]. extended [CGL^{+}16, YH17]. extending [OMSC13]. Extensive [BWBB15, CWHP14, GML^{+}12, MHL^{+}16, PTS12, WGRS^{+}17]. extent [BHC13, BHC14, MRSS12, MLS^{+}14, PMPD13, Tho19]. extents [LKT17]. external [SSYT14]. Extracellular [SCG^{+}19, FNSS15, HBD^{+}16, TMK^{+}13]. extract [KFP^{+}18]. extraction [TSDF^{+}16]. extrapolate [CEPPR14]. Extreme [APB^{+}17, CBK18, GPCJ16, GLF17, JD16, MMB17, OBT^{+}11, PMP^{+}17, PST^{+}13, Sch19, WLL^{+}11]. extremely [Bre14]. extremes [BDU^{+}19]. extrinsic [PGP^{+}14]. exudation [GRRA^{+}17, MSS^{+}18].

Fa [SPB^{+}14]. face [CFD^{+}19, IR16]. facets [GdG11, HT17b]. facilitates [KYR^{+}12, MCYR17]. facilitation [CLN^{+}19]. factor [BPL^{+}19b, ITO^{+}17, SCQ^{+}17]. Factors [BBQ^{+}10, ERA^{+}12, PMA18, SDCF16, TSC^{+}19, YHS^{+}17, All10b, ASH^{+}14, BL13, CBS^{+}17, DJD^{+}14, GWN^{+}12, JZZY18, MMGO^{+}17b, QWRJ10, ŠNZ^{+}14, TGC^{+}10, uGH^{+}11]. falls [LPO^{+}11]. Family [Les19]. Famine [KNA^{+}14]. fan [BJDMH10, RBRH10]. farms [SNG^{+}14]. Fast [CESC14, GHSR^{+}16, LdJMS^{+}13]. Faster [HSB^{+}13, KK11]. Faster-growing [HSB^{+}13]. Fate [EOM16, CFB14, EHT10, MPvBS^{+}18, NTM^{+}10, OEM12, ORGE16, TIF^{+}15, VLDM19, WGCC14]. fates [GMMV19]. Fatty [GLS^{+}13, BISZ17, CPPdAR^{+}13, CWF11, GBB19b, HSTK15, IWF19, JTV^{+}16, KNA^{+}14, MMXC15, MKK15, NBSMN19, TEGL11, dKYH^{+}12]. fauna [CFAE^{+}15, MTU18]. faunal [vOSH12]. faveolata [TEGL11]. favor [LOS12, VTH^{+}18]. Fayetteville [HHM^{+}18]. Fe [PKB^{+}17, RRB^{+}16, TSC^{+}19, WGRS^{+}17]. Fear [BMPF19]. feast [KNA^{+}14]. feature [NSO19, SWD^{+}14]. features [LALM16, dlFN10]. fecal [BIM^{+}16, RK13, SPR^{+}15, WRS13]. fed [AC15, CFAE^{+}15, HC10, HCF^{+}10, HC12, WGCC14]. Feedback [AHH^{+}16, GK10, BKA^{+}14]. feedbacks [HW16]. feeders [MSSH12]. Feeding [GBB^{+}18, PTS^{+}19, SGCI14, SAPI14, VIS^{+}13, WMP^{+}19, CLLH14, ETI^{+}16, GPL11, GK15, HRG^{+}15, KVMA17, KGC^{+}16, KSTA18a, LSK11, MJH^{+}16, PCF14, PVA^{+}19, SBDS^{+}15, TRA19, WD15, XNK18, ZTS13, Ano19c]. feeding-current [GK15, KGC^{+}16, XNK18]. felix [ASR^{+}17]. females [SGCI14]. ferric [XSAHV13]. ferritin [CMS^{+}18]. fertilization [CFD^{+}11, DKG15, DFK^{+}17, MGGS18, MLL^{+}14]. fertilized [JTG^{+}11, OCLW11]. fetch [vEG10]. Fewer [LPLH18]. fiddler [LDCT11]. Field [RRD14, BBT^{+}10, GBK^{+}18, HLSW^{+}15, INF12, JGR^{+}14, KZR^{+}19, MU17, NSO19, PHPH^{+}16, Spi15, VMMS^{+}13, ZS18]. fields [GMD11, VLMTEW11]. Fight [SFWP12]. filament [TIF^{+}15]. filamentous [FLLH18, ŠSP17, VSP^{+}11]. filter [ACC^{+}17, CGB^{+}18, LSK11, MSSH12]. filter-feeding [LSK11]. filtering [LJL^{+}18]. filtration [JYS18]. find [LKK13]. Fine [GRSD^{+}14, WGJ^{+}19, WJHS18]. Fine-scale [WJHS18]. finely [ŠNZ^{+}14]. finite [MPM^{+}15]. finite-time [MPM^{+}15]. finmarchicus [CBP12, HTL^{+}18, JWGH19, JMN15, MMJ^{+}12, PPT12]. fire [DPG^{+}12]. fire-affected [DPG^{+}12]. First [BD15, AMNU16]. Fish [CA08, PJFJ^{+}15, SH10a, FDP^{+}18, FC11, GM12, GEC^{+}17, GMJW13, GBK^{+}18, HCD19, HCS11, IPGP10, JGR^{+}14, KCH^{+}12, KYR^{+}12, KGRV18, LP10, MG14, MTEM15, MWR17, Meh10, MVNG11, MS13, NZH^{+}11, PCF14, PHDH14, RWF^{+}12, SPS19, SSH^{+}16, SBK18, SVS^{+}19, SNG^{+}14, TDF^{+}17, WJHS18, WDH^{+}17, WS13, XZC^{+}16, ZPK^{+}12]. fishes [CFRL10, CPHD15, FCD12, TWP13]. fishing [SPP^{+}16]. fishponds [ŠGN^{+}19]. Fitness [HL13, HP19, IWF19, PvEF12]. fixation [AFSM17, ASH^{+}14, BAA^{+}13, BDK^{+}17, CvHB^{+}18, CJW^{+}19, FWWF18, GWB^{+}14, GBD^{+}10, Ho13, HVD^{+}18, JSH12, LWE^{+}11, LWrDM^{+}12, LWWE^{+}18, MFK^{+}13, MGL^{+}16, MCGF^{+}11, MBBG^{+}12, RWC16, RKTLM18, SHT^{+}17, SM10, SM11b, SKK^{+}13, SFI^{+}18, TMH^{+}10, WDMF13, WCC^{+}17, ZCY^{+}15]. fixed [CRJ^{+}14, MPvBS^{+}18]. fixing [GBC^{+}17, SK19, YLH^{+}16]. fjord [AGMR14, CHHT18, DHG^{+}17, DJS18, GLKK10, JP10, MGS12, PML^{+}19, RHSD^{+}10, SKJD^{+}14, WCB^{+}10]. fjords [HDDH^{+}17, MWS10, MSK^{+}17, MMD18]. flagellates [ŠGH^{+}18, SBFB17]. flat [PCD^{+}19]. flatfish [BCDR^{+}19]. flats [LSD18]. flea [BBS12, HMV^{+}18]. flexibility [AGML18, DLP13, HSB^{+}13]. flexible [DBMP^{+}11, LN11]. flexural [RN14]. flight [SFWP12]. floating [HZC^{+}13, RCJ15, VPC10, uGH^{+}11]. Flocs [ZSZ12]. flood [BGM^{+}13, CHL10]. flooded [MF19]. flooding [RKWH18, YH17]. Floodplain [SFB12, AHD^{+}18, FBFR13, RHMSE15, SDS^{+}16]. floods [APB^{+}17, RHMSE15]. Florida [AC17, BR17, HTLM18, HCS11, MBLP11, SGA^{+}17, SS12b, SS12c, SL10a]. Flow [LN11, MDS^{+}10, AGML18, AJC15, BLH^{+}13, CT18a, CFAE^{+}15, CFD^{+}19, DIC^{+}18, FPP^{+}19, HCD19, KHH19, KOFN11, MBHG11, MZH15, MMPSB14, MAD^{+}15, OHKC^{+}12, PMY^{+}19b, SWM^{+}10, SGA^{+}17, SCP^{+}16, TBAS14, TRA19, WZTK15, WWS11]. flow-dominated [PMY^{+}19b]. Flow-induced [LN11]. flow-through [MAD^{+}15]. flowing [BHG^{+}18, SAH^{+}19]. Flowpath [CMS17]. flows [BSY^{+}16, DRP^{+}17, NMST18, ZSZ12, vOSH12]. fluctuating [MSS^{+}18]. fluctuation [LSH^{+}17, WXMS10]. fluctuations [FDH^{+}14, LBB18, SGH12, SGG^{+}11, SNK12, SWD11, SMA15, WCS^{+}18]. fluid [SLK^{+}14, SC10, WKSR13]. flume [AFG^{+}16, DIC^{+}18, DBA16, KGC^{+}12]. Fluorescence [FHS10, LHSG15, AC15, FPG11, HSLH^{+}14, HGD14, WMC^{+}15]. Fluorescence-based [LHSG15]. fluorescent [CÁSO^{+}16, CK12, CK13, MCC^{+}10, PFJ10, THH^{+}13, YHS^{+}17]. fluorometer [PFJ10]. Fluvial [BBLN11]. flux [BR17, BLG^{+}15, CHS^{+}18, DVDB16, FLP^{+}10, HLJ12, JBT11, JTG^{+}11, KB15, Kus14, LGC16, MAC^{+}10, OY10, RMH^{+}17, RK13, SSFF12, SSG^{+}17, SBM16, SPR^{+}15, SL10b, SC10, SSS^{+}19, TIS^{+}13, WRS13, YKT^{+}15, YH17, ZCZ^{+}18]. Fluxes [MdBKL13, AWK^{+}17, AP12, ACD10, ASL16, BMF^{+}16, BHB^{+}19, BLH^{+}13, BMD17, CMW^{+}19, CT18b, CKB^{+}16, DFWPK16, EMO^{+}11, GCSO14, GNHGM13, GSPM13, GJR^{+}19, GN16, HCW^{+}10, HCLS11, HTLM18, HEBS10, HSC^{+}11, HBM11, JMJ^{+}19, KKH11, KBM^{+}14, MCYR17, NSG^{+}16, NHS^{+}12, OBT^{+}11, OFGF12, PSB^{+}16, RKL14, SLK^{+}14, SH10a, SS12b, SS12c, TEZ^{+}18, WLHW13, ZHD^{+}16]. focusing [MMFBB18]. following [AEH19, BDC^{+}14, MGL^{+}13, TMH^{+}18, VLWV14]. Food [KKB^{+}18, KGL^{+}16, MBK^{+}11, VMC^{+}13, Ano19c, BCC^{+}12, BMPF19, CPPdAR^{+}13, CBF11, CPHD15, CS12, CBF10, DFK^{+}17, DML17, DHK11, DvOR^{+}16, DRP^{+}17, FGBS^{+}18, FHR^{+}15, FCRW^{+}16, FPSL18, FWvD^{+}18, GLS^{+}13, GFDC11, GBB^{+}18, GRDPL14, HOD^{+}17, HDDH^{+}17, JTV^{+}16, JC14, KBA^{+}12, KLM^{+}17, KWB^{+}16, LEN^{+}15, LPLH18, LJL^{+}18, LHLT13, LWWC^{+}16, MKB^{+}19, MDB16, MDF^{+}14, MPK^{+}13, MBLD15, MMJ^{+}12, NB17, PvEF12, PCF14, PH13, PDER10, PWF18, PLE^{+}17, RHV^{+}13, SCF^{+}15, ŠNZ^{+}14, ŠGH^{+}18, ŠGN^{+}19, SWP11, SMG12, SBA^{+}11, SL10b, SCP^{+}16, TIF^{+}15, TYX^{+}19, TSK13, VMF^{+}11, VABMS^{+}12, WDJF12, WRO^{+}11, WD15, WMP^{+}19, ZTS13, ZPK^{+}12, vOSH12]. food-web [HOD^{+}17, WDJF12]. food-webs [SCP^{+}16]. foodwebs [GBS17, KS16]. Foraging [XNK18, CMG^{+}15, GMD11, MJH^{+}16, SAS^{+}11]. foraminifer [HONR11]. Foraminifera [NCT^{+}14, NCT^{+}15, RSTP12, SHKU11, UA10]. force [RN14]. forced [SLK^{+}10]. forces [JD16, KCH^{+}12, VP15b]. forcing [AA11, BPPF12, BLM^{+}10, FZL^{+}14, GWN^{+}12, GMGM^{+}13, KHTO13, KWRS13, PMLC^{+}10, RPG13, RGM15, SRAB10, SRA10, WFL^{+}12]. forecasting [LC12]. forest [LWS^{+}17, RMDK10, vEG10]. forested [LBR^{+}12]. forests [OCR10, RDZ^{+}13]. foretaste [RLB^{+}10]. form [SRM^{+}18]. Formation [BBR^{+}14, YAC^{+}19, BNW^{+}14a, CFD^{+}19, HNZ^{+}16, JSK^{+}15, KZR^{+}16, LFH^{+}12, Lee18, LBS17, MDB19, MA18, MAF19, MCC^{+}10, NO17, OCB^{+}18, RLC^{+}11, Sch19, ŠSP17, TT12, WBS^{+}10]. former [MMN^{+}10]. forming [ARB^{+}19, GK10, WBG^{+}16, WdBJF16, WGM16, YKBJL12]. Formosa [IR16]. Formulation [MCH12, Lat14]. fornicata [NBDM16]. fosters [WGDAA19]. Founder [HMV^{+}18]. four [ELJ^{+}16, SMN^{+}15, TW10a]. fraction [ZD18]. fractionated [LYH17, SPGRP^{+}17]. fractionation [Ano10, BC10, BLJ13, CFD15, CMB10, CKCEP10, CCC10, DT16, DLBF17, MC16, SES18, WMBR13, WYL16]. fragilissimus [TJJ^{+}15]. fragmentation [ARML10]. fragments [ZXM^{+}11]. Fram [GRT^{+}14]. framework [FFA13]. France [PDER10]. franciscana [BPL^{+}19a, JW14]. Francisco [CJS^{+}17, GOD^{+}18]. free [MWBM19, MVT^{+}17, NXL^{+}18, SES18, SRCL^{+}13, TCG^{+}17, WCB^{+}10]. free-living [MVT^{+}17, TCG^{+}17]. freeze [HMFB16]. freeze-coring [HMFB16]. frequency [AJC15, BCRW15, GGC^{+}14, IH18, PMLC^{+}10, SLHA19, SDS^{+}11, vH19]. fresh [LHSBP18, OCB^{+}18, SSL^{+}12]. freshwater [ASW^{+}19, BB11, BSM17, CEES14, FHS10, GWSEA10, GMS^{+}18, GdG11, HW16, HCL^{+}18, HKU^{+}10, HMFF12, JMJ^{+}19, KNA^{+}14, KSY11, KOFN11, LZR^{+}17, LJL^{+}18, LFC17, LRG16, MTU18, MPK^{+}13, MXWC11, NTA14, àNTS13, PE16a, PBL^{+}18, PBV16, PSZ^{+}13, RSJ^{+}18, RJFMG17, SYdTP^{+}11, SGA10, ŠNZ^{+}14, ŠGH^{+}18, SWD11, SS19, TW10a, TW11, THH^{+}13, VPC10, VP15a, XXZ^{+}19]. freshwater-marine [BSM17]. freshwater-tidal [HMFF12]. freshwaters [AAC^{+}19, DBFL11, YJO^{+}19]. friends [BBMS17]. frigida [AJ15]. fringing [BMF^{+}16, CSU13, NLHAA^{+}17, WFL^{+}12, WLHW13]. front [CLFW17, NLO^{+}12]. frontal [MS13]. fronts [TB18, WGJ^{+}19, WMT^{+}12]. frozen [BBC^{+}13]. Frustule [DMB^{+}12]. Frustule-related [DMB^{+}12]. frustules [WHH^{+}11]. Fryxell [SSS^{+}16]. Fuchskuhle [CCC10]. fucoid [BMD17]. Fucus [ARB^{+}19, RCJ15]. fuel [OSB^{+}15, TMK^{+}13]. fueled [OCB^{+}18, TLR^{+}13]. fueling [CPPdAR^{+}13]. fuels [GBD^{+}10, XBR^{+}18]. Fukami [NUH^{+}12]. Fukami-ike [NUH^{+}12]. fully [WRH^{+}17]. function [BGM^{+}13, GBR14, LFGK10, MMB17, PDFS14, PDP^{+}10, PJFJ^{+}15, VZJ^{+}17]. Functional [MMWR17, TW10b, BDS11, KSTA18a, RSJ^{+}18, SMMF19, TBLG14, WKAM^{+}19, ZCL^{+}19]. functionally [ASSG12]. functioning [RGB^{+}19, SNG^{+}14]. functions [BSCC15]. fundyense [BVSR^{+}15, HLSW^{+}15, MTH^{+}11]. Fungal [FWvD^{+}18, MM11, MU17]. fungi [KOFN11, MKW^{+}19]. future [BSFH10, CUW11, FCC11, PE16a, PCPZ18, RLB^{+}10].

gain [CEES14, JLG11]. gaining [DBA16]. galeata [SZH^{+}10]. galeata-hyalina [SZH^{+}10]. gammaproteobacterial [OMB^{+}16]. gas [BBJ^{+}19, GKS12, KBJ^{+}18, Kus14, LVDM19, LVM^{+}10, LDL^{+}19, MQJG13, OBT^{+}11, RMH^{+}17, SBM16, SSU^{+}16, SOM17, SSB^{+}16, TBK15, TSDF^{+}16, VPC10, vBBM^{+}19]. gases [BWB^{+}10, WKB^{+}10]. Gasterosteus [KKHP14]. gastropod [HA16, NBDM16]. gastropods [SGG^{+}11]. Gdańsk [PSZ^{+}13]. GDGT [ZKMT^{+}13]. GDGT-based [ZKMT^{+}13]. gelatinous [HRMD19, RWB^{+}19]. gene [CMS^{+}18, DMB^{+}12, HTL^{+}18, HBB^{+}11, RSJ^{+}18, SSS^{+}16, TAE^{+}18]. General [SL10a]. generalist [LGV13, TMK^{+}13]. Generalizations [SdlFdlF^{+}10]. generated [GTPB^{+}11, HD19]. generation [LF19, PPL10]. generations [GNWDL19]. Genetic [All10b, MXWC11, ELJ^{+}16, HMV^{+}18, IBPG17, JB19, LLL10, MNW^{+}19, PSS^{+}14, PMP^{+}12]. Genetically [CR16, GRSD^{+}14]. Geneva [CLB19]. genotype [PMP^{+}12]. genotype-specific [PMP^{+}12]. genus [LDCT11]. Geochemical [LFB^{+}10, SAP^{+}11, YWY^{+}15, DSM^{+}18, GdVT^{+}11, HSP^{+}16]. geochemistry [CF10, DHZ^{+}19, MWC^{+}16, NEH^{+}19]. Geodia [LKF^{+}18]. Geographic [BGP^{+}15, WVL^{+}18]. geographical [YHS^{+}17, YYMN13]. geography [ASW^{+}19]. geologies [RAB^{+}17]. geomorphic [CSU13, GSBR11]. geomorphically [EMS16]. geomorphology [DHZ^{+}19]. Geophysical [MHH^{+}17]. George [SWD^{+}14]. Georges [MBBG^{+}12]. Georgia [JMM14, LHSBP18]. Gephyrocapsa [THFG16, ZBSR15]. Germany [BSSW11, WBS^{+}10]. get [BBMS17]. gets [MDB16]. Getting [LHLT13]. Giant [KZR^{+}16, KZR^{+}19, BRNS18, DPM18, MBLP11, MJH^{+}16, MRB11, PMLC^{+}10, RCH^{+}15, SDS^{+}11, WMP^{+}19]. gigas [BHW^{+}12, BMC^{+}16]. Giling [Ano21a]. Gill [FCD12, TRA19]. glacial [MSAM18, PJUR15, SS12a, VZJ^{+}17]. glacialis [FNSS15, JMN15, PPT12]. glaciated [FBFR13]. glacier [CFAE^{+}15, FHR^{+}15]. glacier-fed [CFAE^{+}15]. Gladioferens [HAL17]. glass [KYC^{+}15]. glauca [VdSLC^{+}16]. Glibert [CJC^{+}12]. gliders [SBM^{+}15]. Global [BM16, MRE18, RBG^{+}10, dGD13, VP15a, CÁSO^{+}16, ESMS13, KKH11, MJJMM17, MRSS12, ML19, SHSK14, SCG^{+}19, WDJF12, WGM16, WVL^{+}18]. globe [SBR^{+}13]. globosa [LG10]. glomalin [AWG^{+}12]. glycolipids [WCV^{+}12]. goby [TB18]. GOCI [QHVM18]. golden [LLW^{+}18]. Goldman [HSB^{+}13]. gondii [SSL^{+}12]. Gonyostomum [LFH^{+}12]. Gorges [RBY^{+}17]. governed [ABS^{+}19, RVvdP^{+}17]. governing [TSC^{+}19]. Gracilaria [GSPM13]. gracilis [GPL11, KNA^{+}14]. gradient [ARML10, BSA^{+}16, CHH^{+}17, CJS^{+}17, DJS18, FOT^{+}15, GEC^{+}17, LZR^{+}17, MvdPK^{+}15, MHPW18, PSS^{+}14, RBM14, RLL^{+}10, SSU^{+}16, SLBH^{+}19, SPGRP^{+}17, SBH^{+}11, WSUC^{+}18, WMP^{+}19, YP18]. gradients [ABD^{+}17, BVC^{+}14, FWO^{+}18, GRSD^{+}14, HS10, JKKM13, LV16, PMY^{+}19b, SHSK14]. grained [CHW14]. Grand [HYK^{+}15, SOO^{+}17]. grass [PCPZ18]. gravel [MAD^{+}15]. gravels [TMH^{+}10]. gravitational [SSGM18]. gray [RWM^{+}19]. Grazer [BTJ^{+}12, LG10, BH13, HMD11, HNZ^{+}16, HCL^{+}18, SFWP12]. Grazer-induced [BTJ^{+}12, LG10, HNZ^{+}16, HCL^{+}18, SFWP12]. grazers [JLC^{+}15, RRD14, SMMF19]. Grazing [LFH^{+}12, MDSG18, BTJ^{+}12, CL10, CLHL12, CLLH14, CSS^{+}16, CPF16, EB12, GLMG15, GNWDL19, KYC^{+}15, KBL^{+}10, Lat14, LFL17, MAV^{+}13, MSAM18, MDS^{+}10, PS17, SRM^{+}18, SNM^{+}15, ŠSP17, WKK^{+}11]. Great [BBM11, FPD^{+}10, RSE^{+}17, RDB^{+}18, SOO^{+}17, ZNVF16, BDB^{+}14, BPL^{+}19a, BWS10, BGW^{+}15, CUW11, DC15, FLP^{+}10, FVSL19, JAS^{+}15, JW14, LÁSDC18, MLCD13, OWFS11, RGG^{+}10, SSH^{+}16, UA10]. greater [HAA^{+}19]. Green [HHM^{+}18, HZC^{+}13, HCL^{+}18, LDT^{+}11, RWM^{+}19, VFS^{+}15, WXF^{+}15, YLH^{+}16, ZXM^{+}11]. Greenhouse [SBM16, BWB^{+}10, BBJ^{+}19, KBJ^{+}18, LVM^{+}10, LDL^{+}19, WKB^{+}10, vBBM^{+}19]. Greenland [ACW^{+}18, AGMR14, FGMN17, HNSM12, MSAM18, MGS12, PML^{+}19, RHV^{+}13, SNO^{+}16, SKJD^{+}14]. gross [BPB^{+}17, DdG10, QS19]. Grosse [CCC10]. grounds [SVS^{+}19]. Groundwater [LDL^{+}19, MSGS^{+}13, WSM^{+}19, DB11, GSZL13, KDGL19, KKH11, KSG^{+}10, LKS^{+}16, LKLH10, LSH^{+}17, LCH^{+}14, LSD18, MGT15, OBL^{+}19, PVLMT^{+}16, RDP^{+}17, RGM15, SS12b, SS12c, VLMTEW11, WGC^{+}13, WGCC14]. groundwater-borne [SS12b, SS12c]. Groundwater-derived [MSGS^{+}13, WGC^{+}13]. groundwater-dominated [KSG^{+}10]. groundwater-fed [WGCC14]. group [BDS11]. groups [ASSG12, BSFH10, KPV^{+}11, LCM^{+}17, MMPSB14, OCLW11, SPP10, SDMK10]. growing [HSB^{+}13, RLSC^{+}13, SNK12]. grown [THFG16]. Growth [CRB^{+}17, LLB17, TBSR13, ADCH18, AA18, BYD19, BPW^{+}19, BBTK^{+}16, BWD^{+}11, BWD^{+}12, BPL^{+}19b, Bre10, BVSR^{+}15, CL10, CL11, CH11, ETKL12, ETKL15, ETKL16, FRA^{+}17, Fie13, FDS^{+}14, FDBW16, GBL13, HST^{+}14, HLG15, HLSW^{+}15, HCK11, IH11, KG18, KMP^{+}11, KWGS18, KWGN^{+}10, KTL17, LLL10, Lat14, LGV13, LBHS13, LFL17, LWWC^{+}16, LCBC16, MCH12, MCWB10, MM11, MWC^{+}16, MGS12, MDE11, NBSMN19, RSTS^{+}18, SLU11, SASB^{+}15, SJM11, SDS^{+}11, ŠGH^{+}18, SW11, SNM^{+}15, SSGL19, THFG16, UA10, WAB^{+}17, WCI^{+}14, XPQ^{+}10, dGCB^{+}11]. Guanabara [CKB^{+}16]. guano [WGRS^{+}17]. Guaymas [LBNT11]. guided [YAC^{+}19]. guild [MAB^{+}17]. Gulf [LGC13b, LBNT11, OrIA10, PHJ12, vdHHC^{+}19, BPA12, BSC^{+}15, BLLB12, CPPdAR^{+}13, DCCB17, FCRW^{+}16, GdVT^{+}11, GDD^{+}16, GCR^{+}10, GNHGM13, GBMG12, HXS^{+}10, HCC^{+}13, KZB^{+}10, KMP^{+}11, LLH^{+}15, Les16, LGC13a, MPM^{+}15, MMC^{+}10, MTH^{+}11, PGB^{+}19, PSZ^{+}13, RG13, SSFF12, STC^{+}11, SFB12, TGGZS^{+}10, TKK^{+}17, VHV10, WWC^{+}13, WSB^{+}13, ZMS^{+}18]. guts [TGG^{+}11]. gyre [DBH^{+}16, FLPL13, FMM^{+}14, KBVW12, CPHD15, DDK10, HPCD13, HDP15, LWB^{+}17]. Gyres [DBV^{+}11].

H [KRR16, BCF^{+}17]. H-Print [BCF^{+}17]. Håkon [FWFB10, LFB^{+}10]. Habitat [BSRP^{+}12, CJWS15, PHCD14, ARB^{+}19, JPH^{+}18, LDCT11, VdSLC^{+}16, WGDAA19, WdBJF16, WGM16, WDH^{+}17, XZC^{+}16]. habitat-forming [WdBJF16, WGM16]. habitats [DRP^{+}17, EMO^{+}11, EMS16, FWFB10, FLM^{+}19, GYP^{+}18, HCD19, HW16, MF19, MHH^{+}17, SAPI14, SSH^{+}16, SPB^{+}14, SGS18]. half [CLLH14]. half-saturation [CLLH14]. halide [FYT^{+}12]. Halimeda [VFS^{+}15]. halocline [FDS^{+}18]. Halomethane [JBPM15]. Halophila [CvHB^{+}18, SLS^{+}11]. Hampshire [SBM16]. haploid [KS13]. Haptophyta [WRH^{+}18]. Harbor [KDGL19]. hard [FLP^{+}10, FVSL19, JAD^{+}13, MKBSK19, MAS^{+}16]. hard-substrate [MKBSK19]. hard-water [FLP^{+}10, FVSL19, JAD^{+}13]. hardwater [SSGL19]. harengus [DDH^{+}19, KKHP14, KMH^{+}17]. Harmful [RGM^{+}11, BVSR^{+}15, HMD11, JLG11, KG18, KSWFG13, RKBA14, RKLH11, SBM^{+}15, SMN^{+}15, SS12b, SS12c, SHF^{+}11, TWWY18]. harpacticoid [KJKS18]. Hastigerinella [HONR11]. hatching [All10b, BSBK13]. Hatteras [MBBG^{+}12]. HAUSGARTEN [MKBSK19]. Hawaii [KDGL19, KSG^{+}10, PCD^{+}19, GCH^{+}18]. Hawaiian [RDP^{+}17]. Hawai‘ [PGRR^{+}19]. HCO [KRR16]. head [CSGW18]. head-capsules [CSGW18]. headland [MFL11]. headwater [DVSV13, HHE^{+}19, HAA^{+}19, JBLJ12, JTH^{+}13, MTT17, PCO^{+}15, REDW10, SBM16, TBSL17]. Headwaters [XDK^{+}17, VZJ^{+}17]. healthier [MDB16]. heat [RKL14, SSW19, SC10, WVL^{+}18]. heatwaves [PSD^{+}17]. heavily [PJFJ^{+}15]. height [SVLS^{+}16]. helgolandicus [MHA^{+}18, MMJ^{+}12]. help [BBMS17]. Hemiscylliidae [WLS^{+}11]. Hemisphere [BPRG^{+}18, FRA^{+}17, LHS19]. Hemoglobin [SZH^{+}10]. Herbicides [NFRU11]. herbivore [PvEF12]. herbivory [GNWDL19, MMD15, MMD18, PJFJ^{+}15]. herdmani [LRY12]. herring [DDH^{+}19, KKHP14, KMH^{+}17, NZH^{+}11]. Heterocapsa [KMF10, MPAS17]. heterocyst [WCV^{+}12]. heterocystous [GWSEA10]. heterogeneities [MG14]. heterogeneity [TLH^{+}11, VBC^{+}12]. Heterogeneous [SWM^{+}10]. Heterosigma [HMD11]. Heterotrophic [PS17, TYX^{+}19, TEGL11, ATP^{+}15, BLMS17, DLBF17, ETI^{+}16, FDS^{+}14, KHG^{+}13, MAFCD^{+}18, MLK11, MW15, MDE11, MSD^{+}14, RGGL^{+}12, RGLM^{+}12, SKJD^{+}14, VF10]. heterotrophs [CL17]. heterotrophy [BS18b, FPPA^{+}11, GBR14, HCK10, HCH^{+}19, JTV^{+}16, SSJR^{+}10]. Hg [Kus14, AHD^{+}18, RQC^{+}15]. hierarchical [CAQS16]. High [AMQ^{+}11, HH14, KJKS18, LDY^{+}16, Lee18, MLD^{+}16, MWS10, NXL^{+}18, OCB^{+}18, PHPH^{+}16, RHMSE15, TCG^{+}17, TDM^{+}13, TAV^{+}10, ÅCA^{+}18, AJC15, Ano19c, ASA^{+}18, ABS^{+}19, BPA12, BCRW15, BCVAn10, CFAE^{+}15, CJHR19, DHG^{+}17, DMSHC16, FGBS^{+}18, GBD^{+}10, GBB^{+}18, HVJ^{+}19, HCK14, IH18, JD16, KCB^{+}17, KvdPVB13, LdlSB^{+}12, LCW17a, LCCF10, MRKR^{+}14, MSK^{+}17, MCYR17, MRH^{+}15, MGJH18, PRS^{+}18, PHG13, RQC^{+}15, RKMN^{+}13, SMM11, SNM11, SSU^{+}16, SCPE15, SDS^{+}11, SIW^{+}11, SPO^{+}18, TRA19, VBBR17, WHD10, WCG^{+}17, vH19, DL11, SJ11]. high-Arctic [ÅCA^{+}18]. high-elevation [SMM11, SNM11]. high-frequency [AJC15, BCRW15, IH18, SDS^{+}11, vH19]. high-irradiance-induced [KvdPVB13]. high-latitude [MGJH18, RKMN^{+}13, WHD10]. high-light [SCPE15]. high-molecular-weight [LCW17a]. high-mountain [DMSHC16]. high-quality [Ano19c, FGBS^{+}18, GBB^{+}18, WCG^{+}17]. High-resolution [TDM^{+}13, ABS^{+}19, HCK14]. High-turbidity [NXL^{+}18, VBBR17]. higher [WHL^{+}11]. highlights [JAD^{+}13]. highly [BAY^{+}14, DBMP^{+}11, EM13, GMBL16, GHS14, NSV^{+}14, RNG^{+}13, SFLB16, SW14, TCFP19, TTV^{+}13, TDF^{+}17]. highly-productive [GHS14]. Hii [MKG^{+}15]. hill [DRP^{+}17]. Historical [BR17, TWP13, RRAS17]. histories [GM12, WJHS18]. history [BH16, BMDC10, LAM12, LJ18, SBDS^{+}15]. Hjort [FDP^{+}18]. HMMV [FWFB10]. HNLC [MRH^{+}15, NO17]. holobiont [DJD^{+}14]. Holothuroidea [SVG^{+}18]. Homeostasis [NCC14, HS18, SZH^{+}10]. homogenizes [ZCL^{+}19]. hopanoids [ZTW^{+}11]. Horizon [FCRW^{+}16]. horizontal [JGR^{+}14, MRSE14, OrIA10, PHJ12, PH15, WMI^{+}17]. horneri [LLW^{+}18]. host [MMWR17, PGRR^{+}19]. hosting [HRPW15, SHKU11]. hot [GGL^{+}15, SFLB16, WMBR13]. hotspots [BVvB^{+}19, Man10, MFL11, TGG^{+}11]. houses [NTI^{+}15]. Hovsgol [KZR^{+}16]. Howe [WHD10]. hsp [TAE^{+}18]. Huaihe [ZZW16]. Huanghe [WLG^{+}16]. Hudson [ACD10, CS12, HMFF10, HMFF12, MGSM10]. human [BBK^{+}15, BDC^{+}14, CHH^{+}17, SDS^{+}16]. human-impacted [SDS^{+}16]. human-mediated [BDC^{+}14]. humans [TWP13]. Humboldt [uGH^{+}11]. humic [HS11, JBLJ12, OCB^{+}18, RJFMG17]. Huron [CSU13, NHP17]. Hurricane [CWHP14]. hurricanes [SLG10]. huxleyi [ARW^{+}10, BRS11, BSCC15, FRA^{+}17, Fie13, FCC11, FAF^{+}12, KS13, LCCF10, MMWR17, MLGZ16, RR12, SES18, SBFC18, WA14, WRH^{+}18, ZKL^{+}14]. hyalina [SZH^{+}10]. hybridization [PPT12]. hybridizing [RKG^{+}11]. hydraulic [DB13, VPWW10]. hydrocarbons [GPS15, ZZW16]. hydrochemistry [MAD^{+}15]. Hydrodynamic [HHA18, AFG^{+}16, JD16, MMFBB18, SNG^{+}14, WP14, WHAM15]. Hydrodynamics [KCL^{+}14, GWN^{+}12, RMK^{+}16, RMLVK12, TDM^{+}13]. Hydrogen [DVSV13, KBA^{+}12, VHV10]. Hydrologic [HCF^{+}10, BDU^{+}19, MBH^{+}15, MAD^{+}15, SBB^{+}18, SRAB10, SRA10]. Hydrological [Dem19, EKS^{+}18, HSP^{+}16, MHPW18]. Hydrology [FUS^{+}16, RAB^{+}17]. hydrolysis [BB11]. hydromedusae [SGCC16]. hydropeaking [HYK^{+}15]. hydrothermal [CSC^{+}11, SPB^{+}14]. hyperboreus [JMNG^{+}13, VGJ17]. hypereutrophic [CSD10, DBFL11, VTH^{+}18]. Hypersaline [GM12, ASL16, DL11, NEH^{+}19]. hypertrophic [ŠGN^{+}19]. Hypolimnetic [MMN^{+}10, CT18a, DHW11, JAD^{+}13, SSB^{+}18, UCOG16]. hypolimnion [BSN^{+}14, NRL15]. hyporheic [FUS^{+}16, LTH^{+}12, SC10]. hypothesis [FDP^{+}18, IH11, Lan14, Lat14, LJL^{+}18, MMFBB18, PWWF18]. Hypoxia [HJB^{+}12, TK12, CG17, CWRX19, HD19, JABZ19, JAD^{+}13, KT13, NPT11, PMPD13, RSG11, Scu16, Sha10, VSD10, WCJ^{+}17, ZSM14]. hypoxia-driven [VSD10]. Hypoxia-induced [TK12]. Hypoxic [REE^{+}12, BSC^{+}15, HT17a, LWS^{+}17, RRB^{+}16, SSGB^{+}17]. hysteresis [CSME13].

Iberia [IR16]. Iberian [VMCM^{+}17, CMM^{+}11, TAV^{+}10]. Ice [MWR17, NHP17, SLA^{+}18, VMAS^{+}16, YAC^{+}19, AJG13, AMNU16, AJ15, BBC^{+}13, BJ15, BAY^{+}14, BCF^{+}17, BCRW15, CDW^{+}16, CMS17, DTKMK15, EM13, FLPL13, GRT^{+}14, GVS^{+}10, HGD14, HKS^{+}15, JSK^{+}15, JLR^{+}17, KIH^{+}15, KFJ13, KGL^{+}16, KZR^{+}16, KZR^{+}19, LKT17, LHS19, MKLKP16, NXL^{+}18, OBI12, PHB^{+}10, RKL14, RVvdP^{+}17, SS16, SPSG14, SSS^{+}16, SKV^{+}19, SAPI14, SPO^{+}18, SMA15, UVGS10, VLWV14, WCB^{+}10]. ice-break [SS16]. ice-covered [CDW^{+}16, CMS17, DTKMK15, FLPL13, HGD14, JLR^{+}17, MKLKP16, RKL14, SSS^{+}16, SPO^{+}18, SMA15, WCB^{+}10]. ice-free [NXL^{+}18, WCB^{+}10]. ice-out [AJG13, BJ15]. ice-walled [SMA15]. Iceland [MAC^{+}10, PCY^{+}10]. Identification [FAF^{+}12, HMFB16, HZC^{+}13, HNL^{+}13]. identified [HML^{+}14]. identifies [SKV^{+}19]. Identifying [BAY^{+}14]. identity [HSTK15]. II [RRB^{+}16]. III [JBT11]. ike [NUH^{+}12]. Illinois [CF10]. Illuminated [SSC^{+}10, MBB^{+}18]. imagery [LAM12, WSTD10]. imaging [AJC15, HSLH^{+}14, JTG^{+}11, PFJ10]. imbalance [GHS14, LTPA17]. Imberger [PHJ12]. immersion [BMD17]. immune [PDP^{+}10]. Impact [BHS^{+}16, HVJ^{+}19, HVD^{+}18, MSAM18, SYW18, YAC^{+}19, AGML18, AA18, BCDR^{+}19, CHH^{+}17, HEB^{+}19, JCS^{+}18, LYH17, PRS^{+}18, PLS^{+}16, RETS16, SBC^{+}17, SBB^{+}18, Tho19, WTC^{+}17, vdHHC^{+}19]. impacted [KGRV18, SDS^{+}16]. Impacts [SPP^{+}16, TNI19, HQB^{+}18, KBHT19, MPSA17, PS17, SFS^{+}16, SSFR19, WLO^{+}19]. impairs [HNZ^{+}16]. Implication [DVC^{+}17]. Implications [AP12, BHW^{+}12, BMM^{+}13, BOT^{+}15, BIS^{+}10, BDC^{+}14, BBS12, BBQ^{+}10, CZB^{+}18, CUW11, CHPH13, CSME13, DBV^{+}11, EM13, GDD^{+}16, KKH11, KTS^{+}14, KPJ12, LCS^{+}19, LPLH18, MTU18, MMB17, MWSB18, NCT^{+}15, PE16a, PHPH^{+}16, RASD10, RHV^{+}13, SMF10, SAS^{+}11, SIW^{+}11, SLBH^{+}19, SS12a, SSB^{+}16, SH11, WRB^{+}19, WMC^{+}15, ZKMT^{+}13, ZTW^{+}11, AMMH^{+}13, BLWV10, CEES14, CA08, ESMS13, FZL^{+}14, GWD^{+}16, HL13, HST^{+}14, HLJ12, Hir12, JGR^{+}14, LK14, MAC^{+}10, MRSS12, MBP^{+}17, MAFCD^{+}18, MSR16, RBD18, SBT^{+}19, SH10a, SCL^{+}19, SSN12, TCFP19, VPC10, WC17]. Importance [EMS16, JC14, MCGF^{+}11, WM12, BBT^{+}10, BSG14, BDS11, GRRA^{+}17, GVS^{+}10, JW14, KGL^{+}16, KBE^{+}17, LBC^{+}18, MDB19, MAS^{+}16, OALD10, SSU^{+}16, SKK^{+}13, SSYT14, VMI13, vOSH12]. Important [CSJ^{+}14, AHH^{+}16, GCH^{+}18, GYP^{+}18, KSFT13, LRS^{+}10, PSD^{+}17, TSB^{+}19, WKG^{+}16, ZXM^{+}11]. imposed [JD16]. imprint [PJUR15]. improper [Lat14]. Improved [LRG16, MD10, ZWA^{+}14]. improves [WS18]. Improving [KTH^{+}19]. IMS101 [BWB^{+}10, WKB^{+}10]. in-lake [SBvH^{+}15]. in-stream [CRCGG^{+}17]. inactivation [HBD^{+}11]. incidence [MWS10]. Inconsequential [RKWH18]. Inconsistency [PWWF18]. Incorporation [LWrDM^{+}12, GBS17]. Increase [HBB^{+}11, APB^{+}17, ETI^{+}16, GMBL16, JTV^{+}16, KRB^{+}18, KRR16, MSS^{+}18, NFRU11, OPA^{+}14, ZCK^{+}16]. Increased [BRS^{+}13, BLS^{+}16, CCV^{+}18, DIC^{+}18, HST^{+}14, HBD^{+}11, KGRV18, KBVW12, WBB^{+}17, KJG10, LRG16, PSG^{+}16, VMF^{+}11, WHH^{+}11]. increases [CF13b, CF14, GBK^{+}18, KSP^{+}12, NWT^{+}19, SMLC^{+}18, SBFB17]. increasing [BR17, CESC13, KK13, MMGO^{+}17a, SKV^{+}19, WCS^{+}18, WE19, WdBJF16, WSUC^{+}18]. incubations [CESC14]. independent [MBC^{+}16]. index [SGG^{+}11]. India [MGW^{+}13]. Indian [MLS^{+}18, RBCS16, WBG^{+}16]. indicate [SVS^{+}19]. indicated [CF10]. indicates [BCF^{+}17, JTG^{+}11, SGG^{+}11, WTC^{+}17]. indications [MdBKL13]. indicators [BISZ17, FC11, VHR^{+}11, WLV18, WDH^{+}17]. indices [IBPG17, LEK^{+}18, Tho19]. Indirect [BH13, HC10, SBA^{+}11]. Individual [BPW^{+}19, BSH16, HMFB16, KTRK11, PE13, WAB^{+}17]. Individual-based [BPW^{+}19]. Indonesia [KCM^{+}10, OBL^{+}19]. induce [PSNE15]. induced [BTJ^{+}12, BBR12, DHZ^{+}19, DHK11, EKS^{+}18, GLS^{+}13, GRT^{+}14, HS11, HBB^{+}11, HFP10, HNZ^{+}16, HCL^{+}18, IGP^{+}12, KWGN^{+}10, KvdPVB13, LN11, LG10, MGS12, MTW12, MSR16, NRL15, RM14, Rie15, RG19, RCV^{+}14, SGH12, SLK^{+}10, SHKU11, SFWP12, SMC^{+}10, SW11, TAE^{+}18, TK12, THA17, VPWW10, XDC^{+}19]. inducers [KS16]. inducible [KS13]. Induction [GBB^{+}19a, KM10, SBDS^{+}15]. inedible [FWvD^{+}18]. inermis [CTA^{+}19, MEM^{+}17]. inertial [Aus13, CTH15, VBBR15]. infaunal [CH11, HHA18, SPP^{+}16]. infection [PS13, ŠSP17, USB^{+}10]. infective [RBRH10]. infer [CJC^{+}12, LGR^{+}12]. inferences [SL10b]. inferred [ALL^{+}10a, BBB^{+}17, CPHD15, FSBT16, FDS^{+}18, LLW^{+}18, VdSLC^{+}16, WLHW13]. Inferring [HCK14, TBSL17]. infiltration [BRF^{+}17]. Inflow [LACI10, BGB^{+}14, SFMF15]. inflows [LDL^{+}19]. Influence [CWF11, CFB14, CSU13, DM17, FDS^{+}18, GCSO14, KWGS18, LG10, MGW^{+}13, RAV^{+}17, RPL16, SBDS^{+}15, VLDM19, VML^{+}19, VBGG^{+}13, All10b, AAC^{+}19, BSRP^{+}12, BGP^{+}15, CF13a, DTPP12, DSM^{+}18, DMB^{+}12, HDK^{+}12, HJMD13, HHS^{+}18, HBBM19, HLH13, KCH^{+}12, KGvdH16, LRG16, MMH^{+}18, MAD^{+}15, MMD15, NSG^{+}16, RPI^{+}12, RDC^{+}19, RMK^{+}16, SLA^{+}15, SFB12, SvKP^{+}18, SMC^{+}10, SS19, WCB^{+}10, WCJ^{+}17, WCC^{+}17, WHD10, WFR10, WDL^{+}17, YJO^{+}19]. influenced [BJF18, FB12, HHHT19, KMF10, MACM11, MKG^{+}15, NLHAA^{+}17, PGRR^{+}19, SVLS^{+}16]. influences [BSM17, BHM^{+}17, DBC^{+}13, HCF^{+}10, HMFF12, LJL^{+}18, LS14, MMHT10, NEH^{+}19, SRM^{+}18, TZD^{+}15]. influencing [BBQ^{+}10, LHS19, SDCF16, TBSL17]. Information [Ano17g, Ano17h, Ano17i, Ano17j, Ano17k, Ano17a, Ano17b, Ano17c, Ano17d, Ano17e, Ano17f, Ano18h, Ano18i, Ano18j, Ano18k, Ano18c, Ano18d, Ano18e, Ano18f, Ano18g, Ano19d, Ano19e, Ano19f, Ano19g, Ano19h, Ano19i, Ano19j, Ano19k, Ano19l, Ano19m, Ano19o, Ano19n, Ano19p, Ano19q, Ano19r, Ano19s, Ano19t, CESC19, GBB^{+}19c, KSTA18b, LF17a, SHT^{+}18, ZXZ17a, KFP^{+}18]. infragravity [MP17]. infrared [KDGL19, LAM12, RDT^{+}14, SW14]. infrastructure [RWM^{+}19]. ingestion [GOD^{+}18]. inherent [LLH^{+}15, SCQ^{+}17]. inhibit [KG18]. inhibited [PKWS19]. inhibition [HS18, LCS^{+}19, MTW12, SNM^{+}15]. initiation [WXF^{+}15]. inland [LFGK10, LRG16, MRH^{+}15, SOH^{+}18, SL10a]. Inlet [CHHT18]. inlets [FEW^{+}14]. innate [PDP^{+}10]. inner [GFDC11, WS13]. Inorganic [CMM^{+}11, MMC^{+}10, TFH17, WDCH18, vdHHC^{+}19, DTPP12, ETI^{+}16, FLP^{+}10, HLG15, HTLM18, LCH^{+}14, LZC^{+}14, MSGS^{+}13, OALD10, RvSM17, SKLG10, SCR^{+}12, STC^{+}11, WWC^{+}13, WLG^{+}16, WKG^{+}16]. inornata [CRB^{+}17]. input [CBF10, DvOR^{+}16, LV16, SCAB^{+}16]. inputs [AC15, BBJ^{+}19, HGvB^{+}13, KWRS13, KSG^{+}10, NLO^{+}12, PSH^{+}11, TWP13, XPQ^{+}10]. inshore [CBS^{+}17, dGCB^{+}11]. inshore-offshore [dGCB^{+}11]. insight [LGW^{+}19, MD10, WS18, WCB^{+}10]. Insights [BRS18, BMD17, CFAE^{+}15, CBS^{+}17, DIC^{+}18, DDF^{+}10, HS18, IBPG17, KWB^{+}16, LLW^{+}18, MGHS18, MH16, MBLD15, OWS^{+}17, PE13, RMDK10, RDZ^{+}13, SKK^{+}15, WWC^{+}13, WSM^{+}19, ÁSNCÁ^{+}13, KPP^{+}18, SLC^{+}16, TCFP19, ZMWM11]. Instabilities [RGM15]. instability [Sch19]. Instantaneous [TT14]. Instr [Ano17g, Ano17h, Ano17i, Ano17j, Ano17k, Ano18h, Ano18i, Ano18j, Ano18k, Ano19j, Ano19k, Ano19l, Ano19m, CESC19, GBB^{+}19c, KSTA18b, LF17a, SHT^{+}18, ZXZ17a]. insufficient [HBD^{+}11]. insularity [LV16]. intact [BHS^{+}16, BHB^{+}12, GKS12]. integrated [SMM11, WXF^{+}15]. Integrating [WFB^{+}11, Ano19c, GBB^{+}18]. Integration [KDGL19, SBM^{+}15]. intense [PKWS19, VLMTEW11, YH17]. intensification [JHLK^{+}19]. Intensity [BSM17, LKLH10, ZBSR15]. Intensive [GML^{+}12]. Inter [CB19, GGPM^{+}10, GMD11, BCM^{+}17, RPL16, Scu16]. Inter-annual [CB19, GGPM^{+}10, GMD11, RPL16, Scu16]. inter-specific [BCM^{+}17]. Interaction [PvEF12, RDZ^{+}13, CHL^{+}17, DBSP^{+}16, ETKL16, RCIB14, TBSR13]. Interactions [HAC^{+}11, MCWB10, BBT^{+}10, BSSR10, BRT^{+}10, CEB^{+}17, CdC^{+}11, FT11, HHW^{+}19, HMD11, HJMD13, LEG^{+}10, LS14, MTM^{+}16, MMWR17, QFH18, RKLH11, RSJ^{+}18, STB^{+}16, SD10, SHF^{+}12, ŠSP17, TDM^{+}13, VCPC^{+}16, WGDAA19, WRO^{+}11, XFH14]. Interactive [CL10, CRS^{+}17, LJ18, LTX^{+}17, PSD^{+}17, SLH^{+}15, SH11, GLMG15, PSNE15]. interactively [MVNG11, WZBW^{+}11]. Interannual [MDB19, RVvdP^{+}17, SPSG14, BAG^{+}17, CB12, DDF^{+}10, KTK^{+}13, PMP^{+}12, PMPD13, RKBA14, XDC^{+}19, ZSM14]. interception [CHS^{+}18]. interface [BKD^{+}16, CDW^{+}16, GLI^{+}15, HT17a, LG16, RWB^{+}19, SLK^{+}14, SWD11, WCG^{+}17]. interflow [HMHI13]. interior [UFW^{+}18]. intermediate [ÁSNCÁ^{+}13]. Intermittent [KHG^{+}13, dCGS19]. Internal [DB13, PPL10, RDC^{+}19, SWL11, SSN12, WKSR13, ABS^{+}19, BK13, CTH15, DHH15, GJR^{+}19, NI10, NRL15, PMRRA19, SI10, SSYT14, VPMrI12, VMMS^{+}13, VMI13, VMCM^{+}17, vH19, LWS^{+}17]. internalization [SMLC^{+}18]. Interplay [uGH^{+}11, LFGK10, OIS10, PHJ12, RPG13]. Interpolating [LC12]. interpret [SHK13]. interpretation [BLG^{+}15, SSC^{+}17]. interreplacement [BMM^{+}13]. Interspecies [TW10a]. Intertidal [BRM^{+}19, CTG15, VPG^{+}19, WKG^{+}16, ALG^{+}13, BHV^{+}17, BGP^{+}15, DPG^{+}12, FEW^{+}14, GML^{+}12, GSPM13, JD16, MBH^{+}15, MMPSB14, MPvBS^{+}18, PPPA14, PLS^{+}16, RWC16, SWE^{+}18, TMK^{+}13]. intra [MGJH18]. intra-annual [MGJH18]. intracellular [BRS^{+}13]. Intraspecific [Hir12, SWP11, WHR18]. intrathermocline [Lee18]. Intrinsic [PGP^{+}14]. introduced [CBP10]. introduction [FSBT16]. intrusion [PVLMT^{+}16]. intrusions [LACI10]. invaded [PCPZ18]. Invasion [SOM^{+}15, BBS12, GGC^{+}14, LFH^{+}12, MGL^{+}13, OBM^{+}11, PWWF18, TB18, TMH^{+}18]. invasions [BBCM^{+}13, DBRB^{+}15]. Invasive [WLV17, BBB^{+}17, HJT^{+}13a, HJT^{+}13b, HSR15, JTH^{+}11, KKB^{+}18, MMB17, PSS^{+}14, RSTS^{+}18, RAV^{+}17, SBA^{+}11, TB18, WLV18]. inventories [LWE^{+}11]. Inventory [KZR^{+}16]. inverse [SL10b]. invertebrate [JC14, KM10, KPP^{+}18, MWS10, VMC^{+}13, WLV17, WGJ^{+}19]. invertebrate-chemoautotroph [MWS10]. Invertebrates [BBM11, BSM17, BSRP^{+}12, HLGA17, MKBSK19, MSK^{+}17, PWWF18, PMA18]. investigate [KDGL19]. investigated [KGM14]. Investigating [DvOR^{+}16, TB18]. investigation [CLB19, FJBP15, JAD^{+}13, SS12b, SS12c]. investments [BAB^{+}16]. invisibility [GRDPL14]. invisible [PFJ10]. ion [FNSS15, MMH^{+}18, SES18]. Ircinia [ASR^{+}17]. Irene [CWHP14]. Iron [CEB^{+}17, CMS^{+}18, LBHS13, MVL^{+}10, OSB^{+}15, RETS16, STB^{+}16, SHF^{+}12, VGM14, WDMF13, ATP^{+}15, BS18a, BTC^{+}19, BIS^{+}10, BRS^{+}13, BG10b, BBB^{+}14, CFD^{+}11, CBF11, CWF11, CFB14, CJ17, DMB^{+}12, EBMR12, FDS^{+}14, FDBW16, IHSS^{+}19, JTG^{+}11, JLR^{+}17, KWM^{+}19, KBHT19, LJ18, MBH^{+}15, MBC^{+}18, MEM^{+}17, MVG^{+}15, NO17, NLO^{+}12, NHS^{+}12, àNTS13, NSV^{+}14, OCLW11, PK14, RNK^{+}16, RLC^{+}11, RLSC^{+}13, RLL^{+}10, RKMN^{+}13, SDSC12, SAS^{+}11, SIW^{+}11, SAP^{+}11, SMH^{+}11, SH11, TSC^{+}19, TNMV^{+}10, WHH^{+}11, WGRS^{+}17, XSAHV13, XFH14, JBT11]. iron-binding [BBB^{+}14]. Iron-dependent [RETS16]. iron-fertilized [JTG^{+}11]. Iron-light [SHF^{+}12]. iron-limited [BRS^{+}13, DMB^{+}12, FDBW16, STB^{+}16]. iron-mediated [SAP^{+}11]. Iron-poor [OSB^{+}15]. iron-rich [àNTS13, RLC^{+}11]. irradiance [ASA^{+}18, BPRG^{+}18, GRGL^{+}13, HSLH^{+}14, KvdPVB13, SLS^{+}11, SSPK^{+}12, THFG16, WHD10, XFH14]. Iseo [VPMrI12, VMI13, HMHI13]. Island [GLMG15, GBT^{+}17, RPMK17, SWD^{+}14, VW17, DCRC16, WHD10]. islandica [MWC^{+}16]. Islands [KH16, KKH11, BAA^{+}13, CHH^{+}17, ELJ^{+}16, UIY^{+}11, VML^{+}19]. Islet [YWY^{+}15]. isolate [HLG15]. isolated [FYT^{+}12, GJR^{+}19]. isolates [CEB^{+}17]. Isolating [WHAM15]. isolation [MXWC11]. Isopora [YLH^{+}16]. isoprene [ESMS13]. isoprenoid [BAY^{+}14]. isoscape [WRB^{+}19]. Isotope [DT16, OCR10, AHD^{+}18, BJDMH10, BSCG17, BGB^{+}14, BTH^{+}16, CS12, CKCEP10, CCC10, CBF10, DTM18, EWB12, EED10, FC11, GLS^{+}13, GMMV19, GCH^{+}18, GRE^{+}16, GVS^{+}10, HPCD13, HHM^{+}18, HOD^{+}17, JSB^{+}14, KGL^{+}16, KWB^{+}16, LRM17, MZH15, MD15, OLC18, SES18, SMG12, TG17, VTH^{+}18, WYL16, WFK^{+}16, WGCC14, ZLLM10]. isotopes [CPPdAR^{+}13, CFD^{+}11, CSGW18, FDS^{+}18, KBA^{+}12, KLM^{+}17, LKLH10, MTEM15, MBLD15, MQJG13, RS16, RHV^{+}13, SSYT14, TMO^{+}18, VHR^{+}11, WLHW13, KBA^{+}14]. Isotopic [CFRL10, GRDPL14, Ano10, Ano21c, BWBB15, BC10, BSMC12, CFD15, CMB10, CPHD15, HSC^{+}11, KFP^{+}18, MC16, MGW^{+}13, NCT^{+}15, RPMK17, RSTP12, SBvH^{+}15, SRAB10, SRA10, TFLS14, WM12, WRWPG19, WSB^{+}13, ZMWM11, WKAM^{+}19]. isotopomer [WFK^{+}16]. Israel [AES11]. Issue [Ano17g, Ano17h, Ano17i, Ano17j, Ano17k, Ano17a, Ano17b, Ano17c, Ano17d, Ano17e, Ano17f, Ano18h, Ano18i, Ano18j, Ano18k, Ano18c, Ano18d, Ano18e, Ano18f, Ano18g, Ano19d, Ano19e, Ano19f, Ano19g, Ano19h, Ano19i, Ano19j, Ano19k, Ano19l, Ano19m, Ano19o, Ano19n, Ano19p, Ano19q, Ano19r, Ano19s, Ano19t, CESC19, GBB^{+}19c, KSTA18b, LF17a, SHT^{+}18, ZXZ17a]. Italy [VPMrI12, VMI13].

J [ACD10, PHJ12]. Janeiro [CKB^{+}16, NEH^{+}19]. January [SWD^{+}14]. Japan [KK13, KJKS18, MTU18, NUH^{+}12, SOM^{+}15, SSYT14, TNI19, THH^{+}13]. Japanese [UIY^{+}11]. Jelly [LdJMS^{+}13, LPO^{+}11]. Jellyfish [SCP^{+}16, DJS18, KTRK11]. Jersey [CMW^{+}19]. Jet [NA17]. joint [DMSHC16]. jointly [LdlSB^{+}12]. July [PCY^{+}10]. Jumping [JJ17]. jumps [DB13]. Junk [MDB16]. juvenile [CH11, HCD19, TIN^{+}14, VdSLC^{+}16].

Kāne’ohe [PCD^{+}19]. Kaiser [Ano21b]. Karenia [HST^{+}14]. Karlsson [ACC^{+}19]. karst [HC12, YMB^{+}18]. Kelp [FGBS^{+}18, uGH^{+}11, BRNS18, CDA16, DPM18, HRN11, LCS^{+}19, LWS^{+}17, MRB11, MP17, PMLC^{+}10, RPMK17, RCH^{+}15, RMDK10, RDZ^{+}13, dBWL^{+}13]. kelp-derived [RPMK17]. Kenya [WKJS^{+}14]. Kenya/Tanzania [WKJS^{+}14]. Key [SSS^{+}19, ARB^{+}19, BPL^{+}19b, ITO^{+}17, KTK^{+}13, LTH^{+}12, PJFJ^{+}15, SCG^{+}19]. Keys [MBLP11]. keystone [PvEF12]. kill [PS13]. Killer [SLG10]. killifish [HHA18]. kilometer [HHS^{+}18]. kindtii [MXWC11, VMC^{+}13]. kinetic [KBVW12, LBR^{+}12]. Kinetics [RSM13, BKD^{+}16, CLLH14, LALGM18, MEM^{+}17, WLR17, SFWP12]. King [SWD^{+}14]. Kinneret [AES11]. Kivu [MRC^{+}16, RDB^{+}18, SBe10, SSW19]. know [BB10]. knows [LVDM19]. Kobbefjord [MGS12]. Kona [KSG^{+}10]. Kongsfjorden [DHG^{+}17, KvdPB18]. Korea [LKLH10]. Krill [KYRMD18, KK11, BPW^{+}19, CTA^{+}19, MPM^{+}15, RNK^{+}16, RK13, RHDTS^{+}11, SAS^{+}11, SAPI14, TT14, TGGZS^{+}10].

L. [MZH15, MMBP18]. lab [GRRA^{+}17]. labeling [DTM18, OEMB10, OEM12, ORGE16, OLC18, ZLLM10, dKYH^{+}12]. labile [DBA16, HT17a, WM12, vOSH12]. lability [CK12, CK13]. laboratory [BC19, CESC14, DBA16, ESMS13, RMDK10, RDZ^{+}13, ZWA^{+}14]. Lack [GWB^{+}14]. lacustrine [DML17, PE17, SCF^{+}15]. Lagoon [TDF^{+}17, CHL10, CSD10, IR16, KKHP14, RGG^{+}10, SWZ^{+}15, SLC18, SGA^{+}17, TvBR^{+}19]. lagoonal [CAS^{+}17]. lagoons [NEH^{+}19, dlFN10]. Lagrangian [CLB19, KCB^{+}17]. lagunensis [KG18]. Lake [AMQ^{+}11, BPL^{+}19a, CR11, CF10, EMH12, EP14, GNHGM13, HHM^{+}18, KHTO13, LVDM19, NXL^{+}18, OrIA10, PHJ12, PE16b, SM11b, SBK18, VHM^{+}10, WSTD10, AJG13, Aus19, BCC^{+}12, BL13, BK13, BPRG^{+}18, BNW^{+}14a, BBR12, BVvB^{+}19, BGB^{+}14, BKA^{+}14, BCRW15, BCVAn10, BSSW11, CDW^{+}16, CT18a, CKP^{+}15, CKD^{+}16, CJHR19, CPOMA15, CGT16, CCC10, CMS17, CJWS15, DKG15, DFK^{+}17, DB13, DHZ^{+}19, EKS^{+}18, EWB12, FBV11, FSCB11, FVSL19, FLPL13, GTPB^{+}11, GDCM13, GPH^{+}13, GKS12, GSBR11, GBP^{+}12, HHE^{+}19, HBR^{+}14, HGD14, HNHS^{+}15, HPS10b, HSR15, HPL11, HKU^{+}10, IH18, JAD^{+}13, JTV^{+}16, JLR^{+}17, KKB^{+}18, KBA^{+}12, KWRS13, KBJ^{+}18, KKP^{+}19, KFP^{+}18, KBM^{+}14, KBE^{+}17, KHP18, LACI10, LRM17, LG16, LH19, LSHK11, LC12, MSSH12, MMGP^{+}12, MHRH11, MGL^{+}16, MRSS12, MAF19, MKK15, MRC^{+}16, MKLKP16, NSG^{+}16, NAH^{+}11, àNTS13]. lake [NWT^{+}19, OBT^{+}11, OY10, OMB^{+}16, OBI12, PCJK13, PCW19, PDER10, PHL^{+}18, PSB^{+}16, PPL10, PMRRA19, PS17, RMF11, RSJ^{+}18, RGO^{+}11, SMM11, SNM11, SFS^{+}16, SPP10, SPFP11, SS16, SPSG14, SJB^{+}19, SFMF15, SBvH^{+}15, SOM17, SSB^{+}18, SK19, SI10, SWL11, SLPM15, SBS^{+}13, SWD11, SMG12, SKKV11, SBB^{+}18, SPO^{+}18, SCBR12, SCL^{+}19, SPG^{+}11, SSYT14, TGC^{+}10, TMF^{+}14, TAE^{+}18, THH^{+}13, TTV^{+}13, TST^{+}19, TMO^{+}18, UCOG16, UIY^{+}11, VLDM19, VPMrI12, VBBR17, VTH^{+}18, VSP^{+}11, VHR^{+}11, VMMS^{+}13, VCPC^{+}16, VMI13, WS18, WKSR13, WLV17, WCM19, WYL16, WBZ^{+}13, WFK^{+}16, WCJ16, WSUC^{+}18, WCCP14, WRH^{+}17, WKJS^{+}14, XXZ^{+}19, ZOB^{+}15, ZPK^{+}12, ZCK^{+}16, dKNL^{+}15, AES11, AMB^{+}11, AA11, Aus13, BVC^{+}14, BNW^{+}14b, BRT^{+}10, BS18b, BLS^{+}16, BSSW11, CCK^{+}12, CLB19, CCC10, CSU13, CMK^{+}10, GTPB^{+}11, HMHI13, JABZ19, JHLK^{+}19, JW14, KCM^{+}10]. Lake [KIH^{+}15, KYG^{+}12, KWB^{+}16, LEK^{+}18, LCM^{+}12, LK14, LZK18, LG16, MQP^{+}16, MRSE14, MWBM19, MMGP^{+}12, MGW^{+}13, MRC^{+}16, NUH^{+}12, NHP17, OWS^{+}17, PRS^{+}18, PSH^{+}11, PE13, PDER10, PHL^{+}18, PFH^{+}17, QHVM18, RDB^{+}18, RMNZ12, RAV^{+}17, RPH^{+}10, SSS^{+}16, SLK^{+}10, SBe10, SSW19, SMA15, SSYT14, TGC^{+}10, TCG^{+}17, THH^{+}13, TA14, TSDF^{+}16, VPMrI12, VBBR15, VLWV14, VAH11, VHrI10, VMI13, WP14, WZG^{+}14, WBS^{+}10, WKJS^{+}14, WMI^{+}17, XXZ^{+}19, XPQ^{+}10, YAC^{+}19, ZWL^{+}14, ZMWM11, dKYH^{+}12]. Lake-size [KHTO13]. lake-water [VHR^{+}11]. lake-watershed [JLR^{+}17]. Lakes [WBZ^{+}14, APS^{+}19, ACD10, ACW^{+}18, Ano21a, AHD^{+}18, AA18, BHC13, BHC14, BHS^{+}16, BMF^{+}16, BHB^{+}19, BPGE13, BBC^{+}13, BJ15, BBK^{+}15, BL18, Bre14, BGB^{+}14, BBQ^{+}10, BLWV10, BSY^{+}16, CA08, CTH15, CJHR19, CKCEP10, CR10, DKG15, DBSP^{+}16, DMMV15, DB11, DTKMK15, DKK^{+}14, DMSHC16, FWS^{+}14, FWO^{+}18, FPD^{+}10, FLP^{+}10, FOT^{+}15, FMP^{+}13, GSG^{+}17, GJWS14, GJWS16, HMO^{+}18, HGdG^{+}19, HS11, HATF17, HNL^{+}13, HSTK15, HFP10, HML^{+}14, JSH12, JLR^{+}17, KHTO13, KBA^{+}14, KKS10, KFJ13, KBT16, KHVS11, KMC^{+}15, KZR^{+}16, KZR^{+}19, LBC^{+}18, LPLH18, LZR^{+}17, LCW^{+}17b, LS15, LHS19, MNW^{+}19, MLD^{+}16, MMN^{+}10, MJJMM17, MRSS12, MLS^{+}14, Meh10, MMH^{+}18, MBE^{+}13, MAD^{+}15, MW15, MMFBB18, MSD^{+}14, MMG16, OSC14, OSB^{+}15, OWFS11, OSHS19, PSH^{+}11, PTS^{+}19, PH15, PHG13, RR13, RKG^{+}11, RKWH18, RLB^{+}10, RHV^{+}13, RAKE05, RKL14]. lakes [SHSK14, SJM11, SNO^{+}16, SBvH^{+}15, SM10, SP11, SLP^{+}14, SLA^{+}15, SBK18, SPG^{+}13, SH10a, SdlFdlF^{+}10, SS12a, SDH^{+}14, SBR^{+}13, SS19, SSJR^{+}10, SRAB10, SRA10, SSM^{+}19, SHL^{+}18, SSGL19, TLG^{+}11, TPM^{+}14, UIY^{+}11, VSdG17, VP15b, VBC^{+}12, WWC^{+}18, WMC^{+}15, WXMS10, WCP^{+}15, WVL^{+}18, ZZY^{+}10, ZHN^{+}10, ZCL^{+}19, ZZW16, ZZAC13, ZHD^{+}16, vEG10, BGW^{+}15, BBM11, DC15, RSE^{+}17, SSH^{+}16, SOO^{+}17, ZNVF16]. laminae [HMFB16]. laminated [TBK15]. Land [DCCB17, BSM17, GTR^{+}13, Ker17, KSG^{+}10, KGvdH16, LLH^{+}15, LMR14, MHRH11, SLE10, TT12, WC17, WYW^{+}10, WCG^{+}17, YJO^{+}19, YWY^{+}15, ZTW^{+}11]. land-based [LMR14]. Land-use [DCCB17, KGvdH16, MHRH11]. Landscape [VZJ^{+}17, BSRP^{+}12, FSCB11, FWO^{+}18, FLM^{+}19, FJBP15, KHTO13, Rie15, VCPC^{+}16, WS18, WBZ^{+}14]. landscape-based [WS18]. landscape-scale [BSRP^{+}12, FJBP15]. landward [KJKS18]. Lanice [BBR^{+}14]. Lanyu [YWY^{+}15]. Large [GMS^{+}18, KPW^{+}11, SSH^{+}14, SBKO18, WC17, WKK^{+}11, WHR18, YYMN13, APB^{+}17, BPPF12, BBR12, BCRW15, BSSW11, CT18a, CTH15, Clo18, CHL10, CKB^{+}16, FWO^{+}18, FLM^{+}19, GMD11, GK14, HDK^{+}12, HC12, HCK14, HSTK15, HCC^{+}13, JAD^{+}13, Ker17, LACI10, LL11, LBS17, MSSH12, MAB^{+}17, MAF19, MWC^{+}16, MSD^{+}14, MRC^{+}16, NAH^{+}11, OY10, PRL18, PPL10, QFH18, RBI^{+}10, SCF^{+}15, SFMF15, SBS^{+}13, SVMT15, SSM^{+}19, TGC^{+}10, TCG^{+}17, TB18, THH^{+}13, TTV^{+}13, VBBR17, VP15b, VAH11, WDL^{+}17, YLJ11, ZXM^{+}11]. Large-scale [YYMN13, BPPF12, MWC^{+}16, PRL18, RBI^{+}10, SCF^{+}15, TB18, VAH11, ZXM^{+}11]. larger [HAA^{+}19, SHKU11, WCI^{+}14]. largest [GTPB^{+}11, GKT^{+}15, TSB^{+}19, WXF^{+}15, XXZ^{+}19]. larva [JTH^{+}11]. larvae [FDP^{+}18, FGBS^{+}18, HCS11, IPGP10, LDCT11, MCT^{+}14, RCV^{+}14, RLPL14, SGA10, SWP11, SPPS10, WRB^{+}19, WGJ^{+}19, WMC^{+}18, WHAM15]. Larval [MCT^{+}14, MFL11, MFM^{+}12, BCDR^{+}19, DDH^{+}19, FRP^{+}14, GBMG12, HNHS^{+}15, HPS^{+}10a, HCS11, KPP^{+}18, LRS^{+}10, MF19, MSM^{+}17, MS13, PRL18, RNG^{+}13, RPL16, SGG^{+}11, WJHS18, ZSZ12, ZPK^{+}12]. laser [PFJ10]. last [JAD^{+}13]. Latasa [Lan14]. late [HVJ^{+}19, KYRMD18, MDB19, PCY^{+}10]. latitude [LCBC16, MRKR^{+}14, MGJH18, RKMN^{+}13, WHD10]. Latitudinal [CNL^{+}15, HP19, HLH13, LV16, MLS^{+}18, MCGF^{+}11, SvKP^{+}18]. lato [RSE^{+}17]. Lau [SPB^{+}14]. Laurentian [BBM11, RSE^{+}17, SOO^{+}17, ZNVF16]. law [MD10]. Lawrence [BPW^{+}19, FLM^{+}19, HT17a, vdHHC^{+}19, GdVT^{+}11, MPM^{+}15]. layer [ANP^{+}14, BNW^{+}14a, BBB^{+}14, CT18a, DHH15, HWZ13, KT13, OIS10, PHJ12, SMLC^{+}18, SNK12, SBNC^{+}19, SWL11]. layered [SBNC^{+}19]. layers [BBMS17, CHPH13, CFW^{+}14, LBS17, SBBNM14, TWWY18]. lead [CSJ^{+}14, MBC^{+}18, SMLC^{+}18]. leads [MHL^{+}16, SHKU11]. Leaf [SCPE15, KLM^{+}17, KOFN11]. leaf-eating [KLM^{+}17]. leakage [BOT^{+}15]. learn [SOM17, Sha10]. lee [MFL11]. legacies [PJUR15]. legacy [DKK^{+}14, FVSL19, MHRH11]. leidyi [CMG^{+}15, HJT^{+}13a, HJT^{+}13b, JTH^{+}11, JCS^{+}18]. Length [PH15, BTJ^{+}12, BPL^{+}19b, Hir12, SSGM18]. Length-scale [PH15]. lens [KZR^{+}19]. lens-like [KZR^{+}19]. lentic [HT17b]. Leptodora [MXWC11, VMC^{+}13]. lessen [SMH^{+}11]. Lessons [PCW19]. leucine [HWZ13, dGCB^{+}11]. Levant [GHSR^{+}16]. level [ETKL16, GTPB^{+}11, GGL^{+}18, LAM12, LH19, LALGM18, LBB18, SJB^{+}19, TG17, VCPC^{+}16, WXMS10, XZC^{+}16, YMB^{+}18]. levels [BHW^{+}12, BB10, CFRL10, CUW11, HATF17, RHV^{+}13, WA14, WHL^{+}11]. Liaohe [LYL^{+}17]. LiDAR [KYG^{+}12]. Life [JD16, BH16, BMDC10, DdD^{+}10, GYP^{+}18, LAM12, LJ18, RR12, TRA19, WRH^{+}18]. life-cycle [GYP^{+}18, RR12, WRH^{+}18]. ligands [BBB^{+}14]. Light [AJ15, BRT^{+}10, BVvB^{+}19, BS18b, ETKL15, KYG^{+}12, LWB^{+}17, PvDM^{+}13, PKB^{+}17, ŚF19, AHH^{+}16, BPB^{+}17, BLH^{+}13, BG10b, CUW11, ETKL16, EM13, EBMR12, GFH13, GSB^{+}17, GBR14, HYK^{+}15, HQB^{+}18, HBZ12, JGR^{+}14, KGRV18, KWGS18, LBC^{+}18, Les19, LAC^{+}19, MSS^{+}18, MZH15, MVNG11, MU17, MBE^{+}13, NWT^{+}19, PE16b, PE17, RBD18, RR12, RNT^{+}19, RDT^{+}14, RAB^{+}17, RGGL^{+}12, SLK^{+}10, SKV^{+}19, SCPE15, SYdTP^{+}11, SNK12, SLH^{+}15, SBF18, SHF^{+}12, SH11, TBHM^{+}13, WPL^{+}14, XSAHV13, XLS^{+}19, ZBSR15, ZD18, vEG10]. light-dependency [KWGS18]. lignin [BBS^{+}18]. like [KMF10, KZR^{+}19, TAV^{+}10, UVGS10]. Limfjorden [JP10]. limit [SLS^{+}11, VP15b, uGH^{+}11]. Limitation [ASA^{+}18, Alo17, BMBI12, BLMS17, BSA^{+}16, CAQS16, CJ17, DHK11, FDS^{+}14, GWSEA10, GBB19b, HST^{+}14, Ho13, JJ17, KTK^{+}13, KBHT19, LGV13, LAC^{+}19, MEM^{+}17, MdBKL13, MVT^{+}17, NCC14, PK14, PWF16, REDW10, SLK^{+}10, SPHVA19, SIW^{+}11, SW11, SBH^{+}11, SMH^{+}11, SH11, TSC^{+}19, THFG16, WHH^{+}11, WZC13]. limitations [LBHS13, VIS^{+}13]. Limited [MFM^{+}12, BRS^{+}13, CBF11, DMB^{+}12, FDBW16, GBL13, KvdPB18, MFK^{+}13, MAB^{+}17, NG13, PSG^{+}16, STB^{+}16, SSPK^{+}12, SVS^{+}19, VTH^{+}18]. Limiting [BB10, LC11, MCWB10]. limits [CG17, CJWS15, SL10a]. Limnol [Ano21b]. limnological [FVSL19, SLHA19, WCM19]. Limnology [Ano21a, SPO^{+}18, Xen19]. Lincoln [RMJ^{+}18]. Line [CHH^{+}17]. lineages [Piw19]. linear [MA18]. lined [FYVU17]. linkage [XZC^{+}16]. linkages [AGCA16, TKB18]. Linked [SSB^{+}16, BWS10, BBJ^{+}19, CBP12, FSBT16, SHSK14, SSFF12]. Linking [LV16, MPM^{+}15, MWC^{+}16, SNM11]. links [BJ15, BKA^{+}14, RG13]. Lions [GBMG12, KZB^{+}10]. Lipid [JWGH19, BAY^{+}14, KGL^{+}16, LGV13, SGVR16, VGJ17]. lipids [BHS^{+}16, BHB^{+}12, ZKMT^{+}13]. lithogenic [DTPP12]. litter [KOFN11, MH16, MM11]. little [BBMS17, MTT17]. Littoral [HFP10, HMFF12, CMK^{+}10, LBB18, WXMS10]. Littoral-zone [HMFF12]. lived [GPA^{+}14, MS13, nVOH12]. Living [HPS10b, NBSMN19, TRA19, MVT^{+}17, TCG^{+}17]. lno.10504 [Ano21a]. load [SL10a]. loaded [NXL^{+}18]. loading [ES13, GWN^{+}12, JSH12, KJG10, KHVS11, LMR14, SK19, ZSM14]. loadings [SSYT14]. loads [BSA^{+}16, CBK18, LdlSB^{+}12, RAV^{+}17, WTN^{+}15]. lobate [CMG^{+}15, JCS^{+}18]. lobster [WRB^{+}19]. Local [HSR15, MCT^{+}14, HLH13, JPH^{+}18, NA17, PBV16, SvKP^{+}18, XZC^{+}16]. Locating [TRD^{+}14]. location [HZC^{+}13]. loch [CBP12, HGvB^{+}13]. Lofoten [WB19]. Lombok [OBL^{+}19]. Lonar [MGW^{+}13]. Long [APS^{+}19, BGW^{+}15, DC15, EP14, HSCM19, KMC^{+}15, KHK^{+}19, MKG^{+}15, MSR16, PCW19, PJUR15, RG13, SK19, SSFR19, VKC18, VvO11, WVV^{+}11, WB19, Xen19, ZHN^{+}10, AAIA14a, AAIA14b, BvBB^{+}16, BMB^{+}18, CJS^{+}17, Clo19, DB13, GPA^{+}14, LSDW18, LC12, MKBSK19, OEMB10, nVOH12, RWM^{+}19, RKWH18, RGO^{+}11, RNT^{+}19, RMNZ12, Sha10, TNI19, TCFP19, WCM19, ZWL^{+}14, GLMG15, GBT^{+}17, MKBSK19, VW17]. long-distance [BMB^{+}18]. long-lived [GPA^{+}14, nVOH12]. long-standing [LSDW18]. long-studied [Clo19]. Long-term [APS^{+}19, DC15, EP14, HSCM19, KMC^{+}15, KHK^{+}19, MKG^{+}15, MSR16, PCW19, PJUR15, RG13, SK19, SSFR19, VKC18, VvO11, WVV^{+}11, WB19, Xen19, ZHN^{+}10, AAIA14a, AAIA14b, BvBB^{+}16, CJS^{+}17, LC12, MKBSK19, OEMB10, RWM^{+}19, RKWH18, RGO^{+}11, RNT^{+}19, RMNZ12, TNI19, TCFP19, WCM19, ZWL^{+}14, MKBSK19]. long-time [Sha10]. longevities [HBCK10]. longicornis [SNTK15]. longimanus [BSBK13, BBB^{+}17, BBS12, WLV17, WLV18]. longispina [FSST11, PMP^{+}12, PTS12]. longitudinal [HCK14, PMP^{+}17, SPFP11]. loop [BRS18, BKA^{+}14, PD11]. López [CL11]. Lophelia [LGC13a, LGC13b, MKB^{+}19]. Lord [WHD10]. losing [DBA16, SC10]. loss [CRJ^{+}14, CBS^{+}17, DIC^{+}18, GFT^{+}14, GML^{+}12, JLG11, KRB^{+}18, KvdPVB13, MHL^{+}16, OEMB10, SBB^{+}18, TTTM^{+}19, TT12, VdRA^{+}19, WC17, WVL^{+}18]. losses [VCPC^{+}16]. lotic [FBFR13]. Louisiana [EBMR12]. Low [CGP^{+}19, CJW^{+}19, HWZ13, KvdPVB13, SLC18, ASSG12, Bre10, CT18a, CTG15, CT18b, CF10, GBD^{+}10, HATF17, LL11, LFC17, MSS^{+}18, MCYR17, MRH^{+}15, OBNP^{+}10, OSB^{+}15, PMLC^{+}10, PKWS19, SSS^{+}19, WA14, WHD10]. low-energy [CT18a]. low-frequency [PMLC^{+}10]. low-molecular-weight [LFC17]. low-oxygen [LL11]. low-phosphorus [OBNP^{+}10]. low-tide [CTG15, CT18b]. lower [GCH^{+}18, GPS15, HSB^{+}13, LCZ^{+}19, WD15, ZKMT^{+}13, ZZW16, OPZ13]. lowland [SGA10, SWP11]. LR [DMS^{+}18]. lucius [MF19]. Lugano [BNW^{+}14b, WZG^{+}14]. Luminescence [TLB^{+}16]. lunar [HLFM^{+}10, OMSC13]. Lyapunov [MPM^{+}15]. lysis [CPF16, EB12, PD11]. lysogenic [PS13]. lytic [PS13].

M [ACD10, OMSC13, vH19]. M. [RF13]. Mackenzie [TLG^{+}11]. macroaggregates [SSL^{+}12]. macroalga [ARB^{+}19, GSPM13, PSD^{+}17, RSTS^{+}18]. Macroalgae [WCS^{+}18, CHPH13, HAC^{+}11, HBM^{+}15, MRB11, SKV^{+}19]. macroalgal [BPV^{+}19, HLH13, LCS^{+}19, TDS^{+}10]. macrobenthic [SBT^{+}19]. Macrocystis [DPM18, MRB11, PMLC^{+}10, RCH^{+}15, RMDK10, RDZ^{+}13]. macrofauna [TLR^{+}13]. macrofaunal [RPB17]. macroinvertebrate [HT17b, KHH19, KGvdH16, WXMS10, WWS11, ZCL^{+}19]. macroinvertebrates [Ano19c, GPCJ16, GBB^{+}18]. macronutrients [MRE18]. Macrophyte [VCPC^{+}16, GAM^{+}19, SLK^{+}10]. macrophyte-derived [GAM^{+}19]. Macrophytes [dKNL^{+}15, BSRP^{+}12, CLN^{+}19, LBB18, VP15b, ZLLM10]. macroscale [SCL^{+}19]. macroscales [WS18]. macrotidal [LDY^{+}16]. macrozooplankton [KKS10]. made [SSB^{+}16, WXF^{+}15]. Madeira [BMF^{+}16]. maenas [GGC^{+}14, MCT^{+}14]. magna [MCWB10, PDP^{+}10, SW11, TYX^{+}19]. magnitude [GTR^{+}13, PKB^{+}17]. Magnitudes [FLP^{+}10]. main [AGCA16, JWS15]. Maine [BJ15, Les16, MTH^{+}11]. Major [MGGS18, ASSG12, CH11, CJ17, DKSA19, GMMV19, HLFM^{+}10, MMPSB14, MMH^{+}18, SAPI14, ŠGN^{+}19, TDF^{+}17, TNK^{+}14, TZD^{+}15, VSdG17, vBBM^{+}19]. major-ion [MMH^{+}18]. make [TDF^{+}17]. Malay [TLB^{+}16]. man [SSB^{+}16]. man-made [SSB^{+}16]. management [BDC^{+}14, SM10, TCFP19, WZR19]. manganese [HS18, JLR^{+}17, MBC^{+}18, MdBKL13, NLO^{+}12, RSM13, TNMV^{+}10]. mangrove [ARML10, CCW^{+}19, KLM^{+}17, MSGS^{+}13, OCR10, OLC18, PCPZ18, RMH^{+}17, SML^{+}19]. mangrove-dominated [ARML10]. mangroves [Alo17]. manipulation [LEN^{+}15, OVRJ13, PCJK13]. maple [MM11]. Marenzelleria [RF13]. margin [BK11, GVS^{+}10, NTK^{+}18, SFLB16]. margins [FB12]. Marguerite [RVvdP^{+}17]. marina [AHJS15, DIC^{+}18, FJBP15, HHHT19, HBM11, MZH15, MHH^{+}17, MMBP18, RBM14, VPWW10]. Marine [CBP10, FSBT16, LCZ^{+}19, AWG^{+}12, AACS11, AMMH^{+}13, BYD19, BTJ^{+}12, BMC^{+}16, BSM17, BWP^{+}10, BSMC12, BMB^{+}18, CBFK19, CSÁS^{+}10, CWF11, CLHL12, CFB14, CL17, CEB^{+}17, CMS^{+}18, CdC^{+}11, DdD^{+}10, DT16, DMN15, DdG10, ETKL15, ESMS13, FTC10, FT11, FYT^{+}12, GMD11, GBL13, GBB^{+}19a, GBMG12, GRDPL14, GRPB^{+}17, HBB^{+}11, HBCK10, JBPM15, KP13, KPW^{+}11, KS13, KPJ12, KWGN^{+}10, LRY12, LF16, LF17b, LBHS13, LK14, LKK13, LALGM18, LRG16, MCLT15, MTM^{+}16, MJJMM17, MB10, MHT13, MTEM15, MCC^{+}10, MMPSB14, MZB^{+}15, MBO^{+}16, NTA14, NLM^{+}12, NRS16, NB17, ORC^{+}17, PBA^{+}15, PLS^{+}16, PSD^{+}17, RASD10, RS19, RSM13, RBI^{+}10, RCSÁS^{+}10, RSTS^{+}18, SLC^{+}16, SGCI14, SASB^{+}15, SPS19, SBDS^{+}15, SEYJ11, SSL^{+}12, SFLB16, SMW^{+}18, SYW18, SHF^{+}11, SH11, TJJ^{+}15, USB^{+}10, VMAS^{+}16, VBGG^{+}13, VF10, VPWW10, WWC^{+}13, WOC^{+}18]. marine [WKAM^{+}19, XBR^{+}18, XSAM12, ZYZ19, vdJFS^{+}18]. marine-derived [LRG16]. marinus [VIS^{+}13]. maritimus [AMMH^{+}13]. markers [WJHS18]. marsh [ALG^{+}13, AC17, BvBB^{+}16, CEES14, CZB^{+}18, CF13b, CF14, FYVU17, KJG10, KOFN11, LHSBP18, PE16a, PCPZ18, SKGT17, SBNC^{+}19, SSP^{+}18, SGS18, SVMT15, TMH^{+}18, WDCH18]. marsh-dominated [WDCH18]. marsh-lined [FYVU17]. marshes [GGL^{+}18, SHM^{+}19, WKG^{+}16]. Mass [MMB17, RDT^{+}14, CFD15, CL11, CR10, EBMR12, Hir12, HLGA17, LRM^{+}19, NLM^{+}12, RBY^{+}17, RSN16, RAKE05, RN14, SSC^{+}10, WGC^{+}13]. Mass-specific [RDT^{+}14, EBMR12, NLM^{+}12]. Massachusetts [MDE11]. masses [ÁSNCÁ^{+}13, IHSS^{+}19, MVT^{+}17, RMJ^{+}18]. massive [LCBC16, PKB^{+}17, TLB^{+}16]. master [SPR^{+}15]. Masthead [Ano19n]. mat [VLJ^{+}10]. Matano [KCM^{+}10]. material [DTPP12, WM12]. maternal [PvEF12]. Mating [SNTK15, KSY11, LRY12]. Mats [GSPM13, HGD14, LFB^{+}10, MDF^{+}14]. matter [ÁSNCÁ^{+}13, BVSM15, BLWV10, CRCGG^{+}17, CSÁS^{+}10, CÁSO^{+}16, CT18b, CPG^{+}10, CHS^{+}18, CDA16, CGT16, CHV^{+}17, CCC10, CK12, CK13, CFF^{+}17, DVC^{+}17, DCCB17, DWDH10, EKS^{+}18, EMB12, EBMR12, FUS^{+}16, FHS10, FPG11, FB12, FEC^{+}16, GKT^{+}15, GMS^{+}18, GAM^{+}19, GPS15, HA16, HKP^{+}16, HEB^{+}19, HT17a, HEH^{+}17, HMFF10, HMFF12, JSK^{+}15, KBA^{+}12, KPW^{+}11, KHCH14, KWRS13, KMC^{+}15, KWB^{+}16, KHK^{+}19, LPO^{+}11, LZK18, LTX^{+}17, LÁSDC18, LBR^{+}12, MGHS18, MPONC^{+}17, MPK^{+}13, MKW^{+}19, MA18, MMXC15, MBLD15, MBAS^{+}17, MCC^{+}10, MSD^{+}14, MBO^{+}16, MGJH18, NNE12, NWT^{+}19, OCB^{+}18, OWFS11, PMY^{+}19b, PCO^{+}15, PML^{+}19, RBCS16, RCSÁS^{+}10, REDW10, RZW11, SLC^{+}16, SHSK14, SKK^{+}15, SCF^{+}15, SEYJ11, SFB12, SFLB16, SBC^{+}17, SSC^{+}10, SYW18, TLG^{+}11, TEZ^{+}18, THH^{+}13, TAV^{+}10, TTV^{+}13, TSDF^{+}16, TZD^{+}15, UVGS10, WM12, WDX^{+}11, WMC^{+}15, WSM^{+}19, WYW^{+}10]. matter [WZBW^{+}11, WSTG18, XSAHV13, YHS^{+}17, YJO^{+}19, ZZY^{+}10, ZZAC13, dCGS19]. Mauritanian [KTS^{+}14]. maxima [LBC^{+}18, LBS17, SVS^{+}19, WCP^{+}15]. maximum [AvSGK18, ETKL12, LCM^{+}17, MSSH12, SS16]. may [BS18b, DKSA19, HCAF18, LFH^{+}12, OBM^{+}11, PBV16, SM10, WCS^{+}18]. McMurdo [DTKMK15, DKK^{+}14, VMAS^{+}16]. meadow [AFG^{+}16, BDP^{+}19, CvHB^{+}18, CB12, CHL^{+}17, HBM11]. meadows [AFSM17, BBS^{+}18, MMGO^{+}17b, MHH^{+}17, PHLSSS19, RASV^{+}17]. mean [CT18a, SRA10, XZC^{+}16]. means [BTH^{+}16]. measured [AGMR14, EWB12, HBM11, PSB^{+}16, RBM14, dKYH^{+}12]. measurement [YWL^{+}17]. measurements [BPB^{+}17, BLH^{+}13, BFD^{+}11, BMD17, BLG^{+}15, DTPP12, EM13, HCK11, JD16, Joh10, KTH^{+}19, LP10, SKK^{+}15, SSB^{+}18, SWL11, VPC10, WSB^{+}13]. measures [KS16]. Measuring [Tho19]. Mechanical [YKBJL12, SvKP^{+}18, WHH^{+}11]. mechanically [CGB^{+}18]. Mechanism [YAC^{+}19, AvSGK18, KZR^{+}16, LdJMS^{+}13, LDCT11, MCC^{+}10, MPAS17, NRL15]. Mechanisms [CF13a, FSBT16, KTK^{+}13, RGO^{+}11, SM11a, HNL^{+}13, IGP^{+}12, KMF10, LCW^{+}17b, MGGS18, MPM^{+}15, MCT^{+}14, PRL18, VBBR17, ZSZ12, dBWL^{+}13]. Mechanistic [SKV^{+}19, LGW^{+}19]. mediate [RRD14]. mediated [BDC^{+}14, CLN^{+}19, JMJ^{+}19, PLE^{+}17, SAP^{+}11]. mediation [HHA18]. Mediterranean [GGPM^{+}10, ALL^{+}10a, AFSM17, ACA^{+}11, BA14, BCVAn10, CNL^{+}15, CdC^{+}11, DMSHC16, FDP^{+}18, FEC^{+}16, GAH11, GBMG12, KZB^{+}10, MGGS18, MPONC^{+}17, MAD^{+}15, PCF14, RBG^{+}10, RGGL^{+}12, RGLM^{+}12, SCAB^{+}16, SWM^{+}10, TvBR^{+}19, VMF^{+}11, WVGB10]. medium [PMRRA19, SPG^{+}13]. medium-size [PMRRA19]. medium-sized [SPG^{+}13]. megacity [BBJ^{+}19]. Meganyctiphanes [BPW^{+}19, CTA^{+}19, KK11]. megatidal [MBO^{+}16]. Meiofauna [NNE12]. Meiofaunal [MTT17]. Mekong [GBD^{+}10]. melanostomus [TB18]. Melosira [KIH^{+}15]. melting [GRT^{+}14]. meltwater [MSAM18, SS12a]. Members [Ano19o, Ano19p, Ano19q, Ano19r, Ano19s, Ano19t]. membranacea [SMF10]. Membranipora [SMF10]. Menten [FFA13]. menu [LTPK^{+}18]. Mercenaria [MAS^{+}16]. Mercury [AHD^{+}18, HGM10, BFD^{+}11, BPRG^{+}18, CMW^{+}19, GBT^{+}17, JSB^{+}14, JW14, KMC^{+}15, Kus14, LF19, OSHS19, RQC^{+}15]. Meretta [AMQ^{+}11]. meromictic [MRC^{+}16, VHM^{+}10, WBZ^{+}13]. meroplankton [IH11, SJ11]. Meso [CSS^{+}16, NSV^{+}14, NWT^{+}19, YYMN13]. Meso- [CSS^{+}16, NSV^{+}14, YYMN13]. meso-eutrophic [NWT^{+}19]. mesocosm [HMH^{+}16, PLE^{+}17, SJB^{+}19, Spi15, WVGB10]. mesocosms [LEN^{+}15, SCF^{+}15]. Mesodinium [JJ17]. mesograzer [HHHT19]. Mesopelagic [HDP15, BIM^{+}16, CPHD15, GCH^{+}18, HONR11, JTG^{+}11, KTRK11, MG17, MVT^{+}17]. mesophotic [PGRR^{+}19]. mesoscale [ATP^{+}15, CHS^{+}18, TNMV^{+}10, WBG^{+}16]. mesotidal [IR16]. mesotrophic [KHP18, PS17, TMO^{+}18]. mesozooplankton [KT13]. meta [FBV11, MJJMM17]. meta-analysis [FBV11, MJJMM17]. Metabolic [AACS11, Ano21a, GSG^{+}17, SAH^{+}19, ZCK^{+}16, AdGAD14, BHG^{+}18, DdG10, FMGR^{+}11, FCD12, HDP15, KTL17, Les16, RSJ^{+}18, SJB^{+}19, WLO^{+}19]. Metabolism [GLF17, AEH19, ARB^{+}19, BDP^{+}19, BWB^{+}10, BMD17, BGM^{+}13, CBFK19, CB12, CG17, CSU13, DKG15, DHG^{+}17, EMO^{+}11, FDS^{+}14, FAF^{+}12, GNHGM13, GTR^{+}13, GN16, HCK10, HEB^{+}19, HBR13, HSBA10, HH14, HBM11, IH18, KBA^{+}12, KB15, MLCD13, OSC14, dGD13, RBM14, RAB^{+}17, SMM11, SJB^{+}19, SGVR16, SKJD^{+}14, SLH^{+}15, SPGRP^{+}17, SSJR^{+}10, SCBR12, TDM^{+}13, VBC^{+}12, WKB^{+}10, dGCB^{+}11]. metabolome [WRH^{+}18]. metacommunities [HS10]. metacommunity [HT17b]. Metagenomic [HNL^{+}13, KHCH14, VLJ^{+}10, BSC^{+}15]. Metal [VF10, ANP^{+}14, BLLB12, HS18, HCW^{+}10, HCLS11, LYH17, ORC^{+}17, WZR19, WFR10]. metalimnetic [KBE^{+}17, WCCP14, WCP^{+}15]. metalimnion [Ano21a, GSG^{+}17]. metalloenzymes [MTW12].

Metaproteomic [WDX^{+}11]. metatranscriptomics [MTK^{+}17]. metazoan [ACA^{+}11]. meteorological [AA11, WSTD10, ZSM14]. meter [SLS^{+}11]. Methane [ÅCA^{+}18, APP12, CDW^{+}16, DPSW16, DBSP^{+}16, GMBL16, HW16, HNHS^{+}15, LKS^{+}16, PSB^{+}16, TMF^{+}14, ZOB^{+}15, BMF^{+}16, BMN16, BNW^{+}14b, BK11, BSSW11, CKB^{+}16, FWFB10, FCRW^{+}16, GMS^{+}18, GAM^{+}19, HFP10, HSP^{+}16, JBB^{+}16, JMJ^{+}19, JP10, KHTO13, KHCH14, LVDM19, LGC16, MLD^{+}16, àNTS13, OMB^{+}16, PHPH^{+}16, RETS16, SBvH^{+}15, SWM^{+}18, SOM17, SDS^{+}16, SAP^{+}11, SSB^{+}16, TSB^{+}19, TLR^{+}13, TMH^{+}18, TSDF^{+}16, TMH^{+}10, UCOG16, VLDM19, VHM^{+}10, WCJ16, XBR^{+}18, ZMS^{+}18]. methane-derived [HNHS^{+}15]. methane-enriched [UCOG16]. methane-rich [KHCH14]. Methanogenesis [MMGP^{+}12, AES11, CGT16, XBR^{+}18]. Methanogenic [CKCEP10, CCC10]. methanotrophic [HMV12, ZOB^{+}15]. methanotrophs [BNW^{+}14a, OMB^{+}16]. methanotrophy [AES11, CGT16, SSS^{+}16, SIH^{+}17, SSB^{+}16, TLR^{+}13, UCOG16]. Methionine [BMM^{+}13]. method [MMPSB14, PSB^{+}16, SW11]. Methodological [KPP^{+}18]. Methyl [BFD^{+}11, FYT^{+}12, CMW^{+}19, HKU^{+}10]. Methylmercury [LF16, LF17b, GBT^{+}17, HGM10, LF19, OCB^{+}18, TBF^{+}13]. Methylotrophic [XBR^{+}18]. metrics [WBZ^{+}14]. Mexican [BJDMH10, MMC^{+}10]. Mexico [BSC^{+}15, CPPdAR^{+}13, DCCB17, FCRW^{+}16, GDD^{+}16, GCR^{+}10, HCC^{+}13, LGC13a, LGC13b, PGB^{+}19, RG13, SFB12, TKK^{+}17, WWC^{+}13, YMB^{+}18, ZMS^{+}18]. Meyer [Ano21b]. Meyer-Kaiser [Ano21b]. Michaelis [FFA13]. Michigan [LG16, RAV^{+}17]. Micro [BNW^{+}14b]. Micro-aerobic [BNW^{+}14b]. microalga [KS13]. microalgae [BVvB^{+}19, CCW^{+}19, ESMS13, HAC^{+}11, KG18, LLB17]. Microalgal [SMLC^{+}18, HSLH^{+}14, RRD14]. Microalgal-driven [SMLC^{+}18]. microbe [GLS^{+}13, VCPC^{+}16]. microbe-induced [GLS^{+}13]. Microbes [TLR^{+}13, BIM^{+}16, FFA13, SCG^{+}19, TYX^{+}19]. Microbial [CVS^{+}10, DTL^{+}19, DHW11, GGPM^{+}10, LF19, MGK15, MG17, MAS^{+}16, RLL^{+}10, ŠGN^{+}19, SPG^{+}11, BHB^{+}12, CCK^{+}12, CPF16, FT11, FBFR13, FEC^{+}16, GGL^{+}15, GBP^{+}12, HMFB16, HGD14, HHS^{+}18, HWZ13, HEH^{+}17, HDDH^{+}17, HSP^{+}16, KHP18, LEN^{+}15, LSH^{+}17, LKF^{+}18, LHLT13, MFMC^{+}10, MC16, MACM11, MGS12, MCYR17, NUH^{+}12, NTM^{+}10, OALD10, PD11, PLE^{+}17, RLC^{+}11, RSJ^{+}18, SCF^{+}15, SMR^{+}17, ŠNZ^{+}14, SBH^{+}11, SGRB10, SSC^{+}17, VSP^{+}11, VLJ^{+}10, VML^{+}19, VMAS^{+}16, WGDAA19, WRO^{+}11, WYW^{+}10, ZTW^{+}11, ZXL^{+}19, vOSH12]. microbially [MBH^{+}15]. microbiome [BBC^{+}13]. microbubble [RMH^{+}17]. microcystin [PHG13, DMS^{+}18]. microcystin-LR [DMS^{+}18]. Microcystis [ALdML^{+}14, BVSM15, FDBW16, GPL11, GOD^{+}18, HL13, LGW^{+}19, MQP^{+}16, PHG13, WKK^{+}11]. microelectrode [HGD14, TGG^{+}11]. microenvironment [WPL^{+}14]. microenvironments [LCS^{+}19]. Microgeographic [JLC^{+}15]. microlayer [TAV^{+}10]. micron [JYS18]. micronekton [GCH^{+}18, RRCH^{+}19]. Micronutrients [Alo17]. microorganisms [SDCF16]. Micropaleontological [GdVT^{+}11]. Microphytobenthic [TMK^{+}13]. microphytobenthos [EOM16, MMPSB14, OCR10, OEM12, ORGE16]. microphytoplankton [AAIA14a, AAIA14b]. microplankton [ACA^{+}11, RBCS16]. microzooplankton [CLHL12, CLLH14, CSS^{+}16, EB12, KKS10, Lat14, LCZ^{+}19, NSV^{+}14, PLE^{+}17, SNM^{+}15, SBFC18]. microzooplankton-dominated [PLE^{+}17]. mid [HGM10, CWHP14, CSC^{+}11, SBBNM14]. mid-Atlantic [HGM10, CWHP14, CSC^{+}11, SBBNM14]. middle [ZZW16, KZR^{+}19]. middle-lower [ZZW16]. midnight [DHG^{+}17]. midsummer [ZXN^{+}12]. Midwater [HPCD13]. migrating [HV16, NL14]. Migration [OR16, BM16, BRF^{+}17, DHG^{+}17, EHW^{+}15, FOT^{+}15, HSR^{+}10, HPS^{+}10a, IPGP10, JSFC18, KTRK11, KGM14, MWSB18, OMSC13, PGB^{+}19, PK14, RRCH^{+}19, SMN^{+}15, TGGZS^{+}10, VMC^{+}13, WCB^{+}10, WFB^{+}11]. migrations [Ano17l, HCS11, ZXZ17b]. Mikata [SSYT14]. mild [PKWS19, RLB^{+}10, RPG13]. mildest [PST^{+}13]. miliacea [LHSBP18]. miliaris [TRD^{+}14]. millennial [XZGW17]. millennial-aged [XZGW17]. Mineral [BS18a, JCF^{+}10, PDFS14, PE13]. mineralization [ÁSNCÁ^{+}13, CEPPR14, CGT16, KBT16, LCM^{+}12, LZK18, NNE12]. mineralogy [NEH^{+}19]. minerals [DMN15, LK15, SKLG10]. minerogenic [EP14]. minimum [DTFR12, FCD12, KBE^{+}17, LEN^{+}15, MMC^{+}10, NHS^{+}12, TSB^{+}19, VGM14]. mining [RSJ^{+}18, SLBH^{+}19, VB17]. minuses [GWD^{+}16]. mirror [JGR^{+}14]. mirror-based [JGR^{+}14]. Misconduct [How15b]. missing [HLFM^{+}10, MSGS^{+}13]. Mississippi [OPZ13, GCH^{+}12, HHS^{+}18, SFB12, TT12]. mitigate [OMB^{+}16, WCS^{+}18]. Mitochondrial [PTS12]. mixed [ANP^{+}14, AdBVA10, Ano19c, GBB^{+}18, HWZ13, LBS17, SNK12]. Mixing [HMHI13, KCM^{+}10, PMRRA19, WSB^{+}13, YAC^{+}19, AGLM17, AGML18, BBR12, CFD15, CSD10, DHH15, FYVU17, GGTC^{+}18, HD19, HGvB^{+}13, IHSS^{+}19, JMM14, JSB^{+}14, KGT12, KWGS18, LBS17, LC11, MBH^{+}15, NA17, OrIA10, PHJ12, RCV^{+}14, SBBNM14, SLPM15, VLDM19, VMCM^{+}17, XDC^{+}19]. mixotrophic [HLG15, JB19, PS17]. mixotrophs [HGdG^{+}19]. Mixotrophy [MPAS17, DKSA19]. Mnemiopsis [CMG^{+}15, HJT^{+}13a, HJT^{+}13b, JTH^{+}11, JCS^{+}18]. mobile [MH16]. mobility [SWD11]. Mobilization [WLL^{+}11, XZGW17]. mode [KNL10, SSS^{+}16]. mode-water [KNL10]. Model [BLG^{+}15, YAC^{+}19, BRR^{+}13, BSCC15, BBR^{+}14, BDC^{+}14, BG10b, CRB^{+}17, CAQS16, FDL17, HV16, HBR^{+}14, HE10, HRPW15, HSBA10, HH14, HRN11, KGM14, LEN^{+}15, MCH12, NTA14, RSG11, RAKE05, RN14, RMDK10, RDZ^{+}13, SKV^{+}19, SBNC^{+}19, SSPK^{+}12, SBF18, TBHM^{+}13, TBLG14, WP14, WDJF12, CR10]. Model-based [BLG^{+}15]. Model-guided [YAC^{+}19]. modeled [SPR^{+}15]. Modeling [EO13, FLM^{+}19, GBR14, KGT12, RGB^{+}19, SPG^{+}13, SMA15, BPW^{+}19, CLB19, GAH11, JHD^{+}11, KGC^{+}12, KFJ13, LEG^{+}10, MMFBB18, RAV^{+}17, SOM17, Scu16, SPMW11, SCQ^{+}17, WGC^{+}13, WSB^{+}13]. models [BMW10, CEPPR14, DMS^{+}18, ESMS13, FFA13, FYVU17, JSB^{+}14, LHLT13, MA18, SBT^{+}19, SRAB10, SRA10, SL10a, SL10b, SC10]. moderate [WGM16]. moderates [SBK18]. moderating [BPL^{+}19b]. Modern [WKJS^{+}14, BWBB15, PE16a, RSTP12]. modes [BL18, SWL11]. modifications [BVC^{+}14]. modified [MD10, RHMSE15, TCFP19]. modulate [ŠGH^{+}18, WZBW^{+}11]. modulated [TDM^{+}13]. modulates [MLGZ16, ŠNZ^{+}14, TAE^{+}18, VCPC^{+}16]. modulating [ZBSR15]. modulation [RBD18, RR12, RGGL^{+}12]. Molecular [DSM^{+}18, JAZ^{+}10, MPONC^{+}17, SSC^{+}10, VdRA^{+}19, ASSG12, FSST11, KWB^{+}16, LCW17a, LFC17, LLW^{+}18, RGM^{+}11, SOM^{+}15]. Molybdenum [GWSEA10]. Monica [HMV12]. monimolimnion [JW14]. monitoring [LC12, TCFP19]. monomictic [NUH^{+}12]. monoxide [ZXN^{+}12]. Montastraea [TEGL11]. Monterey [HONR11]. moored [SBM^{+}15]. moorings [Joh10]. morphofunctional [FBL15]. Morphological [PSS^{+}14, HCL^{+}18, KPV^{+}11, LdlSB^{+}12, LLW^{+}18, RG19]. morphology [TPM^{+}14, TRA19, WBS^{+}10, WHH^{+}11]. morphometric [SvKP^{+}18]. morphometry [SBK18, VMMS^{+}13]. mortalit [VSD10]. Mortality [MHA^{+}18, ADCH18, CBS^{+}17, HMD11, MMB17, MGS12, PST^{+}13, SJM11, TIS^{+}13, WD15]. mosaic [TSC^{+}19, GKT^{+}15]. Mosby [FWFB10, LFB^{+}10]. most [PST^{+}13]. Motile [vSGAK17, BHV^{+}17]. motion [HRN11, HPL11, KYRMD18, MP17, PHLSSS19]. motions [OIS10, PHJ12]. mountain [BCVAn10, BLWV10, DMSHC16, FOT^{+}15, KHVS11, HML^{+}14]. mountainous [WGH^{+}10]. mountaintop [VB17]. movement [HMD11, HBBM19, KYRMD18, TT14]. movements [MHT13, SSH^{+}16]. moving [JCS^{+}18]. mucus [HA16]. Mud [FWFB10, LFB^{+}10]. muddy [SBNC^{+}19]. mudflat [BvBB^{+}16]. mudflats [GSPM13]. multi [BGW^{+}15, BL18, BCM^{+}17, CHHT18, CS12]. multi-armed [BL18]. multi-isotope [CS12]. multi-scale [BCM^{+}17]. multi-sensor [BGW^{+}15]. multi-year [CHHT18]. multibasin [ILPL13]. Multidecadal [DHZ^{+}19, HHE^{+}19]. Multifaceted [MPSA17]. Multiple [KS16, Ker17, MA18, PSNE15, TLG^{+}11, FYVU17, GNWDL19, HCD19, HT17b, LACI10, MMWR17, MMBP18, SSFR19, WMT^{+}12]. multipopulation [FSST11]. Multiscale [FSCB11, LDT^{+}11]. multiseries [SHF^{+}11]. multispecific [WZTK15]. multispectral [KYG^{+}12]. multivariate [RBI^{+}10]. Murderkill [FYVU17]. murky [LPLH18]. mussel [Les16, PSS^{+}14, PLS^{+}16, SGA10, TRA19, WCS^{+}18]. Mussels [NA17, CS12, KKB^{+}18, KKS10, RAV^{+}17]. muta [MBLP11]. Mya [BMDC10]. Mycosporine [KMF10, TAV^{+}10, UVGS10]. Mycosporine-like [KMF10, TAV^{+}10, UVGS10]. myctophid [CFRL10]. myriad [FMP^{+}13]. mysid [BRT^{+}10]. Mytilus [Les16].

N [PFH^{+}17, BMBI12, BCF^{+}17, BLJ13, CHHT18, CvHB^{+}18, CAQS16, CHL^{+}17, DLP13, DLBF17, DFK^{+}17, EOM16, HSB^{+}13, JSH12, JTH^{+}13, KK13, KBJ^{+}18, LWrDM^{+}12, LWWE^{+}18, MBB^{+}18, MHL^{+}16, MCGF^{+}11, NTK^{+}18, OHKC^{+}12, SMR^{+}17, THA17, TG17, UA10, WE19, WFK^{+}16, YJO^{+}19, ZCZ^{+}18, ZLLM10, ZCY^{+}15]. nano [MBO^{+}16]. nano- [MBO^{+}16]. nanocyanobacterium [MFK^{+}13]. Nanofibrils [SH10b]. nanoflagellates [PS17]. nanomolar [ZF17]. nanophytoplankton [Piw19]. nanoSIMS [BBTK^{+}16]. narrow [DB13, LDT^{+}11]. national [BGB^{+}14]. native [SSFR19]. Natural [PPT12, SPTS15, ASA^{+}18, BHC13, BHC14, BS18a, BSMC12, CEB^{+}17, GC16, HSC^{+}14, JTG^{+}11, KM10, MLS^{+}14, MBTK18, MBP^{+}17, NCC14, PvDM^{+}13, PDP^{+}10, RDT^{+}14, RLL^{+}10, SLC18, SDH^{+}14, TSDF^{+}16, WXMS10]. naturally [BHW^{+}12, BCC^{+}12, CGP^{+}19, MRH^{+}15, OCLW11, SCF^{+}15]. nature [RWB^{+}19, SHM^{+}19]. nature-based [SHM^{+}19]. nauplii [JMNG^{+}13, SGCI14, VIS^{+}13]. neap [VMCM^{+}17]. Near [VBBR15, Aus13, BHW^{+}12, CTH15, PHPH^{+}16, PMRRA19, RDT^{+}14, SFMF15, SW14, VPC10, VML^{+}19, ZXN^{+}12]. near-field [PHPH^{+}16]. Near-inertial [VBBR15, Aus13, CTH15]. near-infrared [RDT^{+}14, SW14]. near-shelf [TSC^{+}19]. near-shore [VML^{+}19]. near-surface [VPC10]. near-term [BHW^{+}12]. Nearshore [GWN^{+}12, CDA16, FZL^{+}14, HCD19, JHD^{+}11, MF19, MSM^{+}17, OLF^{+}11, OFGF12, PRL18, SNvD^{+}10, SSH^{+}16, SPG^{+}13]. negative [BHW^{+}12]. negatively [GOD^{+}18, WGM16]. neglecta [RF13]. Negro [BMF^{+}16]. nekton [ALG^{+}13]. nematode [MGT15]. nematodes [GVS^{+}10]. Neogobius [TB18]. nepheloid [BNW^{+}14a]. net [BRNS18, BS18b, CF13b, GSPM13, HEBS10, HCH^{+}19, KEH^{+}14, KTS^{+}14, LWWC^{+}16, SPGRP^{+}17, SSJR^{+}10]. network [MBB^{+}18, RGM^{+}11, SSU^{+}16]. neustonic [MTK^{+}17]. Nevada [SMM11]. newly [OLC18]. next [Edm15]. Nhecolândia [FMP^{+}13]. niche [CTA^{+}19, FA10, ITO^{+}17, MAB^{+}17, MCYR17, PWWF18, WOC^{+}18]. niches [ABD^{+}17, BVP^{+}15, CFRL10, INF12, WKAM^{+}19]. Nickel [Ho13, MBC^{+}18, TNMV^{+}10]. night [DHG^{+}17, GSB^{+}17, KK11, MKLKP16]. Niñ [VLWV14]. Ningaloo [FDH^{+}14]. Niño [MMHT10, SCAB^{+}16]. Niskin [SSC^{+}17]. Nitrate [MCH12, MD15, NCT^{+}14, BSCC15, BCRC16, BSMC12, DBFL11, DSS^{+}11, FDS^{+}18, GWD^{+}16, HC10, HCF^{+}10, HKU^{+}10, Joh10, KSFT13, KJG10, KvdPB18, LTH^{+}12, MCGF^{+}11, MAS^{+}16, MRH^{+}15, QFH18, RS16, RRB^{+}16, RDB^{+}18, SMR^{+}17, SBC^{+}17, SPPS10, SYW18, TFLS14, TG17, TMO^{+}18, WBG^{+}16, WZG^{+}14, WGC^{+}13, WGCC14]. nitrate-low [MRH^{+}15]. Nitric [SSKdB14]. nitricline [WTC^{+}17]. nitrification [AMMH^{+}13, DTM18, MHL^{+}16, PFvO^{+}18, SSG^{+}17, SWE^{+}18, SBS^{+}13, SDCF16]. nitrifying [BSMC12]. nitrite [BSC^{+}15, BC10, BSMC12, MCH12, MC16]. Nitrogen [ASH^{+}14, ACC^{+}17, BGM^{+}13, CF13b, EWB12, FWWF18, JWS15, LK14, LWWC^{+}16, MC16, OHKC^{+}12, OWS^{+}17, PFvO^{+}18, RBY^{+}17, RSTP12, RBRH10, SM10, TFLS14, VFME18, WTN^{+}15, XPQ^{+}10, AFSM17, AHJS15, ACW^{+}18, BSR^{+}17, BMW10, BAA^{+}13, BBTK^{+}16, BDK^{+}17, BHM^{+}17, BLWV10, BSA^{+}16, CRJ^{+}14, CPHD15, CMM^{+}11, DTL^{+}19, DHH15, EED10, EMS16, EOM16, FWO^{+}18, GFT^{+}14, GML^{+}12, GLKK10, GWSEA10, GWD^{+}16, GWB^{+}14, GBC^{+}17, GHS14, GRE^{+}16, GBD^{+}10, GN16, HAC^{+}11, HCK14, HOD^{+}17, HJB^{+}12, HLGA17, Ho13, HTLM18, HVD^{+}18, JKKM13, JHLK^{+}19, JSH12, JTH^{+}13, KRB^{+}18, KMF10, LWE^{+}11, LKF^{+}18, LCCF10, LWrDM^{+}12, LBB18, LMR14, MTSG18, MFK^{+}13, MGL^{+}16, MTEM15, MGW^{+}13, MBP^{+}17, MKG^{+}15, MCYR17, MDSG18, MAS^{+}16, NPT11, NCT^{+}15, OSB^{+}15, OVRJ13, PSH^{+}11, PK14, PGP^{+}14, RS16, RvSM17, REE^{+}12, RWC16, SCR^{+}12, SM11b, SK19, SLH^{+}15, SKK^{+}13]. nitrogen [SFI^{+}18, SRM^{+}18, SS12a, SSYT14, TK12, TKB18, THFG16, VCPC^{+}16, WDMF13, WCC^{+}17, WE19, WGCC14, WRH^{+}18, WLHW13, XXZ^{+}19, XLS^{+}19, YLH^{+}16, vdHHC^{+}19, SM11b]. nitrogen-cycle [JKKM13]. nitrogen-enriched [GWD^{+}16]. nitrogen-fixing [GBC^{+}17, SK19, YLH^{+}16]. nitrogen-limited [MFK^{+}13]. nitrogen-rich [OVRJ13, SS12a]. nitrogen-to-phosphorus [BMW10, OSB^{+}15]. Nitrosopumilus [AMMH^{+}13]. nitrous [BSN^{+}14, DHW11, SPPS10, WGC^{+}13, XXZ^{+}19]. Nitzschia [AJ15, LBHS13, SHF^{+}11]. NMR [SKK^{+}15]. No [Meh10, JTH^{+}11, WTC^{+}17]. noble [TBK15]. noble-gas [TBK15]. Noctiluca [TRD^{+}14, VdRA^{+}19]. nodosa [IOB^{+}11]. nodule [SLBH^{+}19]. noltii [LdlSB^{+}12]. Non [OWM^{+}18, BHV^{+}17, BSY^{+}16, CRB^{+}17, PCY^{+}10, RWC16, SPP^{+}16, SSFR19]. non-bloom [PCY^{+}10]. non-cohesive [SPP^{+}16]. non-motile [BHV^{+}17]. non-native [SSFR19]. Non-seagrass [OWM^{+}18]. non-seasonal [BSY^{+}16]. non-vegetated [RWC16]. non-zooxanthellate [CRB^{+}17]. nonconsumptive [MHA^{+}18]. nonindigenous [SMF10]. noninvasive [SDS^{+}11]. Nonlinear [OR16, CLLH14, FWS^{+}14, LFL17, SMMF19, VMI13]. nonmotile [GK15, HSR15]. nonnative [GSPM13]. nonphotochemical [DVSV13, RGG^{+}10]. nonpoint [JSH12]. nonpoint-source [JSH12]. nonrepresentative [MLS^{+}14]. noon [OPZ13]. Nordic [BPW^{+}19]. Norfolk [SSN12]. normalized [ESMS13]. norms [PZHD18, ZKL^{+}14]. North [CVS^{+}10, DBH^{+}16, VPG^{+}19, GJWS14, GJWS16, MvdPK^{+}15, VBC^{+}12, WLV17, ZCK^{+}16, ÁSNCÁ^{+}13, BFW^{+}13, BLW15, BPA12, BGW^{+}15, BHB^{+}12, BSA^{+}16, BTH^{+}16, CTA^{+}19, CR16, CPHD15, CSS^{+}16, DDK10, DBV^{+}11, DKSA19, DvOR^{+}16, FPP^{+}19, FMM^{+}14, HPCD13, HDP15, HLG15, HLJ12, HQB^{+}18, HEBS10, JWGH19, KBVW12, LKS^{+}16, LWB^{+}17, MRKR^{+}14, MMD15, MBBG^{+}12, PFvO^{+}18, PWS^{+}11, RWM^{+}14, RKMN^{+}13, SLG10, SMR^{+}17, SSB^{+}16, SS17, TBLG14, UFW^{+}18, WCC^{+}17, WB19, WM17]. North-Atlantic [CTA^{+}19]. north-south [MvdPK^{+}15]. north-temperate [GJWS14, GJWS16, WLV17, ZCK^{+}16]. northeast [ATP^{+}15, ÁSNCÁ^{+}13, BSB^{+}10, CG17, HLH13, KEH^{+}14, MLL^{+}14, PNR19, RLL^{+}10, SGG^{+}11, STB^{+}16, SSH^{+}14, SDCF16, WRS13, CEB^{+}17, KCL^{+}14, MvdPK^{+}15]. northeastern [dGCB^{+}11]. northern [APS^{+}19, BBSK18, BPPF12, Bre14, BLLB12, DCCB17, DBSP^{+}16, FLP^{+}10, FVSL19, FLM^{+}19, KH16, LGC13a, MF19, MWC^{+}16, PGB^{+}19, PMPD13, RG13, RVvdP^{+}17, SLA^{+}15, TKK^{+}17, VSdG17, WAB^{+}17, XDC^{+}19, dlFN10, FPD^{+}10, IGP^{+}12, LHS19, LGC13b, RCJ15]. northward [HZC^{+}13]. northwest [ALL^{+}10a, ACA^{+}11, BA14, GBMG12, JAZ^{+}10, LCBC16, MMHT10, PCF14, GMGM^{+}13, PPT12]. northwestern [GAH11, KK13, RGGL^{+}12, RGLM^{+}12, ZCY^{+}15]. norvegica [BPW^{+}19, CTA^{+}19, KK11]. Norway [GLKK10, JAS^{+}15, MWC^{+}16]. Norwegian [HATF17]. Nostoc [SJM11]. nourishes [MSSH12]. novel [SSS^{+}16, TLR^{+}13, YWL^{+}17, YLJ11]. novo [LWWE^{+}18]. NPQ [BHV^{+}17]. NNO_{3}^{−} [FYVU17]. nuance [FDP^{+}18]. nudibranch [SGG^{+}11]. null [Lat14].

number [GBK^{+}18, SdlFdlF^{+}10]. Numerical [FRP^{+}14, BH13, CLB19, DMS^{+}18, ZWA^{+}14]. nurseries [TDF^{+}17]. nursery [FLM^{+}19, WDH^{+}17]. Nutrient [ALG^{+}13, DRE^{+}10, GLI^{+}15, GJR^{+}19, GBB19b, HDK^{+}12, HHS^{+}18, KSG^{+}10, OBL^{+}19, TIN^{+}14, ZCL^{+}19, ZSM14, AP12, ARML10, ASR^{+}17, AC15, AJ15, BBT^{+}10, BMW10, BMBI12, BSA^{+}16, CBF10, CJ17, ETKL12, ES13, EMO^{+}11, ETI^{+}16, FFA13, FDS^{+}18, FDBW16, GCSO14, GC16, GNHGM13, GSPM13, GAM^{+}19, GSZL13, GvBBB17, HSC^{+}14, JSH12, JJ17, JWS15, KGRV18, KWRS13, KHK^{+}19, KOFN11, KvdPVB13, LdlSB^{+}12, LEN^{+}15, LWE^{+}11, LTPA17, LAC^{+}19, LG10, MAB^{+}17, MZB^{+}15, NCC14, OWS^{+}17, OFGF12, OSB^{+}15, PvDM^{+}13, PSG^{+}16, RDC^{+}19, SS12b, SS12c, SvKP^{+}18, Spi15, SL10a, TWP13, VLJ^{+}10, VMCM^{+}17, WS18, WGM16, WC17, WZBW^{+}11, WFL^{+}12, WLHW13, ZLLM10]. nutrient-depleted [FDBW16]. nutrient-limited [MAB^{+}17]. nutrient-replete [FDBW16]. Nutrients [BPGE13, SGA^{+}17, AFG^{+}16, CCV^{+}18, CL10, DC15, DMSHC16, FBV11, GLMG15, GLF18, HLG15, JM16, KHH19, LKLH10, LC11, MCWB10, MVNG11, MBE^{+}13, SLU11, SGRB10, UA10, WS18]. nutrition [SZH^{+}10]. Nutritional [GVS^{+}10, BISZ17, FBFR13, GCH^{+}18, JLG10, PWF16]. nuttallii [ZLLM10]. NW [VMCM^{+}17, GGPM^{+}10, IHSS^{+}19, KZB^{+}10]. Nyanza [GNHGM13].

O [HH14, BDP^{+}19, CHHT18, HH14, MQJG13, TG17, VHR^{+}11, WFK^{+}16, ZCZ^{+}18]. oases [ÅCA^{+}18]. Obelia [SGCC16]. objects [SGH12]. obliquus [HNZ^{+}16, HCL^{+}18]. observation [NL14]. observational [SMA13]. Observations [Aus13, Aus19, CT18a, EMH12, GAH11, JHD^{+}11, SVMT15, WYL16, ABS^{+}19, BGW^{+}15, KZR^{+}19, QHVM18, TIF^{+}15, UFW^{+}18, WSM^{+}19]. Observatory [CVS^{+}10, MKBSK19, GGPM^{+}10]. Observed [AMB^{+}11, GPH^{+}13, LSDW18, SBM^{+}15]. Observing [Joh10, RGM^{+}11]. occupy [RHV^{+}13]. occurrence [SLBH^{+}19, VHrI10]. occurring [HZC^{+}13, LKLH10, SCF^{+}15]. Ocean [CVS^{+}10, HRG^{+}15, KH16, Man10, MLGZ16, SW14, WCI^{+}14, AWK^{+}17, BRS11, BPB^{+}17, BMW10, BHW^{+}12, BG10a, BSFH10, BIS^{+}10, BWD^{+}11, BWD^{+}12, BVP^{+}15, CÁSO^{+}16, CLHL12, CSJ^{+}14, CHV^{+}17, CESC14, CHPH13, CAS^{+}17, CSME13, Edm11, FB12, FCC11, GGC^{+}14, GDD^{+}16, GBC^{+}17, GLF18, HVJ^{+}19, HCH^{+}19, JMNG^{+}13, KSG^{+}10, KLEH16, KBHT19, KRR16, LCW17a, LCH^{+}14, LUM15, MCLT12, MBC^{+}16, MAC^{+}10, MRE18, NBDM16, OMSC13, nVOH12, RSTP12, RSTS^{+}18, RGM^{+}11, RPL16, SPTS15, SFLB16, SSH^{+}14, SCG^{+}19, TIN^{+}14, TSB^{+}19, TBSR13, UFW^{+}18, VLMTEW11, VZJ^{+}17, VFS^{+}15, WCS^{+}18, WGH^{+}16, WKG^{+}16, WDJF12, WC17, WGH^{+}10, WBB^{+}17, WZC13, XFH14, XLS^{+}19, ZBSR15, ZHG15, ATP^{+}15, ABB^{+}14, AdBVA10, ABD^{+}17, BAG^{+}14, BPA12, BAG^{+}17, CFD^{+}11, CLJ^{+}19, CFRL10, CG17, CEB^{+}17, DVDB16, EB12, FYC^{+}18, HOD^{+}17, HWZ13, HQB^{+}18, JBB^{+}16, JWGH19, JTG^{+}11, KYRMD18]. Ocean [KK13, KHCH14, KGL^{+}16, LKT17, LKS^{+}16, LAC^{+}19, LGC16, MVL^{+}10, MLS^{+}18, MEM^{+}17, MVT^{+}17, MvdPK^{+}15, MCGF^{+}11, NRS16, NLO^{+}12, OCLW11, PvDM^{+}13, PNR19, PFvO^{+}18, PSNE15, RS16, RBCS16, RS19, RDB^{+}16, RZW11, RWM^{+}14, RKMN^{+}13, RKTLM18, SSFF12, SSG^{+}17, SDSC12, SHT^{+}17, SGG^{+}11, STB^{+}16, SMR^{+}17, SFI^{+}18, SDCF16, SMH^{+}11, SHF^{+}12, SSS^{+}19, WMBR13, WBG^{+}16, WGRS^{+}17, YHS^{+}17, YYMN13]. ocean-reef [GLF18]. oceanic [ASK^{+}11, BBMS17, BRS^{+}13, CHS^{+}18, CLFW17, FDS^{+}14, HWZ13, IBPG17, KKH11, KvdPVB13, NMST18, NLHAA^{+}17, PRL18, dGD13, WD15]. oceanica [AFSM17, BRS^{+}13, CB12, CB19, GPA^{+}14, IOB^{+}11, MMGO^{+}17b, THFG16, ZBSR15]. Oceanogr [Ano21b]. Oceanographic [GDD^{+}16, HNSM12, NEH^{+}19, WFL^{+}12, CHH^{+}17, Joh10, Tho19, VML^{+}19]. Oceanography [Ano21a, MMC^{+}10, Xen19]. oceans [CL10, HW16, KKH11, NG13, PTS^{+}19, WLL^{+}11, XDK^{+}17, BCRC16]. OCPs [ZZW16]. Odum [HBR13]. Off [WMBR13, AAIA14a, AAIA14b, FCD12, GFT^{+}14, GRE^{+}16, GAK^{+}19, JAZ^{+}10, JHD^{+}11, MQJG13, RPMK17, SKGT17, TAV^{+}10, VGM14, WCJ^{+}17]. offer [MDF^{+}14]. offset [CCW^{+}19, HCAF18, SM11b]. offshore [BSA^{+}16, PMA18, WTC^{+}17, dGCB^{+}11]. offspring [LRY12]. Oikopleura [LTPK^{+}18, LBR^{+}13, LSK11]. oil [FCRW^{+}16]. Oithona [AACS11, SGCI14, VIS^{+}13, ZTS13]. Okely [PHJ12]. Old [GBS17]. oligomesotrophic [SPP10]. oligopeptide [ALdML^{+}14]. oligopeptide-based [ALdML^{+}14]. oligotrich [JB19]. oligotrophic [CPOMA15, GSZL13, HS18, HCH^{+}19, HML^{+}14, JYS18, KP13, KSFT13, KKH11, LCW17a, MBE^{+}13, SNM11, SJM11, SBS^{+}13]. oligotrophy [MFMC^{+}10]. oliogohaline [TMH^{+}18]. Olympia [Car10]. omega [IWF19]. omega-3 [IWF19]. omega-6 [IWF19]. Omnivorous [ŠGN^{+}19]. omnivory [SD10]. oncaeid [NTI^{+}15]. one [SDS^{+}11]. one-year [SDS^{+}11]. Onondaga [EMH12]. onset [KIH^{+}15, SLPM15, TF11, ZXN^{+}12]. Ontario [BRT^{+}10, RPH^{+}10]. onto [LK15]. Ontogenetic [Hir12, HLGA17, IPGP10, WLS^{+}11]. ontogeny [HBBM19]. oocysts [SSL^{+}12]. oomycetes [MKW^{+}19]. open [BSFH10, BVP^{+}15, KB15, MRE18]. open-ocean [BSFH10]. opens [FHS10]. Optical [DVC^{+}17, HKP^{+}16, HE10, JTG^{+}11, NRS16, RSN16, RS19, SOH^{+}18, WBZ^{+}14, AGCA16, ASK^{+}11, BDB^{+}14, BSG14, BFD^{+}11, CDA16, DCRC16, GDCM13, LLH^{+}15, MGHS18, OR16, SCQ^{+}17, TRD^{+}14, USB^{+}10, UVGS10, WSTG18]. Optically [CMW^{+}19, RNT^{+}19]. optics [EP14, JLRK12]. Optimal [XZC^{+}16, HV16, HV19, THA17]. Optimality [SPMW11, TBHM^{+}13]. Optimality-based [SPMW11]. optimization [SMMF19]. Orbicella [Edm15]. order [HHS^{+}18, SCQ^{+}17]. Oregon [APP12]. organ [PCPZ18]. Organic [KLEH16, KWB^{+}16, LÁSDC18, NB17, PMY^{+}19b, SVLS^{+}16, VW17, ALL^{+}10a, AHJS15, ÁSNCÁ^{+}13, BSCG17, BBLN11, BMBI12, BHD^{+}17, BVSM15, BLWV10, CEPPR14, CPPdAR^{+}13, CRCGG^{+}17, CSÁS^{+}10, CÁSO^{+}16, CKP^{+}15, CRJ^{+}14, CTG15, CT18b, CPG^{+}10, CDA16, CGT16, CHV^{+}17, CCC10, CK12, CK13, CFF^{+}17, DFWPK16, DIC^{+}18, DTL^{+}19, DVC^{+}17, DCCB17, DBA16, Dem19, DWDH10, DvOR^{+}16, EKS^{+}18, EMB12, EBMR12, FUS^{+}16, FHS10, FPG11, FHR^{+}15, FB12, FLP^{+}10, FEC^{+}16, GKT^{+}15, GJWS14, GJWS16, GMS^{+}18, GAM^{+}19, GBP^{+}12, GdG11, HGG^{+}17, HA16, HKP^{+}16, HBR^{+}14, HEB^{+}19, HLG15, HT17a, HLJ12, HEH^{+}17, HSTK15, HMH^{+}16, HGT^{+}18, HDDH^{+}17, HMFF10, HMFF12, JMM14, JTH^{+}13, JP10, JSK^{+}15, KBA^{+}12, KZB^{+}10, KPW^{+}11, KKH11, KHCH14, KWRS13, KBT16, KMC^{+}15, KHK^{+}19, LTH^{+}12, LHSG15, LPO^{+}11, LZK18, LTX^{+}17, LBR^{+}12, MSGS^{+}13, MGHS18, MPONC^{+}17, MPK^{+}13]. organic [MKW^{+}19, MA18, MMXC15, MBLD15, MBAS^{+}17, MCC^{+}10, MHH^{+}17, MSD^{+}14, MBO^{+}16, MGSM10, MGJH18, NNE12, NWT^{+}19, OALD10, OCB^{+}18, OWFS11, OVRJ13, PCO^{+}15, PML^{+}19, PBL^{+}18, PHLSSS19, RRAS17, RR13, RM14, RCH^{+}15, RASV^{+}17, RCSÁS^{+}10, REDW10, RZW11, RHSD^{+}10, RHDTS^{+}11, SLC^{+}16, SHSK14, SSFF12, SKK^{+}15, SCF^{+}15, SCR^{+}12, SLP^{+}14, SLA^{+}15, SEYJ11, SFB12, SFLB16, SBC^{+}17, SSC^{+}10, SHL^{+}18, SYW18, SSS^{+}19, TGC^{+}10, TLG^{+}11, TEZ^{+}18, THH^{+}13, TAV^{+}10, TTV^{+}13, TSDF^{+}16, TZD^{+}15, UFW^{+}18, WM12, WDX^{+}11, WWC^{+}18, WMC^{+}15, WCJ^{+}15, WSM^{+}19, WGH^{+}10, WYW^{+}10, WZBW^{+}11, WDL^{+}17, WSTG18, XSAHV13, XZGW17, YHS^{+}17, YJO^{+}19, ZZY^{+}10, ZHN^{+}10, ZZAC13, ZCK^{+}16, dCGS19, vEG10, JBT11]. organic-aggregate-associated [TGC^{+}10]. organic-iron [JBT11]. organically [SMH^{+}11]. organics [ASSG12]. organisms [CHL^{+}17, SPMW11]. organize [BBMS17]. organochlorine [ZZW16]. orientation [NMST18]. origin [CCV^{+}18]. original [HZC^{+}13]. origins [ZKMT^{+}13]. orthophosphate [IGP^{+}12]. oscillate [BDU^{+}19]. Oscillation [MMHT10, SCAB^{+}16, WB19, HLJ12, MMHT10]. Oscillatory [VPWW10]. ostracodes [CF10]. Ostreococcus [CLFW17]. ONO_{3}^{−} [FYVU17]. other

[SH10b]. otolith [GM12]. otoliths [MWR17, WJHS18]. our [GMMV19]. outbreaks [SMF10, SLG10]. outflow [PFH^{+}17]. outgassing [APB^{+}17, SML^{+}19]. output [CRB^{+}17]. outwelling [SML^{+}19]. overcoming [JJ17]. overestimation [HCH^{+}19]. overlap [BL13, TCG^{+}17]. overnight [SHSK14]. oversaturation [TMF^{+}14]. Overwinter [BPL^{+}19a]. overwintering [JWGH19, LLL10, SAPI14, WB19]. oxic [BKD^{+}16, OMB^{+}16]. oxidation [BPA12, BNW^{+}14b, BC10, BK11, CDW^{+}16, CMB10, DTL^{+}19, FDL17, GFT^{+}14, HNHS^{+}15, HQB^{+}18, NFW13, àNTS13, RSM13, RRB^{+}16, RDB^{+}18, RETS16, SAP^{+}11, TSB^{+}19, TMH^{+}10, WBZ^{+}13, XLS^{+}19, ZOB^{+}15, Ano10]. oxidative [SMC^{+}10, TGGZS^{+}10]. oxide [BSN^{+}14, DHW11, SSKdB14, SPPS10, WGC^{+}13, XXZ^{+}19]. oxidizer [NFW13]. oxidizers [MBP^{+}17, UMHH^{+}14]. oxidizing [AMMH^{+}13, BPA12, JAZ^{+}10, MACM11, PWS^{+}11, SDCF16, VFME18]. oxyclines [KBM^{+}14]. Oxygen [BC10, BSMC12, CMB10, DMMV15, IR16, JMM14, KTS^{+}14, AWK^{+}17, BPB^{+}17, BLH^{+}13, BDU^{+}19, BWS^{+}14, BLM^{+}10, BMB^{+}18, CRJ^{+}14, CSGW18, CWRX19, CF10, DTFR12, FWFB10, FCD12, GRT^{+}14, GLF17, HSLH^{+}14, HGD14, HSBA10, HQB^{+}18, HBM11, IH18, JHD^{+}11, Joh10, KB15, KBM^{+}14, KBE^{+}17, LL11, LCM^{+}12, LRM^{+}19, MC16, MMC^{+}10, MMN^{+}10, NHS^{+}12, NCT^{+}15, ORC^{+}17, QWRJ10, RS16, RLB^{+}10, RMNZ12, SWE^{+}18, SSB^{+}18, SSGB^{+}17, Sha10, SHK13, TKB18, TSB^{+}19, TMO^{+}18, VGM14, VHR^{+}11, WBG^{+}16, WDCH18, WFK^{+}16, WMM18, WCP^{+}15, WGCC14, WSB^{+}13, YMB^{+}18, ZF17, Ano10]. oxygen-deficient [WFK^{+}16]. oxygen-depleted [NCT^{+}15]. oxygenase [nVOH12]. oxygenated [LK14, LZK18, SWM^{+}18, TMF^{+}14, TMH^{+}10]. oxygenation [GdVT^{+}11, SWM^{+}10]. Oyashio [IHSS^{+}19]. oyster [BHW^{+}12, BMC^{+}16, BGP^{+}15, Car10, WHAM15]. oysters [PKWS19].

P [ACD10, GRPB^{+}17, PHJ12, BMBI12, CAQS16, DKSA19, HSB^{+}13, KK13, PFH^{+}17, SKJD^{+}14, SH10a, THA17, VABMS^{+}12, WZC13]. P-depleted [DKSA19]. P-enrichment [VABMS^{+}12]. P-limitation [WZC13]. P. [HCK11]. pace [Clo19]. Pacific [BPA12, CPHD15, CEB^{+}17, CVS^{+}10, DDK10, DBV^{+}11, HPCD13, HDP15, HQB^{+}18, LWB^{+}17, RDB^{+}16, SMR^{+}17, ATP^{+}15, BHW^{+}12, BMC^{+}16, BCRC16, BBTK^{+}16, CRJ^{+}14, CJW^{+}19, CLJ^{+}19, CG17, DTFR12, DSLLL19, DBH^{+}16, DvOR^{+}16, FMM^{+}14, HS18, HLH13, HOD^{+}17, HEBS10, HCH^{+}19, IHSS^{+}19, JKKM13, KEH^{+}14, KK13, KBL^{+}10, KBVW12, LWE^{+}11, LKS^{+}16, Man10, MVL^{+}10, MMC^{+}10, MLL^{+}14, MMHT10, NO17, PNR19, RS16, RZW11, RLL^{+}10, SSG^{+}17, SGG^{+}11, STB^{+}16, SKK^{+}13, SSH^{+}14, SDCF16, SSN12, SL10b, SS17, SSS^{+}19, UFW^{+}18, VGM14, WMM18, WTC^{+}17, WRS13, YHS^{+}17, YYMN13, dGCB^{+}11]. pack [PHB^{+}10]. Page [Ano19a, Ano19d, Ano19e, Ano19f, Ano19g, Ano19h, Ano19i]. PAHs [ZZW16]. paired [Spi15]. Palau [TDM^{+}13, WTC^{+}17]. Palearctic [PTS12]. paleoclimate [SRAB10]. paleoecological [PDER10]. paleoisotopic [PDER10]. paleolimnology [FSST11, HML^{+}14]. palifera [YLH^{+}16]. pallida [HRPW15]. pan [SJB^{+}19]. pan-European [SJB^{+}19]. Pantanal [FMP^{+}13]. Paracalanus [TSK13]. paradigm [WFB^{+}11]. Paradox [TMF^{+}14, ŠSP17]. Paraíba [PMP^{+}17]. paralytic [BMDC10, MMHT10]. parameter [LHLT13]. Parameterization [GC16, ZWA^{+}14]. Parameterizing [SdlFdlF^{+}10]. parameters [KLEH16, LBHS13, MZH15, PvEF12, TW10a]. parasites [FWvD^{+}18]. parasitism [VP15a]. Pareto [SP11]. partial [CESC13]. partially [HBB^{+}11]. Particle [OIS10, PHJ12, ASK^{+}11, AAC^{+}19, BIM^{+}16, CSJ^{+}14, CGB^{+}18, EP14, HPCD13, KCL^{+}14, KGM14, MAC^{+}10, MVT^{+}17, NMST18, NLM^{+}12, NRS16, PE13, RSN16, SSGM18, TCG^{+}17, USB^{+}10]. particle-associated [MVT^{+}17]. particle-attached [TCG^{+}17]. particle-reactive [CSJ^{+}14]. particle-tracking [KGM14]. particles [ALL^{+}10a, BIS^{+}10, BVvB^{+}19, DM17, FTC10, GCH^{+}18, HCLS11, JYS18, JTG^{+}11, LBNT11, MB10, NLM^{+}12, PDFS14, PE13, PE17, PFJ10, RLC^{+}11, RSN16, RDT^{+}14, SKLG10, SCQ^{+}17]. Particulate [CHS^{+}18, MLS^{+}18, RBCS16, WLHW13, BA14, BBLN11, BDK^{+}17, CTG15, CT18b, CFF^{+}17, DTPP12, DWDH10, EBMR12, GAH11, GPS15, GLF18, HMFF10, HMFF12, LG16, MMXC15, MBO^{+}16, PFH^{+}17, RZW11, SSFF12, SEYJ11, SYW18, SSS^{+}19, TEZ^{+}18, UVGS10, WM12, WGH^{+}10]. partition [OLC18]. Partitioning [GKS12, MRB11, WZG^{+}14, BC19, CTA^{+}19, EWB12, FA10, KLM^{+}17, MAB^{+}17, PE16b, TJJ^{+}15]. parvus [TSK13]. passage [CWHP14]. passive [HPL11, SWM^{+}10]. past [BPRG^{+}18, RKG^{+}11]. Patagonia [HPM^{+}10, VBGG^{+}13]. Patagonian [BDB^{+}14, CBP10]. Patch [GMJW13, CFD^{+}19, FJBP15]. patchiness [BSSR10, DOD10]. patchy [CLN^{+}19]. Paternal [BAB^{+}16]. paternity [SNTK15]. path [NTM^{+}10]. pathogen [FSBT16]. pathway [CKCEP10]. Pathways [CFW^{+}14, CGT16, GLKK10, GMMV19, JMJ^{+}19, JSB^{+}14, LTH^{+}12, MBB^{+}18, MMGP^{+}12, RvSM17, RLPL14, WLS^{+}11, ZOB^{+}15]. pattern [BK13, NBSMN19, WWC^{+}18]. Patterns [AWG^{+}12, AMNU16, BTC^{+}19, BWS^{+}14, Clo19, DPM18, LBC^{+}18, RCH^{+}15, RPB17, RNT^{+}19, SBM16, WWS11, ALdML^{+}14, BR17, BRNS18, BM16, CLWD13, FNSS15, GDD^{+}16, HMFB16, HHS^{+}18, IBPG17, JM16, KTRK11, LZR^{+}17, LDL^{+}19, MXWC11, MRE18, OSC14, PE16a, PJUR15, PRL18, dGD13, RSJ^{+}18, RG19, RK13, RAV^{+}17, SNM11, SHSK14, SPP10, SPFP11, SJ11, SKKV11, TW11, TLB^{+}16, TB18, VLDM19, VZJ^{+}17, VW17, WVGB10, WE19, WMT^{+}12, ZHN^{+}10, dGCB^{+}11]. Pb [SMLC^{+}18]. PbTx [KPSW10]. PbTx-2 [KPSW10]. PCBs [CMW^{+}19]. pCO [AMB^{+}11, BDP^{+}19, BPL^{+}19b, FVSL19, HRPW15, MRH^{+}15, SSU^{+}16, SHF^{+}11]. Peace [RKWH18]. peaks [PFJ10]. Pearl [CWRX19, KDGL19]. peat [KMC^{+}15]. peatlands [MLD^{+}16]. pectinatus [HAL17]. Pedro [CVS^{+}10]. Pelagic [DFK^{+}17, MLCD13, MBLD15, QS19, RLC^{+}11, BVvB^{+}19, BSY^{+}16, CMK^{+}10, GGL^{+}15, HBCK10, Hir12, HLGA17, JGR^{+}14, KBA^{+}14, KBM^{+}14, LHLT13, LBR^{+}13, MDF^{+}14, RSG11, RCV^{+}14, SAS^{+}11, SNTK15, SBA^{+}11, TW10b, VdSLC^{+}16, VIS^{+}13, WZG^{+}14, WZBW^{+}11, WS13]. Pelagic-benthic [MBLD15]. Pelagophyceae [KG18, KSWFG13]. pellet [SPR^{+}15, WRS13]. pellets [BIM^{+}16, RK13]. penetration [LCM^{+}12, SWE^{+}18]. Peninsula [RVvdP^{+}17, VMCM^{+}17, CMM^{+}11, GAK^{+}19, HVJ^{+}19, MMD18, TLB^{+}16, TSSH19, TAV^{+}10, VCM13, ZCZ^{+}18]. Perceiving [GK15]. perception [AvSGK18, KGC^{+}16, PJ16]. perennial [ARB^{+}19]. perennially [SMA15]. performance [THFG16]. Perils [CJC^{+}12, LGR^{+}12]. period [BS18b, HZC^{+}13]. Periodic [LSH^{+}17, BBLN11]. periodically [RRB^{+}16]. periods [HEB^{+}19]. periphyton [dKNL^{+}15]. peritidal [RPB17]. permafrost [DMMV15, KMC^{+}15, LVM^{+}10, MW15]. permanently [HHM^{+}18, MKLKP16, SSS^{+}16, SPO^{+}18]. permeable [AWK^{+}17, BLH^{+}13, CPG^{+}10, CSME13, DMS^{+}18, GML^{+}12, GCR^{+}10, KGC^{+}12, MHL^{+}16, SSKdB14, SBNC^{+}19]. peroxide [DVSV13, VHV10]. perpetuate [GHS14]. persistence [BMM^{+}13, CEES14, FSBT16, HT17a, MQP^{+}16, uGH^{+}11]. Persistent [BH16, DBRB^{+}15, VSP^{+}11]. persistently [KCM^{+}10]. perspective [HPCD13, HBM^{+}15, LWE^{+}19, LBR^{+}13]. perspectives [DSM^{+}18, GMMV19, HW16, HSCM19]. perturbation [DLP13]. pertusa [LGC13a, LGC13b, MKB^{+}19]. Peru [GRE^{+}16, VGM14]. Perumytilus [PLS^{+}16]. Peruvian [NHS^{+}12]. pesticides [ZZW16]. pH [CGP^{+}19, FNSS15, HAL17, ITO^{+}17, KH16, LK15, LRG16, NLHAA^{+}17, SMLC^{+}18, WCS^{+}18, WYL16, XSAM12]. Phaeocystis [KBHT19, LG10]. Phaeodactylum [CSJ^{+}14, RLSC^{+}13]. Phaeodaria [SBKO18]. Phagocytosis [LKF^{+}18]. Phase [PT11, dCGS19]. phases [GYP^{+}18, XDC^{+}19]. phenolic [RLSC^{+}13]. phenology [AJG13, TNI19]. phenoloxidase [PDP^{+}10]. phenotypic [TB18]. Philippine [LYH17]. phosphatase [DM17, DDK10, DBV^{+}11, GFPSG13, LDY^{+}16, MLS^{+}18]. Phosphatases [SBH^{+}11]. Phosphate [MLK11, BAG^{+}14, BVvB^{+}19, DM17, LDY^{+}16, SHF^{+}11, WMBR13]. phosphate-replete [DM17]. Phosphonate [BWB^{+}10, WKB^{+}10]. Phosphorus [KHVS11, LZK18, LBNT11, PHG13, WRH^{+}18, ACC^{+}17, BMW10, BBS^{+}18, BCVAn10, BSA^{+}16, CR11, CBK18, DBV^{+}11, FSCB11, FWS^{+}14, FPD^{+}10, GFH13, GBL13, GHS14, HSR^{+}10, JHLK^{+}19, JBLJ12, JLR^{+}17, KBH^{+}19, KFJ13, KHG^{+}13, KRB^{+}18, LG16, LJ18, MAFCD^{+}18, MKG^{+}15, NHS^{+}12, OWS^{+}17, OALD10, OBNP^{+}10, OSB^{+}15, OVRJ13, PK14, PWF16, RSG11, RAV^{+}17, SWZ^{+}15, SWP11, SWD11, SS19, TK12, TNMV^{+}10, WDMF13, WZC13, XPQ^{+}10]. phosphorus-limited [GBL13]. phosphorus-rich [FPD^{+}10]. photic [HAC^{+}11]. photo [GC16, HBD^{+}11, SGME11]. photo-acclimatory [SGME11]. photo-inactivation [HBD^{+}11]. photo-physiology [GC16]. Photoacclimation [LAC^{+}19, SLS^{+}11]. photoacclimatization [PGRR^{+}19]. Photoadaptation [MKLKP16, TBHM^{+}13]. photochemical [BVSM15, KBT16, SSC^{+}10, VBGG^{+}13]. photochemically [RM14]. Photodegradation [KPSW10, TBF^{+}13]. Photodissolution [SMW^{+}18, EMB12]. Photoinhibition [MBP^{+}17, AdBVA10, ARW^{+}10, GBR14, HS11, HBB^{+}11, LCCF10]. photoinhibition-driven [GBR14]. photolysis [KTS^{+}14]. photolytic [HKP^{+}16]. photon [RM14]. photons [Kir13]. photoperiod [SNK12, ŚF19]. photophysiological [FMM^{+}14, SBC^{+}17]. Photophysiology [KvdPB18, KBHT19, MRKR^{+}14]. Photoprotection [KMF10, EHW^{+}15, PHB^{+}10, TAE^{+}18]. photoprotective [SCPE15]. photoreactivity [OWFS11]. photoresponse [SMN^{+}15]. photosynthesis [BPB^{+}17, BWD^{+}11, BWD^{+}12, HPM^{+}10, HBZ12, LLB17, MRC^{+}16, PvDM^{+}13, RPI^{+}12, RSTS^{+}18, SBF18]. Photosynthetic [MRKR^{+}14, CF13a, FRA^{+}17, GFH13, HGD14, HXS^{+}10, HNZ^{+}16, JJ17, LBHS13, RDB^{+}16, RKTLM18, SBdB10, SHT^{+}17, SSPK^{+}12]. phototrophic [FDL17]. Phylogenetic [Les19, ASSG12, SPP10, YLJ11]. phylogeny [LDCT11, NTA14, SASB^{+}15]. phylotypes [BWD^{+}11, BWD^{+}12]. Physical [BLW15, CMMKH12, CWRX19, DMS^{+}18, HMV12, KCH^{+}12, KBM^{+}14, NAH^{+}11, OLF^{+}11, RR13, RNG^{+}13, WCP^{+}15, ZCY^{+}15, ZNVF16, BSBK13, FBV11, FSBT16, GRT^{+}14, JLR^{+}17, KHTO13, KBE^{+}17, LDT^{+}11, MTH^{+}11, QWRJ10, Scu16, SPO^{+}18, SSM^{+}19]. Physical-biological [ZCY^{+}15]. physical-induced [GRT^{+}14]. Physicochemical [KEH^{+}14, BVC^{+}14, RPG13]. physiochemical [ZCK^{+}16]. Physiological [BVC^{+}14, LCCF10, WdBJF16, BRNS18, LBHS13, MBHG11, MDE11, NBDM16, PLS^{+}16, SBF18, SGRB10, THFG16]. Physiology [vHOM^{+}19, BRT^{+}10, GC16, HTL^{+}18, HXS^{+}10, KHPIP^{+}14, SHF^{+}11]. phytoflagellates [SMN^{+}15]. phytoplankter [WKK^{+}11]. Phytoplankton [ETKL16, HPM^{+}10, INF12, IH11, KPV^{+}11, KSP^{+}12, LTPA17, MVNG11, MvdPK^{+}15, OPZ13, PNR19, SRCL^{+}13, SSPK^{+}12, SBFC18, WRO^{+}11, AdBVA10, ADCH18, AJC15, ASA^{+}18, ABD^{+}17, BSG14, BLW15, BYD19, BDS11, BL13, BMM^{+}13, BISZ17, BAG^{+}17, BSSR10, BSFH10, Bre14, BVP^{+}15, BRS^{+}13, BG10b, BCVAn10, CSÁS^{+}10, CL10, CL11, CWF11, CLHL12, CL17, CBS^{+}17, Clo18, CHL10, DVC^{+}17, DBFL11, DMSHC16, DVDB16, ETKL12, ETKL15, EB12, FBV11, FMGR^{+}11, FPP^{+}19, FBL15, FLLH18, FMM^{+}14, GLMG15, GNWDL19, GC16, GWD^{+}16, GGTC^{+}18, GBT^{+}17, GBD^{+}10, GvBBB17, GBB19b, GLF18, HGG^{+}17, HS11, HVJ^{+}19, HLJ12, HSB^{+}13, HSTK15, HXS^{+}10, HVD^{+}18, HKS^{+}15, IHSS^{+}19, IGP^{+}12, JM16, KWRS13, KTK^{+}13, KLEH16, KMP^{+}11, KWGS18, KCB^{+}17, KTL17, KvdPVB13, KvdPB18, LRM17, Lat14, LCM^{+}17, LF16, LF17b, LWB^{+}17, LAC^{+}19, LFC17, LCZ^{+}19, LUM15, LDT^{+}11, MCH12]. phytoplankton [MCLT12, MCLT15, MVL^{+}10, MRB11, MPAS17, MBE^{+}13, MDS^{+}10, MSD^{+}14, MKLKP16, MMD15, MRE18, OWS^{+}17, OCLW11, OFGF12, PJ16, PvDM^{+}13, PKB^{+}17, PCM^{+}16, RPMK17, RS19, RCSÁS^{+}10, RVvdP^{+}17, RKMN^{+}13, RGM^{+}11, SLU11, SASB^{+}15, SS16, SNvD^{+}10, SLK^{+}10, SWD^{+}14, SAPI14, SYdTP^{+}11, SLA^{+}18, STB^{+}16, SGA^{+}17, SNK12, SRM^{+}18, SS12a, SPGRP^{+}17, SNM^{+}15, SMH^{+}11, SHF^{+}12, SSYT14, SSGL19, Tad10, TF11, TFLS14, TBSR13, VBGG^{+}13, VMCM^{+}17, WYL16, WCJ^{+}15, WSUC^{+}18, WCG^{+}17, XPQ^{+}10, XSAM12, XFH14, YP18, ZD18]. phytoplankton-bacteria [DMSHC16]. phytoplankton-mass [CL11]. phytoplankton-zooplankton [BCVAn10]. PIC [FCC11]. pico [MBO^{+}16]. pico-size [MBO^{+}16]. picocyanobacterial [GRSD^{+}14]. picoeukaryotes [RDB^{+}16]. picoeukaryotic [PBA^{+}15]. picophytoplankton [CLWD13, CFVU11, SBF18, SL10b]. picoplankton [CBFK19, MPSA17]. pigment [VvO11]. pigmentation [BSH16]. pigmented [DKSA19]. pigments [HKS^{+}15, MBLD15]. pike [FLM^{+}19, MF19]. pipes [CRCGG^{+}17, WE19]. piscivorous [NZH^{+}11]. pistillata [HRG^{+}15, SIW^{+}11]. pit [MAD^{+}15]. plagiosum [WLS^{+}11]. Plain [vOSH12, DRP^{+}17]. Plains [FPD^{+}10, FLP^{+}10, FVSL19, OWFS11]. planar [PFJ10]. planktivorous [GMJW13, MG14]. Plankton [SJ11, AdGAD14, ASW^{+}19, APF^{+}18, CCK^{+}12, CPG^{+}10, FTC10, GRGL^{+}13, GBB^{+}19a, KVA18, LYH17, MFMC^{+}10, MAV^{+}13, MLL^{+}14, PSG^{+}16, PTS^{+}19, Rie15, RPG13, SKJD^{+}14, SPHVA19, SKKV11, TIF^{+}15, TBLG14, VP15a, VMF^{+}11]. plankton-derived [CPG^{+}10]. planktonic [AvSGK18, DdD^{+}10, DdG10, FPSL18, vSGAK17, HJMD13, LPLH18, MTK^{+}17, NG13, PDER10, dGD13, RSTP12, RBI^{+}10, SD10, SPMW11, SBA^{+}11, TIF^{+}15, ZS18]. Planktothrix [GPH^{+}13, VSP^{+}11]. plant [CFD^{+}19, GAM^{+}19, GK10, GK14, GN16, JMJ^{+}19, KGvdH16, MBK^{+}11, MACM11, PCPZ18, SJ11, VCPC^{+}16]. plant-influenced [MACM11]. plant-mediated [JMJ^{+}19]. plant-microbe [VCPC^{+}16]. Plants [CFD^{+}19, GN16, SSP^{+}18]. Plasticity [SSP^{+}18, THA17, BTJ^{+}12, PGRR^{+}19, SvKP^{+}18, TAE^{+}18]. plastidic [FPP^{+}19]. plastids [JB19]. Plateau [LCW^{+}17b, ZZY^{+}10, MNW^{+}19, SHL^{+}18]. platform [GPH^{+}13, GLF17]. play [DKSA19]. Pleistocene [MXWC11]. Pleistocene-driven [MXWC11]. Plesné [KKP^{+}19]. plume [CSS^{+}16, GBD^{+}10, GCH^{+}12, HDK^{+}12, HCC^{+}13, LWWC^{+}16, PHPH^{+}16, WFR10, WDL^{+}17]. plumes [MAF19]. plumosus [SPPS10]. Pluses [GWD^{+}16]. Plußsee [RMNZ12]. POC [FCC11]. Pocillopora [WHD10]. pockmarks [BSSW11, HSP^{+}16, WBS^{+}10]. Poeobius [CHS^{+}18]. Poincaré [BBR12, CTH15]. point [CESC13, LHS19, PVLMT^{+}16, VLMTEW11]. point-source [PVLMT^{+}16, VLMTEW11]. polar [BHS^{+}16, BHB^{+}12, DHG^{+}17, LLB17, MKLKP16]. pollen [MBK^{+}11, MPK^{+}13]. pollution [BJDMH10, FPD^{+}10]. polonium [CSJ^{+}14]. polyamines [KHP18, MWBM19]. polychaete [BBR^{+}14, CHS^{+}18, CH11]. polychaetes [HHA18]. polychlorinated [CMW^{+}19]. Polycyclic [ZZW16, GPS15]. polykrikoides [JLG10, JLG11]. polymeric [MKW^{+}19, TMK^{+}13]. polymetallic [SLBH^{+}19]. polymictic [OSB^{+}15]. polymorpha [KKS10]. Polynya [SSPK^{+}12]. polynyas [PKB^{+}17]. Polyphosphate [DBH^{+}16, OBNP^{+}10, MLS^{+}18]. polyunsaturated [IWF19]. POM [LRG16, MBO^{+}16]. pond [MCCA18, MDF^{+}14, VZJ^{+}17, vBBM^{+}19]. ponds [DBSP^{+}16, LVM^{+}10, MM11, OCB^{+}18, SGS18]. pool [SM10, SM11b, SKK^{+}13, SC10, Tho19, WTC^{+}17]. pool-riffle-pool [SC10]. pools [BBB^{+}14, LWS^{+}17, SWZ^{+}15]. poor [OSB^{+}15]. Population [BBS12, Car10, MGT15, AA18, BRM^{+}19, CNL^{+}15, CRB^{+}17, Edm15, KTRK11, KSY11, KTL17, MCWB10, MTK^{+}17, NG13, PDP^{+}10, SMF10, SGG^{+}11, SVG^{+}18, WKK^{+}11, ZKL^{+}14]. populations [BMDC10, CGP^{+}19, CBFK19, CR16, GRSD^{+}14, HLSW^{+}15, KP13, MACM11, MBP^{+}17, MMJ^{+}12, OMSC13, PvDM^{+}13, PWF18, SPFP11, TDF^{+}17, WB19]. Porcupine [vOSH12]. Pore [FEW^{+}14, AFG^{+}16, AES11, RPI^{+}12, SCR^{+}12, SBdB10, TMH^{+}18, YKT^{+}15, ZZAC13]. Pore-water [FEW^{+}14, AFG^{+}16, AES11, RPI^{+}12, SCR^{+}12, SBdB10, TMH^{+}18, YKT^{+}15, ZZAC13]. porewater [VPWW10]. Porites [CHH^{+}17, Edm11, LCBC16, MPSA17, TLB^{+}16, TEGL11]. pose [GM12]. poses [JTH^{+}11]. Posidonia [AFSM17, CB12, CB19, GPA^{+}14, HCK11, IOB^{+}11, MMGO^{+}17b]. Possible [MNW^{+}19, XSAM12, MCC^{+}10, WGH^{+}16]. postglacial [MXWC11]. Potential [AAC^{+}19, GGL^{+}18, HCH^{+}19, RETS16, TMH^{+}10, ARB^{+}19, BCRC16, BWS^{+}14, DJD^{+}14, DBC^{+}13, HST^{+}14, HMV12, HNL^{+}13, KP13, KBA^{+}14, KZR^{+}16, KNL10, KWF^{+}17, KLM^{+}17, Lat14, LSHK11, MDF^{+}14, MAFCD^{+}18, PRL18, SSG^{+}17, WYL16, WCV^{+}12, YLH^{+}16, uGH^{+}11, vdJFS^{+}18]. potentially [GRPB^{+}17]. potentials [RSJ^{+}18]. pothole [ZZAC13]. prairie [OWFS11, WWS11, ZZAC13]. pre [GPH^{+}13]. pre-alpine [GPH^{+}13]. prealpine [SPFP11]. precipitation [CBK18, DMB^{+}12, KWGN^{+}10, SRA10]. precision [SSC^{+}10]. preconditioning [GGTC^{+}18]. Predation [KKHP14, KMH^{+}17, LRY12, PKWS19, vSGAK17, HHA18, HBCK10, LBS17, Rie15, SBFB17, ŠSP17, VMC^{+}13, ZEXH15]. Predator [DML17, BMPF19, BSH16, GMD11, HJMD13, KMH^{+}17, LWE^{+}19, MAB^{+}17, MWSB18, SGCI14, SBDS^{+}15, SD10, SBA^{+}11, ŠSP17, VMC^{+}13, WLV17]. predator-derived [BMPF19]. predator-prey [HJMD13, SD10]. predators [CFRL10, DRE^{+}10, KM10, Meh10, SBFC18, TIS^{+}13]. predatory [BBB^{+}17, CMG^{+}15, JCS^{+}18]. predict [KIH^{+}15, MA18, PCPZ18]. predictability [KSP^{+}12, PHL^{+}18]. predicted [KPV^{+}11, NBDM16]. Predicting [MZH15, WLO^{+}19, ZHG15, ML19]. Prediction [TPM^{+}14]. predictions [BMW10, MD10, WS18]. predictive [SRAB10, SRA10]. predictors [BPGE13]. predicts [GGC^{+}14, WAB^{+}17]. preference [RBRH10]. preferential [NMST18]. preindustrial [OSHS19]. presence [BC19, FLLH18, KCB^{+}17, SMLC^{+}18]. present [CGP^{+}19, RKG^{+}11]. presented [Bre10]. preservation [NTM^{+}10]. pressure [CESC13, LBS17, MMGO^{+}17a, MMGO^{+}17b, ZMS^{+}18]. pressures [BDC^{+}14]. Prevalence [YLH^{+}16]. prevent [PSH^{+}11]. Prey [AvSGK18, BBMS17, CBP10, DPLG^{+}19, KGC^{+}16, MF19, SGCC16, DML17, GMD11, GNWDL19, GBK^{+}18, GK15, HJMD13, HBBM19, HPS^{+}10a, LSK11, MG14, Meh10, MWSB18, NSO19, SGCI14, SBDS^{+}15, ŠGH^{+}18, SD10]. Primary [SHT^{+}17, SFLQ^{+}19, WSUC^{+}18, AGMR14, BRNS18, BPRG^{+}18, CvHB^{+}18, CB19, DRE^{+}10, DdG10, EM13, FPGR^{+}13, GJWS14, GJWS16, GSB^{+}17, HYK^{+}15, HC10, HAA^{+}19, KEH^{+}14, KTK^{+}13, LFB^{+}10, LMR14, MRB11, OY10, QS19, SLA^{+}15, SKK^{+}13, SSM^{+}19, WHL^{+}11, WTC^{+}17, WTN^{+}15]. prime [dCGS19]. Priming [HA16, CKP^{+}15]. Prince [VML^{+}19]. principal [KFP^{+}18]. principles [HESU13]. Print [BCF^{+}17]. Prionace [VdSLC^{+}16]. prism [BGP^{+}15]. prized [TDF^{+}17]. probabilistic [BMN16]. probabilistic-survey [BMN16]. probability [HPS^{+}10a]. Proboscia [MEM^{+}17]. process [DHW11, SOM17]. process-based [SOM17]. Processes [ADS^{+}17, OrIA10, PHJ12, Ano21a, BK11, Clo19, DJS18, GSG^{+}17, HHW^{+}19, HSP^{+}16, HZC^{+}13, JKKM13, KBH^{+}19, LFB^{+}10, MDB19, MHA^{+}18, MBH^{+}15, MT11, NO17, OLF^{+}11, SLHA19, Scu16, TBSL17, VSdG17, WSB^{+}13]. processing [ASR^{+}17, GBP^{+}12, HJB^{+}12, MM11, MAS^{+}16, OEMB10]. processors [SBM^{+}15]. Prochlorococcus [BCRC16, CGL^{+}16, DNH^{+}18, GRRA^{+}17, HS18]. produced [BSMC12, FPSL18, KLEH16, KGL^{+}16, VLDM19]. producers [GSB^{+}17, KTK^{+}13, SSM^{+}19, WSUC^{+}18]. producing [HHW^{+}19, HLSW^{+}15]. Production [CSÁS^{+}10, HT17a, KNL10, MSD^{+}14, RCSÁS^{+}10, WRB^{+}19, ARW^{+}10, AGMR14, BA14, BRNS18, BWB^{+}10, CB19, CJWS15, CFF^{+}17, DHW11, DML17, DVSV13, DHK11, DdG10, EM13, ESMS13, FBV11, FPGR^{+}13, FYT^{+}12, GRGL^{+}13, GJWS14, GJWS16, GMS^{+}18, GCH^{+}12, HBD^{+}16, HC10, HLH13, HAA^{+}19, HEBS10, HCAF18, HCC^{+}13, HML^{+}14, JTH^{+}13, JBPM15, JP10, KEH^{+}14, KKH11, KTS^{+}14, LEK^{+}18, LRY12, LTPA17, LFB^{+}10, LM12, LMR14, MTT17, MBTK18, ML19, MCC^{+}10, MRB11, MW15, MQJG13, NTK^{+}18, OEMB10, OPZ13, PD11, PWF18, QS19, RCH^{+}15, RGG^{+}10, SLC^{+}16, SPS19, SBT^{+}19, SLA^{+}15, SBDS^{+}15, SPTS15, SKK^{+}13, SFLQ^{+}19, SSYT14, SHF^{+}11, SCG^{+}19, TST^{+}19, VHV10, WCC^{+}17, WKB^{+}10, WWS11, WCG^{+}17, WDL^{+}17, YYMN13, ZTS13, ZMS^{+}18, dBWL^{+}13]. productive [BCC^{+}12, GHS14, JHD^{+}11, SFLB16]. Productivity [WCJ16, BPRG^{+}18, BAG^{+}17, CvHB^{+}18, DBSP^{+}16, DTM18, GWD^{+}16, HYK^{+}15, HVJ^{+}19, HCK11, KHH19, LDT^{+}11, MVL^{+}10, OY10, PH13, RASD10, RDB^{+}16, RAV^{+}17, SHSK14, SCF^{+}15, SGJB14, SAH^{+}19, SHT^{+}17, SHD^{+}11, SS12a, SH11, Tad10, VCM13, VB17, WS18, WLO^{+}19, WHL^{+}11, WAB^{+}17, WTC^{+}17, WTN^{+}15]. products [BB11]. profile [PMP^{+}17, RLSC^{+}13, SGME11]. profiles [AES11, RLB^{+}10, RHSD^{+}10, YKT^{+}15]. profiling [HGD14, HCK14, SGG^{+}11, TGG^{+}11, VLJ^{+}10]. proglacial [FHR^{+}15]. projections [Edm15]. prokaryotes [MSD^{+}14]. prokaryotic [ORC^{+}17, SKKV11, TFLS14, YYMN13]. prolifera [HZC^{+}13, ZXM^{+}11]. prolong [LWS^{+}17]. prolonged [BHM^{+}17]. promising [SW11]. promote [KWM^{+}19, PBV16, RCIB14]. promotes [ŠSP17]. promoting [SK19]. pronounced [ZHN^{+}10]. Propagule [PBV16, BDC^{+}14]. propagules [TDS^{+}10]. properties [AGCA16, BDB^{+}14, BSG14, BGP^{+}15, CDA16, DVC^{+}17, HKP^{+}16, HE10, JPH^{+}18, LdlSB^{+}12, LLH^{+}15, MTH^{+}11, PE13, SCQ^{+}17, USB^{+}10, UVGS10, WSTG18, YKBJL12]. protactinium [CSJ^{+}14]. protect [SBFC18]. protected [GBMG12]. protection [SHM^{+}19, SGVR16]. protein [AWG^{+}12, MRKR^{+}14, nVOH12, SZH^{+}10, SBFB17]. proteins [NTM^{+}10]. proteobacteria [FYT^{+}12]. Proteomic [UCOG16, DWDH10]. proteomics [NTM^{+}10]. protist [PCM^{+}16]. protistan [CVS^{+}10, HMD11, MMD15, MMD18, ŠGN^{+}19]. protists [BSB^{+}10, DLBF17, GRPB^{+}17, HJMD13, WOC^{+}18]. protons [LK15]. protozoan [GRDPL14]. provenance [ZMWM11]. provide [BJDMH10, FWvD^{+}18, KGRV18, PCJK13, WDH^{+}17]. provides [FGBS^{+}18, WS18]. province [JHW^{+}19]. Proxies [BA14, WZC13, ZKMT^{+}13]. proxy [TRD^{+}14, WYL16]. Prydz [FYC^{+}18]. Prymnesiophyceae [KS13, LG10]. prymnesiophyte [BSCC15]. Pseudo [LBHS13, SHF^{+}11]. Pseudo-nitzschia [LBHS13, SHF^{+}11]. Pseudocalanus [TSK13]. Pseudodiaptomus [VIS^{+}13]. Pseudomonas [FYT^{+}12]. pseudonana [FAF^{+}12, MEM^{+}17, SLC^{+}16, Sch19, SLH^{+}15]. pteropod [TSSH19]. Publications [How15b]. Puget [KT13, UMHH^{+}14]. pulex [LLL10, SOM^{+}15]. pulse [EOM16, MMPSB14]. pulse-chase [EOM16, MMPSB14]. pulsed [OHKC^{+}12]. pulses [Dem19]. pump [ALL^{+}10a, HLFM^{+}10, OSB^{+}15, SCR^{+}12, VGJ17]. pumping [FEW^{+}14, XDC^{+}19]. purity [GDCM13]. purpuratus [PLS^{+}16]. putative [HCH^{+}19]. puzzles [LSDW18]. pyrifera [DPM18, MRB11, PMLC^{+}10, RCH^{+}15, RMDK10, RDZ^{+}13]. pyrosomes [DSLLL19].

Qiandaohu [ZWL^{+}14]. Qinghai [MNW^{+}19]. quadrat [LAM12]. quadrat-level [LAM12]. quagga [RAV^{+}17]. qualitative [BRS18]. quality [Ano19c, BH13, BMPF19, BISZ17, BGB^{+}14, Clo19, CFF^{+}17, FGBS^{+}18, FPD^{+}10, FDB^{+}15, GBB^{+}18, HEB^{+}19, JC14, MBK^{+}11, MTEM15, MW15, PMP^{+}17, PMY^{+}19b, PvEF12, PWF18, RWM^{+}19, SCL^{+}19, VABMS^{+}12, WLV18, WCM19, WDH^{+}17, WCG^{+}17, ZHG15]. Quantification [BHC14, BPA12, CMW^{+}19, HSBA10, RMH^{+}17]. quantified [BFD^{+}11]. Quantifying [AES11, BYD19, HBR^{+}14, JTH^{+}13, KGC^{+}12, KPP^{+}18, RPMK17, TSDF^{+}16, TMO^{+}18, YKBJL12, DHW11, KYRMD18]. Quantitative [VSP^{+}11, BBTK^{+}16]. Quantity [CFF^{+}17, PMY^{+}19b, ŚF19, VABMS^{+}12]. quantum [EMB12, KBT16]. quarter [Edm15]. quarter-century [Edm15]. quartet [HGvB^{+}13]. Questioning [WLO^{+}19]. questions [GMMV19]. Quintana [YMB^{+}18]. quotas [TNMV^{+}10].

Ra [HGvB^{+}13, LKLH10, SBNC^{+}19]. Radiance [SGH12]. radiant [Kir13]. Radiation [FSST11, BSH16, BCVAn10, CRS^{+}17, EHW^{+}15, FOT^{+}15, HS11, HVJ^{+}19, HBB^{+}11, HKS^{+}15, KMF10, RWF^{+}12, SEYJ11, SMC^{+}10, VMF^{+}11, VABMS^{+}12, WFB^{+}11]. radiation-induced [HS11, SMC^{+}10]. radiative [HE10]. radiatively [Aus19]. Radiocarbon [DKK^{+}14, ZMWM11, KWB^{+}16, SKK^{+}15]. radionuclides [CSJ^{+}14]. radium [BTH^{+}16, HGvB^{+}13]. radon [DB11, KDGL19, WSM^{+}19, WGC^{+}13]. radon-222 [DB11, KDGL19]. rafts [uGH^{+}11, vHOM^{+}19]. rain [CH11, OBT^{+}11]. rainfall [CTG15, CT18b]. rainforest [BBLN11]. range [SES18]. ranging [KYG^{+}12]. Rapid [BVSR^{+}15, FDB^{+}15, JLG11, OEMB10, PDP^{+}10, PHB^{+}10, BDC^{+}14]. rapidly [SCP^{+}16]. rare [LZR^{+}17, LCW^{+}17b, SJM11]. raschii [BPW^{+}19, CTA^{+}19]. Rate [EMB12, BYD19, CL11, CH11, ETKL12, Fie13, HST^{+}14, HBD^{+}11, KGT12, SDS^{+}11, WCI^{+}14]. Rates [CGT16, MBBG^{+}12, SBS^{+}13, AES11, AACS11, AvSGK18, AGMR14, AA18, BPA12, BAA^{+}13, CB12, CJW^{+}19, DHW11, FRA^{+}17, FWWF18, GLKK10, GMMV19, GSPM13, GHS14, HH14, HRG^{+}15, KRB^{+}18, KTL17, Lat14, LBHS13, LFL17, MDE11, MMD18, MQJG13, NFW13, RN14, SGCI14, SDCF16, Tad10, TDS^{+}10, WC17, WRH^{+}17, ZTS13]. rather [DPLG^{+}19, PKB^{+}17, SGVR16, WYL16]. ratio [BRR^{+}13, BD15, JM16, THA17, YJO^{+}19, Joh10]. ratios [AHD^{+}18, HBCK10, KK13, LRM17, MGW^{+}13, MZB^{+}15, MQJG13, OSB^{+}15, SGCI14, UA10, WDMF13, WC17, WGCC14, WLHW13]. Reach [RAB^{+}17]. Reach-scale [RAB^{+}17]. reaches [ZZW16]. Reaching [LHS19]. reaction [GBK^{+}18, PZHD18, ZKL^{+}14]. reactive [CSJ^{+}14, HQB^{+}18]. reactivity [IR16]. reactors [CRCGG^{+}17, WE19]. reaeration [HSBA10]. real [Joh10, SDH^{+}14]. Realizing [KWF^{+}17]. really [BB10]. reassessment [SL10b]. recolonization [MXWC11]. Reconciling [KTL17]. reconfiguration [LN11]. reconsidered [TMF^{+}14]. Reconstructing [GdG11, SSH^{+}16, VSdG17, GM12]. reconstructions [CSGW18]. record [ACW^{+}18, MWR17]. recorded [JAS^{+}15, RK13]. recorder [TIF^{+}15]. records [BBB^{+}17, LYL^{+}17, Sha10, WCM19]. Recovering [LFL17]. Recovery [MJJMM17, SVG^{+}18, AMQ^{+}11, AEH19, MACM11, WRWPG19, Ano21c]. Recruitment [MKBSK19, MSK^{+}17, MFL11, CGP^{+}19, FJBP15, HAL17, LCS^{+}19, MFM^{+}12, TCFP19, WAB^{+}17, WDH^{+}17, WMT^{+}12, ZEXH15]. recurrence [GGPM^{+}10]. recurrent [SS12b, SS12c]. recurring [SWZ^{+}15]. recycling [GHS14, LZK18, REE^{+}12, VLJ^{+}10, WGRS^{+}17]. Red [MM11, HST^{+}14, LKLH10, OLF^{+}11, SW14, CvHB^{+}18, WMP^{+}19]. red-tides [LKLH10]. Redfield [Joh10]. redistribution [CTG15, KYR^{+}12]. redox [EO13, HHM^{+}18, SWD11, SBH^{+}11]. redox-gradient [SBH^{+}11]. redox-stratified [HHM^{+}18]. reduce [KKS10, PSH^{+}11, WHL^{+}11]. Reduced [PRL18, PHLSSS19, MMN^{+}10, SMH^{+}11]. reduces [BTJ^{+}12, HRG^{+}15, KvdPVB13]. reduction [BSA^{+}16, DSS^{+}11, JP10, KJG10, LTH^{+}12, MC16, MAS^{+}16, RRB^{+}16, RDB^{+}18, TG17, WZG^{+}14, ZMS^{+}18]. reductions [SKV^{+}19]. Reef [UA10, BGP^{+}15, CPPdAR^{+}13, CESC13, CESC14, CRS^{+}17, Edm11, ELJ^{+}16, FZL^{+}14, FDH^{+}14, GFPSG13, GJR^{+}19, GSZL13, GLF17, GLF18, HGT^{+}18, HCS11, IPGP10, JLRK12, KCH^{+}12, KTH^{+}19, LSD18, MRB11, MDS^{+}10, NLHAA^{+}17, OBL^{+}19, PGRR^{+}19, PCD^{+}19, PJFJ^{+}15, RPI^{+}12, RCH^{+}15, RDC^{+}19, RMK^{+}16, SLC18, SPTS15, SHD^{+}11, TDM^{+}13, WGDAA19, WHD10, WDH^{+}17, WFL^{+}12, WLHW13, YLH^{+}16, BWS10, CUW11, LÁSDC18, MLCD13, RGG^{+}10]. reef-building [CRS^{+}17, ELJ^{+}16, GFPSG13, JLRK12, PGRR^{+}19, YLH^{+}16]. reef-scale [RCH^{+}15]. reefs [GDD^{+}16, KYC^{+}15, LGC13a, LGC13b, Man10, RDP^{+}17, YKT^{+}15]. reflect [PHDH14, SWP11]. reflectance [SW14]. reflects [SBvH^{+}15]. refuge [LL11]. refugial [MXWC11]. regenerated [XLS^{+}19]. regeneration [BSR^{+}17, CMM^{+}11, MGL^{+}16, NSV^{+}14]. regime [AP12, GMGM^{+}13, HNSM12, PCJK13, SGCC16, VLWV14, WVV^{+}11]. regimes [ASK^{+}11, BHG^{+}18, CZB^{+}18, LS15, SAH^{+}19, SBBNM14, SBB^{+}18, TBHM^{+}13]. region [BSB^{+}10, CMM^{+}11, GLI^{+}15, HEBS10, IHSS^{+}19, MRH^{+}15, RDT^{+}14, STC^{+}11, SSS^{+}19, TSC^{+}19, ZHG15]. Regional [ACW^{+}18, BRNS18, FWS^{+}14, MGJH18, SWCL12, SLP^{+}14, WAB^{+}17, FWO^{+}18, HHHT19, LSHK11, MRSS12, MHPW18]. Regional-scale [SWCL12, SLP^{+}14, HHHT19]. regions [KSP^{+}12, RDB^{+}16, VPG^{+}19]. regression [MA18]. regressions [WS18]. Regular [SS16]. regulate [MSM^{+}17, WCJ16]. regulated [DBSP^{+}16]. regulates [GNWDL19, KKB^{+}18]. regulating [PCF14, RWF^{+}12]. Regulation [AWK^{+}17, BLWV10, BBS12, DDK10, MMG16, RKTLM18, TST^{+}19, WGJ^{+}19]. regulations [LTX^{+}17]. rejection [LSK11]. Rejoinder [LGR^{+}12]. related [AWG^{+}12, CH11, DCCB17, DMB^{+}12, GOD^{+}18, KMC^{+}15, LRG16, SGVR16, SZH^{+}10, TGC^{+}10, VLMTEW11]. relates [GAM^{+}19, HLGA17]. Relating [SWD11]. Relation [KSG^{+}10, LLH^{+}15, MSK^{+}17, BAG^{+}14, BSH16, HTL^{+}18, HKS^{+}15, LTPA17, MvdPK^{+}15, MRE18, OY10, RKMN^{+}13, SDS^{+}11]. Relations [BL13]. Relationship [BHB^{+}12, CL11, ETI^{+}16, JM16, LKLH10, SBK18, VPC10, YP18]. Relationships [CL10, CJC^{+}12, FWS^{+}14, KGL^{+}16, LGR^{+}12, LSDW18, RSN16, SPTS15, SSFR19, SLBNG11, TCFP19]. Relative [BBT^{+}10, HQB^{+}18, AAIA14a, AAIA14b, BDS11, LBC^{+}18, MMHT10, MAS^{+}16, SSYT14, VSdG17]. relaxation [MFM^{+}12]. release [CHV^{+}17, HFP10, SRCL^{+}13, SFLQ^{+}19, WBG^{+}16, WFL^{+}12]. released [HCAF18]. relevance [IWF19]. relevant [AHJS15, SM10]. relieve [SBH^{+}11]. rely [FEC^{+}16]. remains [VHR^{+}11]. Remineralization [BIS^{+}10, PPPA14, BIM^{+}16, BK11, HA16, JMM14, SSGM18, TNK^{+}14]. Remote [GMGM^{+}13, BPRG^{+}18, KCH^{+}12, SBM^{+}15]. Removal [DMS^{+}18, HSP^{+}16, LZC^{+}14, ACC^{+}17, CGB^{+}18, FDB^{+}15, HCF^{+}10, HCK14, PPPA14, VB17]. removes [LTPK^{+}18]. renewal [HRN11, JMM14, WP14]. reniformis [ETI^{+}16, SWM^{+}10]. Reorganization [RPH^{+}10]. repair [HBD^{+}11, SSPK^{+}12]. repeated [UFW^{+}18]. replete [DM17, FDBW16, LDY^{+}16]. Reply [KGC^{+}16, ACC^{+}19, CL11]. represent [WRH^{+}17]. representation [SBF18, TBLG14]. representative [NMST18]. representing [GRRA^{+}17, MF19]. represents [JBT11]. reproduced [LEN^{+}15]. reproducibility [PCW19]. reproduction [BPW^{+}19, HRPW15, SGVR16]. reproductive [CRB^{+}17, HP19, MAC^{+}10, MBHG11, PCF14]. Republic [KKP^{+}19]. requirement [HVD^{+}18, TW10a]. requirements [AMMH^{+}13, SMH^{+}11]. Research [MKBSK19, HSCM19, SCL^{+}19, SOO^{+}17, SSFR19]. reservoir [BSN^{+}14, BMN16, BBS12, DFWPK16, DHW11, DHH15, DPSW16, ILPL13, KGRV18, OIS10, PHJ12, ŠNZ^{+}14, WMI^{+}17, XFLM14, ZWL^{+}14, RBY^{+}17]. reservoirs [CFW^{+}14, GHS14, HMV^{+}18, MRSS12, RQC^{+}15]. residence [BGB^{+}14, CF14, SHSK14]. residency [FC11]. resident [KMH^{+}17]. Residual [NI10]. Resilience [BJF18, JCS^{+}18, SBdB10, WHR18, GDD^{+}16, KGRV18, KJKS18]. resistance [BMDC10, JLG11, LFH^{+}12, LCCF10, WGH^{+}16]. resolution [ABS^{+}19, BPA12, HCK14, JD16, PHPH^{+}16, SSH^{+}14, SPO^{+}18, TDM^{+}13]. Resolving [LSDW18, RKBA14]. Resource [MKB^{+}19, MCLT15, MZB^{+}15, SPHVA19, BH13, BLMS17, CLWD13, CLN^{+}19, CJWS15, GEC^{+}17, OPA^{+}14, REDW10, SWP11, TYX^{+}19]. Resource-driven [BH13]. Resources [BCC^{+}12, CPHD15, GFDC11, MCLT12, WCCP14]. respect [NLM^{+}12]. respiration [AdGAD14, Dem19, ERA^{+}12, HWZ13, HEH^{+}17, HH14, HCH^{+}19, KCL^{+}14, KRB^{+}18, MGK15, MG17, RPI^{+}12, SNM11, SHSK14, SBR^{+}13, SFLQ^{+}19, SVG^{+}18, TSDF^{+}16, TKK^{+}17, WMP^{+}19]. responds [BG10a]. Response [ATP^{+}15, AHS11, ACD10, BHC14, FBFR13, JMNG^{+}13, KBA^{+}14, KVA18, KWRS13, Lan14, LWE^{+}11, LBR^{+}13, SM11b, SSM^{+}19, AP12, ARW^{+}10, BH13, BVC^{+}14, BPPF12, BPL^{+}19b, BGM^{+}13, BLM^{+}10, CHH^{+}17, CT18b, DFK^{+}17, DC15, EHW^{+}15, FCC11, GTPB^{+}11, GHSR^{+}16, HLJ12, HPS^{+}10a, ILPL13, KK13, KBJ^{+}18, LTH^{+}12, LEN^{+}15, LMR14, MFMC^{+}10, MMB17, MLL^{+}14, MMBP18, MZB^{+}15, MP17, PMLC^{+}10, PRL18, RPH^{+}10, SGME11, SCPE15, SGG^{+}11, SPHVA19, SK19, ŠGH^{+}18, SSH^{+}14, SSPK^{+}12, SMC^{+}10, SBF18, SRAB10, SRA10, TIN^{+}14, TW10b, TFLS14, VPMrI12, VABMS^{+}12, WHAM15, XZC^{+}16, XNK18, ZWL^{+}14, ZBSR15]. Responses [AJC15, IHSS^{+}19, QFH18, SIW^{+}11, YH17, CJHR19, CESC13, KSTA18a, KRR16, LLL10, LABJ18, LH19, NBDM16, PNR19, RR13, SMMF19, SGJB14, SFWP12, SGA^{+}17, Spi15, SGRB10, WCJ^{+}15, WdBJF16, WBZ^{+}14, WHR18, WD15, WRH^{+}18, XFH14, ZCK^{+}16]. Resting [DHK11, BSBK13, SM11a]. restoration [CZB^{+}18, SGA^{+}17]. restored [LHSG15, LH17]. restructures [BSY^{+}16]. resulting [SWM^{+}10]. Results [GJWS14, GJWS16, GVS^{+}10, KK13, LK14, MKBSK19, PGP^{+}14, SCF^{+}15, SGRB10, WCCP14]. resurgence [GK14]. resuspended [NXL^{+}18, SMW^{+}18]. Resuspension [KYR^{+}12, KFJ13, KHG^{+}13, VBBR17]. retention [CMS17, FTC10, GHS14, IH11, JWS15, KGM14, MFL11, MS13, OEMB10, PPPA14, RBY^{+}17, RGB^{+}19, SS19]. Retracted [ZXZ17b]. Retraction [Ano17l]. retreating [MWR17]. reveal [ALdML^{+}14, RHV^{+}13, WCM19, WRH^{+}18]. revealed [AJC15, BCRW15, CHL^{+}17, CBF10, DTM18, GFDC11, KGL^{+}16, LVDM19, MCYR17, OHKC^{+}12, OEMB10, ORGE16, PFJ10, SSC^{+}10, VKC18, WFK^{+}16]. Revealing [CPPdAR^{+}13]. reveals [BSCC15, CPF16, GCH^{+}18, HEB^{+}19, KVMA17, LdJMS^{+}13, MDB19, MPM^{+}15, MTK^{+}17, OVRJ13, PvEF12, SSFR19, WZR19, WGRS^{+}17]. reverses [WLV17]. reversing [KGRV18]. review [FHS10, JBB^{+}16, MHT13]. revisited [HSB^{+}13]. Revisiting [HBR13, MRBR10]. Rhodophyta [SGME11]. rhythms [SAH^{+}19, SMN^{+}15]. Ria [IR16, VMCM^{+}17]. ribbon [LDT^{+}11]. Ribulose [nVOH12]. Ribulose-1 [nVOH12]. Rica [GRSD^{+}14, ANP^{+}14]. rich [FPD^{+}10, KHCH14, àNTS13, OVRJ13, RLC^{+}11, SS12a]. richness [BBCM^{+}13, LTPA17, LV16]. Ridge [HSP^{+}16, SPB^{+}14, CSC^{+}11, SSN12]. riffle [SC10]. right [LHLT13]. rigidity [RN14]. ringed [BAY^{+}14]. rings [KZR^{+}16, KZR^{+}19, Lee18, NA17]. Rio [CKB^{+}16, NEH^{+}19]. rip [FRP^{+}14]. rip-channeled [FRP^{+}14]. riparian [SCAB^{+}16]. rippled [KGC^{+}12]. rise [NTK^{+}18]. Rising [RWB^{+}19, NFRU11]. risk [BTJ^{+}12, DIC^{+}18, vSGAK17, LRY12, Rie15, ZS18]. risky [MWSB18]. rival [HSC^{+}14]. River [APP12, CWRX19, HMFF12, LZC^{+}14, PMP^{+}17, QS19, REE^{+}12, APB^{+}17, CFW^{+}14, CAS^{+}17, DPSW16, FB12, FLM^{+}19, FEC^{+}16, GLI^{+}15, GKT^{+}15, HDK^{+}12, HC10, HCF^{+}10, HCC^{+}13, KHH19, Ker17, KZB^{+}10, KCB^{+}17, LG16, MU17, NXL^{+}18, PHPH^{+}16, PD11, RAB^{+}17, SAH^{+}19, SFMF15, SSKdB14, SGA10, SDS^{+}16, TLH^{+}11, TMH^{+}10, TT12, TZD^{+}15, WGH^{+}10, WFR10, WDL^{+}17, CS12, CSS^{+}16, CFF^{+}17, DBRB^{+}15, EHT10, FDB^{+}15, GPS15, GBD^{+}10, GCH^{+}12, HYK^{+}15, HPM^{+}10, HHS^{+}18, HSC^{+}11, HMFF10, LWWC^{+}16, MKG^{+}15, OPZ13, PHPH^{+}16, SFB12, SRAB10, SSC^{+}10, TEZ^{+}18, TT12, WLG^{+}16, WCC^{+}17, XZGW17, ZKMT^{+}13, ZZW16]. river-borne [KZB^{+}10]. river-dominated [PHPH^{+}16]. river-estuary-ocean [CAS^{+}17]. river-floodplain [SDS^{+}16]. river-influenced [FB12]. river-loaded [NXL^{+}18]. river-scale [KCB^{+}17]. Riverine [MGSM10, PLS^{+}16, AEH19, CCV^{+}18, Ker17, KHG^{+}13, MT11]. rivers [BMF^{+}16, BBSK18, BLJ13, CRCGG^{+}17, ERA^{+}12, HC12, HCK14, HEH^{+}17, HAA^{+}19, IH11, LS15, MH16, MGJH18, SIH^{+}17, VB17, dCGS19]. rock [WRB^{+}19]. rockfishes [WAB^{+}17]. rocky [JD16, LAM12, MBBW11, PPPA14, SMF10]. rocky-shore [LAM12, MBBW11]. Role [EBMR12, JP10, SOM17, TGGZS^{+}10, VMCM^{+}17, ACA^{+}11, BIM^{+}16, CKD^{+}16, CBS^{+}17, CSJ^{+}14, CPOMA15, DHG^{+}17, DSLLL19, DJD^{+}14, DKSA19, FOT^{+}15, FGMN17, HEBS10, IH11, KBH^{+}19, LRS^{+}10, LALGM18, ML19, PBA^{+}15, PCF14, RWF^{+}12, RMLVK12, STCS10, SAPI14, SFI^{+}18, SSW19, SBKO18, SSGM18, SSS^{+}19, TIS^{+}13, VHrI10, WB19, WGH^{+}10, WXMS10, WYW^{+}10, XSAM12]. Roles [BCVAn10, BFW^{+}13, GLMG15, JMM14, LBS17, MKW^{+}19, SSG^{+}17]. Roo [YMB^{+}18]. root [HCK11, MSS^{+}18]. Ross [BHD^{+}17, SJ11, SSPK^{+}12]. rotation [BK13]. rotation-affected [BK13]. rotifer [ZEXH15]. rotifers [FA10, RG19]. rotundata [MPAS17]. round [TB18]. rRNA [SSS^{+}16]. rubescens [GPH^{+}13, VSP^{+}11]. RuBisCO [nVOH12, HBB^{+}11]. rubrum [JJ17]. run [DPSW16]. run-of-the-river [DPSW16]. Rusinga [GNHGM13].

S [SSS^{+}16]. Saanich [CHHT18]. Saginaw [NHP17]. Saharan [MFMC^{+}10, vdJFS^{+}18]. Sakinaw [VHM^{+}10]. Sal [CL11]. salina [GBL13]. saline [MMH^{+}18, MAD^{+}15]. salinity [CJS^{+}17, GPCJ16, HAL17, PMY^{+}19b, PKWS19, VLMTEW11, WVGB10, WVV^{+}11]. salinization [Ker17]. salmon [AP12]. salmonids [CBP10]. salp [DPLG^{+}19, SSH^{+}14]. Salpa [HP19]. Salt [BPL^{+}19a, CF14, JW14, ALG^{+}13, AC17, BvBB^{+}16, CZB^{+}18, CF13b, KJG10, REE^{+}12, SCR^{+}12, SSW19, SGS18, WKG^{+}16]. salt-wedge [REE^{+}12]. saltmarsh [SML^{+}19]. saltwater [PVLMT^{+}16]. salty [dlFN10]. same [GN16]. sample [MLS^{+}14, SBM^{+}15, SSC^{+}17]. samples [ALL^{+}10a, WCV^{+}12]. sampling [GPH^{+}13, PHPH^{+}16, SJ11, SCL^{+}19]. San [CVS^{+}10, CJS^{+}17, GOD^{+}18]. sands [CSME13, KGC^{+}12, OEM12, RPI^{+}12, SWE^{+}18]. sandy [IR16, OEMB10]. Santa [BSG14, HMV12]. Sapphirinid [TIS^{+}13]. Sargasso [DNH^{+}18, HBD^{+}16, KNL10, MLK11, NFW13, OALD10, OBNP^{+}10, RLPL14, TFLS14, TNMV^{+}10, ZXN^{+}12]. Sargassum [LLW^{+}18, vHOM^{+}19]. satellite [BGW^{+}15, SBM^{+}15, WSTD10, ZD18]. satellite-derived [ZD18]. saturating [BMW10]. saturation [Bre10, CLLH14, JCF^{+}10, LGC13a, LGC13b, NLHAA^{+}17, SW11]. saxitoxin [HLSW^{+}15]. saxitoxin-producing [HLSW^{+}15]. scalable [WBZ^{+}14]. Scale [CdC^{+}11, BBT^{+}10, BPPF12, BSRP^{+}12, BCM^{+}17, CAQS16, DTM18, FJBP15, GMGM^{+}13, HHHT19, KCB^{+}17, KPP^{+}18, MWC^{+}16, NLO^{+}12, PH15, PRL18, RCH^{+}15, RKL14, RBI^{+}10, RAB^{+}17, SCF^{+}15, SWCL12, SCPE15, SSB^{+}18, SLP^{+}14, SI10, SSGM18, SS17, TB18, TBSL17, TLH^{+}11, VPMrI12, VAH11, WGJ^{+}19, WJHS18, WTN^{+}15, YYMN13, ZXM^{+}11]. Scale-dependent [CdC^{+}11, TBSL17]. scaled [RMDK10, RDZ^{+}13]. scales [BRM^{+}19, HCD19, HLH13, HSC^{+}14, LACI10, MG14, MMWR17, PST^{+}13, SGRB10]. scaling [CFAE^{+}15, ETKL12, Hir12, HLGA17, KTL17, SLU11, SASB^{+}15]. Scallop [JAS^{+}15]. Scandinavian [HJT^{+}13b, HJT^{+}13a]. scandium [MBC^{+}18]. scanning [CMG^{+}15]. scarce [TFH17]. scattering [BBMS17, EM13, PE16b, PE17]. Scavenging [DJS18, FTC10, CSJ^{+}14, OSHS19, VLJ^{+}10, vdJFS^{+}18]. scenarios [FCC11]. Scenedesmus [HNZ^{+}16, HCL^{+}18]. Scientific [How15b]. scintillans [VdRA^{+}19]. Scleractinia [CRB^{+}17]. scleractinian [ETI^{+}16, HRG^{+}15, TFH17, TEGL11, GBR14]. sclerochronology [SDS^{+}11]. SCM1 [AMMH^{+}13]. scope [IH18]. Scotia [BIM^{+}16]. Scotland [HGvB^{+}13]. Scoured [BDU^{+}19]. scyphomedusae [RG13]. Sea [BJDMH10, BIM^{+}16, FDH^{+}14, LFB^{+}10, SSFF12, VPG^{+}19, ÅCA^{+}18, AMNU16, APF^{+}18, AJ15, BAY^{+}14, BCF^{+}17, CBP12, EM13, FGBS^{+}18, GLI^{+}15, GEC^{+}17, GRT^{+}14, GVS^{+}10, HRMD19, HEBS10, JSK^{+}15, KVA18, LKT17, LKLH10, LRG16, MWR17, MKBSK19, MRH^{+}15, NFRU11, PHB^{+}10, RBRH10, RNG^{+}13, RWB^{+}19, RVvdP^{+}17, SLU11, SLG10, SKV^{+}19, SAPI14, SLBH^{+}19, SVG^{+}18, TvBR^{+}19, TGG^{+}11, TAV^{+}10, TBSL17, UVGS10, WCB^{+}10, WCG^{+}17, YMB^{+}18, ZCZ^{+}18, ZTW^{+}11, ZXL^{+}19, ABD^{+}17, ASL16, ABS^{+}19, ACC^{+}17, BA14, BHD^{+}17, BRS18, BKD^{+}16, BHB^{+}12, BSA^{+}16, BTH^{+}16, CFD15, CvHB^{+}18, CRJ^{+}14, CMMKH12, DNH^{+}18, DDH^{+}19, DWDH10, EO13, GGPM^{+}10, GML^{+}12, GLI^{+}15, GAH11, GBD^{+}10, HBD^{+}16, HJB^{+}12, HPS^{+}10a, HCW^{+}10, HCLS11, HZC^{+}13, JZZY18, KBH^{+}19, KK13, KKHP14, KNL10, LYH17, LLW^{+}18, LGC16, MGGS18, MGHS18, MPONC^{+}17, MBLD15, MLK11, MdBKL13, MVG^{+}15, MDSG18, MSR16, NTK^{+}18]. Sea [NFW13, OALD10, OBNP^{+}10, PKB^{+}17, PWS^{+}11, Piw19, RBG^{+}10, RF13, RCJ15, RGLM^{+}12, RLPL14, SLE10, SWM^{+}18, SJ11, STC^{+}11, SW14, SFI^{+}18, SSPK^{+}12, SFLQ^{+}19, SBH^{+}11, SSB^{+}16, SCQ^{+}17, TBLG14, TRD^{+}14, Tho19, TFLS14, TNMV^{+}10, VPG^{+}19, WDX^{+}11, WXF^{+}15, WMP^{+}19, WLR17, XDC^{+}19, ZXN^{+}12, ZCY^{+}15, ZYZ19, ZXL^{+}19]. sea-ice [PHB^{+}10]. Seabed [SAS^{+}11, RBD18]. seabird [WGRS^{+}17]. seafloor [JBB^{+}16, SCP^{+}16, SSS^{+}19, ZS18]. seagrass [AHH^{+}16, AFG^{+}16, AFSM17, AHJS15, BBT^{+}10, BDP^{+}19, CB12, CB19, CHL^{+}17, CDA16, CUW11, EMO^{+}11, EMS16, GPA^{+}14, HE10, HLH13, HCAF18, HBZ12, LdlSB^{+}12, MMGO^{+}17a, MHH^{+}17, MMBP18, OWM^{+}18, RASV^{+}17, RRD14, SVLS^{+}16, SM11a, SWCL12, SCPE15, SLS^{+}11, SvKP^{+}18, TTTM^{+}19, ZWA^{+}14]. seagrasses [CF13a, HCK11, MSS^{+}18, RMLVK12]. seals [BAY^{+}14]. Seas [PvDM^{+}13, SSB^{+}16, SLA^{+}18]. seascapes [KEH^{+}14]. Season [TKB18, GBD^{+}10, MSK^{+}17, PS17]. Season-specific [TKB18]. Seasonal [ALdML^{+}14, ABS^{+}19, AGMR14, ARB^{+}19, BSR^{+}17, BK13, BLMS17, CB12, CH11, CVS^{+}10, DBMP^{+}11, DDF^{+}10, FNSS15, FMM^{+}14, HKU^{+}10, JM16, KVMA17, LRM17, MPvBS^{+}18, MMD18, NUH^{+}12, PVLMT^{+}16, PMP^{+}12, PMPD13, PSZ^{+}13, PWF18, RMF11, RRGGCA19, RK13, RKTLM18, SKJD^{+}14, SJ11, STC^{+}11, SI10, SvKP^{+}18, TAE^{+}18, TST^{+}19, vBBM^{+}19, ADS^{+}17, BNW^{+}14a, BMD17, BSY^{+}16, CFVU11, CSD10, FDH^{+}14, HTL^{+}18, HV19, HNL^{+}13, HSC^{+}11, LKT17, MDB19, MAV^{+}13, MKK15, MGS12, QWRJ10, RMJ^{+}18, SCAB^{+}16, SSU^{+}16, SWD11, SMA15, TGGZS^{+}10, UMHH^{+}14, VSdG17, VBGG^{+}13, WLS^{+}11, WVGB10]. Seasonality [FVSL19, HONR11, JC14, SBT^{+}19, WRH^{+}17, APF^{+}18, BPW^{+}19, CvHB^{+}18, GTR^{+}13, KVA18, REDW10, WKJS^{+}14]. seasonally [BBC^{+}13, CHHT18, MF19, SSGB^{+}17, SSB^{+}16, VSP^{+}11, WCB^{+}10]. seasons [CCW^{+}19, JMJ^{+}19, SHM^{+}19]. seawater [BOT^{+}15, JHD^{+}11, KPSW10, KK13, LTPK^{+}18, LF19, LM12, Man10, PCD^{+}19, RSN16, SH10b, TvBR^{+}19, ZYZ19]. seaweed [FDH^{+}14, WGM16]. seaweeds [WdBJF16]. Secchi [LSDW18]. Second [SCQ^{+}17]. Second-order [SCQ^{+}17]. secondary [DML17, LEK^{+}18, SBT^{+}19, VB17, WCG^{+}17]. sections [GKT^{+}15]. sector [RBCS16]. Sediment [FEC^{+}16, JPH^{+}18, SFMF15, VBBR17, ALL^{+}10a, AHH^{+}16, AHS11, ACW^{+}18, AC17, BHB^{+}19, BC19, BLG^{+}15, BBB^{+}17, BRF^{+}17, BLM^{+}10, BSSW11, CZB^{+}18, CR11, CCW^{+}19, CKCEP10, CCC10, CF10, EMH12, EMO^{+}11, GCSO14, GKS12, GMS^{+}18, HHM^{+}18, HT17a, HCAF18, JLR^{+}17, KFJ13, LVDM19, LHSG15, LK15, MDB19, MTT17, MMN^{+}10, MAF19, MMPSB14, MW15, MMFBB18, NNE12, àNTS13, OWM^{+}18, SLK^{+}14, SSKdB14, SPP^{+}16, SWZ^{+}15, SHD^{+}11, SWD11, SSB^{+}16, TTTM^{+}19, VKC18, VvO11, XBR^{+}18, YKT^{+}15, ZZAC13]. sediment-based [MMN^{+}10]. sediment-dwelling [SHD^{+}11]. sediment-trap [EMH12]. Sedimentary [LYL^{+}17, RSG11, SSGB^{+}17, CGT16, DCCB17, EMB12, FMGR^{+}11, HA16, HMFB16, MBLD15, RK13, SKK^{+}15, WCM19, WKJS^{+}14]. Sedimentation [MAC^{+}10, OY10, ALdML^{+}14, MDB19, PE16a]. Sedimented [ZLLM10]. sedimentological [KBH^{+}19, NEH^{+}19]. sediments [AWG^{+}12, AES11, AWK^{+}17, ASH^{+}14, AHD^{+}18, BLH^{+}13, BPRG^{+}18, BPV^{+}19, BGR14, BBM11, BK11, BMB^{+}18, CEPPR14, CHW14, CPG^{+}10, CKD^{+}16, DIC^{+}18, DT16, DMS^{+}18, EOM16, GML^{+}12, GLKK10, GCR^{+}10, GBP^{+}12, HNHS^{+}15, HGM10, HSP^{+}16, JAZ^{+}10, JBB^{+}16, JAD^{+}13, JPH^{+}18, JBT11, JP10, KYR^{+}12, KPW^{+}11, KHCH14, KPJ12, KJG10, LTH^{+}12, LCM^{+}12, LK14, LZK18, MBB^{+}18, MHL^{+}16, MMGO^{+}17a, MT11, MGW^{+}13, MPvBS^{+}18, MAS^{+}16, NPT11, NXL^{+}18, OEMB10, OSB^{+}15, RSTP12, RSM13, RRB^{+}16, RETS16, RWC16, SVLS^{+}16, SM11a, SBdB10, SPP^{+}16, SEYJ11, SBNC^{+}19, SMW^{+}18, SBH^{+}11, SPG^{+}11, TMK^{+}13, TBK15, VLDM19, VPWW10, WKJS^{+}14, ZZW16, ZMS^{+}18, ZXL^{+}19]. seed [RMLVK12, ZXM^{+}11]. seeding [SBM^{+}15]. seedling [CZB^{+}18, IOB^{+}11]. seedlings [OLC18]. seep [CBF10, TLR^{+}13]. seepage [IR16, WMC^{+}15]. seeps [ÅCA^{+}18, VFS^{+}15]. segment [TLH^{+}11]. segregation [MACM11]. Seiche [BL18, SWL11]. seiches [BK13]. seiching [DHH15]. Seine [RGB^{+}19]. selected [DTL^{+}19]. selection [BMDC10, CLWD13, FBL15, KM10, MF19, SBDS^{+}15, SNTK15]. Selective [MJH^{+}16, GPL11, SRCL^{+}13]. Selectivity [MG14, GNWDL19]. selenium [JSB^{+}14]. self [CGB^{+}18, OBM^{+}11, SSP^{+}18]. self-adapting [SSP^{+}18]. self-cleaning [CGB^{+}18]. self-utilization [OBM^{+}11]. semen [LFH^{+}12]. semi [BMBI12, GEC^{+}17, Ker17]. semi-arid [BMBI12, Ker17]. semi-enclosed [GEC^{+}17]. semiarid [BGM^{+}13, YH17]. semienclosed [SGA^{+}17]. senescence [FAF^{+}12]. sensing [SBM^{+}15]. sensitive [CESC14, WD15]. Sensitivity [AA11, TBSR13, CF13a, CL17, ES13, HS11, LCZ^{+}19, WP14]. sensor [BGW^{+}15]. sensors [WBZ^{+}14]. Sensory [CMG^{+}15, TSK13, BRT^{+}10]. Sensory-scanning [CMG^{+}15]. sensu [RSE^{+}17]. sentinel [WBZ^{+}14]. separately [LdlSB^{+}12]. separates [OCR10]. separation [MCYR17]. sequence [PTS12, SSS^{+}16, SC10]. sequestration [BOT^{+}15, CG17, HHE^{+}19, HLFM^{+}10, KYC^{+}15, MHH^{+}17, SSGM18, WWC^{+}18]. Sereda [ACD10]. series [CHHT18, FSBT16, GK14, KH16, WRS13]. services [SWCL12]. seston [BVC^{+}14, MKK15]. set [BPGE13, SHSK14, WMT^{+}12]. setting [CSU13]. settled [MTSG18]. Settlement [TDS^{+}10, ZXM^{+}11, FDP^{+}18, PRL18, ZSZ12]. settlement-stage [FDP^{+}18]. settling [AJ15, DMN15, SKLG10, vdJFS^{+}18]. seven [SBF18]. several [RHV^{+}13]. Severe [WHL^{+}11, JHW^{+}19]. sewage [AEH19, BJDMH10]. sex [BD15, DdD^{+}10, HBCK10]. sexes [MHA^{+}18]. Sexual [FJBP15, BVSR^{+}15, SNTK15]. shading [BJF18]. shadows [SIH^{+}17]. Shallow [SGS18, AHS11, ASH^{+}14, BPV^{+}19, BLG^{+}15, BVvB^{+}19, BKA^{+}14, BBQ^{+}10, BGM^{+}13, BSY^{+}16, CWHP14, CAS^{+}17, DKG15, GTPB^{+}11, GCR^{+}10, GSBR11, HCD19, HAC^{+}11, HSP^{+}16, HKU^{+}10, JMJ^{+}19, JHD^{+}11, KFJ13, LRM17, LRM^{+}19, LALM16, LSD18, MMC^{+}10, OEMB10, OEM12, QS19, RDC^{+}19, REE^{+}12, SJB^{+}19, SK19, SI10, SSB^{+}16, SSYT14, TGC^{+}10, TCG^{+}17, TTV^{+}13, TST^{+}19, VP15b, WSM^{+}19, WMI^{+}17, XFLM14, ZCL^{+}19, ZHD^{+}16, ZPK^{+}12, dKNL^{+}15, dlFN10, vH19]. shallow-water [AHS11, ASH^{+}14, LALM16]. shape [ABD^{+}17, CPF16, HCD19, Hir12, PBV16, VZJ^{+}17, ZEXH15, ZD18]. shaped [KTK^{+}13, SPFP11]. shapes [FDH^{+}14, RKG^{+}11, UMHH^{+}14]. shaping [WYW^{+}10]. shark [VdSLC^{+}16, WLS^{+}11]. Shear [CTH15, NMST18, SVS^{+}19]. Shelf [ABD^{+}17, FYC^{+}18, RMJ^{+}18, VMAS^{+}16, APF^{+}18, BRR^{+}13, BPPF12, BGR14, GMGM^{+}13, GFDC11, HRMD19, HDK^{+}12, HGM10, HCC^{+}13, JAZ^{+}10, JBT11, KSFT13, KVA18, LCBC16, LDT^{+}11, MRBR10, MAF19, MBAS^{+}17, NTK^{+}18, PGB^{+}19, RNG^{+}13, SS12b, SS12c, TSC^{+}19, WBG^{+}16, WS13, WDL^{+}17, BDB^{+}14, CBP10]. Shelf-basin [ABD^{+}17]. shelf-slope [BPPF12]. shelf/near [TSC^{+}19]. shelf/near-shelf [TSC^{+}19]. shell [BMDC10, MWC^{+}16, PLS^{+}16, WGH^{+}16]. shellfish [BMDC10, MMHT10]. shells [JAS^{+}15]. shelves [FDS^{+}18]. Shield [AA18]. shift [AAIA14a, AAIA14b, CBF10, HNSM12, PCJK13, SLK^{+}10, THA17, VLWV14]. shifting [LAC^{+}19]. Shifts [LC11, SCF^{+}15, GMGM^{+}13, HMFF10, MGL^{+}13, RMF11, SMA13, TK12, Tho19, WCJ^{+}15, WMC^{+}18]. Shimizu [PHJ12]. ships [BBM11]. shoaling [NI10]. shoals [LP10]. shore [LAM12, MBBW11, MCT^{+}14, VML^{+}19]. shoreline [SHM^{+}19]. Short [ADCH18, BvBB^{+}16, GPH^{+}13, JHD^{+}11, NPT11, CESC14, ETI^{+}16, MS13, PD11, RPG13, SSS^{+}19, YMB^{+}18]. short-circuit [PD11]. short-lived [MS13]. Short-term [ADCH18, BvBB^{+}16, GPH^{+}13, JHD^{+}11, NPT11, CESC14, ETI^{+}16, RPG13, SSS^{+}19, YMB^{+}18]. shortwave [SW14]. shotgun [DWDH10]. show [AEH19, WE19, WHH^{+}11]. showing [LCZ^{+}19]. shows [BHW^{+}12, PSH^{+}11, SM11b]. shredder [MM11]. shrimp [BPL^{+}19a, JW14]. Shutdown [TF11]. Si [CFD^{+}11]. Siberia [PRS^{+}18]. sibling [RF13]. siderophore [BTC^{+}19]. Sierra [SMM11]. signal [ACW^{+}18, MDB19]. signature [HBZ12, Les19]. signatures [CBF10, EMH12, HMFB16, HSC^{+}11, KPJ12, RPMK17, SMG12, SSC^{+}10]. Significance [SH10b, WLS^{+}11]. significant [ASR^{+}17, BPGE13, DVSV13, GRPB^{+}17]. Silica [MRBR10, CF13b, CF14, DMB^{+}12, HSC^{+}11, KBL^{+}10, KNL10, LYL^{+}17, MBTK18, WLR17]. Silicic [DBC^{+}13, KBVW12, MVL^{+}10]. siliciclastic [PMY19a]. silicification [DBC^{+}13]. silico [GRRA^{+}17]. Silicon [CFD^{+}11, GRE^{+}16, LALM16, LALGM18, CFD15, HSC^{+}11, KBL^{+}10, MRBR10, MEM^{+}17, PRS^{+}18, WLR17]. silvery [JGR^{+}14]. Similar [LCW^{+}17b, RF13]. similarities [LSHK11]. similarity [MMD18]. similis [MNW^{+}19]. simple [HBR^{+}14]. simplifies [NB17]. simulated [DBA16, SNK12]. simulation [IH18, LBS17, OBI12]. Simulations [SRA10, FRP^{+}14, GBMG12, LBR^{+}13]. Simultaneous [HSBA10, Joh10, GSBR11, SFWP12]. since [WWC^{+}18]. Single [MDE11, HNL^{+}13, HPS10b, RGLM^{+}12, SDMK10, YWL^{+}17]. Single-cell [MDE11, HPS10b, RGLM^{+}12, SDMK10]. single-stranded [HNL^{+}13]. sink [BSN^{+}14, CCV^{+}18, CAS^{+}17, MSGS^{+}13, MMGO^{+}17b, TSB^{+}19, WLHW13]. Sinking [DVDB16, ALL^{+}10a, DM17, HCLS11, KGT12, LdJMS^{+}13, LKK13, LBNT11, MB10, MLGZ16, SSFF12, SSG^{+}17, SSGM18, TIS^{+}13, TNK^{+}14]. sinkings [MD10]. sinks [PMY19a, PHPH^{+}16]. sinuosa [HCK11]. site [DCRC16]. sites [LKS^{+}16, WC17]. Sitka [BLLB12]. situ [ALL^{+}10a, AFSM17, BFD^{+}11, BMC^{+}16, BMD17, DHW11, FPGR^{+}13, GOD^{+}18, GJWS14, GJWS16, GKS12, HMH^{+}16, KYRMD18, KDGL19, KFJ13, KTH^{+}19, LP10, MMPSB14, MCYR17, NLM^{+}12, SDS^{+}11, SPTS15, TIF^{+}15, WPL^{+}14]. six [BCDR^{+}19]. Size [CBFK19, SPGRP^{+}17, SPR^{+}15, BFW^{+}13, BPW^{+}19, BBQ^{+}10, CL10, CL11, CLWD13, DPLG^{+}19, DOD10, DML17, Edm15, EHW^{+}15, GC16, KHTO13, KBL^{+}10, KTL17, LYH17, LUM15, MCLT12, MCLT15, MRSS12, Meh10, MRE18, MBO^{+}16, NTA14, NLM^{+}12, OR16, PBV16, Piw19, PMRRA19, RSN16, SGCI14, SLU11, SASB^{+}15, SP11, SH10b, SS17, VMC^{+}13, WM12, WDJF12, WSTG18]. size-based [VMC^{+}13]. size-dependent [WSTG18]. size-distribution [SP11]. size-fractionated [LYH17]. size-scaling [KTL17, SLU11]. size-structure [LUM15]. size-structured [WDJF12]. sized [SPG^{+}13]. sizes [DRE^{+}10]. Skagerrak [BBT^{+}10]. skeletal [WTC^{+}17]. skeleton [YLH^{+}16]. slope [BPPF12, FB12, HGM10, JAZ^{+}10, KJKS18, NTK^{+}18, NI10, PCF14]. Slovenia [MMGP^{+}12]. Slow [WGH^{+}16, SJM11]. Small [DKSA19, BK13, CKD^{+}16, DHW11, DB11, HSC^{+}14, MCCA18, MGW^{+}13, MFL11, OIS10, PHJ12, PSB^{+}16, RR13, SNO^{+}16, SBvH^{+}15, SSB^{+}18, SBK18, SRAB10, SRA10, TB18, TSK13, VLDM19, WGH^{+}10, WCCP14, vEG10]. small-scale [SSB^{+}18]. snow [LKK13, RLC^{+}11, vdJFS^{+}18]. snowmelt [CMS17]. Society [ELJ^{+}16]. soft [BMDC10, HATF17, NNE12]. soft-shell [BMDC10]. softer [HATF17]. Soil [PBL^{+}18, AWG^{+}12, KZB^{+}10, YJO^{+}19]. Solar [BPRG^{+}18, SEYJ11, HVJ^{+}19, SSPK^{+}12]. solids [TRA19]. Solimões [BMF^{+}16]. solitary [CNL^{+}15, CRB^{+}17]. Solubility [SDSC12]. soluble [JBT11]. solute [CHW14, HC12, SLK^{+}14]. somatic [MCWB10]. Some [KVA18, HCAF18, APF^{+}18]. Sound [KT13, MSAM18, UMHH^{+}14, BBMS17, BAY^{+}14, BCF^{+}17, GLMG15, GBT^{+}17, VW17]. sounder [RK13]. soundscapes [TLH^{+}11]. Source [FPG11, MMXC15, WLHW13, AHJS15, BSN^{+}14, CAS^{+}17, FBFR13, GCH^{+}18, GRPB^{+}17, HFP10, JSH12, JBT11, KLM^{+}17, LHSG15, PVLMT^{+}16, RASV^{+}17, RBRH10, RHDTS^{+}11, SCR^{+}12, SLH^{+}15, VLMTEW11, WKG^{+}16, WSM^{+}19, WCG^{+}17, ZLLM10, ZXM^{+}11, ZPK^{+}12, dKYH^{+}12]. Source-age [MMXC15]. Source-sink [WLHW13, CAS^{+}17]. Sources [MHA^{+}18, PMY19a, WGCC14, ZZAC13, Ano19c, BTH^{+}16, CPPdAR^{+}13, FYVU17, FEC^{+}16, GMMV19, GBB^{+}18, HCW^{+}10, HCLS11, HSC^{+}14, KSFT13, LRM17, LWWC^{+}16, LGC16, LDL^{+}19, MGGS18, MMGO^{+}17a, OCR10, OLC18, PLS^{+}16, PHPH^{+}16, SKK^{+}15, SMR^{+}17, TLG^{+}11, TTV^{+}13, THFG16, WLG^{+}16, ZZY^{+}10]. South [RMJ^{+}18, BNW^{+}14b, MvdPK^{+}15, WBZ^{+}13, WZG^{+}14, BBTK^{+}16, CCV^{+}18, CRJ^{+}14, CJW^{+}19, DTFR12, DTL^{+}19, DWDH10, DBV^{+}11, GBD^{+}10, HCW^{+}10, HCLS11, JKKM13, MWBM19, NLO^{+}12, PSS^{+}14, RDB^{+}16, VGM14, WDX^{+}11, WMM18, XDC^{+}19, ZCY^{+}15, ZXL^{+}19]. south-Alpine [BNW^{+}14b, WBZ^{+}13]. Southeast [TLB^{+}16, AAIA14a, AAIA14b]. southeastern [SSFF12, STC^{+}11, SRM^{+}18]. Southern [HMV12, MMHT10, BWS^{+}14, BGR14, Car10, DLP13, HZC^{+}13, LKLH10, NSO19, PMLC^{+}10, REDW10, SCAB^{+}16, SWD^{+}14, SBM^{+}15, TBLG14, vEG10, AdBVA10, BAG^{+}14, BPRG^{+}18, CFD^{+}11, CFRL10, CDA16, EB12, FYC^{+}18, FRA^{+}17, HSC^{+}14, JTG^{+}11, KYRMD18, MEM^{+}17, MVT^{+}17, MQJG13, OCLW11, OFGF12, RBCS16, RKTLM18, SNvD^{+}10, SDSC12, SMH^{+}11, SHF^{+}12, WGRS^{+}17, YYMN13, vdHHC^{+}19]. Southward [Ano17l, ZXZ17b]. southwest [LWE^{+}11, LG16, SNO^{+}16]. Southwestern [IR16, ERA^{+}12]. sp [CHS^{+}18, KMF10, QFH18, RG13, RJFMG17, RWM^{+}14, TSK13]. space [ITO^{+}17]. sparse [CJHR19]. Spartina [CZB^{+}18, TMH^{+}18]. Spatial [ASH^{+}14, BL13, BHB^{+}12, FA10, GNHGM13, GBT^{+}17, GFDC11, HL13, HOD^{+}17, MBBW11, MBO^{+}16, PE16a, PRS^{+}18, RS16, RMH^{+}17, RKMN^{+}13, VBC^{+}12, WLS^{+}11, WRWPG19, AFSM17, BRM^{+}19, BRT^{+}10, CLWD13, DOD10, DTKMK15, EED10, FZL^{+}14, FMP^{+}13, GRSD^{+}14, GAK^{+}19, HS10, JHW^{+}19, KT13, MG14, MACM11, MMWR17, NAH^{+}11, OWS^{+}17, PST^{+}13, RPMK17, RNT^{+}19, RAV^{+}17, SKKV11, SPO^{+}18, TLH^{+}11, VPMrI12, VMMS^{+}13, WWC^{+}18, ZYZ19, dlFN10, Ano21c]. spatially [BMN16, SVS^{+}19, Tho19]. Spatio [CKB^{+}16, NSG^{+}16, SKGT17]. Spatio-temporal [CKB^{+}16, NSG^{+}16, SKGT17]. Spatiotemporal [JZZY18, KHP18, Man10, OY10, PCD^{+}19, SPP10, TCG^{+}17, JAD^{+}13, MSAM18]. spatiotemporally [FLM^{+}19]. spawner [BMC^{+}16]. spawning [BMC^{+}16, FLM^{+}19, SVS^{+}19]. specialist [TMK^{+}13]. speciation [AHD^{+}18, BAG^{+}14, JKKM13, LG16, VGM14]. Species [CCK^{+}12, BCDR^{+}19, BYD19, BDS11, BBCM^{+}13, CTA^{+}19, CUW11, ETKL16, GYP^{+}18, GBD^{+}10, GN16, HHW^{+}19, HMD11, HVJ^{+}19, HS10, HSR15, ITO^{+}17, KSP^{+}12, LABJ18, LTPA17, MNW^{+}19, MMB17, PZHD18, RBG^{+}10, RF13, RVvdP^{+}17, SRCL^{+}13, SMF10, SMA13, SGME11, SSFR19, TDS^{+}10, Tho19, TGGZS^{+}10, TBSR13, WZC13, YKBJL12]. Species-dependent [CCK^{+}12]. species-specific [SRCL^{+}13]. specific [BBCM^{+}13, BCM^{+}17, CL11, DNH^{+}18, EBMR12, Fie13, HOD^{+}17, MBTK18, NLM^{+}12, PBA^{+}15, PMP^{+}12, Piw19, RDT^{+}14, SRCL^{+}13, TKB18, WRH^{+}18]. spectra [SW14]. Spectral [GRGL^{+}13, NBG17, PE13, FB12, HS11, RM14, RNT^{+}19, RDT^{+}14, ZD18]. spectrometry [SSC^{+}10]. spectroscopy [AC15, FHS10, SKK^{+}15]. spectrum [BKD^{+}16]. speed [LdJMS^{+}13]. speeds [FDP^{+}18, MD10]. spinuligerum [vHOM^{+}19]. spiny [BBS12]. spiralis [ZLLM10]. Spitsbergen [KvdPB18]. splash [MBBW11]. sponge [ASR^{+}17, KYC^{+}15, LKF^{+}18, MBLP11, MJH^{+}16]. Sponges [KSFT13, FT11, HGT^{+}18, HTLM18, LALM16, LALGM18, MRBR10, MCYR17, SWM^{+}10, WMP^{+}19]. sporadic [KMC^{+}15]. spot [WMBR13]. spots [GGL^{+}15, SFLB16]. spp [Edm11, HKS^{+}15, IWF19, KHPIP^{+}14, MQP^{+}16, RG13, SGCC16, TGG^{+}11]. sprat [HPS^{+}10a, NZH^{+}11]. Spreading [SPB^{+}14]. Spring [KTK^{+}13, AC15, BJ15, CR16, GLMG15, GGTC^{+}18, HC10, HCF^{+}10, HC12, HCC^{+}13, HKS^{+}15, IHSS^{+}19, JZZY18, KYRMD18, KIH^{+}15, MMD15, SS16, SPSG14, SNvD^{+}10, SLA^{+}18, ŠNZ^{+}14, SLG^{+}14, SFLQ^{+}19, TF11, VMCM^{+}17, ZXN^{+}12]. spring-fed [AC15, HC10, HCF^{+}10, HC12]. springs [BR17, KGvdH16]. St [GdVT^{+}11, MPM^{+}15, vdHHC^{+}19]. St. [BPW^{+}19, FLM^{+}19, HT17a]. stability [ABS^{+}19, DBMP^{+}11, MGJH18]. stabilize [DML17]. Stable [BSCG17, BWBB15, BGB^{+}14, CCC10, GMMV19, GCH^{+}18, RHV^{+}13, VHR^{+}11, AHD^{+}18, BJDMH10, BTH^{+}16, CPPdAR^{+}13, CBF10, EED10, FC11, GLS^{+}13, HPCD13, HHM^{+}18, JBT11, JSB^{+}14, KBA^{+}12, KBA^{+}14, KGL^{+}16, KLM^{+}17, KWB^{+}16, LRM17, MTEM15, MBLD15, SBvH^{+}15, SMG12, SSYT14, VTH^{+}18, WFK^{+}16, WGCC14, WLHW13, ZMWM11, ZHD^{+}16]. stable-isotope [CBF10]. stable-isotope-addition [EED10]. stage [BBCM^{+}13, FDP^{+}18]. stages [AACS11, RR12]. staining [FAF^{+}12]. standardized [SJB^{+}19]. standing [KKH11, KOFN11, LSDW18, MRBR10]. starvation [WRH^{+}18]. state [BHS^{+}16, BLS^{+}16, FBFR13, FMM^{+}14, GBL13, NEH^{+}19, RCIB14, SZH^{+}10, Spi15, SRA10, ZZY^{+}10]. States [BHC13, JCF^{+}10, LGC13a, LGC13b, MA18, ZHD^{+}16, BHC14, BGB^{+}14, MRSS12, MLS^{+}14, SDH^{+}14, WWC^{+}13]. station [AAIA14a, AAIA14b, BDK^{+}17, DBH^{+}16, GWB^{+}14, MGK15, MG17]. statistical [HSBA10]. status [CR11, JHW^{+}19, PS13, SJB^{+}19, SvKP^{+}18, Tad10]. Steady [GBL13]. Steady-state [GBL13]. steelhead [CBP10]. steep [BBLN11, NSO19, SSU^{+}16]. step [KM10]. steps [GRDPL14]. sterol [CWF11]. sterols [RASD10]. stickleback [KKHP14]. stimulate [REE^{+}12]. stimulated [TTTM^{+}19]. stimulation [SSGL19]. stipulacea [CvHB^{+}18, SLS^{+}11]. stochastic [SRA10]. stock [SAS^{+}11]. Stockholm [TBK15]. stocks [BBS^{+}18, MRBR10, PHLSSS19, TTTM^{+}19]. Stoichiometric [GHS14, SMC^{+}10, BMPF19, BISZ17, HSB^{+}13]. stoichiometry [BMW10, BK11, CJ17, FWWF18, HBBM19, HESU13, JSH12, KBHT19, MRKR^{+}14, MVNG11, MEM^{+}17, PFH^{+}17, PWF18, RBCS16, SD10, SWP11, THA17, WZBW^{+}11, YJO^{+}19]. Stokes [MD10]. stony [MBHG11, MPSA17]. storage [CMS^{+}18, GHS14, HCAF18, JWS15, MKB^{+}19]. storm [GPA^{+}14, GGL^{+}18, SVMT15, WLL^{+}11]. storms [FSBT16, SLG10]. Strain [PBA^{+}15, DNH^{+}18, FRA^{+}17, HS18]. Strain-specific [PBA^{+}15]. strains [SBF18]. Strait [GRT^{+}14, JMM14, MVT^{+}17]. Straits [HCS11]. stranded [HNL^{+}13]. strategies [Ano19c, GBB^{+}18, LLL10, MAC^{+}10, SMH^{+}11, WZR19]. strategy [BFW^{+}13, CMG^{+}15, PVA^{+}19]. Stratification [MRSE14, SBBNM14, ASL16, ABS^{+}19, BCRW15, CR10, CSD10, IGP^{+}12, KGT12, LBC^{+}18, MvdPK^{+}15, PRL18, RAKE05, RMNZ12, RVvdP^{+}17, SNO^{+}16, SLPM15, VLDM19, VCM13, WP14]. stratification-induced [IGP^{+}12]. stratified [Ano21a, BHB^{+}19, BSN^{+}14, CR10, CFW^{+}14, FDL17, GSG^{+}17, HHM^{+}18, HD19, KKB^{+}18, KCM^{+}10, KKS10, LBS17, MMGP^{+}12, OSC14, QWRJ10, RAKE05, SI10, SdlFdlF^{+}10, SPO^{+}18, SCBR12, SSB^{+}16, VPMrI12]. stratifying [APF^{+}18, KVA18]. Stratigraphic [SLK^{+}14]. Stream [KB15, Ano19c, BMBI12, BDU^{+}19, CRCGG^{+}17, CLN^{+}19, DRE^{+}10, Dem19, DVSV13, FUS^{+}16, FHR^{+}15, GTR^{+}13, GSB^{+}17, GBB^{+}18, HHS^{+}18, JC14, KRB^{+}18, LSHK11, LBR^{+}12, MACM11, MBP^{+}17, OVRJ13, PH13, PGP^{+}14, SCAB^{+}16, SSU^{+}16, SGRB10, SC10, TBAS14, TBSL17, TBF^{+}13, WWS11, WYW^{+}10]. streamflow [DBA16]. streamlined [HS18]. streams [BLJ13, BLMS17, CFAE^{+}15, CFD^{+}19, HEB^{+}19, HEH^{+}17, HAA^{+}19, HH14, JBLJ12, JTH^{+}13, LHSG15, LH17, LDL^{+}19, MTT17, PCO^{+}15, PJUR15, RWM^{+}19, RvSM17, REDW10, SBM16, SWP11]. streamwater [CK12, CK13]. strength [BCVAn10, SSM^{+}19, WHH^{+}11, WDH^{+}17]. stress [CRS^{+}17, FWO^{+}18, GvBBB17, GHSR^{+}16, LABJ18, Les19, RLSC^{+}13, RKMN^{+}13, SHKU11, SCPE15, SMC^{+}10, TGGZS^{+}10, WHD10]. stressors [Les16, MMBP18]. stromatolites [RPB17]. Strong [AGCA16, LBR^{+}13]. strongly [BHB^{+}19, BG10a, ILPL13, NLHAA^{+}17, VBC^{+}12]. Strontium [MAC^{+}10]. Structural [YLJ11]. Structure [CFVU11, AA11, BSG14, BAG^{+}17, BRS18, BBQ^{+}10, CPHD15, CVS^{+}10, DDF^{+}10, FMM^{+}14, GRSD^{+}14, HMV^{+}18, HHHT19, HLJ12, HOD^{+}17, JPH^{+}18, KCH^{+}12, KT13, KHH19, KMP^{+}11, KZR^{+}16, KBL^{+}10, LBC^{+}18, LSH^{+}17, LJL^{+}18, LUM15, LDT^{+}11, MCLT12, MCLT15, MvdPK^{+}15, MDE11, MRE18, NB17, PMP^{+}12, PHCD14, Piw19, PFJ10, RBCS16, RGO^{+}11, RRGGCA19, SCF^{+}15, SWCL12, SFI^{+}18, TA14, VPMrI12, VMCM^{+}17, WRB^{+}19, WWS11, XFLM14, ZWL^{+}14]. structured [LGC13a, LGC13b, WDJF12]. structures [BBR^{+}14, CWRX19, GBB^{+}19a]. structuring [CPOMA15, WXMS10]. studied [Clo19, MBB^{+}18]. studies [APS^{+}19, BLWV10, IH18, KYG^{+}12, PCW19, RHV^{+}13, RGM^{+}11, SRAB10, WP14, WLG^{+}16, Xen19, ZTW^{+}11]. Study [YAC^{+}19, AFG^{+}16, BC19, BBTK^{+}16, BBR^{+}14, BAY^{+}14, CSJ^{+}14, CJS^{+}17, CFD^{+}19, EMH12, EP14, EOM16, GYP^{+}18, GBL13, HMH^{+}16, IGP^{+}12, LBS17, MBC^{+}18, MGS12, MU17, MCYR17, OEM12, PDER10, PHL^{+}18, PLE^{+}17, RASD10, RF13, RMNZ12, RAV^{+}17, SDS^{+}11, Scu16, SW14, VTH^{+}18, WWC^{+}13, WXF^{+}15, WVGB10, WRS13]. Stylophora [HRG^{+}15, SIW^{+}11]. sub [GDCM13, GSB^{+}17, KT13, ORGE16, PMA18, RNT^{+}19, RAB^{+}17, VML^{+}19]. sub-alpine [GDCM13, GSB^{+}17]. sub-Antarctic [PMA18, VML^{+}19]. sub-Arctic [ORGE16]. sub-catchment [RAB^{+}17]. sub-estuary [KT13, RNT^{+}19]. subalpine [EKS^{+}18, EWB12, HEB^{+}19, VMI13]. subannual [HMFB16]. subarctic [ATP^{+}15, DMMV15, FMM^{+}14, HEBS10, MLD^{+}16, MLL^{+}14, MGS12, NO17, PNR19, PFvO^{+}18, RMF11, RLL^{+}10, RHSD^{+}10, SKJD^{+}14, STB^{+}16, UFW^{+}18]. subduction [SSGM18]. subfossil [CSGW18]. subject [GLF17]. subjected [BBS^{+}18]. sublittoral [GCR^{+}10]. Submarine [KKH11, RDP^{+}17, BOT^{+}15, GSZL13, KDGL19, KCL^{+}14, KSG^{+}10, LKS^{+}16, LKLH10, LSH^{+}17, LCH^{+}14, LSD18, OBL^{+}19, PVLMT^{+}16, RGM15, SS12b, SS12c, VLMTEW11]. submerged [NBG17, VP15b, WZTK15, ZLLM10]. submergence [MBBW11]. submerging [SKGT17]. submersed [GK10, GK14]. submesoscale [MS13]. submicron [JYS18]. Subpolar [JWGH19, FPP^{+}19, MMD15]. subpopulations [ALdML^{+}14]. subsea [BOT^{+}15]. subsequent [DMN15]. subsidence [KMC^{+}15]. subsidies [ALG^{+}13, BLWV10, MSM^{+}17, dKNL^{+}15]. subsidizes [HMH^{+}16]. subsidy [MDF^{+}14]. substances [MMN^{+}10, RJFMG17, TMK^{+}13]. substantial [MNW^{+}19, SSFF12]. Substrate [BKD^{+}16, AFG^{+}16, MKBSK19, XLS^{+}19]. substrates [LTH^{+}12]. Subsurface [SBM^{+}15, SVS^{+}19, ZXL^{+}19]. subterranean [PMY^{+}19b, PMY19a]. subtidal [EOM16, MHL^{+}16, OEMB10, OEM12, SMF10]. Subtropical [CPHD15, DDK10, DBV^{+}11, HPCD13, HDP15, LWB^{+}17, ARML10, ASSG12, BCRC16, BSB^{+}10, CCW^{+}19, DBH^{+}16, DKSA19, EOM16, HC10, HCF^{+}10, HEBS10, JMJ^{+}19, KBVW12, LCZ^{+}19, OEMB10, OEM12, OY10, PZHD18, PD11, RQC^{+}15, RHMSE15, SOM17, SHD^{+}11, UFW^{+}18, WE19, YH17, ZWL^{+}14]. success [BSBK13, LRY12, LFH^{+}12, PWWF18, SNTK15]. succession [FGMN17, GYP^{+}18, HV19, KVMA17, MAV^{+}13, MTM^{+}16, SPB^{+}14, ZXM^{+}11]. successions [TB18]. sudden [BLS^{+}16, OLF^{+}11]. sufficient [HS10]. suffocated [BDU^{+}19]. suggest [GLS^{+}13, SPTS15]. suggests [PTS12]. Sul [PMP^{+}17]. sulfate [JP10, ZMS^{+}18]. Sulfide [VSD10, ARW^{+}10, FDL17, GFT^{+}14, KWM^{+}19, WBZ^{+}13, ZYZ19]. sulfide-dependent [WBZ^{+}13]. sulfoniopropionate [ARW^{+}10]. Sulfur [FC11, DJD^{+}14, GRPB^{+}17, JZZY18, MBH^{+}15, NUH^{+}12, TNMV^{+}10, YLH^{+}16]. Summer [FPPA^{+}11, WRH^{+}17, AA11, CWRX19, HVJ^{+}19, KBVW12, OBT^{+}11, PSG^{+}16, PCY^{+}10, RBCS16, RMNZ12, RVvdP^{+}17, SNvD^{+}10, SLA^{+}18, SK19, VCM13, ZCY^{+}15]. summers [BWS10]. summertime [GMBL16]. sun [DHG^{+}17]. sunlight [GRGL^{+}13]. sunscreens [HKS^{+}15]. Superior [AMB^{+}11, AA11, Aus13, BVC^{+}14, BS18b, KYG^{+}12, KWB^{+}16, LCM^{+}12, LK14, LZK18, TA14, VLWV14, YAC^{+}19, ZMWM11]. superoxide [HBD^{+}16, RGG^{+}10, SCG^{+}19]. supersaturation [SWM^{+}18]. supplied [DBC^{+}13]. supplies [BBS^{+}18]. supply [CFD^{+}11, GSZL13, KCL^{+}14, MCLT15, MSM^{+}17, MZB^{+}15, THA17, UFW^{+}18, VMCM^{+}17, WDMF13, ZSZ12]. support [ANP^{+}14, BBSK18, CS12, DTM18, KBA^{+}12, KBA^{+}14, LJL^{+}18, WCG^{+}17]. supported [Bre10, DKSA19]. supporting [BCC^{+}12, GFDC11, MMGO^{+}17a, vOSH12]. supports [BBTK^{+}16, SSS^{+}16]. suppresses [JSFC18, LRY12]. Suppression [MTH^{+}11, SSGL19]. Surf [MSM^{+}17, FRP^{+}14]. Surface [BDB^{+}14, BHS^{+}16, BMW10, BHB^{+}12, BBB^{+}14, CEES14, CL10, CFW^{+}14, FCD12, HLGA17, HCW^{+}10, HRN11, LWE^{+}11, ML19, NFRU11, OIS10, OLF^{+}11, PHJ12, PH15, PHL^{+}18, RS19, RETS16, RGGL^{+}12, RGLM^{+}12, SGH12, SBM^{+}15, SBFB17, TAV^{+}10, TPM^{+}14, TFLS14, UIY^{+}11, VPC10, XBR^{+}18, YHS^{+}17, ZMS^{+}18, ZXL^{+}19, vHOM^{+}19]. surfaces [YWL^{+}17]. surge [GGL^{+}18, MBBW11, SVMT15]. surpass [MRBR10]. surrounded [CKB^{+}16, SML^{+}19]. surrounding [BBJ^{+}19, MHH^{+}17]. survey [BMN16, HEH^{+}17, OWS^{+}17]. Survival [CGL^{+}16, HPS^{+}10a, IGP^{+}12, MKB^{+}19, All10b, BPL^{+}19a, LdlSB^{+}12, VvO11]. Susceptibility [WZR19]. Suspended [GPS15, ALL^{+}10a, BVvB^{+}19, CT18b, GCH^{+}18, HPCD13, HMFF10, TRA19]. suspension [USB^{+}10]. sustainability [PE16a]. sustaining [VSdG17]. Svalbard [HSP^{+}16, DHG^{+}17, GLKK10, HDDH^{+}17, KPW^{+}11]. swarms [TT14]. Sweden [JBLJ12, vEG10]. Swedish [Bre14, SLP^{+}14]. swell [MP17]. swim [KK11]. Swimming [FDP^{+}18, LWE^{+}19]. switch [WE19]. Switzerland [BNW^{+}14b, RLB^{+}10]. SwmA [SBFB17]. sydowii [RBRH10]. Symbiodiniaceae [Les19]. Symbiodinium [BWD^{+}11, BWD^{+}12, KHPIP^{+}14, WHD10]. symbiont [JLRK12, PGRR^{+}19, UA10]. symbiont-bearing [JLRK12, UA10]. symbionts [HBD^{+}11, HRPW15, LKF^{+}18, TFH17]. symbioses [MWS10]. symbiosis [DBMP^{+}11]. symbiotic [FPPA^{+}11, FPGR^{+}13, FT11, GRPB^{+}17]. sympatric [BPW^{+}19]. symphony [ŠNZ^{+}14]. Synchronized [BPPF12]. Synchronous [WCJ^{+}15]. synchrony [ASW^{+}19, BRM^{+}19, BCM^{+}17]. Synechococcus [ANP^{+}14, BWP^{+}10, DNH^{+}18, DMN15, JBPM15, MZB^{+}15, SBFB17, TNMV^{+}10]. Synergistic [ARML10, ŠSP17]. synoptic [HEH^{+}17, SS17]. synoptic-scale [SS17]. synthase [BMM^{+}13]. synthesis [ASA^{+}18, ETKL16, HBR13, LWWE^{+}18]. system [AHS11, BSCC15, BPPF12, DBSP^{+}16, DJD^{+}14, DPG^{+}12, DTM18, EO13, EED10, FZL^{+}14, HDK^{+}12, KSG^{+}10, Man10, MFM^{+}12, MBO^{+}16, MGSM10, MGT15, NSO19, PJFJ^{+}15, RSG11, RMK^{+}16, STC^{+}11, SFB12, WM12, WWC^{+}13, WGM16, WAB^{+}17, WDL^{+}17, YH17, BTC^{+}19, TSC^{+}19]. systems [AP12, FDBW16, GMMV19, HAC^{+}11, HCH^{+}19, JHD^{+}11, PE17, SGA^{+}17, SRM^{+}18, VPC10, VLJ^{+}10, WSM^{+}19, WGH^{+}10].

T. [CTA^{+}19, ETI^{+}16]. tactics [GMD11]. Tahoe [PHL^{+}18]. Taihu [CCK^{+}12, MQP^{+}16, XPQ^{+}10, dKYH^{+}12, QHVM18, TGC^{+}10, TCG^{+}17, XXZ^{+}19]. Taiwan [JHD^{+}11, YWY^{+}15]. tale [CRCGG^{+}17]. tallgrass [WWS11]. tamarense [SFWP12]. tanakai [ITO^{+}17]. Tanganyika [CMK^{+}10, VAH11]. Tanzania [WKJS^{+}14]. taurine [CHV^{+}17]. taxa [BVP^{+}15, RCIB14, SK19, ŠGN^{+}19, WMT^{+}12, XFH14]. taxifolia [EMO^{+}11, OBM^{+}11, RSTS^{+}18]. Taxon [BBCM^{+}13, MBTK18]. Taxon- [BBCM^{+}13]. Taxon-specific [MBTK18]. Taxonomic [HSTK15, ETKL12, ETKL15, RKG^{+}11, WKAM^{+}19, ZCL^{+}19]. taxonomy [DPLG^{+}19]. technique [HBM11]. techniques [KB15, SMMF19]. telemetry [VdSLC^{+}16]. Temora [SNTK15]. temperate [BWBB15, BDP^{+}19, BK13, BKA^{+}14, CF13b, CRB^{+}17, CGP^{+}19, CHL10, DPSW16, DBMP^{+}11, FPPA^{+}11, FPGR^{+}13, GLMG15, GJWS14, GJWS16, HS11, HNL^{+}13, HAA^{+}19, JM16, JC14, KOFN11, LACI10, LLB17, MA18, MGL^{+}16, MEM^{+}17, MRB11, MDE11, PBL^{+}18, PHLSSS19, RR13, RRGGCA19, SPS19, SBvH^{+}15, SWZ^{+}15, SWL11, SLPM15, VBC^{+}12, VSP^{+}11, WLV17, WS13, ZCK^{+}16]. Temperature [KTL17, LUM15, MCLT12, RKBA14, SLU11, SHKU11, SNK12, SLG^{+}14, SW11, WZBW^{+}11, BHB^{+}19, BYD19, BPGE13, CSGW18, CL10, CL17, DBSP^{+}16, ETKL16, ETI^{+}16, Fie13, FDH^{+}14, GdVT^{+}11, GLMG15, GLF17, GBR14, HYK^{+}15, HSLH^{+}14, HBB^{+}11, HQB^{+}18, JHLK^{+}19, JC14, KNA^{+}14, KSY11, LABJ18, LAM12, MKK15, MSK^{+}17, MMJ^{+}12, NFRU11, PvEF12, PMP^{+}12, PSG^{+}16, PHL^{+}18, PWF18, RLB^{+}10, Rie15, SJB^{+}19, SDS^{+}11, SPHVA19, SIW^{+}11, SPPS10, SH11, TJJ^{+}15, THA17, TPM^{+}14, VLWV14, WLO^{+}19, WHD10, XXZ^{+}19, YWL^{+}17, ZMS^{+}18]. Temperature- [KTL17, SW11]. Temperature-induced [SHKU11, THA17]. temperatures [BBQ^{+}10, RMNZ12, SNO^{+}16, SMF10, WdBJF16, WRH^{+}17]. Temporal [AFSM17, BCRC16, BDK^{+}17, CFAE^{+}15, CDA16, CBF10, EED10, GFT^{+}14, PRL18, TvBR^{+}19, TEZ^{+}18, VMMS^{+}13, WWC^{+}18, ZYZ19, dlFN10, BRM^{+}19, CKB^{+}16, DB11, GBT^{+}17, HLH13, IBPG17, JD16, MHRH11, MTM^{+}16, MBO^{+}16, NSG^{+}16, PE16a, RS16, RPMK17, RMH^{+}17, RAV^{+}17, RKMN^{+}13, SKGT17, VPMrI12, WVL^{+}18]. Temporal-spatial [WWC^{+}18]. temporary [MM11, TBAS14, WVGB10]. tentative [WLR17]. term [AAIA14a, AAIA14b, APS^{+}19, ADCH18, BHW^{+}12, BGW^{+}15, BvBB^{+}16, CJS^{+}17, CESC14, DC15, EP14, ETI^{+}16, GPH^{+}13, HSCM19, JHD^{+}11, KMC^{+}15, KHK^{+}19, LC12, MKBSK19, MKG^{+}15, MSR16, NPT11, OEMB10, PCW19, PJUR15, RWM^{+}19, RKWH18, RGO^{+}11, RG13, RPG13, RNT^{+}19, RMNZ12, SK19, SSFR19, SSS^{+}19, TNI19, TCFP19, VKC18, VvO11, WCM19, WVV^{+}11, WB19, Xen19, YMB^{+}18, ZHN^{+}10, ZWL^{+}14, MKBSK19]. terminates [ŚF19]. termination [BRF^{+}17]. terms [KPV^{+}11]. Terrestrial [AWG^{+}12, BBSK18, CS12, HMH^{+}16, KBA^{+}12, KHVS11, AAC^{+}19, BLWV10, TYX^{+}19, WLS^{+}11, KBA^{+}14]. terrestrially [SBC^{+}17]. terrestrially-derived [SBC^{+}17]. terrigenous [FB12, MBAS^{+}17]. test [LLL10, LAM12, MMFBB18]. Testing [SMA13, Lan14]. testudinum [BBS^{+}18, BJF18, HE10, SCPE15]. tetraether [ZKMT^{+}13]. Th [SBNC^{+}19].

Thai [TLB^{+}16]. Thai-Malay [TLB^{+}16]. Thalassia [BBS^{+}18, BJF18, HE10, SCPE15]. Thalassiosira [BRS^{+}13, FAF^{+}12, HBB^{+}11, MEM^{+}17, SLC^{+}16, Sch19, SLH^{+}15, TJJ^{+}15]. Thaumarchaeal [SSG^{+}17, HQB^{+}18]. Thaumarchaeota [PWS^{+}11]. thaw [DMMV15, LVM^{+}10]. Their [SBKO18, TIS^{+}13, BBMS17, BBM11, CBF11, CFRL10, DJD^{+}14, FT11, GK15, GAK^{+}19, LOS12, LTH^{+}12, LALGM18, MDB19, MSAM18, MSR16, NCT^{+}15, PZHD18, PE17, RASD10, RLC^{+}11, RMJ^{+}18, RMK^{+}16, RR12, SSG^{+}17, SRCL^{+}13, SGA^{+}17, SMN^{+}15, SKK^{+}13, SSN12, Tad10, VBBR17, uGH^{+}11]. theoretical [VTH^{+}18, ZF17]. theories [APF^{+}18, KVA18]. theory [KTL17, SMA13, WLO^{+}19, WFB^{+}11]. there [CL17]. Thermal [CUW11, JLRK12, PZHD18, SASB^{+}15, SNO^{+}16, XFLM14, ZWL^{+}14, AA11, BCRW15, CRS^{+}17, FZL^{+}14, GSBR11, GHSR^{+}16, LBC^{+}18, Les19, LS15, LCZ^{+}19, PMP^{+}12, PST^{+}13, PRL18, RDC^{+}19, RGO^{+}11, SFS^{+}16, SBB^{+}18, TA14, ZKL^{+}14]. thermally [BSN^{+}14, CR10, KKS10, RAKE05]. thermocline [PPL10, SGJB14]. Thermodynamics [RMK^{+}16]. thermography [KDGL19]. Thermohaline [ASL16]. thermohydrodynamics [dlFN10]. thermokarst [MLD^{+}16]. Thiamin [FLLH18]. thin [SBBNM14, TWWY18]. things [MTT17]. thiols [LFC17]. thiotrophic [LFB^{+}10]. third [XXZ^{+}19]. thompsoni [HP19]. thorium [CSJ^{+}14]. those [MRBR10]. threat [JTH^{+}11]. threatened [LABJ18]. Three [MMFBB18, OBI12, EMS16, HE10, HH14, IH11, LWE^{+}19, LWWC^{+}16, OSC14, RF13, SPP10, TGGZS^{+}10, WE19, WOC^{+}18, RBY^{+}17]. Three-dimensional [MMFBB18, OBI12, HE10, LWE^{+}19]. threespine [KKHP14]. Threshold [SMA13, Bre10]. Threshold-driven [SMA13]. Thresholds [DdG10, SW11]. through-flow [OHKC^{+}12]. throughout [EB12, HPM^{+}10, MQP^{+}16, SHM^{+}19]. Thysanoessa [BPW^{+}19, CTA^{+}19]. Tibetan [MNW^{+}19, SHL^{+}18]. Tidal [KGM14, RBD18, VLMTEW11, WMC^{+}18, ADCH18, AC17, BFD^{+}11, BMD17, BGP^{+}15, CEES14, CF14, CMW^{+}19, DCCB17, DTM18, FEW^{+}14, HPM^{+}10, HMFF12, LSH^{+}17, LHSBP18, MSGS^{+}13, MCT^{+}14, PE16a, RGM15, SML^{+}19, SHM^{+}19, SSP^{+}18, SVMT15, WGC^{+}13]. tidally [GJR^{+}19, VPG^{+}19]. tide [CTG15, CT18b, GLF17, GLF18, HCD19, HST^{+}14, KG18, KSWFG13, LWS^{+}17, OPA^{+}14, OLF^{+}11, WXF^{+}15, ZXM^{+}11]. tide-dominated [GLF17, GLF18]. tide-driven [HCD19]. tides [LKLH10, LLW^{+}18, VMCM^{+}17]. tiered [OMSC13]. tightly [VCPC^{+}16]. time [BGB^{+}14, CHHT18, CF14, DdD^{+}10, FSBT16, GBS17, GK14, IBPG17, Joh10, KH16, MPM^{+}15, SHSK14, SLHA19, Sha10, SSPK^{+}12, WRS13]. time-dependent [SSPK^{+}12]. time-frequency [SLHA19]. time-series [CHHT18, FSBT16, KH16, WRS13]. times [BM16]. timescale [ASW^{+}19, BSB^{+}18]. timescales [LH17, SM10, SHK13]. Timing [BMC^{+}16, LHS19, MDB19]. tip [ZCZ^{+}18]. tipping [CESC13]. tissue [BLJ13, JLRK12]. Tohoku [KJKS18]. Tokyo [KHK^{+}19, TNI19]. tolerance [CUW11, IOB^{+}11, LGW^{+}19, WA14, WHD10]. tolerances [PMP^{+}12]. tolerates [VFS^{+}15]. Tonga [SPB^{+}14]. tonsa [DHK11, JLG10, JLG11, TW10b]. tool [BGB^{+}14]. tools [KWF^{+}17, MH16]. top [Meh10, PDER10]. top-down [Meh10, PDER10]. Topographic [AAO^{+}19]. Total [YKT^{+}15, FSCB11, Kir13, TRA19]. Tower [CSC^{+}11]. toxic [BH16, BRF^{+}17, FWvD^{+}18, GNWDL19, JLC^{+}15, LGW^{+}19, XNK18]. toxicity [DBFL11, HST^{+}14, HLSW^{+}15, JHLK^{+}19, WZR19]. toxicokinetics [HHW^{+}19]. toxin [DMS^{+}18, MMHT10, SBDS^{+}15]. toxins [BMDC10]. Toxoplasma [SSL^{+}12]. Trace [BLLB12, HCW^{+}10, HCLS11, LYH17, ANP^{+}14, CJ17, MMH^{+}18, ORC^{+}17, SH10b, TNK^{+}14, WFR10]. traced [LKLH10]. tracer [DTM18, EWB12, FB12, GVS^{+}10, JTH^{+}13, RF13]. tracers [BTH^{+}16, GMMV19, RASD10, TLG^{+}11, TMO^{+}18]. Tracing [MMPSB14]. tracked [CFD^{+}11]. Tracking [CLB19, KGM14, LHSG15]. tracks [PWS^{+}11]. trails [LKK13]. Trait [FFA13, FPP^{+}19, TBLG14, KVMA17, KWF^{+}17, SPR^{+}15, WGH^{+}16, WKAM^{+}19]. Trait-based [FFA13, FPP^{+}19, TBLG14, KWF^{+}17]. traits [ALdML^{+}14, BCDR^{+}19, BH16, BSM17, CPOMA15, ETKL12, FBL15, GBB^{+}19a, HV19, KVMA17, PLS^{+}16, SYdTP^{+}11, SvKP^{+}18]. trajectories [RCJ15]. transboundary [TZD^{+}15]. transcription [DMB^{+}12]. transcriptional [SDCF16]. transcriptome [KBHT19]. Transcriptomic [LGW^{+}19]. transect [HHS^{+}18, MLS^{+}18, MCGF^{+}11, ZTW^{+}11, dGCB^{+}11]. transfer [ACW^{+}18, GBT^{+}17, HE10, KWB^{+}16, MDSG18, RN14, RMH^{+}17, SSU^{+}16, SAS^{+}11, VPC10]. Transformation [OEM12, RJFMG17, EKS^{+}18, GAM^{+}19, MKW^{+}19, ORGE16, PML^{+}19, RBY^{+}17]. Transformations [CAS^{+}17, KPJ12, KHG^{+}13, OHKC^{+}12, RGB^{+}19]. transformed [CJC^{+}12, LGR^{+}12]. transient [JWS15]. transition [MD15, RRD14]. transitional [KMH^{+}17]. transitions [BVSR^{+}15, PT11]. translocation [PGP^{+}14]. transmission [HNL^{+}13]. transparency [FOT^{+}15, SNO^{+}16, WFB^{+}11]. transparent [AAC^{+}19]. Transport [FYC^{+}18, FWFB10, BBK^{+}15, BBCM^{+}13, BDC^{+}14, BMB^{+}18, CHW14, CLB19, EWB12, FRP^{+}14, GPS15, HSR^{+}10, HC12, HCS11, IR16, KYR^{+}12, KMC^{+}15, KBE^{+}17, LKS^{+}16, MMFBB18, PHPH^{+}16, RBY^{+}17, RCV^{+}14, RKTLM18, SFMF15, SHT^{+}17, SGG^{+}11, SBNC^{+}19, SKK^{+}13, SC10, TBK15, WKSR13, WBG^{+}16, WLL^{+}11]. transported [BBM11]. transregional [WWC^{+}13]. trap [ALL^{+}10a, EMH12]. treated [KCB^{+}17, LM12]. treatment [MACM11]. tree [KKP^{+}19]. Trench [KJKS18]. Trend [RMNZ12, JHW^{+}19]. trends [DC15, HLH13, LKT17, MKG^{+}15, MSR16, PHDH14, TKB18]. tributaries [PFH^{+}17]. Trichodesmium [BS18a, BAA^{+}13, BWB^{+}10, BRS^{+}13, GWB^{+}14, HBD^{+}16, Ho13, KWGN^{+}10, OALD10, OBNP^{+}10, RWM^{+}14, WKB^{+}10]. tricornutum [CSJ^{+}14, RLSC^{+}13]. trigger [FMGR^{+}11, FPSL18, GMD11]. triggered [BS18b, DMSHC16, WLL^{+}11]. triggers [BMC^{+}16, WHAM15]. triphosphate [MLK11]. triple [MQJG13, TMO^{+}18, WSB^{+}13]. triplet [MA18]. Trophic [ACA^{+}11, CTA^{+}19, CHL^{+}17, CPHD15, HHW^{+}19, KLM^{+}17, MTEM15, MCYR17, SLBNG11, BBT^{+}10, BHS^{+}16, BFW^{+}13, BSSR10, CR11, CFRL10, DLP13, DLBF17, DC15, FPSL18, GAH11, GBT^{+}17, GRDPL14, JSB^{+}14, JC14, KCH^{+}12, KVMA17, KWRS13, LS14, PGRR^{+}19, PS13, PLE^{+}17, PMA18, RHV^{+}13, RCIB14, RRCH^{+}19, RRGGCA19, SJB^{+}19, SD10, Spi15, WLV17, WB19, WHL^{+}11, WWS11, WKAM^{+}19, XZC^{+}16, ZZY^{+}10]. Tropical [ASA^{+}18, DSLLL19, WMM18, AJC15, BBLN11, BPA12, BSY^{+}16, CF13a, CEPPR14, CRJ^{+}14, CJW^{+}19, CPOMA15, CSS^{+}16, CKB^{+}16, DSS^{+}11, HE10, HOD^{+}17, JKKM13, KCM^{+}10, LWE^{+}11, LJL^{+}18, LÁSDC18, MRBR10, MBC^{+}16, MSS^{+}18, MLS^{+}18, MSD^{+}14, MRC^{+}16, PZHD18, RBG^{+}10, SWP11, SPGRP^{+}17, VGM14, VFS^{+}15, WLS^{+}11, WCC^{+}17, WLL^{+}11, WMI^{+}17, XFLM14, ZOB^{+}15, dKNL^{+}15]. trout [CBP10]. trumps [GWB^{+}14]. tube [BBR^{+}14]. tube-building [BBR^{+}14]. tunas [LRS^{+}10]. tundra [DHZ^{+}19, MDF^{+}14, MW15]. tuned [ŠNZ^{+}14]. tunicate [LBR^{+}13]. tunnel [PBV16]. turbid [BVvB^{+}19, SW14]. Turbidity [HYK^{+}15, LS14, NXL^{+}18, SVLS^{+}16, TRA19, VBBR17]. Turbulence [CLN^{+}19, FBL15, MCCA18, AGML18, CT18a, GTPB^{+}11, JCS^{+}18, LP10, PPL10, RMDK10, VPC10]. Turbulence-mediated [CLN^{+}19]. Turbulent [vH19, CSD10, LBS17, PTS^{+}19, TDS^{+}10, TF11, WVL^{+}18]. turf [LCS^{+}19]. turn [RPH^{+}10]. turn-of-the-century [RPH^{+}10]. turnover [ADS^{+}17, BSSW11, GHSR^{+}16, HS10, LWWE^{+}18, MWBM19, REDW10, SSKdB14, WMC^{+}15]. Twenty [AEH19]. twilight [SBKO18]. Two [KM10, LABJ18, PST^{+}13, BBK^{+}15, CBK18, CBS^{+}17, CGT16, CUW11, CKCEP10, CRS^{+}17, FEW^{+}14, HSLH^{+}14, HHW^{+}19, HNL^{+}13, JLC^{+}15, JWS15, LDY^{+}16, LKS^{+}16, LALM16, LSD18, MTT17, MEM^{+}17, PK14, PE16b, RDP^{+}17, SMA13, SWM^{+}10, SBNC^{+}19, SIW^{+}11, SRM^{+}18, SSM^{+}19, TEGL11, TSK13, VBC^{+}12, VIS^{+}13, WFK^{+}16, XFH14]. two-component [PE16b]. two-dimensional [HSLH^{+}14]. two-layer [SBNC^{+}19]. Two-step [KM10]. type [TBLG14]. types [SOH^{+}18]. Typha [KOFN11]. typhoon [WCJ^{+}17]. typical [PSZ^{+}13].

U.K. [YJO^{+}19]. U.S. [HEH^{+}17, JCF^{+}10, LPLH18, OWFS11]. U.S.A [PHL^{+}18]. U.S.A. [AJG13, CWHP14, GMBL16, HHM^{+}18, HTLM18, KT13, KSG^{+}10, LHSBP18, QS19, VdRA^{+}19]. ubiquitous [NMST18]. Uca [LDCT11]. Ugandan [PHG13]. Uinta [HML^{+}14]. Ulleung [NTK^{+}18]. ultimate [RKLH11]. Ultrahigh [PD11, SSC^{+}10]. ultraviolet [BSH16, BCVAn10, CRS^{+}17, FOT^{+}15, HS11, HBB^{+}11, HKS^{+}15, KMF10, RWF^{+}12, SSPK^{+}12, SMC^{+}10, UVGS10, VMF^{+}11, WFB^{+}11]. Ulva [HZC^{+}13, ZXM^{+}11]. un-stratified [LBS17]. Unbalanced [BSA^{+}16]. Uncertainty [KBA^{+}14]. Uncoupling [BB11, WLV18]. Under-ice [BCRW15, BBC^{+}13, EM13, KIH^{+}15]. Underestimation [SNM^{+}15]. underlie [LCW^{+}17b]. underlying [OLF^{+}11]. Understanding [RNK^{+}16, RvSM17, SLC^{+}16]. understory [CHPH13, MRB11]. underwater [JGR^{+}14, SGH12, SPO^{+}18, TLH^{+}11]. Unexpected [GK14, MD15]. unicellular [ABB^{+}14, BAA^{+}13, GFH13, MFK^{+}13, SMLC^{+}18, SPHVA19, YLJ11]. Uniform [ZHD^{+}16]. Unimodal [CL11, SASB^{+}15]. uninvaded [EMO^{+}11]. unique [ANP^{+}14, DSLLL19, WM12, WGDAA19]. United [MLS^{+}14, BHC13, BHC14, BGB^{+}14, MRSS12, SDH^{+}14, WWC^{+}13]. universal [BDC^{+}14]. unmanaged [KKP^{+}19]. Unprecedented [SJM11]. unproductive [KBA^{+}14, RKG^{+}11]. Unraveling [WYW^{+}10, WSTG18, YWL^{+}17]. unrestored [LHSG15, LH17]. Untersee [SMA15]. upgrade [TYX^{+}19]. upland [MMFBB18]. uplift [MNW^{+}19]. upon [GMS^{+}18, HVD^{+}18]. upper [BIM^{+}16, BIS^{+}10, GOD^{+}18, GCH^{+}18, RS16, SWM^{+}18, SSH^{+}14, SRAB10, ZXN^{+}12]. Uptake [SMH^{+}11, ASSG12, BSR^{+}17, BMW10, BMBI12, BC19, BDS11, BB11, BLM^{+}10, CFD^{+}11, CLN^{+}19, FFA13, GWD^{+}16, GLF18, HS18, HCC^{+}13, JJ17, KRR16, KBVW12, LF16, LF17b, MFK^{+}13, MBP^{+}17, MEM^{+}17, MLK11, MSD^{+}14, NCT^{+}14, OALD10, PFvO^{+}18, RvSM17, SBC^{+}17, TJJ^{+}15, TKK^{+}17, WM17, WMM18, WFL^{+}12, XSAM12]. Upward [HSR^{+}10]. Upwelling [BWS10, CMK^{+}10, HLH13, ANP^{+}14, CMM^{+}11, CFVU11, GFT^{+}14, GRE^{+}16, HDK^{+}12, HHHT19, JAS^{+}15, KTS^{+}14, LS15, MFM^{+}12, SdlFdlF^{+}10, VFME18, VMCM^{+}17, WM12, WCJ^{+}15, WAB^{+}17, WDL^{+}17, ZCZ^{+}18, ZCY^{+}15]. upwelling-driven [WCJ^{+}15]. Upwelling-influence [HLH13]. upwelling-influenced [HHHT19]. urban [CKB^{+}16, FPG11, LHSG15, LH17, MH16, vBBM^{+}19]. Urbanization [PCO^{+}15, SLE10]. urchin [FGBS^{+}18]. urchins [SLG10]. urea [DBFL11, FPD^{+}10, SMR^{+}17]. Urrutia [CL11]. USA [CF10, FYVU17, KH16, KDGL19, SBM16]. USC [CVS^{+}10]. Use [HH14, KFP^{+}18, WCCP14, BMM^{+}13, BSM17, CF13a, DCCB17, GEC^{+}17, GTR^{+}13, HNHS^{+}15, Ker17, KSG^{+}10, KGvdH16, LLH^{+}15, LKK13, MHRH11, NCT^{+}14, OPA^{+}14, RASD10, SLE10, TT12, WYW^{+}10, ZTW^{+}11]. Using [GBMG12, HGvB^{+}13, LAM12, NTA14, NTM^{+}10, OLC18, PCPZ18, RASV^{+}17, SSB^{+}18, SHK13, SSYT14, AES11, AC15, BFD^{+}11, BBTK^{+}16, BAY^{+}14, CPPdAR^{+}13, CAQS16, CCK^{+}12, DWDH10, DTM18, FPP^{+}19, GPH^{+}13, GKS12, HBR^{+}14, HMFB16, HGD14, HESU13, HSBA10, HML^{+}14, JSH12, JD16, JTH^{+}13, KBA^{+}14, KGM14, LC12, LBR^{+}12, MH16, MMXC15, MMPSB14, OHKC^{+}12, PHL^{+}18, PFJ10, RKBA14, RGM^{+}11, SBT^{+}19, SJ11, SBNC^{+}19, SPO^{+}18, SC10, TIF^{+}15, TB18, TPM^{+}14, TMO^{+}18, WGCC14, WSTD10, ZLLM10]. using-small [TB18]. Utah [BPL^{+}19a, HML^{+}14]. utilization [BS18a, ETKL12, HRPW15, MKB^{+}19, OBM^{+}11, PRS^{+}18, SYdTP^{+}11]. UV [AdGAD14, EHW^{+}15, SGVR16, TAE^{+}18, VABMS^{+}12]. UV-A [EHW^{+}15]. UV-enhanced [AdGAD14]. UV-induced [TAE^{+}18]. UV-protection [SGVR16]. UVR [CCV^{+}18, DMSHC16].

V [CFD^{+}19]. VA [QS19]. Valleys [DKK^{+}14, DTKMK15]. Vallisneria [ZLLM10]. Valu [SPB^{+}14]. value [JLG10, TYX^{+}19]. values [DLP13, LHLT13]. valve [WHH^{+}11]. valvometry [SDS^{+}11]. vanadium [JBPM15]. vanadium-dependent [JBPM15]. Vancouver [DCRC16, RPMK17]. Variability [ASK^{+}11, HLJ12, LVM^{+}10, MB10, RLSC^{+}13, AFSM17, ACW^{+}18, ASR^{+}17, ADS^{+}17, AMB^{+}11, BA14, BLW15, BJ15, BDK^{+}17, BLS^{+}16, CJS^{+}17, Clo19, CH11, CKB^{+}16, DCRC16, EED10, FWS^{+}14, FMP^{+}13, GNHGM13, KH16, KTK^{+}13, KHK^{+}19, LRM17, MCLT15, ML19, MWC^{+}16, MGL^{+}13, MBBW11, MHPW18, NSG^{+}16, NLM^{+}12, PCD^{+}19, PSS^{+}14, PVLMT^{+}16, PBL^{+}18, RKBA14, RNK^{+}16, RG13, RNT^{+}19, RMH^{+}17, RVvdP^{+}17, RGGL^{+}12, RPL16, SHSK14, SPSG14, SLE10, STC^{+}11, SBR^{+}13, SCQ^{+}17, TNI19, TvBR^{+}19, TEZ^{+}18, TST^{+}19, VLMTEW11, VML^{+}19, VMMS^{+}13, WB19, WTC^{+}17, WJHS18, XFLM14, XDC^{+}19, ZSM14]. Variable [BLM^{+}10, BAY^{+}14, BCF^{+}17, CGP^{+}19, DBMP^{+}11, FLM^{+}19, GMS^{+}18, HGD14, MKB^{+}19, SLC18]. variables [CJC^{+}12, DC15, LGR^{+}12, VBGG^{+}13]. Variation [BMD17, BDC^{+}14, NCT^{+}15, PFH^{+}17, YJO^{+}19, Ano21c, BBCM^{+}13, BGP^{+}15, CDA16, DPM18, ELJ^{+}16, ETKL12, ETKL15, EKS^{+}18, FZL^{+}14, GFDC11, JPH^{+}18, KSY11, LAM12, LH19, LGW^{+}19, MDB19, NLHAA^{+}17, PMPD13, PTS12, RS16, RR13, RMJ^{+}18, RBI^{+}10, SSU^{+}16, SLP^{+}14, SvKP^{+}18, WRWPG19, WKK^{+}11, YYMN13, vBBM^{+}19]. Variations [AC17, TNMV^{+}10, USB^{+}10, CNL^{+}15, CB12, CB19, GLF17, HP19, HKU^{+}10, HSC^{+}11, LCW17a, OY10, RPMK17, RGO^{+}11, Scu16, SRA10, SH10b, WLS^{+}11, WVL^{+}18, YMB^{+}18, ZKL^{+}14]. varies [LDCT11, MTEM15, RCIB14]. variety [UIY^{+}11]. various [GdG11]. varved [JAD^{+}13]. varying [BLH^{+}13, GAH11, SSM^{+}19, Tho19, THFG16]. vector [BBCM^{+}13]. vector-specific [BBCM^{+}13]. vegetated [RWC16]. vegetation [LN11, NBG17, SKGT17]. vegetative [WZTK15, ZXM^{+}11]. vehicle [SPO^{+}18]. velocities [BM16, RMH^{+}17]. velocity [MB10, SVS^{+}19, VPC10, VLMTEW11]. vent [CGP^{+}19, SPB^{+}14]. vents [BOT^{+}15]. vermiculophylla [GSPM13]. Vertical [AGLM17, BRF^{+}17, HCW^{+}10, HCS11, IPGP10, LKT17, LCW17a, OSC14, OR16, SPFP11, AGML18, BM16, BLG^{+}15, BSB^{+}18, DHG^{+}17, EHW^{+}15, FOT^{+}15, HSR^{+}10, HPS^{+}10a, HPL11, JSFC18, KWGS18, LMR14, MWSB18, MvdPK^{+}15, OMSC13, OFGF12, PGB^{+}19, PK14, PFJ10, RCV^{+}14, RRCH^{+}19, RRGGCA19, RWF^{+}12, RHSD^{+}10, SAS^{+}11, SMN^{+}15, SVS^{+}19, TGGZS^{+}10, VMC^{+}13, WCB^{+}10, WMC^{+}18, WFB^{+}11]. Vertically [JSFC18, HV16, NL14]. vesiculosus [ARB^{+}19, RCJ15]. Vestnesa [HSP^{+}16]. via [KJG10, SMMF19, TMH^{+}10]. viability [KvdPVB13]. vicinity [GDD^{+}16]. Victoria [GNHGM13, MRSE14, PHJ12, OrIA10]. video [TIF^{+}15]. view [CAQS16, GvBBB17]. Vigo [VMCM^{+}17]. Viral [EB12, CPF16, PS13, PD11, USB^{+}10]. Virginia [EHT10]. viridis [RF13]. virus [HNL^{+}13, MGS12, MMWR17]. virus-host [MMWR17]. virus-induced [MGS12]. Viruses [LTPK^{+}18, BSB^{+}10, LTX^{+}17]. viscous [SGCC16]. visual [GDCM13]. Vitamin [BWP^{+}10, PBA^{+}15, CEB^{+}17, FLLH18, KMP^{+}11]. vitamins [KSWFG13]. volatile [HGG^{+}17]. volcanic [MLL^{+}14, MBE^{+}13]. Volcano [FWFB10, LFB^{+}10]. volume [GBK^{+}18]. vs [BAA^{+}13, CFD15, DBRB^{+}15, GPCJ16, KMF10, MCGF^{+}11, NXL^{+}18, PMY^{+}19b, SML^{+}19, WRB^{+}19, WMC^{+}18]. vulnerabilities [SFS^{+}16]. vulnerability [KS16, NFRU11, SBFB17, WKK^{+}11]. vulpes [HCD19].

Wadden [GML^{+}12, VPG^{+}19]. walled [SMA15]. wane [TBAS14]. Waquoit [MDE11]. warm [BWS10, MDB16, RGO^{+}11, SKK^{+}13, TIF^{+}15, VLWV14, WRB^{+}19, WTC^{+}17]. warm-core [WRB^{+}19]. Warmer [SHD^{+}11, CEPPR14, KSP^{+}12, WBB^{+}17]. Warming [GBK^{+}18, YAC^{+}19, AHS11, CLHL12, DMSHC16, GHSR^{+}16, MLGZ16, NRL15, NBDM16, PNR19, QFH18, RSE^{+}17, RCIB14, RSTS^{+}18, RPH^{+}10, SMF10, Sch19, Tad10, VMF^{+}11, WSUC^{+}18, WRH^{+}17, XFH14, ZNVF16, ZHG15]. warnings [PCJK13]. Washington [KT13]. Wastewater [BHM^{+}17, KCB^{+}17, MACM11]. Water [CFD15, FCRW^{+}16, HHM^{+}18, Kus14, PMP^{+}17, RMJ^{+}18, Rie15, UA10, YAC^{+}19, AFG^{+}16, AES11, AdBVA10, AHS11, ÁSNCÁ^{+}13, ASH^{+}14, BHB^{+}19, BPGE13, BC19, BGW^{+}15, BNW^{+}14b, BGB^{+}14, BBS12, CDW^{+}16, CKP^{+}15, Clo19, CKCEP10, CFW^{+}14, DFWPK16, DWDH10, EHT10, ERA^{+}12, FEW^{+}14, FSCB11, FPD^{+}10, FLP^{+}10, FVSL19, FDB^{+}15, FBFR13, FDS^{+}18, FYT^{+}12, GTPB^{+}11, GGC^{+}14, GdVT^{+}11, GDD^{+}16, GGL^{+}18, GAM^{+}19, HCK10, HMV^{+}18, HT17a, HJB^{+}12, HCW^{+}10, HCLS11, HMHI13, HD19, HGvB^{+}13, IHSS^{+}19, JCS^{+}18, JAD^{+}13, JMM14, JLR^{+}17, KYC^{+}15, KCL^{+}14, Kir13, KB15, KKP^{+}19, KFP^{+}18, KZR^{+}16, KNL10, LKF^{+}18, LFGK10, LH19, LALM16, LBB18, LGC13a, LGC13b, MKB^{+}19, MGW^{+}13, MVT^{+}17, MMH^{+}18, MW15, MRC^{+}16, NO17, NWT^{+}19, OBT^{+}11, PMP^{+}12, PH15, PHLSSS19, RPI^{+}12, RSG11, RWM^{+}19, RHSD^{+}10, SVLS^{+}16, SCR^{+}12, SNO^{+}16, SLK^{+}14, SJB^{+}19, SAPI14]. water [SBdB10, SBK18, SWD11, SCL^{+}19, SSB^{+}16, TIF^{+}15, TvBR^{+}19, TBSL17, TMH^{+}18, UCOG16, UIY^{+}11, VLDM19, VLWV14, VHR^{+}11, WP14, WLV18, WCM19, WDX^{+}11, WCJ^{+}17, WBZ^{+}13, WXMS10, WJHS18, WFB^{+}11, WSB^{+}13, YKT^{+}15, YWL^{+}17, ZZAC13, ZOB^{+}15, ZHG15]. water-column [ERA^{+}12, SSB^{+}16]. water-level [GTPB^{+}11]. water-quality [Clo19]. Waterborne [HNZ^{+}16]. waterlouse [FA10]. waters [ACA^{+}11, ADS^{+}17, BDB^{+}14, BHS^{+}16, BSCG17, BHG^{+}18, BHB^{+}12, CCV^{+}18, CDA16, CWHP14, EBMR12, GMBL16, GM12, HJT^{+}13a, HMV12, HATF17, HSC^{+}14, JYS18, JM16, JHW^{+}19, JBT11, KP13, KMH^{+}17, LDY^{+}16, LK15, LCZ^{+}19, LÁSDC18, MLCD13, MBBG^{+}12, OMB^{+}16, PSZ^{+}13, Piw19, RS19, RGGL^{+}12, RGLM^{+}12, SAH^{+}19, SWD^{+}14, SWM^{+}18, SHK13, SW14, SOH^{+}18, SL10a, SDMK10, TMF^{+}14, TFLS14, WMBR13, WSM^{+}19, WM17, YHS^{+}17, vH19, HJT^{+}13b]. Watershed [TT12, BMBI12, DTM18, HAA^{+}19, JLR^{+}17, PFH^{+}17, SHL^{+}18, UIY^{+}11]. watersheds [CBK18, TWP13]. watsonii [GFH13, MFK^{+}13]. Wave [HCD19, HFP10, RCV^{+}14, VP15b, BBR12, IOB^{+}11, JD16, KFJ13, MMGO^{+}17b, MBBW11, MP17, NBG17, SVLS^{+}16, SPG^{+}13, VMMS^{+}13, WKSR13, WZTK15, ZWA^{+}14, vH19]. wave-driven [WKSR13, WZTK15]. wave-exposed [MBBW11]. wave-imposed [JD16]. Wave-induced [HFP10, RCV^{+}14]. Waves [LdlSB^{+}12, ABS^{+}19, CTH15, NI10, NRL15, PPL10, RDC^{+}19, RDZ^{+}13, SGH12, SPG^{+}13, SI10, SSP^{+}18, SSN12, VPMrI12, VBBR15, VMI13, VMCM^{+}17]. wax [CBP12, PT11, TBAS14]. Weak [KBJ^{+}18, NZH^{+}11, XSAM12]. weakly [BHB^{+}19, RSJ^{+}18]. weather [BJ15, JLR^{+}17, KTK^{+}13]. weather-climate [BJ15]. web [BCC^{+}12, CPPdAR^{+}13, CS12, DFK^{+}17, DvOR^{+}16, FHR^{+}15, FCRW^{+}16, FPSL18, HOD^{+}17, KGL^{+}16, KWB^{+}16, LEN^{+}15, LJL^{+}18, LHLT13, NB17, PH13, PDER10, PLE^{+}17, RHV^{+}13, SCF^{+}15, ŠNZ^{+}14, SMG12, SL10b, VMF^{+}11, VMC^{+}13, WDJF12, WRO^{+}11, WD15, vOSH12]. webs [CBF10, DML17, DRP^{+}17, GLS^{+}13, GFDC11, GRDPL14, HDDH^{+}17, JTV^{+}16, KBA^{+}12, LPLH18, LWWC^{+}16, MDF^{+}14, MPK^{+}13, MBLD15, ŠGN^{+}19, SBA^{+}11, SCP^{+}16, TIF^{+}15]. webs-evidence [JTV^{+}16]. Weddell [MdBKL13]. Wedderburn [SdlFdlF^{+}10]. wedge [REE^{+}12, SCR^{+}12]. week [SSH^{+}14]. Weekly [Piw19, RMNZ12]. weight [ASSG12, LCW17a, LFC17]. weissflogii [HBB^{+}11, TJJ^{+}15]. well [TMF^{+}14, TMH^{+}10]. well-oxygenated [TMF^{+}14, TMH^{+}10]. West [BBTK^{+}16, RVvdP^{+}17, RPMK17, SS12b, SS12c, VdRA^{+}19, RHV^{+}13]. Western [FGMN17, HNSM12, WMI^{+}17, BSR^{+}17, BBK^{+}15, CLJ^{+}19, CSS^{+}16, DDH^{+}19, DKSA19, FMM^{+}14, HCH^{+}19, JABZ19, JHLK^{+}19, Ker17, LEK^{+}18, NRS16, NO17, OBL^{+}19, PE13, RS19, RWM^{+}14, SKK^{+}13, SFI^{+}18, SBC^{+}17, SSN12, TLG^{+}11, UFW^{+}18, VCM13, WCC^{+}17, WTC^{+}17, HVJ^{+}19, LYH17, MMD18, NXL^{+}18, PTS12, TSSH19, WSTD10]. wet [dCGS19]. wet-phase [dCGS19]. Wetland [HMFF10, JMJ^{+}19, BFD^{+}11, BSB^{+}18, EED10, FSCB11, MF19, OHKC^{+}12, RHMSE15, SSH^{+}16, TT12, TZD^{+}15, VZJ^{+}17, WVGB10]. Wetland-driven [HMFF10]. wetlands [ARML10, CSU13, FMP^{+}13, MA18, PBL^{+}18, SML^{+}19]. WH8102 [MZB^{+}15]. whales [BCF^{+}17, CdC^{+}11, NSO19]. Where [AHH^{+}16]. which [DKSA19]. Who [WXF^{+}15]. whole [DKG15, DFK^{+}17, DTM18, EED10, GKS12, KBJ^{+}18, NSG^{+}16, OHKC^{+}12, PCJK13, PCW19, ZCK^{+}16]. whole-ecosystem [OHKC^{+}12]. whole-lake [GKS12, PCW19, ZCK^{+}16]. whole-system [EED10]. wide [Meh10]. widens [IH18]. Widespread [SHSK14, PSD^{+}17, SDH^{+}14]. will [PSH^{+}11, SPTS15]. Willamette [APP12]. Winam [PHJ12, OrIA10]. Wind [LS15, WMI^{+}17, ABS^{+}19, BSSR10, CFVU11, HSR15, HCC^{+}13, ILPL13, KFJ13, MAF19, PBV16, SPG^{+}13, SSM^{+}19, VPMrI12, XDC^{+}19]. Wind-driven [LS15, BSSR10, CFVU11, HSR15, HCC^{+}13, MAF19]. wind-induced [XDC^{+}19]. wind-wave [KFJ13]. Windermere [SLPM15]. windows [FHS10]. windy [MFL11]. Winnipeg [BLS^{+}16]. Winter [GGTC^{+}18, JLR^{+}17, KIH^{+}15, TA14, FPPA^{+}11, GLMG15, MQP^{+}16, MDB19, MPAS17, NHP17, OBI12, PWS^{+}11, RLB^{+}10, RVvdP^{+}17, RGLM^{+}12, SSFF12, VLWV14]. Winter-mixing [GGTC^{+}18]. Wintertime [BJ15, VCM13]. Wisconsin [WMC^{+}15]. Within [UIY^{+}11, AJG13, BYD19, BVC^{+}14, BLLB12, CBFK19, DTKMK15, EM13, EED10, FLM^{+}19, FDH^{+}14, IR16, JMJ^{+}19, LCM^{+}17, LCS^{+}19, MPM^{+}15, MAB^{+}17, MBH^{+}15, MRC^{+}16, NRS16, PTS12, RHV^{+}13, RS19, RPB17, RMDK10, SC10, TNMV^{+}10, WBG^{+}16, WZTK15, WKK^{+}11, WMM18, WJHS18, ZKL^{+}14]. Within-lake [UIY^{+}11]. within-population [ZKL^{+}14]. without [LWE^{+}19, SK19, WCJ^{+}17]. Woods [RPH^{+}10]. woody [WLL^{+}11]. world [WXF^{+}15]. worms [RF13]. Wyoming [BBK^{+}15].

xanthophyll [BHV^{+}17, KMF10]. Xestospongia [MBLP11].

Yangtze [GLI^{+}15, ZZW16]. Yarra [REE^{+}12]. year [AMNU16, CHHT18, EMH12, SDS^{+}11, SGRB10, WRS13]. years [AEH19, BPRG^{+}18, CSC^{+}11, DKG15, DBRB^{+}15, FVSL19, JAD^{+}13, KIH^{+}15, LYL^{+}17, SSGL19]. Yellow [HZC^{+}13, LZC^{+}14, LLW^{+}18, SW14, SCQ^{+}17, TEZ^{+}18, WXF^{+}15, WLR17, XZGW17]. yield [BRF^{+}17, EMB12, KBT16]. York [HHM^{+}18, EMH12, EP14, EHT10, HMFF10, HMFF12, PE16b, QS19]. Young [MSAM18]. Younger [Ano17l, LOS12, ZXZ17b]. Yungui [LCW^{+}17b, ZZY^{+}10].

Zealand [MWS10]. zebra [CS12]. zetterstedtii [SJM11]. Zhejiang [JHW^{+}19]. Zinc [SES18, HS18, TNMV^{+}10, XSAM12]. Zizaniopsis [LHSBP18]. zone [BG10a, BVvB^{+}19, BSC^{+}15, CTG15, DTFR12, FUS^{+}16, FCD12, FRP^{+}14, HMFF12, JD16, KBH^{+}19, KMC^{+}15, KBL^{+}10, LKT17, MSSH12, MGK15, MG17, MBH^{+}15, MD15, NHS^{+}12, OBT^{+}11, RGB^{+}19, SBKO18, SC10, TSB^{+}19, VFME18, VGM14, WMM18]. zones [CRJ^{+}14, CMK^{+}10, GCH^{+}18, GYP^{+}18, HFP10, JWS15, MSM^{+}17, ORC^{+}17, SBdB10, SPG^{+}13, YYMN13]. zooplankter [BH16, WLV18]. zooplanktivorous [GBK^{+}18]. Zooplanktivory [Edm11]. Zooplankton [GTPB^{+}11, GNWDL19, Kiø13, PCJK13, WRS13, AAO^{+}19, AA18, BBSK18, BSSR10, BBQ^{+}10, BSH16, BCVAn10, BSY^{+}16, BCM^{+}17, CPOMA15, CHV^{+}17, DHG^{+}17, DLP13, DBRB^{+}15, EHW^{+}15, FWvD^{+}18, FLLH18, GLMG15, GMJW13, GBT^{+}17, HRMD19, HPCD13, HDP15, HV16, HLFM^{+}10, HSTK15, HMH^{+}16, HPL11, IBPG17, KVMA17, KGM14, KSTA18a, KGT12, KBL^{+}10, LL11, LWrDM^{+}12, LWWC^{+}16, LV16, MBK^{+}11, MF19, MKK15, MXWC11, MGL^{+}13, MSM^{+}17, MDSG18, NZH^{+}11, NL14, OR16, PGB^{+}19, PBV16, PVA^{+}19, RMF11, RKLH11, RRGGCA19, SMMF19, SGJB14, SWM^{+}18, SRM^{+}18, SHL^{+}18, VKC18, VABMS^{+}12, WRB^{+}19, WCB^{+}10, WCCP14, WFB^{+}11, WBB^{+}17, YP18, dKYH^{+}12, dKNL^{+}15]. zooxanthellate [CRB^{+}17]. Zostera [AHJS15, DIC^{+}18, EMO^{+}11, FJBP15, HHHT19, HBM11, LdlSB^{+}12, MZH15, MHH^{+}17, MMBP18, RBM14].

References

Austin:2011:SSL

[AA11] Jay A. Austin and Joshua Allen. Sensitivity of summer Lake Superior thermal structure to meteorological forcing. Limnology and Oceanography, 57(4):1141–1154, May 2011. CODEN LIOCAH. ISSN 0024-3590.

Azan:2018:ICD

[AA18] Shakira S. E. Azan and Shelley E. Arnott. The impact of calcium decline on population growth rates of crustacean zooplankton in Canadian Shield lakes. Limnology and Oceanography, 63(2): 602–616, March 2018. CODEN LIOCAH. ISSN 0024-3590.

Attermeyer:2019:PTI

[AAC^{+}19] Katrin Attermeyer, Sara Andersson, Núria Catalán, Karolina Einarsdottir, Marloes Groeneveld, Anna J. Székely, and Lars J. Tranvik. Potential terrestrial influence on transparent exopolymer particle concentrations in boreal freshwaters. Limnology and Oceanography, 64(6):2455–2466, November 2019. CODEN LIOCAH. ISSN 0024-3590.

Almeda:2011:MRC

[AACS11] Rodrigo Almeda, Miquel Alcaraz, Albert Calbet, and Enric Saiz. Metabolic rates and carbon budget of early developmental stages of the marine cyclopoid copepod Oithona davisae. Limnology and Oceanography, 56(1):403–414, January 2011. CODEN LI-OCAH. ISSN 0024-3590.

Ajani:2014:DDR

[AAIA14a] Penelope A. Ajani, Andrew P. Allen, Tim Ingleton, and Leanne Armand. A decadal decline in relative abundance and a shift in microphytoplankton composition at a long-term coastal station off southeast Australia. Limnology and Oceanography, 59(2): 519–531, March 2014. CODEN LIOCAH. ISSN 0024-3590.

Ajani:2014:EDD

[AAIA14b] Penelope A. Ajani, Andrew P. Allen, Tim Ingleton, and Leanne Armand. Erratum: a decadal decline in relative abundance and a shift in microphytoplankton composition at a long-term coastal station off southeast Australia. Limnology and Oceanography, 61(1):2240–2242, November 2014. CODEN LIOCAH. ISSN 0024-3590.

Aarflot:2019:CBD

[AAO^{+}19] Johanna M. Aarflot, Dag L. Aksnes, Anders F. Opdal, Hein Rune Skjoldal, and Øyvind Fiksen. Caught in broad daylight: Topographic constraints of zooplankton depth distributions. Limnology and Oceanography, 64(3):849–859, May 2019. CODEN LIOCAH. ISSN 0024-3590.

Agawin:2014:DUC

[ABB^{+}14] Nona S. R. Agawin, Mar Benavides, Antonio Busquets, Pere Ferriol, Lucas J. Stal, and Javier Arístegui. Dominance of unicellular cyanobacteria in the diazotrophic community in the Atlantic Ocean. Limnology and Oceanography, 59(6):623–637, March 2014. CODEN LIOCAH. ISSN 0024-3590.

Ardyna:2017:SBG

[ABD^{+}17] M. Ardyna, M. Babin, E. Devred, A. Forest, M. Gosselin, P. Raimbault, and J.-É. Tremblay. Shelf-basin gradients shape ecological phytoplankton niches and community composition in the coastal Arctic Ocean (Beaufort Sea). Limnology and Oceanography, 65(9):2113–2132, September 2017. CODEN LI-OCAH. ISSN 0024-3590.

Arnon:2019:SDI

[ABS^{+}19] Ali Arnon, Steve Brenner, John S. Selker, Isaac Gertman, and Nadav G. Lensky. Seasonal dynamics of internal waves governed by stratification stability and wind: Analysis of high-resolution observations from the Dead Sea. Limnology and Oceanography, 64(5):1864–1882, September 2019. CODEN LIOCAH. ISSN 0024-3590.

Arellano:2015:ACN

[AC15] Ana R. Arellano and Paula G. Coble. Assessing carbon and nutrient inputs in a spring-fed estuary using fluorescence spectroscopy and discriminatory classification. Limnology and Oceanography, 60(3):789–804, May 2015. CODEN LIOCAH. ISSN 0024-3590.

Arriola:2017:VCB

[AC17] Jill M. Arriola and Jaye E. Cable. Variations in carbon burial and sediment accretion along a tidal creek in a Florida salt marsh. Limnology and Oceanography, 62(S1):S15–S28, November 2017. CODEN LIOCAH. ISSN 0024-3590.

Almeda:2011:TRC

[ACA^{+}11] Rodrigo Almeda, Albert Calbet, Miquel Alcaraz, Enric Saiz, Isabel Trepat, Laura Arin, Juancho Movilla, and Violeta Saló. Trophic role and carbon budget of metazoan microplankton in northwest Mediterranean coastal waters. Limnology and Oceanography, 56(1):415–430, January 2011. CODEN LIOCAH. ISSN 0024-3590.

Aastrom:2018:MCS

[ÅCA^{+}18] Emmelie K. L. Åström, Michael L. Carroll, William G. Ambrose, Jr., Arunima Sen, Anna Silyakova, and JoLynn Carroll. Methane cold seeps as biological oases in the high-Arctic deep sea. Limnology and Oceanography, 63(S1):S209–S231, March 2018. CODEN LIOCAH. ISSN 0024-3590.

Asmala:2017:ECF

[ACC^{+}17] Eero Asmala, Jacob Carstensen, Daniel J. Conley, Caroline P. Slomp, Johanna Stadmark, and Maren Voss. Efficiency of the coastal filter: Nitrogen and phosphorus removal in the Baltic Sea. Limnology and Oceanography, 62(S1):S222–S238, November 2017. CODEN LIOCAH. ISSN 0024-3590.

Asmala:2019:RCK

[ACC^{+}19] Eero Asmala, Jacob Carstensen, Daniel J. Conley, Caroline P. Slomp, Johanna Stadmark, and Maren Voss. A reply to the comment by Karlsson et al. Limnology and Oceanography, 64 (4):1832–1833, July 2019. CODEN LIOCAH. ISSN 0024-3590. See [KBH^{+}19].

Amyot:2010:RCE

[ACD10] Marc Amyot, Saad Chidam, and Éric Demers. Response to “Comparative estimate of P fluxes in lakes” by J. M. Sereda and J. J. Hudson. Limnology and Oceanography, 55(1):466–468, January 2010. CODEN LIOCAH. ISSN 0024-3590. See [SH10a].

Anderson:2018:RVA

[ACW^{+}18] N. J. Anderson, C. J. Curtis, E. J. Whiteford, V. J. Jones, S. Mc-Gowan, G. L. Simpson, and J. Kaiser. Regional variability in the atmospheric nitrogen deposition signal and its transfer to the sediment record in Greenland lakes. Limnology and Oceanography, 64(4):2250–2265, September 2018. CODEN LIOCAH. ISSN 0024-3590.

Alderkamp:2010:CPC

[AdBVA10] Anne-Carlijn Alderkamp, Hein J. W. de Baar, Ronald J. W. Visser, and Kevin R. Arrigo. Can photoinhibition control phytoplankton abundance in deeply mixed water columns of the Southern Ocean? Limnology and Oceanography, 55(3):1248– 1264, May 2010. CODEN LIOCAH. ISSN 0024-3590.

Anderson:2018:STE

[ADCH18] Sean R. Anderson, Quintin P. Diou-Cass, and Elizabeth L. Harvey. Short-term estimates of phytoplankton growth and mortality in a tidal estuary. Limnology and Oceanography, 63(6): 2411–2422, November 2018. CODEN LIOCAH. ISSN 0024-3590.

Agusti:2014:CUE

[AdGAD14] Susana Agustí, Aurore Regaudie de Gioux, Jesús M. Arrieta, and Carlos M. Duarte. Consequences of UV-enhanced community respiration for plankton metabolic balance. Limnology and Oceanography, 59(1):223–232, January 2014. CODEN LIOCAH. ISSN 0024-3590.

Asher:2017:PDS

[ADS^{+}17] E. C. Asher, J. W. H. Dacey, M. Stukel, M. C. Long, and P. D. Tortell. Processes driving seasonal variability in DMS, DMSP, and DMSO concentrations and turnover in coastal Antarctic waters. Limnology and Oceanography, 62(1):104–124, January 2017. CODEN LIOCAH. ISSN 0024-3590.

Arroita:2019:TYD

[AEH19] Maite Arroita, Arturo Elosegi, and Robert O. Hall, Jr. Twenty years of daily metabolism show riverine recovery following sewage abatement. Limnology and Oceanography, 64(S1):S77– S92, January 2019. CODEN LIOCAH. ISSN 0024-3590.

Adler:2011:QRM

[AES11] Michal Adler, Werner Eckert, and Orit Sivan. Quantifying rates of methanogenesis and methanotrophy in Lake Kinneret sediments (Israel) using pore-water profiles. Limnology and Oceanography, 56(4):1525–1535, July 2011. CODEN LIOCAH. ISSN 0024-3590.

Adhitya:2016:EDN

[AFG^{+}16] A. Adhitya, A. M. Folkard, L. L. Govers, M. M. van Katwijk, H. H. de Iongh, P. M. J. Herman, and T. J. Bouma. The exchange of dissolved nutrients between the water column and substrate pore-water due to hydrodynamic adjustment at seagrass meadow edges: a flume study. Limnology and Oceanography, 61(6):2286–2295, November 2016. CODEN LIOCAH. ISSN 0024-3590.

Agawin:2017:TSV

[AFSM17] Nona S. R. Agawin, Pere Ferriol, Eva Sintes, and Gabriel Moyà. Temporal and spatial variability of in situ nitrogen fixation activities associated with the Mediterranean seagrass Posidonia oceanica meadows. Limnology and Oceanography, 62(6):2575– 2592, November 2017. CODEN LIOCAH. ISSN 0024-3590.

Amaral:2016:SLB

[AGCA16] Valentina Amaral, Daniel Graeber, Danilo Calliari, and Cecilia Alonso. Strong linkages between DOM optical properties and main clades of aquatic bacteria. Limnology and Oceanography, 61(3):906–918, May 2016. CODEN LIOCAH. ISSN 0024-3590.

Abdolahpour:2017:VMC

[AGLM17] Maryam Abdolahpour, Marco Ghisalberti, Paul Lavery, and Kathryn McMahon. Vertical mixing in coastal canopies. Limnology and Oceanography, 62(1):26–42, January 2017. CODEN LIOCAH. ISSN 0024-3590.

Abdolahpour:2018:IFF

[AGML18] Maryam Abdolahpour, Marco Ghisalberti, Kathryn McMahon, and Paul S. Lavery. The impact of flexibility on flow, turbulence, and vertical mixing in coastal canopies. Limnology and Oceanography, 63(6):2777–2792, November 2018. CODEN LI-OCAH. ISSN 0024-3590.

Attard:2014:SRB

[AGMR14] Karl M. Attard, Ronnie N. Glud, Daniel F. McGinnis, and Søren Rysgaard. Seasonal rates of benthic primary production in a Greenland fjord measured by aquatic eddy correlation. Limnology and Oceanography, 59(5):1555–1569, September 2014. CO-DEN LIOCAH. ISSN 0024-3590.

Araujo:2018:MSH

[AHD^{+}18] Beatriz Ferreira Araujo, Holger Hintelmann, Brian Dimock, Rodrigo de Lima Sobrinho, Marcelo Correa Bernardes, Marcelo Gomes de Almeida, Alex V. Krusche, Thiago Pessanha Rangel, Fabiano Thompson, and Carlos Eduardo de Rezende.

Mercury speciation and Hg stable isotope ratios in sediments from Amazon floodplain lakes — Brazil. Limnology and Oceanography, 63(3):1134–1145, May 2018. CODEN LIOCAH. ISSN 0024-3590.

Adams:2016:FBS

[AHH^{+}16] Matthew P. Adams, Renae K. Hovey, Matthew R. Hipsey, Louise C. Bruce, Marco Ghisalberti, Ryan J. Lowe, Renee K. Gruber, Leonardo Ruiz-Montoya, Paul S. Maxwell, David P. Callaghan, Gary A. Kendrick, and Katherine R. O’Brien. Feedback between sediment and light for seagrass: Where is it important? Limnology and Oceanography, 61(6):1937–1955, November 2016. CODEN LIOCAH. ISSN 0024-3590.

Alexandre:2015:DON

[AHJS15] Ana Alexandre, Paul W. Hill, Davey L. Jones, and Rui Santos. Dissolved organic nitrogen: a relevant, complementary source of nitrogen for the seagrass Zostera marina. Limnology and Oceanography, 60(5):1477–1483, September 2015. CODEN LI-OCAH. ISSN 0024-3590.

Alsterberg:2011:RSW

[AHS11] Christian Alsterberg, Stefan Hulth, and Kristina Sundbäck. Response of a shallow-water sediment system to warming. Limnology and Oceanography, 56(6):2147–2160, November 2011. CO-DEN LIOCAH. ISSN 0024-3590.

Aumack:2015:LNE

[AJ15] C. F. Aumack and A. R. Juhl. Light and nutrient effects on the settling characteristics of the sea ice diatom Nitzschia frigida. Limnology and Oceanography, 60(3):765–776, May 2015. CO-DEN LIOCAH. ISSN 0024-3590.

Angles:2015:RCP

[AJC15] Sílvia Anglès, Antoni Jordi, and Lisa Campbell. Responses of the coastal phytoplankton community to tropical cyclones revealed by high-frequency imaging flow cytometry. Limnology and Oceanography, 60(5):1562–1576, September 2015. CODEN LIOCAH. ISSN 0024-3590.

Arp:2013:RLI

[AJG13] Christopher D. Arp, Benjamin M. Jones, and Guido Grosse. Recent lake ice-out phenology within and among lake districts of Alaska, U.S.A. Limnology and Oceanography, 58(6):2013–2028, November 2013. CODEN LIOCAH. ISSN 0024-3590.

Agha:2014:SDS

[ALdML^{+}14] Ramsy Agha, Marià Àngeles Lezcano, Marià del Mar Labrador, Samuel. Cirés, and Antonio Quesada. Seasonal dynamics and sedimentation patterns of Microcystis oligopeptide-based chemotypes reveal subpopulations with different ecological traits. Limnology and Oceanography, 59(3):861–871, May 2014. CODEN LIOCAH. ISSN 0024-3590.

Allen:2013:NSN

[ALG^{+}13] Dennis M. Allen, Stacy A. Luthy, Jason A. Garwood, Robert F. Young, and Richard F. Dame. Nutrient subsidies from nekton in salt marsh intertidal creeks. Limnology and Oceanography, 58 (3):1048–1060, May 2013. CODEN LIOCAH. ISSN 0024-3590.

Abramson:2010:EBS

[ALL^{+}10a] Lynn Abramson, Cindy Lee, Zhanfei Liu, Stuart G. Wakeham, and Jennifer Szlosek. Exchange between suspended and sinking particles in the northwest Mediterranean as inferred from the organic composition of in situ pump and sediment trap samples. Limnology and Oceanography, 55(2):725–739, March 2010. CODEN LIOCAH. ISSN 0024-3590.

Allen:2010:GEF

[All10b] Michael R. Allen. Genetic and environmental factors influence survival and hatching of diapausing eggs. Limnology and Oceanography, 55(2):549–559, March 2010. CODEN LIOCAH. ISSN 0024-3590.

Alongi:2017:MME

[Alo17] Daniel M. Alongi. Micronutrients and mangroves: Experimental evidence for copper limitation. Limnology and Oceanography, 62 (6):2759–2772, November 2017. CODEN LIOCAH. ISSN 0024-3590.

Atilla:2011:OVL

[AMB^{+}11] Nazan Atilla, Galen A. McKinley, Val Bennington, Matthew Baehr, Noel Urban, Michael DeGrandpre, Ankur R. Desai, and Chin Wu. Observed variability of Lake Superior pCO_{2}. Limnology and Oceanography, 56(3):775–786, May 2011. CODEN LIOCAH. ISSN 0024-3590.

Amin:2013:CRA

[AMMH^{+}13] Shady A. Amin, James W. Moffett, Willm Martens-Habbena, Jeremy E. Jacquot, Yang Han, Allan Devol, Anitra E. Ingalls, David A. Stahl, and E. Virginia Armbrust. Copper requirements of the ammonia-oxidizing archaeon Nitrosopumilus maritimus SCM1 and implications for nitrification in the marine environment. Limnology and Oceanography, 58(6):2037–2045, November 2013. CODEN LIOCAH. ISSN 0024-3590.

Aslam:2016:PDC

[AMNU16] Shazia N. Aslam, Christine Michel, Andrea Niemi, and Graham J. C. Underwood. Patterns and drivers of carbohydrate budgets in ice algal assemblages from first year Arctic sea ice. Limnology and Oceanography, 61(3):919–937, May 2016. CODEN LIOCAH. ISSN 0024-3590.

Antoniades:2011:CEA

[AMQ^{+}11] Dermot Antoniades, Neal Michelutti, Roberto Quinlan, Jules M. Blais, Sylvia Bonilla, Marianne S. V. Douglas, Reinhard Pienitz, John P. Smol, and Warwick F. Vincent. Cultural eutrophication, anoxia, and ecosystem recovery in Meretta Lake, High Arctic Canada. Limnology and Oceanography, 56(2):639–650, March 2011. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:2010:EOI

[Ano10] Anonymous. Erratum: Oxygen isotopic exchange and fractionation during bacterial ammonia oxidation. Limnology and Oceanography, 55(4):1805, July 2010. CODEN LIOCAH. ISSN 0024-3590. See [CMB10].

Anonymous:2015:E

[Ano15] Anonymous. Erratum. Limnology and Oceanography, 60(3): 1102, May 2015. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:2017:IIa

[Ano17a] Anonymous. Issue information. Limnology and Oceanography, 62(1):1–2, January 2017. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:2017:IIb

[Ano17b] Anonymous. Issue information. Limnology and Oceanography, 62(3):863–864, May 2017. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:2017:IIc

[Ano17c] Anonymous. Issue information. Limnology and Oceanography, 62(4):1329–1330, July 2017. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:2017:IId

[Ano17d] Anonymous. Issue information. Limnology and Oceanography, 62(5):1797–1798, September 2017. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:2017:IIe

[Ano17e] Anonymous. Issue information. Limnology and Oceanography, 62(6):2343–2344, November 2017. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:2017:IIf

[Ano17f] Anonymous. Issue information. Limnology and Oceanography, 62(S1):S1–S2, November 2017. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:2017:IIIa

[Ano17g] Anonymous. Issue information — instr to contrib. Limnology and Oceanography, 62(1):394–396, January 2017. CODEN LI-OCAH. ISSN 0024-3590.

Anonymous:2017:IIIb

[Ano17h] Anonymous. Issue information — instr to contrib. Limnology and Oceanography, 62(2):859–861, March 2017. CODEN LI-OCAH. ISSN 0024-3590.

Anonymous:2017:IIIc

[Ano17i] Anonymous. Issue information — instr to contrib. Limnology and Oceanography, 62(5):1794–1796, July 2017. CODEN LI-OCAH. ISSN 0024-3590.

Anonymous:2017:IIId

[Ano17j] Anonymous. Issue information — instr to contrib. Limnology and Oceanography, 62(6):2879–2881, November 2017. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:2017:IIIe

[Ano17k] Anonymous. Issue information — instr to contrib. Limnology and Oceanography, 62(S1):S400–S402, November 2017. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:2017:RSM

[Ano17l] Anonymous. Retraction of Southward migrations of the Atlantic Equatorial Currents during the Younger Dryas. Limnology and Oceanography, 65(9):2338, September 2017. CODEN LIOCAH. ISSN 0024-3590. See [ZXZ17b].

Anonymous:2018:Ea

[Ano18a] Anonymous. Erratum. Limnology and Oceanography, 63(3): 1444, May 2018. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:2018:Eb

[Ano18b] Anonymous. Erratum. Limnology and Oceanography, 63(4): 1832, July 2018. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:2018:IIa

[Ano18c] Anonymous. Issue information. Limnology and Oceanography, 63(1):1–2, January 2018. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:2018:IIb

[Ano18d] Anonymous. Issue information. Limnology and Oceanography, 63(2):507–508, March 2018. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:2018:IIc

[Ano18e] Anonymous. Issue information. Limnology and Oceanography, 63(3):1039–1040, May 2018. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:2018:IId

[Ano18f] Anonymous. Issue information. Limnology and Oceanography, 63(6):2305–2306, November 2018. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:2018:IIe

[Ano18g] Anonymous. Issue information. Limnology and Oceanography, 63(S1):S1–S2, March 2018. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:2018:IIIa

[Ano18h] Anonymous. Issue information — instr to contrib. Limnology and Oceanography, 63(4):503–505, January 2018. CODEN LI-OCAH. ISSN 0024-3590.

Anonymous:2018:IIIb

[Ano18i] Anonymous. Issue information — instr to contrib. Limnology and Oceanography, 63(2):1036–1038, March 2018. CODEN LI-OCAH. ISSN 0024-3590.

Anonymous:2018:IIIc

[Ano18j] Anonymous. Issue information — instr to contrib. Limnology and Oceanography, 63(6):2887–2889, November 2018. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:2018:IIId

[Ano18k] Anonymous. Issue information — instr to contrib. Limnology and Oceanography, 63(S1):S496–S498, March 2018. CODEN LI-OCAH. ISSN 0024-3590.

Anonymous:2019:AP

[Ano19a] Anonymous. ASLO page. Limnology and Oceanography, 64(S1): i–iii, January 2019. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:2019:E

[Ano19b] Anonymous. Erratum. Limnology and Oceanography, 64(3): 1422, May 2019. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:2019:EFS

[Ano19c] Anonymous. Erratum: Feeding strategies for the acquisition of high-quality food sources in stream macroinvertebrates: Collecting, integrating, and mixed feeding. Limnology and Oceanography, 64(4):1834, July 2019. CODEN LIOCAH. ISSN 0024-3590. See [GBB^{+}18].

Anonymous:2019:IICa

[Ano19d] Anonymous. Issue information — copyright page. Limnology and Oceanography, 64(1):iii, January 2019. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:2019:IICb

[Ano19e] Anonymous. Issue information — copyright page. Limnology and Oceanography, 64(2):iii, March 2019. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:2019:IICc

[Ano19f] Anonymous. Issue information — copyright page. Limnology and Oceanography, 64(3):iii, May 2019. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:2019:IICd

[Ano19g] Anonymous. Issue information — copyright page. Limnology and Oceanography, 64(4):iii, July 2019. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:2019:IICe

[Ano19h] Anonymous. Issue information — copyright page. Limnology and Oceanography, 64(5):iii, September 2019. CODEN LI-OCAH. ISSN 0024-3590.

Anonymous:2019:IICf

[Ano19i] Anonymous. Issue information — copyright page. Limnology and Oceanography, 64(6):iii, November 2019. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:2019:IIIa

[Ano19j] Anonymous. Issue information — instr to contrib. Limnology and Oceanography, 64(1):iv–vi, January 2019. CODEN LI-OCAH. ISSN 0024-3590.

Anonymous:2019:IIIb

[Ano19k] Anonymous. Issue information — instr to contrib. Limnology and Oceanography, 64(2):i–iii, March 2019. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:2019:IIIc

[Ano19l] Anonymous. Issue information — instr to contrib. Limnology and Oceanography, 66(4):i–iii, September 2019. CODEN LI-OCAH. ISSN 0024-3590.

Anonymous:2019:IIId

[Ano19m] Anonymous. Issue information — instr to contrib. Limnology and Oceanography, 64(6):2771–2773, November 2019. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:2019:IIMb

[Ano19n] Anonymous. Issue information — masthead. Limnology and Oceanography, 64(1):ii, January 2019. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:2019:IIMa

[Ano19o] Anonymous. Issue information — members. Limnology and Oceanography, 64(1):i, January 2019. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:2019:IIMc

[Ano19p] Anonymous. Issue information — members. Limnology and Oceanography, 64(2):i–ii, March 2019. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:2019:IIMd

[Ano19q] Anonymous. Issue information — members. Limnology and Oceanography, 64(3):i–ii, May 2019. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:2019:IIMe

[Ano19r] Anonymous. Issue information — members. Limnology and Oceanography, 64(4):i–ii, July 2019. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:2019:IIMf

[Ano19s] Anonymous. Issue information — members. Limnology and Oceanography, 64(5):i–ii, September 2019. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:2019:IIMg

[Ano19t] Anonymous. Issue information — members. Limnology and Oceanography, 64(6):i–ii, November 2019. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:2021:CGL

[Ano21a] Anonymous. Corrigendum for Giling et al. (2017). Delving deeper: Metabolic processes in the metalimnion of stratified lakes. Limnology and Oceanography, 62, 1288–1306 (doi: 10.1002/lno.10504). Limnology and Oceanography, 66(5):2088– 2092, May 2021. CODEN LIOCAH. ISSN 0024-3590. See [GSG^{+}17].

Anonymous:2021:CMK

[Ano21b] Anonymous. Corrigendum for Meyer-Kaiser et al. 2019 [Limnol. Oceanogr. 64(5) 1924–1938]. Limnology and Oceanography, 66 (8):3253, August 2021. CODEN LIOCAH. ISSN 0024-3590. See [MKBSK19].

Anonymous:2021:CSV

[Ano21c] Anonymous. Corrigendum: Spatial variation in the biochemical and isotopic composition of corals during bleaching and recovery. Limnology and Oceanography, 65(3):1611–1612, April 2021. CODEN LIOCAH. ISSN 0024-3590. See [WRWPG19].

Ahlgren:2014:UTM

[ANP^{+}14] Nathan A. Ahlgren, Abigail Noble, Allison P. Patton, Kathryn Roache-Johnson, Laurel Jackson, Daniela Robinson, Cedar McKay, Lisa R. Moore, Mak A. Saito, and Gabrielle Rocap. The unique trace metal and mixed layer conditions of the Costa Rica upwelling dome support a distinct and dense community of Synechococcus. Limnology and Oceanography, 61(1):2166–2184, November 2014. CODEN LIOCAH. ISSN 0024-3590.

Norethi:2013:AOM

[àNTS13] Katrin à. Nordi, Bo Thamdrup, and Carsten J. Schubert. Anaerobic oxidation of methane in an iron-rich Danish freshwater lake sediment. Limnology and Oceanography, 58(2):546–554, March 2013. CODEN LIOCAH. ISSN 0024-3590.

Albers:2012:ERS

[AP12] Sam J. Albers and Ellen L. Petticrew. Ecosystem response to a salmon disturbance regime: Implications for downstream nutrient fluxes in aquatic systems. Limnology and Oceanography, 57 (1):113–123, January 2012. CODEN LIOCAH. ISSN 0024-3590.

Almeida:2017:EFI

[APB^{+}17] Rafael M. Almeida, Felipe S. Pacheco, Nathan Barros, Emma Rosi, and Fábio Roland. Extreme floods increase CO_{2} outgassing from a large Amazonian river. Limnology and Oceanography, 63(3):989–999, May 2017. CODEN LIOCAH. ISSN 0024-3590.

Atkinson:2018:CWD

[APF^{+}18] Angus Atkinson, Luca Polimene, Elaine S. Fileman, Claire E. Widdicombe, Andrea J. McEvoy, Tim J. Smyth, Nicolas Djeghri, Sévrine F. Sailley, and Louise E. Cornwell. Comment. What drives plankton seasonality in a stratifying shelf sea? Some competing and complementary theories. Limnology and Oceanography, 63(6):2877–2884, November 2018. CODEN LIOCAH. ISSN 0024-3590. See response [KVA18].

Anthony:2012:MDW

[APP12] Sara E. Anthony, Fredrick G. Prahl, and Tawnya D. Peterson. Methane dynamics in the Willamette River, Oregon. Limnology and Oceanography, 58(1):1517–1530, September 2012. CODEN LIOCAH. ISSN 0024-3590.

Amundsen:2019:LTE

[APS^{+}19] Per-Arne Amundsen, Raul Primicerio, Aslak Smalås, Eirik H. Henriksen, Rune Knudsen, Roar Kristoffersen, and Anders Klemetsen. Long-term ecological studies in northern lakes — challenges, experiences, and accomplishments. Limnology and Oceanography, 64(S1):S11–S21, January 2019. CODEN LI-OCAH. ISSN 0024-3590.

Attard:2019:SMC

[ARB^{+}19] K. M. Attard, I. F. Rodil, P. Berg, J. Norkko, A. Norkko, and R. N. Glud. Seasonal metabolism and carbon export potential of a key coastal habitat: The perennial canopy-forming macroalga Fucus vesiculosus. Limnology and Oceanography, 64(1):149–164, January 2019. CODEN LIOCAH. ISSN 0024-3590.

Allgeier:2010:SNC

[ARML10] Jacob E. Allgeier, Amy D. Rosemond, Andrew S. Mehring, and Craig A. Layman. Synergistic nutrient colimitation across a gradient of ecosystem fragmentation in subtropical mangrovedominated wetlands. Limnology and Oceanography, 55(6):2660– 2668, November 2010. CODEN LIOCAH. ISSN 0024-3590.

Archer:2010:DSD

[ARW^{+}10] Stephen D. Archer, Maria Ragni, Richard Webster, Ruth L. Airs, and Richard J. Geider. Dimethyl sulfoniopropionate and dimethyl sulfide production in response to photoinhibition in Emiliania huxleyi. Limnology and Oceanography, 55(4):1579– 1589, July 2010. CODEN LIOCAH. ISSN 0024-3590.

Archer:2018:LDS

[ASA^{+}18] Stephen D. Archer, Jacqueline Stefels, Ruth L. Airs, Tracy Lawson, Timothy J. Smyth, Andrew P. Rees, and Richard J. Geider. Limitation of dimethylsulfoniopropionate synthesis at high irradiance in natural phytoplankton communities of the tropical Atlantic. Limnology and Oceanography, 63(1):227–242, January 2018. CODEN LIOCAH. ISSN 0024-3590.

Andersson:2014:NFS

[ASH^{+}14] Björn Andersson, Kristina Sundbäck, Maria Hellman, Sara Hallin, and Christian Alsterberg. Nitrogen fixation in shallowwater sediments: Spatial distribution and controlling factors. Limnology and Oceanography, 59(6):1932–1944, November 2014. CODEN LIOCAH. ISSN 0024-3590.

Antoine:2011:VOP

[ASK^{+}11] David Antoine, David A. Siegel, Tihomir Kostadinov, Stéphane Maritorena, Norm B. Nelson, Bernard Gentili, Vincenzo Vellucci, and Nathalie Guillocheau. Variability in optical particle backscattering in contrasting bio-optical oceanic regimes. Limnology and Oceanography, 57(4):955–973, May 2011. CODEN LIOCAH. ISSN 0024-3590.

Arnon:2016:TSD

[ASL16] A. Arnon, J. S. Selker, and N. G. Lensky. Thermohaline stratification and double diffusion diapycnal fluxes in the hypersaline Dead Sea. Limnology and Oceanography, 61(4):1214–1231, July 2016. CODEN LIOCAH. ISSN 0024-3590.

Alvarez-Salgado:2013:NIM

[ÁSNCÁ^{+}13] X. A. Álvarez-Salgado, M. Nieto-Cid, M. Álvarez, F. F. Pérez, P. Morin, and H. Mercier. New insights on the mineralization of dissolved organic matter in central, intermediate, and deep water masses of the northeast North Atlantic. Limnology and Oceanography, 58(2):681–696, March 2013. CODEN LIOCAH. ISSN 0024-3590.

Archer:2017:ACD

[ASR^{+}17] Stephanie K. Archer, Julia L. Stevens, Ryann E. Rossi, Kenan O. Matterson, and Craig A. Layman. Abiotic conditions drive significant variability in nutrient processing by a common Caribbean sponge, Ircinia felix. Limnology and Oceanography, 62(5):1783–1793, July 2017. CODEN LIOCAH. ISSN 0024-3590.

Alonso-Saez:2012:BUL

[ASSG12] Laura Alonso-Sáez, Olga Sánchez, and Josep M. Gasol. Bacterial uptake of low molecular weight organics in the subtropical Atlantic: Are major phylogenetic groups functionally different? Limnology and Oceanography, 57(3):798–808, May 2012. CO-DEN LIOCAH. ISSN 0024-3590.

Anderson:2019:DST

[ASW^{+}19] Thomas L. Anderson, Lawrence W. Sheppard, Jonathan A. Walter, Susan P. Hendricks, Todd D. Levine, David S. White, and Daniel C. Reuman. The dependence of synchrony on timescale and geography in freshwater plankton. Limnology and Oceanography, 64(2):483–502, March 2019. CODEN LIOCAH. ISSN 0024-3590.

Adly:2015:RHB

[ATP^{+}15] Carol L. Adly, Jean-Eric Tremblay, Rodney T. Powell, Evelyn Armstrong, Graham Peers, and Neil M. Price. Response of heterotrophic bacteria in a mesoscale iron enrichment in the northeast subarctic Pacific Ocean. Limnology and Oceanography, 60 (1):136–148, January 2015. CODEN LIOCAH. ISSN 0024-3590.

Austin:2013:ONI

[Aus13] Jay Austin. Observations of near-inertial energy in Lake Superior. Limnology and Oceanography, 58(2):715–728, March 2013. CODEN LIOCAH. ISSN 0024-3590.

Austin:2019:ORD

[Aus19] Jay A. Austin. Observations of radiatively driven convection in a deep lake. Limnology and Oceanography, 66(4):2152–2160, September 2019. CODEN LIOCAH. ISSN 0024-3590.

Almeda:2018:PPM

[AvSGK18] Rodrigo Almeda, Hans van Someren Gréve, and Thomas Kiørboe. Prey perception mechanism determines maximum clearance rates of planktonic copepods. Limnology and Oceanography, 63(6):2695–2707, November 2018. CODEN LIOCAH. ISSN 0024-3590.

Adame:2012:TMC

[AWG^{+}12] Maria Fernanda Adame, Sara F. Wright, Alistair Grinham, Kellie Lobb, Claire E. Reymond, and Catherine E. Lovelock. Terrestrial–marine connectivity: Patterns of terrestrial soil carbon deposition in coastal sediments determined by analysis of glomalin related soil protein. Limnology and Oceanography, 58 (1):1492–1502, September 2012. CODEN LIOCAH. ISSN 0024-3590.

Ahmerkamp:2017:RBO

[AWK^{+}17] Soeren Ahmerkamp, Christian Winter, Knut Krämer, Dirk de Beer, Felix Janssen, Jana Friedrich, Marcel M. M. Kuypers, and Moritz Holtappels. Regulation of benthic oxygen fluxes in permeable sediments of the coastal ocean. Limnology and Oceanography, 62(5):1935–1954, September 2017. CODEN LI-OCAH. ISSN 0024-3590.

Barnes:2014:PCP

[BA14] Morvan Barnes and David Antoine. Proxies of community production derived from the diel variability of particulate attenuation and backscattering coefficients in the northwest Mediterranean Sea. Limnology and Oceanography, 61(1):2133–2149, November 2014. CODEN LIOCAH. ISSN 0024-3590.

Benavides:2013:ENF

[BAA^{+}13] Mar Benavides, Javier Arístegui, Nona S. R. Agawin, José López Cancio, and Santiago Hernàndez-León. Enhancement of nitrogen fixation rates by unicellular diazotrophs vs. Trichodesmium after a dust deposition event in the Canary Islands. Limnology and Oceanography, 58(2):267–275, January 2013. CODEN LIOCAH. ISSN 0024-3590.

Bjaerke:2016:PEI

[BAB^{+}16] Oda Bjærke, Tom Andersen, Kjersti S. Bækkedal, Marius Nordbotten, Lars F. Skau, and Josefin Titelman. Paternal energetic investments in copepods. Limnology and Oceanography, 61(2): 508–517, March 2016. CODEN LIOCAH. ISSN 0024-3590.

Baars:2014:DCS

[BAG^{+}14] Oliver Baars, Wafa Abouchami, Stephen J. G. Galer, Marie Boye, and Peter L. Croot. Dissolved cadmium in the Southern Ocean: Distribution, speciation, and relation to phosphate. Limnology and Oceanography, 59(2):385–399, March 2014. CO-DEN LIOCAH. ISSN 0024-3590.

Blais:2017:CIC

[BAG^{+}17] Marjolaine Blais, Mathieu Ardyna, Michel Gosselin, Dany Dumont, Simon Bélanger, Jean-Éric Tremblay, Yves Gratton, Christian Marchese, and Michel Poulin. Contrasting interannual changes in phytoplankton productivity and community structure in the coastal Canadian Arctic Ocean. Limnology and Oceanography, 62(6):2480–2497, November 2017. CODEN LI-OCAH. ISSN 0024-3590.

Brown:2014:IVS

[BAY^{+}14] T. A. Brown, C. Alexander, D. J. Yurkowski, S. H. Ferguson, and S. T. Belt. Identifying variable sea ice carbon contributions to the Arctic ecosystem: a case study using highly branched isoprenoid lipid biomarkers in Cumberland Sound ringed seals. Limnology and Oceanography, 59(5):1581–1589, September 2014. CODEN LIOCAH. ISSN 0024-3590.

Becker:2010:LLE

[BB10] Claes Becker and Maarten Boersma. Limiting levels of eicosapentaenoic acid: What do we really know? Limnology and Oceanography, 55(1):459–462, January 2010. CODEN LIOCAH. ISSN 0024-3590.

Beier:2011:UCA

[BB11] Sara Beier and Stefan Bertilsson. Uncoupling of chitinase activity and uptake of hydrolysis products in freshwater bacterioplankton. Limnology and Oceanography, 56(4):1179–1188, July 2011. CODEN LIOCAH. ISSN 0024-3590.

Bundy:2014:DPD

[BBB^{+}14] Randelle M. Bundy, Dondra V. Biller, Kristen N. Buck, Kenneth W. Bruland, and Katherine A. Barbeau. Distinct pools of dissolved iron-binding ligands in the surface and benthic boundary layer of the California Current. Limnology and Oceanography, 59(3):769–787, May 2014. CODEN LIOCAH. ISSN 0024-3590.

Branstrator:2017:CDI

[BBB^{+}17] Donn K. Branstrator, Ashley E. Beranek, Meghan E. Brown, Leif K. Hembre, and Daniel R. Engstrom. Colonization dynamics of the invasive predatory cladoceran, Bythotrephes longimanus, inferred from sediment records. Limnology and Oceanography, 63(3):1096–1110, May 2017. CODEN LIOCAH. ISSN 0024-3590.

Bertilsson:2013:UIM

[BBC^{+}13] Stefan Bertilsson, Amy Burgin, Cayelan C. Carey, Samuel B. Fey, Hans-Peter Grossart, Lorena M. Grubisic, Ian D. Jones, Georgiy Kirillin, Jay T. Lennon, Ashley Shade, and Robyn L. Smyth. The under-ice microbiome of seasonally frozen lakes. Limnology and Oceanography, 58(6):1998–2012, November 2013. CODEN LIOCAH. ISSN 0024-3590.

Briski:2013:TVS

[BBCM^{+}13] Elizabeta Briski, Sarah A. Bailey, Oscar Casas-Monroy, Claudio DiBacco, Irena Kaczmarska, Janice E. Lawrence, Jonas Leichsenring, Colin Levings, Michael L. MacGillivary, Christopher W. McKindsey, Leslie E. Nasmith, Marie Parenteau, Grace E. Piercey, Richard B. Rivkin, Andre Rochon, Suzanne Roy, Nathalie Simard, Bei Sun, Candice Way, Andrea M. Weise, and Hugh J. MacIsaac. Taxon- and vector-specific variation in species richness and abundance during the transport stage of biological invasions. Limnology and Oceanography, 58(4):1361– 1372, July 2013. CODEN LIOCAH. ISSN 0024-3590.

Brigham:2019:AIC

[BBJ^{+}19] Brian A. Brigham, Jeffrey A. Bird, Andrew R. Juhl, Christopher J. Zappa, Angel D. Montero, and Gregory D. O’Mullan. Anthropogenic inputs from a coastal megacity are linked to greenhouse gas concentrations in the surrounding estuary. Limnology and Oceanography, 64(6):2497–2511, November 2019. CODEN LIOCAH. ISSN 0024-3590.

Brahney:2015:ECT

[BBK^{+}15] J. Brahney, A. P. Ballantyne, P. Kociolek, P. R. Leavitt, G. L. Farmer, and J. C. Neff. Ecological changes in two contrasting lakes associated with human activity and dust transport in western Wyoming. Limnology and Oceanography, 60(2):678– 695, March 2015. CODEN LIOCAH. ISSN 0024-3590.

Bass:2011:FDD

[BBLN11] Adrian M. Bass, Michael I. Bird, Michael J. Liddell, and Paul N. Nelson. Fluvial dynamics of dissolved and particulate organic carbon during periodic discharge events in a steep tropical rainforest catchment. Limnology and Oceanography, 56(6):2282– 2292, November 2011. CODEN LIOCAH. ISSN 0024-3590.

Briski:2011:ITD

[BBM11] Elizabeta Briski, Sarah A. Bailey, and Hugh J. MacIsaac. Invertebrates and their dormant eggs transported in ballast sediments of ships arriving to the Canadian coasts and the Laurentian Great Lakes. Limnology and Oceanography, 56(5):1929–1939, September 2011. CODEN LIOCAH. ISSN 0024-3590.

Benoit-Bird:2017:POS

[BBMS17] Kelly J. Benoit-Bird, Mark A. Moline, and Brandon L. Southall. Prey in oceanic sound scattering layers organize to get a little help from their friends. Limnology and Oceanography, 62(6): 2788–2798, November 2017. CODEN LIOCAH. ISSN 0024-

Brucet:2010:FIZ

[BBQ^{+}10] Sandra Brucet, Dani Boix, Xavier D. Quintana, Elisabeth Jensen, Louise W. Nathansen, Carolina Trochine, Mariana Meerhoff, Stáphanie Gascón, and Erik Jeppesen. Factors influencing zooplankton size structure at contrasting temperatures in coastal shallow lakes: Implications for effects of climate change. Limnology and Oceanography, 55(4):1697–1711, July 2010. CO-DEN LIOCAH. ISSN 0024-3590.

Bouffard:2012:PWI

[BBR12] Damien Bouffard, Leon Boegman, and Yerubandi R. Rao. Poincaré wave–induced mixing in a large lake. Limnology and Oceanography, 57(4):1201–1216, July 2012. CODEN LIOCAH. ISSN 0024-3590.

Borsje:2014:FEB

[BBR^{+}14] Bas W. Borsje, Tjeerd J. Bouma, Marijn Rabaut, Peter M. J. Herman, and Suzanne J. M. H. Hulscher. Formation and erosion of biogeomorphological structures: a model study on the tubebuilding polychaete Lanice conchilega. Limnology and Oceanography, 59(4):1297–1309, July 2014. CODEN LIOCAH. ISSN 0024-3590.

Brown:2012:PRS

[BBS12] Meghan E. Brown, Donn K. Branstrator, and Lyle J. Shannon. Population regulation of the spiny water flea (Bythotrephes longimanus) in a reservoir: Implications for invasion. Limnology and Oceanography, 57(1):251–271, January 2012. CODEN LIOCAH. ISSN 0024-3590.

Barry:2018:CBC

[BBS^{+}18] Savanna C. Barry, Thomas S. Bianchi, Michael R. Shields, Jack A. Hutchings, Charles A. Jacoby, and Thomas K. Frazer. Characterizing blue carbon stocks in Thalassia testudinum meadows subjected to different phosphorus supplies: a lignin biomarker approach. Limnology and Oceanography, 63(6):2630– 2646, November 2018. CODEN LIOCAH. ISSN 0024-3590.

Berggren:2018:TSZ

[BBSK18] M. Berggren, P. Bengtson, A. R. A. Soares, and J. Karlsson. Terrestrial support of zooplankton biomass in northern rivers. Limnology and Oceanography, 63(6):2479–2492, November 2018. CODEN LIOCAH. ISSN 0024-3590.

Baden:2010:RIT

[BBT^{+}10] Susanne Baden, Christoffer Boström, Stefan Tobiasson, Heidi Arponen, and Per-Olav Moksnes. Relative importance of trophic interactions and nutrient enrichment in seagrass ecosystems: a broad-scale field experiment in the Baltic–Skagerrak area. Limnology and Oceanography, 55(3):1435–1448, May 2010. CODEN LIOCAH. ISSN 0024-3590.

Bonnet:2016:DDN

[BBTK^{+}16] Sophie Bonnet, Hugo Berthelot, Kendra Turk-Kubo, Véronique Cornet-Barthaux, Sarah Fawcett, Ilana Berman-Frank, Aude Barani, Gérald Grégori, Julien Dekaezemacker, Mar Benavides, and Douglas G. Capone. Diazotroph derived nitrogen supports diatom growth in the South West Pacific: a quantitative study using nanoSIMS. Limnology and Oceanography, 61(5):1549– 1562, September 2016. CODEN LIOCAH. ISSN 0024-3590.

Buchwald:2010:OIF

[BC10] Carolyn Buchwald and Karen L. Casciotti. Oxygen isotopic fractionation and exchange during bacterial nitrite oxidation. Limnology and Oceanography, 55(3):1064–1074, May 2010. CODEN LIOCAH. ISSN 0024-3590.

Becker:2019:CUS

[BC19] Amani E. Becker and David Copplestone. Cadmium uptake from sediment by Cylindrotheca closterium and the effect of diatom presence on partitioning of cadmium between sediment and water: a laboratory study. Limnology and Oceanography, 64 (6):2550–2568, November 2019. CODEN LIOCAH. ISSN 0024-3590.

Batt:2012:RSF

[BCC^{+}12] Ryan D. Batt, Stephen R. Carpenter, Jonathan J. Cole, Michael L. Pace, Timothy J. Cline, Robert A. Johnson, and David A. Seekell. Resources supporting the food web of a naturally productive lake. Limnology and Oceanography, 58(1):1443– 1452, September 2012. CODEN LIOCAH. ISSN 0024-3590.

Barbut:2019:HLT

[BCDR^{+}19] Léo Barbut, Clara Groot Crego, Sophie Delerue-Ricard, Sara Vandamme, Filip A. M. Volckaert, and Geneviève Lacroix. How larval traits of six flatfish species impact connectivity. Limnology and Oceanography, 64(3):1150–1171, May 2019. CODEN LIOCAH. ISSN 0024-3590.

Brown:2017:CCB

[BCF^{+}17] T. A. Brown, E. Chrystal, S. H. Ferguson, D. J. Yurkowski, C. Watt, N. E. Hussey, T. C. Kelley, and S. T. Belt. Coupled changes between the H-Print biomarker and δ ^{15}N indicates a variable sea ice carbon contribution to the diet of Cumberland Sound beluga whales. Limnology and Oceanography, 62(4):1606– 1619, July 2017. CODEN LIOCAH. ISSN 0024-3590.

Buttay:2017:EMS

[BCM^{+}17] Lucie Buttay, Bernard Cazelles, Ana Miranda, Gerardo Casas, Enrique Nogueira, and Rafael González-Quirós. Environmental multi-scale effects on zooplankton inter-specific synchrony. Limnology and Oceanography, 62(4):1355–1365, July 2017. CODEN LIOCAH. ISSN 0024-3590.

Berube:2016:TDP

[BCRC16] Paul M. Berube, Allison Coe, Sara E. Roggensack, and Sallie W. Chisholm. Temporal dynamics of Prochlorococcus cells with the potential for nitrate assimilation in the subtropical Atlantic and Pacific Oceans. Limnology and Oceanography, 61(2):482–495, March 2016. CODEN LIOCAH. ISSN 0024-3590.

Bruesewitz:2015:UIT

[BCRW15] Denise A. Bruesewitz, Cayelan C. Carey, David C. Richardson, and Kathleen C. Weathers. Under-ice thermal stratification dynamics of a large, deep lake revealed by high-frequency data. Limnology and Oceanography, 60(2):347–359, March 2015. CO-DEN LIOCAH. ISSN 0024-3590.

Bullejos:2010:RPU

[BCVAn10] Francisco José Bullejos, Presentación Carrillo, Manuel Villar-Argaiz, and Juan Manuel Medina-Sá ncheza. Roles of phosphorus and ultraviolet radiation in the strength of phytoplanktonzooplankton coupling in a Mediterranean high mountain lake. Limnology and Oceanography, 55(6):2549–2562, November 2010. CODEN LIOCAH. ISSN 0024-3590.

Burris:2015:FEB

[BD15] Z. P. Burris and H. G. Dam. First evidence of biased sex ratio at birth in a calanoid copepod. Limnology and Oceanography, 60 (2):722–731, March 2015. CODEN LIOCAH. ISSN 0024-3590.

Balch:2014:SBC

[BDB^{+}14] W. M. Balch, D. T. Drapeau, B. C. Bowler, E. R. Lyczkowski, L. C. Lubelczyk, S. C. Painter, and A. J. Poulton. Surface biological, chemical, and optical properties of the Patagonian Shelf coccolithophore bloom, the brightest waters of the Great Calcite Belt. Limnology and Oceanography, 59(5):1715–1732, September 2014. CODEN LIOCAH. ISSN 0024-3590.

Briski:2014:VPC

[BDC^{+}14] Elizabeta Briski, D. Andrew R. Drake, Farrah T. Chan, Sarah A. Bailey, and Hugh J. MacIsaac. Variation in propagule and colonization pressures following rapid human-mediated transport: Implications for a universal assemblage-based management model. Limnology and Oceanography, 59(6):2068–2076, November 2014. CODEN LIOCAH. ISSN 0024-3590.

Bottjer:2017:TVN

[BDK^{+}17] Daniela Böttjer, John E. Dore, David M. Karl, Ricardo M. Letelier, Claire Mahaffey, Samuel T. Wilson, Jonathan Zehr, and Matthew J. Church. Temporal variability of nitrogen fixation and particulate nitrogen export at Station ALOHA. Limnology and Oceanography, 62(1):200–216, January 2017. CODEN LIOCAH. ISSN 0024-3590.

Berg:2019:DBM

[BDP^{+}19] Peter Berg, Marie Lise Delgard, Pierre Polsenaere, Karen J. McGlathery, Scott C. Doney, and Amelie C. Berger. Dynamics of benthic metabolism, O_{2}, and pCO_{2} in a temperate seagrass meadow. Limnology and Oceanography, 64(6):2586–2604, November 2019. CODEN LIOCAH. ISSN 0024-3590.

Behl:2011:RIS

[BDS11] Stephan Behl, Anne Donval, and Herwig Stibor. The relative importance of species diversity and functional group diversity on carbon uptake in phytoplankton communities. Limnology and Oceanography, 56(2):683–694, March 2011. CODEN LIOCAH. ISSN 0024-3590.

Blaszczak:2019:SSU

[BDU^{+}19] Joanna R. Blaszczak, Joseph M. Delesantro, Dean L. Urban, Martin W. Doyle, and Emily S. Bernhardt. Scoured or suffocated: Urban stream ecosystems oscillate between hydrologic and dissolved oxygen extremes. Limnology and Oceanography, 64(3):877–894, May 2019. CODEN LIOCAH. ISSN 0024-3590.

Bergamaschi:2011:MMD

[BFD^{+}11] B. A. Bergamaschi, J. A. Fleck, B. D. Downing, E. Boss, B. Pellerin, N. K. Ganju, D. H. Schoellhamer, A. A. Byington, W. A. Heim, M. Stephenson, and R. Fujii. Methyl mercury dynamics in a tidal wetland quantified using in situ optical measurements. Limnology and Oceanography, 56(4):1355–1371, July 2011. CO-DEN LIOCAH. ISSN 0024-3590.

Barton:2013:RCS

[BFW^{+}13] Andrew D. Barton, Zoe V. Finkel, Ben A. Ward, David G. Johns, and Michael J. Follows. On the roles of cell size and trophic strategy in North Atlantic diatom and dinoflagellate communities. Limnology and Oceanography, 58(2):254–266, January 2013. CODEN LIOCAH. ISSN 0024-3590.

Borgesa:2010:CCC

[BG10a] Alberto V. Borgesa and Nathalie Gypensb. Carbonate chemistry in the coastal zone responds more strongly to eutrophication than ocean acidification. Limnology and Oceanography, 55(1): 346–353, January 2010. CODEN LIOCAH. ISSN 0024-3590.

Buitenhuis:2010:MPA

[BG10b] Erik T. Buitenhuis and Richard J. Geider. A model of phytoplankton acclimation to iron–light colimitation. Limnology and Oceanography, 55(2):714–724, March 2010. CODEN LIOCAH. ISSN 0024-3590.

Brooks:2014:SIE

[BGB^{+}14] J. Renée Brooks, John J. Gibson, S. Jean Birks, Marc H. Weber, Kent D. Rodecap, and John L. Stoddard. Stable isotope estimates of evaporation : inflow and water residence time for lakes across the United States as a tool for national lake water quality assessments. Limnology and Oceanography, 61(1):2150–2165, November 2014. CODEN LIOCAH. ISSN 0024-3590.

Bruesewitz:2013:EEF

[BGM^{+}13] Denise A. Bruesewitz, Wayne S. Gardner, Rae F. Mooney, Lindsey Pollard, and Edward J. Buskey. Estuarine ecosystem function response to flood and drought in a shallow, semiarid estuary: Nitrogen cycling and ecosystem metabolism. Limnology and Oceanography, 58(6):2293–2309, November 2013. CODEN LIOCAH. ISSN 0024-3590.

Byers:2015:GVI

[BGP^{+}15] James E. Byers, Jonathan H. Grabowski, Michael F. Piehler, A. Randall Hughes, Heidi W. Weiskel, Jennafer C. Malek, and David L. Kimbro. Geographic variation in intertidal oyster reef properties and the influence of tidal prism. Limnology and Oceanography, 60(3):1051–1063, May 2015. CODEN LIOCAH. ISSN 0024-3590.

Brin:2014:ECA

[BGR14] Lindsay D. Brin, Anne E. Giblin, and Jeremy J. Rich. Environmental controls of anammox and denitrification in southern New England estuarine and shelf sediments. Limnology and Oceanography, 59(3):851–860, May 2014. CODEN LIOCAH. ISSN 0024-3590.

Binding:2015:LTW

[BGW^{+}15] Caren E. Binding, Tracie A. Greenberg, Sue B. Watson, Shannah Rastin, and Jessica Gould. Long term water clarity changes in North America’s Great Lakes from multi-sensor satellite observations. Limnology and Oceanography, 60(6):1976–1995, November 2015. CODEN LIOCAH. ISSN 0024-3590.

BarreiroFelpeto:2013:IBC

[BH13] Aldo Barreiro Felpeto and Nelson G. Hairston, Jr. Indirect bottom-up control of consumer-resource dynamics: Resourcedriven algal quality alters grazer numerical response. Limnology and Oceanography, 58(3):827–838, May 2013. CODEN LI-OCAH. ISSN 0024-3590.

Beyer:2016:PDE

[BH16] Jessica E. Beyer and K. David Hambright. Persistent and delayed effects of toxic cyanobacteria exposure on life history traits of a common zooplankter. Limnology and Oceanography, 61(2): 587–595, March 2016. CODEN LIOCAH. ISSN 0024-3590.

Brandsma:2012:SDI

[BHB^{+}12] Joost Brandsma, Ellen C. Hopmans, Corina P. D. Brussaard, Harry J. Witte, Stefan Schouten, and Jaap S. Sinninghe Damsté. Spatial distribution of intact polar lipids in North Sea surface waters: Relationship with environmental conditions and microbial community composition. Limnology and Oceanography, 57 (4):959–973, July 2012. CODEN LIOCAH. ISSN 0024-3590.

Barrett:2019:CAC

[BHB^{+}19] P. M. Barrett, E. A. Hull, K. Burkart, O. Hargrave, J. McLean, V. F. Taylor, B. P. Jackson, J. E. Gawel, and R. B. Neumann. Contrasting arsenic cycling in strongly and weakly stratified contaminated lakes: Evidence for temperature control on sediment– water arsenic fluxes. Limnology and Oceanography, 64(3):1333– 1346, May 2019. CODEN LIOCAH. ISSN 0024-3590.

Bachmann:2013:ENL

[BHC13] Roger W. Bachmann, Mark V. Hoyer, and Daniel E. Canfield, Jr. The extent that natural lakes in the United States of America have been changed by cultural eutrophication. Limnology and Oceanography, 58(3):945–950, May 2013. CODEN LIOCAH. ISSN 0024-3590. See comment [SDH^{+}14].

Bachmann:2014:RCQ

[BHC14] Roger W. Bachmann, Mark V. Hoyer, and Daniel E. Canfield, Jr. Response to comments: Quantification of the extent of cultural eutrophication of natural lakes in the United States. Limnology and Oceanography, 61(1):2231–2239, November 2014. CODEN LIOCAH. ISSN 0024-3590.

Bercovici:2017:DOC

[BHD^{+}17] Sarah K. Bercovici, Bruce A. Huber, Hans B. DeJong, Robert B. Dunbar, and Dennis A. Hansell. Dissolved organic carbon in the Ross Sea: Deep enrichment and export. Limnology and Oceanography, 62(6):2593–2603, November 2017. CODEN LI-OCAH. ISSN 0024-3590.

Bernhardt:2018:MRF

[BHG^{+}18] E. S. Bernhardt, J. B. Heffernan, N. B. Grimm, E. H. Stanley, J. W. Harvey, M. Arroita, A. P. Appling, M. J. Cohen, W. H. McDowell, R. O. Hall, Jr., J. S. Read, B. J. Roberts, E. G. Stets, and C. B. Yackulic. The metabolic regimes of flowing waters. Limnology and Oceanography, 63(S1):S99–S118, March 2018. CODEN LIOCAH. ISSN 0024-3590.

Bruesewitz:2017:WIN

[BHM^{+}17] Denise A. Bruesewitz, Timothy J. Hoellein, Rae F. Mooney, Wayne S. Gardner, and Edward J. Buskey. Wastewater influences nitrogen dynamics in a coastal catchment during a prolonged drought. Limnology and Oceanography, 62(S1):S239– S257, November 2017. CODEN LIOCAH. ISSN 0024-3590.

Bale:2016:ITS

[BHS^{+}16] Nicole J. Bale, Ellen C. Hopmans, Petra L. Schoon, Anna de Kluijver, John A. Downing, Jack J. Middelburg, Jaap S. Sinninghe Damsté, and Stefan Schouten. Impact of trophic state on the distribution of intact polar lipids in surface waters of lakes. Limnology and Oceanography, 61(3):1065–1077, May 2016. CO-DEN LIOCAH. ISSN 0024-3590.

Blommaert:2017:CND

[BHV^{+}17] Lander Blommaert, Marie J. J. Huysman, Wim Vyverman, Johann Lavaud, and Koen Sabbe. Contrasting NPQ dynamics and xanthophyll cycling in a motile and a non-motile intertidal benthic diatom. Limnology and Oceanography, 62(4):1466–1479, July 2017. CODEN LIOCAH. ISSN 0024-3590.

Barton:2012:POC

[BHW^{+}12] Alan Barton, Burke Hales, George G. Waldbusser, Chris Langdon, and Richard A. Feely. The Pacific oyster, Crassostrea gigas, shows negative correlation to naturally elevated carbon dioxide levels: Implications for near-term ocean acidification effects. Limnology and Oceanography, 57(3):698–710, May 2012. CO-DEN LIOCAH. ISSN 0024-3590.

Belcher:2016:RPA

[BIM^{+}16] Anna Belcher, Morten Iversen, Clara Manno, Stephanie A. Henson, Geraint A. Tarling, and Richard Sanders. The role of particle associated microbes in remineralization of fecal pellets in the upper mesopelagic of the Scotia Sea, Antarctica. Limnology and Oceanography, 61(3):1049–1064, May 2016. CODEN LIOCAH. ISSN 0024-3590.

Boyd:2010:RUO

[BIS^{+}10] P. W. Boyd, E. Ibisanmi, S. G. Sander, K. A. Hunter, and G. A. Jackson. Remineralization of upper ocean particles: Implications for iron biogeochemistry. Limnology and Oceanography, 55 (3):1271–1288, May 2010. CODEN LIOCAH. ISSN 0024-3590.

Bi:2017:EDC

[BISZ17] Rong Bi, Stefanie Ismar, Ulrich Sommer, and Meixun Zhao. Environmental dependence of the correlations between stoichiometric and fatty acid-based indicators of phytoplankton nutritional quality. Limnology and Oceanography, 62(1):334–347, January 2017. CODEN LIOCAH. ISSN 0024-3590.

Beyene:2015:WWC

[BJ15] Mussie T. Beyene and Shaleen Jain. Wintertime weatherclimate variability and its links to early spring ice-out in Maine lakes. Limnology and Oceanography, 60(6):1890–1905, November 2015. CODEN LIOCAH. ISSN 0024-3590.

Baker:2010:SFC

[BJDMH10] David M. Baker, Eric Jordán-Dahlgren, Miguel Angel Maldonado, and C. Drew Harvell. Sea fan corals provide a stable isotope baseline for assessing sewage pollution in the Mexican Caribbean. Limnology and Oceanography, 55(5):2139–2149, September 2010. CODEN LIOCAH. ISSN 0024-3590.

Barry:2018:RSI

[BJF18] Savanna C. Barry, Charles A. Jacoby, and Thomas K. Frazer. Resilience to shading influenced by differential allocation of biomass in Thalassia testudinum. Limnology and Oceanography, 63(4):1817–1831, July 2018. CODEN LIOCAH. ISSN 0024-3590.

Burdige:2011:AOM

[BK11] David J. Burdige and Tomoko Komada. Anaerobic oxidation of methane and the stoichiometry of remineralization processes in continental margin sediments. Limnology and Oceanography, 56(5):1781–1796, September 2011. CODEN LIOCAH. ISSN 0024-3590.

Bernhardt:2013:SPR

[BK13] Juliane Bernhardt and Georgiy Kirillin. Seasonal pattern of rotation-affected internal seiches in a small temperate lake. Limnology and Oceanography, 58(4):1344–1360, July 2013. CODEN LIOCAH. ISSN 0024-3590.

Brothers:2014:FLL

[BKA^{+}14] S. Brothers, J. Köhler, K. Attermeyer, H. P. Grossart, T. Mehner, N. Meyer, K. Scharnweber, and S. Hilt. A feedback loop links brownification and anoxia in a temperate, shallow lake. Limnology and Oceanography, 59(4):1388–1398, July 2014. CODEN LIOCAH. ISSN 0024-3590.

Bonaglia:2016:DDB

[BKD^{+}16] Stefano Bonaglia, Isabell Klawonn, Loreto De Brabandere, Barbara Deutsch, Bo Thamdrup, and Volker Brüchert. Denitrification and DNRA at the Baltic Sea oxic–anoxic interface: Substrate spectrum and kinetics. Limnology and Oceanography, 62 (3):1900–1915, September 2016. CODEN LIOCAH. ISSN 0024-3590.

Beisner:2013:SOL

[BL13] Beatrix E. Beisner and Maria Lorena Longhi. Spatial overlap in lake phytoplankton: Relations with environmental factors and consequences for diversity. Limnology and Oceanography, 58(4): 1419–1430, July 2013. CODEN LIOCAH. ISSN 0024-3590.

Brenner:2018:SMM

[BL18] S. D. Brenner and B. E. Laval. Seiche modes in multi-armed lakes. Limnology and Oceanography, 63(6):2717–2726, November 2018. CODEN LIOCAH. ISSN 0024-3590.

Brand:2015:MBI

[BLG^{+}15] Andreas Brand, Jessica R. Lacy, Steve Gladding, Rusty Holleman, and Mark Stacey. Model-based interpretation of sediment concentration and vertical flux measurements in a shallow estuarine environment. Limnology and Oceanography, 60(2):463– 481, March 2015. CODEN LIOCAH. ISSN 0024-3590.

Berg:2013:ECM

[BLH^{+}13] Peter Berg, Matthew H. Long, Markus Huettel, Jennie E. Rheuban, Karen J. McGlathery, Robert W. Howarth, Kenneth H. Foreman, Anne E. Giblin, and Roxanne Marino. Eddy correlation measurements of oxygen fluxes in permeable sediments exposed to varying current flow and light. Limnology and Oceanography, 58(4):1329–1343, July 2013. CODEN LIOCAH. ISSN 0024-3590.

Bunn:2013:DTF

[BLJ13] Stuart E. Bunn, Catherine Leigh, and Timothy D. Jardine. Diettissue fractionation of δ ^{15}N by consumers from streams and rivers. Limnology and Oceanography, 58(3):765–773, May 2013. CODEN LIOCAH. ISSN 0024-3590.

Brown:2012:TMD

[BLLB12] Matthew T. Brown, Sherry M. Lippiatt, Maeve C. Lohan, and Kenneth W. Bruland. Trace metal distributions within a Sitka eddy in the northern Gulf of Alaska. Limnology and Oceanography, 57(3):503–518, March 2012. CODEN LIOCAH. ISSN 0024-3590.

Bryant:2010:VSO

[BLM^{+}10] Lee D. Bryant, Claudia Lorrai, Daniel F. McGinnis, Andreas Brand, Alfred Wü est, and John C. Little. Variable sediment oxygen uptake in response to dynamic forcing. Limnology and Oceanography, 55(4):950–964, March 2010. CODEN LIOCAH. ISSN 0024-3590.

Burrows:2017:SRL

[BLMS17] Ryan M. Burrows, Hjalmar Laudon, Brendan G. McKie, and Ryan A. Sponseller. Seasonal resource limitation of heterotrophic biofilms in boreal streams. Limnology and Oceanography, 62(1):164–176, January 2017. CODEN LIOCAH. ISSN 0024-3590.

Bunting:2016:IVS

[BLS^{+}16] L. Bunting, P. R. Leavitt, G. L. Simpson, B. Wissel, K. R. Laird, B. F. Cumming, A. St.Amand, and D. R. Engstrom. Increased variability and sudden ecosystem state change in Lake Winnipeg, Canada, caused by 20^{th} century agriculture. Limnology and Oceanography, 61(6):2090–2107, November 2016. CO-DEN LIOCAH. ISSN 0024-3590.

Barton:2015:PCV

[BLW15] Andrew D. Barton, M. Susan Lozier, and Richard G. Williams. Physical controls of variability in North Atlantic phytoplankton communities. Limnology and Oceanography, 60(1):181–197, January 2015. CODEN LIOCAH. ISSN 0024-3590.

Bunting:2010:RNB

[BLWV10] Lynda Bunting, Peter R. Leavitt, R. Paul Weidman, and Rolf D. Vinebrooke. Regulation of the nitrogen biogeochemistry of mountain lakes by subsidies of terrestrial dissolved organic matter and the implications for climate studies. Limnology and Oceanography, 55(1):333–345, January 2010. CODEN LIOCAH. ISSN 0024-3590.

Bianchi:2016:GPD

[BM16] Daniele Bianchi and K. A. S. Mislan. Global patterns of diel vertical migration times and velocities from acoustic data. Limnology and Oceanography, 61(1):353–364, January 2016. CODEN LIOCAH. ISSN 0024-3590.

Burdorf:2018:EOA

[BMB^{+}18] Laurine D. W. Burdorf, Sairah Y. Malkin, Jesper T. Bjerg, Pieter van Rijswijk, Francis Criens, Anton Tramper, and Filip J. R. Meysman. The effect of oxygen availability on longdistance electron transport in marine sediments. Limnology and Oceanography, 63(4):1799–1816, July 2018. CODEN LIOCAH. ISSN 0024-3590.

Bechtold:2012:EPO

[BMBI12] Heather A. Bechtold, Amy M. Marcarelli, Colden V. Baxter, and Richard S. Inouye. Effects of N, P, and organic carbon on stream biofilm nutrient limitation and uptake in a semiarid watershed. Limnology and Oceanography, 58(1):1544–1554, September 2012. CODEN LIOCAH. ISSN 0024-3590.

Bernard:2016:SSM

[BMC^{+}16] I. Bernard, J.-C. Massabuau, P. Ciret, M. Sow, A. Sottolichio, S. Pouvreau, and D. Tran. In situ spawning in a marine broadcast spawner, the Pacific oyster Crassostrea gigas: Timing and environmental triggers. Limnology and Oceanography, 61(2): 635–647, March 2016. CODEN LIOCAH. ISSN 0024-3590.

Bordeyne:2017:VFC

[BMD17] François Bordeyne, Aline Migné, and Dominique Davoult. Variation of fucoid community metabolism during the tidal cycle: Insights from in situ measurements of seasonal carbon fluxes during emersion and immersion. Limnology and Oceanography, 62 (6):2418–2430, November 2017. CODEN LIOCAH. ISSN 0024-3590.

Bricelj:2010:ESR

[BMDC10] V. M. Bricelj, S. P. MacQuarrie, J. A. E. Doane, and L. B. Connell. Evidence of selection for resistance to paralytic shellfish toxins during the early life history of soft-shell clam, Mya arenaria, populations. Limnology and Oceanography, 55(6):2463– 2475, November 2010. CODEN LIOCAH. ISSN 0024-3590.

Barbosa:2016:DMF

[BMF^{+}16] Pedro M. Barbosa, John M. Melack, Vinicius F. Farjalla, João Henrique F. Amaral, Vinicius Scofield, and Bruce R. Forsberg. Diffusive methane fluxes from Negro, Solimões and Madeira rivers and fringing lakes in the Amazon basin. Limnology and Oceanography, 61(S1):S221–S237, November 2016. CODEN LI-OCAH. ISSN 0024-3590.

Bertrand:2013:MSI

[BMM^{+}13] Erin M. Bertrand, Dawn M. Moran, Matthew R. McIlvin, Jeffrey M. Hoffman, Andrew E. Allen, and Mak A. Saito. Methionine synthase interreplacement in diatom cultures and communities: Implications for the persistence of B_{12} use by eukaryotic phytoplankton. Limnology and Oceanography, 58(4):1431–1450, July 2013. CODEN LIOCAH. ISSN 0024-3590.

Beaulieu:2016:ERM

[BMN16] Jake J. Beaulieu, Michael G. McManus, and Christopher T. Nietch. Estimates of reservoir methane emissions based on a spatially balanced probabilistic-survey. Limnology and Oceanography, 61(S1):S27–S40, November 2016. CODEN LIOCAH. ISSN 0024-3590.

Bell:2019:FFE

[BMPF19] Alex T. C. Bell, Dennis L. Murray, Clay Prater, and Paul C. Frost. Fear and food: Effects of predator-derived chemical cues and stoichiometric food quality on Daphnia. Limnology and Oceanography, 64(4):1706–1715, July 2019. CODEN LIOCAH. ISSN 0024-3590.

Ballantyne:2010:DBP

[BMW10] Ford Iv Ballantyne, Duncan N. L. Menge, and Joshua S. Weitz. A discrepancy between predictions of saturating nutrient uptake models and nitrogen-to-phosphorus stoichiometry in the surface ocean. Limnology and Oceanography, 55(3):997–1008, May 2010. CODEN LIOCAH. ISSN 0024-3590.

Blees:2014:BMD

[BNW^{+}14a] Jan Blees, Helge Niemann, Christine B. Wenk, Jakob Zopfi, Carsten J. Schubert, Joël S. Jenzer, Mauro Veronesi, and Moritz F. Lehmann. Bacterial methanotrophs drive the formation of a seasonal anoxic benthic nepheloid layer in an alpine lake. Limnology and Oceanography, 59(4):1410–1420, July 2014. CODEN LIOCAH. ISSN 0024-3590.

Blees:2014:MAB

[BNW^{+}14b] Jan Blees, Helge Niemann, Christine B. Wenk, Jakob Zopfi, Carsten J. Schubert, Mathias K. Kirf, Mauro L. Veronesi, Carmen Hitz, and Moritz F. Lehmann. Micro-aerobic bacterial methane oxidation in the chemocline and anoxic water column of deep south-alpine Lake Lugano (Switzerland). Limnology and Oceanography, 59(2):311–324, March 2014. CODEN LIOCAH. ISSN 0024-3590.

Botnen:2015:ESC

[BOT^{+}15] Helle Augdal Botnen, Abdirahman M. Omar, Ingunn Thorseth, Truls Johannessen, and Guttorm Alendal. The effect of submarine CO_{2} vents on seawater: Implications for detection of subsea carbon sequestration leakage. Limnology and Oceanography, 60 (2):402–410, March 2015. CODEN LIOCAH. ISSN 0024-3590.

Beman:2012:QAO

[BPA12] J. Michael Beman, Brian N. Popp, and Susan E. Alford. Quantification of ammonia oxidation rates and ammonia-oxidizing archaea and bacteria at high resolution in the Gulf of California and eastern tropical North Pacific Ocean. Limnology and Oceanography, 57(3):711–726, May 2012. CODEN LIOCAH. ISSN 0024-3590.

Bailleul:2017:DML

[BPB^{+}17] Benjamin Bailleul, Jisoo Park, Christopher M. Brown, Kay D. Bidle, Sang Hoon Lee, and Paul G. Falkowski. Direct measurements of the light dependence of gross photosynthesis and oxygen consumption in the ocean. Limnology and Oceanography, 63 (3):1066–1079, May 2017. CODEN LIOCAH. ISSN 0024-3590.

Beaulieu:2013:NWT

[BPGE13] Marieke Beaulieu, Frances Pick, and Irene Gregory-Eaves. Nutrients and water temperature are significant predictors of cyanobacterial biomass in a 1147 lakes data set. Limnology and Oceanography, 58(5):1736–1746, September 2013. CODEN LI-OCAH. ISSN 0024-3590.

Belovsky:2019:OSC

[BPL^{+}19a] Gary E. Belovsky, Clay Perschon, Chad Larson, Chad Mellison, Jennifer Slade, Heidi Mahon, Hannah Appiah-Madson, John Luft, Ryan Mosley, John Neill, Kyle Stone, Ashley Kijowski, and James Van Leeuwen. Overwinter survival of crustacean diapausing cysts: Brine shrimp (Artemia franciscana) in Great Salt Lake, Utah. Limnology and Oceanography, 64(6): 2538–2549, November 2019. CODEN LIOCAH. ISSN 0024-3590.

Bretherton:2019:DLK

[BPL^{+}19b] Laura Bretherton, Alex J. Poulton, Tracy Lawson, Nita Rukminasari, Cecilia Balestreri, Declan Schroeder, C. Mark Moore, and David J. Suggett. Day length as a key factor moderating the response of coccolithophore growth to elevated pCO_{2}. Limnology and Oceanography, 64(3):1284–1296, May 2019. CODEN LIOCAH. ISSN 0024-3590.

Bi:2012:CAC

[BPPF12] Hongsheng Bi, William T. Peterson, Jay O. Peterson, and Jennifer L. Fisher. A comparative analysis of coastal and shelf-slope copepod communities in the northern California Current system: Synchronized response to large-scale forcing? Limnology and Oceanography, 58(1):1467–1478, September 2012. CODEN LIOCAH. ISSN 0024-3590.

Biester:2018:SIP

[BPRG^{+}18] Harald Biester, Marta Pérez-Rodríguez, Benjamin-Silas Gilfedder, Antonio Martínez Cortizas, and Yvonne-Marie Hermanns. Solar irradiance and primary productivity controlled mercury accumulation in sediments of a remote lake in the Southern Hemisphere during the past 4000 years. Limnology and Oceanography, 63(2):540–549, March 2018. CODEN LIOCAH. ISSN 0024-3590.

Braeckman:2019:DMD

[BPV^{+}19] U. Braeckman, F. Pasotti, S. Vázquez, K. Zacher, R. Hoffmann, M. Elvert, H. Marchant, C. Buckner, M. L. Quartino, W. Mác Cormack, K. Soetaert, F. Wenzhöfer, and A. Vanreusel. Degradation of macroalgal detritus in shallow coastal Antarctic sediments. Limnology and Oceanography, 64(4):1423–1441, July 2019. CODEN LIOCAH. ISSN 0024-3590.

Benkort:2019:IBM

[BPW^{+}19] D. Benkort, S. Plourde, G. Winkler, J. Cabrol, A. Ollier, L.- E. Cope, and F. Maps. Individual-based modeling explains the contrasted seasonality in size, growth, and reproduction of the sympatric Arctic (Thysanoessa raschii) and Nordic krill (Meganyctiphanes norvegica) in the St. Lawrence Estuary, eastern Canada. Limnology and Oceanography, 64(1):217–237, January 2019. CODEN LIOCAH. ISSN 0024-3590.

Barrera:2017:HPA

[BR17] Kira E. Barrera and Lisa L. Robbins. Historical patterns of acidification and increasing CO_{2} flux associated with Florida springs. Limnology and Oceanography, 62(6):2404–2417, November 2017. CODEN LIOCAH. ISSN 0024-3590.

Brett:2010:LEG

[Bre10] Michael T. Brett. Is a low EPA growth saturation threshold supported by the data presented in Becker and boersma (2005)? Limnology and Oceanography, 55(1):455–458, January 2010. CODEN LIOCAH. ISSN 0024-3590.

Brett:2014:PNS

[Bre14] Michael T. Brett. Are phytoplankton in northern Swedish lakes extremely ^{13}C depleted? Limnology and Oceanography, 59(5): 1795–1799, September 2014. CODEN LIOCAH. ISSN 0024-3590. See response [KBA^{+}14].

Brosnahan:2017:BTT

[BRF^{+}17] Michael L. Brosnahan, David K. Ralston, Alexis D. Fischer, Andrew R. Solow, and Donald M. Anderson. Bloom termination of the toxic dinoflagellate Alexandrium catenella: Vertical migration behavior, sediment infiltration, and benthic cyst yield. Limnology and Oceanography, 62(6):2829–2849, November 2017. CODEN LIOCAH. ISSN 0024-3590.

Barber:2019:ICE

[BRM^{+}19] Julie S. Barber, Casey P. Ruff, James T. McArdle, Lindy L. Hunter, Camille A. Speck, Douglas W. Rogers, and Courtney M. Greiner. Intertidal clams exhibit population synchrony across spatial and temporal scales. Limnology and Oceanography, 64 (S1):S284–S300, January 2019. CODEN LIOCAH. ISSN 0024-

Bell:2018:RPP

[BRNS18] Tom W. Bell, Daniel C. Reed, Norman B. Nelson, and David A. Siegel. Regional patterns of physiological condition determine giant kelp net primary production dynamics. Limnology and Oceanography, 63(4):472–483, January 2018. CODEN LIOCAH. ISSN 0024-3590.

Baird:2013:DMC

[BRR^{+}13] Mark E. Baird, Peter J. Ralph, Farhan Rizwi, Karen Wild-Allen, and Andrew D. L. Steven. A dynamic model of the cellular carbon to chlorophyll ratio applied to a batch culture and a continental shelf ecosystem. Limnology and Oceanography, 58 (4):1215–1226, July 2013. CODEN LIOCAH. ISSN 0024-3590.

Bach:2011:DBE

[BRS11] Lennart Thomas Bach, Ulf Riebesell, and Kai Georg Schulz. Distinguishing between the effects of ocean acidification and ocean carbonation in the coccolithophore Emiliania huxleyi. Limnology and Oceanography, 56(6):2040–2050, November 2011. CO-DEN LIOCAH. ISSN 0024-3590.

Bucciarelli:2013:IIC

[BRS^{+}13] Eva Bucciarelli, Céline Ridame, William G. Sunda, Céline Dimier-Hugueney, Marie Cheize, and Sauveur Belviso. Increased intracellular concentrations of DMSP and DMSO in iron-limited oceanic phytoplankton Thalassiosira oceanica and Trichodesmium erythraeum. Limnology and Oceanography, 58 (5):1667–1679, September 2013. CODEN LIOCAH. ISSN 0024-3590.

Bodini:2018:IEB

[BRS18] Antonio Bodini, Marta Rocchi, and Marco Scotti. Insights into the ecology of the Black Sea through the qualitative loop analysis of the community structure. Limnology and Oceanography, 63 (2):968–984, March 2018. CODEN LIOCAH. ISSN 0024-3590.

Boscarino:2010:LEA

[BRT^{+}10] Brent T. Boscarino, Lars G. Rudstam, Jill Tirabassi, John Janssen, and Ellis R. Loew. Light effects on alewife-mysid interactions in Lake Ontario: a combined sensory physiology, behavioral, and spatial approach. Limnology and Oceanography, 55(5):2061–2072, September 2010. CODEN LIOCAH. ISSN 0024-3590.

Basu:2018:MIU

[BS18a] Subhajit Basu and Yeala Shaked. Mineral iron utilization by natural and cultured Trichodesmium and associated bacteria. Limnology and Oceanography, 63(6):2307–2320, November 2018. CODEN LIOCAH. ISSN 0024-3590.

Brothers:2018:LMT

[BS18b] Soren Brothers and Paul Sibley. Light may have triggered a period of net heterotrophy in Lake Superior. Limnology and Oceanography, 63(4):1785–1798, July 2018. CODEN LIOCAH. ISSN 0024-3590.

Burson:2016:URN

[BSA^{+}16] Amanda Burson, Maayke Stomp, Larissa Akil, Corina P. D. Brussaard, and Jef Huisman. Unbalanced reduction of nutrient loads has created an offshore gradient from phosphorus to nitrogen limitation in the North Sea. Limnology and Oceanography, 61(3):869–888, May 2016. CODEN LIOCAH. ISSN 0024-3590.

Boras:2010:EVP

[BSB^{+}10] Julia A. Boras, M. Montserrat Sala, Federico Baltar, Javier Arístegui, Carlos M. Duarte, and Dolors Vaqué. Effect of viruses and protists on bacteria in eddies of the Canary Current region (subtropical northeast Atlantic). Limnology and Oceanography, 55(4):885–898, March 2010. CODEN LIOCAH. ISSN 0024-3590.

Breithaupt:2018:ATB

[BSB^{+}18] Joshua L. Breithaupt, Joseph M. Smoak, Robert H. Byrne, Matthew N. Waters, Ryan P. Moyer, and Christian J. Sanders. Avoiding timescale bias in assessments of coastal wetland vertical change. Limnology and Oceanography, 63(S1):S477–S495, March 2018. CODEN LIOCAH. ISSN 0024-3590.

Branstrator:2013:ECP

[BSBK13] Donn K. Branstrator, Lyle J. Shannon, Meghan E. Brown, and Marte T. Kitson. Effects of chemical and physical conditions on hatching success of Bythotrephes longimanus resting eggs. Limnology and Oceanography, 58(6):2171–2184, November 2013. CODEN LIOCAH. ISSN 0024-3590.

Bristow:2015:BMA

[BSC^{+}15] Laura A. Bristow, Neha Sarode, John Cartee, Alejandro Caro-Quintero, Bo Thamdrup, and Frank J. Stewart. Biogeochemical and metagenomic analysis of nitrite accumulation in the Gulf of Mexico hypoxic zone. Limnology and Oceanography, 60(5): 1733–1750, September 2015. CODEN LIOCAH. ISSN 0024-3590.

Bartal:2015:MSE

[BSCC15] Roy Bartal, Bingyan Shi, William P. Cochlan, and Edward J. Carpenter. A model system elucidating calcification functions in the prymnesiophyte Emiliania huxleyi reveals dependence of nitrate acquisition on coccoliths. Limnology and Oceanography, 60(1):149–158, January 2015. CODEN LIOCAH. ISSN 0024-3590.

Barber:2017:SIA

[BSCG17] Andrew Barber, Maude Sirois, Gwénaëlle Chaillou, and Yves Gélinas. Stable isotope analysis of dissolved organic carbon in Canada’s eastern coastal waters. Limnology and Oceanography, 62(S1):S71–S84, November 2017. CODEN LIOCAH. ISSN 0024-3590.

Boyd:2010:ECO

[BSFH10] Philip W. Boyd, Robert Strzepek, Feixue Fu, and David A. Hutchins. Environmental control of open-ocean phytoplankton groups: Now and in the future. Limnology and Oceanography, 55 (3):1353–1376, May 2010. CODEN LIOCAH. ISSN 0024-3590.

Barron:2014:EIP

[BSG14] Rebecca K. Barrón, David A. Siegel, and Nathalie Guillocheau. Evaluating the importance of phytoplankton community structure to the optical properties of the Santa Barbara Channel, California. Limnology and Oceanography, 59(3):927–946, May 2014. CODEN LIOCAH. ISSN 0024-3590.

Brusin:2016:ICZ

[BSH16] Martin Brüsin, P. Andreas Svensson, and Samuel Hylander. Individual changes in zooplankton pigmentation in relation to ultraviolet radiation and predator cues. Limnology and Oceanography, 61(4):1337–1344, July 2016. CODEN LIOCAH. ISSN 0024-3590.

Bierschenk:2017:ICL

[BSM17] Antje M. Bierschenk, Candida Savage, and Christoph D. Matthaei. Intensity of catchment land use influences biological traits of benthic invertebrates along a freshwater-marine continuum. Limnology and Oceanography, 62(S1):S292–S308, November 2017. CODEN LIOCAH. ISSN 0024-3590.

Buchwald:2012:OIC

[BSMC12] Carolyn Buchwald, Alyson E. Santoro, Matthew R. McIlvin, and Karen L. Casciotti. Oxygen isotopic composition of nitrate and nitrite produced by nitrifying cocultures and natural marine assemblages. Limnology and Oceanography, 58(1):1361– 1375, September 2012. CODEN LIOCAH. ISSN 0024-3590.

Beaulieu:2014:DAB

[BSN^{+}14] Jake J. Beaulieu, Rebecca L. Smolenski, Christopher T. Nietch, Amy Townsend-Small, Michael S. Elovitz, and Joseph P. Schubauer-Berigan. Denitrification alternates between a source and sink of nitrous oxide in the hypolimnion of a thermally stratified reservoir. Limnology and Oceanography, 59(2):495– 506, March 2014. CODEN LIOCAH. ISSN 0024-3590.

Baer:2017:SNU

[BSR^{+}17] Steven E. Baer, Rachel E. Sipler, Quinn N. Roberts, Patricia L. Yager, Marc E. Frischer, and Deborah A. Bronk. Seasonal nitrogen uptake and regeneration in the western coastal Arctic. Limnology and Oceanography, 62(6):2463–2479, November 2017. CODEN LIOCAH. ISSN 0024-3590.

Britton-Simmons:2012:HBI

[BSRP^{+}12] Kevin H. Britton-Simmons, Alison L. Rhoades, Robert E. Pacunski, Aaron W. E. Galloway, Alexander T. Lowe, Elizabeth A. Sosik, Megan N. Dethier, and David O. Duggins. Habitat and bathymetry influence the landscape-scale distribution and abundance of drift macrophytes and associated invertebrates. Limnology and Oceanography, 57(1):176–184, January 2012. CO-DEN LIOCAH. ISSN 0024-3590.

Blukacz:2010:EEW

[BSSR10] E. A. Blukacz, W. G. Sprules, B. J. Shuter, and J. P. Richards. Evaluating the effect of wind-driven patchiness on trophic interactions between zooplankton and phytoplankton. Limnology and Oceanography, 55(4):1590–1600, July 2010. CODEN LI-OCAH. ISSN 0024-3590.

Bussmann:2011:APL

[BSSW11] Ingeborg Bussmann, Stefan Schlömer, Michael Schlüter, and Martin Wessels. Active pockmarks in a large lake (Lake Constance, Germany): Effects on methane distribution and turnover in the sediment. Limnology and Oceanography, 56(1):379–393, January 2011. CODEN LIOCAH. ISSN 0024-3590.

Burian:2016:BPC

[BSY^{+}16] Alfred Burian, Michael Schagerl, Andrew Yasindi, Gabriel Singer, Mary Nakabungo Kaggwa, and Monika Winder. Benthic-pelagic coupling drives non-seasonal zooplankton blooms and restructures energy flows in shallow tropical lakes. Limnology and Oceanography, 61(3):795–805, May 2016. CO-DEN LIOCAH. ISSN 0024-3590.

Boiteau:2019:PIS

[BTC^{+}19] Rene M. Boiteau, Claire P. Till, Tyler H. Coale, Jessica N. Fitzsimmons, Kenneth W. Bruland, and Daniel J. Repeta. Patterns of iron and siderophore distributions across the California Current System. Limnology and Oceanography, 64(1):376–389, January 2019. CODEN LIOCAH. ISSN 0024-3590.

Burt:2016:CSN

[BTH^{+}16] W. J. Burt, H. Thomas, M. Hagens, J. Pätsch, N. M. Clargo, L. A. Salt, V. Winde, and M. E. Böttcher. Carbon sources in the North Sea evaluated by means of radium and stable carbon isotope tracers. Limnology and Oceanography, 61(2):666–683, March 2016. CODEN LIOCAH. ISSN 0024-3590.

Bergkvist:2012:GIC

[BTJ^{+}12] Johanna Bergkvist, Peter Thor, Hans Henrik Jakobsen, Sten-Åke Wängberg, and Erik Selander. Grazer-induced chain length plasticity reduces grazing risk in a marine diatom. Limnology and Oceanography, 57(1):318–324, January 2012. CODEN LI-OCAH. ISSN 0024-3590.

Bouma:2016:STM

[BvBB^{+}16] T. J. Bouma, J. van Belzen, T. Balke, J. van Dalen, P. Klaassen, A. M. Hartog, D. P. Callaghan, Z. Hu, M. J. F. Stive, S. Temmerman, and P. M. J. Herman. Short-term mudflat dynamics drive long-term cyclic salt marsh dynamics. Limnology and Oceanography, 61(6):2261–2275, November 2016. CODEN LI-OCAH. ISSN 0024-3590.

Bellinger:2014:PMS

[BVC^{+}14] Brent J. Bellinger, Benjamin A. S. Van Mooy, James B. Cotner, Helen F. Fredricks, Claudia R. Benitez-Nelson, Jo Thompson, Anne Cotter, Michael L. Knuth, and Casey M. Godwin. Physiological modifications of seston in response to physicochemical gradients within Lake Superior. Limnology and Oceanography, 59(3):1011–1026, May 2014. CODEN LIOCAH. ISSN 0024-3590.

Brun:2015:ENO

[BVP^{+}15] Philipp Brun, Meike Vogt, Mark R. Payne, Nicolas Gruber, Colleen J. O’Brien, Erik T. Buitenhuis, Corinne Le Quéré, Karine Leblanc, and Ya-Wei Luo. Ecological niches of open ocean phytoplankton taxa. Limnology and Oceanography, 60 (3):1020–1038, May 2015. CODEN LIOCAH. ISSN 0024-3590.

Bittar:2015:CBP

[BVSM15] Thais B. Bittar, Armando A. H. Vieira, Aron Stubbins, and Kenneth Mopper. Competition between photochemical and biological degradation of dissolved organic matter from the cyanobacteria Microcystis aeruginosa. Limnology and Oceanography, 60(4):1172–1194, July 2015. CODEN LIOCAH. ISSN 0024-3590.

Brosnahan:2015:RGC

[BVSR^{+}15] Michael L. Brosnahan, Lourdes Velo-Suárez, David K. Ralston, Sophia E. Fox, Taylor R. Sehein, Alexi Shalapyonok, Heidi M. Sosik, Robert J. Olson, and Donald M. Anderson. Rapid growth and concerted sexual transitions by a bloom of the harmful dinoflagellate Alexandrium fundyense (Dinophyceae). Limnology and Oceanography, 60(6):2059–2078, November 2015. CODEN LIOCAH. ISSN 0024-3590.

Brinkmann:2019:BHP

[BVvB^{+}19] Bregje W. Brinkmann, J. Arie Vonk, Sebastiaan A. M. van Beusekom, Maria Ibanez, Miguel A. de Lucas Pardo, Ruurd Noordhuis, Erik M. M. Manders, Jolanda M. H. Verspagen, and Harm G. van der Geest. Benthic hotspots in the pelagic zone: Light and phosphate availability alter aggregates of microalgae and suspended particles in a shallow turbid lake. Limnology and Oceanography, 64(2):585–596, March 2019. CODEN LIOCAH. ISSN 0024-3590.

Beversdorf:2010:PMT

[BWB^{+}10] L. J. Beversdorf, A. E. White, K. M. Björkman, R. M. Letelier, and D. M. Karl. Phosphonate metabolism by Trichodesmium IMS101 and the production of greenhouse gases. Limnology and Oceanography, 55(4):1768–1778, July 2010. CODEN LIOCAH. ISSN 0024-3590.

Barrie:2015:EEM

[BWBB15] Gemma M. Barrie, Richard H. Worden, Craig D. Barrie, and Adrian J. Boyce. Extensive evaporation in a modern temperate estuary: Stable isotopic and compositional evidence. Limnology and Oceanography, 60(4):1241–1250, July 2015. CODEN LIOCAH. ISSN 0024-3590.

Brading:2011:DEO

[BWD^{+}11] Patrick Brading, Mark E. Warner, Phillip Davey, David J. Smith, Eric P. Achterberg, and David J. Suggett. Differential effects of ocean acidification on growth and photosynthesis among phylotypes of Symbiodinium (Dinophyceae). Limnology and Oceanography, 56(3):927–938, May 2011. CODEN LIOCAH. ISSN 0024-3590.

Brading:2012:EDE

[BWD^{+}12] Patrick Brading, Mark E. Warner, Phillip Davey, David J. Smith, Eric P. Achterberg, and David J. Suggett. Erratum: Differential effects of ocean acidification on growth and photosynthesis among phylotypes of Symbiodinium (Dinophyceae). Limnology and Oceanography, 57(4):1255, July 2012. CODEN LIOCAH. ISSN 0024-3590.

Bonnet:2010:VBE

[BWP^{+}10] Sophie Bonnet, Eric A. Webb, Caterina Panzeca, David M. Karl, Douglas G. Capone, and Sergio A. Sañudo Wilhelmy. Vitamin B_{12} excretion by cultures of the marine cyanobacteria Crocosphaera and Synechococcus. Limnology and Oceanography, 55 (5):1959–1964, September 2010. CODEN LIOCAH. ISSN 0024-3590.

Berkelmans:2010:ULW

[BWS10] Ray Berkelmans, Scarla J. Weeks, and Craig R. Steinberg. Upwelling linked to warm summers and bleaching on the Great Barrier Reef. Limnology and Oceanography, 55(6):2634–2644, November 2010. CODEN LIOCAH. ISSN 0024-3590.

Booth:2014:PPD

[BWS^{+}14] J. A. T. Booth, C. B. Woodson, M. Sutula, F. Micheli, S. B. Weisberg, S. J. Bograd, A. Steele, J. Schoen, and L. B. Crowder. Patterns and potential drivers of declining oxygen content along the southern California coast. Limnology and Oceanography, 59 (4):1127–1138, July 2014. CODEN LIOCAH. ISSN 0024-3590.

Barton:2019:QTD

[BYD19] Samuel Barton and Gabriel Yvon-Durocher. Quantifying the temperature dependence of growth rate in marine phytoplankton within and across species. Limnology and Oceanography, 66 (4):2081–2091, September 2019. CODEN LIOCAH. ISSN 0024-3590.

Chidami:2008:FDB

[CA08] Saad Chidami and Marc Amyot. Fish decomposition in boreal lakes and biogeochemical implications. Limnology and Oceanography, 53(5):1988–1996, September 2008. CODEN LIOCAH. ISSN 0024-3590.

Cha:2016:CSV

[CAQS16] YoonKyung Cha, Ibrahim Alameddine, Song S. Qian, and Craig A. Stow. A cross-scale view of N and P limitation using a Bayesian hierarchical model. Limnology and Oceanography, 61(6):2276–2285, November 2016. CODEN LIOCAH. ISSN 0024-3590.

Carson:2010:PCO

[Car10] Henry S. Carson. Population connectivity of the Olympia oyster in southern California. Limnology and Oceanography, 55(1):134– 148, January 2010. CODEN LIOCAH. ISSN 0024-3590.

Crosswell:2017:CBS

[CAS^{+}17] Joseph R. Crosswell, Iris C. Anderson, Jennifer W. Stanhope, Bryce Van Dam, Mark J. Brush, Scott Ensign, Michael F. Piehler, Brent McKee, Molly Bost, and Hans W. Paerl. Carbon budget of a shallow, lagoonal estuary: Transformations and source-sink dynamics along the river-estuary-ocean continuum. Limnology and Oceanography, 62(S1):S29–S45, November 2017. CODEN LIOCAH. ISSN 0024-3590.

Catala:2016:DFD

[CÁSO^{+}16] T. S. Catalá, X. A. Álvarez-Salgado, J. Otero, F. Iuculano, B. Companys, B. Horstkotte, C. Romera-Castillo, M. Nieto-Cid, M. Latasa, X. A. G. Morán, J. M. Gasol, C. Marrasé, C. A. Stedmon, and I. Reche. Drivers of fluorescent dissolved organic matter in the global epipelagic ocean. Limnology and Oceanography, 61(3):1101–1119, May 2016. CODEN LIOCAH. ISSN 0024-3590.

Champenois:2012:SIV

[CB12] W. Champenois and A. V. Borges. Seasonal and interannual variations of community metabolism rates of a Posidonia oceanica seagrass meadow. Limnology and Oceanography, 57(1):347– 361, January 2012. CODEN LIOCAH. ISSN 0024-3590.

Champenois:2019:IAV

[CB19] W. Champenois and A. V. Borges. Inter-annual variations over a decade of primary production of the seagrass Posidonia oceanica. Limnology and Oceanography, 64(1):32–45, January 2019. CODEN LIOCAH. ISSN 0024-3590.

Cordes:2010:TSN

[CBF10] Erik E. Cordes, Erin L. Becker, and Charles R. Fisherb. Temporal shift in nutrient input to cold-seep food webs revealed by stable-isotope signatures of associated communities. Limnology and Oceanography, 55(6):2537–2548, November 2010. CODEN LIOCAH. ISSN 0024-3590.

Chen:2011:CCL

[CBF11] Xi Chen, Stephen B. Baines, and Nicholas S. Fisher. Can copepods be limited by the iron content of their food? Limnology and Oceanography, 56(2):451–460, March 2011. CODEN LIOCAH. ISSN 0024-3590.

Casey:2019:SDM

[CBFK19] John R. Casey, Karin M. Björkman, Sara Ferrón, and David M. Karl. Size dependence of metabolism within marine picoplankton populations. Limnology and Oceanography, 64(4):1819– 1827, July 2019. CODEN LIOCAH. ISSN 0024-3590.

Carpenter:2018:EPP

[CBK18] Stephen R. Carpenter, Eric G. Booth, and Christopher J. Kucharik. Extreme precipitation and phosphorus loads from two agricultural watersheds. Limnology and Oceanography, 63 (3):1221–1233, May 2018. CODEN LIOCAH. ISSN 0024-3590.

Ciancio:2010:MEI

[CBP10] Javier Ciancio, David A. Beauchamp, and Miguel Pascual. Marine effect of introduced salmonids: Prey consumption by exotic steelhead and anadromous brown trout in the Patagonian Continental Shelf. Limnology and Oceanography, 55(5):2181–2192, September 2010. CODEN LIOCAH. ISSN 0024-3590.

Clark:2012:CWE

[CBP12] K. A. J. Clark, A. S. Brierley, and D. W. Pond. Composition of wax esters is linked to diapause behavior of Calanus finmarchicus in a sea loch environment. Limnology and Oceanography, 57 (1):65–75, January 2012. CODEN LIOCAH. ISSN 0024-3590.

Choi:2017:ILF

[CBS^{+}17] Chang Jae Choi, Michael L. Brosnahan, Taylor R. Sehein, Donald M. Anderson, and Deana L. Erdner. Insights into the loss factors of phytoplankton blooms: The role of cell mortality in the decline of two inshore Alexandrium blooms. Limnology and Oceanography, 62(5):1742–1753, July 2017. CODEN LIOCAH. ISSN 0024-3590.

Conrad:2010:SIF

[CCC10] Ralf Conrad, Peter Claus, and Peter Casperb. Stable isotope fractionation during the methanogenic degradation of organic matter in the sediment of an acidic bog lake, Lake Grosse Fuchskuhle. Limnology and Oceanography, 55(5):1932–1942, September 2010. CODEN LIOCAH. ISSN 0024-3590.

Chen:2012:SDE

[CCK^{+}12] Feizhou Chen, Meijun Chen, Fanxiang Kong, Xiaodong Wu, and Qinglong L. Wu. Species-dependent effects of crustacean plankton on a microbial community, assessed using an enclosure experiment in Lake Taihu, China. Limnology and Oceanography, 57(6):1711–1720, November 2012. CODEN LIOCAH. ISSN 0024-3590.

Cabrerizo:2018:INA

[CCV^{+}18] Marco J. Cabrerizo, Presentación Carrillo, Virginia E. Villafañe, Juan Manuel Medina-Sánchez, and E. Walter Helbling. Increased nutrients from aeolian-dust and riverine origin decrease the CO_{2}-sink capacity of coastal South Atlantic waters under UVR exposure. Limnology and Oceanography, 63(3):1191–1203, May 2018. CODEN LIOCAH. ISSN 0024-3590.

Chen:2019:BMO

[CCW^{+}19] Shunyang Chen, Gail L. Chmura, Yu Wang, Dan Yu, Danyun Ou, Bin Chen, Yong Ye, and Guangcheng Chen. Benthic microalgae offset the sediment carbon dioxide emission in subtropical mangrove in cold seasons. Limnology and Oceanography, 64 (3):1297–1308, May 2019. CODEN LIOCAH. ISSN 0024-3590.

Clark:2016:TVO

[CDA16] Catherine D. Clark, Warren J. De Bruyn, and Paige D. Aiona. Temporal variation in optical properties of chromophoric dissolved organic matter (CDOM) in Southern California coastal waters with nearshore kelp and seagrass. Limnology and Oceanography, 61(1):32–46, January 2016. CODEN LIOCAH. ISSN 0024-3590.

Cotte:2011:SDI

[CdC^{+}11] Cèdric Cotte, Francesco d’Ovidio, Alexis Chaigneau, Marina Lèvy, Isabelle Taupier-Letage, Bruce Mate, and Christophe Guinet. Scale-dependent interactions of Mediterranean whales with marine dynamics. Limnology and Oceanography, 56(1): 219–232, January 2011. CODEN LIOCAH. ISSN 0024-3590.

Canelhas:2016:MOW

[CDW^{+}16] Monica Ricão Canelhas, Blaize A. Denfeld, Gesa A. Weyhenmeyer, David Bastviken, and Stefan Bertilsson. Methane oxidation at the water–ice interface of an ice-covered lake. Limnology and Oceanography, 61(S1):S78–S90, November 2016. CODEN LIOCAH. ISSN 0024-3590.

Cohen:2017:IVI

[CEB^{+}17] Natalie R. Cohen, Kelsey A. Ellis, Wilton G. Burns, Robert H. Lampe, Nina Schuback, Zackary Johnson, Sergio Sañudo-Wilhelmy, and Adrian Marchetti. Iron and vitamin interactions in marine diatom isolates and natural assemblages of the Northeast Pacific Ocean. Limnology and Oceanography, 65(9):2076– 2096, September 2017. CODEN LIOCAH. ISSN 0024-3590.

Cadol:2014:EDS

[CEES14] Daniel Cadol, Katharina Engelhardt, Andrew Elmore, and Geoffrey Sanders. Elevation-dependent surface elevation gain in a tidal freshwater marsh and implications for marsh persistence. Limnology and Oceanography, 59(3):1065–1080, May 2014. CO-DEN LIOCAH. ISSN 0024-3590.

Cardoso:2014:DMO

[CEPPR14] Simone J. Cardoso, Alex Enrich-Prast, Michael L. Pace, and Fábio Roland. Do models of organic carbon mineralization extrapolate to warmer tropical sediments? Limnology and Oceanography, 59(1):48–54, January 2014. CODEN LIOCAH. ISSN 0024-3590.

Comeau:2013:REC

[CESC13] S. Comeau, P. J. Edmunds, N. B. Spindel, and R. C. Carpenter. The responses of eight coral reef calcifiers to increasing partial pressure of CO_{2} do not exhibit a tipping point. Limnology and Oceanography, 58(3):388–398, January 2013. CODEN LIOCAH. ISSN 0024-3590.

Comeau:2014:FCR

[CESC14] S. Comeau, P. J. Edmunds, N. B. Spindel, and R. C. Carpenter. Fast coral reef calcifiers are more sensitive to ocean acidification in short-term laboratory incubations. Limnology and Oceanography, 59(3):1081–1091, May 2014. CODEN LIOCAH. ISSN 0024-3590.

Comeau:2019:III

[CESC19] S. Comeau, P. J. Edmunds, N. B. Spindel, and R. C. Carpenter. Issue information — instr to contrib. Limnology and Oceanography, 59(3):i–iii, May 2019. CODEN LIOCAH. ISSN 0024-3590.

Curry:2010:ELD

[CF10] B. Brandon Curry and Gabriel M. Filippelli. Episodes of low dissolved oxygen indicated by ostracodes and sediment geochemistry at Crystal Lake, Illinois, USA. Limnology and Oceanography, 55(6):2403–2423, November 2010. CODEN LIOCAH. ISSN 0024-3590.

Campbell:2013:MBU

[CF13a] Justin E. Campbell and James W. Fourqurean. Mechanisms of bicarbonate use influence the photosynthetic carbon dioxide sensitivity of tropical seagrasses. Limnology and Oceanography, 58(3):839–848, May 2013. CODEN LIOCAH. ISSN 0024-3590.

Carey:2013:NEI

[CF13b] J. C. Carey and R. W. Fulweiler. Nitrogen enrichment increases net silica accumulation in a temperate salt marsh. Limnology and Oceanography, 58(2):99–111, January 2013. CODEN LI-OCAH. ISSN 0024-3590.

Carey:2014:SMT

[CF14] J. C. Carey and R. W. Fulweiler. Salt marsh tidal exchange increases residence time of silica in estuaries. Limnology and Oceanography, 59(4):1203–1212, July 2014. CODEN LIOCAH. ISSN 0024-3590.

Cauvy-Fraunie:2015:TSH

[CFAE^{+}15] Sophie Cauvy-Fraunié, Patricio Andino, Rodrigo Espinosa, Dean Jacobsen, and Olivier Dangles. Temporal scaling of high flow effects on benthic fauna: Insights from equatorial glacierfed streams. Limnology and Oceanography, 60(5):1836–1847, September 2015. CODEN LIOCAH. ISSN 0024-3590.

Chen:2014:IAI

[CFB14] Xi Chen, Nicholas S. Fisher, and Stephen B. Baines. Influence of algal iron content on the assimilation and fate of iron and carbon in a marine copepod. Limnology and Oceanography, 59 (1):129–140, January 2014. CODEN LIOCAH. ISSN 0024-3590.

Cavagna:2011:SUS

[CFD^{+}11] A.-J. Cavagna, F. Fripiat, F. Dehairs, D. Wolf-Gladrow, B. Cisewski, N. Savoye, L. Andrè, and D. Cardinal. Silicon uptake and supply during a Southern Ocean iron fertilization experiment (EIFEX) tracked by Si isotopes. Limnology and Oceanography, 56(1):147–160, January 2011. CODEN LIOCAH. ISSN 0024-3590.

Cao:2015:DSI

[CFD15] Zhimian Cao, Martin Frank, and Minhan Dai. Dissolved silicon isotopic compositions in the East China Sea: Water mass mixing vs. biological fractionation. Limnology and Oceanography, 60 (5):1619–1633, September 2015. CODEN LIOCAH. ISSN 0024-3590.

Cornacchia:2019:PFF

[CFD^{+}19] Loreta Cornacchia, Andrew Folkard, Grieg Davies, Robert C. Grabowski, Johan van de Koppel, Daphne van der Wal, Geraldene Wharton, Sara Puijalon, and Tjeerd J. Bouma. Plants face the flow in V formation: a study of plant patch alignment in streams. Limnology and Oceanography, 64(3):1087–1102, May 2019. CODEN LIOCAH. ISSN 0024-3590.

Crump:2017:QQP

[CFF^{+}17] Byron C. Crump, Lindy M. Fine, Caroline S. Fortunato, Lydie Herfort, Joseph A. Needoba, Sheryl Murdock, and Fredrick G. Prahl. Quantity and quality of particulate organic matter controls bacterial production in the Columbia River estuary. Limnology and Oceanography, 62(6):2713–2731, November 2017. CODEN LIOCAH. ISSN 0024-3590.

Cherel:2010:INT

[CFRL10] Yves Cherel, Camille Fontaine, Pierre Richard, and Jean-Philippe Labat. Isotopic niches and trophic levels of myctophid fishes and their predators in the Southern Ocean. Limnology and Oceanography, 55(1):324–332, January 2010. CODEN LI-OCAH. ISSN 0024-3590.

Collado-Fabbri:2011:SSD

[CFVU11] Silvana Collado-Fabbri, Daniel Vaulot, and Osvaldo Ulloa. Structure and seasonal dynamics of the eukaryotic picophytoplankton community in a wind-driven coastal upwelling ecosystem. Limnology and Oceanography, 56(6):2334–2346, November 2011. CODEN LIOCAH. ISSN 0024-3590.

Cortes:2014:PRW

[CFW^{+}14] A. Cortés, W. E. Fleenor, M. G. Wells, I. de Vicente, and F. J. Rueda. Pathways of river water to the surface layers of stratified reservoirs. Limnology and Oceanography, 59(1):233–250, January 2014. CODEN LIOCAH. ISSN 0024-3590.

Chu:2017:ELA

[CG17] Jackson W. F. Chu and Katie S. P. Gale. Ecophysiological limits to aerobic metabolism in hypoxia determine epibenthic distributions and energy sequestration in the northeast Pacific Ocean. Limnology and Oceanography, 62(1):59–74, January 2017. CO-DEN LIOCAH. ISSN 0024-3590.

Conley:2018:SCB

[CGB^{+}18] Keats R. Conley, Brad J. Gemmell, Jean-Marie Bouquet, Eric M. Thompson, and Kelly R. Sutherland. A self-cleaning biological filter: How appendicularians mechanically control particle adhesion and removal. Limnology and Oceanography, 63 (2):927–938, March 2018. CODEN LIOCAH. ISSN 0024-3590.

Coe:2016:SPE

[CGL^{+}16] Allison Coe, Julie Ghizzoni, Kristen LeGault, Steven Biller, Sara E. Roggensack, and Sallie W. Chisholm. Survival of Prochlorococcus in extended darkness. Limnology and Oceanography, 61(4):1375–1388, July 2016. CODEN LIOCAH. ISSN 0024-3590.

Caroselli:2019:LVP

[CGP^{+}19] Erik Caroselli, Francesca Gizzi, Fiorella Prada, Chiara Marchini, Valentina Airi, Jaap Kaandorp, Giuseppe Falini, Zvy Dubinsky, and Stefano Goffredo. Low and variable pH decreases recruitment efficiency in populations of a temperate coral naturally present at a CO_{2} vent. Limnology and Oceanography, 64 (3):1059–1069, May 2019. CODEN LIOCAH. ISSN 0024-3590.

Clayer:2016:RPS

[CGT16] François Clayer, Charles Gobeil, and André Tessier. Rates and pathways of sedimentary organic matter mineralization in two basins of a boreal lake: Emphasis on methanogenesis and methanotrophy. Limnology and Oceanography, 61(S1):S131– S149, November 2016. CODEN LIOCAH. ISSN 0024-3590.

Colvin:2011:SVJ

[CH11] Marienne A. Colvin and Brian T. Hentschel. Seasonal variability in the juvenile growth rate of an infaunal polychaete is related to major rain events. Limnology and Oceanography, 56(6):2095– 2102, November 2011. CODEN LIOCAH. ISSN 0024-3590.

Carilli:2017:PCR

[CHH^{+}17] Jessica E. Carilli, Aaron C. Hartmann, Scott F. Heron, John M. Pandolfi, Kim Cobb, Hussein Sayani, Robert Dunbar, and Stuart A. Sandin. Porites coral response to an oceanographic and human impact gradient in the Line Islands. Limnology and Oceanography, 62(6):2850–2863, November 2017. CODEN LI-OCAH. ISSN 0024-3590.

Capelle:2018:MYT

[CHHT18] David W. Capelle, Alyse K. Hawley, Steven J. Hallam, and Philippe D. Tortell. A multi-year time-series of N_{2}O dynamics in a seasonally anoxic fjord: Saanich Inlet, British Columbia. Limnology and Oceanography, 63(2):524–539, March 2018. CODEN LIOCAH. ISSN 0024-3590.

Cook:2010:EFE

[CHL10] Perran L. M. Cook, Daryl P. Holland, and Andrew R. Longmore. Effect of a flood event on the dynamics of phytoplankton and biogeochemistry in a large temperate Australian lagoon. Limnology and Oceanography, 55(3):1123–1133, May 2010. CODEN LIOCAH. ISSN 0024-3590.

Choi:2017:TIA

[CHL^{+}17] Bohyung Choi, Sun-Yong Ha, Jae Seong Lee, Yoshito Chikaraishi, Naohiko Ohkouchi, and Kyung-Hoon Shin. Trophic interaction among organisms in a seagrass meadow ecosystem as revealed by bulk δ ^{13}C and amino acid δ ^{15}N analyses. Limnology and Oceanography, 62(4):1426–1435, July 2017. CODEN LIOCAH. ISSN 0024-3590.

Cornwall:2013:CBL

[CHPH13] Christopher E. Cornwall, Christopher D. Hepburn, Conrad A. Pilditch, and Catriona L. Hurd. Concentration boundary layers around complex assemblages of macroalgae: Implications for the effects of ocean acidification on understory coralline algae. Limnology and Oceanography, 58(2):121–130, January 2013. CO-DEN LIOCAH. ISSN 0024-3590.

Christiansen:2018:PMF

[CHS^{+}18] Svenja Christiansen, Henk-Jan Hoving, Florian Schütte, Helena Hauss, Johannes Karstensen, Arne Körtzinger, Simon-Martin Schröder, Lars Stemmann, Bernd Christiansen, Marc Picheral, Peter Brandt, Bruce Robison, Reinhard Koch, and Rainer Kiko. Particulate matter flux interception in oceanic mesoscale eddies by the polychaete Poeobius sp. Limnology and Oceanography, 64(4):2093–2109, September 2018. CODEN LIOCAH. ISSN 0024-3590.

Clifford:2017:CZR

[CHV^{+}17] Elisabeth L. Clifford, Dennis A. Hansell, Marta M. Varela, Mar Nieto-Cid, Gerhard J. Herndl, and Eva Sintes. Crustacean zooplankton release copious amounts of dissolved organic matter as taurine in the ocean. Limnology and Oceanography, 62(6):2745– 2758, November 2017. CODEN LIOCAH. ISSN 0024-3590.

Cheng:2014:EES

[CHW14] Chiu H. Cheng, Markus Huettel, and Richard A. Wildman. Ebullition-enhanced solute transport in coarse-grained sediments. Limnology and Oceanography, 59(5):1733–1748, September 2014. CODEN LIOCAH. ISSN 0024-3590.

Cunningham:2017:EIL

[CJ17] Brady R. Cunningham and Seth G. John. The effect of iron limitation on cyanobacteria major nutrient and trace element stoichiometry. Limnology and Oceanography, 62(2):846–858, March 2017. CODEN LIOCAH. ISSN 0024-3590.

Cloern:2012:PCC

[CJC^{+}12] James E. Cloern, Alan D. Jassby, Jacob Carstensen, William A. Bennett, Wim Kimmerer, Ralph Mac Nally, David H. Schoellhamer, and Monika Winder. Perils of correlating CUSUMtransformed variables to infer ecological relationships (Breton et al. 2006; Glibert 2010). Limnology and Oceanography, 57(3): 665–668, March 2012. CODEN LIOCAH. ISSN 0024-3590. See rejoinder [LGR^{+}12].

Christianson:2019:ELC

[CJHR19] Kyle R. Christianson, Brett M. Johnson, Mevin B. Hooten, and James J. Roberts. Estimating lake–climate responses from sparse data: an application to high elevation lakes. Limnology and Oceanography, 64(3):1371–1385, May 2019. CODEN LIOCAH. ISSN 0024-3590.

Cloern:2017:EVA

[CJS^{+}17] James E. Cloern, Alan D. Jassby, Tara S. Schraga, Erica Nejad, and Charles Martin. Ecosystem variability along the estuarine salinity gradient: Examples from long-term study of San Francisco Bay. Limnology and Oceanography, 62(S1):S272–S291, November 2017. CODEN LIOCAH. ISSN 0024-3590.

Chang:2019:LRD

[CJW^{+}19] Bonnie X. Chang, Amal Jayakumar, Brittany Widner, Peter Bernhardt, Calvin W. Mordy, Margaret R. Mulholland, and Bess B. Ward. Low rates of dinitrogen fixation in the eastern tropical South Pacific. Limnology and Oceanography, 64(5): 1913–1923, September 2019. CODEN LIOCAH. ISSN 0024-

Craig:2015:HRA

[CJWS15] Nicola Craig, Stuart E. Jones, Brian C. Weidel, and Christopher T. Solomon. Habitat, not resource availability, limits consumer production in lake ecosystems. Limnology and Oceanography, 60(6):2079–2089, November 2015. CODEN LIOCAH. ISSN 0024-3590.

Cory:2012:BLS

[CK12] Rose M. Cory and Louis A. Kaplan. Biological lability of streamwater fluorescent dissolved organic matter. Limnology and Oceanography, 57(5):1347–1360, September 2012. CODEN LIOCAH. ISSN 0024-3590. See erratum [CK13].

Cory:2013:EBL

[CK13] Rose M. Cory and Louis A. Kaplan. Erratum: Biological lability of streamwater fluorescent dissolved organic matter. Limnology and Oceanography, 58(3):428, January 2013. CODEN LIOCAH. ISSN 0024-3590. See [CK12].

Cotovicz:2016:STV

[CKB^{+}16] Luiz C. Cotovicz, Jr., Bastiaan A. Knoppers, Nilva Brandini, Dominique Poirier, Suzan J. Costa Santos, and Gwenaël Abril. Spatio-temporal variability of methane (CH_{4}) concentrations and diffusive fluxes from a tropical coastal embayment surrounded by a large urban area (Guanabara Bay, Rio de Janeiro, Brazil). Limnology and Oceanography, 61(S1):S238– S252, November 2016. CODEN LIOCAH. ISSN 0024-3590.

Conrad:2010:MPI

[CKCEP10] Ralf Conrad, Melanie Klose, Peter Claus, and Alex Enrich-Prast. Methanogenic pathway, ^{13}C isotope fractionation, and archaeal community composition in the sediment of two clearwater lakes of Amazonia. Limnology and Oceanography, 55(2): 689–702, March 2010. CODEN LIOCAH. ISSN 0024-3590.

Chmiel:2016:RSC

[CKD^{+}16] Hannah E. Chmiel, Jovana Kokic, Blaize A. Denfeld, Karólína Einarsdóttir, Marcus B. Wallin, Birgit Koehler, Anastasija Isidorova, David Bastviken, Marie-Ève Ferland, and Sebastian Sobek. The role of sediments in the carbon budget of a small boreal lake. Limnology and Oceanography, 62(3):1814–1825, September 2016. CODEN LIOCAH. ISSN 0024-3590.

Catalan:2015:APE

[CKP^{+}15] Núria Catalán, Anne M. Kellerman, Hannes Peter, Francesc Carmona, and Lars J. Tranvik. Absence of a priming effect on dissolved organic carbon degradation in lake water. Limnology and Oceanography, 60(1):159–168, January 2015. CODEN LIOCAH. ISSN 0024-3590.

Chen:2010:RBP

[CL10] Bingzhang Chen and Hongbin Liu. Relationships between phytoplankton growth and cell size in surface oceans: Interactive effects of temperature, nutrients, and grazing. Limnology and Oceanography, 55(3):965–972, May 2010. CODEN LIOCAH. ISSN 0024-3590.

Chen:2011:CUR

[CL11] Bingzhang Chen and Hongbin Liu. Comment: Unimodal relationship between phytoplankton-mass–specific growth rate and size: a reply to the comment by Sal and López–Urrutia (2011). Limnology and Oceanography, 56(5):1956–1958, September 2011. CODEN LIOCAH. ISSN 0024-3590. See [SLU11].

Chen:2017:TDT

[CL17] Bingzhang Chen and Edward A. Laws. Is there a difference of temperature sensitivity between marine phytoplankton and heterotrophs? Limnology and Oceanography, 62(2):806–817, March 2017. CODEN LIOCAH. ISSN 0024-3590.

Cimatoribus:2019:TLT

[CLB19] A. A. Cimatoribus, U. Lemmin, and D. A. Barry. Tracking Lagrangian transport in Lake Geneva: a 3D numerical modeling investigation. Limnology and Oceanography, 64(3):1252–1269, May 2019. CODEN LIOCAH. ISSN 0024-3590.

Clayton:2017:CED

[CLFW17] Sophie Clayton, Yun-Chi Lin, Michael J. Follows, and Alexandra Z. Worden. Co-existence of distinct Ostreococcus ecotypes at an oceanic front. Limnology and Oceanography, 62(1):75–88, January 2017. CODEN LIOCAH. ISSN 0024-3590.

Chen:2012:DWE

[CLHL12] Bingzhang Chen, Michael R. Landry, Bangqin Huang, and Hongbin Liu. Does warming enhance the effect of microzooplankton grazing on marine phytoplankton in the ocean? Limnology and Oceanography, 57(3):519–526, March 2012. CODEN LIOCAH. ISSN 0024-3590.

Chen:2019:BDD

[CLJ^{+}19] Mingming Chen, Yangyang Lu, Nianzhi Jiao, Jiwei Tian, Shuh-Ji Kao, and Yao Zhang. Biogeographic drivers of diazotrophs in the western Pacific Ocean. Limnology and Oceanography, 64(3): 1403–1421, May 2019. CODEN LIOCAH. ISSN 0024-3590.

Chen:2014:EMG

[CLLH14] Bingzhang Chen, Edward A. Laws, Hongbin Liu, and Bangqin Huang. Estimating microzooplankton grazing half-saturation constants from dilution experiments with nonlinear feeding kinetics. Limnology and Oceanography, 59(3):639–644, May 2014. CODEN LIOCAH. ISSN 0024-3590.

Cornacchia:2019:TMF

[CLN^{+}19] Loreta Cornacchia, Sofia Licci, Heidi Nepf, Andrew Folkard, Daphne van der Wal, Johan van de Koppel, Sara Puijalon, and Tjeerd J. Bouma. Turbulence-mediated facilitation of resource uptake in patchy stream macrophytes. Limnology and Oceanography, 64(2):714–727, March 2019. CODEN LIOCAH. ISSN 0024-3590.

Cloern:2018:WLC

[Clo18] James E. Cloern. Why large cells dominate estuarine phytoplankton. Limnology and Oceanography, 63(S1):S392–S409, March 2018. CODEN LIOCAH. ISSN 0024-3590.

Cloern:2019:PPP

[Clo19] James E. Cloern. Patterns, pace, and processes of water-quality variability in a long-studied estuary. Limnology and Oceanography, 64(S1):S192–S208, January 2019. CODEN LIOCAH. ISSN 0024-3590.

Clark:2013:ESR

[CLWD13] James R. Clark, Timothy M. Lenton, Hywel T. P. Williams, and Stuart J. Daines. Environmental selection and resource allocation determine spatial patterns in picophytoplankton cell size. Limnology and Oceanography, 58(3):1008–1022, May 2013. CODEN LIOCAH. ISSN 0024-3590.

Casciotti:2010:OIE

[CMB10] Karen L. Casciotti, Matthew McIlvin, and Carolyn Buchwald. Oxygen isotopic exchange and fractionation during bacterial ammonia oxidation. Limnology and Oceanography, 55(2):753–762, March 2010. CODEN LIOCAH. ISSN 0024-3590. See erratum [Ano10].

Colin:2015:EPE

[CMG^{+}15] Sean P. Colin, Roshena MacPherson, Brad Gemmell, John H. Costello, Kelly Sutherland, and Cornelia Jaspers. Elevating the predatory effect: Sensory-scanning foraging strategy by the lobate ctenophore Mnemiopsis leidyi. Limnology and Oceanography, 60(1):100–109, January 2015. CODEN LIOCAH. ISSN 0024-3590.

Corman:2010:UCC

[CMK^{+}10] J. R. Corman, P. B. McIntyre, B. Kuboja, W. Mbemba, D. Fink, C. W. Wheeler, C. Gans, E. Michel, and A. S. Flecker. Upwelling couples chemical and biological dynamics across the littoral and pelagic zones of Lake Tanganyika, East Africa. Limnology and Oceanography, 55(1):214–224, January 2010. CODEN LIOCAH. ISSN 0024-3590.

Clark:2011:INA

[CMM^{+}11] Darren R. Clark, Peter I. Miller, E. Malcolm, S. Woodward, and Andrew P. Rees. Inorganic nitrogen assimilation and regeneration in the coastal upwelling region of the Iberian Peninsula. Limnology and Oceanography, 56(5):1689–1702, September 2011. CODEN LIOCAH. ISSN 0024-3590.

Chollett:2012:PEC

[CMMKH12] Iliana Chollett, Peter J. Mumby, Frank E. Müller-Karger, and Chuanmin Hu. Physical environments of the Caribbean Sea. Limnology and Oceanography, 57(4):1233–1244, July 2012. CO-DEN LIOCAH. ISSN 0024-3590.

Cortes:2017:FRS

[CMS17] Alicia Cortés, Sally MacIntyre, and Steven Sadro. Flowpath and retention of snowmelt in an ice-covered Arctic lake. Limnology and Oceanography, 65(9):2023–2044, September 2017. CODEN LIOCAH. ISSN 0024-3590.

Cohen:2018:ISC

[CMS^{+}18] Natalie R. Cohen, Elizabeth Mann, Brooke Stemple, Carly M. Moreno, Sara Rauschenberg, Jeremy E. Jacquot, William G.

Sunda, Benjamin S. Twining, and Adrian Marchetti. Iron storage capacities and associated ferritin gene expression among marine diatoms. Limnology and Oceanography, 63(4):1677–1691, July 2018. CODEN LIOCAH. ISSN 0024-3590.

Chang:2019:OBQ

[CMW^{+}19] G. Chang, T. Martin, K. Whitehead, C. Jones, and F. Spada. Optically based quantification of fluxes of mercury, methyl mercury, and polychlorinated biphenyls (PCBs) at Berry’s Creek tidal estuary, New Jersey. Limnology and Oceanography, 64(1): 93–108, January 2019. CODEN LIOCAH. ISSN 0024-3590.

Caroselli:2015:LVB

[CNL^{+}15] Erik Caroselli, Valentina Nanni, Oren Levy, Giuseppe Falini, Zvy Dubinsky, and Stefano Goffredo. Latitudinal variations in biometry and population density of a Mediterranean solitary coral. Limnology and Oceanography, 60(4):1356–1370, July 2015. CODEN LIOCAH. ISSN 0024-3590.

Cram:2016:DRH

[CPF16] Jacob A. Cram, Alma E. Parada, and Jed A. Fuhrman. Dilution reveals how viral lysis and grazing shape microbial communities. Limnology and Oceanography, 61(3):889–905, May 2016. CODEN LIOCAH. ISSN 0024-3590.

Chipman:2010:DPD

[CPG^{+}10] Lindsay Chipman, David Podgorski, Stefan Green, Joel Kostka, William Cooper, and Markus Huettel. Decomposition of plankton-derived dissolved organic matter in permeable coastal sediments. Limnology and Oceanography, 55(4):857–871, March 2010. CODEN LIOCAH. ISSN 0024-3590.

Choy:2015:TSF

[CPHD15] C. Anela Choy, Brian N. Popp, Cecelia C. S. Hannides, and Jeffrey C. Drazen. Trophic structure and food resources of epipelagic and mesopelagic fishes in the North Pacific Subtropical Gyre ecosystem inferred from nitrogen isotopic compositions. Limnology and Oceanography, 60(4):1156–1171, July 2015. CO-DEN LIOCAH. ISSN 0024-3590.

Ciros-Perez:2015:REB

[CPOMA15] Jorge Ciros-Pérez, Elizabeth Ortega-Mayagoitia, and Javier Alcocer. The role of ecophysiological and behavioral traits in structuring the zooplankton assemblage in a deep, oligotrophic, tropical lake. Limnology and Oceanography, 60(6):2158–2172, November 2015. CODEN LIOCAH. ISSN 0024-3590.

Carreon-Palau:2013:ROC

[CPPdAR^{+}13] Laura Carreón-Palau, Christopher C. Parrish, Jorge A. del Angel-Rodríguez, Horacio Pérez-España, and Sergio Aguiñiga-García. Revealing organic carbon sources fueling a coral reef food web in the Gulf of Mexico using stable isotopes and fatty acids. Limnology and Oceanography, 58(2):593–612, March 2013. CODEN LIOCAH. ISSN 0024-3590.

Cook:2010:ECS

[CR10] Freeman Cook and Alon Rimmer. Erratum: Chemical stratification in thermally stratified lakes: a chloride mass balance model. Limnology and Oceanography, 55(3):1463–1465, May 2010. CO-DEN LIOCAH. ISSN 0024-3590. See [RAKE05].

Carey:2011:LTS

[CR11] Cayelan C. Carey and Emil Rydin. Lake trophic status can be determined by the depth distribution of sediment phosphorus. Limnology and Oceanography, 56(6):2051–2063, November 2011. CODEN LIOCAH. ISSN 0024-3590.

Chen:2016:GDP

[CR16] Gang Chen and Tatiana A. Rynearson. Genetically distinct populations of a diatom co-exist during the North Atlantic spring bloom. Limnology and Oceanography, 61(6):2165–2179, November 2016. CODEN LIOCAH. ISSN 0024-3590.

Caroselli:2017:GPD

[CRB^{+}17] Erik Caroselli, Francesco Ricci, Viviana Brambilla, Chiara Marchini, Giada Tortorelli, Valentina Airi, Guido Mattioli, Oren Levy, Giuseppe Falini, Zvy Dubinsky, and Stefano Goffredo. Growth, population dynamics, and reproductive output model of the non-zooxanthellate temperate solitary coral Caryophyllia inornata (Scleractinia, Caryophylliidae). Limnology and Oceanography, 63(3):1111–1121, May 2017. CODEN LIOCAH. ISSN 0024-3590.

Casas-Ruiz:2017:TPR

[CRCGG^{+}17] Joan P. Casas-Ruiz, Núria Catalán, Lluís Gómez-Gener, Daniel von Schiller, Biel Obrador, Dolly N. Kothawala, Pilar López, Sergi Sabater, and Rafael Marcé. A tale of pipes and reactors: Controls on the in-stream dynamics of dissolved organic matter in rivers. Limnology and Oceanography, 62(S1):S85–S94, November 2017. CODEN LIOCAH. ISSN 0024-3590.

Chang:2014:EOC

[CRJ^{+}14] Bonnie X. Chang, Jeremy R. Rich, Amal Jayakumar, Hema Naik, Anil K. Pratihary, Richard G. Keil, Bess B. Ward, and Allan H. Devol. The effect of organic carbon on fixed nitrogen loss in the eastern tropical South Pacific and Arabian Sea oxygen deficient zones. Limnology and Oceanography, 59(4):1267–1274, July 2014. CODEN LIOCAH. ISSN 0024-3590.

Courtial:2017:IEU

[CRS^{+}17] Lucile Courtial, Stéphane Roberty, J. Malcolm Shick, Fanny Houlbrèque, and Christine Ferrier-Pagès. Interactive effects of ultraviolet radiation and thermal stress on two reef-building corals. Limnology and Oceanography, 63(3):1000–1013, May 2017. CODEN LIOCAH. ISSN 0024-3590.

Cole:2012:TSZ

[CS12] Jonathan J. Cole and Christopher T. Solomon. Terrestrial support of zebra mussels and the Hudson River food web: a multiisotope, Bayesian analysis. Limnology and Oceanography, 57(6): 1802–1815, November 2012. CODEN LIOCAH. ISSN 0024-3590.

Castillo:2010:PCD

[CSÁS^{+}10] Cristina Romera Castillo, Hugo Sarmento, Xosé Antón Álvarez-Salgado, Josep M. Gasol, and Celia Marraséa. Production of chromophoric dissolved organic matter by marine phytoplankton. Limnology and Oceanography, 55(1):446–454, January 2010. CODEN LIOCAH. ISSN 0024-3590.

Cuvelier:2011:CDY

[CSC^{+}11] Daphne Cuvelier, Jozée Sarrazin, Ana Colaço, Jon T. Copley, Adrian G. Glover, Paul A. Tyler, Ricardo Serrão Santos, and Daniel Desbruyères. Community dynamics over 14 years at the Eiffel Tower hydrothermal edifice on the Mid-Atlantic Ridge. Limnology and Oceanography, 56(5):1624–1640, September 2011. CODEN LIOCAH. ISSN 0024-3590.

Cousins:2010:ESS

[CSD10] Mary Cousins, Mark T. Stacey, and Jeana L. Drake. Effects of seasonal stratification on turbulent mixing in a hypereutrophic coastal lagoon. Limnology and Oceanography, 55(1):172–186, January 2010. CODEN LIOCAH. ISSN 0024-3590.

Chang:2018:TMA

[CSGW18] Jie C. Chang, James Shulmeister, Darren R. Gröcke, and Craig A. Woodward. Toward more accurate temperature reconstructions based on oxygen isotopes of subfossil chironomid head-capsules in Australia. Limnology and Oceanography, 63(1): 295–307, January 2018. CODEN LIOCAH. ISSN 0024-3590.

Chuang:2014:IRB

[CSJ^{+}14] Chia-Ying Chuang, Peter H. Santschi, Yuelu Jiang, Yi-Fang Ho, Antonietta Quigg, Laodong Guo, Marin Ayranov, and Dorothea Schumann. Important role of biomolecules from diatoms in the scavenging of particle-reactive radionuclides of thorium, protactinium, lead, polonium, and beryllium in the ocean: a case study with Phaeodactylum tricornutum. Limnology and Oceanography, 59(4):1256–1266, July 2014. CODEN LIOCAH. ISSN 0024-3590.

Cyronak:2013:CCH

[CSME13] Tyler Cyronak, Isaac R. Santos, Ashly McMahon, and Bradley D. Eyre. Carbon cycling hysteresis in permeable carbonate sands over a diel cycle: Implications for ocean acidification. Limnology and Oceanography, 58(2):131–143, January 2013. CODEN LIOCAH. ISSN 0024-3590.

Conroy:2016:MMG

[CSS^{+}16] Brandon J. Conroy, Deborah K. Steinberg, Michael R. Stukel, Joaquim I. Goes, and Victoria J. Coles. Meso- and microzooplankton grazing in the Amazon River plume and western tropical North Atlantic. Limnology and Oceanography, 61(3):825– 840, May 2016. CODEN LIOCAH. ISSN 0024-3590.

Cooper:2013:IGS

[CSU13] Matthew J. Cooper, Alan D. Steinman, and Donald G. Uzarski. Influence of geomorphic setting on the metabolism of Lake Huron fringing wetlands. Limnology and Oceanography, 58(2): 452–464, March 2013. CODEN LIOCAH. ISSN 0024-3590.

Cannon:2018:OTM

[CT18a] David J. Cannon and Cary D. Troy. Observations of turbulence and mean flow in the low-energy hypolimnetic boundary layer of a large lake. Limnology and Oceanography, 63(6):2762–2776, November 2018. CODEN LIOCAH. ISSN 0024-3590.

Chen:2018:BCF

[CT18b] Si Chen and Raymond Torres. Biogeochemical characteristics and fluxes of suspended particulate organic matter in response to low-tide rainfall. Limnology and Oceanography, 63(S1):S307– S323, March 2018. CODEN LIOCAH. ISSN 0024-3590.

Cabrol:2019:TNP

[CTA^{+}19] J. Cabrol, T. Trombetta, S. Amaudrut, F. Aulanier, R. Sage, R. Tremblay, C. Nozais, M. Starr, S. Plourde, and G. Winkler. Trophic niche partitioning of dominant North-Atlantic krill species, Meganyctiphanes norvegica, Thysanoessa inermis, and T. raschii. Limnology and Oceanography, 64(1):165–181, January 2019. CODEN LIOCAH. ISSN 0024-3590.

Chen:2015:IZP

[CTG15] Si Chen, Raymond Torres, and Miguel A. Goñi. Intertidal zone particulate organic carbon redistribution by low-tide rainfall. Limnology and Oceanography, 60(3):1088–1101, May 2015. CO-DEN LIOCAH. ISSN 0024-3590.

Choi:2015:SDN

[CTH15] Jun M. Choi, Cary D. Troy, and Nathan Hawley. Shear dispersion from near-inertial internal Poincaré waves in large lakes. Limnology and Oceanography, 60(6):2222–2235, November 2015. CODEN LIOCAH. ISSN 0024-3590.

Collier:2011:TTT

[CUW11] Catherine J. Collier, Sven Uthicke, and Michelle Waycott. Thermal tolerance of two seagrass species at contrasting light levels: Implications for future distribution in the Great Barrier Reef. Limnology and Oceanography, 56(6):2200–2210, November 2011. CODEN LIOCAH. ISSN 0024-3590.

Cardini:2018:FPP

[CvHB^{+}18] U. Cardini, N. van Hoytema, V. N. Bednarz, M. M. D. Al-Rshaidat, and C. Wild. N_{2} fixation and primary productivity in a Red Sea Halophila stipulacea meadow exposed to seasonality. Limnology and Oceanography, 63(2):786–798, March 2018. CODEN LIOCAH. ISSN 0024-3590.

Countway:2010:SAP

[CVS^{+}10] Peter D. Countway, Patrick D. Vigil, Astrid Schnetzer, Stefanie D. Moorthi, and David A. Caron. Seasonal analysis of protistan community structure and diversity at the USC Microbial Observatory (San Pedro Channel, North Pacific Ocean). Limnology and Oceanography, 55(6):2381–2396, November 2010. CODEN LIOCAH. ISSN 0024-3590.

Chen:2011:IIF

[CWF11] Xi Chen, Stuart G. Wakeham, and Nicholas S. Fisher. Influence of iron on fatty acid and sterol composition of marine phytoplankton and copepod consumers. Limnology and Oceanography, 56(2):716–724, March 2011. CODEN LIOCAH. ISSN 0024-3590.

Crosswell:2014:ECE

[CWHP14] Joseph R. Crosswell, Michael S. Wetz, Burke Hales, and Hans W. Paerl. Extensive CO_{2} emissions from shallow coastal waters during passage of Hurricane Irene (August 2011) over the Mid-Atlantic Coast of the U.S.A. Limnology and Oceanography, 59(5):1651–1665, September 2014. CODEN LIOCAH. ISSN 0024-3590.

Cui:2019:PDS

[CWRX19] Yongsheng Cui, Jiaxue Wu, Jie Ren, and Jie Xu. Physical dynamics structures and oxygen budget of summer hypoxia in the Pearl River Estuary. Limnology and Oceanography, 64(1):131– 148, January 2019. CODEN LIOCAH. ISSN 0024-3590.

Cao:2018:ESD

[CZB^{+}18] Haobing Cao, Zhenchang Zhu, Thorsten Balke, Liquan Zhang, and Tjeerd J. Bouma. Effects of sediment disturbance regimes on Spartina seedling establishment: Implications for salt marsh creation and restoration. Limnology and Oceanography, 63(2): 647–659, March 2018. CODEN LIOCAH. ISSN 0024-3590.

Dimova:2011:EGD

[DB11] Natasha T. Dimova and William C. Burnett. Evaluation of groundwater discharge into small lakes based on the temporal distribution of radon-222. Limnology and Oceanography, 56(2): 486–494, March 2011. CODEN LIOCAH. ISSN 0024-3590.

Dorostkar:2013:IHJ

[DB13] Abbas Dorostkar and Leon Boegman. Internal hydraulic jumps in a long narrow lake. Limnology and Oceanography, 58(2):153– 172, January 2013. CODEN LIOCAH. ISSN 0024-3590.

DeFalco:2016:BLD

[DBA16] Natalie De Falco, Fulvio Boano, and Shai Arnon. Biodegradation of labile dissolved organic carbon under losing and gaining streamflow conditions simulated in a laboratory flume. Limnology and Oceanography, 62(3):1839–1852, September 2016. CO-DEN LIOCAH. ISSN 0024-3590.

Durkin:2013:SAS

[DBC^{+}13] Colleen A. Durkin, Sara J. Bender, Kit Yu Karen Chan, Kelsey Gaessner, Daniel Grünbaum, and E. Virginia Armbrust. Silicic acid supplied to coastal diatom communities influences cellular silicification and the potential export of carbon. Limnology and Oceanography, 58(5):1707–1726, September 2013. CODEN LIOCAH. ISSN 0024-3590.

Donald:2011:CEU

[DBFL11] Derek B. Donald, Matthew J. Bogard, Kerri Finlay, and Peter R. Leavitt. Comparative effects of urea, ammonium, and nitrate on phytoplankton abundance, community composition, and toxicity in hypereutrophic freshwaters. Limnology and Oceanography, 56(6):2161–2175, November 2011. CODEN LIOCAH. ISSN 0024-3590.

Diaz:2016:PDS

[DBH^{+}16] Julia M. Diaz, Karin M. Björkman, Sheean T. Haley, Ellery D. Ingall, David M. Karl, Amelia F. Longo, and Sonya T. Dyhrman. Polyphosphate dynamics at Station ALOHA, North Pacific subtropical gyre. Limnology and Oceanography, 61(1):227–239, January 2016. CODEN LIOCAH. ISSN 0024-3590.

Dimond:2011:SSF

[DBMP^{+}11] J. L. Dimond, B. L. Bingham, G. Muller-Parker, K. Wuesthoff, and L. Francis. Seasonal stability of a flexible algal–cnidarian symbiosis in a highly variable temperate environment. Limnology and Oceanography, 56(6):2233–2242, November 2011. CO-DEN LIOCAH. ISSN 0024-3590.

Dexter:2015:PVE

[DBRB^{+}15] Eric Dexter, Stephen M. Bollens, Gretchen Rollwagen-Bollens, Josh Emerson, and Julie Zimmerman. Persistent vs. ephemeral invasions: 8.5 years of zooplankton community dynamics in the Columbia River. Limnology and Oceanography, 60(2):527–539, March 2015. CODEN LIOCAH. ISSN 0024-3590.

DelSontro:2016:MED

[DBSP^{+}16] Tonya DelSontro, Lennie Boutet, Annick St-Pierre, Paul A. del Giorgio, and Yves T. Prairie. Methane ebullition and diffusion from northern ponds and lakes regulated by the interaction between temperature and system productivity. Limnology and Oceanography, 61(S1):S62–S77, November 2016. CODEN LI-OCAH. ISSN 0024-3590.

Duhamel:2011:CAP

[DBV^{+}11] Solange Duhamel, Karin M. Björkman, France Van Wambeke, Thierry Moutin, and David M. Karl. Characterization of alkaline phosphatase activity in the North and South Pacific Subtropical Gyres: Implications for phosphorus cycling. Limnology and Oceanography, 56(4):1244–1254, July 2011. CODEN LIOCAH. ISSN 0024-3590.

deBettignies:2013:CMD

[dBWL^{+}13] Thibaut de Bettignies, Thomas Wernberg, Paul S. Lavery, Mathew A. Vanderklift, and Margaret B. Mohring. Contrasting mechanisms of dislodgement and erosion contribute to production of kelp detritus. Limnology and Oceanography, 58(5):1680– 1688, September 2013. CODEN LIOCAH. ISSN 0024-3590.

Dove:2015:LTT

[DC15] Alice Dove and Steven C. Chapra. Long-term trends of nutrients and trophic response variables for the Great Lakes. Limnology and Oceanography, 60(2):696–721, March 2015. CODEN LIOCAH. ISSN 0024-3590.

Darrow:2017:LUR

[DCCB17] Elizabeth S. Darrow, Ruth H. Carmichael, Kevin R. Calci, and William Burkhardt III. Land-use related changes to sedimentary organic matter in tidal creeks of the northern Gulf of Mexico. Limnology and Oceanography, 62(2):686–705, March 2017. CODEN LIOCAH. ISSN 0024-3590.

delCampo:2019:DPC

[dCGS19] Rubén del Campo, Rosa Gómez, and Gabriel Singer. Dry phase conditions prime wet-phase dissolved organic matter dynamics in intermittent rivers. Limnology and Oceanography, 66(8):1966– 1979, September 2019. CODEN LIOCAH. ISSN 0024-3590.

DelBelBelluz:2016:BOV

[DCRC16] Justin Del Bel Belluz, Maycira Costa, Gregor Reid, and Stephen Cross. Bio-optical variability at a Vancouver Island aquaculture site. Limnology and Oceanography, 62(3):1686–1704, September 2016. CODEN LIOCAH. ISSN 0024-3590.

DAlelio:2010:TSB

[DdD^{+}10] Domenico D’Alelio, Maurizio Ribera d’Alcalà, Laurent Dubroca, Diana Sarn, Adriana Zingone, and Marina Montresor. The time for sex: a biennial life cycle in a marine planktonic diatom. Limnology and Oceanography, 55(1):106–114, January 2010. CO-DEN LIOCAH. ISSN 0024-3590.

Douglass:2010:SIC

[DDF^{+}10] James G. Douglass, James G. Douglass, Kristin E. France, Kristin E. France, J. Paul Richardson, and J. Emmett Duffy. Seasonal and interannual change in a Chesapeake Bay eelgrass community: Insights into biotic and abiotic control of community structure. Limnology and Oceanography, 55(4):1499–1520, July 2010. CODEN LIOCAH. ISSN 0024-3590.

Duarte:2010:ETG

[DdG10] Carlos M. Duarte and Aurore Regaudie de Gioux. Erratum: Thresholds of gross primary production for the metabolic balance of marine planktonic communities. Limnology and Oceanography, 55(6):2723–2725, November 2010. CODEN LI-OCAH. ISSN 0024-3590.

Dodson:2019:EDL

[DDH^{+}19] J. J. Dodson, G. Daigle, C. Hammer, P. Polte, P. Kotterba, G. Winkler, and C. Zimmermann. Environmental determinants of larval herring (Clupea harengus) abundance and distribution in the western Baltic Sea. Limnology and Oceanography, 64(1): 317–329, January 2019. CODEN LIOCAH. ISSN 0024-3590.

Duhamel:2010:APA

[DDK10] Solange Duhamel, Sonya T. Dyhrman, and David M. Karl. Alkaline phosphatase activity and regulation in the North Pacific Subtropical Gyre. Limnology and Oceanography, 55(3):1414– 1425, May 2010. CODEN LIOCAH. ISSN 0024-3590.

Demars:2019:HPB

[Dem19] Benoît O. L. Demars. Hydrological pulses and burning of dissolved organic carbon by stream respiration. Limnology and Oceanography, 64(1):406–421, January 2019. CODEN LIOCAH. ISSN 0024-3590.

Deininger:2017:PFW

[DFK^{+}17] A. Deininger, C. L. Faithfull, J. Karlsson, M. Klaus, and A.-K. Bergström. Pelagic food web response to whole lake N fertilization. Limnology and Oceanography, 62(4):1498–1511, July 2017. CODEN LIOCAH. ISSN 0024-3590.

Dadi:2016:BDO

[DFWPK16] Tallent Dadi, Kurt Friese, Katrin Wendt-Potthoff, and Matthias Koschorreck. Benthic dissolved organic carbon fluxes in a drinking water reservoir. Limnology and Oceanography, 61(2):445– 459, March 2016. CODEN LIOCAH. ISSN 0024-3590.

delGiorgio:2011:CPB

[dGCB^{+}11] Paul A. del Giorgio, Robert Condon, Thierry Bouvier, Krista. Longnecker, Corinne. Bouvier, Evelyn Sherr, and Josep M. Gasol. Coherent patterns in bacterial growth, growth efficiency, and leucine metabolism along a northeastern Pacific inshoreoffshore transect. Limnology and Oceanography, 56(1):1–16, January 2011. CODEN LIOCAH. ISSN 0024-3590.

Regaudie-de-Gioux:2013:GPO

[dGD13] Aurore Regaudie de Gioux and Carlos M. Duarte. Global patterns in oceanic planktonic metabolism. Limnology and Oceanography, 58(3):977–986, May 2013. CODEN LIOCAH. ISSN 0024-3590.

Darnis:2017:PNM

[DHG^{+}17] G. Darnis, L. Hobbs, M. Geoffroy, J. C. Grenvald, P. E. Renaud, J. Berge, F. Cottier, S. Kristiansen, M. Daase, J. E. Søreide, A. Wold, N. Morata, and T. Gabrielsen. From polar night to midnight sun: Diel vertical migration, metabolism and biogeochemical role of zooplankton in a high Arctic fjord (Kongsfjorden, Svalbard). Limnology and Oceanography, 62(4):1586–1605, July 2017. CODEN LIOCAH. ISSN 0024-3590.

Deemer:2015:CMB

[DHH15] Bridget R. Deemer, Stephen M. Henderson, and John A. Harrison. Chemical mixing in the bottom boundary layer of a eutrophic reservoir: The effects of internal seiching on nitrogen dynamics. Limnology and Oceanography, 60(5):1642–1655, September 2015. CODEN LIOCAH. ISSN 0024-3590.

Drillet:2011:REP

[DHK11] Guillaume Drillet, Benni W. Hansen, and Thomas Kiørboe. Resting egg production induced by food limitation in the calanoid copepod Acartia tonsa. Limnology and Oceanography, 56(6):2064–2070, November 2011. CODEN LIOCAH. ISSN 0024-3590.

Deemer:2011:MDN

[DHW11] Bridget R. Deemer, John A. Harrison, and Elliott W. Whitling. Microbial dinitrogen and nitrous oxide production in a small eutrophic reservoir: an in situ approach to quantifying hypolimnetic process rates. Limnology and Oceanography, 56(4):1189– 1199, July 2011. CODEN LIOCAH. ISSN 0024-3590.

Drake:2019:MCI

[DHZ^{+}19] Travis W. Drake, Robert M. Holmes, Alexander V. Zhulidov, Tatiana Gurtovaya, Peter A. Raymond, James W. McClelland, and Robert G. M. Spencer. Multidecadal climate-induced changes in Arctic tundra lake geochemistry and geomorphology. Limnology and Oceanography, 64(S1):S179–S191, January 2019. CODEN LIOCAH. ISSN 0024-3590.

Dahl:2018:ICF

[DIC^{+}18] Martin Dahl, Eduardo Infantes, Rosanna Clevesjö, Hans W. Linderholm, Mats Björk, and Martin Gullström. Increased current flow enhances the risk of organic carbon loss from Zostera marina sediments: Insights from a flume experiment. Limnology and Oceanography, 63(6):2793–2805, November 2018. CODEN LIOCAH. ISSN 0024-3590.

Deschaseaux:2014:EEF

[DJD^{+}14] E. S. M. Deschaseaux, G. B. Jones, M. A. Deseo, K. M. Shepherd, R. P. Kiene, H. B. Swan, P. L. Harrison, and B. D. Eyre. Effects of environmental factors on dimethylated sulfur compounds and their potential role in the antioxidant system of the coral holobiont. Limnology and Oceanography, 59(3):758–768, May 2014. CODEN LIOCAH. ISSN 0024-3590.

Dunlop:2018:SPJ

[DJS18] Kathy M. Dunlop, Daniel O. B. Jones, and Andrew K. Sweetman. Scavenging processes on jellyfish carcasses across a fjord depth gradient. Limnology and Oceanography, 63(3):1146–1155, May 2018. CODEN LIOCAH. ISSN 0024-3590.

Daniels:2015:BCM

[DKG15] William C. Daniels, George W. Kling, and Anne E. Giblin. Benthic community metabolism in deep and shallow Arctic lakes during 13 years of whole–lake fertilization. Limnology and Oceanography, 60(5):1604–1618, September 2015. CODEN LI-OCAH. ISSN 0024-3590.

Doran:2014:RDE

[DKK^{+}14] Peter T. Doran, Fabien Kenig, Jennifer Lawson Knoepfle, Jill A. Mikucki, and W. Berry Lyons. Radiocarbon distribution and the effect of legacy in lakes of the McMurdo Dry Valleys, Antarctica. Limnology and Oceanography, 59(3):811–826, May 2014. CODEN LIOCAH. ISSN 0024-3590.

deKluijver:2015:MPC

[dKNL^{+}15] A. de Kluijver, J. Ning, Z. Liu, E. Jeppesen, R. D. Gulati, and J. J. Middelburg. Macrophytes and periphyton carbon subsidies to bacterioplankton and zooplankton in a shallow eutrophic lake in tropical China. Limnology and Oceanography, 60(2):375–385, March 2015. CODEN LIOCAH. ISSN 0024-3590.

Duhamel:2019:SPE

[DKSA19] Solange Duhamel, Eunsoo Kim, Ben Sprung, and O. Roger Anderson. Small pigmented eukaryotes play a major role in carbon cycling in the P-depleted western subtropical North Atlantic, which may be supported by mixotrophy. Limnology and Oceanography, 64(6):2424–2440, November 2019. CODEN LI-OCAH. ISSN 0024-3590.

deKluijver:2012:CCS

[dKYH^{+}12] Anna de Kluijver, Jinlei Yu, Marco Houtekamer, Jack J. Middelburg, and Zhengwen Liu. Cyanobacteria as a carbon source for zooplankton in eutrophic Lake Taihu, China, measured by ^{13}C labeling and fatty acid biomarkers. Limnology and Oceanography, 57(4):1245–1254, July 2012. CODEN LIOCAH. ISSN 0024-3590.

Dugan:2011:CDH

[DL11] Hilary A. Dugan and Scott F. Lamoureux. The chemical development of a hypersaline coastal basin in the High Arctic. Limnology and Oceanography, 56(2):495–507, March 2011. CODEN LIOCAH. ISSN 0024-3590.

Decima:2017:ATF

[DLBF17] Moira Décima, Michael R. Landry, Christina J. Bradley, and Marilyn L. Fogel. Alanine δ ^{15}N trophic fractionation in heterotrophic protists. Limnology and Oceanography, 65(9):2308– 2322, September 2017. CODEN LIOCAH. ISSN 0024-3590.

delaFuente:2010:TSF

[dlFN10] Alberto de la Fuente and Yarko Ninóo. Temporal and spatial features of the thermohydrodynamics of shallow salty lagoons in northern Chile. Limnology and Oceanography, 55(1):279–288, January 2010. CODEN LIOCAH. ISSN 0024-3590.

Decima:2013:EPE

[DLP13] Moira Décima, Michael R. Landry, and Brian N. Popp. Environmental perturbation effects on baseline δ ^{15}N values and zooplankton trophic flexibility in the southern California Current ecosystem. Limnology and Oceanography, 58(2):624–634, March 2013. CODEN LIOCAH. ISSN 0024-3590.

Davis:2017:EAP

[DM17] Clare E. Davis and Claire Mahaffey. Elevated alkaline phosphatase activity in a phosphate-replete environment: Influence of sinking particles. Limnology and Oceanography, 62(6):2389– 2403, November 2017. CODEN LIOCAH. ISSN 0024-3590.

Durkin:2012:FRG

[DMB^{+}12] Colleen A. Durkin, Adrian Marchetti, Sara J. Bender, Tiffany Truong, Rhonda Morales, Thomas Mock, and E. Virginia Armbrust. Frustule-related gene transcription and the influence of diatom community composition on silica precipitation in an iron-limited environment. Limnology and Oceanography, 57(6): 1619–1633, November 2012. CODEN LIOCAH. ISSN 0024-3590.

Detmer:2017:PDC

[DML17] Thomas M. Detmer, James H. McCutchan, Jr., and William M. Lewis, Jr. Predator driven changes in prey size distribution stabilize secondary production in lacustrine food webs. Limnology and Oceanography, 62(2):592–605, March 2017. CODEN LIOCAH. ISSN 0024-3590.

Deshpande:2015:ODP

[DMMV15] Bethany N. Deshpande, Sally MacIntyre, Alex Matveev, and Warwick F. Vincent. Oxygen dynamics in permafrost thaw lakes: Anaerobic bioreactors in the Canadian subarctic. Limnology and Oceanography, 60(5):1656–1670, September 2015. CO-DEN LIOCAH. ISSN 0024-3590.

Deng:2015:ECM

[DMN15] Wei Deng, Logan Monks, and Susanne Neuer. Effects of clay minerals on the aggregation and subsequent settling of marine Synechococcus. Limnology and Oceanography, 60(3):805–816, May 2015. CODEN LIOCAH. ISSN 0024-3590.

Danner:2018:RAT

[DMS^{+}18] Kelsey M. Danner, Megan A. Mave, Audrey H. Sawyer, Seungjun Lee, and Jiyoung Lee. Removal of the algal toxin microcystin-LR in permeable coastal sediments: Physical and numerical models. Limnology and Oceanography, 63(4):1593– 1604, July 2018. CODEN LIOCAH. ISSN 0024-3590.

Duran:2016:CPB

[DMSHC16] Cristina Durán, Juan Manuel Medina-Sánchez, Guillermo Herrera, and Presentación Carrillo. Changes in the phytoplanktonbacteria coupling triggered by joint action of UVR, nutrients, and warming in Mediterranean high-mountain lakes. Limnology and Oceanography, 61(2):413–429, March 2016. CODEN LIOCAH. ISSN 0024-3590.

DeMartini:2018:CSS

[DNH^{+}18] Francesca De Martini, Susanne Neuer, Demetra Hamill, Julie Robidart, and Michael W. Lomas. Clade and strain specific contributions of Synechococcus and Prochlorococcus to carbon export in the Sargasso Sea. Limnology and Oceanography, 63 (S1):S448–S457, March 2018. CODEN LIOCAH. ISSN 0024-3590.

Decima:2010:BSD

[DOD10] Moira Décima, Mark D. Ohman, and Alex De Robertis. Body size dependence of euphausiid spatial patchiness. Limnology and Oceanography, 55(4):777–788, March 2010. CODEN LIOCAH. ISSN 0024-3590.

Dittmar:2012:DDB

[DPG^{+}12] Thorsten Dittmar, Jiyoung Paeng, Thomas M. Gihring, I. G. N. A. Suryaputra, and Markus Huettel. Discharge of dissolved black carbon from a fire-affected intertidal system. Limnology and Oceanography, 57(4):1171–1181, July 2012. CODEN LI-OCAH. ISSN 0024-3590.

Dadon-Pilosof:2019:PTR

[DPLG^{+}19] Ayelet Dadon-Pilosof, Fabien Lombard, Amatzia Genin, Kelly R. Sutherland, and Gitai Yahel. Prey taxonomy rather than size determines salp diets. Limnology and Oceanography, 66(8):1996–2010, September 2019. CODEN LIOCAH. ISSN 0024-3590.

Drobnitch:2018:PDV

[DPM18] Sarah Tepler Drobnitch, Taylor Pochron, and Celine Miranda. Patterns and drivers of δ ^{13}C variation in the giant kelp, Macrocystis pyrifera. Limnology and Oceanography, 63(2):871–885, March 2018. CODEN LIOCAH. ISSN 0024-3590.

DelSontro:2016:MDD

[DPSW16] Tonya DelSontro, Karina K. Perez, Sébastien Sollberger, and Bernhard Wehrli. Methane dynamics downstream of a temperate run-of-the-river reservoir. Limnology and Oceanography, 61(S1):S188–S203, November 2016. CODEN LIOCAH. ISSN 0024-3590.

Davis:2010:NED

[DRE^{+}10] John M. Davis, Amy D. Rosemond, Susan L. Eggert, Wyatt F. Cross, and J. Bruce Wallace. Nutrient enrichment differentially affects body sizes of primary consumers and predators in a detritus-based stream. Limnology and Oceanography, 55 (6):2305–2316, November 2010. CODEN LIOCAH. ISSN 0024-

Durden:2017:DCF

[DRP^{+}17] Jennifer M. Durden, Henry A. Ruhl, Corinne Pebody, Sabena J. Blackbird, and Dick van Oevelen. Differences in the carbon flows in the benthic food webs of abyssal hill and plain habitats. Limnology and Oceanography, 62(5):1771–1782, July 2017. CODEN LIOCAH. ISSN 0024-3590.

Decima:2019:UER

[DSLLL19] Moira Décima, Michael R. Stukel, Lucía López-López, and Michael R. Landry. The unique ecological role of pyrosomes in the Eastern tropical Pacific. Limnology and Oceanography, 64 (2):728–743, March 2019. CODEN LIOCAH. ISSN 0024-3590.

Drake:2018:MGP

[DSM^{+}18] Jeana L. Drake, Morgan F. Schaller, Tali Mass, Linda Godfrey, Athena Fu, Robert M. Sherrell, Yair Rosenthal, and Paul G. Falkowski. Molecular and geochemical perspectives on the influence of CO_{2} on calcification in coral cell cultures. Limnology and Oceanography, 63(1):107–121, January 2018. CODEN LI-OCAH. ISSN 0024-3590.

Dong:2011:DRN

[DSS^{+}11] Liang F. Dong, Milika Naqasima Sobey, Cindy J. Smith, Iman Rusmana, Wayne Phillips, Andrew Stott, A. Mark Osborn, and David B. Nedwell. Dissimilatory reduction of nitrate to ammonium, not denitrification or anammox, dominates benthic nitrate reduction in tropical estuaries. Limnology and Oceanography, 56(1):279–291, January 2011. CODEN LIOCAH. ISSN 0024-3590.

Dahnke:2016:IFI

[DT16] Kirstin Dähnke and Bo Thamdrup. Isotope fractionation and isotope decoupling during anammox and denitrification in marine sediments. Limnology and Oceanography, 61(2):610–624, March 2016. CODEN LIOCAH. ISSN 0024-3590.

Dalsgaard:2012:ADO

[DTFR12] Tage Dalsgaard, Bo Thamdrup, Laura Farías, and Niels Peter Revsbech. Anammox and denitrification in the oxygen minimum zone of the eastern South Pacific. Limnology and Oceanography, 57(5):1331–1346, September 2012. CODEN LIOCAH. ISSN 0024-3590.

Dolhi:2015:DSD

[DTKMK15] Jenna M. Dolhi, Amber G. Teufel, Weidong Kong, and Rachael M. Morgan-Kiss. Diversity and spatial distribution of autotrophic communities within and between ice-covered Antarctic lakes (McMurdo Dry Valleys). Limnology and Oceanography, 60(3):977–991, May 2015. CODEN LIOCAH. ISSN 0024-3590.

Damashek:2019:MOS

[DTL^{+}19] Julian Damashek, Bradley B. Tolar, Qian Liu, Aimee O. Okotie-Oyekan, Natalie J. Wallsgrove, Brian N. Popp, and James T. Hollibaugh. Microbial oxidation of selected organic nitrogen compounds in the South Atlantic Bight. Limnology and Oceanography, 64(3):982–995, May 2019. CODEN LIOCAH. ISSN 0024-3590.

Duernberger:2018:ESN

[DTM18] Kimberley A. Duernberger, Craig R. Tobias, and Michael A. Mallin. Ecosystem scale nitrification and watershed support of tidal creek productivity revealed using whole system isotope tracer labeling. Limnology and Oceanography, 64(4):2110–2125, September 2018. CODEN LIOCAH. ISSN 0024-3590.

Daniels:2012:ILM

[DTPP12] Chris J. Daniels, Toby Tyrrell, Alex J. Poulton, and Laura Pettit. The influence of lithogenic material on particulate inorganic carbon measurements of coccolithophores in the Bay of Biscay. Limnology and Oceanography, 57(1):145–153, January 2012. CODEN LIOCAH. ISSN 0024-3590.

Danhiez:2017:OPC

[DVC^{+}17] F. P. Danhiez, V. Vantrepotte, A. Cauvin, E. Lebourg, and H. Loisel. Optical properties of chromophoric dissolved organic matter during a phytoplankton bloom. Implication for DOC estimates from CDOM absorption. Limnology and Oceanography, 62(4):1409–1425, July 2017. CODEN LIOCAH. ISSN 0024-3590.

Durkin:2016:SPA

[DVDB16] Colleen A. Durkin, Benjamin A. S. Van Mooy, Sonya T. Dyhrman, and Ken O. Buesseler. Sinking phytoplankton associated with carbon flux in the Atlantic Ocean. Limnology and Oceanography, 61(4):1172–1187, July 2016. CODEN LIOCAH. ISSN 0024-3590.

Dunlop:2016:CCD

[DvOR^{+}16] Katherine M. Dunlop, Dick van Oevelen, Henry A. Ruhl, Christine L. Huffard, Linda A. Kuhnz, and Kenneth L. Smith, Jr. Carbon cycling in the deep eastern North Pacific benthic food web: Investigating the effect of organic carbon input. Limnology and Oceanography, 61(6):1956–1968, November 2016. CODEN LIOCAH. ISSN 0024-3590.

Dixon:2013:HPD

[DVSV13] Taylor C. Dixon, Andrew W. Vermilyea, Durelle T. Scott, and Bettina M. Voelker. Hydrogen peroxide dynamics in an agricultural headwater stream: Evidence for significant nonphotochemical production. Limnology and Oceanography, 58(6):2133–2144, November 2013. CODEN LIOCAH. ISSN 0024-3590.

Dong:2010:CPO

[DWDH10] Hong-Po Dong, Da-Zhi Wang, Minhan Dai, and Hua-Sheng Hong. Characterization of particulate organic matter in the water column of the South China Sea using a shotgun proteomic approach. Limnology and Oceanography, 55(4):1565–1578, July 2010. CODEN LIOCAH. ISSN 0024-3590.

Evans:2012:VLM

[EB12] Claire Evans and Corina P. D. Brussaard. Viral lysis and microzooplankton grazing of phytoplankton throughout the Southern Ocean. Limnology and Oceanography, 57(6):1826–1837, November 2012. CODEN LIOCAH. ISSN 0024-3590.

Estapa:2012:RIO

[EBMR12] Margaret L. Estapa, Emmanuel Boss, Lawrence M. Mayer, and Collin S. Roesler. Role of iron and organic carbon in massspecific light absorption by particulate matter from Louisiana coastal waters. Limnology and Oceanography, 57(1):97–112, January 2012. CODEN LIOCAH. ISSN 0024-3590.

Edmunds:2011:ZAE

[Edm11] Peter J. Edmunds. Zooplanktivory ameliorates the effects of ocean acidification on the reef coral Porites spp. Limnology and Oceanography, 56(6):2402–2410, November 2011. CODEN LIOCAH. ISSN 0024-3590.

Edmunds:2015:QCD

[Edm15] P. J. Edmunds. A quarter-century demographic analysis of the Caribbean coral, Orbicella annularis, and projections of population size over the next century. Limnology and Oceanography, 60(3):840–855, May 2015. CODEN LIOCAH. ISSN 0024-3590.

Erler:2010:TSV

[EED10] Dirk V. Erler, Bradley D. Eyre, and Leigh Davison. Temporal and spatial variability in the cycling of nitrogen within a constructed wetland: a whole-system stable-isotope-addition experiment. Limnology and Oceanography, 55(3):1172–1187, May 2010. CODEN LIOCAH. ISSN 0024-3590.

Elliott:2010:DWF

[EHT10] David T. Elliott, Courtney K. Harris, and Kam W. Tang. Dead in the water: The fate of copepod carcasses in the York River estuary, Virginia. Limnology and Oceanography, 55(5):1821–1834, September 2010. CODEN LIOCAH. ISSN 0024-3590.

Ekvall:2015:DVM

[EHW^{+}15] Mikael T. Ekvall, Samuel Hylander, Tim Walles, Xi Yang, and Lars-Anders Hansson. Diel vertical migration, size distribution and photoprotection in zooplankton as response to UV-A radiation. Limnology and Oceanography, 60(6):2048–2058, November 2015. CODEN LIOCAH. ISSN 0024-3590.

Ejarque:2018:CIH

[EKS^{+}18] Elisabet Ejarque, Samiullah Khan, Gertraud Steniczka, Jakob Schelker, Martin J. Kainz, and Tom J. Battin. Climate-induced hydrological variation controls the transformation of dissolved organic matter in a subalpine lake. Limnology and Oceanography, 63(3):1355–1371, May 2018. CODEN LIOCAH. ISSN 0024-3590.

Edmunds:2016:EGV

[ELJ^{+}16] Peter J. Edmunds, James J. Leichter, Erika C. Johnston, Eric J. Tong, and Robert J. Toonen. Ecological and genetic variation in reef-building corals on four Society Islands. Limnology and Oceanography, 61(2):543–557, March 2016. CODEN LIOCAH. ISSN 0024-3590.

Ehn:2013:ALA

[EM13] Jens K. Ehn and C. J. Mundy. Assessment of light absorption within highly scattering bottom sea ice from under-ice light measurements: Implications for Arctic ice algae primary production. Limnology and Oceanography, 58(3):893–902, May 2013. CO-DEN LIOCAH. ISSN 0024-3590.

Estapa:2012:RAQ

[EMB12] Margaret L. Estapa, Lawrence M. Mayer, and Emmanuel Boss. Rate and apparent quantum yield of photodissolution of sedimentary organic matter. Limnology and Oceanography, 57(6): 1743–1756, November 2012. CODEN LIOCAH. ISSN 0024-3590.

Effler:2012:DSO

[EMH12] Steven W. Effler, David A. Matthews, and Craig A. Hurteau. Deposition signatures in Onondaga Lake, New York: Observations from a 28-year sediment-trap study. Limnology and Oceanography, 58(1):1531–1543, September 2012. CODEN LI-OCAH. ISSN 0024-3590.

Eyre:2011:DBM

[EMO^{+}11] Bradley D. Eyre, Damien Maher, Joanne M. Oakes, Dirk V. Erler, and Tim M. Glasby. Differences in benthic metabolism, nutrient fluxes, and denitrification in Caulerpa taxifolia communities compared to uninvaded bare sediment and seagrass (Zostera capricorni) habitats. Limnology and Oceanography, 56(5):1737– 1750, September 2011. CODEN LIOCAH. ISSN 0024-3590.

Eyre:2016:CDB

[EMS16] Bradley D. Eyre, Damien T. Maher, and Christian Sanders. The contribution of denitrification and burial to the nitrogen budgets of three geomorphically distinct Australian estuaries: Importance of seagrass habitats. Limnology and Oceanography, 61(3):1144–1156, May 2016. CODEN LIOCAH. ISSN 0024-3590.

Edman:2013:MDC

[EO13] Moa Edman and Anders Omstedt. Modeling the dissolved CO_{2} system in the redox environment of the Baltic Sea. Limnology and Oceanography, 58(2):74–92, January 2013. CODEN LI-OCAH. ISSN 0024-3590.

Eyre:2016:FMN

[EOM16] Bradley D. Eyre, Joanne M. Oakes, and Jack J. Middelburg. Fate of microphytobenthos nitrogen in subtropical subtidal sediments: A ^{15}N pulse-chase study. Limnology and Oceanography, 61(6):2108–2121, November 2016. CODEN LIOCAH. ISSN 0024-3590.

Effler:2014:LTS

[EP14] Steven W. Effler and Feng Peng. Long-term study of minerogenic particle optics in Cayuga Lake, New York. Limnology and Oceanography, 59(2):325–339, March 2014. CODEN LIOCAH. ISSN 0024-3590.

Ellis:2012:FCW

[ERA^{+}12] Erin E. Ellis, Jeffrey E. Richey, Anthony K. Aufdenkampe, Alex V. Krusche, Paul D. Quay, Cleber Salimon, and Hilandia Brandão da Cunha. Factors controlling water-column respiration in rivers of the central and southwestern Amazon Basin. Limnology and Oceanography, 57(3):527–540, March 2012. CO-DEN LIOCAH. ISSN 0024-3590.

Evans:2013:EEE

[ES13] Mary Anne Evans and Donald Scavia. Exploring estuarine eutrophication sensitivity to nutrient loading. Limnology and Oceanography, 58(2):569–578, March 2013. CODEN LIOCAH. ISSN 0024-3590.

Exton:2013:CNI

[ESMS13] D. A. Exton, D. J. Suggett, T. J. McGenity, and M. Steinke. Chlorophyll-normalized isoprene production in laboratory cultures of marine microalgae and implications for global models. Limnology and Oceanography, 58(4):1301–1311, July 2013. CO-DEN LIOCAH. ISSN 0024-3590.

Ezzat:2016:RBH

[ETI^{+}16] Leïla Ezzat, Erica Towle, Jean-Olivier Irisson, Chris Langdon, and Christine Ferrier-Pagès. The relationship between heterotrophic feeding and inorganic nutrient availability in the scleractinian coral T. reniformis under a short-term temperature increase. Limnology and Oceanography, 61(1):89–102, January 2016. CODEN LIOCAH. ISSN 0024-3590.

Edwards:2012:AST

[ETKL12] Kyle F. Edwards, Mridul K. Thomas, Christopher A. Klausmeier, and Elena Litchman. Allometric scaling and taxonomic variation in nutrient utilization traits and maximum growth rate of phytoplankton. Limnology and Oceanography, 57(3):554–566, March 2012. CODEN LIOCAH. ISSN 0024-3590.

Edwards:2015:LGM

[ETKL15] Kyle F. Edwards, Mridul K. Thomas, Christopher A. Klausmeier, and Elena Litchman. Light and growth in marine phytoplankton: allometric, taxonomic, and environmental variation. Limnology and Oceanography, 60(2):540–552, March 2015. CO-DEN LIOCAH. ISSN 0024-3590.

Edwards:2016:PGI

[ETKL16] Kyle F. Edwards, Mridul K. Thomas, Christopher A. Klausmeier, and Elena Litchman. Phytoplankton growth and the interaction of light and temperature: a synthesis at the species and community level. Limnology and Oceanography, 61(4):1232– 1244, July 2016. CODEN LIOCAH. ISSN 0024-3590.

Epstein:2012:NPT

[EWB12] David M. Epstein, Wayne A. Wurtsbaugh, and Michelle A. Baker. Nitrogen partitioning and transport through a subalpine lake measured with an isotope tracer. Limnology and Oceanography, 58(1):1503–1516, September 2012. CODEN LIOCAH. ISSN 0024-3590.

Fontaneto:2010:SNP

[FA10] Diego Fontaneto and Roberto Ambrosini. Spatial niche partitioning in epibiont rotifers on the waterlouse Asellus aquaticus. Limnology and Oceanography, 55(3):1327–1337, May 2010. CO-DEN LIOCAH. ISSN 0024-3590.

Franklin:2012:ISD

[FAF^{+}12] Daniel J. Franklin, Ruth L. Airs, Michelle Fernandes, Thomas G. Bell, Roy J. Bongaerts, John A. Berges, and Gill Malin. Identification of senescence and death in Emiliania huxleyi and Thalassiosira pseudonana: Cell staining, chlorophyll alterations, and dimethylsulfoniopropionate (DMSP) metabolism. Limnology and Oceanography, 57(1):305–317, January 2012. CODEN LIOCAH. ISSN 0024-3590.

Fichot:2012:SSC

[FB12] Cédric G. Fichot and Ronald Benner. The spectral slope coefficient of chromophoric dissolved organic matter (S_{275−−295}) as a tracer of terrigenous dissolved organic carbon in river-influenced ocean margins. Limnology and Oceanography, 58(1):1453–1466, September 2012. CODEN LIOCAH. ISSN 0024-3590.

Freimann:2013:RLM

[FBFR13] Remo Freimann, Helmut Bürgmann, Stuart E. G. Findlay, and Christopher T. Robinson. Response of lotic microbial communities to altered water source and nutritional state in a glaciated alpine floodplain. Limnology and Oceanography, 58(3):951–965, May 2013. CODEN LIOCAH. ISSN 0024-3590.

Fraisse:2015:TEP

[FBL15] Stéphane Fraisse, Myriam Bormans, and Yvan Lagadeuc. Turbulence effects on phytoplankton morphofunctional traits selection. Limnology and Oceanography, 60(3):872–884, May 2015. CODEN LIOCAH. ISSN 0024-3590.

Faithfull:2011:ENP

[FBV11] C. L. Faithfull, A.-K. Bergström, and T. Vrede. Effects of nutrients and physical lake characteristics on bacterial and phytoplankton production: a meta-analysis. Limnology and Oceanography, 56(5):1703–1713, September 2011. CODEN LIOCAH. ISSN 0024-3590.

Fry:2011:SSI

[FC11] Brian Fry and Matthew M. Chumchal. Sulfur stable isotope indicators of residency in estuarine fish. Limnology and Oceanography, 56(5):1563–1576, September 2011. CODEN LIOCAH. ISSN 0024-3590.

Findlay:2011:DPP

[FCC11] Helen S. Findlay, Piero Calosi, and Katharine Crawfurd. Determinants of the PIC : POC response in the coccolithophore Emiliania huxleyi under future ocean acidification scenarios. Limnology and Oceanography, 57(4):1168–1178, May 2011. CODEN LIOCAH. ISSN 0024-3590.

Friedman:2012:GSA

[FCD12] Jason R. Friedman, Nicole E. Condon, and Jeffrey C. Drazen. Gill surface area and metabolic enzyme activities of demersal fishes associated with the oxygen minimum zone off California. Limnology and Oceanography, 57(6):1701–1710, November 2012. CODEN LIOCAH. ISSN 0024-3590.

Fernandez-Carrera:2016:DWH

[FCRW^{+}16] A. Fernández-Carrera, K. L. Rogers, S. C. Weber, J. P. Chanton, and J. P. Montoya. Deep Water Horizon oil and methane carbon entered the food web in the Gulf of Mexico. Limnology and Oceanography, 61(S1):S387–S400, November 2016. CODEN LIOCAH. ISSN 0024-3590.

Foley:2015:RWQ

[FDB^{+}15] Melissa M. Foley, Jeffrey J. Duda, Matthew M. Beirne, Rebecca Paradis, Andrew Ritchie, and Jonathan A. Warrick. Rapid water quality change in the Elwha River estuary complex during dam removal. Limnology and Oceanography, 60(5):1719–1732, September 2015. CODEN LIOCAH. ISSN 0024-3590.

Fujii:2016:CCG

[FDBW16] M. Fujii, T. C. Dang, M. W. Bligh, and T. D. Waite. Cellular characteristics and growth behavior of iron-limited Microcystis aeruginosa in nutrient-depleted and nutrient-replete chemostat systems. Limnology and Oceanography, 61(6):2151–2164, November 2016. CODEN LIOCAH. ISSN 0024-3590.

Fulton:2014:STS

[FDH^{+}14] Christopher J. Fulton, Martial. Depczynski, Thomas H. Holmes, Mae M. Noble, Ben Radford, Thomas Wernberg, and Shaun K. Wilson. Sea temperature shapes seasonal fluctuations in seaweed biomass within the Ningaloo coral reef ecosystem. Limnology and Oceanography, 59(1):156–166, January 2014. CODEN LIOCAH. ISSN 0024-3590.

Findlay:2017:MPS

[FDL17] Alyssa J. Findlay, Dominic M. Di Toro, and George W. Luther III. A model of phototrophic sulfide oxidation in a stratified estuary. Limnology and Oceanography, 62(5):1853–1867, September 2017. CODEN LIOCAH. ISSN 0024-3590.

Faillettaz:2018:SSM

[FDP^{+}18] Robin Faillettaz, Elysanne Durand, Claire B. Paris, Philippe Koubbi, and Jean-Olivier Irisson. Swimming speeds of Mediterranean settlement-stage fish larvae nuance Hjort’s aberrant drift hypothesis. Limnology and Oceanography, 63(2):509–523, March 2018. CODEN LIOCAH. ISSN 0024-3590.

Fourquez:2014:EIL

[FDS^{+}14] Marion Fourquez, Aurélie Devez, Annick Schaumann, Audrey Guéneuguès, Thierry Jouenne, Ingrid Obernosterer, and Stéphane Blain. Effects of iron limitation on growth and carbon metabolism in oceanic and coastal heterotrophic bacteria. Limnology and Oceanography, 59(2):349–360, March 2014. CODEN LIOCAH. ISSN 0024-3590.

Fripiat:2018:IBS

[FDS^{+}18] F. Fripiat, M. Declercq, C. J. Sapart, L. G. Anderson, V. Bruechert, F. Deman, D. Fonseca-Batista, C. Humborg, A. Roukaerts, I. P. Semiletov, and F. Dehairs. Influence of the bordering shelves on nutrient distribution in the Arctic halocline inferred from water column nitrate isotopes. Limnology and Oceanography, 64(4):2154–2170, September 2018. CODEN LIOCAH. ISSN 0024-3590.

Freixa:2016:SMC

[FEC^{+}16] A. Freixa, E. Ejarque, S. Crognale, S. Amalfitano, S. Fazi, A. Butturini, and A. M. Romaní. Sediment microbial communities rely on different dissolved organic matter sources along a Mediterranean river continuum. Limnology and Oceanography, 61(4):1389–1405, July 2016. CODEN LIOCAH. ISSN 0024-3590.

Faber:2014:PWE

[FEW^{+}14] Peter A. Faber, Victor Evrard, Ryan J. Woodland, Ian C. Cartwright, and Perran L. M. Cook. Pore-water exchange driven by tidal pumping causes alkalinity export in two intertidal inlets. Limnology and Oceanography, 59(5):1749–1763, September 2014. CODEN LIOCAH. ISSN 0024-3590.

Fiksen:2013:TBM

[FFA13] Øyvind Fiksen, Michael J. Follows, and Dag L. Aksnes. Trait-based models of nutrient uptake in microbes extend the Michaelis–Menten framework. Limnology and Oceanography, 58 (2):193–202, January 2013. CODEN LIOCAH. ISSN 0024-3590.

Feehan:2018:KDP

[FGBS^{+}18] Colette J. Feehan, Beatrice C. Grauman-Boss, Richard R. Strathmann, Megan N. Dethier, and David O. Duggins. Kelp detritus provides high-quality food for sea urchin larvae. Limnology and Oceanography, 63(S1):S299–S306, March 2018. CODEN LIOCAH. ISSN 0024-3590.

Frank-Gopolos:2017:REC

[FGMN17] Thomas Frank-Gopolos, Eva Friis Møller, and Torkel Gissel Nielsen. The role of egg cannibalism for the Calanus succession in the Disko Bay, Western Greenland. Limnology and Oceanography, 62(3):865–883, May 2017. CODEN LIOCAH. ISSN 0024-3590.

Fellman:2015:EAA

[FHR^{+}15] Jason B. Fellman, Eran Hood, Peter A. Raymond, John Hudson, Maura Bozeman, and Mayumi Arimitsu. Evidence for the assimilation of ancient glacier organic carbon in a proglacial stream food web. Limnology and Oceanography, 60(4):1118–1128, July 2015. CODEN LIOCAH. ISSN 0024-3590.

Fellman:2010:FSO

[FHS10] Jason B. Fellman, Eran Hood, and Robert G. M. Spencer. Fluorescence spectroscopy opens new windows into dissolved organic matter dynamics in freshwater ecosystems: a review. Limnology and Oceanography, 55(6):2452–2462, November 2010. CODEN LIOCAH. ISSN 0024-3590.

Fielding:2013:EHS

[Fie13] Samuel R. Fielding. Emiliania huxleyi specific growth rate dependence on temperature. Limnology and Oceanography, 58(2): 663–666, March 2013. CODEN LIOCAH. ISSN 0024-3590.

Furman:2015:SRZ

[FJBP15] Bradley T. Furman, Lisa J. Jackson, Eric Bricker, and Bradley J. Peterson. Sexual recruitment in Zostera marina: a patch to landscape-scale investigation. Limnology and Oceanography, 60(2):584–599, March 2015. CODEN LIOCAH. ISSN 0024-3590.

Fridolfsson:2018:TVB

[FLLH18] Emil Fridolfsson, Elin Lindehoff, Catherine Legrand, and Samuel Hylander. Thiamin (vitamin B_{1}) content in phytoplankton and zooplankton in the presence of filamentous cyanobacteria. Limnology and Oceanography, 63(6):2423–2435, November 2018. CODEN LIOCAH. ISSN 0024-3590.

Foubert:2019:MES

[FLM^{+}19] Aline Foubert, Céline Le Pichon, Marc Mingelbier, John M. Farrell, Jean Morin, and Frédéric Lecomte. Modeling the effective spawning and nursery habitats of northern pike within a large spatiotemporally variable river landscape (St. Lawrence River, Canada). Limnology and Oceanography, 64(2):803–819, March 2019. CODEN LIOCAH. ISSN 0024-3590.

Finlay:2010:MCO

[FLP^{+}10] K. Finlay, P. R. Leavitt, A. Patoine, A. Patoine, and B. Wissel. Magnitudes and controls of organic and inorganic carbon flux through a chain of hard-water lakes on the northern Great Plains. Limnology and Oceanography, 55(4):1551–1564, July 2010. CODEN LIOCAH. ISSN 0024-3590.

Forrest:2013:CGI

[FLPL13] Alexander L. Forrest, Bernard E. Laval, Roger Pieters, and Darlene S. S. Lim. A cyclonic gyre in an ice-covered lake. Limnology and Oceanography, 58(3):363–375, January 2013. CODEN LI-OCAH. ISSN 0024-3590.

Fietz:2011:CPC

[FMGR^{+}11] Susanne Fietz, Alfredo Martínez-Garcia, Gemma Rueda, Vicky L. Peck, Carme Huguet, Marina Escala, and Antoni Rosell-Melé. Crenarchaea and phytoplankton coupling in sedimentary archives: Common trigger or metabolic dependence? Limnology and Oceanography, 56(5):1907–1916, September 2011. CODEN LIOCAH. ISSN 0024-3590.

Fujiki:2014:SCP

[FMM^{+}14] Tetsuichi Fujiki, Kazuhiko Matsumoto, Yoshihisa Mino, Kosei Sasaoka, Masahide Wakita, Hajime Kawakami, Makio C. Honda, Shuichi Watanabe, and Toshiro Saino. Seasonal cycle of phytoplankton community structure and photophysiological state in the western subarctic gyre of the North Pacific. Limnology and Oceanography, 59(3):887–900, May 2014. CODEN LIOCAH. ISSN 0024-3590.

Furian:2013:CDS

[FMP^{+}13] Sônia Furian, Elisângela Rosemeri Curti Martins, Tatiana Mascari Parizotto, Ary Tavares Rezende-Filho, Reynaldo Luiz Victoria, and Laurent Barbiero. Chemical diversity and spatial variability in myriad lakes in Nhecolândia in the Pantanal wetlands of Brazil. Limnology and Oceanography, 58(6):2249–2261, November 2013. CODEN LIOCAH. ISSN 0024-3590.

Freese:2015:SPE

[FNSS15] Daniela Freese, Barbara Niehoff, Janne E. Søreide, and Franz Josef Sartoris. Seasonal patterns in extracellular ion concentrations and pH of the Arctic copepod Calanus glacialis. Limnology and Oceanography, 60(6):2121–2129, November 2015. CODEN LIOCAH. ISSN 0024-3590.

Fischer:2015:DVM

[FOT^{+}15] Janet M. Fischer, Mark H. Olson, Nora Theodore, Craig E. Williamson, Kevin C. Rose, and Jin Hwang. Diel vertical migration of copepods in mountain lakes: The changing role of ultraviolet radiation across a transparency gradient. Limnology and Oceanography, 60(1):252–262, January 2015. CODEN LIOCAH. ISSN 0024-3590.

Finlay:2010:EEP

[FPD^{+}10] Kerri Finlay, Alain Patoine, Derek B. Donald, Matthew J. Bogard, and Peter R. Leavitt. Experimental evidence that pollution with urea can degrade water quality in phosphorus-rich lakes of the Northern Great Plains. Limnology and Oceanography, 55 (3):1213–1230, May 2010. CODEN LIOCAH. ISSN 0024-3590.

Fellman:2011:SBC

[FPG11] Jason B. Fellman, Kevin C. Petrone, and Pauline F. Grierson. Source, biogeochemical cycling, and fluorescence characteristics of dissolved organic matter in an agro-urban estuary. Limnology and Oceanography, 56(1):243–256, January 2011. CODEN LIOCAH. ISSN 0024-3590.

Ferrier-Pages:2013:SAD

[FPGR^{+}13] C. Ferrier-Pagès, F. Gevaert, S. Reynaud, E. Beraud, D. Menu, M.-A. Janquin, S. Cocito, and A. Peirano. In situ assessment of the daily primary production of the temperate symbiotic coral Cladocora caespitosa. Limnology and Oceanography, 58(4):1409– 1418, July 2013. CODEN LIOCAH. ISSN 0024-3590.

Fragoso:2019:TBA

[FPP^{+}19] Glaucia Moreira Fragoso, Alex James Poulton, Nicola Jane Pratt, Geir Johnsen, and Duncan Alastair Purdie. Trait-based analysis of subpolar North Atlantic phytoplankton and plastidic ciliate communities using automated flow cytometer. Limnology and Oceanography, 64(4):1763–1778, July 2019. CODEN LIOCAH. ISSN 0024-3590.

Ferrier-Pages:2011:SAW

[FPPA^{+}11] C. Ferrier-Pagès, A. Peirano, M. Abbate, S. Cocito, A. Negri, C. Rottier, P. Riera, R. Rodolfo-Metalpa, and S. Reynaud. Summer autotrophy and winter heterotrophy in the temperate symbiotic coral Cladocora caespitosa. Limnology and Oceanography, 56(4):1429–1438, July 2011. CODEN LIOCAH. ISSN 0024-3590.

Franze:2018:DPA

[FPSL18] Gayantonia Franzè, James J. Pierson, Diane K. Stoecker, and Peter J. Lavrentyev. Diatom-produced allelochemicals trigger trophic cascades in the planktonic food web. Limnology and Oceanography, 63(3):1093–1108, May 2018. CODEN LIOCAH. ISSN 0024-3590.

Feng:2017:ECG

[FRA^{+}17] Yuanyuan Feng, Michael Y. Roleda, Evelyn Armstrong, Philip W. Boyd, and Catriona L. Hurd. Environmental controls on the growth, photosynthetic and calcification rates of a Southern Hemisphere strain of the coccolithophore Emiliania huxleyi. Limnology and Oceanography, 62(2):519–540, March 2017. CODEN LIOCAH. ISSN 0024-3590.

Fujimura:2014:NSL

[FRP^{+}14] Atsushi G. Fujimura, Ad J. H. M. Reniers, Claire B. Paris, Alan L. Shanks, Jamie H. MacMahan, and Steven G. Morgan. Numerical simulations of larval transport into a rip-channeled surf zone. Limnology and Oceanography, 59(4):1434–1447, July 2014. CODEN LIOCAH. ISSN 0024-3590.

Feehan:2016:MEL

[FSBT16] Colette J. Feehan, Robert E. Scheibling, Michael S. Brown, and Keith R. Thompson. Marine epizootics linked to storms: Mechanisms of pathogen introduction and persistence inferred from coupled physical and biological time-series. Limnology and Oceanography, 61(1):316–329, January 2016. CODEN LIOCAH. ISSN 0024-3590.

Fergus:2011:MLW

[FSCB11] C. Emi Fergus, Patricia A. Soranno, Kendra Spence Cheruvelil, and Mary T. Bremigan. Multiscale landscape and wetland drivers of lake total phosphorus and water color. Limnology and Oceanography, 56(6):2127–2146, November 2011. CODEN LI-OCAH. ISSN 0024-3590.

Faustova:2011:REE

[FSST11] Markéta Faustova, Veronika Sacherovà, Jan-Erik Svensson, and Derek J. Taylor. Radiation of European Eubosmina (Cladocera) from Bosmina (E.) longispina — concordance of multipopulation molecular data with paleolimnology. Limnology and Oceanography, 56(2):440–450, March 2011. CODEN LIOCAH. ISSN 0024-3590.

Freeman:2011:CIB

[FT11] Christopher J. Freeman and Robert W. Thacker. Complex interactions between marine sponges and their symbiotic microbial communities. Limnology and Oceanography, 56(5):1577–1586, September 2011. CODEN LIOCAH. ISSN 0024-3590.

Fowler:2010:SRB

[FTC10] Scott W. Fowler, Jean-Louis Teyssie, and Thomas M. Church. Scavenging and retention of bismuth by marine plankton and biogenic particles. Limnology and Oceanography, 55(3):1093– 1104, May 2010. CODEN LIOCAH. ISSN 0024-3590.

Fasching:2016:HCD

[FUS^{+}16] Christina Fasching, Amber J. Ulseth, Jakob Schelker, Gertraud Steniczka, and Tom J. Battin. Hydrology controls dissolved organic matter export and composition in an alpine stream and its hyporheic zone. Limnology and Oceanography, 61(2):558–571, March 2016. CODEN LIOCAH. ISSN 0024-3590.

Finlay:2019:SPH

[FVSL19] Kerri Finlay, Richard J. Vogt, Gavin L. Simpson, and Peter R. Leavitt. Seasonality of pCO_{2} in a hard-water lake of the northern Great Plains: The legacy effects of climate and limnological conditions over 36 years. Limnology and Oceanography, 64(S1): S118–S129, January 2019. CODEN LIOCAH. ISSN 0024-3590.

Felden:2010:TCO

[FWFB10] J. Felden, F. Wenzhöfer, T. Feseker, and A. Boetius. Transport and consumption of oxygen and methane in different habitats of the Håkon Mosby Mud Volcano (HMMV). Limnology and Oceanography, 55(6):2366–2380, November 2010. CODEN LI-OCAH. ISSN 0024-3590.

Filstrup:2018:ERN

[FWO^{+}18] Christopher T. Filstrup, Tyler Wagner, Samantha K. Oliver, Craig A. Stow, Katherine E. Webster, Emily H. Stanley, and John A. Downing. Evidence for regional nitrogen stress on chlorophyll-a in lakes across large landscape and climate gradients. Limnology and Oceanography, 63(S1):S324–S339, March 2018. CODEN LIOCAH. ISSN 0024-3590.

Filstrup:2014:RVA

[FWS^{+}14] Christopher T. Filstrup, Tyler Wagner, Patricia A. Soranno, Emily H. Stanley, Craig A. Stow, Katherine E. Webster, and John A. Downing. Regional variability among nonlinear chlorophyll–phosphorus relationships in lakes. Limnology and Oceanography, 59(5):1691–1703, September 2014. CODEN LI-OCAH. ISSN 0024-3590.

Frenken:2018:FPT

[FWvD^{+}18] Thijs Frenken, Joren Wierenga, Ellen van Donk, Steven A. J. Declerck, Lisette N. de Senerpont Domis, Thomas Rohrlack, and Dedmer B. Van de Waal. Fungal parasites of a toxic inedible cyanobacterium provide food to zooplankton. Limnology and Oceanography, 63(6):2384–2393, November 2018. CODEN LIOCAH. ISSN 0024-3590.

Follett:2018:NFR

[FWWF18] Christopher L. Follett, Angelicque E. White, Samuel T. Wilson, and Michael J. Follows. Nitrogen fixation rates diagnosed from diurnal changes in elemental stoichiometry. Limnology and Oceanography, 63(5):1911–1923, September 2018. CODEN LI-OCAH. ISSN 0024-3590.

Fang:2018:TDB

[FYC^{+}18] Ziming Fang, Weifeng Yang, Min Chen, Aron Stubbins, Haoyang Ma, Renming Jia, Qi Li, and Qianna Chen. Transport of dissolved black carbon from the Prydz Bay Shelf, Antarctica to the deep Southern Ocean. Limnology and Oceanography, 64 (4):2179–2190, September 2018. CODEN LIOCAH. ISSN 0024-3590.

Fujimori:2012:MHP

[FYT^{+}12] Takuya Fujimori, Yuki Yoneyama, Gen Taniai, Michiko Kurihara, Hideyuki Tamegai, and Shinya Hashimoto. Methyl halide production by cultures of marine proteobacteria Erythrobacter and Pseudomonas and isolated bacteria from brackish water. Limnology and Oceanography, 57(1):154–162, January 2012. CODEN LIOCAH. ISSN 0024-3590.

Fischer:2017:DBM

[FYVU17] Sarah J. Fischer, Joanna K. York, Yoana G. Voynova, and William J. Ullman. Distance-based mixing models of δ ^{18}NNO_{3}^{−}

and δ ^{18}ONO_{3}^{−} in a marsh-lined estuary with multiple, distinct

sources (Murderkill Estuary, Delaware, USA). Limnology and Oceanography, 62(2):408–420, March 2017. CODEN LIOCAH. ISSN 0024-3590.

Falter:2014:ADS

[FZL^{+}14] James L. Falter, Zhenlin Zhang, Ryan J. Lowe, Frazer McGregor, John Keesing, and Malcolm T. McCulloch. Assessing the drivers of spatial variation in thermal forcing across a nearshore reef system and implications for coral bleaching. Limnology and Oceanography, 59(4):1241–1255, July 2014. CODEN LIOCAH. ISSN 0024-3590.

Gernez:2011:DCP

[GAH11] Pierre Gernez, David Antoine, and Yannick Huot. Diel cycles of the particulate beam attenuation coefficient under varying trophic conditions in the northwestern Mediterranean Sea: Observations and modeling. Limnology and Oceanography, 56(1): 17–36, January 2011. CODEN LIOCAH. ISSN 0024-3590.

Gutt:2019:BCT

[GAK^{+}19] Julian Gutt, Janina Arndt, Casper Kraan, Boris Dorschel, Michael Schröder, Astrid Bracher, and Dieter Piepenburg. Benthic communities and their drivers: a spatial analysis off the Antarctic Peninsula. Limnology and Oceanography, 64(6):2341– 2357, November 2019. CODEN LIOCAH. ISSN 0024-3590.

Grasset:2019:TMD

[GAM^{+}19] Charlotte Grasset, Gwenaël Abril, Raquel Mendonça, Fabio Roland, and Sebastian Sobek. The transformation of macrophyte-derived organic matter to methane relates to plant water and nutrient contents. Limnology and Oceanography, 64 (4):1737–1749, July 2019. CODEN LIOCAH. ISSN 0024-3590.

Guo:2018:FSA

[GBB^{+}18] Fen Guo, Stuart E. Bunn, Michael T. Brett, Brian Fry, Hannes Hager, Xiaoguang Ouyang, and Martin J. Kainz. Feeding strategies for the acquisition of high-quality food sources in stream macroinvertebrates: Collecting, integrating, and mixed feeding. Limnology and Oceanography, 63(5):1964–1978, September 2018. CODEN LIOCAH. ISSN 0024-3590. See erratum [Ano19c].

Grebner:2019:IDT

[GBB^{+}19a] Wiebke Grebner, E. Carina Berglund, Fredrik Berggren, Johan Eklund, Sara Hardadóttir, Mats X. Andersson, and Erik Selander. Induction of defensive traits in marine plankton — new copepodamide structures. Limnology and Oceanography, 64(2): 820–831, March 2019. CODEN LIOCAH. ISSN 0024-3590.

Grosse:2019:NLD

[GBB19b] J. Grosse, C. P. D. Brussaard, and H. T. S. Boschker. Nutrient limitation driven dynamics of amino acids and fatty acids in coastal phytoplankton. Limnology and Oceanography, 64(1): 302–316, January 2019. CODEN LIOCAH. ISSN 0024-3590.

Guo:2019:III

[GBB^{+}19c] Fen Guo, Stuart E. Bunn, Michael T. Brett, Brian Fry, Hannes Hager, Xiaoguang Ouyang, and Martin J. Kainz. Issue information — instr to contrib. Limnology and Oceanography, 63(5): i–iii, July 2019. CODEN LIOCAH. ISSN 0024-3590.

Gradoville:2017:DAN

[GBC^{+}17] Mary R. Gradoville, Deniz Bombar, Byron C. Crump, Ricardo M. Letelier, Jonathan P. Zehr, and Angelicque E. White. Diversity and activity of nitrogen-fixing communities across ocean basins. Limnology and Oceanography, 62(5):1895–1909, September 2017. CODEN LIOCAH. ISSN 0024-3590.

Grosse:2010:MRP

[GBD^{+}10] Julia Grosse, Deniz Bombar, Hai Nhu Doan, Lam Ngoc Nguyen, and Maren Voss. The Mekong River plume fuels nitrogen fixation and determines phytoplankton species distribution in the South China Sea during low and high discharge season. Limnology and Oceanography, 55(4):1668–1680, July 2010. CODEN LIOCAH. ISSN 0024-3590.

Gliwicz:2018:WIN

[GBK^{+}18] Z. Maciej Gliwicz, Ewa Babkiewicz, Rajeev Kumar, Selvaraj Kunjiappan, and Konrad Leniowski. Warming increases the number of apparent prey in reaction field volume of zooplanktivorous fish. Limnology and Oceanography, 63(S1):S30–S43, March 2018. CODEN LIOCAH. ISSN 0024-3590.

Grant:2013:SSB

[GBL13] Scott R. Grant, Paul K. Bienfang, and Edward A. Laws. Steadystate bioassay approach applied to phosphorus-limited continuous cultures: a growth study of the marine chlorophyte Dunaliella salina. Limnology and Oceanography, 58(2):314–324, January 2013. CODEN LIOCAH. ISSN 0024-3590.

Guizien:2012:ULD

[GBMG12] K. Guizien, M. Belharet, P. Marsaleix, and J. M. Guarini. Using larval dispersal simulations for marine protected area design: Application to the Gulf of Lions (northwest Mediterranean). Limnology and Oceanography, 57(4):1099–1112, July 2012. CO-DEN LIOCAH. ISSN 0024-3590.

Gudasz:2012:CMP

[GBP^{+}12] Cristian Gudasz, David Bastviken, Katrin Premke, Kristin Steger, and Lars J. Tranvik. Constrained microbial processing of allochthonous organic carbon in boreal lake sediments. Limnology and Oceanography, 57(1):163–175, January 2012. CODEN LIOCAH. ISSN 0024-3590.

Gustafsson:2014:MPD

[GBR14] Malin S. M. Gustafsson, Mark E. Baird, and Peter J. Ralph. Modeling photoinhibition-driven bleaching in Scleractinian coral as a function of light, temperature, and heterotrophy. Limnology and Oceanography, 59(6):603–622, March 2014. CODEN LIOCAH. ISSN 0024-3590.

Guillemette:2017:OBY

[GBS17] François Guillemette, Thomas S. Bianchi, and Robert G. M. Spencer. Old before your time: Ancient carbon incorporation in contemporary aquatic foodwebs. Limnology and Oceanography, 62(4):1682–1700, July 2017. CODEN LIOCAH. ISSN 0024-3590.

Gosnell:2017:STT

[GBT^{+}17] Kathleen J. Gosnell, Prentiss H. Balcom, Craig R. Tobias, William P. Gilhooly III, and Robert P. Mason. Spatial and temporal trophic transfer dynamics of mercury and methylmercury into zooplankton and phytoplankton of Long Island Sound. Limnology and Oceanography, 63(3):1122–1138, May 2017. CO-DEN LIOCAH. ISSN 0024-3590.

Giannini:2016:PNP

[GC16] Maria Fernanda Colo Giannini and Áurea Maria Ciotti. Parameterization of natural phytoplankton photo-physiology: Effects of cell size and nutrient concentration. Limnology and Oceanography, 61(4):1495–1512, July 2016. CODEN LIOCAH. ISSN 0024-3590.

Guo:2012:CDC

[GCH^{+}12] Xianghui Guo, Wei-Jun Cai, Wei-Jen Huang, Yongchen Wang, Feizhou Chen, Michael C. Murrell, Steven E. Lohrenz, Li-Qing Jiang, Minhan Dai, Justin Hartmann, Qi Lin, and Randy Culp. Carbon dynamics and community production in the Mississippi River plume. Limnology and Oceanography, 57(1):1–17, January 2012. CODEN LIOCAH. ISSN 0024-3590.

Gloeckler:2018:SIA

[GCH^{+}18] Kristen Gloeckler, C. Anela Choy, Cecelia C. S. Hannides, Hilary G. Close, Erica Goetze, Brian N. Popp, and Jeffrey C. Drazen. Stable isotope analysis of micronekton around Hawaii reveals suspended particles are an important nutritional source in the lower mesopelagic and upper bathypelagic zones. Limnology and Oceanography, 63(3):1168–1180, May 2018. CODEN LIOCAH. ISSN 0024-3590.

Gihring:2010:DSS

[GCR^{+}10] Thomas M. Gihring, Andy Canion, Ashley Riggs, Markus Huettel, and Joel E. Kostka. Denitrification in shallow, sublittoral Gulf of Mexico permeable sediments. Limnology and Oceanography, 55(1):43–54, January 2010. CODEN LIOCAH. ISSN 0024-3590.

Gao:2014:ICB

[GCSO14] Yonghui Gao, Jeffrey C. Cornwell, Diane K. Stoecker, and Michael S. Owens. Influence of cyanobacteria blooms on sediment biogeochemistry and nutrient fluxes. Limnology and Oceanography, 59(3):959–971, May 2014. CODEN LIOCAH. ISSN 0024-3590.

Gall:2013:EVC

[GDCM13] Mark P. Gall, Rob J. Davies-Colley, and Rob A. Merrilees. Exceptional visual clarity and optical purity in a sub-alpine lake. Limnology and Oceanography, 58(2):443–451, March 2013. CO-DEN LIOCAH. ISSN 0024-3590.

Georgian:2016:OPC

[GDD^{+}16] Samuel E. Georgian, Danielle DeLeo, Alanna Durkin, Carlos E. Gomez, Melissa Kurman, Jay J. Lunden, and Erik E. Cordes. Oceanographic patterns and carbonate chemistry in the vicinity of cold-water coral reefs in the Gulf of Mexico: Implications for resilience in a changing ocean. Limnology and Oceanography, 61 (2):648–665, March 2016. CODEN LIOCAH. ISSN 0024-3590.

Guillemette:2011:RVF

[GdG11] François Guillemette and Paul A. del Giorgio. Reconstructing the various facets of dissolved organic carbon bioavailability in freshwater ecosystems. Limnology and Oceanography, 56(2):734– 748, March 2011. CODEN LIOCAH. ISSN 0024-3590.

Genovesi:2011:RCB

[GdVT^{+}11] Linda Genovesi, Anne de Vernal, Benoît Thibodeau, Claude Hillaire-Marcel, Alfonso Mucci, and Denis Gilbert. Recent changes in bottom water oxygenation and temperature in the Gulf of St. Lawrence: Micropaleontological and geochemical evidence. Limnology and Oceanography, 56(4):1319–1329, July 2011. CODEN LIOCAH. ISSN 0024-3590.

Giraldo:2017:DGR

[GEC^{+}17] Carolina Giraldo, Bruno Ernande, Pierre Cresson, Dorothée Kopp, Marie Cachera, Morgane Travers-Trolet, and Sébastien Lefebvre. Depth gradient in the resource use of a fish community from a semi-enclosed sea. Limnology and Oceanography, 65 (9):2213–2226, September 2017. CODEN LIOCAH. ISSN 0024-3590.

Grippo:2011:SVB

[GFDC11] Mark A. Grippo, John W. Fleeger, Stanislas F. Dubois, and Richard Condrey. Spatial variation in basal resources supporting benthic food webs revealed for the inner continental shelf. Limnology and Oceanography, 56(3):841–856, May 2011. CO-DEN LIOCAH. ISSN 0024-3590.

Garcia:2013:CUP

[GFH13] Nathan S. Garcia, Fei-Xue Fu, and David A. Hutchins. Colimitation of the unicellular photosynthetic diazotroph Crocosphaera watsonii by phosphorus, light, and carbon dioxide. Limnology and Oceanography, 58(4):1501–1512, July 2013. CODEN LI-OCAH. ISSN 0024-3590.

Godinot:2013:APA

[GFPSG13] C. Godinot, C. Ferrier-Pagès, S. Sikorski, and R. Grover. Alkaline phosphatase activity of reef-building corals. Limnology and Oceanography, 58(2):227–234, January 2013. CODEN LIOCAH. ISSN 0024-3590.

Galan:2014:TDN

[GFT^{+}14] Alexander Galán, Juan Faúndez, Bo Thamdrup, Juan Francisco Santibáñez, and Laura Farías. Temporal dynamics of nitrogen loss in the coastal upwelling ecosystem off central Chile: Evidence of autotrophic denitrification through sulfide oxidation. Limnology and Oceanography, 59(6):1865–1878, November 2014. CODEN LIOCAH. ISSN 0024-3590.

Garside:2014:FCC

[GGC^{+}14] C. J. Garside, T. M. Glasby, M. A. Coleman, B. P. Kelaher, and M. J. Bishop. The frequency of connection of coastal water bodies to the ocean predicts Carcinus maenas invasion. Limnology and Oceanography, 59(4):1288–1296, July 2014. CODEN LIOCAH. ISSN 0024-3590.

Glud:2015:CCM

[GGL^{+}15] Ronnie N. Glud, Hans-Peter Grossart, Morten Larsen, Kam W. Tang, Kristine E. Arendt, Søren Rysgaard, Bo Thamdrup, and Torkel Gissel Nielsen. Copepod carcasses as microbial hot spots for pelagic denitrification. Limnology and Oceanography, 60(6): 2026–2036, November 2015. CODEN LIOCAH. ISSN 0024-

Glass:2018:PMA

[GGL^{+}18] Emma M. Glass, Juan L. Garzon, Seth Lawler, Eleonore Paquier, and Celso M. Ferreira. Potential of marshes to attenuate storm surge water level in the Chesapeake Bay. Limnology and Oceanography, 63(2):951–967, March 2018. CODEN LIOCAH. ISSN 0024-3590.

Galand:2010:IAR

[GGPM^{+}10] Pierre E. Galand, Carmen Gutiérrez-Provecho, Ramon Massana, Josep M. Gasol, and Emilio O. Casamayor. Inter-annual recurrence of archaeal assemblages in the coastal NW Mediterranean Sea (Blanes Bay Microbial Observatory). Limnology and Oceanography, 55(5):2117–2125, September 2010. CODEN LI-OCAH. ISSN 0024-3590.

Gonzalez-Gil:2018:WMP

[GGTC^{+}18] Ricardo González-Gil, Fernando González Taboada, Carlos Cáceres, John L. Largier, and Ricardo Anadón. Winter-mixing preconditioning of the spring phytoplankton bloom in the Bay of Biscay. Limnology and Oceanography, 63(3):1264–1282, May 2018. CODEN LIOCAH. ISSN 0024-3590.

Grantz:2014:SIR

[GHS14] Erin M. Grantz, Brian E. Haggard, and J. Thad Scott. Stoichiometric imbalance in rates of nitrogen and phosphorus retention, storage, and recycling can perpetuate nitrogen deficiency in highly-productive reservoirs. Limnology and Oceanography, 61 (1):2203–2216, November 2014. CODEN LIOCAH. ISSN 0024-3590.

Guy-Haim:2016:CTR

[GHSR^{+}16] Tamar Guy-Haim, Jacob Silverman, Stefanie Raddatz, Martin Wahl, Alvaro Israel, and Gil Rilov. The carbon turnover response to thermal stress of a dominant coralline alga on the fast warming Levant coast. Limnology and Oceanography, 61 (3):1120–1133, May 2016. CODEN LIOCAH. ISSN 0024-3590.

Green:2019:NFI

[GJR^{+}19] Rebecca H. Green, Nicole L. Jones, Matthew D. Rayson, Ryan J. Lowe, Cynthia E. Bluteau, and Gregory N. Ivey. Nutrient fluxes into an isolated coral reef atoll by tidally driven internal bores. Limnology and Oceanography, 64(2):461–473, March 2019. CO-DEN LIOCAH. ISSN 0024-3590.

Godwin:2014:DOC

[GJWS14] Sean C. Godwin, Stuart E. Jones, Brian C. Weidel, and Christopher T. Solomon. Dissolved organic carbon concentration controls benthic primary production: Results from in situ chambers in north-temperate lakes. Limnology and Oceanography, 59(6): 2112–2120, November 2014. CODEN LIOCAH. ISSN 0024-3590.

Godwin:2016:EDO

[GJWS16] Sean C. Godwin, Stuart E. Jones, Brian C. Weidel, and Christopher T. Solomon. Erratum dissolved organic carbon concentration controls benthic primary production: Results from in situ chambers in north-temperate lakes. Limnology and Oceanography, 61(1):407, January 2016. CODEN LIOCAH. ISSN 0024-3590.

Gruber:2010:FEC

[GK10] Renee K. Gruber and W. Michael Kemp. Feedback effects in a coastal canopy-forming submersed plant bed. Limnology and Oceanography, 55(6):2285–2298, November 2010. CODEN LI-OCAH. ISSN 0024-3590.

Gurbisz:2014:URL

[GK14] Cassie Gurbisz and W. Michael Kemp. Unexpected resurgence of a large submersed plant bed in Chesapeake Bay: Analysis of time series data. Limnology and Oceanography, 59(2):482–494, March 2014. CODEN LIOCAH. ISSN 0024-3590.

Goncalves:2015:PAH

[GK15] Rodrigo J. Gonçalves and Thomas Kiørboe. Perceiving the algae: How feeding-current feeding copepods detect their nonmotile prey. Limnology and Oceanography, 60(4):1286–1297, July 2015. CODEN LIOCAH. ISSN 0024-3590. See reply [PJ16].

Grantz:2012:PWL

[GKS12] Erin M. Grantz, Aki Kogo, and J. Thad Scott. Partitioning whole-lake denitrification using in situ dinitrogen gas accumulation and intact sediment core experiments. Limnology and Oceanography, 57(4):925–935, July 2012. CODEN LIOCAH. ISSN 0024-3590.

Gladyshev:2015:DOM

[GKT^{+}15] Michail I. Gladyshev, Olesia V. Kolmakova, Alexander P. Tolomeev, Olesia V. Anishchenko, Olesia N. Makhutova, Anzhelika A. Kolmakova, Elena S. Kravchuk, Larisa A. Glushchenko, Vladimir I. Kolmakov, and Nadezhda N. Sushchik. Differences in organic matter and bacterioplankton between sections of the largest Arctic river: Mosaic or continuum? Limnology and Oceanography, 60(4):1314–1331, July 2015. CODEN LIOCAH. ISSN 0024-3590.

Gruber:2017:MTD

[GLF17] Renee K. Gruber, Ryan J. Lowe, and James L. Falter. Metabolism of a tide-dominated reef platform subject to extreme diel temperature and oxygen variations. Limnology and Oceanography, 62(4):1701–1717, July 2017. CODEN LIOCAH. ISSN 0024-3590.

Gruber:2018:BUP

[GLF18] Renee K. Gruber, Ryan J. Lowe, and James L. Falter. Benthic uptake of phytoplankton and ocean-reef exchange of particulate nutrients on a tide-dominated reef. Limnology and Oceanography, 63(4):1545–1561, July 2018. CODEN LIOCAH. ISSN 0024-3590.

Gao:2015:NDA

[GLI^{+}15] Lei Gao, Daoji Li, Joji Ishizaka, Yanwei Zhang, Haibo Zong, and Laodong Guo. Nutrient dynamics across the river–sea interface in the Changjiang (Yangtze River) estuary — East China Sea region. Limnology and Oceanography, 60(6):2207–2221, November 2015. CODEN LIOCAH. ISSN 0024-3590.

Gihring:2010:DDN

[GLKK10] Thomas M. Gihring, Gaute Lavik, Marcel M. M. Kuypers, and Joel E. Kostka. Direct determination of nitrogen cycling rates and pathways in Arctic fjord sediments (Svalbard, Norway). Limnology and Oceanography, 55(2):740–752, March 2010. CO-DEN LIOCAH. ISSN 0024-3590.

George:2015:IRT

[GLMG15] Jennifer A. George, Darcy J. Lonsdale, Lucas R. Merlo, and Christopher J. Gobler. The interactive roles of temperature, nutrients, and zooplankton grazing in controlling the winter– spring phytoplankton bloom in a temperate, coastal ecosystem, Long Island Sound. Limnology and Oceanography, 60(1):110– 126, January 2015. CODEN LIOCAH. ISSN 0024-3590.

Galloway:2013:FAS

[GLS^{+}13] A. W. E. Galloway, A. T. Lowe, E. A. Sosik, J. S. Yeung, and D. O. Duggins. Fatty acid and stable isotope biomarkers suggest microbe-induced differences in benthic food webs between depths. Limnology and Oceanography, 58(4):1451–1462, July 2013. CODEN LIOCAH. ISSN 0024-3590.

Gillanders:2012:HWP

[GM12] Bronwyn M. Gillanders and Andrew R. Munro. Hypersaline waters pose new challenges for reconstructing environmental histories of fish based on otolith chemistry. Limnology and Oceanography, 57(4):1136–1148, July 2012. CODEN LIOCAH. ISSN 0024-3590.

Gelesh:2016:MCI

[GMBL16] Lauren Gelesh, Kathleen Marshall, William Boicourt, and Laura Lapham. Methane concentrations increase in bottom waters during summertime anoxia in the highly eutrophic estuary, Chesapeake Bay, U.S.A. Limnology and Oceanography, 61(S1): S253–S266, November 2016. CODEN LIOCAH. ISSN 0024-3590.

Garthe:2011:IAC

[GMD11] Stefan Garthe, William A. Montevecchi, and Gail K. Davoren. Inter-annual changes in prey fields trigger different foraging tactics in a large marine predator. Limnology and Oceanography, 56(3):802–812, May 2011. CODEN LIOCAH. ISSN 0024-3590.

Greene:2013:RCF

[GMGM^{+}13] Charles H. Greene, Erin Meyer-Gutbrod, Bruce C. Monger, Louise P. McGarry, Andrew J. Pershing, Igor M. Belkin, Paula S. Fratantoni, David G. Mountain, Robert S. Pickart, Andrey Proshutinsky, Rubao Ji, James J. Bisagni, Sirpa M. A. Hakkinen, Dale B. Haidvogel, Jia Wang, Erica Head, Peter Smith, Philip C. Reid, and Alessandra Conversi. Remote climate forcing of decadal-scale regime shifts in Northwest Atlantic shelf ecosystems. Limnology and Oceanography, 58(3):803–816, May 2013. CODEN LIOCAH. ISSN 0024-3590.

Gliwicz:2013:PEP

[GMJW13] Z. Maciej Gliwicz, Piotr Maszczyk, Jedrzej Jabłoński, and Dariusz Wrzosek. Patch exploitation by planktivorous fish and the concept of aggregation as an antipredation defense in zooplankton. Limnology and Oceanography, 58(5):1621–1639, September 2013. CODEN LIOCAH. ISSN 0024-3590.

Gao:2012:IEN

[GML^{+}12] Hang Gao, Maciej Matyka, Bo Liu, Arzhang Khalili, Joel E. Kostka, Gavin Collins, Stefan Jansen, Moritz Holtappels, Marlene M. Jensen, Thomas H. Badewien, Melanie Beck, Maik Grunwald, Dirk de Beer, Gaute Lavik, and Marcel M. M. Kuypers. Intensive and extensive nitrogen loss from intertidal permeable sediments of the Wadden Sea. Limnology and Oceanography, 57(1):185–198, January 2012. CODEN LIOCAH. ISSN 0024-3590.

Glibert:2019:SIT

[GMMV19] Patricia M. Glibert, Jack J. Middelburg, James W. McClelland, and M. Jake Vander Zanden. Stable isotope tracers: Enriching our perspectives and questions on sources, fates, rates, and pathways of major elements in aquatic systems. Limnology and Oceanography, 64(3):950–981, May 2019. CODEN LIOCAH. ISSN 0024-3590.

Grasset:2018:LVM

[GMS^{+}18] Charlotte Grasset, Raquel Mendonça, Gabriella Villamor Saucedo, David Bastviken, Fabio Roland, and Sebastian Sobek. Large but variable methane production in anoxic freshwater sediment upon addition of allochthonous and autochthonous organic matter. Limnology and Oceanography, 63(4):1488–1501, July 2018. CODEN LIOCAH. ISSN 0024-3590.

Gustafsson:2016:APS

[GN16] Camilla Gustafsson and Alf Norkko. Not all plants are the same: Exploring metabolism and nitrogen fluxes in a benthic community composed of different aquatic plant species. Limnology and Oceanography, 62(3):1787–1799, September 2016. CODEN LI-OCAH. ISSN 0024-3590.

Gikuma-Njuru:2013:SVN

[GNHGM13] P. Gikuma-Njuru, R. E. Hecky, S. J. Guildford, and S. MacIntyre. Spatial variability of nutrient concentrations, fluxes, and ecosystem metabolism in Nyanza Gulf and Rusinga Channel, Lake Victoria (East Africa). Limnology and Oceanography, 58 (3):774–789, May 2013. CODEN LIOCAH. ISSN 0024-3590.

Ger:2019:ZGS

[GNWDL19] Kemal Ali Ger, Suzanne Naus-Wiezer, Luc De Meester, and Miquel Lürling. Zooplankton grazing selectivity regulates herbivory and dominance of toxic phytoplankton over multiple prey generations. Limnology and Oceanography, 64(3):1214–1227, May 2019. CODEN LIOCAH. ISSN 0024-3590.

Ger:2018:SIM

[GOD^{+}18] Kemal Ali Ger, Timothy G. Otten, Rita DuMais, Toni Ignoffo, and Wim Kimmerer. In situ ingestion of Microcystis is negatively related to copepod abundance in the upper San Francisco Estuary. Limnology and Oceanography, 63(6):2394–2410, November 2018. CODEN LIOCAH. ISSN 0024-3590.

Gera:2014:ECS

[GPA^{+}14] Alessandro Gera, Jordi F. Pagès, Rohan Arthur, Simone Farina, Guillem Roca, Javier Romero, and Teresa Alcoverro. The effect of a centenary storm on the long-lived seagrass Posidonia oceanica. Limnology and Oceanography, 59(6):1910–1918, November 2014. CODEN LIOCAH. ISSN 0024-3590.

Galvan:2016:AVE

[GPCJ16] Cristina Galván, Araceli Puente, Sonia Castanedo, and José A. Juanes. Average vs. extreme salinity conditions: Do they equally affect the distribution of macroinvertebrates in estuarine environments? Limnology and Oceanography, 61(3):984–1000, May 2016. CODEN LIOCAH. ISSN 0024-3590.

Garneau:2013:STD

[GPH^{+}13] Marie-Ève Garneau, Thomas Posch, Gregory Hitz, François Pomerleau, Cédric Pradalier, Roland Siegwart, and Jakob Pernthaler. Short-term displacement of Planktothrix rubescens (cyanobacteria) in a pre-alpine lake observed using an autonomous sampling platform. Limnology and Oceanography, 58 (5):1892–1906, September 2013. CODEN LIOCAH. ISSN 0024-3590.

Ger:2011:CAM

[GPL11] Kemal Ali Ger, Renata Panosso, and Miquel Lürling. Consequences of acclimation to Microcystis on the selective feeding behavior of the calanoid copepod Eudiaptomus gracilis. Limnology and Oceanography, 56(6):2103–2114, November 2011. CO-DEN LIOCAH. ISSN 0024-3590.

Gregg:2015:SPM

[GPS15] Tiffany Gregg, Fredrick G. Prahl, and Bernd R. T. Simoneit. Suspended particulate matter transport of polycyclic aromatic hydrocarbons in the lower Columbia River and its estuary. Limnology and Oceanography, 60(6):1935–1949, November 2015. CODEN LIOCAH. ISSN 0024-3590.

Gutierrez-Rodriguez:2014:IIP

[GRDPL14] Andrés Gutiérrez-Rodríguez, Moira Décima, Brian N. Popp, and Michael R. Landry. Isotopic invisibility of protozoan trophic steps in marine food webs. Limnology and Oceanography, 59 (5):1590–1598, September 2014. CODEN LIOCAH. ISSN 0024-3590.

Grasse:2016:SNC

[GRE^{+}16] Patricia Grasse, Evgenia Ryabenko, Claudia Ehlert, Mark A. Altabet, and Martin Frank. Silicon and nitrogen cycling in the upwelling area off Peru: a dual isotope approach. Limnology and Oceanography, 62(3):1661–1676, September 2016. CODEN LIOCAH. ISSN 0024-3590.

Gali:2013:SID

[GRGL^{+}13] Martí Gali, Clara Ruiz-Gonzàlez, Thomas Lefort, Josep M. Gasol, Clara Cardelús, Cristina Romera-Castillo, and Rafel Simó. Spectral irradiance dependence of sunlight effects on plankton dimethylsulfide production. Limnology and Oceanography, 58(2):489–504, March 2013. CODEN LIOCAH. ISSN 0024-3590.

Gutierrez-Rodriguez:2017:DSC

[GRPB^{+}17] Andres Gutierrez-Rodriguez, Loic Pillet, Tristan Biard, Ward Said-Ahmad, Alon Amrani, Rafel Simó, and Fabrice Not. Dimethylated sulfur compounds in symbiotic protists: a potentially significant source for marine DMS(P). Limnology and Oceanography, 63(3):1139–1154, May 2017. CODEN LIOCAH. ISSN 0024-3590.

Grossowicz:2017:PLS

[GRRA^{+}17] Michal Grossowicz, Dalit Roth-Rosenberg, Dikla Aharonovich, Jacob Silverman, Michael J. Follows, and Daniel Sher. Prochlorococcus in the lab and in silico: The importance of representing exudation. Limnology and Oceanography, 62(2): 818–835, March 2017. CODEN LIOCAH. ISSN 0024-3590.

Gutierrez-Rodriguez:2014:FSS

[GRSD^{+}14] Andrés Gutiérrez-Rodríguez, Gillian Slack, Emy F. Daniels, Karen E. Selph, Brian Palenik, and Michael R. Landry. Fine spatial structure of genetically distinct picocyanobacterial populations across environmental gradients in the Costa Rica Dome. Limnology and Oceanography, 59(3):705–723, May 2014. CO-DEN LIOCAH. ISSN 0024-3590.

Glud:2014:BPI

[GRT^{+}14] Ronnie N. Glud, Søren Rysgaard, Gavin Turner, Daniel F. McGinnis, and Raymond J. G. Leakey. Biological- and physicalinduced oxygen dynamics in melting sea ice of the Fram Strait. Limnology and Oceanography, 59(4):1097–1111, July 2014. CO-DEN LIOCAH. ISSN 0024-3590.

Grubisic:2017:ALN

[GSB^{+}17] Maja Grubisic, Gabriel Singer, M. Cristina Bruno, Roy H. A. van Grunsven, Alessandro Manfrin, Michael T. Monaghan, and Franz Hölker. Artificial light at night decreases biomass and alters community composition of benthic primary producers in a sub-alpine stream. Limnology and Oceanography, 62(6):2799– 2810, November 2017. CODEN LIOCAH. ISSN 0024-3590.

Griffiths:2011:ESC

[GSBR11] Jennifer R. Griffiths, Daniel E. Schindler, Laurie S. Balistrieri, and Gregory T. Ruggerone. Effects of simultaneous climate change and geomorphic evolution on thermal characteristics of a shallow Alaskan lake. Limnology and Oceanography, 56(1): 193–205, January 2011. CODEN LIOCAH. ISSN 0024-3590.

Giling:2017:DDM

[GSG^{+}17] Darren P. Giling, Peter A. Staehr, Hans Peter Grossart, Mikkel René Andersen, Bertram Boehrer, Carmelo Escot, Fatih Evrendilek, Lluís Gómez-Gener, Mark Honti, Ian D. Jones, Nusret Karakaya, Alo Laas, Enrique Moreno-Ostos, Karsten Rinke, Ulrike Scharfenberger, Silke R. Schmidt, Michael Weber, R. Iestyn Woolway, Jacob A. Zwart, and Biel Obrador. Delving deeper: Metabolic processes in the metalimnion of stratified lakes. Limnology and Oceanography, 63(3):1288–1306, May 2017. CODEN LIOCAH. ISSN 0024-3590. See corrigendum [Ano21a].

Gonzalez:2013:MNM

[GSPM13] Dana J. Gonzalez, Ashley R. Smyth, Michael F. Piehler, and Karen J. McGlathery. Mats of the nonnative macroalga, Gracilaria vermiculophylla, alter net denitrification rates and nutrient fluxes on intertidal mudflats. Limnology and Oceanography, 58(6):2101–2108, November 2013. CODEN LIOCAH. ISSN 0024-3590.

Greenwood:2013:ESG

[GSZL13] J. E. Greenwood, G. Symonds, L. Zhong, and M. Lourey. Evidence of submarine groundwater nutrient supply to an oligotrophic barrier reef. Limnology and Oceanography, 58(5):1834– 1842, September 2013. CODEN LIOCAH. ISSN 0024-3590.

G-Toth:2011:ZCR

[GTPB^{+}11] Làszló G.-Tóth, Laura Parpala, Csilla Balogh, Istvàn Tàtrai, and Eszter Baranyai. Zooplankton community response to enhanced turbulence generated by water-level decrease in Lake Balaton, the largest shallow lake in Central Europe. Limnology and Oceanography, 56(6):2211–2222, November 2011. CODEN LIOCAH. ISSN 0024-3590.

Griffiths:2013:ALU

[GTR^{+}13] Natalie A. Griffiths, Jennifer L. Tank, Todd V. Royer, Sarah S. Roley, Emma J. Rosi-Marshall, Matt R. Whiles, Jake J. Beaulieu, and Laura T. Johnson. Agricultural land use alters the seasonality and magnitude of stream metabolism. Limnology and Oceanography, 58(4):1513–1529, July 2013. CODEN LIOCAH. ISSN 0024-3590.

Grosse:2017:BVN

[GvBBB17] Julia Grosse, Peter van Breugel, Corina P. D. Brussaard, and Henricus T. S. Boschker. A biosynthesis view on nutrient stress in coastal phytoplankton. Limnology and Oceanography, 62(2): 490–506, March 2017. CODEN LIOCAH. ISSN 0024-3590.

Guilini:2010:NIB

[GVS^{+}10] Katja Guilini, Dick Van Oevelen, Karline Soetaert, Jack J. Middelburg, and Ann Vanreusel. Nutritional importance of benthic bacteria for deep-sea nematodes from the Arctic ice margin: Results of an isotope tracer experiment. Limnology and Oceanography, 55(5):1977–1989, September 2010. CODEN LIOCAH. ISSN 0024-3590.

Gradoville:2014:DTA

[GWB^{+}14] Mary R. Gradoville, Angelicque E. White, Daniela Böttjer, Matthew J. Church, and Ricardo M. Letelier. Diversity trumps acidification: Lack of evidence for carbon dioxide enhancement of Trichodesmium community nitrogen or carbon fixation at Station ALOHA. Limnology and Oceanography, 59(3):645–659, May 2014. CODEN LIOCAH. ISSN 0024-3590.

Glibert:2016:PMA

[GWD^{+}16] Patricia M. Glibert, Frances P. Wilkerson, Richard C. Dugdale, John A. Raven, Christopher L. Dupont, Peter R. Leavitt, Alexander E. Parker, JoAnn M. Burkholder, and Todd M. Kana. Pluses and minuses of ammonium and nitrate uptake and assimilation by phytoplankton and implications for productivity and community composition, with emphasis on nitrogenenriched conditions. Limnology and Oceanography, 61(1):165– 197, January 2016. CODEN LIOCAH. ISSN 0024-3590.

Ge:2012:NHL

[GWN^{+}12] Zhongfu Ge, Richard L. Whitman, Meredith B. Nevers, Mantha S. Phanikumar, and Muruleedhara N. Byappanahalli. Nearshore hydrodynamics as loading and forcing factors for Escherichia coli contamination at an embayed beach. Limnology and Oceanography, 57(1):362–381, January 2012. CODEN LI-OCAH. ISSN 0024-3590.

Glass:2010:MNC

[GWSEA10] Jennifer B. Glass, Felisa Wolfe-Simon, James J. Elser, and Ariel D. Anbar. Molybdenum–nitrogen co-limitation in freshwater and coastal heterocystous cyanobacteria. Limnology and Oceanography, 55(2):667–676, March 2010. CODEN LIOCAH. ISSN 0024-3590.

Godrijan:2018:CZI

[GYP^{+}18] Jelena Godrijan, Jeremy R. Young, Daniela Marić Pfannkuchen, Robert Precali, and Martin Pfannkuchen. Coastal zones as important habitats of coccolithophores: a study of species diversity, succession, and life-cycle phases. Limnology and Oceanography, 63(4):1692–1710, July 2018. CODEN LIOCAH. ISSN 0024-3590.

Hannides:2016:PEB

[HA16] Angelos K. Hannides and Robert C. Aller. Priming effect of benthic gastropod mucus on sedimentary organic matter remineralization. Limnology and Oceanography, 62(3):1640–1650, September 2016. CODEN LIOCAH. ISSN 0024-3590.

Hosen:2019:EPP

[HAA^{+}19] J. D. Hosen, K. S. Aho, A. P. Appling, E. C. Creech, J. H. Fair, R. O. Hall, Jr., E. D. Kyzivat, R. S. Lowenthal, S. Matt, J. Morrison, J. E. Saiers, J. B. Shanley, L. C. Weber, B. Yoon, and P. A. Raymond. Enhancement of primary production during drought in a temperate watershed is greater in larger rivers than headwater streams. Limnology and Oceanography, 64(4):1458– 1472, July 2019. CODEN LIOCAH. ISSN 0024-3590.

Hardison:2011:CND

[HAC^{+}11] A. K. Hardison, I. C. Anderson, E. A. Canuel, C. R. Tobias, and B. Veuger. Carbon and nitrogen dynamics in shallow photic systems: Interactions between macroalgae, microalgae, and bacteria. Limnology and Oceanography, 56(4):1489–1503, July 2011. CODEN LIOCAH. ISSN 0024-3590.

Hemraj:2017:CSP

[HAL17] Deevesh A. Hemraj, Laetitia Allais, and Sophie C. Leterme. A combination of salinity and pH affects the recruitment of Gladioferens pectinatus (Brady) (Copepoda; Calanoida). Limnology and Oceanography, 62(5):1799–1809, September 2017. CODEN LIOCAH. ISSN 0024-3590.

Hessen:2017:WSW

[HATF17] D. O. Hessen, T. Andersen, K. Tominaga, and A. G. Finstad. When soft waters becomes softer; drivers of critically low levels of Ca in Norwegian lakes. Limnology and Oceanography, 62(1): 289–298, January 2017. CODEN LIOCAH. ISSN 0024-3590.

Helbling:2011:IRA

[HBB^{+}11] E. Walter Helbling, Anita G. J. Buma, Peter Boelen, Han J. van der Strate, M. Valeria Fiorda Giordanino, and Virginia E. Villafañe. Increase in Rubisco activity and gene expression due to elevated temperature partially counteracts ultraviolet radiation–induced photoinhibition in the marine diatom Thalassiosira weissflogii. Limnology and Oceanography, 56(4):1330– 1342, July 2011. CODEN LIOCAH. ISSN 0024-3590.

Herstoff:2019:DPE

[HBBM19] Emily M. Herstoff, Stephen B. Baines, Maarten Boersma, and Cédric L. Meunier. Does prey elemental stoichiometry influence copepod movement over ontogeny? Limnology and Oceanography, 64(6):2467–2477, November 2019. CODEN LIOCAH. ISSN 0024-3590.

Hirst:2010:DPC

[HBCK10] A. G. Hirst, D. Bonnet, D. V. P. Conway, and T. Kiørboe. Does predation controls adult sex ratios and longevities in marine pelagic copepods? Limnology and Oceanography, 55(5):2193– 2206, September 2010. CODEN LIOCAH. ISSN 0024-3590.

Hill:2011:IRD

[HBD^{+}11] Ross Hill, Christopher M. Brown, Katrina DeZeeuw, Douglas A. Campbell, and Peter J. Ralph. Increased rate of D1 repair in coral symbionts during bleaching is insufficient to counter accelerated photo-inactivation. Limnology and Oceanography, 56(1): 139–146, January 2011. CODEN LIOCAH. ISSN 0024-3590.

Hansel:2016:DES

[HBD^{+}16] C. M. Hansel, C. Buchwald, J. M. Diaz, J. E. Ossolinski, S. T. Dyhrman, B. A. S. Van Mooy, and Despo Polyviou. Dynamics of extracellular superoxide production by Trichodesmium colonies from the Sargasso Sea. Limnology and Oceanography, 61(4): 1188–1200, July 2016. CODEN LIOCAH. ISSN 0024-3590.

Hume:2011:DOF

[HBM11] Andrew C. Hume, Peter Berg, and Karen J. McGlathery. Dissolved oxygen fluxes and ecosystem metabolism in an eelgrass (Zostera marina) meadow measured with the eddy correlation technique. Limnology and Oceanography, 56(1):86–96, January 2011. CODEN LIOCAH. ISSN 0024-3590.

Hill:2015:CMC

[HBM^{+}15] Ross Hill, Alecia Bellgrove, Peter I. Macreadie, Katherina Petrou, John Beardall, Andy Steven, and Peter J. Ralph. Can macroalgae contribute to blue carbon? An Australian perspective. Limnology and Oceanography, 60(5):1689–1706, September 2015. CODEN LIOCAH. ISSN 0024-3590.

Hoellein:2013:ROS

[HBR13] Timothy J. Hoellein, Denise A. Bruesewitz, and David C. Richardson. Revisiting Odum (1956): a synthesis of aquatic ecosystem metabolism. Limnology and Oceanography, 58(6): 2089–2100, November 2013. CODEN LIOCAH. ISSN 0024-3590.

Hanson:2014:QLA

[HBR^{+}14] Paul C. Hanson, Ishi Buffam, James A. Rusak, Emily H. Stanley, and Carl Watras. Quantifying lake allochthonous organic carbon budgets using a simple equilibrium model. Limnology and Oceanography, 59(1):167–181, January 2014. CODEN LI-OCAH. ISSN 0024-3590.

Hu:2012:SLA

[HBZ12] Xinping Hu, David J. Burdige, and Richard C. Zimmerman. δ ^{13}C is a signature of light availability and photosynthesis in seagrass. Limnology and Oceanography, 57(2):441–448, March 2012. CODEN LIOCAH. ISSN 0024-3590.

Heffernan:2010:DIC

[HC10] James B. Heffernan and Matthew J. Cohen. Direct and indirect coupling of primary production and diel nitrate dynamics in a subtropical spring-fed river. Limnology and Oceanography, 55 (2):677–688, March 2010. CODEN LIOCAH. ISSN 0024-3590.

Hensley:2012:CST

[HC12] Robert T. Hensley and Matthew J. Cohen. Controls on solute transport in large spring-fed karst rivers. Limnology and Oceanography, 57(4):912–924, July 2012. CODEN LIOCAH. ISSN 0024-3590.

Howard:2018:CRC

[HCAF18] Jason L. Howard, Joel C. Creed, Mariana V. P. Aguiar, and James W. Fourqurean. CO_{2} released by carbonate sediment production in some coastal areas may offset the benefits of seagrass “Blue carbon” storage. Limnology and Oceanography, 63 (1):160–172, January 2018. CODEN LIOCAH. ISSN 0024-3590.

Huang:2013:EWD

[HCC^{+}13] Wei-Jen Huang, Wei-Jun Cai, Renato M. Castelao, Yongchen Wang, and Steven E. Lohrenz. Effects of a wind-driven crossshelf large river plume on biological production and CO_{2} uptake on the Gulf of Mexico during spring. Limnology and Oceanography, 58(5):1727–1735, September 2013. CODEN LIOCAH. ISSN 0024-3590.

Haak:2019:WTD

[HCD19] Christopher R. Haak, Geoffrey W. Cowles, and Andy J. Danylchuk. Wave and tide-driven flow act on multiple scales to shape the distribution of a juvenile fish (Albula vulpes) in shallow nearshore habitats. Limnology and Oceanography, 64 (2):597–615, March 2019. CODEN LIOCAH. ISSN 0024-3590.

Heffernan:2010:HBI

[HCF^{+}10] James B. Heffernan, Matthew J. Cohen, Thomas K. Frazer, Ray G. Thomas, Travis J. Rayfield, Jason Gulley, Jonathan B. Martin, Joseph J. Delfino, and Wendy D. Graham. Hydrologic and biotic influences on nitrate removal in a subtropical springfed river. Limnology and Oceanography, 55(1):249–263, January 2010. CODEN LIOCAH. ISSN 0024-3590.

Huang:2019:POC

[HCH^{+}19] Yibin Huang, Bingzhang Chen, Bangqin Huang, Hui Zhou, and Yongquan Yuan. Potential overestimation of community respiration in the western Pacific boundary ocean: What causes the putative net heterotrophy in oligotrophic systems? Limnology and Oceanography, 66(4):2202–2219, September 2019. CODEN LIOCAH. ISSN 0024-3590.

Hagerthey:2010:AME

[HCK10] Scot E. Hagerthey, Jonathan J. Cole, and Deborah Kilbane. Aquatic metabolism in the Everglades: Dominance of water column heterotrophy. Limnology and Oceanography, 55(2):653–666, March 2010. CODEN LIOCAH. ISSN 0024-3590.

Hovey:2011:DMR

[HCK11] Renae K. Hovey, Marion L. Cambridge, and Gary A. Kendrick. Direct measurements of root growth and productivity in the seagrasses Posidonia australis and P. sinuosa. Limnology and Oceanography, 56(1):394–402, January 2011. CODEN LIOCAH. ISSN 0024-3590.

Hensley:2014:INR

[HCK14] Robert T. Hensley, Matthew J. Cohen, and Larry V. Korhnak. Inferring nitrogen removal in large rivers from high-resolution longitudinal profiling. Limnology and Oceanography, 59(4): 1152–1170, July 2014. CODEN LIOCAH. ISSN 0024-3590.

Huang:2018:EAC

[HCL^{+}18] Yuan Huang, Guilian Cui, Bangping Li, Xuexia Zhu, and Zhou Yang. Elevated atmospheric CO_{2} enhances grazer-induced morphological defense in the freshwater green alga Scenedesmus obliquus. Limnology and Oceanography, 63(2):1004–1014, March 2018. CODEN LIOCAH. ISSN 0024-3590.

Ho:2011:TMC

[HCLS11] Tung-Yuan Ho, Wen-Chen Chou, Hui-Ling Lin, and David D. Sheu. Trace metal cycling in the deep water of the South China Sea: The composition, sources, and fluxes of sinking particles. Limnology and Oceanography, 56(4):1225–1243, July 2011. CO-DEN LIOCAH. ISSN 0024-3590.

Huebert:2011:VMR

[HCS11] Klaus B. Huebert, Robert K. Cowen, and Su Sponaugle. Vertical migrations of reef fish larvae in the Straits of Florida and effects on larval transport. Limnology and Oceanography, 56(5):1653– 1666, September 2011. CODEN LIOCAH. ISSN 0024-3590.

Ho:2010:TMC

[HCW^{+}10] Tung-Yuan Ho, Wen-Chen Chou, Ching-Ling Wei, Fei-Jan Lin, George T. F. Wong, and Hui-Ling Line. Trace metal cycling in the surface water of the South China Sea: Vertical fluxes, composition, and sources. Limnology and Oceanography, 55(5): 1807–1820, September 2010. CODEN LIOCAH. ISSN 0024-3590.

Houghton:2019:AHB

[HD19] Isabel A. Houghton and John O. Dabiri. Alleviation of hypoxia by biologically generated mixing in a stratified water column. Limnology and Oceanography, 66(4):2161–2171, September 2019. CODEN LIOCAH. ISSN 0024-3590.

Holding:2017:AAC

[HDDH^{+}17] Johnna M. Holding, Carlos M. Duarte, Antonio Delgado-Huertas, Karline Soetaert, Jorien E. Vonk, Susana Agustí, Paul Wassmann, and Jack J. Middelburg. Autochthonous and allochthonous contributions of organic carbon to microbial food webs in Svalbard fjords. Limnology and Oceanography, 66(5): 1307–1323, May 2017. CODEN LIOCAH. ISSN 0024-3590.

Han:2012:NDB

[HDK^{+}12] Aiqin Han, Minhan Dai, Shuh-Ji Kao, Jianping Gan, Qing Li, Lifang Wang, Weidong Zhai, and Lei Wang. Nutrient dynamics and biological consumption in a large continental shelf system under the influence of both a river plume and coastal upwelling. Limnology and Oceanography, 57(3):486–502, March 2012. CO-DEN LIOCAH. ISSN 0024-3590.

Hannides:2015:MZM

[HDP15] Cecelia C. S. Hannides, Jeffrey C. Drazen, and Brian N. Popp. Mesopelagic zooplankton metabolic demand in the North Pacific Subtropical Gyre. Limnology and Oceanography, 60(2):419–428, March 2015. CODEN LIOCAH. ISSN 0024-3590.

Hedley:2010:OPC

[HE10] J. Hedley and S. Enríquez. Optical properties of canopies of the tropical seagrass Thalassia testudinum estimated by a threedimensional radiative transfer model. Limnology and Oceanography, 55(4):1537–1550, July 2010. CODEN LIOCAH. ISSN 0024-3590.

Harjung:2019:EER

[HEB^{+}19] Astrid Harjung, Elisabet Ejarque, Tom Battin, Andrea Butturini, Francesc Sabater, Masumi Stadler, and Jakob Schelker. Experimental evidence reveals impact of drought periods on dissolved organic matter quality and ecosystem metabolism in subalpine streams. Limnology and Oceanography, 64(1):46–60, January 2019. CODEN LIOCAH. ISSN 0024-3590.

Howard:2010:RNC

[HEBS10] Evan Howard, Steven Emerson, Seth Bushinsky, and Charles Stump. The role of net community production in air-sea carbon fluxes at the North Pacific subarctic–subtropical boundary region. Limnology and Oceanography, 55(6):2585–2596, November 2010. CODEN LIOCAH. ISSN 0024-3590.

Hill:2017:SSM

[HEH^{+}17] Brian H. Hill, Colleen M. Elonen, Alan T. Herlihy, Terri M. Jicha, and Richard M. Mitchell. A synoptic survey of microbial respiration, organic matter decomposition, and carbon efflux in U.S. streams and rivers. Limnology and Oceanography, 62(S1): S147–S159, November 2017. CODEN LIOCAH. ISSN 0024-

Hessen:2013:ESE

[HESU13] Dag O. Hessen, James J. Elser, Robert W. Sterner, and Jotaro Urabe. Ecological stoichiometry: an elementary approach using basic principles. Limnology and Oceanography, 58(6):2219–2236, November 2013. CODEN LIOCAH. ISSN 0024-3590.

Hofmann:2010:WIR

[HFP10] Hilmar Hofmann, Luisa Federwisch, and Frank Peeters. Waveinduced release of methane: Littoral zones as source of methane in lakes. Limnology and Oceanography, 55(5):1990– 2000, September 2010. CODEN LIOCAH. ISSN 0024-3590.

Hawes:2014:EPA

[HGD14] Ian Hawes, Hilke Giles, and Peter T. Doran. Estimating photosynthetic activity in microbial mats in an ice-covered Antarctic lake using automated oxygen microelectrode profiling and variable chlorophyll fluorescence. Limnology and Oceanography, 59 (3):674–688, May 2014. CODEN LIOCAH. ISSN 0024-3590.

Hansson:2019:EDM

[HGdG^{+}19] Truls Hveem Hansson, Hans-Peter Grossart, Paul A. del Giorgio, Nicolas F. St-Gelais, and Beatrix E. Beisner. Environmental drivers of mixotrophs in boreal lakes. Limnology and Oceanography, 64(4):1688–1705, July 2019. CODEN LIOCAH. ISSN 0024-3590.

Halsey:2017:BCV

[HGG^{+}17] Kimberly H. Halsey, Stephen J. Giovannoni, Martin Graus, Yanlin Zhao, Zachary Landry, J. Cameron Thrash, Kevin L. Vergin, and Joost de Gouw. Biological cycling of volatile organic carbon by phytoplankton and bacterioplankton. Limnology and Oceanography, 62(6):2650–2661, November 2017. CODEN LI-OCAH. ISSN 0024-3590.

Hollweg:2010:MMC

[HGM10] T. A. Hollweg, C. C. Gilmour, and R. P. Mason. Mercury and methylmercury cycling in sediments of the mid-Atlantic continental shelf and slope. Limnology and Oceanography, 55(6): 2703–2722, November 2010. CODEN LIOCAH. ISSN 0024-

Hoer:2018:CDO

[HGT^{+}18] Daniel R. Hoer, Patrick J. Gibson, Jake P. Tommerdahl, Niels L. Lindquist, and Christopher S. Martens. Consumption of dissolved organic carbon by Caribbean reef sponges. Limnology and Oceanography, 63(4):337–351, January 2018. CODEN LI-OCAH. ISSN 0024-3590.

Hsieh:2013:URQ

[HGvB^{+}13] Yu-Te Hsieh, Walter Geibert, Pieter van Beek, Henrik Stahl, Dmitry Aleynik, and Gideon M. Henderson. Using the radium quartet (^{228}Ra, ^{226}Ra, ^{224}Ra, and^{223}Ra) to estimate water mixing and radium inputs in Loch Etive, Scotland. Limnology and Oceanography, 58(3):1089–1102, May 2013. CODEN LIOCAH. ISSN 0024-3590.

Hotchkiss:2014:HRD

[HH14] Erin R. Hotchkiss and Robert O. Hall, Jr. High rates of daytime respiration in three streams: Use of δ ^{18}O_{O2} and O_{2} to model diel ecosystem metabolism. Limnology and Oceanography, 59 (3):798–810, May 2014. CODEN LIOCAH. ISSN 0024-3590.

Hentschel:2018:HMK

[HHA18] Brian T. Hentschel, Nicholas T. Hayman, and Todd W. Anderson. Hydrodynamic mediation of killifish predation on infaunal polychaetes. Limnology and Oceanography, 63(S1):S19–S29, March 2018. CODEN LIOCAH. ISSN 0024-3590.

Hall:2019:MCS

[HHE^{+}19] Britt D. Hall, Raymond H. Hesslein, Craig A. Emmerton, Scott N. Higgins, Patricia Ramlal, and Michael J. Paterson. Multidecadal carbon sequestration in a headwater boreal lake. Limnology and Oceanography, 64(S1):S150–S165, January 2019. CODEN LIOCAH. ISSN 0024-3590.

Hayduk:2019:ERS

[HHHT19] Jennifer L. Hayduk, Sally D. Hacker, Jeremy S. Henderson, and Fiona Tomas. Evidence for regional-scale controls on eelgrass (Zostera marina) and mesograzer community structure in upwelling-influenced estuaries. Limnology and Oceanography, 64 (3):1120–1134, May 2019. CODEN LIOCAH. ISSN 0024-3590.

Havig:2018:WCS

[HHM^{+}18] Jeff R. Havig, Trinity L. Hamilton, Michael McCormick, Brianna McClure, Todd Sowers, Bruce Wegter, and Lee R. Kump. Water column and sediment stable carbon isotope biogeochemistry of permanently redox-stratified Fayetteville Green Lake, New York, U.S.A. Limnology and Oceanography, 63(2):570–587, March 2018. CODEN LIOCAH. ISSN 0024-3590.

Henson:2018:NDS

[HHS^{+}18] Michael W. Henson, Jordan Hanssen, Greg Spooner, Patrick Fleming, Markus Pukonen, Frederick Stahr, and J. Cameron Thrash. Nutrient dynamics and stream order influence microbial community patterns along a 2914 kilometer transect of the Mississippi River. Limnology and Oceanography, 63(5):1837–1855, September 2018. CODEN LIOCAH. ISSN 0024-3590.

Harethardottir:2019:TIT

[HHW^{+}19] Sara Hardardóttir, Ditte Marie Hjort, Sylke Wohlrab, Bernd Krock, Uwe John, Torkel Gissel Nielsen, and Nina Lundholm. Trophic interactions, toxicokinetics, and detoxification processes in a domoic acid-producing diatom and two copepod species. Limnology and Oceanography, 64(3):833–848, May 2019. CO-DEN LIOCAH. ISSN 0024-3590.

Hirst:2012:ISM

[Hir12] Andrew G. Hirst. Intraspecific scaling of mass to length in pelagic animals: Ontogenetic shape change and its implications. Limnology and Oceanography, 58(1):1579–1590, September 2012. CODEN LIOCAH. ISSN 0024-3590.

Hietanen:2012:HNP

[HJB^{+}12] Susanna Hietanen, Helena Jäntti, Christo Buizert, Klaus Jürgens, Matthias Labrenz, Maren Voss, and Jorma Kuparinen. Hypoxia and nitrogen processing in the Baltic Sea water column. Limnology and Oceanography, 57(1):325–337, January 2012. CODEN LIOCAH. ISSN 0024-3590.

Harvey:2013:AAC

[HJMD13] Elizabeth L. Harvey, Hae Jin Jeong, and Susanne Menden-Deuer. Avoidance and attraction: Chemical cues influence predator-prey interactions of planktonic protists. Limnology and Oceanography, 58(4):1176–1184, July 2013. CODEN LIOCAH. ISSN 0024-3590.

Haraldsson:2013:ECI

[HJT^{+}13a] Matilda Haraldsson, Cornelia Jaspers, Peter Tiselius, Dag L. Aksnes, Tom Andersen, and Josefin Titelman. Environmental constraints of the invasive Mnemiopsis leidyi in Scandinavian waters. Limnology and Oceanography, 58(1):37–48, January 2013. CODEN LIOCAH. ISSN 0024-3590. See erratum [HJT^{+}13b].

Haraldsson:2013:EEC

[HJT^{+}13b] Matilda Haraldsson, Cornelia Jaspers, Peter Tiselius, Dag L. Aksnes, Tom Andersen, and Josefin Titelman. Erratum: Environmental constraints of the invasive Mnemiopsis leidyi in Scandinavian waters. Limnology and Oceanography, 58(2): 763, March 2013. CODEN LIOCAH. ISSN 0024-3590. See [HJT^{+}13a].

Hansen:2016:OPD

[HKP^{+}16] Angela M. Hansen, Tamara E. C. Kraus, Brian A. Pellerin, Jacob A. Fleck, Bryan D. Downing, and Brian A. Bergamaschi. Optical properties of dissolved organic matter (DOM): Effects of biological and photolytic degradation. Limnology and Oceanography, 61(3):1015–1032, May 2016. CODEN LIOCAH. ISSN 0024-3590.

Hylander:2015:CSA

[HKS^{+}15] Samuel Hylander, Thomas Kiørboe, Pauline Snoeijs, Ruben Sommaruga, and Torkel Gissel Nielsen. Concentrations of sunscreens and antioxidant pigments in Arctic Calanus spp. in relation to ice cover, ultraviolet radiation, and the phytoplankton spring bloom. Limnology and Oceanography, 60(6):2197–2206, November 2015. CODEN LIOCAH. ISSN 0024-3590.

Hughes:2010:SVC

[HKU^{+}10] Claire Hughes, Anthony J. Kettle, Godwin A. Unazi, Keith Weston, Matthew R. Jones, and Martin T. Johnson. Seasonal variations in the concentrations of methyl and ethyl nitrate in a shallow freshwater lake. Limnology and Oceanography, 55(1): 305–314, January 2010. CODEN LIOCAH. ISSN 0024-3590.

Haney:2013:SAM

[HL13] James F. Haney and Winfried Lampert. Spatial avoidance of Microcystis aeruginosa by Daphnia: Fitness consequences and evolutionary implications. Limnology and Oceanography, 58(6): 2122–2132, November 2013. CODEN LIOCAH. ISSN 0024-3590.

Hernandez-Leon:2010:CSZ

[HLFM^{+}10] Santiago Hernández-León, Gara Franchy, Marta Moyano, Inmaculada Menéndez, Claire Schmoker, and Sébastien Putzeys. Carbon sequestration and zooplankton lunar cycles: Could we be missing a major component of the biological pump? Limnology and Oceanography, 55(6):2503–2512, November 2010. CO-DEN LIOCAH. ISSN 0024-3590.

Hattenrath-Lehmann:2015:CIO

[HLG15] Theresa Hattenrath-Lehmann and Christopher J. Gobler. The contribution of inorganic and organic nutrients to the growth of a North American isolate of the mixotrophic dinoflagellate, Dinophysis acuminata. Limnology and Oceanography, 60(5): 1588–1603, September 2015. CODEN LIOCAH. ISSN 0024-3590.

Hirst:2017:OBM

[HLGA17] A. G. Hirst, M. K. S. Lilley, D. S. Glazier, and D. Atkinson. Ontogenetic body-mass scaling of nitrogen excretion relates to body surface area in diverse pelagic invertebrates. Limnology and Oceanography, 62(1):311–319, January 2017. CODEN LI-OCAH. ISSN 0024-3590.

Hessing-Lewis:2013:UIM

[HLH13] Margot L. Hessing-Lewis and Sally D. Hacker. Upwellinginfluence, macroalgal blooms, and seagrass production; temporal trends from latitudinal and local scales in northeast Pacific estuaries. Limnology and Oceanography, 58(3):1103–1112, May 2013. CODEN LIOCAH. ISSN 0024-3590.

Henson:2012:VPC

[HLJ12] Stephanie Henson, Richard Lampitt, and David Johns. Variability in phytoplankton community structure in response to the North Atlantic Oscillation and implications for organic carbon flux. Limnology and Oceanography, 57(6):1591–1601, November 2012. CODEN LIOCAH. ISSN 0024-3590.

Hattenrath-Lehmann:2015:EEC

[HLSW^{+}15] Theresa K. Hattenrath-Lehmann, Juliette L. Smith, Ryan B. Wallace, Lucas R. Merlo, Florian Koch, Heidi Mittelsdorf, Jennifer A. Goleski, Donald M. Anderson, and Christopher J. Gobler. The effects of elevated CO_{2} on the growth and toxicity of field populations and cultures of the saxitoxin-producing dinoflagellate, Alexandrium fundyense. Limnology and Oceanography, 60(1):198–214, January 2015. CODEN LIOCAH. ISSN 0024-3590.

Harvey:2011:AMM

[HMD11] Elizabeth L. Harvey and Susanne Menden-Deuer. Avoidance, movement, and mortality: The interactions between a protistan grazer and Heterosigma akashiwo, a harmful algal bloom species. Limnology and Oceanography, 56(1):371–378, January 2011. CODEN LIOCAH. ISSN 0024-3590.

Harrison:2016:ISP

[HMFB16] B. K. Harrison, A. Myrbo, B. E. Flood, and J. V. Bailey. Identification of subannual patterns in microbial community signatures from individual sedimentary laminae using a freeze-coring approach. Limnology and Oceanography, 61(2):735–747, March 2016. CODEN LIOCAH. ISSN 0024-3590.

Hunsinger:2010:WDS

[HMFF10] Glendon B. Hunsinger, Siddhartha Mitra, Stuart E. G. Findlay, and David T. Fischer. Wetland-driven shifts in suspended particulate organic matter composition of the Hudson River estuary, New York. Limnology and Oceanography, 55(4):1653–1667, July 2010. CODEN LIOCAH. ISSN 0024-3590.

Hunsinger:2012:LZI

[HMFF12] Glendon B. Hunsinger, Siddhartha Mitra, Stuart E. G. Findlay, and David T. Fischer. Littoral-zone influences on particulate organic matter composition along the freshwater-tidal Hudson River, New York. Limnology and Oceanography, 57(5):1303– 1316, September 2012. CODEN LIOCAH. ISSN 0024-3590.

Hitchcock:2016:TDO

[HMH^{+}16] James N. Hitchcock, Simon M. Mitrovic, Wade L. Hadwen, Daniel L. Roelke, Ivor O. Growns, and Ann-Marie Rohlfs. Terrestrial dissolved organic carbon subsidizes estuarine zooplankton: an in situ mesocosm study. Limnology and Oceanography, 61(1):254–267, January 2016. CODEN LIOCAH. ISSN 0024-

Hogg:2013:MIA

[HMHI13] Charlie A. R. Hogg, Clelia L. Marti, Herbert E. Huppert, and Jörg Imberger. Mixing of an interflow into the ambient water of Lake Iseo. Limnology and Oceanography, 58(2):579–592, March 2013. CODEN LIOCAH. ISSN 0024-3590.

Hundey:2014:RCP

[HML^{+}14] E. J. Hundey, K. A. Moser, F. J. Longstaffe, N. Michelutti, and R. Hladyniuk. Recent changes in production in oligotrophic Uinta Mountain lakes, Utah, identified using paleolimnology. Limnology and Oceanography, 59(6):1987–2001, November 2014. CODEN LIOCAH. ISSN 0024-3590.

Hampton:2018:REC

[HMO^{+}18] Stephanie E. Hampton, Suzanne McGowan, Ted Ozersky, Salvatore G. P. Virdis, Tuong Thuy Vu, Trisha L. Spanbauer, Benjamin M. Kraemer, George Swann, Anson W. Mackay, Stephen M. Powers, Michael F. Meyer, Stephanie G. Labou, Catherine M. O’Reilly, Morgan DiCarlo, Aaron W. E. Galloway, and Sherilyn C. Fritz. Recent ecological change in ancient lakes. Limnology and Oceanography, 64(4):2277–2304, September 2018. CODEN LIOCAH. ISSN 0024-3590.

Heintz:2012:PCM

[HMV12] M. B. Heintz, S. Mau, and D. L. Valentine. Physical control on methanotrophic potential in waters of the Santa Monica Basin, Southern California. Limnology and Oceanography, 57(2):420– 432, March 2012. CODEN LIOCAH. ISSN 0024-3590.

Haileselasie:2018:FED

[HMV^{+}18] Tsegazeabe H. Haileselasie, Joachim Mergeay, Joost Vanoverbeke, Luisa Orsini, and Luc De Meester. Founder effects determine the genetic structure of the water flea Daphnia in Ethiopian reservoirs. Limnology and Oceanography, 63(2):915– 926, March 2018. CODEN LIOCAH. ISSN 0024-3590.

Hershey:2015:MEO

[HNHS^{+}15] Anne E. Hershey, Robert M. Northington, John Hart-Smith, Matthew Bostick, and Stephen C. Whalen. Methane efflux and oxidation, and use of methane-derived carbon by larval chironomini, in Arctic lake sediments. Limnology and Oceanography, 60 (1):276–285, January 2015. CODEN LIOCAH. ISSN 0024-3590.

Hewson:2013:MIS

[HNL^{+}13] Ian Hewson, Gabriel Ng, WenFang Li, Brenna A. LaBarre, Isabel Aguirre, Jorge G. Barbosa, Mya Breitbart, Anthony W. Greco, Colleen M. Kearns, Alexander Looi, Lindsay R. Schaffner, Philip D. Thompson, and Nelson G. Hairston, Jr. Metagenomic identification, seasonal dynamics, and potential transmission mechanisms of a Daphnia-associated singlestranded DNA virus in two temperate lakes. Limnology and Oceanography, 58(5):1605–1620, September 2013. CODEN LI-OCAH. ISSN 0024-3590.

Hansen:2012:ORS

[HNSM12] Marc Overgaard Hansen, Torkel Gissel Nielsen, Colin A. Stedmon, and Peter Munk. Oceanographic regime shift during 1997 in Disko Bay, Western Greenland. Limnology and Oceanography, 57(3):634–644, March 2012. CODEN LIOCAH. ISSN 0024-3590.

Huang:2016:WCI

[HNZ^{+}16] Yuan Huang, Haihong Nan, Xuexia Zhu, Bangping Li, Zeng Zhang, and Zhou Yang. Waterborne copper impairs grazerinduced colony formation and photosynthetic efficiency in Scenedesmus obliquus. Limnology and Oceanography, 61(2):625– 634, March 2016. CODEN LIOCAH. ISSN 0024-3590.

Ho:2013:NLN

[Ho13] Tung-Yuan Ho. Nickel limitation of nitrogen fixation in Trichodesmium. Limnology and Oceanography, 58(2):112–120, January 2013. CODEN LIOCAH. ISSN 0024-3590.

Hetherington:2017:SFW

[HOD^{+}17] Elizabeth D. Hetherington, Robert J. Olson, Jeffrey C. Drazen, Cleridy E. Lennert-Cody, Lisa T. Ballance, Ronald S. Kaufmann, and Brian N. Popp. Spatial food-web structure in the eastern tropical Pacific Ocean based on compound-specific nitrogen isotope analysis of amino acids. Limnology and Oceanography, 62(2):541–560, March 2017. CODEN LIOCAH. ISSN 0024-3590.

Hull:2011:SDD

[HONR11] Pincelli M. Hull, Karen J. Osborn, Richard D. Norris, and Bruce H. Robison. Seasonality and depth distribution of a mesopelagic foraminifer, Hastigerinella digitata, in Monterey Bay, California. Limnology and Oceanography, 56(2):562–576, March 2011. CODEN LIOCAH. ISSN 0024-3590.

Howarth:2015:E

[How15a] Robert Howarth. Editorial. Limnology and Oceanography, 60 (1):1–2, January 2015. CODEN LIOCAH. ISSN 0024-3590.

Howarth:2015:EMS

[How15b] Robert W. Howarth. Editorial: Misconduct in scientific publications. Limnology and Oceanography, 60(4):1103–1104, July 2015. CODEN LIOCAH. ISSN 0024-3590.

Howarth:2019:E

[How19] Robert W. Howarth. Editorial. Limnology and Oceanography, 64(1):1–2, January 2019. CODEN LIOCAH. ISSN 0024-3590.

Henschke:2019:LVS

[HP19] Natasha Henschke and Evgeny A. Pakhomov. Latitudinal variations in Salpa thompsoni reproductive fitness. Limnology and Oceanography, 64(2):575–584, March 2019. CODEN LIOCAH. ISSN 0024-3590.

Hannides:2013:MZS

[HPCD13] Cecelia C. S. Hannides, Brian N. Popp, C. Anela Choy, and Jeffrey C. Drazen. Midwater zooplankton and suspended particle dynamics in the North Pacific Subtropical Gyre: a stable isotope perspective. Limnology and Oceanography, 58(6):1931– 1946, November 2013. CODEN LIOCAH. ISSN 0024-3590.

Huber:2011:APV

[HPL11] Andrea M. R. Huber, Frank Peeters, and Andreas Lorke. Active and passive vertical motion of zooplankton in a lake. Limnology and Oceanography, 56(2):695–706, March 2011. CODEN LIOCAH. ISSN 0024-3590.

Helbling:2010:PDP

[HPM^{+}10] E. Walter Helbling, Daniel E. Pérez, César D. Medina, Marcos G. Lagunas, and Virginia E. Villafañe. Phytoplankton distribution and photosynthesis dynamics in the Chubut River estuary (Patagonia, Argentina) throughout tidal cycles. Limnology and Oceanography, 55(1):55–65, January 2010. CODEN LI-OCAH. ISSN 0024-3590.

Hinrichsen:2010:SPL

[HPS^{+}10a] Hans-Harald Hinrichsen, Myron A. Peck, Jörn Schmidt, Bastian Huwer, and Rudi Voss. Survival probability of larval sprat in response to decadal changes in diel vertical migration behavior and prey abundance in the Baltic Sea. Limnology and Oceanography, 55(4):1485–1498, July 2010. CODEN LIOCAH. ISSN 0024-3590.

Hortnagl:2010:LBC

[HPS10b] Paul Hörtnagl, Mariá Teresa Pérez, and Ruben Sommaruga. Living at the border: a community and single-cell assessment of lake bacterioneuston activity. Limnology and Oceanography, 55 (3):1134–1144, May 2010. CODEN LIOCAH. ISSN 0024-3590.

Horak:2018:RIL

[HQB^{+}18] Rachel E. A. Horak, Wei Qin, Anthony D. Bertagnolli, Alexa Nelson, Katherine R. Heal, Hantten Han, Maija Heller, Andrew J. Schauer, Wade H. Jeffrey, E. Virginia Armbrust, James W. Moffett, Anitra E. Ingalls, David A. Stahl, and Allan H. Devol. Relative impacts of light, temperature, and reactive oxygen on thaumarchaeal ammonia oxidation in the North Pacific Ocean. Limnology and Oceanography, 63(2):741–757, March 2018. CODEN LIOCAH. ISSN 0024-3590.

Houlbreque:2015:OAR

[HRG^{+}15] Fanny Houlbrèque, Stéphanie Reynaud, Claire Godinot, François Oberhänsli, Riccardo Rodolfo-Metalpa, and Christine Ferrier-Pagès. Ocean acidification reduces feeding rates in the scleractinian coral Stylophora pistillata. Limnology and Oceanography, 60(1):89–99, January 2015. CODEN LIOCAH. ISSN 0024-3590.

Haberlin:2019:DGZ

[HRMD19] Damien Haberlin, Robin Raine, Rob McAllen, and Thomas K. Doyle. Distinct gelatinous zooplankton communities across a dynamic shelf sea. Limnology and Oceanography, 64(4):1802– 1818, July 2019. CODEN LIOCAH. ISSN 0024-3590.

Huang:2011:MKB

[HRN11] Ivy Huang, Jeffrey Rominger, and Heidi Nepf. The motion of kelp blades and the surface renewal model. Limnology and Oceanography, 56(4):1453–1462, July 2011. CODEN LIOCAH. ISSN 0024-3590.

Hoadley:2015:DCU

[HRPW15] Kenneth D. Hoadley, Dana Rollison, D. Tye Pettay, and Mark E. Warner. Differential carbon utilization and asexual reproduction under elevated pCO_{2} conditions in the model anemone, Exaiptasia pallida, hosting different symbionts. Limnology and Oceanography, 60(6):2108–2120, November 2015. CODEN LI-OCAH. ISSN 0024-3590.

Heino:2010:CSS

[HS10] Jani Heino and Janne Soininen. Are common species sufficient in describing turnover in aquatic metacommunities along environmental and spatial gradients? Limnology and Oceanography, 55 (6):2397–2402, November 2010. CODEN LIOCAH. ISSN 0024-3590.

Harrison:2011:SSP

[HS11] Joel W. Harrison and Ralph E. H. Smith. The spectral sensitivity of phytoplankton communities to ultraviolet radiationinduced photoinhibition differs among clear and humic temperate lakes. Limnology and Oceanography, 56(6):2115–2126, November 2011. CODEN LIOCAH. ISSN 0024-3590.

Hawco:2018:CIC

[HS18] Nicholas J. Hawco and Mak A. Saito. Competitive inhibition of cobalt uptake by zinc and manganese in a Pacific Prochlorococcus strain: Insights into metal homeostasis in a streamlined oligotrophic cyanobacterium. Limnology and Oceanography, 64 (4):2229–2249, September 2018. CODEN LIOCAH. ISSN 0024-3590.

Hillebrand:2013:GRF

[HSB^{+}13] Helmut Hillebrand, Georg Steinert, Maarten Boersma, Arne Malzahn, Cédric Léo Meunier, Christoph Plum, and Robert Ptacnik. Goldman revisited: Faster-growing phytoplankton has lower N : P and lower stoichiometric flexibility. Limnology and Oceanography, 58(6):2076–2088, November 2013. CODEN LI-OCAH. ISSN 0024-3590.

Holtgrieve:2010:SQA

[HSBA10] Gordon W. Holtgrieve, Daniel E. Schindler, Trevor A. Branch, and Z. Teresa A’mar. Simultaneous quantification of aquatic ecosystem metabolism and reaeration using a Bayesian statistical model of oxygen dynamics. Limnology and Oceanography, 55 (3):1047–1063, May 2010. CODEN LIOCAH. ISSN 0024-3590.

Hughes:2011:ESB

[HSC^{+}11] Harold J. Hughes, Francis Sondag, Christine Cocquyt, Alain Laraque, Albert Pandi, Luc André, and Damien Cardinal. Effect of seasonal biogenic silica variations on dissolved silicon fluxes and isotopic signatures in the Congo River. Limnology and Oceanography, 56(2):551–561, March 2011. CODEN LI-OCAH. ISSN 0024-3590.

Howard:2014:ANS

[HSC^{+}14] Meredith D. A. Howard, Martha Sutula, David A. Caron, Yi Chao, John D. Farrara, Hartmut Frenzel, Burton Jones, George Robertson, Karen McLaughlin, and Ashmita Sengupta. Anthropogenic nutrient sources rival natural sources on small scales in the coastal waters of the Southern California Bight. Limnology and Oceanography, 59(1):285–297, January 2014. CODEN LIOCAH. ISSN 0024-3590.

Hampton:2019:LTP

[HSCM19] Stephanie E. Hampton, Mark D. Scheuerell, Matthew J. Church, and John M. Melack. Long-term perspectives in aquatic research. Limnology and Oceanography, 64(S1):S2–S10, January 2019. CODEN LIOCAH. ISSN 0024-3590.

Hancke:2014:ETI

[HSLH^{+}14] Kasper Hancke, Brian K. Sorell, Lars Chresten Lund-Hansen, Morten Larsen, Torunn Hancke, and Ronnie N. Glud. Effects of temperature and irradiance on a benthic microalgal community: a combined two-dimensional oxygen and fluorescence imaging approach. Limnology and Oceanography, 59(5):1599– 1611, September 2014. CODEN LIOCAH. ISSN 0024-3590.

Hong:2016:RMT

[HSP^{+}16] Wei-Li Hong, Simone Sauer, Giuliana Panieri, William G. Ambrose, Jr., Rachael H. James, Andreia Plaza-Faverola, and Andrea Schneider. Removal of methane through hydrological, microbial, and geochemical processes in the shallow sediments of pockmarks along eastern Vestnesa Ridge (Svalbard). Limnology and Oceanography, 61(S1):S324–S343, November 2016. CODEN LIOCAH. ISSN 0024-3590.

Haupt:2010:UPT

[HSR^{+}10] Florian Haupt, Maria Stockenreiter, Elke S. Reichwaldt, Michaela Baumgartner, Winfried Lampert, Maarten Boersma, and Herwig Stibor. Upward phosphorus transport by Daphnia diel vertical migration. Limnology and Oceanography, 55 (2):529–534, March 2010. CODEN LIOCAH. ISSN 0024-3590.

Hoyer:2015:LDN

[HSR15] Andrea B. Hoyer, S. Geoffrey Schladow, and Francisco J. Rueda. Local dispersion of nonmotile invasive bivalve species by winddriven lake currents. Limnology and Oceanography, 60(2):446– 462, March 2015. CODEN LIOCAH. ISSN 0024-3590.

Hardison:2014:ICB

[HST^{+}14] D. Ransom Hardison, William G. Sunda, Patricia A. Tester, Damian Shea, and Wayne R. Litaker. Increased cellular brevetoxins in the red tide dinoflagellate Karenia brevis under CO_{2} limitation of growth rate: Evolutionary implications and potential effects on bloom toxicity. Limnology and Oceanography, 59 (6):560–577, March 2014. CODEN LIOCAH. ISSN 0024-3590.

Hiltunen:2015:TIP

[HSTK15] Minna Hiltunen, Ursula Strandberg, Sami J. Taipale, and Paula Kankaala. Taxonomic identity and phytoplankton diet affect fatty acid composition of zooplankton in large lakes with differing dissolved organic carbon concentration. Limnology and Oceanography, 60(1):303–317, January 2015. CODEN LIOCAH. ISSN 0024-3590.

Hebert:2017:PPB

[HT17a] Mathieu Hébert and Luc Tremblay. Production and persistence of bacterial and labile organic matter at the hypoxic water– sediment interface of the St. Lawrence Estuary. Limnology and Oceanography, 65(9):2154–2167, September 2017. CODEN LI-OCAH. ISSN 0024-3590.

Heino:2017:EDM

[HT17b] Jani Heino and Kimmo T. Tolonen. Ecological drivers of multiple facets of beta diversity in a lentic macroinvertebrate metacommunity. Limnology and Oceanography, 62(6):2431–2444, November 2017. CODEN LIOCAH. ISSN 0024-3590.

Hafker:2018:CFS

[HTL^{+}18] N. Sören Häfker, M. Teschke, K. S. Last, D. W. Pond, L. Hüppe, and B. Meyer. Calanus finmarchicus seasonal cycle and diapause in relation to gene expression, physiology, and endogenous clocks. Limnology and Oceanography, 63(6):2815–2838, November 2018. CODEN LIOCAH. ISSN 0024-3590.

Hoer:2018:DIN

[HTLM18] Daniel R. Hoer, Jake P. Tommerdahl, Niels L. Lindquist, and Christopher S. Martens. Dissolved inorganic nitrogen fluxes from common Florida Bay (U.S.A.) sponges. Limnology and Oceanography, 63(6):2563–2578, November 2018. CODEN LI-OCAH. ISSN 0024-3590.

Hansen:2016:CEV

[HV16] Agnethe N. Hansen and André W. Visser. Carbon export by vertically migrating zooplankton: an optimal behavior model. Limnology and Oceanography, 61(2):701–710, March 2016. CO-DEN LIOCAH. ISSN 0024-3590.

Hansen:2019:SSO

[HV19] Agnethe Nøhr Hansen and André W. Visser. The seasonal succession of optimal diatom traits. Limnology and Oceanography, 64(4):1442–1457, July 2019. CODEN LIOCAH. ISSN 0024-3590.

Hughes:2018:INA

[HVD^{+}18] David J. Hughes, Deepa Varkey, Martina A. Doblin, Tim Ingleton, Allison Mcinnes, Peter J. Ralph, Virginie van Dongen-Vogels, and David J. Suggett. Impact of nitrogen availability upon the electron requirement for carbon fixation in Australian coastal phytoplankton communities. Limnology and Oceanography, 63(5):1891–1910, September 2018. CODEN LIOCAH. ISSN 0024-3590.

Heiden:2019:IOA

[HVJ^{+}19] Jasmin P. Heiden, Christian Völkner, Elizabeth M. Jones, Willem H. van de Poll, Anita G. J. Buma, Michael P. Meredith, Hein J. W. de Baar, Kai Bischof, Dieter Wolf-Gladrow, and Scarlett Trimborn. Impact of ocean acidification and high solar radiation on productivity and species composition of a late summer phytoplankton community of the coastal Western Antarctic Peninsula. Limnology and Oceanography, 64(4):1716–1736, July 2019. CODEN LIOCAH. ISSN 0024-3590.

Hamdan:2016:MEO

[HW16] Leila J. Hamdan and Kimberly P. Wickland. Methane emissions from oceans, coasts, and freshwater habitats: New perspectives and feedbacks on climate. Limnology and Oceanography, 61(S1): S3–S12, November 2016. CODEN LIOCAH. ISSN 0024-3590.

Hill:2013:LMR

[HWZ13] Polly G. Hill, Phillip E. Warwick, and Mikhail V. Zubkov. Low microbial respiration of leucine at ambient oceanic concentration in the mixed layer of the central Atlantic Ocean. Limnology and Oceanography, 58(5):1597–1604, September 2013. CODEN LIOCAH. ISSN 0024-3590.

Hopkinson:2010:ECP

[HXS^{+}10] Brian M. Hopkinson, Yan Xu, Dalin Shi, Patrick J. McGinn, and François M. M. Morel. The effect of CO_{2} on the photosynthetic physiology of phytoplankton in the Gulf of Alaska. Limnology and Oceanography, 55(5):2011–2024, September 2010. CODEN LIOCAH. ISSN 0024-3590.

Hall:2015:TLT

[HYK^{+}15] Robert O. Hall, Jr., Charles B. Yackulic, Theodore A. Kennedy, Michael D. Yard, Emma J. Rosi-Marshall, Nicholas Voichick, and Kathrine E. Behn. Turbidity, light, temperature, and hydropeaking control primary productivity in the Colorado River, grand canyon. Limnology and Oceanography, 60(2):512–526, March 2015. CODEN LIOCAH. ISSN 0024-3590.

Huo:2013:GAB

[HZC^{+}13] Yuanzi Huo, Jianheng Zhang, Liping Chen, Ming Hu, Kefeng Yu, Qunfang Chen, Qing He, and Peimin He. Green algae blooms caused by Ulva prolifera in the southern Yellow Sea: Identification of the original bloom location and evaluation of biological processes occurring during the early northward floating period. Limnology and Oceanography, 58(6):2206–2218, November 2013. CODEN LIOCAH. ISSN 0024-3590.

Iacchei:2017:ATI

[IBPG17] Matthew Iacchei, Elizabeth Butcher, Elan Portner, and Erica Goetze. It’s about time: Insights into temporal genetic patterns in oceanic zooplankton from biodiversity indices. Limnology and Oceanography, 62(5):1836–1852, September 2017. CODEN LI-OCAH. ISSN 0024-3590.

Ivanccic:2012:SMP

[IGP^{+}12] Ingrid Ivancčić, Jelena Godrijan, Martin Pfannkuchen, Daniela Marić, Blazčenka Gasčparović, Tamara Djakovac, and Mirjana Najdek. Survival mechanisms of phytoplankton in conditions of stratification-induced deprivation of orthophosphate: Northern Adriatic case study. Limnology and Oceanography, 57(6):1721– 1731, November 2012. CODEN LIOCAH. ISSN 0024-3590.

Istvanovic:2011:PGT

[IH11] Vera Istvànovic and Màrk Honti. Phytoplankton growth in three rivers: The role of meroplankton and the benthic retention hypothesis. Limnology and Oceanography, 56(4):1439–1452, July 2011. CODEN LIOCAH. ISSN 0024-3590.

Istvanovics:2018:CSH

[IH18] Vera Istvánovics and Márk Honti. Coupled simulation of highfrequency dynamics of dissolved oxygen and chlorophyll widens the scope of lake metabolism studies. Limnology and Oceanography, 63(1):72–90, January 2018. CODEN LIOCAH. ISSN 0024-3590.

Isada:2019:RPA

[IHSS^{+}19] T. Isada, A. Hattori-Saito, H. Saito, Y. Kondo, J. Nishioka, K. Kuma, H. Hattori, R. M. L. McKay, and K. Suzuki. Responses of phytoplankton assemblages to iron availability and mixing water masses during the spring bloom in the Oyashio region, NW Pacific. Limnology and Oceanography, 64(1):197–216, January 2019. CODEN LIOCAH. ISSN 0024-3590.

Imam:2013:SDB

[ILPL13] Yehya E. Imam, Bernard Laval, Roger Pieters, and Gregory Lawrence. The strongly damped baroclinic response to wind in a multibasin reservoir. Limnology and Oceanography, 58(4): 1243–1258, July 2013. CODEN LIOCAH. ISSN 0024-3590.

Irwin:2012:PNE

[INF12] Andrew J. Irwin, Andrew M. Nelles, and Zoe V. Finkel. Phytoplankton niches estimated from field data. Limnology and Oceanography, 57(3):787–797, May 2012. CODEN LIOCAH. ISSN 0024-3590.

Infantes:2011:POC

[IOB^{+}11] Eduardo Infantes, Alejandro Orfila, Tjeerd J. Bouma, Gonzalo Simarro, and Jorge Terrados. Posidonia oceanica and Cymodocea nodosa seedling tolerance to wave exposure. Limnology and Oceanography, 56(6):2223–2232, November 2011. CODEN LI-OCAH. ISSN 0024-3590.

Irisson:2010:VDO

[IPGP10] Jean-Olivier Irisson, Claire B. Paris, Cédric Guigand, and Serge Planes. Vertical distribution and ontogenetic “migration” in coral reef fish larvae. Limnology and Oceanography, 55(4):909– 919, March 2010. CODEN LIOCAH. ISSN 0024-3590.

Ibanhez:2016:OTR

[IR16] J. Severino P. Ibánhez and Carlos Rocha. Oxygen transport and reactivity within a sandy seepage face in a mesotidal lagoon (Ria Formosa, Southwestern Iberia). Limnology and Oceanography, 61(1):61–77, January 2016. CODEN LIOCAH. ISSN 0024-3590.

Iwabuchi:2017:PKF

[ITO^{+}17] Tsubasa Iwabuchi, Hiroyuki Togashi, Saki Okubo, Yuka Tsuchiyama, Hiroko Yamaguchi, Jun Hidema, and Jotaro Urabe. pH as a key factor defining the niche space of the alpine crustacean species Daphnia tanakai. Limnology and Oceanography, 62(1):189–199, January 2017. CODEN LIOCAH. ISSN 0024-3590.

Ilic:2019:ERO

[IWF19] Maja Ilić, Christian Werner, and Patrick Fink. Equal relevance of omega-3 and omega-6 polyunsaturated fatty acids for the fitness of Daphnia spp. Limnology and Oceanography, 64(6):2512– 2525, November 2019. CODEN LIOCAH. ISSN 0024-3590.

Jabbari:2019:EHW

[JABZ19] Aidin Jabbari, Josef D. Ackerman, Leon Boegman, and Yingming Zhao. Episodic hypoxia in the western basin of Lake Erie. Limnology and Oceanography, 66(4):2220–2236, September 2019. CODEN LIOCAH. ISSN 0024-3590.

Jenny:2013:SIV

[JAD^{+}13] Jean-Philippe Jenny, Fabien Arnaud, Jean-Marcel Dorioz, Charline Giguet Covex, Victor Frossard, Pierre Sabatier, Laurent Millet, Jean-Louis Reyss, Kazuyo Tachikawa, Edouard Bard, Cécile Pignol, Fayçal Soufi, Olivier Romeyer, and Marie-Elodie Perga. A spatiotemporal investigation of varved sediments highlights the dynamics of hypolimnetic hypoxia in a large hardwater lake over the last 150 years. Limnology and Oceanography, 58(4):1395–1408, July 2013. CODEN LIOCAH. ISSN 0024-3590.

Jolivet:2015:CUN

[JAS^{+}15] Aurélie Jolivet, Lars Asplin, Øivind Strand, Julien Thébault, and Laurent Chauvaud. Coastal upwelling in Norway recorded in Great Scallop shells. Limnology and Oceanography, 60(4): 1265–1275, July 2015. CODEN LIOCAH. ISSN 0024-3590.

Jaeschke:2010:MEA

[JAZ^{+}10] Andrea Jaeschke, Ben Abbas, Matthias Zabel, Ellen C. Hopmans, Stefan Schouten, and Jaap S. Sinninghe Damsté. Molecular evidence for anaerobic ammonium-oxidizing (anammox) bacteria in continental shelf and slope sediments off northwest Africa. Limnology and Oceanography, 55(1):365–376, January 2010. CODEN LIOCAH. ISSN 0024-3590.

Johnson:2019:GDP

[JB19] Matthew D. Johnson and David J. Beaudoin. The genetic diversity of plastids associated with mixotrophic oligotrich ciliates. Limnology and Oceanography, 66(4):2187–2201, September 2019. CODEN LIOCAH. ISSN 0024-3590.

James:2016:ECC

[JBB^{+}16] Rachael H. James, Philippe Bousquet, Ingeborg Bussmann, Matthias Haeckel, Rolf Kipfer, Ira Leifer, Helge Niemann, Ilia Ostrovsky, Jacek Piskozub, Gregor Rehder, Tina Treude, Lisa Vielstädte, and Jens Greinert. Effects of climate change on methane emissions from seafloor sediments in the Arctic Ocean: a review. Limnology and Oceanography, 61(S1):S283–S299, November 2016. CODEN LIOCAH. ISSN 0024-3590.

Jansson:2012:BPH

[JBLJ12] Mats Jansson, Martin Berggren, Hjalmar Laudon, and Anders Jonsson. Bioavailable phosphorus in humic headwater streams in boreal Sweden. Limnology and Oceanography, 57(4):1161–1170, July 2012. CODEN LIOCAH. ISSN 0024-3590.

Johnson:2015:HPV

[JBPM15] Todd L. Johnson, Bianca Brahamsha, Brian Palenik, and Jens Mühle. Halomethane production by vanadium-dependent bromoperoxidase in marine Synechococcus. Limnology and Oceanography, 60(5):1823–1835, September 2015. CODEN LI-OCAH. ISSN 0024-3590.

Jones:2011:FSO

[JBT11] Morris E. Jones, Jordon S. Beckler, and Martial Taillefert. The flux of soluble organic-iron(III) complexes from sediments represents a source of stable iron(III) to estuarine waters and to the continental shelf. Limnology and Oceanography, 56(5):1811– 1823, September 2011. CODEN LIOCAH. ISSN 0024-3590.

Junker:2014:STB

[JC14] James R. Junker and Wyatt F. Cross. Seasonality in the trophic basis of a temperate stream invertebrate assemblage: Importance of temperature and food quality. Limnology and Oceanography, 59(2):507–518, March 2014. CODEN LIOCAH. ISSN 0024-3590.

Jiang:2010:CMS

[JCF^{+}10] Li-Qing Jiang, Wei-Jun Cai, Richard A. Feely, Yongchen Wang, Xianghui Guo, Dwight K. Gledhill, Xinping Hu, Felipe Arzayus, Feizhou Chen, Justin Hartmann, and Longjun Zhang. Carbonate mineral saturation states along the U.S. East Coast. Limnology and Oceanography, 55(6):2424–2432, November 2010. CO-DEN LIOCAH. ISSN 0024-3590.

Jaspers:2018:RMW

[JCS^{+}18] Cornelia Jaspers, John H. Costello, Kelly R. Sutherland, Brad Gemmell, Kelsey N. Lucas, Jennifer Tackett, Kara Dodge, and Sean P. Colin. Resilience in moving water: Effects of turbulence on the predatory impact of the lobate ctenophore Mnemiopsis leidyi. Limnology and Oceanography, 63(4):445–458, January 2018. CODEN LIOCAH. ISSN 0024-3590.

Jensen:2016:LEE

[JD16] Megan M. Jensen and Mark W. Denny. Life in an extreme environment: Characterizing wave-imposed forces in the rocky intertidal zone using high temporal resolution hydrodynamic measurements. Limnology and Oceanography, 62(3):1750–1761, September 2016. CODEN LIOCAH. ISSN 0024-3590.

Johnsen:2014:AUH

[JGR^{+}14] Sönke Johnsen, Ewa Gassmann, Rick A. Reynolds, Dariusz Stramski, and Curtis Mobley. The asymmetry of the underwater horizontal light field and its implications for mirror-based camouflage in silvery pelagic fish. Limnology and Oceanography, 59(6):1839–1852, November 2014. CODEN LIOCAH. ISSN 0024-3590.

Jiang:2011:STD

[JHD^{+}11] Zong-Pei Jiang, Jr-Chuan Huang, Minhan Dai, Shuh Ji Kao, David J. Hydes, Wen-Chen Chou, and Sen Jan. Short-term dynamics of oxygen and carbon in productive nearshore shallow seawater systems off Taiwan: Observations and modeling. Limnology and Oceanography, 56(5):1832–1849, September 2011. CODEN LIOCAH. ISSN 0024-3590.

Jankowiak:2019:DEN

[JHLK^{+}19] Jennifer Jankowiak, Theresa Hattenrath-Lehmann, Benjamin J. Kramer, Megan Ladds, and Christopher J. Gobler. Deciphering the effects of nitrogen, phosphorus, and temperature on cyanobacterial bloom intensification, diversity, and toxicity in western Lake Erie. Limnology and Oceanography, 64(3):1347– 1370, May 2019. CODEN LIOCAH. ISSN 0024-3590.

Jiang:2019:ASE

[JHW^{+}19] Q. Jiang, J. He, J. Wu, X. Hu, G. Ye, and G. Christakos. Assessing the severe eutrophication status and spatial trend in the coastal waters of Zhejiang province (China). Limnology and Oceanography, 64(1):3–17, January 2019. CODEN LIOCAH. ISSN 0024-3590.

Jiang:2017:JOD

[JJ17] Houshuo Jiang and Matthew D. Johnson. Jumping and overcoming diffusion limitation of nutrient uptake in the photosynthetic ciliate Mesodinium rubrum. Limnology and Oceanography, 62(2):421–436, March 2017. CODEN LIOCAH. ISSN 0024-3590.

Jacquot:2013:SCA

[JKKM13] Jeremy E. Jacquot, Yoshiko Kondo, Angela N. Knapp, and James W. Moffett. The speciation of copper across active gradients in nitrogen-cycle processes in the eastern tropical South Pacific. Limnology and Oceanography, 58(4):1387–1394, July 2013. CODEN LIOCAH. ISSN 0024-3590.

Jiang:2015:MAT

[JLC^{+}15] Xiaodong Jiang, Huishuang Liang, Ying Chen, Xing Xu, and Di Huang. Microgeographic adaptation to toxic cyanobacteria in two aquatic grazers. Limnology and Oceanography, 60(3): 947–956, May 2015. CODEN LIOCAH. ISSN 0024-3590.

Jiang:2010:DDN

[JLG10] Xiaodong Jiang, Darcy J. Lonsdale, and Christopher J. Gobler. Density-dependent nutritional value of the dinoflagellate Cochlodinium polykrikoides to the copepod Acartia tonsa. Limnology and Oceanography, 55(4):1643–1652, July 2010. CODEN LIOCAH. ISSN 0024-3590.

Jiang:2011:RGL

[JLG11] Xiaodong Jiang, Darcy J. Lonsdale, and Christopher J. Gobler. Rapid gain and loss of evolutionary resistance to the harmful dinoflagellate Cochlodinium polykrikoides in the copepod Acartia tonsa. Limnology and Oceanography, 57(4):947–954, May 2011. CODEN LIOCAH. ISSN 0024-3590.

Joung:2017:WWL

[JLR^{+}17] DongJoo Joung, Meagan Leduc, Benjamin Ramcharitar, Yaoyang Xu, Peter D. F. Isles, Jason D. Stockwell, Gregory K. Druschel, Tom Manley, and Andrew W. Schroth. Winter weather and lake-watershed physical configuration drive phosphorus, iron, and manganese dynamics in water and sediment of ice-covered lakes. Limnology and Oceanography, 62(4):1620– 1635, July 2017. CODEN LIOCAH. ISSN 0024-3590.

Jimenez:2012:TET

[JLRK12] Isabel M. Jimenez, Anthony W. D. Larkum, Peter J. Ralph, and Michael Kühl. Thermal effects of tissue optics in symbiontbearing reef-building corals. Limnology and Oceanography, 57 (6):1816–1825, November 2012. CODEN LIOCAH. ISSN 0024-3590.

Jakobsen:2016:CCR

[JM16] Hans H. Jakobsen and Stiig Markager. Carbon-to-chlorophyll ratio for phytoplankton in temperate coastal waters: Seasonal patterns and relationship to nutrients. Limnology and Oceanography, 62(3):1853–1868, September 2016. CODEN LIOCAH. ISSN 0024-3590.

Jeffrey:2019:WME

[JMJ^{+}19] Luke C. Jeffrey, Damien T. Maher, Scott G. Johnston, Brendan P. Kelaher, Andy Steven, and Douglas R. Tait. Wetland methane emissions dominated by plant-mediated fluxes: Contrasting emissions pathways and seasons within a shallow freshwater subtropical wetland. Limnology and Oceanography, 64(5): 1895–1912, September 2019. CODEN LIOCAH. ISSN 0024-3590.

Johannessen:2014:ODS

[JMM14] Sophia C. Johannessen, Diane Masson, and Robie W. Macdonald. Oxygen in the deep Strait of Georgia, 1951–2009: The roles of mixing, deep-water renewal, and remineralization of organic carbon. Limnology and Oceanography, 59(1):211–222, January 2014. CODEN LIOCAH. ISSN 0024-3590.

Jung-Madsen:2015:EDC

[JMN15] Signe Jung-Madsen and Torkel Gissel Nielsen. Early development of Calanus glacialis and C. finmarchicus. Limnology and Oceanography, 60(3):934–946, May 2015. CODEN LIOCAH. ISSN 0024-3590.

Jung-Madsen:2013:EDC

[JMNG^{+}13] Signe Jung-Madsen, Torkel Gissel Nielsen, Peter Grønkjær, Benni Winding Hansen, and Eva Friis Møller. Early development of Calanus hyperboreus nauplii: Response to a changing ocean. Limnology and Oceanography, 58(6):2109–2121, November 2013. CODEN LIOCAH. ISSN 0024-3590.

Johnson:2010:SMN

[Joh10] Kenneth S. Johnson. Simultaneous measurements of nitrate, oxygen, and carbon dioxide on oceanographic moorings: Observing the Redfield Ratio in real time. Limnology and Oceanography, 55(2):615–627, March 2010. CODEN LIOCAH. ISSN 0024-3590.

Jorgensen:2010:RSR

[JP10] Bo Barker Jørgensen and R. John Parkes. Role of sulfate reduction and methane production by organic carbon degradation in eutrophic fjord sediments (Limfjorden, Denmark). Limnology and Oceanography, 55(3):1338–1352, May 2010. CODEN LI-OCAH. ISSN 0024-3590.

Joensuu:2018:SPB

[JPH^{+}18] M. Joensuu, C. A. Pilditch, R. Harris, S. Hietanen, H. Pettersson, and A. Norkko. Sediment properties, biota, and local habitat structure explain variation in the erodibility of coastal sediments. Limnology and Oceanography, 63(1):173–186, January 2018. CODEN LIOCAH. ISSN 0024-3590.

Jones:2014:ASI

[JSB^{+}14] H. J. Jones, K. M. Swadling, E. C. V. Butler, L. A. Barry, and C. K. Macleod. Application of stable isotope mixing models for defining trophic biomagnification pathways of mercury and selenium. Limnology and Oceanography, 59(4):1181–1192, July 2014. CODEN LIOCAH. ISSN 0024-3590.

Johnson:2018:VCH

[JSFC18] Pieter T. J. Johnson, Daniel E. Stanton, Kenneth J. Forshay, and Dana M. Calhoun. Vertically challenged: How disease suppresses Daphnia vertical migration behavior. Limnology and Oceanography, 63(2):886–896, March 2018. CODEN LIOCAH. ISSN 0024-3590.

Jankowski:2012:ANS

[JSH12] KathiJo Jankowski, Daniel E. Schindler, and Gordon W. Holtgrieve. Assessing nonpoint-source nitrogen loading and nitrogen fixation in lakes using δ ^{15}N and nutrient stoichiometry. Limnology and Oceanography, 57(3):671–683, May 2012. CODEN LIOCAH. ISSN 0024-3590.

Jorgensen:2015:CCB

[JSK^{+}15] Linda Jørgensen, Colin A. Stedmon, Hermanni Kaartokallio, Mathias Middelboe, and David N. Thomas. Changes in the composition and bioavailability of dissolved organic matter during sea ice formation. Limnology and Oceanography, 60(3):817–830, May 2015. CODEN LIOCAH. ISSN 0024-3590.

Jouandet:2011:OIM

[JTG^{+}11] Marie-Paule Jouandet, Thomas W. Trull, Lionel Guidi, Marc Picheral, Friederike Ebersbach, Lars Stemmann, and Stéphane Blain. Optical imaging of mesopelagic particles indicates deep carbon flux beneath a natural iron-fertilized bloom in the Southern Ocean. Limnology and Oceanography, 57(4):1130–1140, May 2011. CODEN LIOCAH. ISSN 0024-3590.

Jaspers:2011:ICM

[JTH^{+}11] Cornelia Jaspers, Josefin Titelman, Lars Johan Hansson, Matilda Haraldsson, and Christine Røllike Ditlefsen. The invasive ctenophore Mnemiopsis leidyi poses no direct threat to Baltic cod eggs and larva. Limnology and Oceanography, 56(2): 431–439, March 2011. CODEN LIOCAH. ISSN 0024-3590.

Johnson:2013:QPD

[JTH^{+}13] Laura T. Johnson, Jennifer L. Tank, Robert O. Hall, Jr., Patrick J. Mulholland, Stephen K. Hamilton, H. Maurice Valett, Jackson R. Webster, Melody J. Bernot, William H. McDowell, Bruce J. Peterson, and Suzanne M. Thomas. Quantifying the production of dissolved organic nitrogen in headwater streams using ^{15}N tracer additions. Limnology and Oceanography, 58(4): 1271–1285, July 2013. CODEN LIOCAH. ISSN 0024-3590.

Johansson:2016:ABI

[JTV^{+}16] Karin S. L. Johansson, Cristina Trigal, Tobias Vrede, Pieter van Rijswijk, Willem Goedkoop, and Richard K. Johnson. Algal blooms increase heterotrophy at the base of boreal lake food webs-evidence from fatty acid biomarkers. Limnology and Oceanography, 61(5):1563–1573, September 2016. CODEN LI-OCAH. ISSN 0024-3590.

Jones:2014:GSL

[JW14] Erin F. Jones and Wayne A. Wurtsbaugh. The Great Salt Lake’s monimolimnion and its importance for mercury bioaccumulation in brine shrimp (Artemia franciscana). Limnology and Oceanography, 59(1):141–155, January 2014. CODEN LIOCAH. ISSN 0024-3590.

Jonasdottir:2019:LCO

[JWGH19] Sigrún H. Jónasdóttir, Robert J. Wilson, Astthor Gislason, and Michael R. Heath. Lipid content in overwintering Calanus finmarchicus across the subpolar Eastern North Atlantic Ocean. Limnology and Oceanography, 66(4):2029–2043, September 2019. CODEN LIOCAH. ISSN 0024-3590.

Johnson:2015:NRM

[JWS15] Zachary C. Johnson, John J. Warwick, and Rina Schumer. Nitrogen retention in the main channel and two transient storage zones during nutrient addition experiments. Limnology and Oceanography, 60(1):57–77, January 2015. CODEN LIOCAH. ISSN 0024-3590.

Jacobi:2018:EFM

[JYS18] Yuval Jacobi, Gitai Yahel, and Noa Shenkar. Efficient filtration of micron and submicron particles by ascidians from oligotrophic waters. Limnology and Oceanography, 63(S1):S267–S279, March 2018. CODEN LIOCAH. ISSN 0024-3590.

Jian:2018:SDC

[JZZY18] Shan Jian, Hong-Hai Zhang, Jing Zhang, and Gui-Peng Yang. Spatiotemporal distribution characteristics and environmental control factors of biogenic dimethylated sulfur compounds in the East China Sea during spring and autumn. Limnology and Oceanography, 63(S1):S280–S298, March 2018. CODEN LI-OCAH. ISSN 0024-3590.

Koopmans:2015:SOF

[KB15] Dirk J. Koopmans and Peter Berg. Stream oxygen flux and metabolism determined with the open water and aquatic eddy covariance techniques. Limnology and Oceanography, 60(4): 1344–1355, July 2015. CODEN LIOCAH. ISSN 0024-3590.

Karlsson:2012:TOM

[KBA^{+}12] Jan Karlsson, Martin Berggren, Jenny Ask, Pär Byström, Anders Jonsson, Hjalmar Laudon, and Mats Jansson. Terrestrial organic matter support of lake food webs: Evidence from lake metabolism and stable hydrogen isotopes of consumers. Limnology and Oceanography, 57(4):1042–1048, July 2012. CODEN LIOCAH. ISSN 0024-3590.

Karlsson:2014:RCT

[KBA^{+}14] J. Karlsson, M. Berggren, J. Ask, P. Byström, A. Jonsson, H. Laudon, and M. Jansson. Response to comment: Terrestrial support of pelagic consumers in unproductive lakes — Uncertainty and potential in assessments using stable isotopes. Limnology and Oceanography, 59(5):1800–1803, September 2014. CODEN LIOCAH. ISSN 0024-3590. See [Bre14].

Kreling:2017:IPT

[KBE^{+}17] Julika Kreling, Jenny Bravidor, Christof Engelhardt, Michael Hupfer, Matthias Koschorreck, and Andreas Lorke. The importance of physical transport and oxygen consumption for the development of a metalimnetic oxygen minimum in a lake. Limnology and Oceanography, 62(1):348–363, January 2017. CO-DEN LIOCAH. ISSN 0024-3590.

Karlsson:2019:RSP

[KBH^{+}19] Olof Magnus Karlsson, Andreas Christoffer Bryhn, Lars Håkanson, Joakim Hållén, Per Jonsson, Jan Mikael Malmaeus, and Emil Rydin. On the role of sedimentological processes controlling phosphorus burial in the coastal zone of the Baltic Sea. Limnology and Oceanography, 64(4):1828–1831, July 2019. CO-DEN LIOCAH. ISSN 0024-3590. See reply [ACC^{+}19].

Koch:2019:IIL

[KBHT19] Florian Koch, Sara Beszteri, Lars Harms, and Scarlett Trimborn. The impacts of iron limitation and ocean acidification on the cellular stoichiometry, photophysiology, and transcriptome of Phaeocystis antarctica. Limnology and Oceanography, 64(1): 357–375, January 2019. CODEN LIOCAH. ISSN 0024-3590.

Klaus:2018:WRG

[KBJ^{+}18] Marcus Klaus, Ann-Kristin Bergström, Anders Jonsson, Anne Deininger, Erik Geibrink, and Jan Karlsson. Weak response of greenhouse gas emissions to whole lake N enrichment. Limnology and Oceanography, 63(S1):S340–S353, March 2018. CODEN LIOCAH. ISSN 0024-3590.

Krause:2010:EBS

[KBL^{+}10] Jeffrey W. Krause, Mark A. Brzezinski, Michael R. Landry, Stephen B. Baines, David M. Nelson, Karen E. Selph, Andrew G. Taylor, and Benjamin S. Twiningf. The effects of biogenic silica detritus, zooplankton grazing, and diatom size structure on silicon cycling in the euphotic zone of the eastern equatorial Pacific. Limnology and Oceanography, 55(6):2608–2622, November 2010. CODEN LIOCAH. ISSN 0024-3590.

Kreling:2014:PCO

[KBM^{+}14] Julika Kreling, Jenny Bravidor, Daniel F. McGinnis, Matthias Koschorreck, and Andreas Lorke. Physical controls of oxygen fluxes at pelagic and benthic oxyclines in a lake. Limnology and Oceanography, 59(5):1637–1650, September 2014. CODEN LIOCAH. ISSN 0024-3590.

Koehler:2016:AQY

[KBT16] Birgit Koehler, Elias Broman, and Lars J. Tranvik. Apparent quantum yield of photochemical dissolved organic carbon mineralization in lakes. Limnology and Oceanography, 61(6):2207– 2221, November 2016. CODEN LIOCAH. ISSN 0024-3590.

Krause:2012:IKE

[KBVW12] Jeffrey W. Krause, Mark A. Brzezinski, Tracy A. Villareal, and Cara Wilson. Increased kinetic efficiency for silicic acid uptake as a driver of summer diatom blooms in the North Pacific subtropical gyre. Limnology and Oceanography, 57(4):1084–1098, July 2012. CODEN LIOCAH. ISSN 0024-3590.

Kraus:2017:RSL

[KCB^{+}17] Tamara E. C. Kraus, Kurt D. Carpenter, Brian A. Bergamaschi, Alexander E. Parker, Elizabeth B. Stumpner, Bryan D. Downing, Nicole M. Travis, Frances P. Wilkerson, Carol Kendall, and Timothy D. Mussen. A river-scale Lagrangian experiment examining controls on phytoplankton dynamics in the presence and absence of treated wastewater effluent high in ammonium. Limnology and Oceanography, 63(3):1234–1253, May 2017. CODEN LIOCAH. ISSN 0024-3590.

Karnauskas:2012:PFI

[KCH^{+}12] Mandy Karnauskas, Laurent M. Chérubin, Brittany E. Huntington, Elizabeth A. Babcock, and Dennis A. Thoney. Physical forces influence the trophic structure of reef fish communities on a remote atoll. Limnology and Oceanography, 58(1):1403–1414, September 2012. CODEN LIOCAH. ISSN 0024-3590.

Khripounoff:2014:DCW

[KCL^{+}14] Alexis Khripounoff, Jean-Claude Caprais, Julie Le Bruchec, Philippe Rodier, Philippe Noel, and Cécile Cathalot. Deep cold-water coral ecosystems in the Brittany submarine canyons (Northeast Atlantic): Hydrodynamics, particle supply, respiration, and carbon cycling. Limnology and Oceanography, 59(1): 87–98, January 2014. CODEN LIOCAH. ISSN 0024-3590.

Katsev:2010:MEB

[KCM^{+}10] Sergei Katsev, Sean A. Crowe, Alfonso Mucci, Bjørn Sundby, Sulung Nomosatryo, G. Douglas Haffner, and David A. Fowle. Mixing and its effects on biogeochemistry in the persistently stratified, deep, tropical Lake Matano, Indonesia. Limnology and Oceanography, 55(4):763–776, March 2010. CODEN LI-OCAH. ISSN 0024-3590.

Kelly:2019:IAI

[KDGL19] Jacque L. Kelly, Henrietta Dulai, Craig R. Glenn, and Paul G. Lucey. Integration of aerial infrared thermography and in situ radon-222 to investigate submarine groundwater discharge to Pearl Harbor, Hawaii, USA. Limnology and Oceanography, 64 (1):238–257, January 2019. CODEN LIOCAH. ISSN 0024-3590.

Kavanaugh:2014:PBC

[KEH^{+}14] Maria T. Kavanaugh, Steven R. Emerson, Burke Hales, Deirdre M. Lockwood, Paul D. Quay, and Ricardo M. Letelier. Physicochemical and biological controls on primary and net community production across northeast Pacific seascapes. Limnology and Oceanography, 59(6):2013–2027, November 2014. CODEN LIOCAH. ISSN 0024-3590.

Kerr:2017:MLU

[Ker17] Jason G. Kerr. Multiple land use activities drive riverine salinization in a large, semi-arid river basin in western Canada. Limnology and Oceanography, 62(4):1331–1345, July 2017. CODEN LIOCAH. ISSN 0024-3590.

Kleeberg:2013:EIC

[KFJ13] Andreas Kleeberg, Andreas Freidank, and Klaus Jöhnk. Effects of ice cover on sediment resuspension and phosphorus entrainment in shallow lakes: Combining in situ experiments and windwave modeling. Limnology and Oceanography, 58(5):1819–1833, September 2013. CODEN LIOCAH. ISSN 0024-3590.

Kopec:2018:UPC

[KFP^{+}18] B. G. Kopec, X. Feng, E. S. Posmentier, J. W. Chipman, and R. A. Virginia. Use of principal component analysis to extract environmental information from lake water isotopic compositions. Limnology and Oceanography, 63(3):1340–1354, May 2018. CODEN LIOCAH. ISSN 0024-3590.

Kang:2018:BTA

[KG18] Yoonja Kang and Christopher J. Gobler. The brown tide algae, Aureococcus anophagefferens and Aureoumbra lagunensis (Pelagophyceae), allelopathically inhibit the growth of competing microalgae during harmful algal blooms. Limnology and Oceanography, 63(2):985–1003, March 2018. CODEN LIOCAH. ISSN 0024-3590.

Kessler:2012:QDR

[KGC^{+}12] Adam J. Kessler, Ronnie N. Glud, M. Bayani Cardenas, Morten Larsen, Michael F. Bourke, and Perran L. M. Cook. Quantifying denitrification in rippled permeable sands through combined flume experiments and modeling. Limnology and Oceanography, 57(4):1217–1232, July 2012. CODEN LIOCAH. ISSN 0024-

Kiorboe:2016:RCP

[KGC^{+}16] Thomas Kiørboe, Rodrigo J. Gonçalves, Damien Couespel, Hans van Someren Gréve, Enric Saiz, and Peter Tiselius. Reply to comment: Prey perception in feeding-current feeding copepods. Limnology and Oceanography, 61(4):1169–1171, July 2016. CO-DEN LIOCAH. ISSN 0024-3590. See [PJ16].

Kohlbach:2016:IIA

[KGL^{+}16] Doreen Kohlbach, Martin Graeve, Benjamin A. Lange, Carmen David, Ilka Peeken, and Hauke Flores. The importance of ice algae-produced carbon in the central Arctic Ocean ecosystem: Food web relationships revealed by lipid and stable isotope analyses. Limnology and Oceanography, 61(6):2027–2044, November 2016. CODEN LIOCAH. ISSN 0024-3590.

Kimmerer:2014:TMR

[KGM14] Wim J. Kimmerer, Edward S. Gross, and Michael L. MacWilliams. Tidal migration and retention of estuarine zooplankton investigated using a particle-tracking model. Limnology and Oceanography, 59(3):901–916, May 2014. CODEN LIOCAH. ISSN 0024-3590.

Kelly:2018:ILA

[KGRV18] Patrick T. Kelly, María J. González, William H. Renwick, and Michael J. Vanni. Increased light availability and nutrient cycling by fish provide resilience against reversing eutrophication in an agriculturally impacted reservoir. Limnology and Oceanography, 63(6):2647–2660, November 2018. CODEN LIOCAH. ISSN 0024-3590.

Kirillin:2012:MSR

[KGT12] Georgiy Kirillin, Hans-Peter Grossart, and Kam W. Tang. Modeling sinking rate of zooplankton carcasses: Effects of stratification and mixing. Limnology and Oceanography, 57(3):881–894, May 2012. CODEN LIOCAH. ISSN 0024-3590.

Knysh:2016:IAL

[KGvdH16] Kyle M. Knysh, Donna J. Giberson, and Michael R. van den Heuvel. The influence of agricultural land-use on plant and macroinvertebrate communities in springs. Limnology and Oceanography, 61(2):518–530, March 2016. CODEN LIOCAH. ISSN 0024-3590.

Kapsenberg:2016:OPT

[KH16] Lydia Kapsenberg and Gretchen E. Hofmann. Ocean pH timeseries and drivers of variability along the northern Channel Islands, California, USA. Limnology and Oceanography, 61(3): 953–968, May 2016. CODEN LIOCAH. ISSN 0024-3590.

Kirchman:2014:MAO

[KHCH14] David L. Kirchman, Thomas E. Hanson, Matthew T. Cottrell, and Leila J. Hamdan. Metagenomic analysis of organic matter degradation in methane-rich Arctic Ocean sediments. Limnology and Oceanography, 59(6):548–559, March 2014. CODEN LIOCAH. ISSN 0024-3590.

Kleeberg:2013:IRR

[KHG^{+}13] Andreas Kleeberg, Michael Hupfer, Giselher Gust, Ivette Salka, Kirsten Pohlmann, and Hans-Peter Grossart. Intermittent riverine resuspension: Effects on phosphorus transformations and heterotrophic bacteria. Limnology and Oceanography, 58(2): 635–652, March 2013. CODEN LIOCAH. ISSN 0024-3590.

Kendrick:2019:DNA

[KHH19] Michael R. Kendrick, Anne E. Hershey, and Alexander D. Huryn. Disturbance, nutrients, and antecedent flow conditions affect macroinvertebrate community structure and productivity in an Arctic river. Limnology and Oceanography, 64(S1):S93– S104, January 2019. CODEN LIOCAH. ISSN 0024-3590.

Kubo:2019:LTV

[KHK^{+}19] Atsushi Kubo, Fuminori Hashihama, Jota Kanda, Naho Horimoto-Miyazaki, and Takashi Ishimaru. Long-term variability of nutrient and dissolved organic matter concentrations in Tokyo Bay between 1989 and 2015. Limnology and Oceanography, 64(S1):S209–S222, January 2019. CODEN LIOCAH. ISSN 0024-3590.

Krempaska:2018:SDM

[KHP18] Natalia Krempaska, Karel Horñák, and Jakob Pernthaler. Spatiotemporal distribution and microbial assimilation of polyamines in a mesotrophic lake. Limnology and Oceanography, 63(2):816–832, March 2018. CODEN LIOCAH. ISSN 0024-3590.

Kemp:2014:CDP

[KHPIP^{+}14] Dustin W. Kemp, Xavier Hernandez-Pech, Roberto Iglesias-Prieto, William K. Fitt, and Gregory W. Schmidt. Community dynamics and physiology of Symbiodinium spp. before, during, and after a coral bleaching event. Limnology and Oceanography, 59(3):788–797, May 2014. CODEN LIOCAH. ISSN 0024-3590.

Kankaala:2013:LSD

[KHTO13] Paula Kankaala, Jussi Huotari, Tiina Tulonen, and Anne Ojala. Lake-size dependent physical forcing drives carbon dioxide and methane effluxes from lakes in a boreal landscape. Limnology and Oceanography, 58(6):1915–1930, November 2013. CODEN LIOCAH. ISSN 0024-3590.

Kopacek:2011:PLM

[KHVS11] Jiří Kopàček, Josef Hejzlar, Jaroslav Vrba, and Evžen Stuchlík. Phosphorus loading of mountain lakes: Terrestrial export and atmospheric deposition. Limnology and Oceanography, 56(4): 1343–1354, July 2011. CODEN LIOCAH. ISSN 0024-3590.

Katz:2015:MYL

[KIH^{+}15] Stephen L. Katz, Lyubov R. Izmest’eva, Stephanie E. Hampton, Ted Ozersky, Kirill Shchapov, Marianne V. Moore, Svetlana V. Shimaraeva, and Eugene A. Silow. The “Melosira years” of Lake Baikal: Winter environmental conditions at ice onset predict under-ice algal blooms in spring. Limnology and Oceanography, 60(6):1950–1964, November 2015. CODEN LIOCAH. ISSN 0024-3590.

Kiorboe:2013:ZBC

[Kiø13] Thomas Kiørboe. Zooplankton body composition. Limnology and Oceanography, 58(5):1843–1850, September 2013. CODEN LIOCAH. ISSN 0024-3590.

Kirk:2013:TRE

[Kir13] John T. O. Kirk. The total radiant energy, and the average depth of all the photons, in the water column. Limnology and Oceanography, 58(2):93–98, January 2013. CODEN LIOCAH. ISSN 0024-3590.

Koop-Jakobsen:2010:EIN

[KJG10] Ketil Koop-Jakobsen and Anne E. Giblin. The effect of increased nitrate loading on nitrate reduction via denitrification and DNRA in salt marsh sediments. Limnology and Oceanography, 55(4):789–802, March 2010. CODEN LIOCAH. ISSN 0024-3590.

Kitahashi:2018:HRH

[KJKS18] Tomo Kitahashi, Robert G. Jenkins, Shigeaki Kojima, and Motohiro Shimanaga. High resilience of harpacticoid copepods in the landward slope of the Japan Trench against disturbance of the 2011 Tohoku Earthquake. Limnology and Oceanography, 63 (6):2751–2761, November 2018. CODEN LIOCAH. ISSN 0024-3590.

Klevjer:2011:KMN

[KK11] Thor A. Klevjer and Stein Kaartvedt. Krill (Meganyctiphanes norvegica) swim faster at night. Limnology and Oceanography, 56(3):765–774, May 2011. CODEN LIOCAH. ISSN 0024-3590.

Kim:2013:CSP

[KK13] Tae-Hoon Kim and Guebuem Kim. Changes in seawater N : P ratios in the northwestern Pacific Ocean in response to increasing atmospheric N deposition: Results from the East (Japan) Sea. Limnology and Oceanography, 58(6):1907–1914, November 2013. CODEN LIOCAH. ISSN 0024-3590.

Karatayev:2018:FDR

[KKB^{+}18] Alexander Y. Karatayev, Vadim A. Karatayev, Lyubov E. Burlakova, Mark D. Rowe, Knut Mehler, and Mark D. Clapsadl. Food depletion regulates the demography of invasive dreissenid mussels in a stratified lake. Limnology and Oceanography, 64 (4):2065–2079, September 2018. CODEN LIOCAH. ISSN 0024-3590.

Kim:2011:SGD

[KKH11] Guebuem Kim, Jong-Sun Kim, and Dong-Woon Hwang. Submarine groundwater discharge from oceanic islands standing in oligotrophic oceans: Implications for global biological production and organic carbon fluxes. Limnology and Oceanography, 56 (2):673–682, March 2011. CODEN LIOCAH. ISSN 0024-3590.

Kotterba:2014:PTS

[KKHP14] Paul Kotterba, Carsten Kühn, Cornelius Hammer, and Patrick Polte. Predation of threespine stickleback (Gasterosteus aculeatus) on the eggs of Atlantic herring (Clupea harengus) in a Baltic Sea lagoon. Limnology and Oceanography, 59(6):578–587, March 2014. CODEN LIOCAH. ISSN 0024-3590.

Kopacek:2019:ETD

[KKP^{+}19] Jiří Kopáček, Jiří Kaña, Petr Porcal, Jaroslav Vrba, and Stephen A. Norton. Effects of tree dieback on lake water acidity in the unmanaged catchment of Plešné Lake, Czech Republic. Limnology and Oceanography, 64(4):1614–1626, July 2019. CO-DEN LIOCAH. ISSN 0024-3590.

Kissman:2010:DMD

[KKS10] Carrie E. H. Kissman, Lesley B. Knoll, and Orlando Sarnelle. Dreissenid mussels (Dreissena polymorpha and Dreissena bugensis) reduce microzooplankton and macrozooplankton biomass in thermally stratified lakes. Limnology and Oceanography, 55(5): 1851–1859, September 2010. CODEN LIOCAH. ISSN 0024-3590.

Ko:2016:OAP

[KLEH16] Young Ho Ko, Kitack Lee, Ki Hyuk Eom, and In-Seong Han. Organic alkalinity produced by phytoplankton and its effect on the computation of ocean carbon parameters. Limnology and Oceanography, 61(4):1462–1471, July 2016. CODEN LIOCAH. ISSN 0024-3590.

Kristensen:2017:TDS

[KLM^{+}17] Erik Kristensen, Shing Yip Lee, Perrine Mangion, Cintia O. Quintana, and Thomas Valdemarsen. Trophic discrimination of stable isotopes and potential food source partitioning by leaf-eating crabs in mangrove environments. Limnology and Oceanography, 65(9):2097–2112, September 2017. CODEN LI-OCAH. ISSN 0024-3590.

Kerfoot:2010:TSD

[KM10] W. Charles Kerfoot and A. Scott McNaught. Two-step dialogue between the cladoceran bosmina and invertebrate predators: Induction and natural selection. Limnology and Oceanography, 55 (1):403–419, January 2010. CODEN LIOCAH. ISSN 0024-3590.

Korosi:2015:LTC

[KMC^{+}15] Jennifer B. Korosi, Jamylynn McDonald, Kristen A. Coleman, Michael J. Palmer, John P. Smol, Myrna J. Simpson, and Jules M. Blais. Long-term changes in organic matter and mercury transport to lakes in the sporadic discontinuous permafrost zone related to peat subsidence. Limnology and Oceanography, 60(5):1550–1561, September 2015. CODEN LIOCAH. ISSN 0024-3590.

Korbee:2010:PMA

[KMF10] Nathalie Korbee, M. Teresa Mata, and Félix L. Figueroa. Photoprotection mechanisms against ultraviolet radiation in Heterocapsa sp. (Dinophyceae) are influenced by nitrogen availability: Mycosporine-like amino acids vs. xanthophyll cycle. Limnology and Oceanography, 55(4):899–908, March 2010. CODEN LIOCAH. ISSN 0024-3590.

Kotterba:2017:PAH

[KMH^{+}17] Paul Kotterba, Dorothee Moll, Cornelius Hammer, Myron A. Peck, Daniel Oesterwind, and Patrick Polte. Predation on Atlantic herring (Clupea harengus) eggs by the resident predator community in coastal transitional waters. Limnology and Oceanography, 62(6):2616–2628, November 2017. CODEN LI-OCAH. ISSN 0024-3590.

Koch:2011:EVB

[KMP^{+}11] Florian Koch, Maria Alejandra Marcoval, Caterina Panzeca, Kenneth W. Bruland, Sergio A. Sañudo-Wilhelmy, and Christopher J. Gobler. The effect of vitamin B_{12} on phytoplankton growth and community structure in the Gulf of Alaska. Limnology and Oceanography, 57(4):1023–1034, May 2011. CODEN LIOCAH. ISSN 0024-3590.

Koussoroplis:2014:FFC

[KNA^{+}14] Apostolos-Manuel Koussoroplis, Julia Nussbaumer, Michael T. Arts, Irina A. Guschina, and Martin J. Kainz. Famine and feast in a common freshwater calanoid: Effects of diet and temperature on fatty acid dynamics of Eudiaptomus gracilis. Limnology and Oceanography, 59(3):947–958, May 2014. CODEN LIOCAH. ISSN 0024-3590.

Krause:2010:PDA

[KNL10] Jeffrey W. Krause, David M. Nelson, and Michael W. Lomas. Production, dissolution, accumulation, and potential export of biogenic silica in a Sargasso Sea mode-water eddy. Limnology and Oceanography, 55(2):569–579, March 2010. CODEN LI-OCAH. ISSN 0024-3590.

Kuehn:2011:CFC

[KOFN11] K. A. Kuehn, B. M. Ohsowski, S. N. Francoeur, and R. K. Neely. Contributions of fungi to carbon flow and nutrient cycling from standing dead Typha angustifolia leaf litter in a temperate freshwater marsh. Limnology and Oceanography, 56(2):529–539, March 2011. CODEN LIOCAH. ISSN 0024-3590.

Kamennaya:2013:DEC

[KP13] Nina A. Kamennaya and Anton F. Post. Distribution and expression of the cyanate acquisition potential among cyanobacterial populations in oligotrophic marine waters. Limnology and Oceanography, 58(6):1959–1971, November 2013. CODEN LI-OCAH. ISSN 0024-3590.

Komada:2012:TCA

[KPJ12] Tomoko Komada, Jonathon A. Polly, and Leah Johnson. Transformations of carbon in anoxic marine sediments: Implications from Δ ^{14}C and δ ^{13}C signatures. Limnology and Oceanography, 57(3):567–581, March 2012. CODEN LIOCAH. ISSN 0024-3590.

Kroll:2018:QES

[KPP^{+}18] Ian R. Kroll, Abigail K. Poray, Brandon J. Puckett, David B. Eggleston, and F. Joel Fodrie. Quantifying estuarine-scale invertebrate larval connectivity: Methodological and ecological insights. Limnology and Oceanography, 64(4):1979–1991, September 2018. CODEN LIOCAH. ISSN 0024-3590.

Kieber:2010:PBP

[KPSW10] Robert J. Kieber, Jaclyn Pitt, Stephen A. Skrabal, and Jeffrey L. C. Wright. Photodegradation of the brevetoxin PbTx-2 in coastal seawater. Limnology and Oceanography, 55(6):2299– 2304, November 2010. CODEN LIOCAH. ISSN 0024-3590.

Kruk:2011:PCC

[KPV^{+}11] C. Kruk, E. T. H. M. Peeters, E. H. Van Nes, V. L. M. Huszar, L. S. Costa, and M. Scheffer. Phytoplankton community composition can be predicted best in terms of morphological groups. Limnology and Oceanography, 56(1):110–118, January 2011. CODEN LIOCAH. ISSN 0024-3590.

Kim:2011:LAO

[KPW^{+}11] J.-H. Kim, F. Peterse, V. Willmott, D. Klitgaard Kristensen, M. Baas, S. Schouten, and J. S. Sinninghe Damsté. Large ancient organic matter contributions to Arctic marine sediments (Svalbard). Limnology and Oceanography, 56(4):1463–1474, July 2011. CODEN LIOCAH. ISSN 0024-3590.

Kominoski:2018:ENP

[KRB^{+}18] John S. Kominoski, Amy D. Rosemond, Jonathan P. Benstead, Vlad Gulis, and David W. P. Manning. Experimental nitrogen and phosphorus additions increase rates of stream ecosystem respiration and carbon loss. Limnology and Oceanography, 63 (1):22–36, January 2018. CODEN LIOCAH. ISSN 0024-3590.

Kottmeier:2016:HDI

[KRR16] Dorothee M. Kottmeier, Sebastian D. Rokitta, and Björn Rost. H^{+}-driven increase in CO_{2} uptake and decrease in HCO_{3} uptake explain coccolithophores’ acclimation responses to ocean acidification. Limnology and Oceanography, 61(6):2045–2057, November 2016. CODEN LIOCAH. ISSN 0024-3590.

Kolb:2013:IAD

[KS13] Amelia Kolb and Suzanne Strom. An inducible antipredatory defense in haploid cells of the marine microalga Emiliania huxleyi (Prymnesiophyceae). Limnology and Oceanography, 58(3): 932–944, May 2013. CODEN LIOCAH. ISSN 0024-3590.

Kerfoot:2016:MIA

[KS16] W. Charles Kerfoot and Susan C. Savage. Multiple inducers in aquatic foodwebs: Counter-measures and vulnerability to exotics. Limnology and Oceanography, 61(1):382–406, January 2016. CODEN LIOCAH. ISSN 0024-3590.

Keesing:2013:SIS

[KSFT13] John K. Keesing, Joanna Strzelecki, Jane Fromont, and Damian Thomson. Sponges as important sources of nitrate on an oligotrophic continental shelf. Limnology and Oceanography, 58(6): 1947–1958, November 2013. CODEN LIOCAH. ISSN 0024-3590.

Knee:2010:NIC

[KSG^{+}10] Karen L. Knee, Joseph H. Street, Eric E. Grossman, Alexandria B. Boehm, and Adina Paytan. Nutrient inputs to the coastal ocean from submarine groundwater discharge in a groundwater-dominated system: Relation to land use (Kona coast, Hawaii, U.S.A.). Limnology and Oceanography, 55(3): 1105–1122, May 2010. CODEN LIOCAH. ISSN 0024-3590.

Kruk:2012:PSP

[KSP^{+}12] C. Kruk, A. M. Segura, E. T. H. M. Peeters, V. L. M. Huszar, L. S. Costa, S. Kosten, G. Lacerot, and M. Scheffer. Phytoplankton species predictability increases towards warmer regions. Limnology and Oceanography, 57(4):1126–1135, July 2012. CO-DEN LIOCAH. ISSN 0024-3590.

Kiorboe:2018:AFB

[KSTA18a] Thomas Kiørboe, Enric Saiz, Peter Tiselius, and Ken H. Andersen. Adaptive feeding behavior and functional responses in zooplankton. Limnology and Oceanography, 63(1):308–321, January 2018. CODEN LIOCAH. ISSN 0024-3590.

Kiorboe:2018:III

[KSTA18b] Thomas Kiørboe, Enric Saiz, Peter Tiselius, and Ken H. Andersen. Issue information — instr to contrib. Limnology and Oceanography, 63(1):??, July 2018. CODEN LIOCAH. ISSN 0024-3590.

Koch:2013:EVB

[KSWFG13] Florian Koch, Sergio A. Sañudo-Wilhelmy, Nicholas S. Fisher, and Christopher J. Gobler. Effect of vitamins B_{1} and B_{12} on bloom dynamics of the harmful brown tide alga, Aureococcus anophagefferens (Pelagophyceae). Limnology and Oceanography, 58(5):1761–1774, September 2013. CODEN LIOCAH. ISSN 0024-3590.

Kramer:2011:EMB

[KSY11] Andrew M. Kramer, Orlando Sarnelle, and Jeannette Yen. The effect of mating behavior and temperature variation on the critical population density of a freshwater copepod. Limnology and Oceanography, 56(2):707–715, March 2011. CODEN LIOCAH. ISSN 0024-3590.

Keister:2013:EBL

[KT13] Julie E. Keister and Loren B. Tuttle. Effects of bottom-layer hypoxia on spatial distributions and community structure of mesozooplankton in a sub-estuary of Puget Sound, Washington, U.S.A. Limnology and Oceanography, 58(2):667–680, March 2013. CODEN LIOCAH. ISSN 0024-3590.

Kuffner:2019:IEC

[KTH^{+}19] Ilsa B. Kuffner, Lauren T. Toth, J. Harold Hudson, William B. Goodwin, Anastasios Stathakopoulos, Lucy A. Bartlett, and Elizabeth M. Whitcher. Improving estimates of coral reef construction and erosion with in situ measurements. Limnology and Oceanography, 66(4):2283–2294, September 2019. CODEN LIOCAH. ISSN 0024-3590.

Klais:2013:SPC

[KTK^{+}13] Riina Klais, Timo Tamminen, Anke Kremp, Kristian Spilling, Byoung Woong An, Susanna Hajdu, and Kalle Olli. Spring phytoplankton communities shaped by interannual weather variability and dispersal limitation: Mechanisms of climate change effects on key coastal primary producers. Limnology and Oceanography, 58(2):753–762, March 2013. CODEN LIOCAH. ISSN 0024-3590.

Kremer:2017:TSS

[KTL17] Colin T. Kremer, Mridul K. Thomas, and Elena Litchman. Temperature- and size-scaling of phytoplankton population growth rates: Reconciling the eppley curve and the metabolic theory of ecology. Limnology and Oceanography, 62(4):1658– 1670, July 2017. CODEN LIOCAH. ISSN 0024-3590.

Kaartvedt:2011:BAD

[KTRK11] Stein Kaartvedt, Josefin Titelman, Anders Røstad, and Thor A. Klevjer. Beyond the average: Diverse individual migration patterns in a population of mesopelagic jellyfish. Limnology and Oceanography, 56(6):2189–2199, November 2011. CODEN LI-OCAH. ISSN 0024-3590.

Kitidis:2014:OPM

[KTS^{+}14] Vassilis Kitidis, Gavin H. Tilstone, Pablo Serret, Timothy J. Smyth, Ricardo Torres, and Carol Robinson. Oxygen photolysis in the Mauritanian upwelling: Implications for net community production. Limnology and Oceanography, 59(2):299–310, March 2014. CODEN LIOCAH. ISSN 0024-3590.

Kuss:2014:WAG

[Kus14] Joachim Kuss. Water–air gas exchange of elemental mercury: an experimentally determined mercury diffusion coefficient for hg^{0} water–air flux calculations. Limnology and Oceanography, 59(5):1461–1467, September 2014. CODEN LIOCAH. ISSN 0024-3590.

Kenitz:2018:RCW

[KVA18] K. M. Kenitz, A. W. Visser, and K. H. Andersen. Response to comment: “What drives plankton seasonality in a stratifying shelf sea? Some competing and complementary theories”. Limnology and Oceanography, 63(6):2885–2886, November 2018. CODEN LIOCAH. ISSN 0024-3590. See [APF^{+}18].

Kulk:2018:PNL

[KvdPB18] Gemma Kulk, Willem H. van de Poll, and Anita G. J. Buma. Photophysiology of nitrate limited phytoplankton communities in Kongsfjorden, Spitsbergen. Limnology and Oceanography, 63 (6):2606–2617, November 2018. CODEN LIOCAH. ISSN 0024-3590.

Kulk:2013:LNA

[KvdPVB13] Gemma Kulk, Willem H. van de Poll, Ronald J. W. Visser, and Anita G. J. Buma. Low nutrient availability reduces highirradiance-induced viability loss in oceanic phytoplankton. Limnology and Oceanography, 58(5):1747–1760, September 2013. CODEN LIOCAH. ISSN 0024-3590.

Kenitz:2017:SSZ

[KVMA17] Kasia M. Kenitz, André W. Visser, Patrizio Mariani, and Ken H. Andersen. Seasonal succession in zooplankton feeding traits reveals trophic trait coupling. Limnology and Oceanography, 63 (3):1184–1197, May 2017. CODEN LIOCAH. ISSN 0024-3590.

Kruger:2016:OMT

[KWB^{+}16] B. R. Kruger, J. P. Werne, D. K. Branstrator, T. R. Hrabik, Y. Chikaraishi, N. Ohkouchi, and E. C. Minor. Organic matter transfer in Lake Superior’s food web: Insights from bulk and molecular stable isotope and radiocarbon analyses. Limnology and Oceanography, 61(1):149–164, January 2016. CODEN LI-OCAH. ISSN 0024-3590.

Kremer:2017:RPT

[KWF^{+}17] Colin T. Kremer, Alicia K. Williams, Michael Finiguerra, Allison A. Fong, Anne Kellerman, Sara F. Paver, Bradley B. Tolar, and Benjamin J. Toscano. Realizing the potential of trait-based aquatic ecology: New tools and collaborative approaches. Limnology and Oceanography, 62(1):253–271, January 2017. CO-DEN LIOCAH. ISSN 0024-3590.

Kranz:2010:CCP

[KWGN^{+}10] Sven A. Kranz, Dieter Wolf-Gladrow, Gernot Nehrke, Gerald Langer, and Björn Rosta. Calcium carbonate precipitation induced by the growth of the marine cyanobacteria Trichodesmium. Limnology and Oceanography, 55(6):2563–2569, November 2010. CODEN LIOCAH. ISSN 0024-3590.

Kohler:2018:IVM

[KWGS18] Jan Köhler, Lan Wang, Alexis Guislain, and Tom Shatwell. Influence of vertical mixing on light-dependency of phytoplankton growth. Limnology and Oceanography, 63(3):1156–1167, May 2018. CODEN LIOCAH. ISSN 0024-3590.

Kessler:2019:CBP

[KWM^{+}19] Adam J. Kessler, Michaela Wawryk, Ugo Marzocchi, Keryn L. Roberts, Wei Wen Wong, Nils Risgaard-Petersen, Filip J. R. Meysman, Ronnie N. Glud, and Perran L. M. Cook. Cable bacteria promote DNRA through iron sulfide dissolution. Limnology and Oceanography, 64(3):1228–1238, May 2019. CODEN LIOCAH. ISSN 0024-3590.

Kissman:2013:RPA

[KWRS13] Carrie E. H. Kissman, Craig E. Williamson, Kevin C. Rose, and Jasmine E. Saros. Response of phytoplankton in an alpine lake to inputs of dissolved organic matter through nutrient enrichment and trophic forcing. Limnology and Oceanography, 58(3): 867–880, May 2013. CODEN LIOCAH. ISSN 0024-3590.

Kahn:2015:BGC

[KYC^{+}15] Amanda S. Kahn, Gitai Yahel, Jackson W. F. Chu, Verena Tunnicliffe, and Sally P. Leys. Benthic grazing and carbon sequestration by deep-water glass sponge reefs. Limnology and Oceanography, 60(1):78–88, January 2015. CODEN LIOCAH. ISSN 0024-3590.

Kerfoot:2012:LDR

[KYG^{+}12] W. Charles Kerfoot, Foad Yousef, Sarah A. Green, Robert Regis, Robert Shuchman, Colin N. Brooks, Mike Sayers, Bruce Sabol, and Mark Graves. Light detection and ranging (LiDAR) and multispectral studies of disturbed Lake Superior coastal environments. Limnology and Oceanography, 57(3):749–771, May 2012. CODEN LIOCAH. ISSN 0024-3590.

Katz:2012:RFF

[KYR^{+}12] Timor Katz, Gitai Yahel, Matt Reidenbach, Verena Tunnicliffe, Barak Herut, John Crusius, Frank Whitney, Paul V. R. Snelgrove, and Boaz Lazar. Resuspension by fish facilitates the transport and redistribution of coastal sediments. Limnology and Oceanography, 57(4):945–958, July 2012. CODEN LIOCAH. ISSN 0024-3590.

Kane:2018:KMS

[KYRMD18] Mary K. Kane, Regina Yopak, Christopher Roman, and Susanne Menden-Deuer. Krill motion in the Southern Ocean: quantifying in situ krill movement behaviors and distributions during the late austral autumn and spring. Limnology and Oceanography, 63(6):2839–2857, November 2018. CODEN LIOCAH. ISSN 0024-3590.

Kim:2010:CRB

[KZB^{+}10] J.-H. Kim, B. Zarzycka, R. Buscail, F. Peterse, J. Bonnin, W. Ludwig, S. Schouten, and J. S. Sinninghe Damsté. Contribution of river-borne soil organic carbon to the Gulf of Lions (NW Mediterranean). Limnology and Oceanography, 55(2):507– 518, March 2010. CODEN LIOCAH. ISSN 0024-3590.

Kouraev:2016:GIR

[KZR^{+}16] Alexei V. Kouraev, Elena A. Zakharova, Frédérique Rémy, Andrey G. Kostianoy, Michail N. Shimaraev, Nicholas M. J. Hall, and Andrey Ya. Suknev. Giant ice rings on lakes Baikal and Hovsgol: Inventory, associated water structure and potential formation mechanism. Limnology and Oceanography, 61(3):1001– 1014, May 2016. CODEN LIOCAH. ISSN 0024-3590.

Kouraev:2019:GIR

[KZR^{+}19] Alexei V. Kouraev, Elena A. Zakharova, Frédérique Rémy, Andrey G. Kostianoy, Mikhail N. Shimaraev, Nicholas M. J. Hall, Roman E. Zdorovennov, and Andrey Ya Suknev. Giant ice rings on lakes and field observations of lens-like eddies in the Middle Baikal (2016–2017). Limnology and Oceanography, 64(6):2738– 2754, November 2019. CODEN LIOCAH. ISSN 0024-3590.

Langdon:2018:TTC

[LABJ18] Chris Langdon, Rebecca Albright, Andrew C. Baker, and Paul Jones. Two threatened Caribbean coral species have contrasting responses to combined temperature and acidification stress.

Limnology and Oceanography, 63(6):2450–2464, November 2018. CODEN LIOCAH. ISSN 0024-3590.

Lewis:2019:PAO

[LAC^{+}19] K. M. Lewis, A. E. Arntsen, P. Coupel, H. Joy-Warren, K. E. Lowry, A. Matsuoka, M. M. Mills, G. L. van Dijken, V. Selz, and K. R. Arrigo. Photoacclimation of Arctic Ocean phytoplankton to shifting light and nutrient limitation. Limnology and Oceanography, 64(1):284–301, January 2019. CODEN LI-OCAH. ISSN 0024-3590.

Laborde:2010:IIM

[LACI10] S. Laborde, J. P. Antenucci, D. Copetti, and J. Imberger. Inflow intrusions at multiple scales in a large temperate lake. Limnology and Oceanography, 55(3):1301–1312, May 2010. CODEN LIOCAH. ISSN 0024-3590.

Lopez-Acosta:2018:SCK

[LALGM18] María López-Acosta, Aude Leynaert, Jacques Grall, and Manuel Maldonado. Silicon consumption kinetics by marine sponges: an assessment of their role at the ecosystem level. Limnology and Oceanography, 63(6):2508–2522, November 2018. CODEN LIOCAH. ISSN 0024-3590.

Lopez-Acosta:2016:SCT

[LALM16] María López-Acosta, Aude Leynaert, and Manuel Maldonado. Silicon consumption in two shallow-water sponges with contrasting biological features. Limnology and Oceanography, 61(6): 2139–2150, November 2016. CODEN LIOCAH. ISSN 0024-3590.

Lathlean:2012:UII

[LAM12] Justin A. Lathlean, David J. Ayre, and Todd E. Minchinton. Using infrared imagery to test for quadrat-level temperature variation and effects on the early life history of a rocky-shore barnacle. Limnology and Oceanography, 57(5):1279–1291, September 2012. CODEN LIOCAH. ISSN 0024-3590.

Landry:2014:DBH

[Lan14] Michael R. Landry. On database biases and hypothesis testing with dilution experiments: Response to comment by latasa. Limnology and Oceanography, 64(3):1095–1096, May 2014. CO-DEN LIOCAH. ISSN 0024-3590.

Lonborg:2018:OMB

[LÁSDC18] Christian Lønborg, Xosé Antón Álvarez-Salgado, Samantha Duggan, and Cátia Carreira. Organic matter bioavailability in tropical coastal waters: The Great Barrier Reef. Limnology and Oceanography, 63(2):1015–1035, March 2018. CODEN LIOCAH. ISSN 0024-3590.

Latasa:2014:CPB

[Lat14] Mikel Latasa. Comment: a potential bias in the databases of phytoplankton growth and microzooplankton grazing rates because of the improper formulation of the null hypothesis in dilution experiments. Limnology and Oceanography, 64(3):1092– 1094, May 2014. CODEN LIOCAH. ISSN 0024-3590.

Lu:2018:EWL

[LBB18] Jing Lu, Stuart E. Bunn, and Michele A. Burford. Effects of water level fluctuations on nitrogen dynamics in littoral macrophytes. Limnology and Oceanography, 63(2):833–845, March 2018. CODEN LIOCAH. ISSN 0024-3590.

Leach:2018:PDD

[LBC^{+}18] Taylor H. Leach, Beatrix E. Beisner, Cayelan C. Carey, Patricia Pernica, Kevin C. Rose, Yannick Huot, Jennifer A. Brentrup, Isabelle Domaizon, Hans-Peter Grossart, Bastiaan W. Ibelings, Stéphan Jacquet, Patrick T. Kelly, James A. Rusak, Jason D. Stockwell, Dietmar Straile, and Piet Verburg. Patterns and drivers of deep chlorophyll maxima structure in 100 lakes: The relative importance of light and thermal stratification. Limnology and Oceanography, 63(2):628–646, March 2018. CODEN LIOCAH. ISSN 0024-3590.

Lelong:2013:ICL

[LBHS13] Aurélie Lelong, Eva Bucciarelli, Hélène Hégaret, and Philippe Soudant. Iron and copper limitations differently affect growth rates and photosynthetic and physiological parameters of the marine diatom Pseudo-nitzschia delicatissima. Limnology and Oceanography, 58(2):613–623, March 2013. CODEN LIOCAH. ISSN 0024-3590.

Lyons:2011:PCS

[LBNT11] Gabrielle Lyons, Claudia R. Benitez-Nelson, and Robert C. Thunell. Phosphorus composition of sinking particles in the Guaymas Basin, Gulf of California. Limnology and Oceanography, 57(4):1093–1105, May 2011. CODEN LIOCAH. ISSN 0024-3590.

Lutz:2012:DDD

[LBR^{+}12] Brian D. Lutz, Emily S. Bernhardt, Brian J. Roberts, Rose M. Cory, and Patrick J. Mulholland. Distinguishing dynamics of dissolved organic matter components in a forested stream using kinetic enrichments. Limnology and Oceanography, 57(1):76–89, January 2012. CODEN LIOCAH. ISSN 0024-3590.

Lobon:2013:RPT

[LBR^{+}13] Carla M. Lobón, Jean-Marie Bouquet, Magnus Reeve, Aliona Novac, José Luis Acuña, Eric M. Thompson, and Christofer Troedsson. Response of the pelagic tunicate appendicularian, Oikopleura dioica to controlled simulations of a strong bloom condition: a bottom-up perspective. Limnology and Oceanography, 58(2):215–226, January 2013. CODEN LIOCAH. ISSN 0024-3590.

Lewis:2017:LES

[LBS17] D. M. Lewis, A. Brereton, and J. T. Siddons. A large eddy simulation study of the formation of deep chlorophyll/biological maxima in un-stratified mixed layers: The roles of turbulent mixing and predation pressure. Limnology and Oceanography, 65(9):2277–2307, September 2017. CODEN LIOCAH. ISSN 0024-3590.

Lui:2011:SLN

[LC11] Hon-Kit Lui and Chen-Tung Arthur Chen. Shifts in limiting nutrients in an estuary caused by mixing and biological activity. Limnology and Oceanography, 57(4):989–998, May 2011. CO-DEN LIOCAH. ISSN 0024-3590.

Lottig:2012:IFL

[LC12] Noah R. Lottig and Stephen R. Carpenter. Interpolating and forecasting lake characteristics using long-term monitoring data. Limnology and Oceanography, 57(4):1113–1125, July 2012. CO-DEN LIOCAH. ISSN 0024-3590.

Lough:2016:EDG

[LCBC16] J. M. Lough, N. E. Cantin, J. A. Benthuysen, and T. F. Cooper. Environmental drivers of growth in massive Porites corals over 16 degrees of latitude along Australia’s northwest shelf. Limnology and Oceanography, 61(2):684–700, March 2016. CODEN LIOCAH. ISSN 0024-3590.

Loebl:2010:PBH

[LCCF10] Martina Loebl, Amanda M. Cockshutt, Douglas A. Campbell, and Zoe V. Finkel. Physiological basis for high resistance to photoinhibition under nitrogen depletion in Emiliania huxleyi. Limnology and Oceanography, 55(5):2150–2160, September 2010. CODEN LIOCAH. ISSN 0024-3590.

Liu:2014:ESG

[LCH^{+}14] Qian Liu, Matthew A. Charette, Paul B. Henderson, Daniel C. McCorkle, William Martin, and Minhan Dai. Effect of submarine groundwater discharge on the coastal ocean inorganic carbon cycle. Limnology and Oceanography, 59(5):1529–1554, September 2014. CODEN LIOCAH. ISSN 0024-3590.

Li:2012:CMO

[LCM^{+}12] Jiying Li, Sean A. Crowe, David Miklesh, Matthew Kistner, Donald E. Canfield, and Sergei Katsev. Carbon mineralization and oxygen dynamics in sediments with deep oxygen penetration, Lake Superior. Limnology and Oceanography, 57(6):1634– 1650, November 2012. CODEN LIOCAH. ISSN 0024-3590.

Latasa:2017:DPG

[LCM^{+}17] Mikel Latasa, Ana María Cabello, Xosé Anxelu G. Morán, Ramon Massana, and Renate Scharek. Distribution of phytoplankton groups within the deep chlorophyll maximum. Limnology and Oceanography, 62(2):665–685, March 2017. CODEN LI-OCAH. ISSN 0024-3590.

Layton:2019:CMW

[LCS^{+}19] Cayne Layton, Matthew J. Cameron, Victor Shelamoff, Pamela A. Fernández, Damon Britton, Catriona L. Hurd, Jeffrey T. Wright, and Craig R. Johnson. Chemical microenvironments within macroalgal assemblages: Implications for the inhibition of kelp recruitment by turf algae. Limnology and Oceanography, 64(4):1600–1613, July 2019. CODEN LIOCAH. ISSN 0024-3590.

Lee:2017:VDD

[LCW17a] Chih-Ping Lee, Chia-Yu Cheng, and Liang-Saw Wen. Vertical distributions and diurnal variations of high-molecular-weight dissolved arsenic in the oligotrophic ocean. Limnology and Oceanography, 65(9):2200–2212, September 2017. CODEN LI-OCAH. ISSN 0024-3590.

Liao:2017:SCA

[LCW^{+}17b] Jingqiu Liao, Xiaofeng Cao, Jie Wang, Lei Zhao, Jinhua Sun, Dalin Jiang, and Yi Huang. Similar community assembly mechanisms underlie similar biogeography of rare and abundant bacteria in lakes on Yungui Plateau, China. Limnology and Oceanography, 62(2):723–735, March 2017. CODEN LIOCAH. ISSN 0024-3590.

Liu:2019:MPS

[LCZ^{+}19] Kailin Liu, Bingzhang Chen, Shuwen Zhang, Mitsuhide Sato, Zhiyuan Shi, and Hongbin Liu. Marine phytoplankton in subtropical coastal waters showing lower thermal sensitivity than microzooplankton. Limnology and Oceanography, 64(3):1103– 1119, May 2019. CODEN LIOCAH. ISSN 0024-3590.

Lopez-Duarte:2011:BMD

[LDCT11] Paola C. López-Duarte, John H. Christy, and Richard A. Tankersley. A behavioral mechanism for dispersal in fiddler crab larvae (genus Uca) varies with adult habitat, not phylogeny. Limnology and Oceanography, 56(5):1879–1892, September 2011. CODEN LIOCAH. ISSN 0024-3590.

Lebrato:2013:JBS

[LdJMS^{+}13] Mario Lebrato, Pedro de Jesus Mendes, Deborah K. Steinberg, Joan E. Cartes, Bethan M. Jones, Laura M. Birsa, Roberto Benavides, and Andreas Oschlies. Jelly biomass sinking speed reveals a fast carbon export mechanism. Limnology and Oceanography, 58(3):1113–1122, May 2013. CODEN LIOCAH. ISSN 0024-3590.

Lupon:2019:GIC

[LDL^{+}19] Anna Lupon, Blaize A. Denfeld, Hjalmar Laudon, Jason Leach, Jan Karlsson, and Ryan A. Sponseller. Groundwater inflows control patterns and sources of greenhouse gas emissions from streams. Limnology and Oceanography, 64(4):1545–1557, July 2019. CODEN LIOCAH. ISSN 0024-3590.

LaNafie:2012:WHN

[LdlSB^{+}12] Yayu A. La Nafie, Carmen B. de los Santos, Fernando G. Brun, Marieke M. van Katwijk, and Tjeerd J. Bouma. Waves and high nutrient loads jointly decrease survival and separately affect morphological and biomechanical properties in the seagrass Zostera noltii. Limnology and Oceanography, 57(6):1664–1672, November 2012. CODEN LIOCAH. ISSN 0024-3590.

Lucas:2011:GRM

[LDT^{+}11] Andrew J. Lucas, Christopher L. Dupont, Vera Tai, John L. Largier, Brian Palenik, and Peter J. S. Franks. The green ribbon: Multiscale physical control of phytoplankton productivity and community structure over a narrow continental shelf. Limnology and Oceanography, 56(2):611–626, March 2011. CODEN LIOCAH. ISSN 0024-3590.

Labry:2016:HAP

[LDY^{+}16] Claire Labry, Daniel Delmas, Agnes Youenou, Julien Quere, Aude Leynaert, Stephane Fraisse, Mélanie Raimonet, and Olivier Ragueneau. High alkaline phosphatase activity in phosphate replete waters: The case of two macrotidal estuaries. Limnology and Oceanography, 61(4):1513–1529, July 2016. CODEN LIOCAH. ISSN 0024-3590.

Lee:2018:HCC

[Lee18] Dongkyu Lee. High concentration chlorophyll-a rings associated with the formation of intrathermocline eddies. Limnology and Oceanography, 63(6):2806–2814, November 2018. CODEN LI-OCAH. ISSN 0024-3590.

Lombard:2010:EME

[LEG^{+}10] Fabien Lombard, Damien Eloire, Angelique Gobet, Lars Stemmann, John R. Dolan, Antoine Sciandra, and Gabriel Gorsky. Experimental and modeling evidence of appendicularian-ciliate interactions. Limnology and Oceanography, 55(1):77–90, January 2010. CODEN LIOCAH. ISSN 0024-3590.

Larson:2018:ABC

[LEK^{+}18] James H. Larson, Mary Anne Evans, Robert J. Kennedy, Sean W. Bailey, Keith A. Loftin, Zachary R. Laughrey, Robin A. Femmer, Jeff S. Schaeffer, William B. Richardson, Timothy T. Wynne, J. C. Nelson, and Joseph W. Duris. Associations between cyanobacteria and indices of secondary production in the western basin of Lake Erie. Limnology and Oceanography, 63 (S1):S232–S243, March 2018. CODEN LIOCAH. ISSN 0024-

Larsen:2015:CRN

[LEN^{+}15] Aud Larsen, Jorun K. Egge, Jens C. Nejstgaard, Iole Di Capua, Runar Thyrhaug, Gunnar Bratbak, and T. Frede Thingstad. Contrasting response to nutrient manipulation in Arctic mesocosms are reproduced by a minimum microbial food web model. Limnology and Oceanography, 60(2):360–374, March 2015. CO-DEN LIOCAH. ISSN 0024-3590.

Lesser:2016:CCS

[Les16] Michael P. Lesser. Climate change stressors cause metabolic depression in the blue mussel, Mytilus edulis, from the Gulf of Maine. Limnology and Oceanography, 62(3):1705–1717, September 2016. CODEN LIOCAH. ISSN 0024-3590.

Lesser:2019:PSL

[Les19] Michael P. Lesser. Phylogenetic signature of light and thermal stress for the endosymbiotic dinoflagellates of corals (family Symbiodiniaceae). Limnology and Oceanography, 64(5):1852– 1863, September 2019. CODEN LIOCAH. ISSN 0024-3590.

Lee:2016:MUD

[LF16] Cheng-Shiuan Lee and Nicholas S. Fisher. Methylmercury uptake by diverse marine phytoplankton. Limnology and Oceanography, 61(5):1626–1639, September 2016. CODEN LIOCAH. ISSN 0024-3590.

Lee:2017:III

[LF17a] Cheng-Shiuan Lee and Nicholas S. Fisher. Issue information — instr to contrib. Limnology and Oceanography, 61(5):1325–1327, May 2017. CODEN LIOCAH. ISSN 0024-3590.

Lee:2017:MUD

[LF17b] Cheng-Shiuan Lee and Nicholas S. Fisher. Methylmercury uptake by diverse marine phytoplankton. Limnology and Oceanography, 66(5):1324, May 2017. CODEN LIOCAH. ISSN 0024-3590.

Lee:2019:MGE

[LF19] Cheng-Shiuan Lee and Nicholas S. Fisher. Microbial generation of elemental mercury from dissolved methylmercury in seawater. Limnology and Oceanography, 64(2):679–693, March 2019. CODEN LIOCAH. ISSN 0024-3590.

Lichtschlag:2010:GPC

[LFB^{+}10] Anna Lichtschlag, Janine Felden, Volker Brüchert, Antje Boetius, and Dirk de Beer. Geochemical processes and chemosynthetic primary production in different thiotrophic mats of the Håkon Mosby Mud Volcano (Barents Sea). Limnology and Oceanography, 55(4):931–949, March 2010. CODEN LIOCAH. ISSN 0024-3590.

Liu:2017:CFP

[LFC17] Fengjie Liu, Claude Fortin, and Peter G. C. Campbell. Can freshwater phytoplankton access cadmium bound to lowmolecular-weight thiols? Limnology and Oceanography, 62(6): 2604–2615, November 2017. CODEN LIOCAH. ISSN 0024-3590.

Lindstrom:2010:IBB

[LFGK10] Eva S. Lindström, Xin Mei Feng, Wilhelm Granéli, and Emma S. Kritzberg. The interplay between bacterial community composition and the environment determining function of inland water bacteria. Limnology and Oceanography, 55(5):2052–2060, September 2010. CODEN LIOCAH. ISSN 0024-3590.

Lebret:2012:GRA

[LFH^{+}12] Karen Lebret, María Fernández Fernández, Camilla H. C. Hagman, Karin Rengefors, and Lars-Anders Hansson. Grazing resistance allows bloom formation and may explain invasion success of Gonyostomum semen. Limnology and Oceanography, 57(3): 727–734, May 2012. CODEN LIOCAH. ISSN 0024-3590.

Li:2017:RGG

[LFL17] Qian P. Li, Peter J. S. Franks, and Michael R. Landry. Recovering growth and grazing rates from nonlinear dilution experiments. Limnology and Oceanography, 62(5):1825–1835, September 2017. CODEN LIOCAH. ISSN 0024-3590.

Lundgren:2010:GID

[LG10] Veronica Lundgren and Edna Granéli. Grazer-induced defence in Phaeocystis globosa (Prymnesiophyceae): Influence of different nutrient conditions. Limnology and Oceanography, 55(5): 1965–1976, September 2010. CODEN LIOCAH. ISSN 0024-

Lin:2016:DCA

[LG16] Peng Lin and Laodong Guo. Dynamic changes in the abundance and chemical speciation of dissolved and particulate phosphorus across the river–lake interface in southwest Lake Michigan. Limnology and Oceanography, 61(2):771–789, March 2016. CODEN LIOCAH. ISSN 0024-3590.

Lunden:2013:ASS

[LGC13a] Jay J. Lunden, Samuel E. Georgian, and Erik E. Cordes. Aragonite saturation states at cold-water coral reefs structured by Lophelia pertusa in the northern Gulf of Mexico. Limnology and Oceanography, 58(2):354–362, January 2013. CODEN LI-OCAH. ISSN 0024-3590.

Lunden:2013:EAS

[LGC13b] Jay J. Lunden, Samuel E. Georgian, and Erik E. Cordes. Erratum: Aragonite saturation states at cold-water coral reefs structured by Lophelia pertusa in the Northern Gulf of Mexico. Limnology and Oceanography, 58(3):1147, May 2013. CODEN LI-OCAH. ISSN 0024-3590.

Lorenson:2016:DMB

[LGC16] Thomas D. Lorenson, Jens Greinert, and Richard B. Coffin. Dissolved methane in the Beaufort Sea and the Arctic Ocean, 1992– 2009; sources and atmospheric flux. Limnology and Oceanography, 61(S1):S300–S323, November 2016. CODEN LIOCAH. ISSN 0024-3590.

Lancelot:2012:RPC

[LGR^{+}12] Christiane Lancelot, Philippe Grosjean, Véronique Rousseau, Elsa Breton, and Patricia M. Glibert. Rejoinder to “Perils of correlating CUSUM-transformed variables to infer ecological relationships (Breton et al. 2006; Glibert 2010)”. Limnology and Oceanography, 57(3):669–670, March 2012. CODEN LIOCAH. ISSN 0024-3590. See [CJC^{+}12].

Lau:2013:CED

[LGV13] Danny C. P. Lau, Willem Goedkoop, and Tobias Vrede. Crossecosystem differences in lipid composition and growth limitation of a benthic generalist consumer. Limnology and Oceanography, 58(4):1149–1164, July 2013. CODEN LIOCAH. ISSN 0024-

Lyu:2019:TAD

[LGW^{+}19] Kai Lyu, Lei Gu, Hui Wang, Xuexia Zhu, Lu Zhang, Yunfei Sun, Yuan Huang, and Zhou Yang. Transcriptomic analysis dissects the mechanistic insight into the Daphnia clonal variation in tolerance to toxic Microcystis. Limnology and Oceanography, 64(1): 272–283, January 2019. CODEN LIOCAH. ISSN 0024-3590.

Larsen:2017:DCC

[LH17] L. G. Larsen and J. W. Harvey. Disrupted carbon cycling in restored and unrestored urban streams: Critical timescales and controls. Limnology and Oceanography, 62(S1):S160–S182, November 2017. CODEN LIOCAH. ISSN 0024-3590.

Lisi:2019:EDD

[LH19] Peter J. Lisi and Catherine L. Hein. Eutrophication drives divergent water clarity responses to decadal variation in lake level. Limnology and Oceanography, 64(S1):S49–S59, January 2019. CODEN LIOCAH. ISSN 0024-3590.

Lignell:2013:GRP

[LHLT13] Risto Lignell, Heikki Haario, Marko Laine, and T. Frede Thingstad. Getting the “right” parameter values for models of the pelagic microbial food web. Limnology and Oceanography, 58(2):301–313, January 2013. CODEN LIOCAH. ISSN 0024-3590.

Lopez:2019:RBP

[LHS19] Lianna S. Lopez, Bailey A. Hewitt, and Sapna Sharma. Reaching a breaking point: How is climate change influencing the timing of ice breakup in lakes across the Northern Hemisphere? Limnology and Oceanography, 64(6):2621–2631, November 2019. CODEN LIOCAH. ISSN 0024-3590.

Li:2018:CDZ

[LHSBP18] Shanze Li, Charles S. Hopkinson, Joseph P. Schubauer-Berigan, and Steven C. Pennings. Climate drivers of Zizaniopsis miliacea biomass in a Georgia, U.S.A. tidal fresh marsh. Limnology and Oceanography, 64(4):2266–2276, September 2018. CODEN LIOCAH. ISSN 0024-3590.

Larsen:2015:FBS

[LHSG15] Laurel Larsen, Jud Harvey, Katherine Skalak, and Marissa Goodman. Fluorescence-based source tracking of organic sediment in restored and unrestored urban streams. Limnology and Oceanography, 60(4):1439–1461, July 2015. CODEN LIOCAH. ISSN 0024-3590.

Lind:2018:IED

[LJ18] Patrick R. Lind and Punidan D. Jeyasingh. Interactive effects of dietary phosphorus and iron on Daphnia life history. Limnology and Oceanography, 63(3):1181–1190, May 2018. CODEN LIOCAH. ISSN 0024-3590.

Liew:2018:BIT

[LJL^{+}18] J. H. Liew, T. D. Jardine, R. B. H. Lim, J. T. B. Kwik, H. H. Tan, Z. Y. Kho, and D. C. J. Yeo. Bottom-up influences on tropical freshwater food web structure support the “environmental filtering” hypothesis. Limnology and Oceanography, 63(5):1877– 1890, September 2018. CODEN LIOCAH. ISSN 0024-3590.

Li:2014:NCD

[LK14] Jiying Li and Sergei Katsev. Nitrogen cycling in deeply oxygenated sediments: Results in Lake Superior and implications for marine sediments. Limnology and Oceanography, 59(2):465– 481, March 2014. CODEN LIOCAH. ISSN 0024-3590.

Lee:2015:DPC

[LK15] Junghyun Lee and Guebuem Kim. Dependence of pH in coastal waters on the adsorption of protons onto sediment minerals. Limnology and Oceanography, 60(3):831–839, May 2015. CO-DEN LIOCAH. ISSN 0024-3590.

Leys:2018:PMS

[LKF^{+}18] S. P. Leys, A. S. Kahn, J. K. H. Fang, T. Kutti, and R. J. Bannister. Phagocytosis of microbial symbionts balances the carbon and nitrogen budget for the deep-water boreal sponge Geodia barretti. Limnology and Oceanography, 63(1):187–202, January 2018. CODEN LIOCAH. ISSN 0024-3590.

Lombard:2013:CUC

[LKK13] F. Lombard, M. Koski, and T. Kiørboe. Copepods use chemical trails to find sinking marine snow aggregates. Limnology and Oceanography, 58(2):185–192, January 2013. CODEN LIOCAH. ISSN 0024-3590.

Lee:2010:RBS

[LKLH10] Yong-Woo Lee, Guebuem Kim, Weol-Ae Lim, and Dong-Woon Hwang. A relationship between submarine groundwater borne nutrients traced by Ra isotopes and the intensity of dinoflagellate red-tides occurring in the southern sea of Korea. Limnology and Oceanography, 55(1):1–10, January 2010. CODEN LIOCAH. ISSN 0024-3590.

Lecher:2016:MTT

[LKS^{+}16] Alanna L. Lecher, John Kessler, Katy Sparrow, Fenix Garcia-Tigreros Kodovska, Natasha Dimova, Joseph Murray, Slawek Tulaczyk, and Adina Paytan. Methane transport through submarine groundwater discharge to the North Pacific and Arctic Ocean at two Alaskan sites. Limnology and Oceanography, 61 (S1):S344–S355, November 2016. CODEN LIOCAH. ISSN 0024-3590.

Laney:2017:EZU

[LKT17] Samuel R. Laney, Richard A. Krishfield, and John. M. Toole. The euphotic zone under Arctic Ocean sea ice: Vertical extents and seasonal trends. Limnology and Oceanography, 62(5):1910– 1934, September 2017. CODEN LIOCAH. ISSN 0024-3590.

Larsson:2011:EEL

[LL11] Petter Larsson and Winfried Lampert. Experimental evidence of a low-oxygen refuge for large zooplankton. Limnology and Oceanography, 56(5):1682–1688, September 2011. CODEN LI-OCAH. ISSN 0024-3590.

Lacour:2017:GCC

[LLB17] Thomas Lacour, Jade Larivière, and Marcel Babin. Growth, chla content, photosynthesis, and elemental composition in polar and temperate microalgae. Limnology and Oceanography, 62(1): 43–58, January 2017. CODEN LIOCAH. ISSN 0024-3590.

Le:2015:RBI

[LLH^{+}15] Chengfeng Le, John C. Lehrter, Chuanmin Hu, Blake Schaeffer, Hugh MacIntyre, James D. Hagy, and David L. Beddick. Relation between inherent optical properties and land use and land cover across Gulf Coast estuaries. Limnology and Oceanography, 60(3):920–933, May 2015. CODEN LIOCAH. ISSN 0024-3590.

Lampert:2010:COS

[LLL10] Winfried Lampert, Kathrin P. Lampert, and Petter Larssona. Coexisting overwintering strategies in Daphnia pulex: a test of genetic differences and growth responses. Limnology and Oceanography, 55(5):1893–1900, September 2010. CODEN LI-OCAH. ISSN 0024-3590.

Liu:2018:ISH

[LLW^{+}18] Feng Liu, Xingfeng Liu, Yu Wang, Zhe Jin, Fiona Wanjiku Moejes, and Song Sun. Insights on the Sargassum horneri golden tides in the Yellow Sea inferred from morphological and molecular data. Limnology and Oceanography, 63(4):1762–1773, July 2018. CODEN LIOCAH. ISSN 0024-3590.

Lin:2012:BPS

[LM12] Chia Yu Lin and Steven L. Manley. Bromoform production from seawater treated with bromoperoxidase. Limnology and Oceanography, 57(6):1857–1866, November 2012. CODEN LI-OCAH. ISSN 0024-3590.

Lyngsgaard:2014:CVD

[LMR14] Maren Moltke Lyngsgaard, Stiig Markager, and Katherine Richardson. Changes in the vertical distribution of primary production in response to land-based nitrogen loading. Limnology and Oceanography, 59(5):1679–1690, September 2014. CODEN LIOCAH. ISSN 0024-3590.

Luhar:2011:FIR

[LN11] Mitul Luhar and Heidi M. Nepf. Flow-induced reconfiguration of buoyant and flexible aquatic vegetation. Limnology and Oceanography, 56(6):2003–2017, November 2011. CODEN LI-OCAH. ISSN 0024-3590.

Laney:2012:DFT

[LOS12] Samuel R. Laney, Robert J. Olson, and Heidi M. Sosik. Diatoms favor their younger daughters. Limnology and Oceanography, 58 (1):1572–1578, September 2012. CODEN LIOCAH. ISSN 0024-3590.

Lorke:2010:SMT

[LP10] Andreas Lorke and W. Nikolaus Probst. In situ measurements of turbulence in fish shoals. Limnology and Oceanography, 55(1): 354–364, January 2010. CODEN LIOCAH. ISSN 0024-3590.

Leech:2018:FBL

[LPLH18] Dina M. Leech, Amina I. Pollard, Stephanie G. Labou, and Stephanie E. Hampton. Fewer blue lakes and more murky lakes across the continental U.S.: Implications for planktonic food webs. Limnology and Oceanography, 63(6):2661–2680, November 2018. CODEN LIOCAH. ISSN 0024-3590.

Lebrato:2011:DAO

[LPO^{+}11] Mario Lebrato, Markus Pahlow, Andreas Oschlies, Kylie A. Pitt, Daniel O. B. Jones, Juan Carlos Molinero, and Robert H. Condon. Depth attenuation of organic matter export associated with jelly falls. Limnology and Oceanography, 56(5):1917–1928, September 2011. CODEN LIOCAH. ISSN 0024-3590.

Lowe:2016:IMD

[LRG16] Alexander T. Lowe, Emily A. Roberts, and Aaron W. E. Galloway. Improved marine-derived POM availability and increased pH related to freshwater influence in an inland sea. Limnology and Oceanography, 61(6):2122–2138, November 2016. CODEN LIOCAH. ISSN 0024-3590.

Lammers:2017:SVP

[LRM17] J. M. Lammers, G. J. Reichart, and J. J. Middelburg. Seasonal variability in phytoplankton stable carbon isotope ratios and bacterial carbon sources in a shallow Dutch lake. Limnology and Oceanography, 62(6):2773–2787, November 2017. CODEN LIOCAH. ISSN 0024-3590.

Long:2019:COM

[LRM^{+}19] Matthew H. Long, Jennie E. Rheuban, Daniel C. McCorkle, David J. Burdige, and Richard C. Zimmerman. Closing the oxygen mass balance in shallow coastal ecosystems. Limnology and Oceanography, 64(6):2694–2708, November 2019. CODEN LIOCAH. ISSN 0024-3590.

Llopiz:2010:DDD

[LRS^{+}10] Joel K. Llopiz, David E. Richardson, Akihiro Shiroza, Sharon L. Smith, and Robert K. Cowen. Distinctions in the diets and distributions of larval tunas and the important role of appendicularians. Limnology and Oceanography, 55(3):983–996, May 2010. CODEN LIOCAH. ISSN 0024-3590.

Lasley-Rasher:2012:PRS

[LRY12] Rachel S. Lasley-Rasher and Jeannette Yen. Predation risk suppresses mating success and offspring production in the coastal marine copepod, Eurytemora herdmani. Limnology and Oceanography, 57(2):433–440, March 2012. CODEN LIOCAH. ISSN 0024-3590.

Lunt:2014:TIT

[LS14] Jessica Lunt and Delbert L. Smee. Turbidity influences trophic interactions in estuaries. Limnology and Oceanography, 59(6): 2002–2012, November 2014. CODEN LIOCAH. ISSN 0024-3590.

Lisi:2015:WDU

[LS15] Peter J. Lisi and Daniel E. Schindler. Wind-driven upwelling in lakes destabilizes thermal regimes of downstream rivers. Limnology and Oceanography, 60(1):169–180, January 2015. CODEN LIOCAH. ISSN 0024-3590.

Lubarsky:2018:ESG

[LSD18] Katie A. Lubarsky, Nyssa J. Silbiger, and Megan J. Donahue. Effects of submarine groundwater discharge on coral accretion and bioerosion on two shallow reef flats. Limnology and Oceanography, 63(4):1660–1676, July 2018. CODEN LIOCAH. ISSN 0024-3590.

Lee:2018:RLS

[LSDW18] Zhongping Lee, Shaoling Shang, Keping Du, and Jianwei Wei. Resolving the long-standing puzzles about the observed Secchi depth relationships. Limnology and Oceanography, 63(6):2321– 2336, November 2018. CODEN LIOCAH. ISSN 0024-3590.

Lee:2017:PCC

[LSH^{+}17] Eunhee Lee, Doyun Shin, Sung Pil Hyun, Kyung-Seok Ko, Hee Sun Moon, Dong-Chan Koh, Kyoochul Ha, and Byung-Yong Kim. Periodic change in coastal microbial community structure associated with submarine groundwater discharge and tidal fluctuation. Limnology and Oceanography, 62(2):437–451, March 2017. CODEN LIOCAH. ISSN 0024-3590.

Lottig:2011:CRS

[LSHK11] Noah R. Lottig, Emily H. Stanley, Paul C. Hanson, and Timothy K. Kratz. Comparison of regional stream and lake chemistry: Differences, similarities, and potential drivers. Limnology and Oceanography, 56(5):1551–1562, September 2011. CODEN LIOCAH. ISSN 0024-3590.

Lombard:2011:APR

[LSK11] F. Lombard, E. Selander, and T. Kiørboe. Active prey rejection in the filter-feeding appendicularian Oikopleura dioica. Limnology and Oceanography, 56(4):1504–1512, July 2011. CODEN LIOCAH. ISSN 0024-3590.

Lansdown:2012:CKP

[LTH^{+}12] K. Lansdown, M. Trimmer, C. M. Heppell, F. Sgouridis, S. Ullah, A. L. Heathwaite, A. Binley, and H. Zhang. Characterization of the key pathways of dissimilatory nitrate reduction and their response to complex organic substrates in hyporheic sediments. Limnology and Oceanography, 57(2):387–400, March 2012. CODEN LIOCAH. ISSN 0024-3590.

Lehtinen:2017:PSR

[LTPA17] Sirpa Lehtinen, Timo Tamminen, Robert Ptacnik, and Tom Andersen. Phytoplankton species richness, evenness, and production in relation to nutrient availability and imbalance. Limnology and Oceanography, 62(4):1393–1408, July 2017. CODEN LIOCAH. ISSN 0024-3590.

Lawrence:2018:VMA

[LTPK^{+}18] Janice Lawrence, Joachim Töpper, Elżbieta Petelenz-Kurdziel, Gunnar Bratbak, Aud Larsen, Eric Thompson, Christofer Troedsson, and Jessica Louise Ray. Viruses on the menu: The appendicularian Oikopleura dioica efficiently removes viruses from seawater. Limnology and Oceanography, 63(S1):S244–S253, March 2018. CODEN LIOCAH. ISSN 0024-3590.

Liu:2017:IRV

[LTX^{+}17] Hongbin Liu, Shangjin Tan, Jie Xu, Wang Guo, Xiaomin Xia, and Shun Yan Cheung. Interactive regulations by viruses and dissolved organic matter on the bacterial community. Limnology and Oceanography, 62(S1):S364–S380, November 2017. CODEN LIOCAH. ISSN 0024-3590.

Lopez-Urrutia:2015:TAS

[LUM15] Ángel López-Urrutia and Xosé Anxelu G. Morán. Temperature affects the size-structure of phytoplankton communities in the ocean. Limnology and Oceanography, 60(3):733–738, May 2015. CODEN LIOCAH. ISSN 0024-3590.

Lyons:2016:LZR

[LV16] Devin A. Lyons and Rolf D. Vinebrooke. Linking zooplankton richness with energy input and insularity along altitudinal and latitudinal gradients. Limnology and Oceanography, 61(3):841– 852, May 2016. CODEN LIOCAH. ISSN 0024-3590.

Langenegger:2019:WBK

[LVDM19] T. Langenegger, D. Vachon, D. Donis, and D. F. McGinnis. What the bubble knows: Lake methane dynamics revealed by sediment gas bubble composition. Limnology and Oceanography, 64(4):1526–1544, July 2019. CODEN LIOCAH. ISSN 0024-3590.

Laurion:2010:VGG

[LVM^{+}10] Isabelle Laurion, Warwick F. Vincent, Sally MacIntyre, Leira Retamal, Christiane Dupont, Pierre Francus, and Reinhard Pienitz. Variability in greenhouse gas emissions from permafrost thaw ponds. Limnology and Oceanography, 55(1):115–133, January 2010. CODEN LIOCAH. ISSN 0024-3590.

Letelier:2017:LAP

[LWB^{+}17] Ricardo M. Letelier, Angelicque E. White, Robert R. Bidigare, Benedetto Barone, Matthew J. Church, and David M. Karl. Light absorption by phytoplankton in the North Pacific Subtropical Gyre. Limnology and Oceanography, 62(4):1526–1540, July 2017. CODEN LIOCAH. ISSN 0024-3590.

Law:2011:RSN

[LWE^{+}11] C. S. Law, E. M. S. Woodward, M. J. Ellwood, A. Marriner, S. J. Bury, and K. A. Safi. Response of surface nutrient inventories and nitrogen fixation to a tropical cyclone in the southwest Pacific. Limnology and Oceanography, 56(4):1372–1385, July 2011. CODEN LIOCAH. ISSN 0024-3590.

Langer:2019:TDP

[LWE^{+}19] Sina M. Langer, Linda C. Weiss, Mikael T. Ekvall, Giuseppe Bianco, Lars-Anders Hansson, and Ralph Tollrian. A threedimensional perspective of Daphnia’s swimming behavior with and without predator cues. Limnology and Oceanography, 64 (4):1515–1525, July 2019. CODEN LIOCAH. ISSN 0024-3590.

Loick-Wilde:2012:INF

[LWrDM^{+}12] Natalie Loick-Wilde, Jö rg Dutz, Anja Miltner, Matthias Gehre, Joseph P. Montoya, and Maren Voss. Incorporation of nitrogen from N_{2} fixation into amino acids of zooplankton. Limnology and Oceanography, 57(1):199–210, January 2012. CODEN LIOCAH. ISSN 0024-3590.

Leary:2017:ITP

[LWS^{+}17] Paul R. Leary, C. Brock Woodson, Michael E. Squibb, Mark W. Denny, Stephen G. Monismith, and Fiorenza Micheli. “Internal tide pools” prolong kelp forest hypoxic events. Limnology and Oceanography, 62(6):2864–2878, November 2017. CODEN LIOCAH. ISSN 0024-3590.

Loick-Wilde:2016:NSN

[LWWC^{+}16] Natalie Loick-Wilde, Sarah C. Weber, Brandon J. Conroy, Douglas G. Capone, Victoria J. Coles, Patricia M. Medeiros, Deborah K. Steinberg, and Joseph P. Montoya. Nitrogen sources and net growth efficiency of zooplankton in three Amazon River plume food webs. Limnology and Oceanography, 61(2):460–481, March 2016. CODEN LIOCAH. ISSN 0024-3590.

Loick-Wilde:2018:NAA

[LWWE^{+}18] Natalie Loick-Wilde, Sarah C. Weber, Elvita Eglite, Iris Liskow, Detlef Schulz-Bull, Norbert Wasmund, Dirk Wodarg, and Joseph P. Montoya. De novo amino acid synthesis and turnover during N_{2} fixation. Limnology and Oceanography, 63(3):1076– 1092, May 2018. CODEN LIOCAH. ISSN 0024-3590.

Liao:2017:TMC

[LYH17] Wen-Hsuan Liao, Shun-Chung Yang, and Tung-Yuan Ho. Trace metal composition of size-fractionated plankton in the Western Philippine Sea: The impact of anthropogenic aerosol deposition. Limnology and Oceanography, 65(9):2243–2259, September 2017. CODEN LIOCAH. ISSN 0024-3590.

Liu:2017:SEE

[LYL^{+}17] Jin Liu, Siyuan Ye, Edward Allen Laws, Chunting Xue, Hongming Yuan, Xigui Ding, Guangming Zhao, Shixiong Yang, Lei He, Jin Wang, Shaofeng Pei, Yongbiao Wang, and Qingyuan Lu. Sedimentary environment evolution and biogenic silica records over 33,000 years in the Liaohe delta, China. Limnology and Oceanography, 62(2):474–489, March 2017. CODEN LIOCAH. ISSN 0024-3590.

Liu:2014:RDI

[LZC^{+}14] Zhiyuan Liu, Longjun Zhang, Wei-Jun Cai, Liang Wang, Ming Xue, and Xiangshang Zhang. Removal of dissolved inorganic carbon in the Yellow River Estuary. Limnology and Oceanography, 59(2):413–426, March 2014. CODEN LIOCAH. ISSN 0024-3590.

Li:2018:PRD

[LZK18] Jiying Li, Yishu Zhang, and Sergei Katsev. Phosphorus recycling in deeply oxygenated sediments in Lake Superior controlled by organic matter mineralization. Limnology and Oceanography, 63(3):1372–1385, May 2018. CODEN LIOCAH. ISSN 0024-3590.

Li:2017:CPD

[LZR^{+}17] Huabing Li, Jin Zeng, Lijuan Ren, Jianjun Wang, Peng Xing, and Qinglong L. Wu. Contrasting patterns of diversity of abundant and rare bacterioplankton in freshwater lakes along an elevation gradient. Limnology and Oceanography, 62(4):1570–1585, July 2017. CODEN LIOCAH. ISSN 0024-3590.

McCabe:2018:MLR

[MA18] Andrew J. McCabe and William A. Arnold. Multiple linear regression models to predict the formation efficiency of triplet excited states of dissolved organic matter in temperate wetlands. Limnology and Oceanography, 64(4):1992–2014, September 2018. CODEN LIOCAH. ISSN 0024-3590.

Matich:2017:ENP

[MAB^{+}17] Philip Matich, Jerald S. Ault, Ross E. Boucek, David R. Bryan, Kirk R. Gastrich, Christine L. Harvey, Michael R. Heithaus, Jeremy J. Kiszka, Valeria Paz, Jennifer S. Rehage, and Adam E. Rosenblatt. Ecological niche partitioning within a large predator guild in a nutrient-limited estuary. Limnology and Oceanography, 63(3):934–953, May 2017. CODEN LIOCAH. ISSN 0024-3590.

Martin:2010:SAC

[MAC^{+}10] Patrick Martin, John T. Allen, Matthew J. Cooper, David G. Johns, Richard S. Lampitt, Richard Sanders, and Damon A. H. Teagle. Sedimentation of acantharian cysts in the Iceland Basin: Strontium as a ballast for deep ocean particle flux, and implications for acantharian reproductive strategies. Limnology and Oceanography, 55(2):604–614, March 2010. CODEN LIOCAH. ISSN 0024-3590.

Merbt:2011:BRW

[MACM11] Stephanie N. Merbt, Jean-Christophe Auguet, Emilio O. Casamayor, and Eugènia Marti. Biofilm recovery in a wastewater treatment plant-influenced stream and spatial segregation of ammonia-oxidizing microbial populations. Limnology and Oceanography, 57(4):1054–1064, May 2011. CODEN LIOCAH. ISSN 0024-3590.

Mollema:2015:IFT

[MAD^{+}15] Pauline N. Mollema, Marco Antonellini, Enrico Dinelli, Nicolas Greggio, and Pieter J. Stuyfzand. The influence of flowthrough saline gravel pit lakes on the hydrologic budget and hydrochemistry of a Mediterranean drainage basin. Limnology and Oceanography, 60(6):2009–2025, November 2015. CODEN LIOCAH. ISSN 0024-3590.

McKinney:2019:WDF

[MAF19] Paul McKinney, Jay Austin, and Gills Fai. The wind-driven formation of cross-shelf sediment plumes in a large lake. Limnology and Oceanography, 64(3):1309–1322, May 2019. CODEN LIOCAH. ISSN 0024-3590.

Meunier:2018:CPH

[MAFCD^{+}18] Cédric L. Meunier, Santiago Alvarez-Fernandez, Alessandra Ö. Cunha-Dupont, Carla Geisen, Arne M. Malzahn, Maarten Boersma, and Karen H. Wiltshire. The craving for phosphorus in heterotrophic dinoflagellates and its potential implications for biogeochemical cycles. Limnology and Oceanography, 63(4): 1774–1784, July 2018. CODEN LIOCAH. ISSN 0024-3590.

Manzello:2010:OAH

[Man10] Derek P. Manzello. Ocean acidification hotspots: Spatiotemporal dynamics of the seawater CO_{2} system of eastern Pacific coral reefs. Limnology and Oceanography, 55(1):239–248, January 2010. CODEN LIOCAH. ISSN 0024-3590.

Murphy:2016:MNP

[MAS^{+}16] Anna E. Murphy, Iris C. Anderson, Ashley R. Smyth, Bongkeun Song, and Mark W. Luckenbach. Microbial nitrogen processing in hard clam (Mercenaria mercenaria) aquaculture sediments: the relative importance of denitrification and dissimilatory nitrate reduction to ammonium (DNRA). Limnology and Oceanography, 61(5):1589–1604, September 2016. CODEN LI-OCAH. ISSN 0024-3590.

Mariani:2013:CPS

[MAV^{+}13] Patrizio Mariani, Ken H. Andersen, André W. Visser, Andrew D. Barton, and Thomas Kiørboe. Control of plankton seasonal succession by adaptive grazing. Limnology and Oceanography, 58(2):173–184, January 2013. CODEN LIOCAH. ISSN 0024-3590.

McDonnell:2010:VAS

[MB10] Andrew M. P. McDonnell and Ken O. Buesseler. Variability in the average sinking velocity of marine particles. Limnology and Oceanography, 55(5):2085–2096, September 2010. CODEN LIOCAH. ISSN 0024-3590.

Medeiros:2017:ETD

[MBAS^{+}17] Patricia M. Medeiros, Lydia Babcock-Adams, Michael Seidel, Renato M. Castelao, Daniela Di Iorio, James T. Hollibaugh, and Thorsten Dittmar. Export of terrigenous dissolved organic matter in a broad continental shelf. Limnology and Oceanography, 62(4):1718–1731, July 2017. CODEN LIOCAH. ISSN 0024-3590.

Magri:2018:BPI

[MBB^{+}18] M. Magri, S. Benelli, C. Bondavalli, M. Bartoli, R. R. Christian, and A. Bodini. Benthic N pathways in illuminated and bioturbated sediments studied with network analysis. Limnology and Oceanography, 63(S1):S68–S84, March 2018. CODEN LIOCAH. ISSN 0024-3590.

Mulholland:2012:RDF

[MBBG^{+}12] M. R. Mulholland, P. W. Bernhardt, J. L. Blanco-Garcia, A. Mannino, K. Hyde, E. Mondragon, K. Turk, P. H. Moisander, and J. P. Zehr. Rates of dinitrogen fixation and the abundance of diazotrophs in North American coastal waters between Cape Hatteras and Georges Bank. Limnology and Oceanography, 57 (4):1067–1083, July 2012. CODEN LIOCAH. ISSN 0024-3590.

Mislan:2011:SVE

[MBBW11] K. A. S. Mislan, Carol A. Blanchette, Bernardo R. Broitman, and Libe Washburn. Spatial variability of emergence, splash, surge, and submergence in wave-exposed rocky-shore ecosystems. Limnology and Oceanography, 56(3):857–866, May 2011. CODEN LIOCAH. ISSN 0024-3590.

Maranon:2016:CCI

[MBC^{+}16] Emilio Marañón, William M. Balch, Pedro Cermeño, Natalia González, Cristina Sobrino, Ana Fernández, María Huete-Ortega, Daffne C. López-Sandoval, Maximino Delgado, Marta Estrada, Marta Álvarez, Elisa Fernández-Guallart, and Carles Pelejero. Coccolithophore calcification is independent of carbonate chemistry in the tropical ocean. Limnology and Oceanography, 61(4):1345–1357, July 2016. CODEN LIOCAH. ISSN 0024-3590.

Mellett:2018:BCI

[MBC^{+}18] Travis Mellett, Matthew T. Brown, P. Dreux Chappell, Carolyn Duckham, Jessica N. Fitzsimmons, Claire P. Till, Robert M. Sherrell, Maria T. Maldonado, and Kristen N. Buck. The biogeochemical cycling of iron, copper, nickel, cadmium, manganese, cobalt, lead, and scandium in a California Current experimental study. Limnology and Oceanography, 63(S1):S425–S447, March 2018. CODEN LIOCAH. ISSN 0024-3590.

Modenutti:2013:EVE

[MBE^{+}13] Beatriz E. Modenutti, Esteban G. Balseiro, James J. Elser, Marcela Bastidas Navarro, Florencia Cuassolo, Cecilia Laspoumaderes, Maria S. Souza, and Verónica Díaz Villanueva. Effect of volcanic eruption on nutrients, light, and phytoplankton in oligotrophic lakes. Limnology and Oceanography, 58(4):1165–1175, July 2013. CODEN LIOCAH. ISSN 0024-3590.

McAllister:2015:DHB

[MBH^{+}15] Sean M. McAllister, Joshua M. Barnett, James W. Heiss, Alyssa J. Findlay, Daniel J. MacDonald, Charles L. Dow, George W. Luther III, Holly A. Michael, and Clara S. Chan. Dynamic hydrologic and biogeochemical processes drive microbially enhanced iron and sulfur cycling within the intertidal mixing zone of a beach aquifer. Limnology and Oceanography, 60 (1):329–345, January 2015. CODEN LIOCAH. ISSN 0024-3590.

Mass:2011:EPR

[MBHG11] Tali Mass, Itzchak Brickner, Erica Hendy, and Amatzia Genin. Enduring physiological and reproductive benefits of enhanced flow for a stony coral. Limnology and Oceanography, 56(6):2176– 2188, November 2011. CODEN LIOCAH. ISSN 0024-3590.

Masclaux:2011:FQA

[MBK^{+}11] Hélène Masclaux, Alexandre Bec, Maiko Kagami, Marie-Elodie Perga, Telesphore Sime-Ngando, Christian Desvilettes, and Gilles Bourdier. Food quality of anemophilous plant pollen for zooplankton. Limnology and Oceanography, 57(4):939–946, May 2011. CODEN LIOCAH. ISSN 0024-3590.

McTigue:2015:PBC

[MBLD15] Nathan D. McTigue, Philip Bucolo, Zhanfei Liu, and Kenneth H. Dunton. Pelagic-benthic coupling, food webs, and organic matter degradation in the Chukchi Sea: Insights from sedimentary pigments and stable carbon isotopes. Limnology and Oceanography, 60(2):429–445, March 2015. CODEN LIOCAH. ISSN 0024-3590.

McMurray:2011:BGB

[MBLP11] Steven E. McMurray, James E. Blum, James J. Leichter, and Joseph R. Pawlik. Bleaching of the giant barrel sponge Xestospongia muta in the Florida Keys. Limnology and Oceanography, 56(6):2243–2250, November 2011. CODEN LI-OCAH. ISSN 0024-3590.

Moynihan:2016:STD

[MBO^{+}16] Molly A. Moynihan, Pierrick Barbier, Frédéric Olivier, Nicolas Toupoint, and Tarik Meziane. Spatial and temporal dynamics of nano- and pico-size particulate organic matter (POM) in a coastal megatidal marine system. Limnology and Oceanography, 61(3):1087–1100, May 2016. CODEN LIOCAH. ISSN 0024-3590.

Merbt:2017:PNA

[MBP^{+}17] Stephanie N. Merbt, Susana Bernal, Lorenzo Proia, Eugenia Martí, and Emilio O. Casamayor. Photoinhibition on natural ammonia oxidizers biofilm populations and implications for nitrogen uptake in stream biofilms. Limnology and Oceanography, 62(1):364–375, January 2017. CODEN LIOCAH. ISSN 0024-3590.

McNair:2018:TSC

[MBTK18] Heather M. McNair, Mark A. Brzezinski, Claire P. Till, and Jeffrey W. Krause. Taxon-specific contributions to silica production in natural diatom assemblages. Limnology and Oceanography, 63(3):1056–1075, May 2018. CODEN LIOCAH. ISSN 0024-3590.

Martin:2016:NOI

[MC16] Taylor S. Martin and Karen L. Casciotti. Nitrogen and oxygen isotopic fractionation during microbial nitrite reduction. Limnology and Oceanography, 61(3):1134–1143, May 2016. CODEN LIOCAH. ISSN 0024-3590.

Milbrandt:2010:EPM

[MCC^{+}10] E. C. Milbrandt, P. G. Coble, R. N. Conmy, A. J. Martignette, and J. J. Siwicke. Evidence for the production of marine fluorescent dissolved organic matter in coastal environments and a possible mechanism for formation and dispersion. Limnology and Oceanography, 55(5):2037–2051, September 2010. CODEN LIOCAH. ISSN 0024-3590.

MacIntyre:2018:TSA

[MCCA18] Sally MacIntyre, Adam. T. Crowe, Alicia Cortés, and Lars Arneborg. Turbulence in a small Arctic pond. Limnology and Oceanography, 63(6):2337–2358, November 2018. CODEN LI-OCAH. ISSN 0024-3590.

Mourino-Carballido:2011:IFV

[MCGF^{+}11] Beatriz Mouriño-Carballido, Rocío Graña, Ana Fernàndez, Antonio Bode, Manuel Varela, J. Francisco Domínguez, José Escànez, Demetrio de Armas, and Emilio Marañón. Importance of N_{2} fixation vs. nitrate eddy diffusion along a latitudinal transect in the Atlantic Ocean. Limnology and Oceanography, 57(4): 999–1007, May 2011. CODEN LIOCAH. ISSN 0024-3590.

Malerba:2012:NND

[MCH12] Martino E. Malerba, Sean R. Connolly, and Kirsten Heimann. Nitrate–nitrite dynamics and phytoplankton growth: Formulation and experimental evaluation of a dynamic model. Limnology and Oceanography, 58(1):1555–1571, September 2012. CO-DEN LIOCAH. ISSN 0024-3590.

Maranon:2012:TRP

[MCLT12] Emilio Marañón, Pedro Cermeño, Mikel Latasa, and Rémy D. Tadonléké. Temperature, resources, and phytoplankton size structure in the ocean. Limnology and Oceanography, 57(5):

1266–1278, September 2012. CODEN LIOCAH. ISSN 0024-3590.

Maranon:2015:RSA

[MCLT15] Emilio Marañón, Pedro Cermeño, Mikel Latasa, and Rémy D. Tadonléké. Resource supply alone explains the variability of marine phytoplankton size structure. Limnology and Oceanography, 60(5):1848–1854, September 2015. CODEN LIOCAH. ISSN 0024-3590.

Moksnes:2014:LBD

[MCT^{+}14] Per-Olav Moksnes, Hanna Corell, Kentaroo Tryman, Robinson Hordoir, and Per R. Jonsson. Larval behavior and dispersal mechanisms in shore crab larvae (Carcinus maenas): Local adaptations to different tidal environments? Limnology and Oceanography, 59(6):588–602, March 2014. CODEN LIOCAH. ISSN 0024-3590.

Martin-Creuzburg:2010:IBL

[MCWB10] Dominik Martin-Creuzburg, Alexander Wacker, and Timo Basen. Interactions between limiting nutrients: Consequences for somatic and population growth of Daphnia magna. Limnology and Oceanography, 55(6):2597–2607, November 2010. CO-DEN LIOCAH. ISSN 0024-3590.

Morganti:2017:TNS

[MCYR17] T. Morganti, R. Coma, G. Yahel, and M. Ribes. Trophic niche separation that facilitates co-existence of high and low microbial abundance sponges is revealed by in situ study of carbon and nitrogen fluxes. Limnology and Oceanography, 62(5):1963–1983, September 2017. CODEN LIOCAH. ISSN 0024-3590.

Miklasz:2010:DSS

[MD10] Kevin A. Miklasz and Mark W. Denny. Diatom sinkings speeds: Improved predictions and insight from a modified Stokes’ law. Limnology and Oceanography, 55(6):2513–2525, November 2010. CODEN LIOCAH. ISSN 0024-3590.

Mobius:2015:NDU

[MD15] Jürgen Möbius and Kirstin Dähnke. Nitrate drawdown and its unexpected isotope effect in the Danube estuarine transition zone. Limnology and Oceanography, 60(3):1008–1019, May 2015. CODEN LIOCAH. ISSN 0024-3590.

Malzahn:2016:JFG

[MDB16] Arne M. Malzahn, Dijana Doerfler, and Maarten Boersma. Junk food gets healthier when it’s warm. Limnology and Oceanography, 62(3):1677–1685, September 2016. CODEN LIOCAH. ISSN 0024-3590.

Maier:2019:IVS

[MDB19] Dominique Béatrice Maier, Sebastian Diehl, and Christian Bigler. Interannual variation in seasonal diatom sedimentation reveals the importance of late winter processes and their timing for sediment signal formation. Limnology and Oceanography, 64 (3):1186–1199, May 2019. CODEN LIOCAH. ISSN 0024-3590.

Middag:2013:FDA

[MdBKL13] Rob Middag, Hein J. W. de Baar, Maarten B. Klunder, and Patrick Laan. Fluxes of dissolved aluminum and manganese to the Weddell Sea and indications for manganese co-limitation. Limnology and Oceanography, 58(2):287–300, January 2013. CODEN LIOCAH. ISSN 0024-3590.

Moran:2011:SCP

[MDE11] Xosè Anxelu G. Moràn, Hugh W. Ducklow, and Matthew Erickson. Single-cell physiological structure and growth rates of heterotrophic bacteria in a temperate estuary (Waquoit Bay, Massachusetts). Limnology and Oceanography, 56(1):37–48, January 2011. CODEN LIOCAH. ISSN 0024-3590.

Mariash:2014:BMO

[MDF^{+}14] Heather L. Mariash, Shawn P. Devlin, Laura Forsström, Roger I. Jones, and Milla Rautio. Benthic mats offer a potential subsidy to pelagic consumers in tundra pond food webs. Limnology and Oceanography, 59(3):733–744, May 2014. CODEN LIOCAH. ISSN 0024-3590.

Monismith:2010:FEB

[MDS^{+}10] Stephen G. Monismith, Kristen A. Davis, Gregory G. Shellenbarger, James L. Hench, Nicholas J. Nidzieko, Alyson E. Santoro, Matthew A. Reidenbach, Johanna H. Rosman, Roi Holtzman, Christopher S. Martens, Niels L. Lindquist, Melissa W. Southwell, and Amatzia Geninf. Flow effects on benthic grazing on phytoplankton by a Caribbean reef. Limnology and Oceanography, 55(5):1881–1892, September 2010. CODEN LIOCAH. ISSN 0024-3590.

Motwani:2018:GCT

[MDSG18] Nisha H. Motwani, Jon Duberg, Jennie B. Svedén, and Elena Gorokhova. Grazing on cyanobacteria and transfer of diazotrophic nitrogen to zooplankton in the Baltic Sea. Limnology and Oceanography, 63(2):672–686, March 2018. CODEN LIOCAH. ISSN 0024-3590.

Mehner:2010:NEE

[Meh10] Thomas Mehner. No empirical evidence for community-wide top-down control of prey fish density and size by fish predators in lakes. Limnology and Oceanography, 55(1):203–213, January 2010. CODEN LIOCAH. ISSN 0024-3590.

Meyerink:2017:EIL

[MEM^{+}17] Scott W. Meyerink, Michael J. Ellwood, William A. Maher, G. Dean Price, and Robert F. Strzepek. Effects of iron limitation on silicon uptake kinetics and elemental stoichiometry in two Southern Ocean diatoms, Eucampia antarctica and Proboscia inermis, and the temperate diatom Thalassiosira pseudonana. Limnology and Oceanography, 62(6):2445–2462, November 2017. CODEN LIOCAH. ISSN 0024-3590.

Massa:2019:PSL

[MF19] Ericka A. Massa and John M. Farrell. Prey selection by larval northern pike (Esox lucius) exposed to different zooplankton assemblages representing seasonally flooded wetland and nearshore bay habitats. Limnology and Oceanography, 64(3): 1200–1213, May 2019. CODEN LIOCAH. ISSN 0024-3590.

Masuda:2013:AUD

[MFK^{+}13] Takako Masuda, Ken Furuya, Taketoshi Kodama, Shigenobu Takeda, and Paul J. Harrison. Ammonium uptake and dinitrogen fixation by the unicellular nanocyanobacterium Crocosphaera watsonii in nitrogen-limited continuous cultures. Limnology and Oceanography, 58(6):2029–2036, November 2013. CODEN LIOCAH. ISSN 0024-3590.

Morgan:2011:LRE

[MFL11] Steven G. Morgan, Jennifer L. Fisher, and John L. Largier. Larval retention, entrainment, and accumulation in the lee of a small headland: Recruitment hotspots along windy coasts. Limnology and Oceanography, 56(1):161–178, January 2011. CO-DEN LIOCAH. ISSN 0024-3590.

Morgan:2012:LRD

[MFM^{+}12] Steven G. Morgan, Jennifer L. Fisher, Skyli T. McAfee, John L. Largier, and Chris M. Halle. Limited recruitment during relaxation events: Larval advection and behavior in an upwelling system. Limnology and Oceanography, 57(2):457–470, March 2012. CODEN LIOCAH. ISSN 0024-3590.

Maranen:2010:DOC

[MFMC^{+}10] Emilio Marañén, Ana Fernández, Beatriz Mouriño-Carballido, Sandra Martínez-García, Eva Teira, Pedro Cermeño, Paloma Chouciño, María Huete-Ortega, Emilio Fernández, Alejandra Calvo-Díaz, Xosé Anxelu G. Morán, Antonio Bode, Enrique Moreno-Ostos, Marta M. Varela, Matthew D. Patey, and Eric P. Achterberg. Degree of oligotrophy controls the response of microbial plankton to Saharan dust. Limnology and Oceanography, 55(6):2339–2352, November 2010. CODEN LIOCAH. ISSN 0024-3590.

Maszczyk:2014:SPF

[MG14] Piotr Maszczyk and Z. Maciej Gliwicz. Selectivity by planktivorous fish at different prey densities, heterogeneities, and spatial scales. Limnology and Oceanography, 59(1):68–78, January 2014. CODEN LIOCAH. ISSN 0024-3590.

Martinez-Garcia:2017:MRM

[MG17] Sandra Martínez-García. Microbial respiration in the mesopelagic zone at Station ALOHA. Limnology and Oceanography, 62(1): 320–333, January 2017. CODEN LIOCAH. ISSN 0024-3590.

Macias:2018:MFS

[MGGS18] D. Macias, E. Garcia-Gorriz, and A. Stips. Major fertilization sources and mechanisms for Mediterranean Sea coastal ecosystems. Limnology and Oceanography, 63(2):897–914, March 2018. CODEN LIOCAH. ISSN 0024-3590.

Margolin:2018:BSD

[MGHS18] Andrew R. Margolin, Margherita Gonnelli, Dennis A. Hansell, and Chiara Santinelli. Black Sea dissolved organic matter dynamics: Insights from optical analyses. Limnology and Oceanography, 63(3):1425–1443, May 2018. CODEN LIOCAH. ISSN 0024-3590.

Mutschlecner:2018:RIA

[MGJH18] Audrey E. Mutschlecner, Jennifer J. Guerard, Jeremy B. Jones, and Tamara K. Harms. Regional and intra-annual stability of dissolved organic matter composition and biolability in highlatitude Alaskan rivers. Limnology and Oceanography, 63(4): 1605–1621, July 2018. CODEN LIOCAH. ISSN 0024-3590.

Martinez-Garcia:2015:MRE

[MGK15] Sandra Martínez-García and David M. Karl. Microbial respiration in the euphotic zone at Station ALOHA. Limnology and Oceanography, 60(3):1039–1050, May 2015. CODEN LIOCAH. ISSN 0024-3590.

Mines:2013:ESZ

[MGL^{+}13] Conor H. Mines, Anas Ghadouani, Pierre Legendre, Norman D. Yan, and Gregory N. Ivey. Examining shifts in zooplankton community variability following biological invasion. Limnology and Oceanography, 58(3):399–408, January 2013. CODEN LI-OCAH. ISSN 0024-3590.

McCarthy:2016:BNR

[MGL^{+}16] Mark J. McCarthy, Wayne S. Gardner, Moritz F. Lehmann, Alexandre Guindon, and David F. Bird. Benthic nitrogen regeneration, fixation, and denitrification in a temperate, eutrophic lake: Effects on the nitrogen budget and cyanobacteria blooms. Limnology and Oceanography, 61(4):1406–1423, July 2016. CO-DEN LIOCAH. ISSN 0024-3590.

Middelboe:2012:BCC

[MGS12] Mathias Middelboe, Ronnie N. Glud, and Mikael K. Sejr. Bacterial carbon cycling in a subarctic fjord: a seasonal study on microbial activity, growth efficiency, and virus-induced mortality in kobbefjord, Greenland. Limnology and Oceanography, 57 (6):1732–1742, November 2012. CODEN LIOCAH. ISSN 0024-3590.

Mundy:2010:REE

[MGSM10] C. J. Mundy, Michel Gosselin, Michel Starr, and Christine Michel. Riverine export and the effects of circulation on dissolved organic carbon in the Hudson Bay system, Canada. Limnology and Oceanography, 55(1):315–323, January 2010. CO-DEN LIOCAH. ISSN 0024-3590.

Muschiol:2015:PDC

[MGT15] Daniel Muschiol, Olav Giere, and Walter Traunspurger. Population dynamics of a cavernicolous nematode community in a chemoautotrophic groundwater system. Limnology and Oceanography, 60(1):127–135, January 2015. CODEN LIOCAH. ISSN 0024-3590.

Menzel:2013:IBW

[MGW^{+}13] Philip Menzel, Birgit Gaye, Martin G. Wiesner, Sushma Prasad, Martina Stebich, Brijraj Krishna Das, Ambili Anoop, Nils Riedel, and Nathani Basavaiah. Influence of bottom water anoxia on nitrogen isotopic ratios and amino acid contributions of recent sediments from small eutrophic Lonar Lake, central India. Limnology and Oceanography, 58(3):1061–1074, May 2013. CODEN LIOCAH. ISSN 0024-3590.

McCormick:2016:ALA

[MH16] Amanda R. McCormick and Timothy J. Hoellein. Anthropogenic litter is abundant, diverse, and mobile in urban rivers: Insights from cross-ecosystem analyses using ecosystem and community ecology tools. Limnology and Oceanography, 62(3): 1718–1734, September 2016. CODEN LIOCAH. ISSN 0024-3590.

Maud:2018:MCH

[MHA^{+}18] Jacqueline L. Maud, Andrew G. Hirst, Angus Atkinson, Penelope K. Lindeque, and Andrea J. McEvoy. Mortality of Calanus helgolandicus: Sources, differences between the sexes and consumptive and nonconsumptive processes. Limnology and Oceanography, 63(4):1741–1761, July 2018. CODEN LIOCAH. ISSN 0024-3590.

Miyajima:2017:GCO

[MHH^{+}17] Toshihiro Miyajima, Masakazu Hori, Masami Hamaguchi, Hiromori Shimabukuro, and Goro Yoshida. Geophysical constraints for organic carbon sequestration capacity of Zostera marina seagrass meadows and surrounding habitats. Limnology and Oceanography, 63(3):954–972, May 2017. CODEN LIOCAH. ISSN 0024-3590.

Marchant:2016:CND

[MHL^{+}16] Hannah K. Marchant, Moritz Holtappels, Gaute Lavik, Soeren Ahmerkamp, Christian Winter, and Marcel M. M. Kuypers.

Coupled nitrification–denitrification leads to extensive N loss in subtidal permeable sediments. Limnology and Oceanography, 61 (3):1033–1048, May 2016. CODEN LIOCAH. ISSN 0024-3590.

Montagna:2018:EHV

[MHPW18] Paul A. Montagna, Xinping Hu, Terence A. Palmer, and Michael Wetz. Effect of hydrological variability on the biogeochemistry of estuaries across a regional climatic gradient. Limnology and Oceanography, 63(6):2465–2478, November 2018. CODEN LI-OCAH. ISSN 0024-3590.

Martin:2011:LUL

[MHRH11] Sherry L. Martin, Daniel B. Hayes, Daniel T. Rutledge, and David W. Hyndman. The land-use legacy effect: Adding temporal context to lake chemistry. Limnology and Oceanography, 56 (6):2362–2370, November 2011. CODEN LIOCAH. ISSN 0024-3590.

McMahon:2013:REA

[MHT13] Kelton W. McMahon, Li Ling Hamady, and Simon R. Thorrold. A review of ecogeochemistry approaches to estimating movements of marine animals. Limnology and Oceanography, 58(2): 697–714, March 2013. CODEN LIOCAH. ISSN 0024-3590.

McMurray:2016:SFG

[MJH^{+}16] Steven E. McMurray, Zackary I. Johnson, Dana E. Hunt, Joseph R. Pawlik, and Christopher M. Finelli. Selective feeding by the giant barrel sponge enhances foraging efficiency. Limnology and Oceanography, 61(4):1271–1286, July 2016. CODEN LIOCAH. ISSN 0024-3590.

McCrackin:2017:RLC

[MJJMM17] Michelle L. McCrackin, Holly P. Jones, Peter C. Jones, and David Moreno-Mateos. Recovery of lakes and coastal marine ecosystems from eutrophication: a global meta-analysis. Limnology and Oceanography, 62(2):507–518, March 2017. CODEN LIOCAH. ISSN 0024-3590.

Maier:2019:SUC

[MKB^{+}19] Sandra R. Maier, Tina Kutti, Raymond John Bannister, Peter van Breugel, Pieter van Rijswijk, and Dick van Oevelen. Survival under conditions of variable food availability: Resource utilization and storage in the cold-water coral Lophelia pertusa.

Limnology and Oceanography, 64(4):1651–1671, July 2019. CO-DEN LIOCAH. ISSN 0024-3590.

Meyer-Kaiser:2019:RAD

[MKBSK19] Kirstin Meyer-Kaiser, Melanie Bergmann, Thomas Soltwedel, and Michael Klages. Recruitment of Arctic deep-sea invertebrates: Results from a long-term hard-substrate colonization experiment at the Long-Term Ecological Research observatory HAUSGARTEN. Limnology and Oceanography, 64(5):1924– 1938, September 2019. CODEN LIOCAH. ISSN 0024-3590. See corrigendum [Ano21b].

Miyazako:2015:LTT

[MKG^{+}15] T. Miyazako, H. Kamiya, T. Godo, Y. Koyama, Y. Nakashima, S. Sato, M. Kishi, A. Fujihara, Y. Tabayashi, and M. Yamamuro. Long-term trends in nitrogen and phosphorus concentrations in the Hii River as influenced by atmospheric deposition from East Asia. Limnology and Oceanography, 60(2):629–640, March 2015. CODEN LIOCAH. ISSN 0024-3590.

McMeans:2015:ESS

[MKK15] Bailey C. McMeans, Apostolos-Manuel Koussoroplis, and Martin J. Kainz. Effects of seasonal seston and temperature changes on lake zooplankton fatty acids. Limnology and Oceanography, 60(2):573–583, March 2015. CODEN LIOCAH. ISSN 0024-3590.

Morgan-Kiss:2016:PPN

[MKLKP16] R. M. Morgan-Kiss, M. P. Lizotte, W. Kong, and J. C. Priscu. Photoadaptation to the polar night by phytoplankton in a permanently ice-covered Antarctic lake. Limnology and Oceanography, 61(1):3–13, January 2016. CODEN LIOCAH. ISSN 0024-3590.

Masigol:2019:CRA

[MKW^{+}19] Hossein Masigol, Seyed Akbar Khodaparast, Jason Nicholas Woodhouse, Keilor Rojas-Jimenez, Jeremy Fonvielle, Forough Rezakhani, Reza Mostowfizadeh-Ghalamfarsa, Darshan Neubauer, Tobias Goldhammer, and Hans-Peter Grossart. The contrasting roles of aquatic fungi and oomycetes in the degradation and transformation of polymeric organic matter. Limnology and Oceanography, 64(6):2662–2678, November 2019. CODEN LIOCAH. ISSN 0024-3590.

McParland:2019:RDD

[ML19] Erin L. McParland and Naomi M. Levine. The role of differential DMSP production and community composition in predicting variability of global surface DMSP concentrations. Limnology and Oceanography, 64(2):757–773, March 2019. CODEN LIOCAH. ISSN 0024-3590.

McKinnon:2013:PMW

[MLCD13] A. D. McKinnon, M. Logan, S. A. Castine, and S. Duggan. Pelagic metabolism in the waters of the Great Barrier Reef. Limnology and Oceanography, 58(4):1227–1242, July 2013. CO-DEN LIOCAH. ISSN 0024-3590.

Matveev:2016:HME

[MLD^{+}16] Alex Matveev, Isabelle Laurion, Bethany N. Deshpande, Najat Bhiry, and Warwick F. Vincent. High methane emissions from thermokarst lakes in subarctic peatlands. Limnology and Oceanography, 61(S1):S150–S164, November 2016. CODEN LI-OCAH. ISSN 0024-3590.

Milner:2016:OWM

[MLGZ16] Sara Milner, Gerald Langer, Michaël Grelaud, and Patrizia Ziveri. Ocean warming modulates the effects of acidification on Emiliania huxleyi calcification and sinking. Limnology and Oceanography, 61(4):1322–1336, July 2016. CODEN LIOCAH. ISSN 0024-3590.

Michelou:2011:PAT

[MLK11] Vanessa K. Michelou, Michael W. Lomas, and David L. Kirchman. Phosphate and adenosine-5’-triphosphate uptake by cyanobacteria and heterotrophic bacteria in the Sargasso Sea. Limnology and Oceanography, 56(1):323–332, January 2011. CODEN LIOCAH. ISSN 0024-3590.

Melancon:2014:ERN

[MLL^{+}14] Josiane Mélançon, Maurice Levasseur, Martine Lizotte, Pierre Delmelle, Jay Cullen, Roberta C. Hamme, Angelica Peña, Kyle G. Simpson, Michael Scarratt, Jean-Éric Tremblay, Jie Zhou, Keith Johnson, Nes Sutherland, Michael Arychuk, Nina Nemcek, and Marie Robert. Early response of the northeast subarctic Pacific plankton assemblage to volcanic ash fertilization. Limnology and Oceanography, 59(1):55–67, January 2014. CODEN LIOCAH. ISSN 0024-3590.

McDonald:2014:CBA

[MLS^{+}14] Cory P. McDonald, Noah R. Lottig, John L. Stoddard, Alan T. Herlihy, Sarah Lehmann, Steven G. Paulsen, David V. Peck, Amina I. Pollard, and R. Jan Stevenson. Comment on Bachmann et al. (2013): A nonrepresentative sample cannot describe the extent of cultural eutrophication of natural lakes in the United States. Limnology and Oceanography, 61(1):2226–2230, November 2014. CODEN LIOCAH. ISSN 0024-3590.

Martin:2018:PPA

[MLS^{+}18] Patrick Martin, Federico M. Lauro, Amit Sarkar, Nathalie Goodkin, Satya Prakash, and P. N. Vinayachandran. Particulate polyphosphate and alkaline phosphatase activity across a latitudinal transect in the tropical Indian Ocean. Limnology and Oceanography, 63(3):1395–1406, May 2018. CODEN LIOCAH. ISSN 0024-3590.

Mehring:2011:RMD

[MM11] Andrew S. Mehring and Timothy J. Maret. Red maple dominance enhances fungal and shredder growth and litter processing in temporary ponds. Limnology and Oceanography, 57(4): 1106–1114, May 2011. CODEN LIOCAH. ISSN 0024-3590.

McDowell:2017:MMD

[MMB17] W. G. McDowell, W. H. McDowell, and J. E. Byers. Mass mortality of a dominant invasive species in response to an extreme climate event: Implications for ecosystem function. Limnology and Oceanography, 62(1):177–188, January 2017. CODEN LI-OCAH. ISSN 0024-3590.

Moreno-Marin:2018:ARM

[MMBP18] Francisco Moreno-Marín, Fernando Guillermo Brun, and Morten Foldager Pedersen. Additive response to multiple environmental stressors in the seagrass Zostera marina L. Limnology and Oceanography, 63(4):1528–1544, July 2018. CODEN LIOCAH. ISSN 0024-3590.

Maske:2010:ICB

[MMC^{+}10] Helmut Maske, Ramón Cajal Medrano, Armando Trasviña Castro, Alejandrina Jiménez Mercado, Cesar O. Almeda Jauregui, Gilberto Gaxiola Castro, and José Ochoa. Inorganic carbon and biological oceanography above a shallow oxygen minimum in the entrance to the Gulf of California in the Mexican Pacific. Limnology and Oceanography, 55(2):481–491, March 2010. CODEN LIOCAH. ISSN 0024-3590.

Morison:2015:ESP

[MMD15] Françoise Morison and Susanne Menden-Deuer. Early spring phytoplankton dynamics in the subpolar North Atlantic: The influence of protistan herbivory. Limnology and Oceanography, 60(4):1298–1313, July 2015. CODEN LIOCAH. ISSN 0024-3590.

Morison:2018:SSR

[MMD18] F. Morison and S. Menden-Deuer. Seasonal similarity in rates of protistan herbivory in fjords along the Western Antarctic Peninsula. Limnology and Oceanography, 63(6):2858–2876, November 2018. CODEN LIOCAH. ISSN 0024-3590.

Morales-Marin:2018:TDH

[MMFBB18] L. A. Morales-Marin, J. R. French, H. Burningham, and R. W. Battarbee. Three-dimensional hydrodynamic and sediment transport modeling to test the sediment focusing hypothesis in upland lakes. Limnology and Oceanography, 63(S1):S156–S176, March 2018. CODEN LIOCAH. ISSN 0024-3590.

Muller:2016:ARC

[MMG16] Beat Müller, Joseph S. Meyer, and René Gächter. Alkalinity regulation in calcium carbonate-buffered lakes. Limnology and Oceanography, 61(1):341–352, January 2016. CODEN LIOCAH. ISSN 0024-3590.

Mazarrasa:2017:DCS

[MMGO^{+}17a] Inés Mazarrasa, Núria Marbà, Jordi Garcia-Orellana, Pere Masqué, Ariane Arias-Ortiz, and Carlos M. Duarte. Dynamics of carbon sources supporting burial in seagrass sediments under increasing anthropogenic pressure. Limnology and Oceanography, 62(4):1451–1465, July 2017. CODEN LIOCAH. ISSN 0024-3590.

Mazarrasa:2017:EEF

[MMGO^{+}17b] Inés Mazarrasa, Núria Marbà, Jordi Garcia-Orellana, Pere Masqué, Ariane Arias-Ortiz, and Carlos M. Duarte. Effect of environmental factors (wave exposure and depth) and anthropogenic pressure in the C sink capacity of Posidonia oceanica meadows. Limnology and Oceanography, 62(4):1436–1450, July 2017. CODEN LIOCAH. ISSN 0024-3590.

Mandic-Mulec:2012:MPS

[MMGP^{+}12] Ines Mandic-Mulec, Katja Gorenc, Marinka Gams Petrišiš, Jadran Faganeli, and Nives Ogrinc. Methanogenesis pathways in a stratified eutrophic alpine lake (Lake Bled, Slovenia). Limnology and Oceanography, 57(3):868–880, May 2012. CODEN LIOCAH. ISSN 0024-3590.

Mochizuki:2018:DTE

[MMH^{+}18] Akihito Mochizuki, Takahiro Murata, Ko Hosoda, Ayuriin Dulmaa, Chananbaatar Ayushsuren, Darmaa Ganchimeg, Valentin V. Drucker, Vladimir A. Fialkov, Tolga Depci, Tijen Üner, Fatih Oğhan, and Masahito Sugiyama. Distribution of trace elements and the influence of major-ion water chemistry in saline lakes. Limnology and Oceanography, 63(3):1253–1263, May 2018. CODEN LIOCAH. ISSN 0024-3590.

Moore:2010:RIN

[MMHT10] Stephanie K. Moore, Nathan J. Mantua, Barbara M. Hickey, and Vera L. Trainer. The relative influences of El Niño–Southern Oscillation and Pacific decadal oscillation on paralytic shellfish toxin accumulation in northwest Pacific shellfish. Limnology and Oceanography, 55(6):2262–2274, November 2010. CODEN LIOCAH. ISSN 0024-3590.

Moller:2012:ECT

[MMJ^{+}12] Eva Friis Møller, Marie Maar, Sigrún H. Jónasdóttir, Torkel Gissel Nielsen, and Kajsa Tönnesson. The effect of changes in temperature and food on the development of Calanus finmarchicus and Calanus helgolandicus populations. Limnology and Oceanography, 57(1):211–220, January 2012. CODEN LIOCAH. ISSN 0024-3590.

Matzinger:2010:HOC

[MMN^{+}10] Andreas Matzinger, Beat Müller, Pius Niederhauser, Martin Schmid, and Alfred Wüest. Hypolimnetic oxygen consumption by sediment-based reduced substances in former eutrophic lakes. Limnology and Oceanography, 55(5):2073–2084, September 2010. CODEN LIOCAH. ISSN 0024-3590.

Miyatake:2014:TCF

[MMPSB14] Tetsuro Miyatake, Tanja C. W. Moerdijk-Poortvliet, Lucas J. Stal, and Henricus T. S. Boschker. Tracing carbon flow from microphytobenthos to major bacterial groups in an intertidal marine sediment by using an in situ ^{13}C pulse-chase method. Limnology and Oceanography, 59(4):1275–1287, July 2014. CO-DEN LIOCAH. ISSN 0024-3590.

Middleton:2017:FDE

[MMWR17] Julia E. Middleton, Joaquín Martínez Martínez, William H. Wilson, and Nicholas R. Record. Functional dynamics of Emiliania huxleyi virus-host interactions across multiple spatial scales. Limnology and Oceanography, 63(3):922–933, May 2017. CO-DEN LIOCAH. ISSN 0024-3590.

McIntosh:2015:SAD

[MMXC15] Hadley A. McIntosh, Ann P. McNichol, Li Xu, and Elizabeth A. Canuel. Source-age dynamics of estuarine particulate organic matter using fatty acid δ ^{13}C and Δ ^{14}C composition. Limnology and Oceanography, 60(2):611–628, March 2015. CODEN LIOCAH. ISSN 0024-3590.

Ma:2019:CDS

[MNW^{+}19] Xiaolin Ma, Yijun Ni, Xiaoyu Wang, Wei Hu, and Mingbo Yin. Clonal diversity and substantial genetic divergence of the Daphnia similis species complex in Chinese lakes: Possible adaptations to the uplift of the Qinghai–Tibetan Plateau. Limnology and Oceanography, 64(6):2725–2737, November 2019. CODEN LIOCAH. ISSN 0024-3590.

Mullarney:2017:DRK

[MP17] Julia C. Mullarney and Conrad A. Pilditch. The differential response of kelp to swell and infragravity wave motion. Limnology and Oceanography, 62(6):2524–2537, November 2017. CODEN LIOCAH. ISSN 0024-3590.

Millette:2017:MHR

[MPAS17] N. C. Millette, J. J. Pierson, A. Aceves, and D. K. Stoecker. Mixotrophy in Heterocapsa rotundata: a mechanism for dominating the winter phytoplankton. Limnology and Oceanography, 62(2):836–845, March 2017. CODEN LIOCAH. ISSN 0024-3590.

Masclaux:2013:HPO

[MPK^{+}13] Hélène Masclaux, Marie-Elodie Perga, Maiko Kagami, Christian Desvilettes, Gilles Bourdier, and Alexandre Bec. How pollen organic matter enters freshwater food webs. Limnology and Oceanography, 58(4):1185–1195, July 2013. CODEN LIOCAH. ISSN 0024-3590.

Maps:2015:LAF

[MPM^{+}15] Frédéric Maps, Stéphane Plourde, Ian H. McQuinn, Simon St-Onge-Drouin, Diane Lavoie, Joël Chassé, and Véronique Lesage. Linking acoustics and finite-time Lyapunov exponents reveals areas and mechanisms of krill aggregation within the Gulf of St. Lawrence, eastern Canada. Limnology and Oceanography, 60 (6):1965–1975, November 2015. CODEN LIOCAH. ISSN 0024-3590.

Martinez-Perez:2017:MCD

[MPONC^{+}17] Alba María Martínez-Pérez, Helena Osterholz, Mar Nieto-Cid, Marta Álvarez, Thorsten Dittmar, and Xosé Antón Álvarez-Salgado. Molecular composition of dissolved organic matter in the Mediterranean Sea. Limnology and Oceanography, 62(6): 2699–2712, November 2017. CODEN LIOCAH. ISSN 0024-3590.

McNally:2017:MIS

[MPSA17] Sean P. McNally, Rachel J. Parsons, Alyson E. Santoro, and Amy Apprill. Multifaceted impacts of the stony coral Porites astreoides on picoplankton abundance and community composition. Limnology and Oceanography, 62(1):217–234, January 2017. CODEN LIOCAH. ISSN 0024-3590.

Moerdijk-Poortvliet:2018:SCB

[MPvBS^{+}18] Tanja C. W. Moerdijk-Poortvliet, Peter van Breugel, Koen Sabbe, Olivier Beauchard, Lucas J. Stal, and Henricus T. S. Boschker. Seasonal changes in the biochemical fate of carbon fixed by benthic diatoms in intertidal sediments. Limnology and Oceanography, 63(2):550–569, March 2018. CODEN LIOCAH. ISSN 0024-3590.

Munro:2013:BPR

[MQJG13] David R. Munro, Paul D. Quay, Laurie W. Juranek, and Ralf Goericke. Biological production rates off the Southern California coast estimated from triple O_{2} isotopes and O_{2} : Ar gas ratios. Limnology and Oceanography, 58(4):1312–1328, July 2013. CO-DEN LIOCAH. ISSN 0024-3590.

Ma:2016:PCM

[MQP^{+}16] Jianrong Ma, Boqiang Qin, Hans W. Paerl, Justin D. Brookes, Nathan S. Hall, Kun Shi, Yongqiang Zhou, Jinsong Guo, Zhe Li, Hai Xu, Tingfeng Wu, and Shengxing Long. The persistence of cyanobacterial (Microcystis spp.) blooms throughout winter in Lake Taihu, China. Limnology and Oceanography, 61(2):711– 722, March 2016. CODEN LIOCAH. ISSN 0024-3590.

Miller:2011:PPP

[MRB11] Robert J. Miller, Daniel C. Reed, and Mark A. Brzezinski. Partitioning of primary production among giant kelp (Macrocystis pyrifera), understory macroalgae, and phytoplankton on a temperate reef. Limnology and Oceanography, 56(1):119–132, January 2011. CODEN LIOCAH. ISSN 0024-3590.

Maldonado:2010:RSB

[MRBR10] Manuel Maldonado, Ana Riesgo, Arianna Bucci, and Klaus Rützler. Revisiting silicon budgets at a tropical continental shelf: Silica standing stocks in sponges surpass those in diatoms. Limnology and Oceanography, 55(5):2001–2010, September 2010. CODEN LIOCAH. ISSN 0024-3590.

Morana:2016:CAP

[MRC^{+}16] Cédric Morana, Fleur A. E. Roland, Sean A. Crowe, Marc Llirós, Alberto V. Borges, François Darchambeau, and Steven Bouillon. Chemoautotrophy and anoxygenic photosynthesis within the water column of a large meromictic tropical lake (Lake Kivu, East Africa). Limnology and Oceanography, 61(4):1424–1437, July 2016. CODEN LIOCAH. ISSN 0024-3590.

Mousing:2018:GPP

[MRE18] Erik Askov Mousing, Katherine Richardson, and Marianne Ellegaard. Global patterns in phytoplankton biomass and community size structure in relation to macronutrients in the open ocean. Limnology and Oceanography, 63(3):1298–1312, May 2018. CODEN LIOCAH. ISSN 0024-3590.

Murray:2015:ISH

[MRH^{+}15] James W. Murray, Emily Roberts, Evan Howard, Michael O’Donnell, Cory Bantam, Emily Carrington, Mike Foy, Barbara Paul, and Amanda Fay. An inland sea high nitrate-low chlorophyll (HNLC) region with naturally high pCO_{2}. Limnology and Oceanography, 60(3):957–966, May 2015. CODEN LIOCAH. ISSN 0024-3590.

Macey:2014:PPS

[MRKR^{+}14] A. I. Macey, T. Ryan-Keogh, S. Richier, C. M. Moore, and T. S. Bibby. Photosynthetic protein stoichiometry and photophysiology in the high latitude North Atlantic. Limnology and Oceanography, 59(6):1853–1864, November 2014. CODEN LI-OCAH. ISSN 0024-3590.

MacIntyre:2014:SHE

[MRSE14] Sally MacIntyre, José R. Romero, Gregory M. Silsbe, and Brian M. Emery. Stratification and horizontal exchange in Lake Victoria, East Africa. Limnology and Oceanography, 59(6):1805– 1838, November 2014. CODEN LIOCAH. ISSN 0024-3590.

McDonald:2012:RAS

[MRSS12] Cory P. McDonald, Jennifer A. Rover, Edward G. Stets, and Robert G. Striegl. The regional abundance and size distribution of lakes and reservoirs in the United States and implications for estimates of global lake extent. Limnology and Oceanography, 57 (3):597–606, March 2012. CODEN LIOCAH. ISSN 0024-3590.

Mullaney:2013:ERC

[MS13] Thomas J. Mullaney and Iain M. Suthers. Entrainment and retention of the coastal larval fish assemblage by a short-lived, submesoscale, frontal eddy of the East Australian Current. Limnology and Oceanography, 58(5):1546–1556, September 2013. CO-DEN LIOCAH. ISSN 0024-3590.

Middelbo:2018:IGM

[MSAM18] Ane Bruun Middelbo, Mikael Kristian Sejr, Kristine Engel Arendt, and Eva Friis Møller. Impact of glacial meltwater on spatiotemporal distribution of copepods and their grazing impact in Young Sound, NE Greenland. Limnology and Oceanography, 63(4):322–336, January 2018. CODEN LIOCAH. ISSN 0024-3590.

Morana:2014:PDO

[MSD^{+}14] Cedric Morana, Hugo Sarmento, Jean-Pierre Descy, Josep M. Gasol, Alberto V. Borges, Steven Bouillon, and François Darchambeau. Production of dissolved organic matter by phytoplankton and its uptake by heterotrophic prokaryotes in large tropical lakes. Limnology and Oceanography, 59(4):1364–1375, July 2014. CODEN LIOCAH. ISSN 0024-3590.

Maher:2013:GDD

[MSGS^{+}13] D. T. Maher, I. R. Santos, L. Golsby-Smith, J. Gleeson, and B. D. Eyre. Groundwater-derived dissolved inorganic and organic carbon exports from a mangrove tidal creek: The missing mangrove carbon sink? Limnology and Oceanography, 58(2): 475–488, March 2013. CODEN LIOCAH. ISSN 0024-3590.

Meyer:2017:RBI

[MSK^{+}17] Kirstin S. Meyer, Andrew K. Sweetman, Piotr Kuklinski, Peter Leopold, Daniel Vogedes, Jørgen Berge, Colin Griffiths, Craig M. Young, and Paul E. Renaud. Recruitment of benthic invertebrates in high Arctic fjords: Relation to temperature, depth, and season. Limnology and Oceanography, 62(6):2732– 2744, November 2017. CODEN LIOCAH. ISSN 0024-3590.

Morgan:2017:SZR

[MSM^{+}17] Steven G. Morgan, Alan L. Shanks, Jamie MacMahan, Ad J. H. M. Reniers, Chris D. Griesemer, Marley Jarvis, and Atsushi G. Fujimura. Surf zones regulate larval supply and zooplankton subsidies to nearshore communities. Limnology and Oceanography, 62(6):2811–2828, November 2017. CODEN LI-OCAH. ISSN 0024-3590.

Muller:2016:LTA

[MSR16] Jens Daniel Müller, Bernd Schneider, and Gregor Rehder. Longterm alkalinity trends in the Baltic Sea and their implications for CO_{2}-induced acidification. Limnology and Oceanography, 61 (6):1984–2002, November 2016. CODEN LIOCAH. ISSN 0024-3590.

Martin:2018:CTS

[MSS^{+}18] Belinda C. Martin, John Statton, Andre R. Siebers, Pauline F. Grierson, Megan H. Ryan, and Gary A. Kendrick. Colonizing tropical seagrasses increase root exudation under fluctuating and continuous low light. Limnology and Oceanography, 63(S1): S381–S391, March 2018. CODEN LIOCAH. ISSN 0024-3590.

Malkin:2012:DCM

[MSSH12] Sairah Y. Malkin, Greg M. Silsbe, Ralph E. H. Smith, and E. Todd Howell. A deep chlorophyll maximum nourishes benthic filter feeders in the coastal zone of a large clear lake. Limnology and Oceanography, 57(3):735–748, May 2012. CODEN LIOCAH. ISSN 0024-3590.

Meiggs:2011:ERD

[MT11] Deidre Meiggs and Martial Taillefert. The effect of riverine discharge on biogeochemical processes in estuarine sediments. Limnology and Oceanography, 56(5):1797–1810, September 2011. CODEN LIOCAH. ISSN 0024-3590.

McMahon:2015:TDN

[MTEM15] Kelton W. McMahon, Simon R. Thorrold, Travis S. Elsdon, and Matthew D. McCarthy. Trophic discrimination of nitrogen stable isotopes in amino acids varies with diet quality in a marine fish. Limnology and Oceanography, 60(3):1076–1087, May 2015. CODEN LIOCAH. ISSN 0024-3590.

McGillicuddy:2011:SAF

[MTH^{+}11] D. J. and,, Jr. McGillicuddy, D. W. Townsend, R. He, B. A. Keafer, J. L. Kleindinst, Y. Li, J. P. Manning, D. G. Mountain, M. A. Thomas, and D. M. Anderson. Suppression of the 2010 Alexandrium fundyense bloom by changes in physical, biological, and chemical properties of the Gulf of Maine. Limnology and Oceanography, 56(6):2411–2426, November 2011. CODEN LIOCAH. ISSN 0024-3590.

Mojib:2017:CMR

[MTK^{+}17] N. Mojib, M. Thimma, M. Kumaran, R. Sougrat, and X. Irigoien. Comparative metatranscriptomics reveals decline of a neustonic planktonic population. Limnology and Oceanography, 62(1):299–310, January 2017. CODEN LIOCAH. ISSN 0024-3590.

Martiny:2016:BIC

[MTM^{+}16] Adam C. Martiny, Agathe Talarmin, Céline Mouginot, Jeanette A. Lee, Jeremy S. Huang, Alyssa G. Gellene, and David A. Caron. Biogeochemical interactions control a temporal succession in the elemental composition of marine communities. Limnology and Oceanography, 61(2):531–542, March 2016. CODEN LIOCAH. ISSN 0024-3590.

Marzocchi:2018:ESD

[MTSG18] Ugo Marzocchi, Bo Thamdrup, Peter Stief, and Ronnie N. Glud. Effect of settled diatom-aggregates on benthic nitrogen cycling. Limnology and Oceanography, 63(4):431–444, January 2018. CODEN LIOCAH. ISSN 0024-3590.

Majdi:2017:LTC

[MTT17] Nabil Majdi, Ilka Threis, and Walter Traunspurger. It’s the little things that count: Meiofaunal density and production in the sediment of two headwater streams. Limnology and Oceanography, 62(1):151–163, January 2017. CODEN LIOCAH. ISSN 0024-3590.

Makino:2018:FFC

[MTU18] Wataru Makino, Akifumi S. Tanabe, and Jotaro Urabe. The fauna of freshwater calanoid copepods in Japan in the early decades of the 21^{st} century: Implications for the assessment and conservation of biodiversity. Limnology and Oceanography, 63 (2):758–772, March 2018. CODEN LIOCAH. ISSN 0024-3590.

Moffett:2012:CII

[MTW12] James W. Moffett, Caroline B. Tuit, and B. B. Ward. Chelatorinduced inhibition of copper metalloenzymes in denitrifying bacteria. Limnology and Oceanography, 57(1):272–280, January 2012. CODEN LIOCAH. ISSN 0024-3590.

Miura:2017:CEF

[MU17] Aya Miura and Jotaro Urabe. Changes in epilithic fungal communities under different light conditions in a river: a field experimental study. Limnology and Oceanography, 62(2):579–591, March 2017. CODEN LIOCAH. ISSN 0024-3590.

Mojica:2015:PCS

[MvdPK^{+}15] Kristina D. A. Mojica, Willem H. van de Poll, Michael Kehoe, Jef Huisman, Klaas R. Timmermans, Anita G. J. Buma, Hans J. van der Woerd, Lisa Hahn-Woernle, Henk A. Dijkstra, and Corina P. D. Brussaard. Phytoplankton community structure in relation to vertical stratification along a north-south gradient in the Northeast Atlantic Ocean. Limnology and Oceanography, 60(5):1498–1521, September 2015. CODEN LIOCAH. ISSN 0024-3590.

Moffett:2015:BIA

[MVG^{+}15] James W. Moffett, Jagruti Vedamati, Tyler J. Goepfert, Anil Pratihary, Mangesh Gauns, and S. W. A. Naqvi. Biogeochemistry of iron in the Arabian Sea. Limnology and Oceanography, 60(5):1671–1688, September 2015. CODEN LIOCAH. ISSN 0024-3590.

Marchetti:2010:ISA

[MVL^{+}10] Adrian Marchetti, Diana E. Varela, Veronica P. Lance, Veronica P. Lance, Matteo Palmucci, Mario Giordano, and E. Virginia Armbrust. Iron and silicic acid effects on phytoplankton productivity, diversity, and chemical composition in the central equatorial Pacific Ocean. Limnology and Oceanography, 55(1): 11–29, January 2010. CODEN LIOCAH. ISSN 0024-3590.

Mette:2011:PCS

[MVNG11] Elizabeth M. Mette, Michael J. Vanni, Jennifer M. Newell, and María J. Gonzàlez. Phytoplankton communities and stoichiometry are interactively affected by light, nutrients, and fish. Limnology and Oceanography, 56(6):1959–1975, November 2011. CODEN LIOCAH. ISSN 0024-3590.

Milici:2017:DCC

[MVT^{+}17] Mathias Milici, Marius Vital, Jürgen Tomasch, Thomas H. Badewien, Helge-A. Giebel, Iris Plumeier, Hui Wang, Dietmar H. Pieper, Irene Wagner-Döbler, and Meinhard Simon. Diversity and community composition of particle-associated and free-living bacteria in mesopelagic and bathypelagic Southern Ocean water masses: Evidence of dispersal limitation in the Bransfield Strait. Limnology and Oceanography, 63(3):1080– 1095, May 2017. CODEN LIOCAH. ISSN 0024-3590.

Moquin:2015:EPD

[MW15] Paul A. Moquin and Frederick J. Wrona. Effects of permafrost degradation on water and sediment quality and heterotrophic bacterial production of Arctic tundra lakes: an experimental approach. Limnology and Oceanography, 60(5):1484–1497, September 2015. CODEN LIOCAH. ISSN 0024-3590.

Madhuri:2019:CTD

[MWBM19] Sumeda Madhuri, Kai Wang, Darren Bade, and Xiaozhen Mou. Concentration and turnover of dissolved free polyamines on the South Coast of Lake Erie. Limnology and Oceanography, 64(4): 1641–1650, July 2019. CODEN LIOCAH. ISSN 0024-3590.

Mette:2016:LLS

[MWC^{+}16] Madelyn J. Mette, Alan D. Wanamaker, Jr., Michael L. Carroll, William G. Ambrose, Jr., and Michael J. Retelle. Linking large-scale climate variability with Arctica islandica shell growth and geochemistry in northern Norway. Limnology and Oceanography, 61(2):748–764, March 2016. CODEN LIOCAH. ISSN 0024-3590.

McMullin:2017:IFO

[MWR17] Rebecca M. McMullin, Stephen R. Wing, and Malcolm R. Reid. Ice fish otoliths record dynamics of advancing and retreating sea ice in Antarctica. Limnology and Oceanography, 62(6):2662– 2673, November 2017. CODEN LIOCAH. ISSN 0024-3590.

McLeod:2010:HII

[MWS10] Rebecca J. McLeod, Stephen R. Wing, and Jennifer E. Skilton. High incidence of invertebrate-chemoautotroph symbioses in benthic communities of the New Zealand fjords. Limnology and Oceanography, 55(5):2097–2106, September 2010. CODEN LIOCAH. ISSN 0024-3590.

Modenutti:2018:WEP

[MWSB18] Beatriz E. Modenutti, Laura Wolinski, Maria S. Souza, and Esteban G. Balseiro. When eating a prey is risky: Implications for predator diel vertical migration. Limnology and Oceanography, 63(2):939–950, March 2018. CODEN LIOCAH. ISSN 0024-3590.

Millette:2011:PDD

[MXWC11] Katie L. Millette, Sen Xu, Jonathan D. S. Witt, and Melania E. Cristescu. Pleistocene-driven diversification in freshwater zooplankton: Genetic patterns of refugial isolation and postglacial recolonization in Leptodora kindtii (Crustacea, Cladocera). Limnology and Oceanography, 56(5):1725–1736, September 2011. CODEN LIOCAH. ISSN 0024-3590.

Mouginot:2015:RAM

[MZB^{+}15] Céline Mouginot, Amy E. Zimmerman, Juan A. Bonachela, Helen Fredricks, Steven D. Allison, Benjamin A. S. Van Mooy, and Adam C. Martiny. Resource allocation by the marine cyanobacterium Synechococcus WH8102 in response to different nutrient supply ratios. Limnology and Oceanography, 60(5):1634–1641, September 2015. CODEN LIOCAH. ISSN 0024-3590.

McPherson:2015:PCI

[MZH15] Meredith L. McPherson, Richard C. Zimmerman, and Victoria. J. Hill. Predicting carbon isotope discrimination in eelgrass (Zostera marina L.) from the environmental parameters — light, flow, and [DIC]. Limnology and Oceanography, 60(6):1875–1889, November 2015. CODEN LIOCAH. ISSN 0024-3590.

Nishizaki:2017:MBR

[NA17] Michael Nishizaki and Josef Daniel Ackerman. Mussels blow rings: Jet behavior affects local mixing. Limnology and Oceanography, 62(1):125–136, January 2017. CODEN LIOCAH. ISSN 0024-3590.

Ng:2011:PCS

[NAH^{+}11] S. M. Y. Ng, J. P. Antenucci, M. R. Hipsey, G. Tibor, and T. Zohary. Physical controls on the spatial evolution of a dinoflagellate bloom in a large lake. Limnology and Oceanography, 56 (6):2265–2281, November 2011. CODEN LIOCAH. ISSN 0024-3590.

Nordstrom:2017:OES

[NB17] Marie C. Nordström and Erik Bonsdorff. Organic enrichment simplifies marine benthic food web structure. Limnology and Oceanography, 65(9):2179–2188, September 2017. CODEN LI-OCAH. ISSN 0024-3590.

Noisette:2016:APR

[NBDM16] Fanny Noisette, François Bordeyne, Dominique Davoult, and Sophie Martin. Assessing the physiological responses of the gastropod Crepidula fornicata to predicted ocean acidification and warming. Limnology and Oceanography, 61(2):430–444, March 2016. CODEN LIOCAH. ISSN 0024-3590.

Nowacki:2017:SWD

[NBG17] Daniel J. Nowacki, Alexis Beudin, and Neil K. Ganju. Spectral wave dissipation by submerged aquatic vegetation in a backbarrier estuary. Limnology and Oceanography, 62(2):736–753, March 2017. CODEN LIOCAH. ISSN 0024-3590.

Nova:2019:LBE

[NBSMN19] Clarice Casa Nova, Reinaldo Luiz Bozelli, Alejandro Spitzy, and Dörthe Müller-Navarra. Living in a browning environment: Effects on Daphnia’s growth and fatty acid pattern. Limnology and Oceanography, 64(1):18–31, January 2019. CODEN LIOCAH. ISSN 0024-3590.

Nifong:2014:HNL

[NCC14] Rachel L. Nifong, Matthew J. Cohen, and Wendell P. Cropper, Jr. Homeostasis and nutrient limitation of benthic autotrophs in natural chemostats. Limnology and Oceanography, 59(6):2101– 2111, November 2014. CODEN LIOCAH. ISSN 0024-3590.

Nomaki:2014:NUF

[NCT^{+}14] Hidetaka Nomaki, Yoshito Chikaraishi, Masashi Tsuchiya, Takashi Toyofuku, Naohiko Ohkouchi, Katsuyuki Uematsu, Akihiro Tame, and Hiroshi Kitazato. Nitrate uptake by Foraminifera and use in conjunction with endobionts under anoxic conditions. Limnology and Oceanography, 59(6):1879– 1888, November 2014. CODEN LIOCAH. ISSN 0024-3590.

Nomaki:2015:VNI

[NCT^{+}15] Hidetaka Nomaki, Yoshito Chikaraishi, Masashi Tsuchiya, Takashi Toyofuku, Hisami Suga, Yoko Sasaki, Katsuyuki Uematsu, Akihiro Tame, and Naohiko Ohkouchi. Variation in the nitrogen isotopic composition of amino acids in benthic Foraminifera: Implications for their adaptation to oxygendepleted environments. Limnology and Oceanography, 60(6): 1906–1916, November 2015. CODEN LIOCAH. ISSN 0024-3590.

Nascimento:2019:OSI

[NEH^{+}19] Gabriela S. Nascimento, Timothy I. Eglinton, Negar Haghipour, Ana Luiza Albuquerque, Anelize Bahniuk, Judith A. McKenzie, and Crisogono Vasconcelos. Oceanographic and sedimentological influences on carbonate geochemistry and mineralogy in hypersaline coastal lagoons, Rio de Janeiro state, Brazil. Limnology and Oceanography, 64(6):2605–2620, November 2019. CO-DEN LIOCAH. ISSN 0024-3590.

Negri:2011:HIV

[NFRU11] Andrew P. Negri, Florita Flores, Till Röthig, and Sven Uthicke. Herbicides increase the vulnerability of corals to rising sea surface temperature. Limnology and Oceanography, 56(2):471–485, March 2011. CODEN LIOCAH. ISSN 0024-3590.

Newell:2013:DDA

[NFW13] Silvia E. Newell, Sarah E. Fawcett, and Bess B. Ward. Depth distribution of ammonia oxidation rates and ammonia-oxidizer community composition in the Sargasso Sea. Limnology and Oceanography, 58(4):1491–1500, July 2013. CODEN LIOCAH. ISSN 0024-3590.

Norton:2013:EDB

[NG13] Emily L. Norton and Erica Goetze. Equatorial dispersal barriers and limited population connectivity among oceans in a planktonic copepod. Limnology and Oceanography, 58(5):1581–1596, September 2013. CODEN LIOCAH. ISSN 0024-3590.

Nguyen:2017:ICW

[NHP17] Tuan D. Nguyen, Nathan Hawley, and Mantha S. Phanikumar. Ice cover, winter circulation, and exchange in Saginaw Bay and Lake Huron. Limnology and Oceanography, 62(1):376–393, January 2017. CODEN LIOCAH. ISSN 0024-3590.

Noffke:2012:BIP

[NHS^{+}12] A. Noffke, C. Hensen, S. Sommer, F. Scholz, L. Bohlen, T. Mosch, M. Graco, and K. Wallmann. Benthic iron and phosphorus fluxes across the Peruvian oxygen minimum zone. Limnology and Oceanography, 57(3):851–867, May 2012. CODEN LIOCAH. ISSN 0024-3590.

Nakayama:2010:RCD

[NI10] K. Nakayama and J. Imberger. Residual circulation due to internal waves shoaling on a slope. Limnology and Oceanography, 55 (3):1009–1023, May 2010. CODEN LIOCAH. ISSN 0024-3590.

Noss:2014:DOB

[NL14] Christian Noss and Andreas Lorke. Direct observation of biomixing by vertically migrating zooplankton. Limnology and Oceanography, 59(3):724–732, May 2014. CODEN LIOCAH. ISSN 0024-3590.

Norzagaray-Lopez:2017:ASP

[NLHAA^{+}17] C. O. Norzagaray-López, J. M. Hernández-Ayón, L. E. Calderon Aguilera, H. Reyes-Bonilla, C. Chapa-Balcorta, and A. Ayala-Bocos. Aragonite saturation and pH variation in a fringing reef are strongly influenced by oceanic conditions. Limnology and Oceanography, 62(6):2375–2388, November 2017. CODEN LI-OCAH. ISSN 0024-3590.

Neukermans:2012:SVM

[NLM^{+}12] Griet Neukermans, Hubert Loisel, Xavier Mériaux, Rosa Astoreca, and David McKee. In situ variability of mass-specific beam attenuation and backscattering of marine particles with respect to particle size, density, and composition. Limnology and Oceanography, 57(1):124–144, January 2012. CODEN LI-OCAH. ISSN 0024-3590.

Noble:2012:BSI

[NLO^{+}12] Abigail E. Noble, Carl H. Lamborg, Dan C. Ohnemus, Phoebe J. Lam, Tyler J. Goepfert, Chris I. Measures, Caitlin H. Frame, Karen L. Casciotti, Giacomo R. DiTullio, Joe Jennings, and Mak A. Saito. Basin-scale inputs of cobalt, iron, and manganese from the Benguela–Angola front to the South Atlantic Ocean. Limnology and Oceanography, 57(4):989–1010, July 2012. CO-DEN LIOCAH. ISSN 0024-3590.

Nayak:2018:EUP

[NMST18] Aditya R. Nayak, Malcolm N. McFarland, James M. Sullivan, and Michael S. Twardowski. Evidence for ubiquitous preferential particle orientation in representative oceanic shear flows. Limnology and Oceanography, 63(1):122–143, January 2018. CO-DEN LIOCAH. ISSN 0024-3590.

Nascimento:2012:MEO

[NNE12] Francisco J. A. Nascimento, Johan Näslund, and Ragnar Elmgren. Meiofauna enhances organic matter mineralization in soft sediment ecosystems. Limnology and Oceanography, 57(1):338– 346, January 2012. CODEN LIOCAH. ISSN 0024-3590.

Nishioka:2017:DID

[NO17] Jun Nishioka and Hajime Obata. Dissolved iron distribution in the western and central subarctic Pacific: HNLC water formation and biogeochemical processes. Limnology and Oceanography, 65(9):2004–2022, September 2017. CODEN LIOCAH. ISSN 0024-3590.

Neubacher:2011:STH

[NPT11] Elke C. Neubacher, Ruth E. Parker, and Mark Trimmer. Shortterm hypoxia alters the balance of the nitrogen cycle in coastal sediments. Limnology and Oceanography, 56(2):651–665, March 2011. CODEN LIOCAH. ISSN 0024-3590.

Nishri:2015:MHW

[NRL15] Aminadav Nishri, Alon Rimmer, and Yury Lechinsky. The mechanism of hypolimnion warming induced by internal waves. Limnology and Oceanography, 60(4):1462–1476, July 2015. CO-DEN LIOCAH. ISSN 0024-3590.

Neukermans:2016:OCC

[NRS16] Griet Neukermans, Rick A. Reynolds, and Dariusz Stramski. Optical classification and characterization of marine particle assemblages within the western Arctic Ocean. Limnology and Oceanography, 61(4):1472–1494, July 2016. CODEN LIOCAH. ISSN 0024-3590.

Natchimuthu:2016:STV

[NSG^{+}16] Sivakiruthika Natchimuthu, Ingrid Sundgren, Magnus Gålfalk, Leif Klemedtsson, Patrick Crill, Åsa Danielsson, and David Bastviken. Spatio-temporal variability of lake CH_{4} fluxes and its influence on annual whole lake emission estimates. Limnology and Oceanography, 61(S1):S13–S26, November 2016. CODEN LIOCAH. ISSN 0024-3590.

Nickels:2019:EPF

[NSO19] Catherine F. Nickels, Linsey M. Sala, and Mark D. Ohman. The euphausiid prey field for blue whales around a steep bathymetric feature in the southern California current system. Limnology and Oceanography, 64(1):390–405, January 2019. CODEN LI-OCAH. ISSN 0024-3590.

Nuester:2014:RHB

[NSV^{+}14] Jochen Nuester, Steve Shema, Alexander Vermont, David M. Fields, and Benjamin S. Twining. The regeneration of highly bioavailable iron by meso- and microzooplankton. Limnology and Oceanography, 59(4):1399–1409, July 2014. CODEN LI-OCAH. ISSN 0024-3590.

Nakov:2014:UPM

[NTA14] Teofil Nakov, Edward C. Theriot, and Andrew J. Alverson. Using phylogeny to model cell size evolution in marine and freshwater diatoms. Limnology and Oceanography, 59(1):79–86, January 2014. CODEN LIOCAH. ISSN 0024-3590.

Nishibe:2015:DDA

[NTI^{+}15] Yuichiro Nishibe, Kazutaka Takahashi, Tadafumi Ichikawa, Kiyotaka Hidaka, Hiroaki Kurogi, Kyohei Segawa, and Hiroaki Saito. Degradation of discarded appendicularian houses by oncaeid copepods. Limnology and Oceanography, 60(3):967–976, May 2015. CODEN LIOCAH. ISSN 0024-3590.

Na:2018:PTD

[NTK^{+}18] Taehee Na, Bo Thamdrup, Bomina Kim, Sung-Han Kim, Verona Vandieken, Dong-Jin Kang, and Jung-Ho Hyun. N_{2} production through denitrification and anammox across the continental margin (shelf–slope–rise) of the Ulleung Basin, East Sea.

Limnology and Oceanography, 63(S1):S410–S424, March 2018. CODEN LIOCAH. ISSN 0024-3590.

Nunn:2010:PPU

[NTM^{+}10] Brook L. Nunn, Ying S. Ting, Lars Malmström, Yihsuan S. Tsai, Angela Squier, David R. Goodlett, and H. Rodger Harvey. The path to preservation: Using proteomics to decipher the fate of diatom proteins during microbial degradation. Limnology and Oceanography, 55(4):1790–1804, July 2010. CODEN LIOCAH. ISSN 0024-3590.

Nakagawa:2012:SCM

[NUH^{+}12] Mayuko Nakagawa, Yuichiro Ueno, Shohei Hattori, Maki Umemura, Akihiko Yagi, Ken Takai, Keisuke Koba, Yuji Sasaki, Akiko Makabe, and Naohiro Yoshida. Seasonal change in microbial sulfur cycling in monomictic Lake Fukami-ike, Japan. Limnology and Oceanography, 57(4):974–988, July 2012. CO-DEN LIOCAH. ISSN 0024-3590.

Orellana:2012:RBC

[nVOH12] Mó nica V. Orellana and Dennis A. Hansell. Ribulose-1,5-bisphosphate carboxylase/oxygenase (RuBisCO): a long-lived protein in the deep ocean. Limnology and Oceanography, 57 (3):826–834, May 2012. CODEN LIOCAH. ISSN 0024-3590.

Nydahl:2019:COM

[NWT^{+}19] Anna C. Nydahl, Marcus B. Wallin, Lars J. Tranvik, Carolin Hiller, Katrin Attermeyer, Julie A. Garrison, Fernando Chaguaceda, Kristin Scharnweber, and Gesa A. Weyhenmeyer. Colored organic matter increases CO_{2} in meso-eutrophic lake water through altered light climate and acidity. Limnology and Oceanography, 64(2):744–756, March 2019. CODEN LIOCAH. ISSN 0024-3590.

Niu:2018:HTE

[NXL^{+}18] Qianru Niu, Meng Xia, Stuart A. Ludsin, Philip Y. Chu, Doran M. Mason, and Edward S. Rutherford. High-turbidity events in Western Lake Erie during ice-free cycles: Contributions of river-loaded vs. resuspended sediments. Limnology and Oceanography, 63(6):2545–2562, November 2018. CODEN LI-OCAH. ISSN 0024-3590.

Nie:2011:WBH

[NZH^{+}11] Xiang-Ping Nie, Jenny Zie, Norbert Häubner, Bo Tallmark, and Pauline Snoeijs. Why Baltic herring and sprat are weak conduits for astaxanthin from zooplankton to piscivorous fish. Limnology and Oceanography, 57(4):1155–1167, May 2011. CODEN LIOCAH. ISSN 0024-3590.

Orchard:2010:DIO

[OALD10] Elizabeth D. Orchard, James W. Ammerman, Michael W. Lomas, and Sonya T. Dyhrman. Dissolved inorganic and organic phosphorus uptake in Trichodesmium and the microbial community: The importance of phosphorus ester in the Sargasso Sea. Limnology and Oceanography, 55(3):1390–1399, May 2010. CODEN LIOCAH. ISSN 0024-3590.

Oveisy:2012:TDS

[OBI12] A. Oveisy, L. Boegman, and J. Imberger. Three-dimensional simulation of lake and ice dynamics during winter. Limnology and Oceanography, 57(1):43–57, January 2012. CODEN LI-OCAH. ISSN 0024-3590.

Oehler:2019:NDS

[OBL^{+}19] Till Oehler, Hendra Bakti, Rachmat Fajar Lubis, Ananta Purwoarminta, Robert Delinom, and Nils Moosdorf. Nutrient dynamics in submarine groundwater discharge through a coral reef (western Lombok, Indonesia). Limnology and Oceanography, 64 (6):2646–2661, November 2019. CODEN LIOCAH. ISSN 0024-3590.

Oakes:2011:CSU

[OBM^{+}11] Joanne M. Oakes, Melissa D. Bautista, Damien Maher, W. Brian Jones, and Bradley D. Eyre. Carbon self-utilization may assist Caulerpa taxifolia invasion. Limnology and Oceanography, 56(5):1824–1831, September 2011. CODEN LIOCAH. ISSN 0024-3590.

Orchard:2010:PTL

[OBNP^{+}10] Elizabeth D. Orchard, Claudia R. Benitez-Nelson, Perry J. Pellechia, Michael W. Lomas, and Sonya T. Dyhrman. Polyphosphate in Trichodesmium from the low-phosphorus Sargasso Sea. Limnology and Oceanography, 55(5):2161–2169, September 2010. CODEN LIOCAH. ISSN 0024-3590.

Ojala:2011:CGF

[OBT^{+}11] Anne Ojala, Jessica López Bellido, Tiina Tulonen, Paula Kankaala, and Jussi Huotari. Carbon gas fluxes from a brownwater and a clear-water lake in the boreal zone during a summer with extreme rain events. Limnology and Oceanography, 56(1): 61–76, January 2011. CODEN LIOCAH. ISSN 0024-3590.

Ortega:2018:HMF

[OCB^{+}18] Sonia Herrero Ortega, Núria Catalán, Erik Björn, Hannes Gröntoft, Torfi Geir Hilmarsson, Stefan Bertilsson, Pianpian Wu, Kevin Bishop, Oded Levanoni, and Andrea G. Bravo. High methylmercury formation in ponds fueled by fresh humic and algal derived organic matter. Limnology and Oceanography, 63 (S1):S44–S53, March 2018. CODEN LIOCAH. ISSN 0024-3590.

Obernosterer:2011:DBG

[OCLW11] Ingrid Obernosterer, Philippe Catala, Philippe Lebaron, and Nyree J. West. Distinct bacterial groups contribute to carbon cycling during a naturally iron fertilized phytoplankton bloom in the Southern Ocean. Limnology and Oceanography, 56(6):2391– 2401, November 2011. CODEN LIOCAH. ISSN 0024-3590.

Oakes:2010:IEM

[OCR10] Joanne M. Oakes, Rod M. Connolly, and Andrew T. Revill. Isotope enrichment in mangrove forests separates microphytobenthos and detritus as carbon sources for animals. Limnology and Oceanography, 55(1):393–402, January 2010. CODEN LIOCAH. ISSN 0024-3590.

Oakes:2012:TFM

[OEM12] Joanne M. Oakes, Bradley D. Eyre, and Jack J. Middelburg. Transformation and fate of microphytobenthos carbon in subtropical shallow subtidal sands: a^{13}C-labeling study. Limnology and Oceanography, 57(6):1846–1856, November 2012. CODEN LIOCAH. ISSN 0024-3590.

Oakes:2010:CPL

[OEMB10] Joanne M. Oakes, Bradley D. Eyre, Jack J. Middelburg, and Henricus T. S. Boschker. Composition, production, and loss of carbohydrates in subtropical shallow subtidal sandy sediments: Rapid processing and long-term retention revealed by 13C-labeling. Limnology and Oceanography, 55(5):2126–2138, September 2010. CODEN LIOCAH. ISSN 0024-3590.

Omand:2012:EVN

[OFGF12] M. M. Omand, F. Feddersen, R. T. Guza, and P. J. S. Franks. Episodic vertical nutrient fluxes and nearshore phytoplankton blooms in Southern California. Limnology and Oceanography, 57 (6):1673–1688, November 2012. CODEN LIOCAH. ISSN 0024-3590.

OBrien:2012:NTT

[OHKC^{+}12] Jonathan M. O’Brien, Stephen K. Hamilton, Lauren E. Kinsman-Costello, Jay T. Lennon, and Nathaniel E. Ostrom. Nitrogen transformations in a through-flow wetland revealed using whole-ecosystem pulsed ^{15}N additions. Limnology and Oceanography, 57(1):221–234, January 2012. CODEN LIOCAH. ISSN 0024-3590.

Okely:2010:PDD

[OIS10] Patricia Okely, Jörg Imberger, and Kenji Shimizu. Particle dispersal due to interplay of motions in the surface layer of a small reservoir. Limnology and Oceanography, 55(2):589–603, March 2010. CODEN LIOCAH. ISSN 0024-3590. See comment [PHJ12].

Ouyang:2018:UIL

[OLC18] Xiaoguang Ouyang, Shing Yip Lee, and Rod M. Connolly. Using isotope labeling to partition sources of CO_{2} efflux in newly established mangrove seedlings. Limnology and Oceanography, 63 (2):731–740, March 2018. CODEN LIOCAH. ISSN 0024-3590.

Omand:2011:PBP

[OLF^{+}11] Melissa M. Omand, James J. Leichter, Peter J. S. Franks, R. T. Guza, Andrew J. Lucas, and Falk Feddersen. Physical and biological processes underlying the sudden surface appearance of a red tide in the nearshore. Limnology and Oceanography, 56(3): 787–801, May 2011. CODEN LIOCAH. ISSN 0024-3590.

Oswald:2016:AGM

[OMB^{+}16] Kirsten Oswald, Jana Milucka, Andreas Brand, Philipp Hach, Sten Littmann, Bernhard Wehrli, Marcel M. M. Kuypers, and Carsten J. Schubert. Aerobic gammaproteobacterial methanotrophs mitigate methane emissions from oxic and anoxic lake waters. Limnology and Oceanography, 61(S1):S101–S118, November 2016. CODEN LIOCAH. ISSN 0024-3590.

Ochoa:2013:DLC

[OMSC13] José Ochoa, H. Maske, J. Sheinbaum, and J. Candela. Diel and lunar cycles of vertical migration extending to below 1000 m in the ocean and the vertical connectivity of depth-tiered populations. Limnology and Oceanography, 58(4):1207–1214, July 2013. CODEN LIOCAH. ISSN 0024-3590.

Olli:2014:ATR

[OPA^{+}14] Kalle Olli, Robert Ptacnik, Tom Andersen, Olga Trikk, Riina Klais, Sirpa Lehtinen, and Timo Tamminen. Against the tide: Recent diversity increase enhances resource use in a coastal ecosystem. Limnology and Oceanography, 59(1):267–274, January 2014. CODEN LIOCAH. ISSN 0024-3590.

Ochs:2013:DBN

[OPZ13] Clifford A. Ochs, Orathai Pongruktham, and Paul V. Zimba. Darkness at the break of noon: Phytoplankton production in the Lower Mississippi River. Limnology and Oceanography, 58 (2):555–568, March 2013. CODEN LIOCAH. ISSN 0024-3590.

Ohman:2016:NEB

[OR16] Mark D. Ohman and Jean-Baptiste Romagnan. Nonlinear effects of body size and optical attenuation on diel vertical migration by zooplankton. Limnology and Oceanography, 61(2): 765–770, March 2016. CODEN LIOCAH. ISSN 0024-3590.

Ohnemus:2017:ETM

[ORC^{+}17] Daniel C. Ohnemus, Sara Rauschenberg, Gregory A. Cutter, Jessica N. Fitzsimmons, Robert M. Sherrell, and Benjamin S. Twining. Elevated trace metal content of prokaryotic communities associated with marine oxygen deficient zones. Limnology and Oceanography, 62(1):3–25, January 2017. CODEN LI-OCAH. ISSN 0024-3590.

Oakes:2016:TFS

[ORGE16] Joanne M. Oakes, Søren Rysgaard, Ronnie N. Glud, and Bradley D. Eyre. The transformation and fate of sub-Arctic microphytobenthos carbon revealed through ^{13}C-labeling. Limnology and Oceanography, 61(6):2296–2308, November 2016. CO-DEN LIOCAH. ISSN 0024-3590.

Okely:2010:PAH

[OrIA10] Patricia Okely, Jö rg Imberger, and Jason P. Antenucci. Processes affecting horizontal mixing and dispersion in Winam Gulf, Lake Victoria. Limnology and Oceanography, 55(5):1865–1880, September 2010. CODEN LIOCAH. ISSN 0024-3590. See comment [PHJ12].

Orihel:2015:NPI

[OSB^{+}15] Diane M. Orihel, David W. Schindler, Nathaniel C. Ballard, Mark D. Graham, David W. O’Connell, Lindsey R. Wilson, and Rolf D. Vinebrooke. The “nutrient pump:” iron-poor sediments fuel low nitrogen-to-phosphorus ratios and cyanobacterial blooms in polymictic lakes. Limnology and Oceanography, 60(3): 856–871, May 2015. CODEN LIOCAH. ISSN 0024-3590.

Obrador:2014:VPM

[OSC14] Biel Obrador, Peter A. Staehr, and Jesper P. C. Christensen. Vertical patterns of metabolism in three contrasting stratified lakes. Limnology and Oceanography, 59(4):1228–1240, July 2014. CODEN LIOCAH. ISSN 0024-3590.

Outridge:2019:ASM

[OSHS19] P. M. Outridge, G. A. Stern, P. B. Hamilton, and H. Sanei. Algal scavenging of mercury in preindustrial Arctic lakes. Limnology and Oceanography, 64(4):1558–1571, July 2019. CODEN LIOCAH. ISSN 0024-3590.

Oviedo-Vargas:2013:DOC

[OVRJ13] Diana Oviedo-Vargas, Todd V. Royer, and Laura T. Johnson. Dissolved organic carbon manipulation reveals coupled cycling of carbon, nitrogen, and phosphorus in a nitrogen-rich stream. Limnology and Oceanography, 58(4):1196–1206, July 2013. CO-DEN LIOCAH. ISSN 0024-3590.

Osburn:2011:DOM

[OWFS11] Christopher L. Osburn, Courtney R. Wigdahl, Sherilyn C. Fritz, and Jasmine E. Saros. Dissolved organic matter composition and photoreactivity in prairie lakes of the U.S. Great Plains. Limnology and Oceanography, 56(6):2371–2390, November 2011. CODEN LIOCAH. ISSN 0024-3590.

Oreska:2018:NSC

[OWM^{+}18] Matthew P. J. Oreska, Grace M. Wilkinson, Karen J. McGlathery, Molly Bost, and Brent A. McKee. Non-seagrass carbon contributions to seagrass sediment blue carbon. Limnology and Oceanography, 63(S1):S3–S18, March 2018. CODEN LIOCAH. ISSN 0024-3590.

ODonnell:2017:NPC

[OWS^{+}17] Daniel R. O’Donnell, Paul Wilburn, Eugene A. Silow, Lev Y. Yampolsky, and Elena Litchman. Nitrogen and phosphorus colimitation of phytoplankton in Lake Baikal: Insights from a spatial survey and nutrient enrichment experiments. Limnology and Oceanography, 62(4):1383–1392, July 2017. CODEN LIOCAH. ISSN 0024-3590.

Ostrovsky:2010:SFL

[OY10] Ilia Ostrovsky and Yosef Z. Yacobi. Sedimentation flux in a large subtropical lake: Spatiotemporal variations and relation to primary productivity. Limnology and Oceanography, 55(5): 1918–1931, September 2010. CODEN LIOCAH. ISSN 0024-3590.

Paerl:2015:VBE

[PBA^{+}15] R. W. Paerl, E. M. Bertrand, A. E. Allen, B. Palenik, and F. Azam. Vitamin B1 ecophysiology of marine picoeukaryotic algae: Strain-specific differences and a new role for bacteria in vitamin cycling. Limnology and Oceanography, 60(1):215–228, January 2015. CODEN LIOCAH. ISSN 0024-3590.

Pearse:2018:SOC

[PBL^{+}18] Alex L. Pearse, Jan L. Barton, Rebecca E. Lester, Atun Zawadzki, and Peter I. Macreadie. Soil organic carbon variability in Australian temperate freshwater wetlands. Limnology and Oceanography, 63(S1):S254–S266, March 2018. CODEN LI-OCAH. ISSN 0024-3590.

Pinceel:2016:PSS

[PBV16] Tom Pinceel, Luc Brendonck, and Bram Vanschoenwinkel. Propagule size and shape may promote local wind dispersal in freshwater zooplankton — a wind tunnel experiment. Limnology and Oceanography, 61(1):122–131, January 2016. CODEN LIOCAH. ISSN 0024-3590.

Page:2019:SVS

[PCD^{+}19] Heather N. Page, Travis A. Courtney, Eric H. De Carlo, Noah M. Howins, Irina Koester, and Andreas J. Andersson. Spatiotemporal variability in seawater carbon chemistry for a coral reef flat in Kāne’ohe Bay, Hawai’i. Limnology and Oceanography, 64 (3):913–934, May 2019. CODEN LIOCAH. ISSN 0024-3590.

Papiol:2014:RFA

[PCF14] Vanesa Papiol, Joan E. Cartes, and Emanuela Fanelli. The role of food availability in regulating the feeding dynamics and reproductive cycles of bathyal benthopelagic fish in the northwest Mediterranean slope. Limnology and Oceanography, 59(5):1779– 1794, September 2014. CODEN LIOCAH. ISSN 0024-3590.

Pace:2013:ZPE

[PCJK13] Michael L. Pace, Stephen R. Carpenter, Robert A. Johnson, and Jason T. Kurtzweil. Zooplankton provide early warnings of a regime shift in a whole lake manipulation. Limnology and Oceanography, 58(2):525–532, March 2013. CODEN LIOCAH. ISSN 0024-3590.

Pearman:2016:BPC

[PCM^{+}16] John K. Pearman, Laura Casas, Tony Merle, Craig Michell, and Xabier Irigoien. Bacterial and protist community changes during a phytoplankton bloom. Limnology and Oceanography, 61(1): 198–213, January 2016. CODEN LIOCAH. ISSN 0024-3590.

Parr:2015:UCC

[PCO^{+}15] Thomas B. Parr, Christopher S. Cronan, Tsutomu Ohno, Stuart E. G. Findlay, Sean M. C. Smith, and Kevin S. Simon. Urbanization changes the composition and bioavailability of dissolved organic matter in headwater streams. Limnology and Oceanography, 60(3):885–900, May 2015. CODEN LIOCAH. ISSN 0024-3590.

Peng:2018:UMO

[PCPZ18] Dan Peng, Luzhen Chen, Steven C. Pennings, and Yihui Zhang. Using a marsh organ to predict future plant communities in a Chinese estuary invaded by an exotic grass and mangrove. Limnology and Oceanography, 63(6):2595–2605, November 2018. CODEN LIOCAH. ISSN 0024-3590.

Pace:2019:LTS

[PCW19] M. L. Pace, S. R. Carpenter, and G. M. Wilkinson. Long-term studies and reproducibility: Lessons from whole-lake experiments. Limnology and Oceanography, 64(S1):S22–S33, January 2019. CODEN LIOCAH. ISSN 0024-3590.

Poulton:2010:CDN

[PCY^{+}10] Alex J. Poulton, Anastasia Charalampopoulou, Jeremy R. Young, Glen A. Tarran, Mike I. Lucas, and Graham D. Quartly. Coccolithophore dynamics in non-bloom conditions during late summer in the central Iceland Basin (July–August 2007). Limnology and Oceanography, 55(4):1601–1613, July 2010. CODEN LIOCAH. ISSN 0024-3590.

Pollard:2011:UBP

[PD11] Peter C. Pollard and Hugh Ducklow. Ultrahigh bacterial production in a eutrophic subtropical Australian river: Does viral lysis short-circuit the microbial loop? Limnology and Oceanography, 57(4):1115–1129, May 2011. CODEN LIOCAH. ISSN 0024-3590.

Perga:2010:CBT

[PDER10] Marie-Elodie Perga, Marc Desmet, Dirk Enters, and Jean-Louis Reyss. A century of bottom-up- and top-down driven changes on a lake planktonic food web: a paleoecological and paleoisotopic study of Lake Annecy, France. Limnology and Oceanography, 55 (4):803–816, March 2010. CODEN LIOCAH. ISSN 0024-3590.

Passow:2014:AFM

[PDFS14] Uta Passow, Christina L. De La Rocha, Caitlin Fairfield, and Katrin Schmidt. Aggregation as a function of and mineral particles. Limnology and Oceanography, 59(6):532–547, March 2014. CODEN LIOCAH. ISSN 0024-3590.

Pauwels:2010:REP

[PDP^{+}10] Kevin Pauwels, Luc De Meester, Stéphanie Put, Ellen Decaestecker, Ellen Decaestecker, and Robby Stoks. Rapid evolution of phenoloxidase expression, a component of innate immune function, in a natural population of Daphnia magna. Limnology and Oceanography, 55(3):1408–1413, May 2010. CODEN LIOCAH. ISSN 0024-3590.

Peng:2013:SAP

[PE13] Feng Peng and Steven W. Effler. Spectral absorption properties of mineral particles in western Lake Erie: Insights from individual particle analysis. Limnology and Oceanography, 58(5): 1775–1789, September 2013. CODEN LIOCAH. ISSN 0024-3590.

Palinkas:2016:STP

[PE16a] Cindy M. Palinkas and Katharina A. M. Engelhardt. Spatial and temporal patterns of modern (≈ 100 yr) sedimentation in a tidal freshwater marsh: Implications for future sustainability. Limnology and Oceanography, 61(1):132–148, January 2016. CODEN LIOCAH. ISSN 0024-3590.

Peng:2016:ATC

[PE16b] Feng Peng and Steven W. Effler. Advancing two-component partitioning of light scattering in Cayuga Lake, New York. Limnology and Oceanography, 61(1):298–315, January 2016. CODEN LIOCAH. ISSN 0024-3590.

Peng:2017:CCP

[PE17] Feng Peng and Steven W. Effler. Characterizations of calcite particles and evaluations of their light scattering effects in lacustrine systems. Limnology and Oceanography, 62(2):645–664, March 2017. CODEN LIOCAH. ISSN 0024-3590.

Prater:2017:VPC

[PFH^{+}17] Clay Prater, Paul C. Frost, E. Todd Howell, Susan B. Watson, Arthur Zastepa, Sarah S. E. King, Richard J. Vogt, and Marguerite A. Xenopoulos. Variation in particulate C : N : P stoichiometry across the Lake Erie watershed from tributaries to its outflow. Limnology and Oceanography, 62(S1):S194–S206, November 2017. CODEN LIOCAH. ISSN 0024-3590.

Prairie:2010:CPI

[PFJ10] Jennifer C. Prairie, Peter J. S. Franks, and Jules S. Jaffe. Cryptic peaks: invisible vertical structure in fluorescent particles revealed using a planar laser imaging fluorometer. Limnology and Oceanography, 55(5):1943–1958, September 2010. CODEN LI-OCAH. ISSN 0024-3590.

Peng:2018:NUN

[PFvO^{+}18] Xuefeng Peng, Sarah E. Fawcett, Nicolas van Oostende, Martin J. Wolf, Dario Marconi, Daniel M. Sigman, and Bess B.

Ward. Nitrogen uptake and nitrification in the subarctic North Atlantic Ocean. Limnology and Oceanography, 63(4):1462–1487, July 2018. CODEN LIOCAH. ISSN 0024-3590.

Parra:2019:ADZ

[PGB^{+}19] Sabrina M. Parra, Adam T. Greer, Jeffrey W. Book, Alison L. Deary, Inia M. Soto, Carla Culpepper, Frank J. Hernandez, and Travis N. Miles. Acoustic detection of zooplankton diel vertical migration behaviors on the northern Gulf of Mexico shelf. Limnology and Oceanography, 66(4):2092–2113, September 2019. CODEN LIOCAH. ISSN 0024-3590.

Peipoch:2014:IED

[PGP^{+}14] Marc Peipoch, Esperança Gacia, Ada Pastor, Miquel Ribot, Joan Ll. Riera, Francesc Sabater, and Eugènia Martí. Intrinsic and extrinsic drivers of autotrophic nitrogen cycling in stream ecosystems: Results from a translocation experiment. Limnology and Oceanography, 59(6):1973–1986, November 2014. CODEN LIOCAH. ISSN 0024-3590.

Padilla-Gamino:2019:EMR

[PGRR^{+}19] Jacqueline L. Padilla-Gamiño, Melissa S. Roth, Lisa J. Rodrigues, Christina J. Bradley, Robert R. Bidigare, Ruth D. Gates, Celia M. Smith, and Heather L. Spalding. Ecophysiology of mesophotic reef-building corals in Hawai‘ i is influenced by symbiont–host associations, photoacclimatization, trophic plasticity, and adaptation. Limnology and Oceanography, 66(8): 1980–1995, September 2019. CODEN LIOCAH. ISSN 0024-3590.

Parker:2013:DPC

[PH13] Stephanie M. Parker and Alexander D. Huryn. Disturbance and productivity as codeterminants of stream food web complexity in the Arctic. Limnology and Oceanography, 58(6):2158–2170, November 2013. CODEN LIOCAH. ISSN 0024-3590.

Peeters:2015:LSD

[PH15] Frank Peeters and Hilmar Hofmann. Length-scale dependence of horizontal dispersion in the surface water of lakes. Limnology and Oceanography, 60(6):1917–1934, November 2015. CODEN LIOCAH. ISSN 0024-3590.

Petrou:2010:RPS

[PHB^{+}10] Katherina Petrou, Ross Hill, Christopher M. Brown, Douglas A. Campbell, Martina A. Doblin, and Peter J. Ralph. Rapid photoprotection in sea-ice diatoms from the East Antarctic pack ice. Limnology and Oceanography, 55(3):1400–1407, May 2010. CODEN LIOCAH. ISSN 0024-3590.

Penczykowski:2014:HSE

[PHCD14] Rachel M. Penczykowski, Spencer R. Hall, David J. Civitello, and Meghan A. Duffy. Habitat structure and ecological drivers of disease. Limnology and Oceanography, 59(2):340–348, March 2014. CODEN LIOCAH. ISSN 0024-3590.

Paterson:2014:CED

[PHDH14] Gordon Paterson, Craig E. Hebert, Ken G. Drouillard, and G. Doug Haffner. Congruent energy density trends of fish and birds reflect ecosystem change. Limnology and Oceanography, 59 (4):1171–1180, July 2014. CODEN LIOCAH. ISSN 0024-3590.

Poste:2013:PEC

[PHG13] A. E. Poste, R. E. Hecky, and S. J. Guildford. Phosphorus enrichment and carbon depletion contribute to high Microcystis biomass and microcystin concentrations in Ugandan lakes. Limnology and Oceanography, 58(3):1075–1088, May 2013. CODEN LIOCAH. ISSN 0024-3590.

Peeters:2012:CPD

[PHJ12] Frank Peeters, Hilmar Hofmann, and Michael Junk. Comment on “Particle dispersal due to interplay of motions in the surface layer of a small reservoir” (by P. Okely, J. Imberger, and K. Shimizu) and “Processes affecting horizontal mixing and dispersion in Winam Gulf, Lake Victoria” (by P. Okely, J. Imberger, and J. P. Antenucci). Limnology and Oceanography, 57(1):382– 386, January 2012. CODEN LIOCAH. ISSN 0024-3590. See [OIS10, OrIA10].

Piccolroaz:2018:PLS

[PHL^{+}18] S. Piccolroaz, N. C. Healey, J. D. Lenters, S. G. Schladow, S. J. Hook, G. B. Sahoo, and M. Toffolon. On the predictability of lake surface temperature using air temperature in a changing climate: a case study for Lake Tahoe (U.S.A.). Limnology and Oceanography, 63(1):243–261, January 2018. CODEN LIOCAH. ISSN 0024-3590.

Prentice:2019:RWM

[PHLSSS19] Carolyn Prentice, Margot Hessing-Lewis, Rhea Sanders-Smith, and Anne K. Salomon. Reduced water motion enhances organic carbon stocks in temperate eelgrass meadows. Limnology and Oceanography, 64(6):2389–2404, November 2019. CODEN LI-OCAH. ISSN 0024-3590.

Pfeiffer-Herbert:2016:HRS

[PHPH^{+}16] A. S. Pfeiffer-Herbert, F. G. Prahl, B. Hales, J. A. Lerczak, S. D. Pierce, and M. D. Levine. High resolution sampling of methane transport in the Columbia River near-field plume: Implications for sources and sinks in a river-dominated estuary. Limnology and Oceanography, 61(S1):S204–S220, November 2016. CODEN LIOCAH. ISSN 0024-3590.

Piwosz:2019:WDA

[Piw19] Kasia Piwosz. Weekly dynamics of abundance and size structure of specific nanophytoplankton lineages in coastal waters (Baltic Sea). Limnology and Oceanography, 66(4):2172–2186, September 2019. CODEN LIOCAH. ISSN 0024-3590.

Paffenhofer:2016:CPP

[PJ16] Gustav-Adolf Paffenhöfer and Houshuo Jiang. Comment: On phytoplankton perception by calanoid copepods. Limnology and Oceanography, 61(4):1163–1168, July 2016. CODEN LIOCAH. ISSN 0024-3590. See [TSK13, GK15] and reply [KGC^{+}16].

Plass-Johnson:2015:FHK

[PJFJ^{+}15] Jeremiah G. Plass-Johnson, Sebastian C. A. Ferse, Jamaluddin Jompa, Christian Wild, and Mirta Teichberg. Fish herbivory as key ecological function in a heavily degraded coral reef system. Limnology and Oceanography, 60(4):1382–1391, July 2015. CODEN LIOCAH. ISSN 0024-3590.

Peric:2015:LTE

[PJUR15] Mirela Sertić Perić, Christa Jolidon, Urs Uehlinger, and Christopher T. Robinson. Long-term ecological patterns of alpine streams: an imprint of glacial legacies. Limnology and Oceanography, 60(3):992–1007, May 2015. CODEN LIOCAH. ISSN 0024-3590.

Peacock:2014:EAV

[PK14] Melissa B. Peacock and Raphael M. Kudela. Evidence for active vertical migration by two dinoflagellates experiencing iron, nitrogen, and phosphorus limitation. Limnology and Oceanography, 59(3):660–673, May 2014. CODEN LIOCAH. ISSN 0024-3590.

Park:2017:LAR

[PKB^{+}17] Jisoo Park, Fedor I. Kuzminov, Benjamin Bailleul, Eun Jin Yang, SangHoon Lee, Paul G. Falkowski, and Maxim Y. Gorbunov. Light availability rather than Fe controls the magnitude of massive phytoplankton bloom in the Amundsen Sea polynyas, Antarctica. Limnology and Oceanography, 65(9):2260– 2276, September 2017. CODEN LIOCAH. ISSN 0024-3590.

Pusack:2019:POI

[PKWS19] Timothy J. Pusack, David L. Kimbro, J. Wilson White, and Christopher D. Stallings. Predation on oysters is inhibited by intense or chronically mild, low salinity events. Limnology and Oceanography, 64(1):81–92, January 2019. CODEN LIOCAH. ISSN 0024-3590.

Pree:2017:DCM

[PLE^{+}17] Bernadette Pree, Aud Larsen, Jorun Karin Egge, Paolo Simonelli, Rakhesh Madhusoodhanan, Tatiana Margo Tsagaraki, Selina Våge, Svein Rune Erga, Gunnar Bratbak, and T. Frede Thingstad. Dampened copepod-mediated trophic cascades in a microzooplankton-dominated microbial food web: a mesocosm study. Limnology and Oceanography, 63(3):1031–1044, May 2017. CODEN LIOCAH. ISSN 0024-3590.

Perez:2016:RDI

[PLS^{+}16] Claudia A. Pérez, Nelson A. Lagos, Gonzalo S. Saldías, George Waldbusser, and Cristian A. Vargas. Riverine discharges impact physiological traits and carbon sources for shell carbonate in the marine intertidal mussel Perumytilus purpuratus. Limnology and Oceanography, 61(3):969–983, May 2016. CODEN LIOCAH. ISSN 0024-3590.

Puccinelli:2018:FAT

[PMA18] Eleonora Puccinelli, Christopher D. McQuaid, and Isabelle J. Ansorge. Factors affecting trophic compositions of offshore benthic invertebrates at a sub-Antarctic archipelago. Limnology and Oceanography, 64(4):2206–2228, September 2018. CODEN LIOCAH. ISSN 0024-3590.

Paulsen:2019:BTA

[PML^{+}19] Maria Lund Paulsen, Oliver Müller, Aud Larsen, Eva Friis Møller, Mathias Middelboe, Mikael K. Sejr, and Colin Stedmon. Biological transformation of Arctic dissolved organic matter in a NE Greenland fjord. Limnology and Oceanography, 64 (3):1014–1033, May 2019. CODEN LIOCAH. ISSN 0024-3590.

Parnell:2010:RGK

[PMLC^{+}10] P. Ed Parnell, Eric F. Miller, Cleridy E. Lennert-Cody, Paul K. Dayton, Melissa L. Carter, and Timothy D. Stebbins. The response of giant kelp (Macrocystis pyrifera) in southern California to low-frequency climate forcing. Limnology and Oceanography, 55(6):2686–2702, November 2010. CODEN LIOCAH. ISSN 0024-3590.

Paul:2012:SIC

[PMP^{+}12] Rüdiger J. Paul, Ansgar Mertenskötter, Olaf Pinkhaus, Ralph Pirow, Ulrike Gigengack, Ina Buchen, Marita Koch, Wolfgang Horn, and Bettina Zeis. Seasonal and interannual changes in water temperature affect the genetic structure of a Daphnia assemblage (D. longispina complex) through genotype-specific thermal tolerances. Limnology and Oceanography, 57(3):619–633, March 2012. CODEN LIOCAH. ISSN 0024-3590.

Pacheco:2017:WQL

[PMP^{+}17] Felipe S. Pacheco, Marcela Miranda, Luciano P. Pezzi, Arcilan Assireu, Marcelo M. Marinho, Márcio Malafaia, André Reis, Matias Sales, Gilsinéia Correia, Patrícia Domingos, Allan Iwama, Conrado Rudorff, Pedro Oliva, and Jean P. Ometto. Water quality longitudinal profile of the Paraíba do Sul River, Brazil during an extreme drought event. Limnology and Oceanography, 62(S1):S131–S146, November 2017. CODEN LI-OCAH. ISSN 0024-3590.

Peterson:2013:SIV

[PMPD13] Jay O. Peterson, Cheryl A. Morgan, William T. Peterson, and Emanuele Di Lorenzo. Seasonal and interannual variation in the extent of hypoxia in the northern California Current from 1998– 2012. Limnology and Oceanography, 58(6):2279–2292, November 2013. CODEN LIOCAH. ISSN 0024-3590.

Priet-Maheo:2019:MID

[PMRRA19] M. C. Priet-Mahéo, C. L. Ramón, F. J. Rueda, and H. Ó. Andradóttir. Mixing and internal dynamics of a medium-size and deep lake near the Arctic circle. Limnology and Oceanography, 64(1):61–80, January 2019. CODEN LIOCAH. ISSN 0024-3590.

Pain:2019:SSC

[PMY19a] Andrea J. Pain, Jonathan B. Martin, and Caitlin R. Young. Sources and sinks of CO_{2} and CH_{4} in siliciclastic subterranean estuaries. Limnology and Oceanography, 64(4):1500–1514, July 2019. CODEN LIOCAH. ISSN 0024-3590.

Pain:2019:OMQ

[PMY^{+}19b] Andrea J. Pain, Jonathan B. Martin, Caitlin R. Young, Laibin Huang, and Arnoldo Valle-Levinson. Organic matter quantity and quality across salinity gradients in conduit- vs. diffuse flowdominated subterranean estuaries. Limnology and Oceanography, 64(3):1386–1402, May 2019. CODEN LIOCAH. ISSN 0024-3590.

Pena:2019:PRW

[PNR19] M. Angelica Peña, Nina Nemcek, and Marie Robert. Phytoplankton responses to the 2014–2016 warming anomaly in the northeast subarctic Pacific Ocean. Limnology and Oceanography, 64(2):515–525, March 2019. CODEN LIOCAH. ISSN 0024-3590.

Preusse:2010:IWG

[PPL10] Martina Preusse, Frank Peeters, and Andreas Lorke. Internal waves and the generation of turbulence in the thermocline of a large lake. Limnology and Oceanography, 55(6):2353–2365, November 2010. CODEN LIOCAH. ISSN 0024-3590.

Pather:2014:ACR

[PPPA14] Santhiska Pather, Catherine A. Pfister, David M. Post, and Mark A. Altabet. Ammonium cycling in the rocky intertidal: Remineralization, removal, and retention. Limnology and Oceanography, 59(2):361–372, March 2014. CODEN LIOCAH. ISSN 0024-3590.

Parent:2012:NHB

[PPT12] Geneviève J. Parent, Stéphane Plourde, and Julie Turgeon. Natural hybridization between Calanus finmarchicus and C. glacialis (Copepoda) in the Arctic and Northwest Atlantic. Limnology and Oceanography, 57(4):1057–1066, July 2012. CODEN LIOCAH. ISSN 0024-3590.

Pineda:2018:RBL

[PRL18] Jesús Pineda, Nathalie Reyns, and Steven J. Lentz. Reduced barnacle larval abundance and settlement in response to largescale oceanic disturbances: Temporal patterns, nearshore thermal stratification, and potential mechanisms. Limnology and Oceanography, 63(6):2618–2629, November 2018. CODEN LI-OCAH. ISSN 0024-3590.

Panizzo:2018:SDD

[PRS^{+}18] V. N. Panizzo, S. Roberts, G. E. A. Swann, S. McGowan, A. W. Mackay, E. Vologina, V. Pashley, and M. S. A. Horstwood. Spatial differences in dissolved silicon utilization in Lake Baikal, Siberia: Examining the impact of high diatom biomass events and eutrophication. Limnology and Oceanography, 63(4):1562– 1578, July 2018. CODEN LIOCAH. ISSN 0024-3590.

Payet:2013:KKB

[PS13] Jérôme P. Payet and Curtis A. Suttle. To kill or not to kill: The balance between lytic and lysogenic viral infection is driven by trophic status. Limnology and Oceanography, 58(2):465–474, March 2013. CODEN LIOCAH. ISSN 0024-3590.

Princiotta:2017:HMN

[PS17] Sarah DeVaul Princiotta and Robert W. Sanders. Heterotrophic and mixotrophic nanoflagellates in a mesotrophic lake: Abundance and grazing impacts across season and depth. Limnology and Oceanography, 62(2):632–644, March 2017. CODEN LI-OCAH. ISSN 0024-3590.

Podgrajsek:2016:MFS

[PSB^{+}16] E. Podgrajsek, E. Sahlée, D. Bastviken, S. Natchimuthu, N. Kljun, H. E. Chmiel, L. Klemedtsson, and A. Rutgersson. Methane fluxes from a small boreal lake measured with the eddy covariance method. Limnology and Oceanography, 61(S1):S41– S50, November 2016. CODEN LIOCAH. ISSN 0024-3590.

PiresGouvea:2017:IEM

[PSD^{+}17] Lidiane Pires Gouvêa, Nadine Schubert, Cintia Dalcuche Leal Martins, Marina Sissini, Fernanda Ramlov, Eva Regina de Oliveira Rodrigues, Eduardo Oliveira Bastos, Vanessa Carvalho Freire, Marcelo Maraschin, José Carlos Simonassi, Donnangelo A. Varela, Davide Franco, Valeria Cassano, Alessandra Larissa Fonseca, Bonomi Barufi J., and Paulo Antunes Horta. Interactive effects of marine heatwaves and eutrophication on the ecophysiology of a widespread and ecologically important macroalga. Limnology and Oceanography, 65(9):2056– 2075, September 2017. CODEN LIOCAH. ISSN 0024-3590.

Paul:2016:EIC

[PSG^{+}16] Carolin Paul, Ulrich Sommer, Jessica Garzke, Maria Moustaka-Gouni, Allanah Paul, and Birte Matthiessen. Effects of increased CO_{2} concentration on nutrient limited coastal summer plankton depend on temperature. Limnology and Oceanography, 61(3): 853–868, May 2016. CODEN LIOCAH. ISSN 0024-3590.

Paterson:2011:CLS

[PSH^{+}11] M. J. Paterson, D. W. Schindler, R. E. Hecky, D. L. Findlay, and K. J. Rondeau. Comment: Lake 227 shows clearly that controlling inputs of nitrogen will not reduce or prevent eutrophication of lakes. Limnology and Oceanography, 56(4):1545–1547, July 2011. CODEN LIOCAH. ISSN 0024-3590. See response [SM11b].

Piontek:2015:MEC

[PSNE15] Judith Piontek, Martin Sperling, Eva-Maria Nöthig, and Anja Engel. Multiple environmental changes induce interactive effects on bacterial degradation activity in the Arctic Ocean. Limnology and Oceanography, 60(4):1392–1410, July 2015. CODEN LIOCAH. ISSN 0024-3590.

Paolucci:2014:MGV

[PSS^{+}14] Esteban M. Paolucci, Paula Sardiña, Francisco Sylvester, Pablo V. Perepelizin, Aibin Zhan, Sara Ghabooli, Melania E. Cristescu, Marcia D. Oliveira, and Hugh J. MacIsaac. Morphological and genetic variability in an alien invasive mussel across an environmental gradient in South America. Limnology and Oceanography, 59(2):400–412, March 2014. CODEN LIOCAH. ISSN 0024-3590.

Pineda:2013:TSS

[PST^{+}13] Jesús Pineda, Victoria Starczak, Ann Tarrant, Jonathan Blythe, Kristen Davis, Tom Farrar, Michael Berumen, and José C. B. da Silva. Two spatial scales in a bleaching event: Corals from the mildest and the most extreme thermal environments escape mortality. Limnology and Oceanography, 58(5):1531–1545, September 2013. CODEN LIOCAH. ISSN 0024-3590.

Piwosz:2013:SDA

[PSZ^{+}13] Kasia Piwosz, Michaela M. Salcher, Michael Zeder, Anetta Ameryk, and Jakob Pernthaler. Seasonal dynamics and activity of typical freshwater bacteria in brackish waters of the Gulf of Gdańsk. Limnology and Oceanography, 58(3):817–826, May 2013. CODEN LIOCAH. ISSN 0024-3590.

Pond:2011:PTW

[PT11] David W. Pond and Geraint A. Tarling. Phase transitions of wax esters adjust buoyancy in diapausing Calanoides acutus. Limnology and Oceanography, 56(4):1310–1318, July 2011. CODEN LIOCAH. ISSN 0024-3590.

Petrusek:2012:MSV

[PTS12] Adam Petrusek, Anne Thielsch, and Klaus Schwenk. Mitochondrial sequence variation suggests extensive cryptic diversity within the Western Palearctic Daphnia longispina complex. Limnology and Oceanography, 57(6):1838–1845, November 2012. CODEN LIOCAH. ISSN 0024-3590.

Pecseli:2019:FPT

[PTS^{+}19] Hans L. Pécseli, Jan K. Trulsen, Jan Erik Stiansen, Svein Sundby, and Petter Fossum. Feeding of plankton in turbulent oceans and lakes. Limnology and Oceanography, 64(3):1034– 1046, May 2019. CODEN LIOCAH. ISSN 0024-3590.

Prowe:2019:BZF

[PVA^{+}19] A. E. Friederike Prowe, André W. Visser, Ken H. Andersen, Sanae Chiba, and Thomas Kiørboe. Biogeography of zooplankton feeding strategy. Limnology and Oceanography, 64(2):661– 678, March 2019. CODEN LIOCAH. ISSN 0024-3590.

Palmer:2013:LNC

[PvDM^{+}13] Molly A. Palmer, Gert L. van Dijken, B. Greg Mitchell, Brian J. Seegers, Kate E. Lowry, Matthew M. Mills, and Kevin R. Arrigo. Light and nutrient control of photosynthesis in natural phytoplankton populations from the Chukchi and Beaufort Seas, Arctic Ocean. Limnology and Oceanography, 58(6):2185–2205, November 2013. CODEN LIOCAH. ISSN 0024-3590.

Pajk:2012:ICF

[PvEF12] Franja Pajk, Eric von Elert, and Patrick Fink. Interaction of changes in food quality and temperature reveals maternal effects on fitness parameters of a keystone aquatic herbivore. Limnology and Oceanography, 57(1):281–292, January 2012. CODEN LIOCAH. ISSN 0024-3590.

Parra:2016:SVS

[PVLMT^{+}16] Sabrina M. Parra, Arnoldo Valle-Levinson, Ismael Mariño-Tapia, Cecilia Enriquez, Julio Candela, and Julio Sheinbaum. Seasonal variability of saltwater intrusion at a point-source submarine groundwater discharge. Limnology and Oceanography, 61 (4):1245–1258, July 2016. CODEN LIOCAH. ISSN 0024-3590.

Prater:2016:ECP

[PWF16] Clay Prater, Nicole D. Wagner, and Paul C. Frost. Effects of calcium and phosphorus limitation on the nutritional ecophysiology of Daphnia. Limnology and Oceanography, 61(1):268–278, January 2016. CODEN LIOCAH. ISSN 0024-3590.

Prater:2018:SEF

[PWF18] Clay Prater, Nicole D. Wagner, and Paul C. Frost. Seasonal effects of food quality and temperature on body stoichiometry, biochemistry, and biomass production in Daphnia populations. Limnology and Oceanography, 63(4):1727–1740, July 2018. CO-DEN LIOCAH. ISSN 0024-3590.

Pitcher:2011:CTW

[PWS^{+}11] Angela Pitcher, Cornelia Wuchter, Kathi Siedenberg, Stefan Schouten, and Jaap S. Sinninghe Damsté. Crenarchaeol tracks winter blooms of ammonia-oxidizing Thaumarchaeota in the coastal North Sea. Limnology and Oceanography, 56(6):2308– 2318, November 2011. CODEN LIOCAH. ISSN 0024-3590.

Pettitt-Wade:2018:INB

[PWWF18] Harri Pettitt-Wade, Kyle W. Wellband, and Aaron T. Fisk. Inconsistency for the niche breadth invasion success hypothesis in aquatic invertebrates. Limnology and Oceanography, 63(1): 144–159, January 2018. CODEN LIOCAH. ISSN 0024-3590.

Pajk:2018:TRN

[PZHD18] Franja Pajk, Jiexiang Zhang, Bo-Ping Han, and Henri J. Dumont. Thermal reaction norms of a subtropical and a tropical species of Diaphanosoma (cladocera) explain their distribution. Limnology and Oceanography, 63(3):1204–1220, May 2018. CO-DEN LIOCAH. ISSN 0024-3590.

Qu:2018:RLC

[QFH18] Pingping Qu, Feixue Fu, and David A. Hutchins. Responses of the large centric diatom Coscinodiscus sp. to interactions between warming, elevated CO_{2}, and nitrate availability. Limnology and Oceanography, 63(3):1407–1424, May 2018. CODEN LIOCAH. ISSN 0024-3590.

Qi:2018:DCC

[QHVM18] Lin Qi, Chuanmin Hu, Petra M. Visser, and Ronghua Ma. Diurnal changes of cyanobacteria blooms in Taihu Lake as derived from GOCI observations. Limnology and Oceanography, 63(4): 1711–1726, July 2018. CODEN LIOCAH. ISSN 0024-3590.

Qin:2019:PCG

[QS19] Qubin Qin and Jian Shen. Pelagic contribution to gross primary production dynamics in shallow areas of York River, VA, U.S.A. Limnology and Oceanography, 64(4):1484–1499, July 2019. CO-DEN LIOCAH. ISSN 0024-3590.

QuinonesRivera:2010:EBP

[QWRJ10] Zoraida J. Quiñones Rivera, Björn Wissel, Nancy N. Rabalais, and Dubravko Justić. Effects of biological and physical factors on seasonal oxygen dynamics in a stratified, eutrophic coastal ecosystem. Limnology and Oceanography, 55(1):289–304, January 2010. CODEN LIOCAH. ISSN 0024-3590.

Rovelli:2017:RSR

[RAB^{+}17] Lorenzo Rovelli, Karl M. Attard, Andrew Binley, Catherine M. Heppell, Henrik Stahl, Mark Trimmer, and Ronnie N. Glud. Reach-scale river metabolism across contrasting sub-catchment geologies: Effect of light and hydrology. Limnology and Oceanography, 62(S1):S381–S399, November 2017. CODEN LIOCAH. ISSN 0024-3590.

Rimmer:2005:CST

[RAKE05] Alon Rimmer, Yasuaki Aota, Michio Kumagai, and Werner Eckert. Chemical stratification in thermally stratified lakes: a chloride mass balance model. Limnology and Oceanography, 50(1): 147–157, January 2005. CODEN LIOCAH. ISSN 0024-3590. See erratum [CR10].

Rampen:2010:CSS

[RASD10] Sebastiaan W. Rampen, Ben A. Abbas, Stefan Schouten, and Jaap S. Sinninghe Damsté. A comprehensive study of sterols in marine diatoms (Bacillariophyta): Implications for their use as tracers for diatom productivity. Limnology and Oceanography, 55(1):91–105, January 2010. CODEN LIOCAH. ISSN 0024-3590.

Reef:2017:UED

[RASV^{+}17] Ruth Reef, Trisha B. Atwood, Jimena Samper-Villarreal, Maria Fernanda Adame, Eugenia M. Sampayo, and Catherine E. Lovelock. Using eDNA to determine the source of organic carbon in seagrass meadows. Limnology and Oceanography, 63(3): 1254–1265, May 2017. CODEN LIOCAH. ISSN 0024-3590.

Rowe:2017:IIQ

[RAV^{+}17] Mark D. Rowe, Eric J. Anderson, Henry A. Vanderploeg, Steven A. Pothoven, Ashley K. Elgin, Jia Wang, and Foad Yousef. Influence of invasive quagga mussels, phosphorus loads, and climate on spatial and temporal patterns of productivity in Lake Michigan: a biophysical modeling study. Limnology and Oceanography, 62(6):2629–2649, November 2017. CODEN LI-OCAH. ISSN 0024-3590.

Rembauville:2016:PMS

[RBCS16] M. Rembauville, S. Blain, J. Caparros, and I. Salter. Particulate matter stoichiometry driven by microplankton community structure in summer in the Indian sector of the Southern Ocean. Limnology and Oceanography, 61(4):1301–1321, July 2016. CO-DEN LIOCAH. ISSN 0024-3590.

Roberts:2018:TMS

[RBD18] Emyr Martyn Roberts, David George Bowers, and Andrew John Davies. Tidal modulation of seabed light and its implications for benthic algae. Limnology and Oceanography, 63(1):91–106, January 2018. CODEN LIOCAH. ISSN 0024-3590.

Raitsos:2010:GCC

[RBG^{+}10] Dionysios E. Raitsos, Gregory Beaugrand, Dimitrios Georgopoulos, Argyro Zenetos, Antonietta M. Pancucci-Papadopoulou, Alexander Theocharis, and Evangelos Papathanassiou. Global climate change amplifies the entry of tropical species into the eastern Mediterranean Sea. Limnology and Oceanography, 55 (4):1478–1484, July 2010. CODEN LIOCAH. ISSN 0024-3590.

Rombouts:2010:MAL

[RBI^{+}10] Isabelle Rombouts, Grégory Beaugrand, Frédéric Ibañez, Stéphane Gasparini, Sanae Chiba, and Louis Legendre. A multivariate approach to large-scale variation in marine planktonic copepod diversity and its environmental correlates. Limnology and Oceanography, 55(5):2219–2229, September 2010. CODEN LIOCAH. ISSN 0024-3590.

Rheuban:2014:EMA

[RBM14] Jennie E. Rheuban, Peter Berg, and Karen J. McGlathery. Ecosystem metabolism along a colonization gradient of eelgrass (Zostera marina) measured by eddy correlation. Limnology and Oceanography, 59(4):1376–1387, July 2014. CODEN LIOCAH. ISSN 0024-3590.

Rivest:2010:NSP

[RBRH10] Emily B. Rivest, David M. Baker, Krystal L. Rypien, and C. Drew Harvell. Nitrogen source preference of Aspergillus sydowii, an infective agent associated with aspergillosis of sea fan corals. Limnology and Oceanography, 55(1):386–392, January 2010. CODEN LIOCAH. ISSN 0024-3590.

Ran:2017:NTT

[RBY^{+}17] Xiangbin Ran, Lex Bouwman, Zhigang Yu, Arthur Beusen, Hongtao Chen, and Qingzhen Yao. Nitrogen transport, transformation, and retention in the Three Gorges Reservoir: a mass balance approach. Limnology and Oceanography, 65(9):2323– 2337, September 2017. CODEN LIOCAH. ISSN 0024-3590.

Reed:2015:PCR

[RCH^{+}15] Daniel C. Reed, Craig A. Carlson, Elisa R. Halewood, J. Clinton Nelson, Shannon L. Harrer, Andrew Rassweiler, and Robert J. Miller. Patterns and controls of reef-scale production of dissolved organic carbon by giant kelp Macrocystis pyrifera. Limnology and Oceanography, 60(6):1996–2008, November 2015. CODEN LIOCAH. ISSN 0024-3590.

Rigosi:2014:IBC

[RCIB14] Anna Rigosi, Cayelan C. Carey, Bas W. Ibelings, and Justin D. Brookes. The interaction between climate warming and eutrophication to promote cyanobacteria is dependent on trophic state and varies among taxa. Limnology and Oceanography, 59(1): 99–114, January 2014. CODEN LIOCAH. ISSN 0024-3590.

Rothausler:2015:ADT

[RCJ15] Eva Rothäusler, Hanna Corell, and Veijo Jormalainen. Abundance and dispersal trajectories of floating Fucus vesiculosus in the Northern Baltic Sea. Limnology and Oceanography, 60(6): 2173–2184, November 2015. CODEN LIOCAH. ISSN 0024-3590.

Romera-Castillo:2010:EPC

[RCSÁS^{+}10] Cristina Romera-Castillo, Hugo Sarmento, Xosé Antón Álvarez-Salgado, Josep M. Gasol, and Celia Marrasé. Erratum: Production of chromophoric dissolved organic matter by marine phytoplankton. Limnology and Oceanography, 50(1):1466, May 2010. CODEN LIOCAH. ISSN 0024-3590.

Rohrs:2014:WIT

[RCV^{+}14] Johannes Röhrs, Kai Håkon Christensen, Frode Vikebø, Svein Sundby, Øyvind Saetra, and Göran Broström. Wave-induced transport and vertical mixing of pelagic eggs and larvae. Limnology and Oceanography, 59(4):1213–1227, July 2014. CODEN LIOCAH. ISSN 0024-3590.

Rii:2016:DPP

[RDB^{+}16] Yoshimi M. Rii, Solange Duhamel, Robert R. Bidigare, David M. Karl, Daniel J. Repeta, and Matthew J. Church. Diversity and productivity of photosynthetic picoeukaryotes in biogeochemically distinct regions of the South East Pacific Ocean. Limnology and Oceanography, 61(3):806–824, May 2016. CO-DEN LIOCAH. ISSN 0024-3590.

Roland:2018:DAA

[RDB^{+}18] Fleur A. E. Roland, François Darchambeau, Alberto V. Borges, Cédric Morana, Loreto De Brabandere, Bo Thamdrup, and Sean A. Crowe. Denitrification, anaerobic ammonium oxidation, and dissimilatory nitrate reduction to ammonium in an East African Great Lake (Lake Kivu). Limnology and Oceanography, 63(2):687–701, March 2018. CODEN LIOCAH. ISSN 0024-3590.

Reid:2019:IWI

[RDC^{+}19] Emma C. Reid, Thomas M. DeCarlo, Anne L. Cohen, George T. F. Wong, Steven J. Lentz, Aryan Safaie, Austin Hall, and Kristen A. Davis. Internal waves influence the thermal and nutrient environment on a shallow coral reef. Limnology and Oceanography, 66(8):1949–1965, September 2019. CODEN LI-OCAH. ISSN 0024-3590.

Richardson:2017:SGD

[RDP^{+}17] Christina M. Richardson, Henrietta Dulai, Brian N. Popp, Kathleen Ruttenberg, and Joseph K. Fackrell. Submarine groundwater discharge drives biogeochemistry in two Hawaiian reefs. Limnology and Oceanography, 62(S1):S348–S363, November 2017. CODEN LIOCAH. ISSN 0024-3590.

Rottgers:2014:MSL

[RDT^{+}14] Rüdiger Röttgers, Cecile Dupouy, Bettina B. Taylor, Astrid Bracher, and Slawomir B. Woźniak. Mass-specific light absorption coefficients of natural aquatic particles in the near-infrared spectral region. Limnology and Oceanography, 59(5):1449–1460, September 2014. CODEN LIOCAH. ISSN 0024-3590.

Rosman:2013:IWC

[RDZ^{+}13] Johanna H. Rosman, Mark W. Denny, Robert B. Zeller, Stephen G. Monismith, and Jeffrey R. Koseff. Interaction of waves and currents with kelp forests (Macrocystis pyrifera): Insights from a dynamically scaled laboratory model. Limnology and Oceanography, 58(3):790–802, May 2013. CODEN LI-OCAH. ISSN 0024-3590.

Romito:2010:ESR

[REDW10] Angela M. Romito, Susan L. Eggert, Jeffrey M. Diez, and J. Bruce Wallace. Effects of seasonality and resource limitation on organic matter turnover by Chironomidae (Diptera) in southern Appalachian headwater streams. Limnology and Oceanography, 55(3):1083–1092, May 2010. CODEN LIOCAH. ISSN 0024-3590.

Roberts:2012:HES

[REE^{+}12] Keryn L. Roberts, Vera M. Eate, Bradley D. Eyre, Daryl P. Holland, and Perran L. M. Cook. Hypoxic events stimulate nitrogen recycling in a shallow salt-wedge estuary: The Yarra River estuary, Australia. Limnology and Oceanography, 58(1): 1427–1442, September 2012. CODEN LIOCAH. ISSN 0024-

Rooze:2016:IDA

[RETS16] Jurjen Rooze, Matthias Egger, Iana Tsandev, and Caroline P. Slomp. Iron-dependent anaerobic oxidation of methane in coastal surface sediments: Potential controls and impact. Limnology and Oceanography, 61(S1):S267–S282, November 2016. CODEN LIOCAH. ISSN 0024-3590.

Renz:2013:SWD

[RF13] Judith R. Renz and Stefan Forster. Are similar worms different? A comparative tracer study on bioturbation in the three sibling species Marenzelleria arctia, M. viridis, and M. neglecta from the Baltic Sea. Limnology and Oceanography, 58(6):2046–2058, November 2013. CODEN LIOCAH. ISSN 0024-3590.

Robinson:2013:LTC

[RG13] Kelly L. Robinson and William M. Graham. Long-term change in the abundances of northern Gulf of Mexico scyphomedusae Chrysaora sp. and Aurelia spp. with links to climate variability. Limnology and Oceanography, 58(2):235–253, January 2013. CODEN LIOCAH. ISSN 0024-3590.

Riessen:2019:DDP

[RG19] Howard P. Riessen and John J. Gilbert. Divergent developmental patterns of induced morphological defenses in rotifers and Daphnia: Ecological and evolutionary context. Limnology and Oceanography, 64(2):541–557, March 2019. CODEN LIOCAH. ISSN 0024-3590.

Romero:2019:MBF

[RGB^{+}19] Estela Romero, Josette Garnier, Gilles Billen, Antsiva Ramarson, Philippe Riou, and Romain Le Gendre. Modeling the biogeochemical functioning of the Seine estuary and its coastal zone: Export, retention, and transformations. Limnology and Oceanography, 64(3):895–912, May 2019. CODEN LIOCAH. ISSN 0024-3590.

Rose:2010:DNS

[RGG^{+}10] Andrew L. Rose, Aurélie Godrant, Aurélie Godrant, Miles Furnas, and T. David Waite. Dynamics of nonphotochemical superoxide production in the Great Barrier Reef lagoon. Limnology and Oceanography, 55(4):1521–1536, July 2010. CODEN LI-OCAH. ISSN 0024-3590.

Ruiz-Gonzalez:2012:AVL

[RGGL^{+}12] Clara Ruiz-González, Martí Galí, Thomas Lefort, Clara Cardelús, Rafel Simó, and Josep M. Gasol. Annual variability in light modulation of bacterial heterotrophic activity in surface northwestern Mediterranean waters. Limnology and Oceanography, 58(1):1376–1388, September 2012. CODEN LIOCAH. ISSN 0024-3590.

Ruiz-Gonzalez:2012:DCB

[RGLM^{+}12] Clara Ruiz-González, Thomas Lefort, Ramon Massana, Rafel Simó, and Josep M. Gasol. Diel changes in bulk and single-cell bacterial heterotrophic activity in winter surface waters of the northwestern Mediterranean Sea. Limnology and Oceanography, 57(1):29–42, January 2012. CODEN LIOCAH. ISSN 0024-3590.

Ryan:2011:HPE

[RGM^{+}11] J. Ryan, D. Greenfield, R. Marin III, C. Preston, B. Roman, S. Jensen, D. Pargett, J. Birch, C. Mikulski, G. Doucette, and C. Scholin. Harmful phytoplankton ecology studies using an autonomous molecular analytical and ocean observing network. Limnology and Oceanography, 56(4):1255–1272, July 2011. CO-DEN LIOCAH. ISSN 0024-3590.

Roper:2015:ISG

[RGM15] Tania Röper, Janek Greskowiak, and Gudrun Massmann. Instabilities of submarine groundwater discharge under tidal forcing. Limnology and Oceanography, 60(1):22–28, January 2015. CO-DEN LIOCAH. ISSN 0024-3590.

Rimmer:2011:MLT

[RGO^{+}11] Alon Rimmer, Gideon Gal, Tamar Opher, Yury Lechinsky, and Yosef Z. Yacobi. Mechanisms of long-term variations in the thermal structure of a warm lake. Limnology and Oceanography, 57(4):974–988, May 2011. CODEN LIOCAH. ISSN 0024-3590.

Ruiz-Halpern:2011:AKS

[RHDTS^{+}11] Sergio Ruiz-Halpern, Carlos M. Duarte, Antonio Tovar-Sanchez, Marcos Pastor, Burkhard Horstkotte, Sebastien Lasternas, and Susana Agustí. Antarctic krill as a source of dissolved organic carbon to the Antarctic ecosystem. Limnology and Oceanography, 56(2):521–528, March 2011. CODEN LIOCAH. ISSN 0024-3590.

Ruiz-Halpern:2015:HCE

[RHMSE15] Sergio Ruiz-Halpern, Damien T. Maher, Isaac R. Santos, and Bradley D. Eyre. High CO_{2} evasion during floods in an Australian subtropical estuary downstream from a modified acidic floodplain wetland. Limnology and Oceanography, 60(1):42–56, January 2015. CODEN LIOCAH. ISSN 0024-3590.

Ruiz-Halpern:2010:AWE

[RHSD^{+}10] Sergio Ruiz-Halpern, Mikael K. Sejr, Carlos M. Duarte, Dorte Krause-Jensen, Tage Dalsgaard, Jordi Dachs, and Søren Rysgaard. Air-water exchange and vertical profiles of organic carbon in a subarctic fjord. Limnology and Oceanography, 55(4): 1733–1740, July 2010. CODEN LIOCAH. ISSN 0024-3590.

Reuss:2013:SIR

[RHV^{+}13] Nina S. Reuss, Ladislav Hamerlík, Gaute Velle, Anders Michelsen, Ole Pedersen, and Klaus P. Brodersen. Stable isotopes reveal that chironomids occupy several trophic levels within West Greenland lakes: Implications for food web studies. Limnology and Oceanography, 58(3):1023–1034, May 2013. CODEN LIOCAH. ISSN 0024-3590.

Riessen:2015:WTA

[Rie15] Howard P. Riessen. Water temperature alters predation risk and the adaptive landscape of induced defenses in plankton communities. Limnology and Oceanography, 60(6):2037–2047, November 2015. CODEN LIOCAH. ISSN 0024-3590.

Rojas-Jimenez:2017:THS

[RJFMG17] Keilor Rojas-Jimenez, Jeremy A. Fonvielle, Hua Ma, and Hans-Peter Grossart. Transformation of humic substances by the freshwater ascomycete Cladosporium sp. Limnology and Oceanography, 62(5):1955–1962, September 2017. CODEN LI-OCAH. ISSN 0024-3590.

Rostad:2013:SDP

[RK13] Anders Røstad and Stein Kaartvedt. Seasonal and diel patterns in sedimentary flux of krill fecal pellets recorded by an echo sounder. Limnology and Oceanography, 58(6):1985–1997, November 2013. CODEN LIOCAH. ISSN 0024-3590.

Ralston:2014:TDE

[RKBA14] David K. Ralston, Bruce A. Keafer, Michael L. Brosnahan, and Donald M. Anderson. Temperature dependence of an estuarine harmful algal bloom: Resolving interannual variability in bloom dynamics using a degree-day approach. Limnology and Oceanography, 59(4):1112–1126, July 2014. CODEN LIOCAH. ISSN 0024-3590.

Rellstab:2011:AES

[RKG^{+}11] Christian Rellstab, Barbara Keller, Stèphanie Girardclos, Flavio S. Anselmetti, and Piet Spaak. Anthropogenic eutrophication shapes the past and present taxonomic composition of hybridizing Daphnia in unproductive lakes. Limnology and Oceanography, 56(1):292–302, January 2011. CODEN LIOCAH. ISSN 0024-3590.

Rizk:2014:BSC

[RKL14] William Rizk, Georgiy Kirillin, and Matti Leppäranta. Basinscale circulation and heat fluxes in ice-covered lakes. Limnology and Oceanography, 59(2):445–464, March 2014. CODEN LI-OCAH. ISSN 0024-3590.

Remmel:2011:EAH

[RKLH11] Emily J. Remmel, Nicole Kohmescher, James H. Larson, and K. David Hambright. An experimental analysis of harmful algae–zooplankton interactions and the ultimate defense. Limnology and Oceanography, 56(2):461–470, March 2011. CODEN LIOCAH. ISSN 0024-3590.

Ryan-Keogh:2013:STD

[RKMN^{+}13] Thomas J. Ryan-Keogh, Anna I. Macey, Maria C. Nielsdóttir, Michael I. Lucas, Sebastian S. Steigenberger, Mark C. Stinchcombe, Eric P. Achterberg, Thomas S. Bibby, and C. Mark Moore. Spatial and temporal development of phytoplankton iron stress in relation to bloom dynamics in the high-latitude North Atlantic Ocean. Limnology and Oceanography, 58(2):533–545, March 2013. CODEN LIOCAH. ISSN 0024-3590.

Ryan-Keogh:2018:SRC

[RKTLM18] Thomas J. Ryan-Keogh, Sandy J. Thomalla, Hazel Little, and Jenna-Rose Melanson. Seasonal regulation of the coupling between photosynthetic electron transport and carbon fixation in the Southern Ocean. Limnology and Oceanography, 63(5):1856– 1876, September 2018. CODEN LIOCAH. ISSN 0024-3590.

Remmer:2018:IEF

[RKWH18] Casey R. Remmer, Wynona H. Klemt, Brent B. Wolfe, and Roland I. Hall. Inconsequential effects of flooding in 2014 on lakes in the Peace–Athabasca Delta (Canada) due to long-term drying. Limnology and Oceanography, 63(4):1502–1518, July 2018. CODEN LIOCAH. ISSN 0024-3590.

Rempfer:2010:EEM

[RLB^{+}10] Johannes Rempfer, David M. Livingstone, Christian Blodau, Richard Forster, Pius Niederhauser, and Rolf Kipfer. The effect of the exceptionally mild European winter of 2006–2007 on temperature and oxygen profiles in lakes in Switzerland: a foretaste of the future? Limnology and Oceanography, 55(5):2170–2180, September 2010. CODEN LIOCAH. ISSN 0024-3590.

Reiche:2011:PBC

[RLC^{+}11] Marco Reiche, Shipeng Lu, Valerian Ciobotă, Thomas R. Neu, Sandor Nietzsche, Petra Rösch, Jürgen Popp, and Kirsten Küsel. Pelagic boundary conditions affect the biological formation of iron-rich particles (iron snow) and their microbial communities. Limnology and Oceanography, 56(4):1386–1398, July 2011. CO-DEN LIOCAH. ISSN 0024-3590.

Royer:2010:MDD

[RLL^{+}10] Sarah-Jeanne Royer, Maurice Levasseur, Martine Lizotte, Michael Arychuk, Michael G. Scarratt, Chi Shing Wong, Connie Lovejoy, Marie Robert, Keith Johnson, Angelica Peña, Sonia Michaud, and Ronald P. Kiened. Microbial dimethylsulfoniopropionate (DMSP) dynamics along a natural iron gradient in the northeast subarctic Pacific. Limnology and Oceanography, 55 (4):1614–1626, July 2010. CODEN LIOCAH. ISSN 0024-3590.

Rypina:2014:DPA

[RLPL14] Irina I. Rypina, Joel K. Llopiz, Lawrence J. Pratt, and M. Susan Lozier. Dispersal pathways of American eel larvae from the Sargasso Sea. Limnology and Oceanography, 59(5):1704–1714, September 2014. CODEN LIOCAH. ISSN 0024-3590.

Rico:2013:VPP

[RLSC^{+}13] Milagros Rico, Aroa López, J. Magdalena Santana-Casiano, Aridane G. Gonzàlez, and Melchor Gonzàlez-Dàvila. Variability of the phenolic profile in the diatom Phaeodactylum tricornutum growing under copper and iron stress. Limnology and Oceanography, 58(2):144–152, January 2013. CODEN LIOCAH. ISSN 0024-3590.

Reader:2014:ESP

[RM14] Heather E. Reader and William L. Miller. The efficiency and spectral photon dose dependence of photochemically induced changes to the bioavailability of dissolved organic carbon. Limnology and Oceanography, 59(1):182–194, January 2014. CO-DEN LIOCAH. ISSN 0024-3590.

Rosman:2010:CTW

[RMDK10] Johanna H. Rosman, Stephen G. Monismith, Mark W. Denny, and Jeffrey R. Koseff. Currents and turbulence within a kelp forest (Macrocystis pyrifera): Insights from a dynamically scaled laboratory model. Limnology and Oceanography, 55(3):1145– 1158, May 2010. CODEN LIOCAH. ISSN 0024-3590.

Rautio:2011:SSB

[RMF11] Milla Rautio, Heather Mariash, and Laura Forsström. Seasonal shifts between autochthonous and allochthonous carbon contributions to zooplankton diets in a subarctic lake. Limnology and Oceanography, 56(4):1513–1524, July 2011. CODEN LIOCAH. ISSN 0024-3590.

Rosentreter:2017:STV

[RMH^{+}17] J. A. Rosentreter, D. T. Maher, D. T. Ho, M. Call, J. G. Barr, and B. D. Eyre. Spatial and temporal variability of CO_{2} and CH_{4} gas transfer velocities and quantification of the CH_{4} microbubble flux in mangrove dominated estuaries. Limnology and Oceanography, 62(2):561–578, March 2017. CODEN LIOCAH. ISSN 0024-3590.

Richardson:2018:WMT

[RMJ^{+}18] L. E. Richardson, J. F. Middleton, N. P. James, T. K. Kyser, and B. N. Opdyke. Water masses and their seasonal variation on the Lincoln Shelf, South Australia. Limnology and Oceanography, 63(5):1944–1963, September 2018. CODEN LIOCAH. ISSN 0024-3590.

Rogers:2016:THA

[RMK^{+}16] Justin S. Rogers, Stephen G. Monismith, David A. Koweek, Walter I. Torres, and Robert B. Dunbar. Thermodynamics and hydrodynamics in an atoll reef system and their influence on coral cover. Limnology and Oceanography, 61(6):2191–2206, November 2016. CODEN LIOCAH. ISSN 0024-3590.

Ruiz-Montoya:2012:RHS

[RMLVK12] L. Ruiz-Montoya, R. J. Lowe, K. P. Van Niel, and G. A. Kendrick. The role of hydrodynamics on seed dispersal in seagrasses. Limnology and Oceanography, 57(5):1257–1265, September 2012. CODEN LIOCAH. ISSN 0024-3590.

Rosner:2012:TAW

[RMNZ12] Ralph Rö Rösner, Dörthe C. Müller-Navarra, and Eduardo Zorita. Trend analysis of weekly temperatures and oxygen concentrations during summer stratification in Lake Plußsee: A long-term study. Limnology and Oceanography, 58(1):1479– 1491, September 2012. CODEN LIOCAH. ISSN 0024-3590.

Rominger:2014:EBF

[RN14] Jeffrey T. Rominger and Heidi M. Nepf. Effects of blade flexural rigidity on drag force and mass transfer rates in model blades. Limnology and Oceanography, 59(6):2028–2041, November 2014. CODEN LIOCAH. ISSN 0024-3590.

Robins:2013:PBC

[RNG^{+}13] Peter E. Robins, Simon P. Neill, Luis Giménez, Stuart R. Jenkins, and Shelagh K. Malham. Physical and biological controls on larval dispersal and connectivity in a highly energetic shelf sea. Limnology and Oceanography, 58(2):505–524, March 2013. CODEN LIOCAH. ISSN 0024-3590.

Ratnarajah:2016:UVI

[RNK^{+}16] Lavenia Ratnarajah, Stephen Nicol, So Kawaguchi, Ashley T. Townsend, Delphine Lannuzel, Klaus M. Meiners, and Andrew R. Bowie. Understanding the variability in the iron concentration of Antarctic krill. Limnology and Oceanography, 62 (3):1651–1660, September 2016. CODEN LIOCAH. ISSN 0024-3590.

Rose:2019:PSS

[RNT^{+}19] Kevin C. Rose, Patrick J. Neale, Maria Tzortziou, Charles L. Gallegos, and Thomas E. Jordan. Patterns of spectral, spatial, and long-term variability in light attenuation in an optically complex sub-estuary. Limnology and Oceanography, 64 (S1):S257–S272, January 2019. CODEN LIOCAH. ISSN 0024-

Rishworth:2017:PDB

[RPB17] Gavin M. Rishworth, Renzo Perissinotto, and Matthew S. Bird. Patterns and drivers of benthic macrofaunal communities dwelling within extant peritidal stromatolites. Limnology and Oceanography, 65(9):2227–2242, September 2017. CODEN LIOCAH. ISSN 0024-3590.

Romero:2013:IBS

[RPG13] Estela Romero, Francesc Peters, and Òscar Guadayol. The interplay between short-term, mild physicochemical forcing and plankton dynamics in a coastal area. Limnology and Oceanography, 58(3):903–920, May 2013. CODEN LIOCAH. ISSN 0024-3590.

Ruhland:2010:RAC

[RPH^{+}10] Kathleen M. Rühland, Andrew M. Paterson, Kathryn Hargan, Ashley Jenkin, Bev J. Clark, and John P. Smol. Reorganization of algal communities in the Lake of the Woods (Ontario, Canada) in response to turn-of-the-century damming and recent warming. Limnology and Oceanography, 55(6):2433–2451, November 2010. CODEN LIOCAH. ISSN 0024-3590.

Rao:2012:IPW

[RPI^{+}12] Alexandra M. F. Rao, Lubos Polerecky, Danny Ionescu, Filip J. R. Meysman, and Dirk de Beer. The influence of pore-water advection, benthic photosynthesis, and respiration on calcium carbonate dynamics in reef sands. Limnology and Oceanography, 57(3):809–825, May 2012. CODEN LIOCAH. ISSN 0024-3590.

Rypina:2016:IOC

[RPL16] Irina I. Rypina, Lawrence J. Pratt, and M. Susan Lozier. Influence of ocean circulation changes on the inter-annual variability of American eel larval dispersal. Limnology and Oceanography, 61(5):1574–1588, September 2016. CODEN LIOCAH. ISSN 0024-3590.

Ramshaw:2017:QST

[RPMK17] Brock C. Ramshaw, Evgeny A. Pakhomov, Russell W. Markel, and Sven Kaehler. Quantifying spatial and temporal variations in phytoplankton and kelp isotopic signatures to estimate the distribution of kelp-derived detritus off the west coast of Vancouver Island, Canada. Limnology and Oceanography, 65(9):

2133–2153, September 2017. CODEN LIOCAH. ISSN 0024-3590.

Razavi:2015:EEM

[RQC^{+}15] N. Roxanna Razavi, Mingzhi Qu, Dongmei Chen, Yang Zhong, Wenwei Ren, Yuxiang Wang, and Linda M. Campbell. Effect of eutrophication on mercury (Hg) dynamics in subtropical reservoirs from a high Hg deposition ecoregion. Limnology and Oceanography, 60(2):386–401, March 2015. CODEN LIOCAH. ISSN 0024-3590.

Rokitta:2012:ECT

[RR12] Sebastian D. Rokitta and Björn Rost. Effects of CO_{2} and their modulation by light in the life-cycle stages of the coccolithophore Emiliania huxleyi. Limnology and Oceanography, 57(3):607–618, March 2012. CODEN LIOCAH. ISSN 0024-3590.

Read:2013:PRS

[RR13] Jordan S. Read and Kevin C. Rose. Physical responses of small temperate lakes to variation in dissolved organic carbon concentrations. Limnology and Oceanography, 58(3):921–931, May 2013. CODEN LIOCAH. ISSN 0024-3590.

Radbourne:2017:HDO

[RRAS17] Alan D. Radbourne, David B. Ryves, N. John Anderson, and Daniel R. Scott. The historical dependency of organic carbon burial efficiency. Limnology and Oceanography, 62(4):1480–1497, July 2017. CODEN LIOCAH. ISSN 0024-3590.

Robertson:2016:DNR

[RRB^{+}16] Elizabeth K. Robertson, Keryn L. Roberts, Laurine D. W. Burdorf, Perran Cook, and Bo Thamdrup. Dissimilatory nitrate reduction to ammonium coupled to Fe(II) oxidation in sediments of a periodically hypoxic estuary. Limnology and Oceanography, 61(1):365–381, January 2016. CODEN LIOCAH. ISSN 0024-3590.

Romero-Romero:2019:DTE

[RRCH^{+}19] Sonia Romero-Romero, C. Anela Choy, Cecelia C. S. Hannides, Brian N. Popp, and Jeffrey C. Drazen. Differences in the trophic ecology of micronekton driven by diel vertical migration. Limnology and Oceanography, 64(4):1473–1483, July 2019. CODEN LIOCAH. ISSN 0024-3590.

Reynolds:2014:FEE

[RRD14] Pamela L. Reynolds, J. Paul Richardson, and J. Emmett Duffy. Field experimental evidence that grazers mediate transition between microalgal and seagrass dominance. Limnology and Oceanography, 59(3):1053–1064, May 2014. CODEN LIOCAH. ISSN 0024-3590.

Romero-Romero:2019:SVD

[RRGGCA19] Sonia Romero-Romero, Ricardo González-Gil, Carlos Cáceres, and José Luis Acuña. Seasonal and vertical dynamics in the trophic structure of a temperate zooplankton assemblage. Limnology and Oceanography, 66(8):1939–1948, September 2019. CODEN LIOCAH. ISSN 0024-3590.

Rafter:2016:SDT

[RS16] Patrick A. Rafter and Daniel M. Sigman. Spatial distribution and temporal variation of nitrate nitrogen and oxygen isotopes in the upper equatorial Pacific Ocean. Limnology and Oceanography, 61(1):14–31, January 2016. CODEN LIOCAH. ISSN 0024-3590.

Reynolds:2019:OCM

[RS19] Rick A. Reynolds and Dariusz Stramski. Optical characterization of marine phytoplankton assemblages within surface waters of the western Arctic Ocean. Limnology and Oceanography, 64 (6):2478–2496, November 2019. CODEN LIOCAH. ISSN 0024-3590.

Reavie:2017:CWC

[RSE^{+}17] Euan D. Reavie, Gerald V. Sgro, Lisa R. Estepp, Andrew J. Bramburger, Victoria L. Shaw Chraïbi, Robert W. Pillsbury, Meijun Cai, Craig A. Stow, and Alice Dove. Climate warming and changes in Cyclotella sensu lato in the Laurentian Great Lakes. Limnology and Oceanography, 62(2):768–783, March 2017. CODEN LIOCAH. ISSN 0024-3590.

Reed:2011:SPD

[RSG11] Daniel C. Reed, Caroline P. Slomp, and Bo G. Gustafsson. Sedimentary phosphorus dynamics and the evolution of bottomwater hypoxia: a coupled benthic–pelagic model of a coastal system. Limnology and Oceanography, 57(4):1075–1092, May 2011. CODEN LIOCAH. ISSN 0024-3590.

Ren:2018:CPF

[RSJ^{+}18] Lijuan Ren, Xingyu Song, Erik Jeppesen, Peng Xing, Lone Liboriussen, Xiangrong Xu, and Qinglong L. Wu. Contrasting patterns of freshwater microbial metabolic potentials and functional gene interactions between an acidic mining lake and a weakly alkaline lake. Limnology and Oceanography, 63(S1):S354–S366, March 2018. CODEN LIOCAH. ISSN 0024-3590.

Richard:2013:KMA

[RSM13] Dominique Richard, Bjørn Sundby, and Alfonso Mucci. Kinetics of manganese adsorption, desorption, and oxidation in coastal marine sediments. Limnology and Oceanography, 58(3):987–996, May 2013. CODEN LIOCAH. ISSN 0024-3590.

Reynolds:2016:OBP

[RSN16] Rick A. Reynolds, Dariusz Stramski, and Griet Neukermans. Optical backscattering by particles in Arctic seawater and relationships to particle mass concentration, size distribution, and bulk composition. Limnology and Oceanography, 62(3):1869– 1890, September 2016. CODEN LIOCAH. ISSN 0024-3590.

Ren:2012:NIC

[RSTP12] Haojia Ren, Daniel M. Sigman, Robert C. Thunell, and Maria G. Prokopenko. Nitrogen isotopic composition of planktonic Foraminifera from the modern ocean and recent sediments. Limnology and Oceanography, 57(4):1011–1024, July 2012. CO-DEN LIOCAH. ISSN 0024-3590.

Roth-Schulze:2018:EWO

[RSTS^{+}18] Alexandra J. Roth-Schulze, Torsten Thomas, Peter Steinberg, Marty R. Deveney, Jason E. Tanner, Kathryn H. Wiltshire, Stephanie Papantoniou, John W. Runcie, and C. Frederico D. Gurgel. The effects of warming and ocean acidification on growth, photosynthesis, and bacterial communities for the marine invasive macroalga Caulerpa taxifolia. Limnology and Oceanography, 63(4):459–471, January 2018. CODEN LIOCAH. ISSN 0024-3590.

Ribot:2017:UPD

[RvSM17] M. Ribot, D. von Schiller, and E. Martí. Understanding pathways of dissimilatory and assimilatory dissolved inorganic nitrogen uptake in streams. Limnology and Oceanography, 63(3): 1166–1183, May 2017. CODEN LIOCAH. ISSN 0024-3590.

Rozema:2017:IVP

[RVvdP^{+}17] P. D. Rozema, H. J. Venables, W. H. van de Poll, A. Clarke, M. P. Meredith, and A. G. J. Buma. Interannual variability in phytoplankton biomass and species composition in northern Marguerite Bay (West Antarctic Peninsula) is governed by both winter sea ice cover and summer stratification. Limnology and Oceanography, 62(1):235–252, January 2017. CODEN LIOCAH. ISSN 0024-3590.

Robinson:2019:RBE

[RWB^{+}19] Tiera-Brandy Robinson, Oliver Wurl, Enno Bahlmann, Klaus Jürgens, and Christian Stolle. Rising bubbles enhance the gelatinous nature of the air–sea interface. Limnology and Oceanography, 64(6):2358–2372, November 2019. CODEN LIOCAH. ISSN 0024-3590.

Russell:2016:BBN

[RWC16] Douglas G. Russell, F. Y. Warry, and Perran L. M. Cook. The balance between nitrogen fixation and denitrification on vegetated and non-vegetated intertidal sediments. Limnology and Oceanography, 61(6):2058–2075, November 2016. CODEN LI-OCAH. ISSN 0024-3590.

Rose:2012:RUR

[RWF^{+}12] Kevin C. Rose, Craig E. Williamson, Janet M. Fischer, Sandra J. Connelly, Mark Olson, Andrew J. Tucker, and Douglas A. Noe. The role of ultraviolet radiation and fish in regulating the vertical distribution of Daphnia. Limnology and Oceanography, 57 (6):1867–1876, November 2012. CODEN LIOCAH. ISSN 0024-3590.

Rouco:2014:TSC

[RWM^{+}14] Mönica Rouco, Hannah Joy Warren, Dennis J. McGillicuddy, Jr., John B. Waterbury, and Sonya T. Dyhrman. Trichodesmium sp. clade distributions in the western North Atlantic Ocean. Limnology and Oceanography, 59(6):1899–1909, November 2014. CODEN LIOCAH. ISSN 0024-3590.

Reisinger:2019:CLT

[RWM^{+}19] Alexander J. Reisinger, Ellen Woytowitz, Emily Majcher, Emma J. Rosi, Kenneth T. Belt, Jonathan M. Duncan, Sujay S. Kaushal, and Peter M. Groffman. Changes in long-term water quality of Baltimore streams are associated with both gray and green infrastructure. Limnology and Oceanography, 64(S1): S60–S76, January 2019. CODEN LIOCAH. ISSN 0024-3590.

Rontani:2011:DPO

[RZW11] J.-F. Rontani, N. Zabeti, and S. G. Wakeham. Degradation of particulate organic matter in the equatorial Pacific Ocean: Biotic or abiotic? Limnology and Oceanography, 56(1):333–349, January 2011. CODEN LIOCAH. ISSN 0024-3590.

Savoy:2019:MRF

[SAH^{+}19] Philip Savoy, Alison P. Appling, James B. Heffernan, Edward G. Stets, Jordan S. Read, Judson W. Harvey, and Emily S. Bernhardt. Metabolic rhythms in flowing waters: an approach for classifying river productivity regimes. Limnology and Oceanography, 64(5):1835–1851, September 2019. CODEN LIOCAH. ISSN 0024-3590.

Sivan:2011:GEI

[SAP^{+}11] Orit Sivan, Michal Adler, Ann Pearson, Faina Gelman, Itay Bar-Or, Seth G. John, and Werner Eckert. Geochemical evidence for iron-mediated anaerobic oxidation of methane. Limnology and Oceanography, 56(4):1536–1544, July 2011. CODEN LIOCAH. ISSN 0024-3590.

Schmidt:2014:FOA

[SAPI14] Katrin Schmidt, Angus Atkinson, David W. Pond, and Louise C. Ireland. Feeding and overwintering of Antarctic krill across its major habitats: The role of sea ice cover, water depth, and phytoplankton abundance. Limnology and Oceanography, 59(1): 17–36, January 2014. CODEN LIOCAH. ISSN 0024-3590.

Schmidt:2011:SFA

[SAS^{+}11] Katrin Schmidt, Angus Atkinson, Sebastian Steigenberger, Sophie Fielding, Margaret C. M. Lindsay, David W. Pond, Geraint A. Tarling, Thor A. Klevjer, Claire S. Allen, Stephen Nicol, and Eric P. Achterberg. Seabed foraging by Antarctic krill: Implications for stock assessment, bentho-pelagic coupling, and the vertical transfer of iron. Limnology and Oceanography, 56(4):1411–1428, July 2011. CODEN LIOCAH. ISSN 0024-3590.

Sal:2015:TAP

[SASB^{+}15] Sofía Sal, Laura Alonso-Sáez, Juan Bueno, Francisca C. García, and Ángel López-Urrutia. Thermal adaptation, phylogeny, and the unimodal size scaling of marine phytoplankton growth. Limnology and Oceanography, 60(4):1212–1221, July 2015. CODEN LIOCAH. ISSN 0024-3590.

Strecker:2011:DIE

[SBA^{+}11] Angela L. Strecker, Beatrix E. Beisner, Shelley E. Arnott, Andrew M. Paterson, Jennifer G. Winter, Ora E. Johannsson, and Norman D. Yan. Direct and indirect effects of an invasive planktonic predator on pelagic food webs. Limnology and Oceanography, 56(1):179–192, January 2011. CODEN LIOCAH. ISSN 0024-3590.

Spence:2018:ILH

[SBB^{+}18] Christopher Spence, Ken Beaty, Paul J. Blanchfield, Lee Hrenchuk, and Murray D. MacKay. The impact of a loss of hydrologic connectivity on boreal lake thermal and evaporative regimes. Limnology and Oceanography, 64(4):2028–2044, September 2018. CODEN LIOCAH. ISSN 0024-3590.

Shroyer:2014:SMR

[SBBNM14] E. L. Shroyer, K. J. Benoit-Bird, J. D. Nash, and J. N. Moum. Stratification and mixing regimes in biological thin layers over the Mid-Atlantic Bight. Limnology and Oceanography, 59(4): 1349–1363, July 2014. CODEN LIOCAH. ISSN 0024-3590.

Sipler:2017:CPI

[SBC^{+}17] Rachel E. Sipler, Steven E. Baer, Tara L. Connelly, Marc E. Frischer, Quinn N. Roberts, Patricia L. Yager, and Deborah A. Bronk. Chemical and photophysiological impact of terrestriallyderived dissolved organic matter on nitrate uptake in the coastal western Arctic. Limnology and Oceanography, 62(5):1881–1894, September 2017. CODEN LIOCAH. ISSN 0024-3590.

Schoon:2010:RPW

[SBdB10] Raphaela Schoon, Andrew Bissett, and Dirk de Beer. Resilience of pore-water chemistry and calcification in photosynthetic zones of calcifying sediments. Limnology and Oceanography, 55(1):377–385, January 2010. CODEN LIOCAH. ISSN 0024-3590.

Senft-Batoh:2015:IPP

[SBDS^{+}15] Christina D. Senft-Batoh, Hans G. Dam, Sandra E. Shumway, Gary H. Wikfors, and Carl D. Schlichting. Influence of predator– prey evolutionary history, chemical alarm-cues, and feeding selection on induction of toxin production in a marine dinoflagellate. Limnology and Oceanography, 60(1):318–328, January 2015. CODEN LIOCAH. ISSN 0024-3590.

Schmid:2010:DDC

[SBe10] Martin Schmid, Myles Busbridge, and Alfred Wü est. Doublediffusive convection in Lake Kivu. Limnology and Oceanography, 55(1):225–238, January 2010. CODEN LIOCAH. ISSN 0024-3590.

Stawiarski:2018:PRS

[SBF18] Beate Stawiarski, Erik T. Buitenhuis, and Mehera Fallens. The physiological response of seven strains of picophytoplankton to light, and its representation in a dynamic photosynthesis model. Limnology and Oceanography, 63(S1):S367–S380, March 2018. CODEN LIOCAH. ISSN 0024-3590.

Strom:2017:SCS

[SBFB17] Suzanne Strom, Kelley Bright, Kerri Fredrickson, and Bianca Brahamsha. The Synechococcus cell surface protein SwmA increases vulnerability to predation by flagellates and ciliates. Limnology and Oceanography, 62(2):784–794, March 2017. CO-DEN LIOCAH. ISSN 0024-3590.

Strom:2018:PDD

[SBFC18] Suzanne L. Strom, Kelley J. Bright, Kerri A. Fredrickson, and Elizabeth C. Cooney. Phytoplankton defenses: Do Emiliania huxleyi coccoliths protect against microzooplankton predators? Limnology and Oceanography, 63(2):617–627, March 2018. CO-DEN LIOCAH. ISSN 0024-3590.

Steenbergh:2011:PRC

[SBH^{+}11] Anne K. Steenbergh, Paul L. E. Bodelier, Hans L. Hoogveld, Caroline P. Slomp, and Hendrikus J. Laanbroek. Phosphatases relieve carbon limitation of microbial activity in Baltic Sea sediments along a redox-gradient. Limnology and Oceanography, 56 (6):2018–2026, November 2011. CODEN LIOCAH. ISSN 0024-3590.

Seekell:2018:LMM

[SBK18] David A. Seekell, Pär Byström, and Jan Karlsson. Lake morphometry moderates the relationship between water color and fish biomass in small boreal lakes. Limnology and Oceanography, 64(4):2171–2178, September 2018. CODEN LIOCAH. ISSN 0024-3590.

Stukel:2018:LPT

[SBKO18] Michael R. Stukel, Tristan Biard, Jeffrey Krause, and Mark D. Ohman. Large Phaeodaria in the twilight zone: Their role in the carbon cycle. Limnology and Oceanography, 63(6):2579–2594, November 2018. CODEN LIOCAH. ISSN 0024-3590.

Seegers:2015:SSS

[SBM^{+}15] Bridget N. Seegers, James M. Birch, Roman Marin III, Chris A. Scholin, David A. Caron, Erica L. Seubert, Meredith D. A. Howard, George L. Robertson, and Burton H. Jones. Subsurface seeding of surface harmful algal blooms observed through the integration of autonomous gliders, moored environmental sample processors, and satellite remote sensing in southern California. Limnology and Oceanography, 60(3):754–764, May 2015. CODEN LIOCAH. ISSN 0024-3590.

Schade:2016:GGF

[SBM16] John D. Schade, Jason Bailio, and William H. McDowell. Greenhouse gas flux from headwater streams in New Hampshire, USA: Patterns and drivers. Limnology and Oceanography, 61(S1): S165–S174, November 2016. CODEN LIOCAH. ISSN 0024-3590.

Shi:2019:DTL

[SBNC^{+}19] Xiangming Shi, Claudia R. Benitez-Nelson, Pinghe Cai, Lijian He, and Willard S. Moore. Development of a two-layer transport model in layered muddy–permeable marsh sediments using 224Ra–^{228}Th disequilibria. Limnology and Oceanography, 64(4): 1672–1687, July 2019. CODEN LIOCAH. ISSN 0024-3590.

Solomon:2013:ERD

[SBR^{+}13] Christopher T. Solomon, Denise A. Bruesewitz, David C. Richardson, Kevin C. Rose, Matthew C. Van de Bogert, Paul C. Hanson, Timothy K. Kratz, Bret Larget, Rita Adrian, Brenda Leroux Babin, Chih-Yu Chiu, David P. Hamilton, Evelyn E. Gaiser, Susan Hendricks, Vera Istvànovics, Alo Laas, David M. O’Donnell, Michael L. Pace, Elizabeth Ryder, Peter A. Staehr, Thomas Torgersen, Michael J. Vanni, Kathleen C. Weathers, and Guangwei Zhu. Ecosystem respiration: Drivers of daily variability and background respiration in lakes around the globe. Limnology and Oceanography, 58(3):849–866, May 2013. CODEN LIOCAH. ISSN 0024-3590.

Small:2013:RCN

[SBS^{+}13] Gaston E. Small, George S. Bullerjahn, Robert W. Sterner, Benjamin F. N. Beall, Sandra Brovold, Jacques C. Finlay, Robert M. L. McKay, and Maitreyee Mukherjee. Rates and controls of nitrification in a large oligotrophic lake. Limnology and Oceanography, 58(2):276–286, January 2013. CODEN LIOCAH. ISSN 0024-3590.

Saulnier:2019:SCM

[SBT^{+}19] Erwan Saulnier, Anik Brind’Amour, Adrien Tableau, Marta M. Rufino, Jean-Claude Dauvin, Christophe Luczak, and Hervé Le Bris. Seasonality in coastal macrobenthic biomass and its implications for estimating secondary production using empirical models. Limnology and Oceanography, 64(3):935–949, May 2019. CODEN LIOCAH. ISSN 0024-3590.

Schilder:2015:SCI

[SBvH^{+}15] Jos Schilder, David Bastviken, Maarten van Hardenbroek, Markus Leuenberger, Päivi Rinta, Tabea Stötter, and Oliver Heiri. The stable carbon isotopic composition of Daphnia ephippia in small, temperate lakes reflects in-lake methane availability. Limnology and Oceanography, 60(3):1064–1075, May 2015. CODEN LIOCAH. ISSN 0024-3590.

Swanson:2010:DHT

[SC10] Travis E. Swanson and M. Bayani Cardenas. Diel heat transport within the hyporheic zone of a pool-riffle-pool sequence of a losing stream and evaluation of models for fluid flux estimation using heat. Limnology and Oceanography, 55(4):1741–1754, July 2010. CODEN LIOCAH. ISSN 0024-3590.

Sanpera-Calbet:2016:NNS

[SCAB^{+}16] Isis Sanpera-Calbet, Vicenç Acuña, Andrea Butturini, Rafael Marcé, and Isabel Muñoz. El Niño southern oscillation and seasonal drought drive riparian input dynamics in a Mediterranean stream. Limnology and Oceanography, 61(1):214–226, January 2016. CODEN LIOCAH. ISSN 0024-3590.

Staehr:2012:EMS

[SCBR12] Peter A. Staehr, Jesper P. A. Christensen, Ryan D. Batt, and Jordan S. Read. Ecosystem metabolism in a stratified lake. Limnology and Oceanography, 57(5):1317–1330, September 2012. CODEN LIOCAH. ISSN 0024-3590.

Sanders:2015:SMF

[SCF^{+}15] Robert W. Sanders, Sandra L. Cooke, Janet M. Fischer, Samuel B. Fey, Adam W. Heinze, Wade H. Jeffrey, Amy L. Macaluso, Robert E. Moeller, Donald P. Morris, Patrick J. Neale, Mark H. Olson, J. Dean Pakulski, Jason A. Porter, Donald M. Schoener, and Craig E. Williamson. Shifts in microbial food web structure and productivity after additions of naturally occurring dissolved organic matter: Results from large-scale lacustrine mesocosms. Limnology and Oceanography, 60(6):2130– 2144, November 2015. CODEN LIOCAH. ISSN 0024-3590.

Sutherland:2019:ESP

[SCG^{+}19] Kevin M. Sutherland, Allison Coe, Rebecca J. Gast, Sydney Plummer, Christopher P. Suffridge, Julia M. Diaz, Jeff S. Bowman, Scott D. Wankel, and Colleen M. Hansel. Extracellular superoxide production by key microbes in the global ocean. Limnology and Oceanography, 64(6):2679–2693, November 2019. CODEN LIOCAH. ISSN 0024-3590.

Schaum:2019:EBF

[Sch19] C.-E. Schaum. Enhanced biofilm formation aids adaptation to extreme warming and environmental instability in the diatom Thalassiosira pseudonana and its associated bacteria. Limnology and Oceanography, 64(2):441–460, March 2019. CODEN LIOCAH. ISSN 0024-3590.

Stanley:2019:BLW

[SCL^{+}19] Emily H. Stanley, Sarah M. Collins, Noah R. Lottig, Samantha K. Oliver, Katherine E. Webster, Kendra S. Cheruvelil, and Patricia A. Soranno. Biases in lake water quality sampling and implications for macroscale research. Limnology and Oceanography, 64(4):1572–1585, July 2019. CODEN LIOCAH. ISSN 0024-3590.

Sweetman:2016:JDS

[SCP^{+}16] Andrew K. Sweetman, Ariella Chelsky, Kylie A. Pitt, Hector Andrade, Dick van Oevelen, and Paul E. Renaud. Jellyfish decomposition at the seafloor rapidly alters biogeochemical cycling and carbon flow through benthic food-webs. Limnology and Oceanography, 61(4):1449–1461, July 2016. CODEN LI-OCAH. ISSN 0024-3590.

Schubert:2015:LCS

[SCPE15] Nadine Schubert, María Florencia Colombo-Pallota, and Susana Enríquez. Leaf and canopy scale characterization of the photoprotective response to high-light stress of the seagrass Thalassia testudinum. Limnology and Oceanography, 60(1):286–302, January 2015. CODEN LIOCAH. ISSN 0024-3590.

Sun:2017:SOV

[SCQ^{+}17] Deyong Sun, Shuguo Chen, Zhongfeng Qiu, Shengqiang Wang, Yu Huan, Yijun He, and Tinglu Zhang. Second-order variability of inherent optical properties of particles in Bohai Sea and Yellow Sea: Driving factor analysis and modeling. Limnology and Oceanography, 63(3):1266–1287, May 2017. CODEN LIOCAH. ISSN 0024-3590.

Santos:2012:SWP

[SCR^{+}12] Isaac R. Santos, Perran L. M. Cook, Louissa Rogers, Jason de Weys, and Bradley D. Eyre. The “salt wedge pump”: Convection-driven pore-water exchange as a source of dissolved organic and inorganic carbon and nitrogen to an estuary. Limnology and Oceanography, 58(1):1415–1426, September 2012. CODEN LIOCAH. ISSN 0024-3590.

Scully:2016:CPP

[Scu16] Malcolm E. Scully. The contribution of physical processes to inter-annual variations of hypoxia in Chesapeake Bay: a 30-yr modeling study. Limnology and Oceanography, 61(6):2243–2260, November 2016. CODEN LIOCAH. ISSN 0024-3590.

Siuda:2010:EOP

[SD10] Amy N. S. Siuda and Hans G. Dam. Effects of omnivory and predator-prey elemental stoichiometry on planktonic trophic interactions. Limnology and Oceanography, 55(5):2107–2116, September 2010. CODEN LIOCAH. ISSN 0024-3590.

Smith:2016:FIN

[SDCF16] Jason M. Smith, Julian Damashek, Francisco P. Chavez, and Christopher A. Francis. Factors influencing nitrification rates and the abundance and transcriptional activity of ammoniaoxidizing microorganisms in the dark northeast Pacific Ocean. Limnology and Oceanography, 61(2):596–609, March 2016. CO-DEN LIOCAH. ISSN 0024-3590.

Smith:2014:CCE

[SDH^{+}14] Val H. Smith, Walter K. Dodds, Karl E. Havens, Daniel R. Engstrom, Hans W. Paerl, Brian Moss, and Gene E. Likens. Comment: Cultural eutrophication of natural lakes in the United States is real and widespread. Limnology and Oceanography, 61 (1):2217–2225, November 2014. CODEN LIOCAH. ISSN 0024-3590. See [BHC13].

Shintani:2010:GWN

[SdlFdlF^{+}10] Tetsuya Shintani, Alberto de la Fuente, Alberto de la Fuente, Yarko Niño, and Jörg Imberger. Generalizations of the Wedderburn number: Parameterizing upwelling in stratified lakes. Limnology and Oceanography, 55(3):1377–1389, May 2010. CO-DEN LIOCAH. ISSN 0024-3590.

Straza:2010:ASC

[SDMK10] Tiffany R. A. Straza, Hugh W. Ducklow, Alison E. Murray, and David L. Kirchman. Abundance and single-cell activity of bacterial groups in Antarctic coastal waters. Limnology and Oceanography, 55(6):2526–2536, November 2010. CODEN LI-OCAH. ISSN 0024-3590.

Schwartzmann:2011:SGC

[SDS^{+}11] Caroline Schwartzmann, Gilles Durrieu, Mohamedou Sow, Pierre Ciret, Claire E. Lazareth, and Jean-Charles Massabuau. In situ giant clam growth rate behavior in relation to temperature: a one-year coupled study of high-frequency noninvasive valvometry and sclerochronology. Limnology and Oceanography, 56(5):1940–1951, September 2011. CODEN LIOCAH. ISSN 0024-3590.

Sieczko:2016:AMD

[SDS^{+}16] Anna Katarzyna Sieczko, Katalin Demeter, Gabriel Andreas Singer, Michael Tritthart, Stefan Preiner, Magdalena Mayr, Karin Meisterl, and Peter Peduzzi. Aquatic methane dynamics in a human-impacted river-floodplain of the Danube. Limnology and Oceanography, 61(S1):S175–S187, November 2016. CODEN LIOCAH. ISSN 0024-3590.

Schlosser:2012:SIS

[SDSC12] Christian Schlosser, Christina L. De La Rocha, Peter Streu, and Peter L. Croot. Solubility of iron in the Southern Ocean. Limnology and Oceanography, 57(3):684–697, May 2012. CODEN LIOCAH. ISSN 0024-3590.

Samanta:2018:ZIF

[SES18] Moneesha Samanta, Michael J. Ellwood, and Robert F. Strzepek. Zinc isotope fractionation by Emiliania huxleyi cultured across a range of free zinc ion concentrations. Limnology and Oceanography, 63(2):660–671, March 2018. CODEN LI-OCAH. ISSN 0024-3590.

Shank:2011:SRE

[SEYJ11] G. Christopher Shank, Anne Evans, Youhei Yamashita, and Rudolf Jaffé. Solar radiation–enhanced dissolution of particulate organic matter from coastal marine sediments. Limnology and Oceanography, 56(2):577–588, March 2011. CODEN LIOCAH. ISSN 0024-3590.

Slusarczyk:2019:LQP

[ŚF19] M. Ślusarczyk and S. Flis. Light quantity, not photoperiod terminates diapause in the crustacean Daphnia. Limnology and Oceanography, 64(1):124–130, January 2019. CODEN LIOCAH. ISSN 0024-3590.

Shen:2012:FID

[SFB12] Yuan Shen, Cédric G. Fichot, and Ronald Benner. Floodplain influence on dissolved organic matter composition and export from the Mississippi–Atchafalaya River system to the Gulf of Mexico. Limnology and Oceanography, 57(4):1149–1160, July 2012. CODEN LIOCAH. ISSN 0024-3590.

Shiozaki:2018:DCS

[SFI^{+}18] Takuhei Shiozaki, Amane Fujiwara, Minoru Ijichi, Naomi Harada, Shigeto Nishino, Shinro Nishi, Toshi Nagata, and Koji Hamasaki. Diazotroph community structure and the role of nitrogen fixation in the nitrogen cycle in the Chukchi Sea (western Arctic Ocean). Limnology and Oceanography, 64(4):2191–2205, September 2018. CODEN LIOCAH. ISSN 0024-3590.

Shen:2016:BHS

[SFLB16] Yuan Shen, Cédric G. Fichot, Sheng-Kang Liang, and Ronald Benner. Biological hot spots and the accumulation of marine dissolved organic matter in a highly productive ocean margin. Limnology and Oceanography, 61(4):1287–1300, July 2016. CO-DEN LIOCAH. ISSN 0024-3590.

Spilling:2019:PPC

[SFLQ^{+}19] Kristian Spilling, Antonio Fuentes-Lema, Daniel Quemaliños, Riina Klais, and Cristina Sobrino. Primary production, carbon release, and respiration during spring bloom in the Baltic Sea. Limnology and Oceanography, 64(4):1779–1789, July 2019. CODEN LIOCAH. ISSN 0024-3590.

Scheu:2015:STD

[SFMF15] K. R. Scheu, D. A. Fong, S. G. Monismith, and O. B. Fringer. Sediment transport dynamics near a river inflow in a large alpine lake. Limnology and Oceanography, 60(4):1195–1211, July 2015. CODEN LIOCAH. ISSN 0024-3590.

Sahoo:2016:CCI

[SFS^{+}16] G. B. Sahoo, A. L. Forrest, S. G. Schladow, J. E. Reuter, R. Coats, and M. Dettinger. Climate change impacts on lake thermal dynamics and ecosystem vulnerabilities. Limnology and Oceanography, 61(2):496–507, March 2016. CODEN LIOCAH. ISSN 0024-3590.

Selander:2012:FFD

[SFWP12] Erik Selander, Tony Fagerberg, Sylke Wohlrab, and Henrik Pavia. Fight and flight in dinoflagellates? Kinetics of simultaneous grazer-induced responses in Alexandrium tamarense. Limnology and Oceanography, 57(1):58–64, January 2012. CODEN LIOCAH. ISSN 0024-3590.

Schwalb:2010:DFM

[SGA10] Astrid Nadine Schwalb, Marcus Garvie, and Josef Daniel Ackerman. Dispersion of freshwater mussel larvae in a lowland river. Limnology and Oceanography, 55(2):628–638, March 2010. CO-DEN LIOCAH. ISSN 0024-3590.

Shangguan:2017:NPS

[SGA^{+}17] Yini Shangguan, Patricia M. Glibert, Jeffrey A. Alexander, Christopher J. Madden, and Susan Murasko. Nutrients and phytoplankton in semienclosed lagoon systems in Florida Bay and their responses to changes in flow from Everglades restoration. Limnology and Oceanography, 62(S1):S327–S347, November 2017. CODEN LIOCAH. ISSN 0024-3590.

Sutherland:2016:PCC

[SGCC16] Kelly R. Sutherland, Brad J. Gemmell, Sean P. Colin, and John H. Costello. Prey capture by the cosmopolitan hydromedusae, Obelia spp., in the viscous regime. Limnology and Oceanography, 61(6):2309–2317, November 2016. CODEN LI-OCAH. ISSN 0024-3590.

Saiz:2014:FRP

[SGCI14] Enric Saiz, Kaiene Griffell, Albert Calbet, and Stamatina Isari. Feeding rates and prey : predator size ratios of the nauplii and adult females of the marine cyclopoid copepod Oithona davisae. Limnology and Oceanography, 59(6):2077–2088, November 2014. CODEN LIOCAH. ISSN 0024-3590.

Schultz:2011:CIR

[SGG^{+}11] Stewart T. Schultz, Jeffrey H. R. Goddard, Terrence M. Gosliner, Douglas E. Mason, William E. Pence, Gary R. Mc-Donald, Vicki B. Pearse, and John S. Pearse. Climate-index response profiling indicates larval transport is driving population fluctuations in nudibranch gastropods from the northeast Pacific Ocean. Limnology and Oceanography, 56(2):749–763, March 2011. CODEN LIOCAH. ISSN 0024-3590.

Sabbah:2012:RFI

[SGH12] Shai Sabbah, Suzanne M. Gray, and Craig W. Hawryshyn. Radiance fluctuations induced by surface waves can enhance the appearance of underwater objects. Limnology and Oceanography, 57(4):1025–1041, July 2012. CODEN LIOCAH. ISSN 0024-3590.

Simek:2018:BPF

[ŠGH^{+}18] Karel Šimek, Vesna Grujčić, Martin W. Hahn, Karel Horñák, Jitka Jezberová, Vojtěch Kasalický, Jiří Nedoma, Michaela M. Salcher, and Tanja Shabarova. Bacterial prey food characteristics modulate community growth response of freshwater bacterivorous flagellates. Limnology and Oceanography, 63(4):484– 502, January 2018. CODEN LIOCAH. ISSN 0024-3590.

Sastri:2014:BPR

[SGJB14] Akash R. Sastri, Joanna Gauthier, Philippe Juneau, and Beatrix E. Beisner. Biomass and productivity responses of zooplankton communities to experimental thermocline deepening. Limnology and Oceanography, 59(1):1–16, January 2014. CODEN LIOCAH. ISSN 0024-3590.

Schubert:2011:PAR

[SGME11] Nadine Schubert, Ernesto García-Mendoza, and Susana Enríquez. Is the photo-acclimatory response of Rhodophyta conditioned by the species carotenoid profile? Limnology and Oceanography, 56(6):2347–2361, November 2011. CODEN LIOCAH. ISSN 0024-3590.

Simek:2019:MFW

[ŠGN^{+}19] Karel Šimek, Vesna Grujčić, Jiří Nedoma, Jitka Jezberová, Michal Šorf, Anna Matoušů, Libor Pechar, Thomas Posch, Estelle P. Bruni, and Jaroslav Vrba. Microbial food webs in hypertrophic fishponds: Omnivorous ciliate taxa are major protistan bacterivores. Limnology and Oceanography, 66(4):2295–2309, September 2019. CODEN LIOCAH. ISSN 0024-3590.

Suberkropp:2010:EPS

[SGRB10] Keller Suberkropp, Vladislav Gulis, Amy D. Rosemond, and Jonathan P. Benstead. Ecosystem and physiological scales of microbial responses to nutrients in a detritus-based stream: Results of a 5-year continuous enrichment. Limnology and Oceanography, 55(1):149–160, January 2010. CODEN LIOCAH. ISSN 0024-3590.

Spivak:2018:SPB

[SGS18] Amanda C. Spivak, Kelsey M. Gosselin, and Sean P. Sylva. Shallow ponds are biogeochemically distinct habitats in salt marsh ecosystems. Limnology and Oceanography, 63(4):1622– 1642, July 2018. CODEN LIOCAH. ISSN 0024-3590.

Schneider:2016:CAC

[SGVR16] Tobias Schneider, Guillaume Grosbois, Warwick F. Vincent, and Milla Rautio. Carotenoid accumulation in copepods is related to lipid metabolism and reproduction rather than to UVprotection. Limnology and Oceanography, 61(4):1201–1213, July 2016. CODEN LIOCAH. ISSN 0024-3590.

Sereda:2010:CEP

[SH10a] Jeff M. Sereda and Jeff J. Hudson. Comparative estimate of P fluxes in lakes: a comment on “Fish decomposition in boreal lakes and biogeochemical implications” by Chidami and Amyot (2008). Limnology and Oceanography, 55(1):463–465, January 2010. CODEN LIOCAH. ISSN 0024-3590. See [CA08] and response [ACD10].

Stolpe:2010:NOC

[SH10b] Björn Stolpe and Martin Hassellöv. Nanofibrils and other colloidal biopolymers binding trace elements in coastal seawater: Significance for variations in element size distributions. Limnology and Oceanography, 55(1):187–202, January 2010. CODEN LIOCAH. ISSN 0024-3590.

Sunda:2011:IEL

[SH11] William G. Sunda and Susan A. Huntsman. Interactive effects of light and temperature on iron limitation in a marine diatom: Implications for marine productivity and carbon cycling. Limnology and Oceanography, 56(4):1475–1488, July 2011. CODEN LIOCAH. ISSN 0024-3590.

Sharp:2010:EOD

[Sha10] Jonathan H. Sharp. Estuarine oxygen dynamics: What can we learn about hypoxia from long-time records in the Delaware Estuary? Limnology and Oceanography, 55(2):535–548, March 2010. CODEN LIOCAH. ISSN 0024-3590.

Sinutok:2011:WMA

[SHD^{+}11] Sutinee Sinutok, Ross Hill, Martina A. Doblin, Richard Wuhrer, and Peter J. Ralph. Warmer more acidic conditions cause decreased productivity and calcification in subtropical coral reef sediment-dwelling calcifiers. Limnology and Oceanography, 56 (4):1200–1212, July 2011. CODEN LIOCAH. ISSN 0024-3590.

Sun:2011:ECP

[SHF^{+}11] Jun Sun, David A. Hutchins, Yuanyuan Feng, Erica L. Seubert, David A. Caron, and Fei-Xue Fu. Effects of changing pCO_{2} and phosphate availability on domoic acid production and physiology of the marine harmful bloom diatom Pseudo-nitzschia multiseries. Limnology and Oceanography, 56(3):829–840, May 2011. CODEN LIOCAH. ISSN 0024-3590.

Strzepek:2012:ILI

[SHF^{+}12] Robert F. Strzepek, Keith A. Hunter, Russell D. Frew, Paul J. Harrison, and Philip W. Boyd. Iron-light interactions differ in Southern Ocean phytoplankton. Limnology and Oceanography, 57(4):1182–1200, July 2012. CODEN LIOCAH. ISSN 0024-

Shen:2013:UTI

[SHK13] Jian Shen, Bo Hong, and Albert Y. Kuo. Using timescales to interpret dissolved oxygen distributions in the bottom waters of Chesapeake Bay. Limnology and Oceanography, 58(6):2237– 2248, November 2013. CODEN LIOCAH. ISSN 0024-3590.

Schmidt:2011:TIS

[SHKU11] Christiane Schmidt, Petra Heinz, Michal Kucera, and Sven Uthicke. Temperature-induced stress leads to bleaching in larger benthic Foraminifera hosting endosymbiotic diatoms. Limnology and Oceanography, 56(5):1587–1602, September 2011. CO-DEN LIOCAH. ISSN 0024-3590.

Su:2018:AAO

[SHL^{+}18] Y. Su, E. Hu, Z. Liu, E. Jeppesen, and J. J. Middelburg. Assimilation of ancient organic carbon by zooplankton in Tibetan Plateau lakes is depending on watershed characteristics. Limnology and Oceanography, 63(6):2359–2371, November 2018. CO-DEN LIOCAH. ISSN 0024-3590.

Schoutens:2019:HET

[SHM^{+}19] Ken Schoutens, Maike Heuner, Vanessa Minden, Tilla Schulte Ostermann, Alexandra Silinski, Jean-Philippe Belliard, and Stijn Temmerman. How effective are tidal marshes as naturebased shoreline protection throughout seasons? Limnology and Oceanography, 64(4):1750–1762, July 2019. CODEN LIOCAH. ISSN 0024-3590.

Sadro:2014:WVO

[SHSK14] Steven Sadro, Gordon W. Holtgrieve, Christopher T. Solomon, and Gregory R. Koch. Widespread variability in overnight patterns of ecosystem respiration linked to gradients in dissolved organic matter, residence time, and productivity in a global set of lakes. Limnology and Oceanography, 59(5):1666–1678, September 2014. CODEN LIOCAH. ISSN 0024-3590.

Schuback:2017:PPC

[SHT^{+}17] Nina Schuback, Clara J. M. Hoppe, Jean-Éric Tremblay, Maria T. Maldonado, and Philippe D. Tortell. Primary productivity and the coupling of photosynthetic electron transport and carbon fixation in the Arctic Ocean. Limnology and Oceanography, 62(3):898–921, May 2017. CODEN LIOCAH. ISSN 0024-

Schuback:2018:III

[SHT^{+}18] Nina Schuback, Clara J. M. Hoppe, Jean-Éric Tremblay, Maria T. Maldonado, and Philippe D. Tortell. Issue information — instr to contrib. Limnology and Oceanography, 62(3): 1445–1447, May 2018. CODEN LIOCAH. ISSN 0024-3590.

Shimizu:2010:SDE

[SI10] Kenji Shimizu and Jörg Imberger. Seasonal differences in the evolution of damped basin-scale internal waves in a shallow stratified lake. Limnology and Oceanography, 55(3):1449–1462, May 2010. CODEN LIOCAH. ISSN 0024-3590.

Shelley:2017:BMR

[SIH^{+}17] Felicity Shelley, Nicola Ings, Alan G. Hildrew, Mark Trimmer, and Jonathan Grey. Bringing methanotrophy in rivers out of the shadows. Limnology and Oceanography, 62(6):2345–2359, November 2017. CODEN LIOCAH. ISSN 0024-3590.

Shick:2011:RIL

[SIW^{+}11] J. Malcolm Shick, Katrina Iglic, Mark L. Wells, Charles G. Trick, Jason Doyle, and Walter C. Dunlap. Responses to iron limitation in two colonies of Stylophora pistillata exposed to high temperature: Implications for coral bleaching. Limnology and Oceanography, 56(3):813–828, May 2011. CODEN LIOCAH. ISSN 0024-3590.

Sewell:2011:SPD

[SJ11] Mary A. Sewell and Jennifer A. Jury. Seasonal patterns in diversity and abundance of the High Antarctic meroplankton: Plankton sampling using a Ross Sea desalination plant. Limnology and Oceanography, 56(5):1667–1681, September 2011. CODEN LIOCAH. ISSN 0024-3590.

Scharfenberger:2019:ETS

[SJB^{+}19] Ulrike Scharfenberger, Erik Jeppesen, Meryem Beklioğlu, Martin Søndergaard, David G. Angeler, Ayşe İdil Çakıroğlu, Stina Drakare, Josef Hejzlar, Aldoushy Mahdy, Eva Papastergiadou, Michal Šorf, Konstantinos Stefanidis, Arvo Tuvikene, Priit Zingel, and Rita Adrian. Effects of trophic status, water level, and temperature on shallow lake metabolism and metabolic balance: a standardized pan-European mesocosm experiment. Limnology and Oceanography, 64(2):616–631, March 2019. CODEN LIOCAH. ISSN 0024-3590.

Sand-Jensen:2011:USG

[SJM11] Kaj Sand-Jensen and Claus Lindskov Møller. Unprecedented slow growth and mortality of the rare colonial cyanobacterium, Nostoc zetterstedtii, in oligotrophic lakes. Limnology and Oceanography, 56(6):1976–1982, November 2011. CODEN LI-OCAH. ISSN 0024-3590.

Shatwell:2019:DNL

[SK19] T. Shatwell and J. Köhler. Decreased nitrogen loading controls summer cyanobacterial blooms without promoting nitrogenfixing taxa: Long-term response of a shallow lake. Limnology and Oceanography, 64(S1):S166–S178, January 2019. CODEN LIOCAH. ISSN 0024-3590.

Schepers:2017:STD

[SKGT17] Lennert Schepers, Matthew Kirwan, Glenn Guntenspergen, and Stijn Temmerman. Spatio-temporal development of vegetation die-off in a submerging coastal marsh. Limnology and Oceanography, 62(1):137–150, January 2017. CODEN LIOCAH. ISSN 0024-3590.

Sejr:2014:SDA

[SKJD^{+}14] Mikael K. Sejr, Dorte Krause-Jensen, Tage Dalsgaard, Sergio Ruiz-Halpern, Carlos M. Duarte, Mathias Middelboe, Ronne N. Glud, Jørgen Bendtsen, Thorsten J. S. Balsby, and Søren Rysgaard. Seasonal dynamics of autotrophic and heterotrophic plankton metabolism and P_{CO2} in a subarctic Greenland fjord. Limnology and Oceanography, 59(5):1764–1778, September 2014. CODEN LIOCAH. ISSN 0024-3590.

Shiozaki:2013:ATD

[SKK^{+}13] Takuhei Shiozaki, Taketoshi Kodama, Satoshi Kitajima, Mitsuhide Sato, and Ken Furuya. Advective transport of diazotrophs and importance of their nitrogen fixation on new and primary production in the western Pacific warm pool. Limnology and Oceanography, 58(2):49–60, January 2013. CODEN LIOCAH. ISSN 0024-3590.

Sanderman:2015:DSO

[SKK^{+}15] Jonathan Sanderman, Evelyn Krull, Thomas Kuhn, Gary Hancock, Janine McGowan, Todd Maddern, Stewart Fallon, and Andy Steven. Deciphering sedimentary organic matter sources: Insights from radiocarbon measurements and NMR spectroscopy. Limnology and Oceanography, 60(3):739–753, May 2015. CODEN LIOCAH. ISSN 0024-3590.

Soininen:2011:DSP

[SKKV11] Janne Soininen, Jenni J. Korhonen, Johanna Karhu, and Adrien Vetterli. Disentangling the spatial patterns in community composition of prokaryotic and eukaryotic lake plankton. Limnology and Oceanography, 56(2):508–520, March 2011. CODEN LIOCAH. ISSN 0024-3590.

Salter:2010:ABB

[SKLG10] I. Salter, A. E. S. Kemp, R. S. Lampitt, and M. Gledhill. The association between biogenic and inorganic minerals and the amino acid composition of settling particles. Limnology and Oceanography, 55(5):2207–2218, September 2010. CODEN LI-OCAH. ISSN 0024-3590.

Scherrer:2019:MMI

[SKV^{+}19] K. J. N. Scherrer, S. Kortsch, Ø. Varpe, G. A. Weyhenmeyer, B. Gulliksen, and R. Primicerio. Mechanistic model identifies increasing light availability due to sea ice reductions as cause for increasing macroalgae cover in the Arctic. Limnology and Oceanography, 64(1):330–341, January 2019. CODEN LIOCAH. ISSN 0024-3590.

Steward:2010:GEM

[SL10a] Joel S. Steward and Edgar F. Lowe. General empirical models for estimating nutrient load limits for Florida’s estuaries and inland waters. Limnology and Oceanography, 55(1):433–445, January 2010. CODEN LIOCAH. ISSN 0024-3590.

Stukel:2010:CPC

[SL10b] Michael R. Stukel and Michael R. Landry. Contribution of picophytoplankton to carbon export in the equatorial Pacific: a reassessment of food web flux inferences from inverse models. Limnology and Oceanography, 55(6):2669–2685, November 2010. CODEN LIOCAH. ISSN 0024-3590.

Seekell:2015:IDO

[SLA^{+}15] David A. Seekell, Jean-François Lapierre, Jenny Ask, Ann-Kristin Bergström, Anne Deininger, Patricia Rodríguez, and Jan Karlsson. The influence of dissolved organic carbon on primary production in northern lakes. Limnology and Oceanography, 60 (4):1276–1285, July 2015. CODEN LIOCAH. ISSN 0024-3590.

Selz:2018:IAC

[SLA^{+}18] Virginia Selz, Samuel Laney, Alexandra E. Arnsten, Kate M. Lewis, Kate E. Lowry, Hannah L. Joy-Warren, Matthew M. Mills, Gert L. van Dijken, and Kevin R. Arrigo. Ice algal communities in the Chukchi and Beaufort Seas in spring and early summer: Composition, distribution, and coupling with phytoplankton assemblages. Limnology and Oceanography, 63(3): 1109–1133, May 2018. CODEN LIOCAH. ISSN 0024-3590.

Simon-Lledo:2019:EPN

[SLBH^{+}19] Erik Simon-Lledó, Brian J. Bett, Veerle A. I. Huvenne, Timm Schoening, Noelie M. A. Benoist, and Daniel O. B. Jones. Ecology of a polymetallic nodule occurrence gradient: Implications for deep-sea mining. Limnology and Oceanography, 64(5):1883– 1894, September 2019. CODEN LIOCAH. ISSN 0024-3590.

Stukel:2011:TCC

[SLBNG11] Michael R. Stukel, Michael R. Landry, Claudia R. Benitez-Nelson, and Ralf Goericke. Trophic cycling and carbon export relationships in the California Current ecosystem. Limnology and Oceanography, 56(5):1866–1878, September 2011. CODEN LIOCAH. ISSN 0024-3590.

Saad:2016:UMD

[SLC^{+}16] Emily M. Saad, Amelia F. Longo, Luke R. Chambers, Rixiang Huang, Claudia Benitez-Nelson, Sonya T. Dyhrman, Julia M. Diaz, Yuanzhi Tang, and Ellery D. Ingall. Understanding marine dissolved organic matter production: Compositional insights from axenic cultures of Thalassiosira pseudonana. Limnology and Oceanography, 61(6):2222–2233, November 2016. CODEN LIOCAH. ISSN 0024-3590.

Shamberger:2018:LVE

[SLC18] Kathryn E. F. Shamberger, Steven J. Lentz, and Anne L. Cohen. Low and variable ecosystem calcification in a coral reef lagoon under natural acidification. Limnology and Oceanography, 63 (2):714–730, March 2018. CODEN LIOCAH. ISSN 0024-3590.

Savage:2010:ELU

[SLE10] Candida Savage, Peter R. Leavitt, and Ragnar Elmgren. Effects of land use, urbanization, and climate variability on coastal eutrophication in the Baltic Sea. Limnology and Oceanography, 55 (3):1033–1046, May 2010. CODEN LIOCAH. ISSN 0024-3590.

Scheibling:2010:KSN

[SLG10] Robert E. Scheibling and Jean-Sébastien Lauzon-Guay. Killer storms: North Atlantic hurricanes and disease outbreaks in sea urchins. Limnology and Oceanography, 55(6):2331–2338, November 2010. CODEN LIOCAH. ISSN 0024-3590.

Simon:2014:TEA

[SLG^{+}14] Heike Simon, Yvonne A. Lipsewers, Helge-Ansgar Giebel, Karen H. Wiltshire, and Meinhard Simon. Temperature effects on aggregation during a spring diatom bloom. Limnology and Oceanography, 59(6):2089–2100, November 2014. CODEN LI-OCAH. ISSN 0024-3590.

Shi:2015:IEL

[SLH^{+}15] Dalin Shi, Weiying Li, Brian M. Hopkinson, Haizheng Hong, Dongmei Li, Shuh-Ji Kao, and Wenfang Lin. Interactive effects of light, nitrogen source, and carbon dioxide on energy metabolism in the diatom Thalassiosira pseudonana. Limnology and Oceanography, 60(5):1805–1822, September 2015. CODEN LIOCAH. ISSN 0024-3590.

Schmidt:2019:DLP

[SLHA19] Silke R. Schmidt, Gunnar Lischeid, Thomas Hintze, and Rita Adrian. Disentangling limnological processes in the timefrequency domain. Limnology and Oceanography, 64(2):423–440, March 2019. CODEN LIOCAH. ISSN 0024-3590.

Schelske:2010:HAD

[SLK^{+}10] Claire L. Schelske, Edgar F. Lowe, William F. Kenney, Lawrence E. Battoe, Mark Brenner, and Michael F. Coveney. How anthropogenic darkening of Lake Apopka induced benthic light limitation and forced the shift from macrophyte to phytoplankton dominance. Limnology and Oceanography, 55(3):1201– 1212, May 2010. CODEN LIOCAH. ISSN 0024-3590.

Sawyer:2014:SCF

[SLK^{+}14] Audrey H. Sawyer, Olesya Lazareva, Kevin D. Kroeger, Kyle Crespo, Clara S. Chan, Thomas Stieglitz, and Holly A. Michael. Stratigraphic controls on fluid and solute fluxes across the sediment–water interface of an estuary. Limnology and Oceanography, 59(3):997–1010, May 2014. CODEN LIOCAH. ISSN 0024-3590.

Seekell:2014:RSV

[SLP^{+}14] David A. Seekell, Jean-François Lapierre, Michael L. Pace, Cristian Gudasz, Sebastian Sobek, and Lars J. Tranvik. Regionalscale variation of dissolved organic carbon concentrations in Swedish lakes. Limnology and Oceanography, 59(5):1612–1620, September 2014. CODEN LIOCAH. ISSN 0024-3590.

Simpson:2015:DMD

[SLPM15] John H. Simpson, Natasha S. Lucas, Ben Powell, and Stephen C. Maberly. Dissipation and mixing during the onset of stratification in a temperate lake, Windermere. Limnology and Oceanography, 60(1):29–41, January 2015. CODEN LIOCAH. ISSN 0024-3590.

Sharon:2011:PSH

[SLS^{+}11] Yoni Sharon, Orly Levitan, Dina Spungin, Ilana Berman-Frank, and Sven Beer. Photoacclimation of the seagrass Halophila stipulacea to the dim irradiance at its 48-meter depth limit. Limnology and Oceanography, 56(1):357–362, January 2011. CODEN LIOCAH. ISSN 0024-3590.

Sal:2011:CTN

[SLU11] Sofía Sal and Ángel López-Urrutia. Comment: Temperature, nutrients, and the size-scaling of phytoplankton growth in the sea. Limnology and Oceanography, 56(5):1952–1955, September 2011. CODEN LIOCAH. ISSN 0024-3590. See comment [CL11].

Scott:2010:NFM

[SM10] J. Thad Scott and Mark J. McCarthy. Nitrogen fixation may not balance the nitrogen pool in lakes over timescales relevant to eutrophication management. Limnology and Oceanography, 55 (3):1265–1270, May 2010. CODEN LIOCAH. ISSN 0024-3590.

Scheef:2011:MCR

[SM11a] Lindsay P. Scheef and Nancy H. Marcus. Mechanisms for copepod resting egg accumulation in seagrass sediments. Limnology and Oceanography, 56(1):363–370, January 2011. CODEN LI-OCAH. ISSN 0024-3590.

Scott:2011:RCN

[SM11b] J. Thad Scott and Mark J. McCarthy. Response to comment: Nitrogen fixation has not offset declines in the Lake 227 nitrogen pool and shows that nitrogen control deserves consideration in aquatic ecosystems. Limnology and Oceanography, 56 (4):1548–1550, July 2011. CODEN LIOCAH. ISSN 0024-3590. See [PSH^{+}11].

Scharfenberger:2013:TDS

[SMA13] Ulrike Scharfenberger, Aldoushy Mahdy, and Rita Adrian. Threshold-driven shifts in two copepod species: Testing ecological theory with observational data. Limnology and Oceanography, 58(2):741–752, March 2013. CODEN LIOCAH. ISSN 0024-3590.

Steel:2015:MCS

[SMA15] H. C. B. Steel, C. P. McKay, and D. T. Andersen. Modeling circulation and seasonal fluctuations in perennially ice-covered and ice-walled Lake Untersee, Antarctica. Limnology and Oceanography, 60(4):1139–1155, July 2015. CODEN LIOCAH. ISSN 0024-3590.

Souza:2010:SDC

[SMC^{+}10] M. S. Souza, B. E. Modenutti, P. Carrillo, M. Villar-Argaiz, J. M. Medina-Sánchez, F. Bullejos, and E. G. Balseiro. Stoichiometric dietary constraints influence the response of copepods to ultraviolet radiation-induced oxidative stress. Limnology and Oceanography, 55(3):1024–1032, May 2010. CODEN LIOCAH. ISSN 0024-3590.

Saunders:2010:IWT

[SMF10] Megan I. Saunders, Anna Metaxas, and Ramón Filgueira. Implications of warming temperatures for population outbreaks of a nonindigenous species (Membranipora membranacea, Bryozoa) in rocky subtidal ecosystems. Limnology and Oceanography, 55 (4):1627–1642, July 2010. CODEN LIOCAH. ISSN 0024-3590.

Smyntek:2012:DCD

[SMG12] Peter M. Smyntek, Stephen C. Maberly, and Jonathan Grey. Dissolved carbon dioxide concentration controls baseline stable carbon isotope signatures of a lake food web. Limnology and Oceanography, 57(5):1292–1302, September 2012. CODEN LI-OCAH. ISSN 0024-3590.

Strzepek:2011:ASS

[SMH^{+}11] Robert F. Strzepek, Maria T. Maldonado, Keith A. Hunter, Russell D. Frew, and Philip W. Boyd. Adaptive strategies by Southern Ocean phytoplankton to lessen iron limitation: Uptake of organically complexed iron and reduced cellular iron requirements. Limnology and Oceanography, 56(6):1983–2002, November 2011. CODEN LIOCAH. ISSN 0024-3590.

Santos:2019:COO

[SML^{+}19] Isaac R. Santos, Damien T. Maher, Reece Larkin, Jackie R. Webb, and Christian J. Sanders. Carbon outwelling and outgassing vs. burial in an estuarine tidal creek surrounded by mangrove and saltmarsh wetlands. Limnology and Oceanography, 64 (3):996–1013, May 2019. CODEN LIOCAH. ISSN 0024-3590.

Sanchez-Marin:2018:MDP

[SMLC^{+}18] Paula Sánchez-Marín, Fengjie Liu, Zhongzhi Chen, Claude Fortin, and Peter G. C. Campbell. Microalgal-driven pH changes in the boundary layer lead to apparent increases in Pb internalization by a unicellular alga in the presence of citrate. Limnology and Oceanography, 63(3):1328–1339, May 2018. CODEN LIOCAH. ISSN 0024-3590.

Sadro:2011:DIE

[SMM11] Steven Sadro, John M. Melack, and Sally MacIntyre. Depthintegrated estimates of ecosystem metabolism in a highelevation lake (Emerald Lake, Sierra Nevada, California). Limnology and Oceanography, 56(5):1764–1780, September 2011. CODEN LIOCAH. ISSN 0024-3590.

Sandhu:2019:ENF

[SMMF19] S. K. Sandhu, A. Yu. Morozov, A. Mitra, and K. Flynn. Exploring nonlinear functional responses of zooplankton grazers in dilution experiments via optimization techniques. Limnology and Oceanography, 64(2):774–784, March 2019. CODEN LIOCAH. ISSN 0024-3590.

Shikata:2015:DVM

[SMN^{+}15] Tomoyuki Shikata, Shigeru Matsunaga, Hiroyo Nishide, Setsuko Sakamoto, Goh Onistuka, and Mineo Yamaguchi. Diurnal vertical migration rhythms and their photoresponse in four phytoflagellates causing harmful algal blooms. Limnology and Oceanography, 60(4):1251–1264, July 2015. CODEN LIOCAH. ISSN 0024-3590.

Shilova:2017:DEN

[SMR^{+}17] I. N. Shilova, M. M. Mills, J. C. Robidart, K. A. Turk-Kubo, K. M. Björkman, Z. Kolber, I. Rapp, G. L. van Dijken, M. J. Church, K. R. Arrigo, E. P. Achterberg, and J. P. Zehr. Differential effects of nitrate, ammonium, and urea as N sources for microbial communities in the North Pacific Ocean. Limnology and Oceanography, 62(6):2550–2574, November 2017. CODEN LIOCAH. ISSN 0024-3590.

Skrabal:2018:PCR

[SMW^{+}18] Stephen A. Skrabal, Alyssa M. McBurney, Linda A. Webb, G. Brooks Avery, Jr., Robert J. Kieber, and Ralph N. Mead. Photodissolution of copper from resuspended coastal marine sediments. Limnology and Oceanography, 63(2):773–785, March 2018. CODEN LIOCAH. ISSN 0024-3590.

Sweetman:2014:BEF

[SNG^{+}14] Andrew K. Sweetman, Karl Norling, Carina Gunderstad, Barbro T. Haugland, and Trine Dale. Benthic ecosystem functioning beneath fish farms in different hydrodynamic environments. Limnology and Oceanography, 59(4):1139–1151, July 2014. CO-DEN LIOCAH. ISSN 0024-3590.

Shatwell:2012:TPE

[SNK12] Tom Shatwell, Andreas Nicklisch, and Jan Köhler. Temperature and photoperiod effects on phytoplankton growing under simulated mixed layer light fluctuations. Limnology and Oceanography, 57(3):541–553, March 2012. CODEN LIOCAH. ISSN 0024-3590.

Sadro:2011:LDP

[SNM11] Steven Sadro, Craig E. Nelson, and John M. Melack. Linking diel patterns in community respiration to bacterioplankton in an oligotrophic high-elevation lake. Limnology and Oceanography, 56(2):540–550, March 2011. CODEN LIOCAH. ISSN 0024-3590.

Stoecker:2015:UMG

[SNM^{+}15] Diane K. Stoecker, Jens C. Nejstgaard, Rakhesh Madhusoodhanan, Georg Pohnert, Stefanie Wolfram, Hans Henrik Jakobsen, Sigitas Šulčius, and Aud Larsen. Underestimation of microzooplankton grazing in dilution experiments due to inhibition of phytoplankton growth. Limnology and Oceanography, 60(4): 1426–1438, July 2015. CODEN LIOCAH. ISSN 0024-3590.

Saros:2016:TSS

[SNO^{+}16] Jasmine E. Saros, Robert M. Northington, Christopher L. Osburn, Benjamin T. Burpee, and Nicholas John Anderson. Thermal stratification in small Arctic lakes of southwest Greenland affected by water transparency and epilimnetic temperatures. Limnology and Oceanography, 61(4):1530–1542, July 2016. CO-DEN LIOCAH. ISSN 0024-3590.

Sichlau:2015:MSS

[SNTK15] Mie H. Sichlau, Einar E. Nielsen, Uffe H. Thygesen, and Thomas Kiørboe. Mating success and sexual selection in a pelagic copepod, Temora longicornis: Evidence from paternity analyses. Limnology and Oceanography, 60(2):600–610, March 2015. CO-DEN LIOCAH. ISSN 0024-3590.

Santoro:2010:CSS

[SNvD^{+}10] Alyson E. Santoro, Nicholas J. Nidzieko, Gert L. van Dijken, Kevin R. Arrigo, and Alexandria B. Boehma. Contrasting spring and summer phytoplankton dynamics in the nearshore Southern California Bight. Limnology and Oceanography, 55(1):264–278, January 2010. CODEN LIOCAH. ISSN 0024-3590.

Simek:2014:FTS

[ŠNZ^{+}14] Karel Šimek, Jirří Nedoma, Petr Znachor, Vojteěch Kasalický, Jan Jezbera, Karel Hornñák, and Jaromír Sed’a. A finely tuned symphony of factors modulates the microbial food web of a freshwater reservoir in spring. Limnology and Oceanography, 59(5): 1477–1492, September 2014. CODEN LIOCAH. ISSN 0024-3590.

Spyrakos:2018:OTI

[SOH^{+}18] Evangelos Spyrakos, Ruth O’Donnell, Peter D. Hunter, Claire Miller, Marian Scott, Stefan G. H. Simis, Claire Neil, Claudio C. F. Barbosa, Caren E. Binding, Shane Bradt, Mariano Bresciani, Giorgio Dall’Olmo, Claudia Giardino, Anatoly A. Gitelson, Tiit Kutser, Lin Li, Bunkei Matsushita, Victor Martinez-Vicente, Mark W. Matthews, Igor Ogashawara, Antonio Ruiz-Verdú, John F. Schalles, Emma Tebbs, Yunlin Zhang, and Andrew N. Tyler. Optical types of inland and coastal waters. Limnology and Oceanography, 63(2):846–870, March 2018. CODEN LIOCAH. ISSN 0024-3590.

So:2015:IME

[SOM^{+}15] Mika So, Hajime Ohtsuki, Wataru Makino, Seiji Ishida, Hitoshi Kumagai, Kenyu G. Yamaki, and Jotaro Urabe. Invasion and molecular evolution of Daphnia pulex in Japan. Limnology and Oceanography, 60(4):1129–1138, July 2015. CODEN LIOCAH. ISSN 0024-3590.

Schmid:2017:RGE

[SOM17] Martin Schmid, Ilia Ostrovsky, and Daniel F. McGinnis. Role of gas ebullition in the methane budget of a deep subtropical lake: What can we learn from process-based modeling? Limnology and Oceanography, 62(6):2674–2698, November 2017. CODEN LIOCAH. ISSN 0024-3590.

Sterner:2017:GCR

[SOO^{+}17] Robert W. Sterner, Peggy Ostrom, Nathaniel E. Ostrom, J. Val Klump, Alan D. Steinman, Erin A. Dreelin, M. Jake Vander Zanden, and Aaron T. Fisk. Grand challenges for research in the Laurentian Great Lakes. Limnology and Oceanography, 62 (6):2510–2523, November 2017. CODEN LIOCAH. ISSN 0024-3590.

Seekell:2011:DPD

[SP11] David A. Seekell and Michael L. Pace. Does the Pareto distribution adequately describe the size-distribution of lakes? Limnology and Oceanography, 56(1):350–356, January 2011. CODEN LIOCAH. ISSN 0024-3590.

Sen:2014:CSH

[SPB^{+}14] Arunima Sen, Elizabeth L. Podowski, Erin L. Becker, Erica A. Shearer, Amy Gartman, Mustafa Yücel, Stéphane Hourdez, George W. Luther III, and Charles R. Fisher. Community succession in hydrothermal vent habitats of the Eastern Lau Spreading Center and Valu Fa Ridge, Tonga. Limnology and Oceanography, 59(5):1510–1528, September 2014. CODEN LI-OCAH. ISSN 0024-3590.

Salcher:2011:VLD

[SPFP11] Michaela M. Salcher, Jakob Pernthaler, Nakita Frater, and Thomas Posch. Vertical and longitudinal distribution patterns of different bacterioplankton populations in a canyon-shaped, deep prealpine lake. Limnology and Oceanography, 56(6):2027– 2039, November 2011. CODEN LIOCAH. ISSN 0024-3590.

Steger:2011:MBC

[SPG^{+}11] Kristin Steger, Katrin Premke, Cristian Gudasz, Ingvar Sundh, and Lars J. Tranvik. Microbial biomass and community composition in boreal lake sediments. Limnology and Oceanography, 56 (2):725–733, March 2011. CODEN LIOCAH. ISSN 0024-3590.

Seibt:2013:MWW

[SPG^{+}13] Christian Seibt, Frank Peeters, Michael Graf, Michael Sprenger, and Hilmar Hofmann. Modeling wind waves and wave exposure of nearshore zones in medium-sized lakes. Limnology and Oceanography, 58(1):23–36, January 2013. CODEN LIOCAH. ISSN 0024-3590.

Soria-Piriz:2017:SFP

[SPGRP^{+}17] Sara Soria-Píriz, Emilio García-Robledo, Sokratis Papaspyrou, Virginia Aguilar, Isabel Seguro, Jenaro Acuña, Álvaro Morales, and Alfonso Corzo. Size fractionated phytoplankton biomass and net metabolism along a tropical estuarine gradient. Limnology and Oceanography, 62(S1):S309–S326, November 2017. CODEN LIOCAH. ISSN 0024-3590.

Serra-Pompei:2019:RLD

[SPHVA19] Camila Serra-Pompei, George I. Hagstrom, André W. Visser, and Ken H. Andersen. Resource limitation determines temperature response of unicellular plankton communities. Limnology and Oceanography, 64(4):1627–1640, July 2019. CODEN LI-OCAH. ISSN 0024-3590.

Spivak:2015:BBR

[Spi15] Amanda C. Spivak. Benthic biogeochemical responses to changing estuary trophic state and nutrient availability: a paired field and mesocosm experiment approach. Limnology and Oceanography, 60(1):3–21, January 2015. CODEN LIOCAH. ISSN 0024-3590.

Smith:2011:OBM

[SPMW11] S. Lan Smith, Markus Pahlow, Agostino Merico, and Kai W. Wirtz. Optimality-based modeling of planktonic organisms. Limnology and Oceanography, 56(6):2080–2094, November 2011. CODEN LIOCAH. ISSN 0024-3590.

Spigel:2018:PLP

[SPO^{+}18] Robert H. Spigel, John C. Priscu, Maciej K. Obryk, William Stone, and Peter T. Doran. The physical limnology of a permanently ice-covered and chemically stratified Antarctic lake using high resolution spatial data from an autonomous underwater vehicle. Limnology and Oceanography, 63(3):1234–1252, May 2018. CODEN LIOCAH. ISSN 0024-3590.

Salcher:2010:SDA

[SPP10] Michaela M. Salcher, Jakob Pernthaler, and Thomas Posch. Spatiotemporal distribution and activity patterns of bacteria from three phylogenetic groups in an oligomesotrophic lake. Limnology and Oceanography, 55(4):846–856, March 2010. CODEN LIOCAH. ISSN 0024-3590.

Sciberras:2016:IBF

[SPP^{+}16] Marija Sciberras, Ruth Parker, Claire Powell, Craig Robertson, Silke Kröger, Stefan Bolam, and Jan Geert Hiddink. Impacts of bottom fishing on the sediment infaunal community and biogeochemistry of cohesive and non-cohesive sediments. Limnology and Oceanography, 61(6):2076–2089, November 2016. CODEN LIOCAH. ISSN 0024-3590.

Stief:2010:CNO

[SPPS10] Peter Stief, Lubos Polerecky, Morten Poulsen, and Andreas Schramm. Control of nitrous oxide emission from Chironomus plumosus larvae by nitrate and temperature. Limnology and Oceanography, 55(4):872–884, March 2010. CODEN LIOCAH. ISSN 0024-3590.

Stamieszkin:2015:SMT

[SPR^{+}15] Karen Stamieszkin, Andrew J. Pershing, Nicholas R. Record, Cynthia H. Pilskaln, Hans G. Dam, and Leah R. Feinberg. Size as the master trait in modeled copepod fecal pellet carbon flux. Limnology and Oceanography, 60(6):2090–2107, November 2015. CODEN LIOCAH. ISSN 0024-3590.

Salter:2019:CCP

[SPS19] Michael A. Salter, Chris T. Perry, and Abigail M. Smith. Calcium carbonate production by fish in temperate marine environments. Limnology and Oceanography, 64(6):2755–2770, November 2019. CODEN LIOCAH. ISSN 0024-3590.

Salonen:2014:IVC

[SPSG14] Kalevi Salonen, Merja Pulkkanen, Pauliina Salmi, and Ross W. Griffiths. Interannual variability of circulation under spring ice in a boreal lake. Limnology and Oceanography, 61(1):2121–2132, November 2014. CODEN LIOCAH. ISSN 0024-3590.

Shaw:2015:NSR

[SPTS15] Emily C. Shaw, Stuart R. Phinn, Bronte Tilbrook, and Andy Steven. Natural in situ relationships suggest coral reef calcium carbonate production will decline with ocean acidification. Limnology and Oceanography, 60(3):777–788, May 2015. CODEN LIOCAH. ISSN 0024-3590.

Steinman:2010:IHRb

[SRA10] Byron A. Steinman, Michael F. Rosenmeier, and Mark B. Abbott. The isotopic and hydrologic response of small, closed-basin lakes to climate forcing from predictive models: Simulations of stochastic and mean state precipitation variations. Limnology and Oceanography, 55(6):2246–2261, November 2010. CODEN LIOCAH. ISSN 0024-3590.

Steinman:2010:IHRa

[SRAB10] Byron A. Steinman, Michael F. Rosenmeier, Mark B. Abbott, and Daniel J. Bain. The isotopic and hydrologic response of small, closed-basin lakes to climate forcing from predictive models: Application to paleoclimate studies in the upper Columbia River basin. Limnology and Oceanography, 55(6):2231–2245, November 2010. CODEN LIOCAH. ISSN 0024-3590.

Sarmento:2013:PSS

[SRCL^{+}13] Hugo Sarmento, Cristina Romera-Castillo, Markus Lindh, Jarone Pinhassi, M. Montserrat Sala, Josep M. Gasol, Cèlia Marrase, and Gordon T. Taylor. Phytoplankton species-specific release of dissolved free amino acids and their selective consumption by bacteria. Limnology and Oceanography, 58(3):1123– 1135, May 2013. CODEN LIOCAH. ISSN 0024-3590.

Sitta:2018:INF

[SRM^{+}18] Kimberly A. Sitta, Michelle Reed, Rebecca Mortensen, Cameron Doll, Timothy Callahan, and Dianne I. Greenfield. The influences of nitrogen form and zooplankton grazing on phytoplankton assemblages in two coastal southeastern systems. Limnology and Oceanography, 63(6):2523–2544, November 2018. CODEN LIOCAH. ISSN 0024-3590.

Slemmons:2012:INR

[SS12a] Krista E. H. Slemmons and Jasmine E. Saros. Implications of nitrogen-rich glacial meltwater for phytoplankton diversity and productivity in alpine lakes. Limnology and Oceanography, 57 (6):1651–1663, November 2012. CODEN LIOCAH. ISSN 0024-3590.

Smith:2012:EIS

[SS12b] Christopher G. Smith and Peter W. Swarzenski. Erratum: an investigation of submarine groundwater-borne nutrient fluxes to the west Florida shelf and recurrent harmful algal blooms. Limnology and Oceanography, 53(2):896, May 2012. CODEN LI-OCAH. ISSN 0024-3590.

Smith:2012:ISG

[SS12c] Christopher G. Smith and Peter W. Swarzenski. An investigation of submarine groundwater-borne nutrient fluxes to the west Florida shelf and recurrent harmful algal blooms. Limnology and Oceanography, 57(2):471–485, March 2012. CODEN LIOCAH. ISSN 0024-3590.

Salmi:2016:RBS

[SS16] Pauliina Salmi and Kalevi Salonen. Regular build-up of the spring phytoplankton maximum before ice-break in a boreal lake. Limnology and Oceanography, 61(1):240–253, January 2016. CODEN LIOCAH. ISSN 0024-3590.

Sugie:2017:CSS

[SS17] Koji Sugie and Koji Suzuki. Characterization of the synopticscale diversity, biogeography, and size distribution of diatoms in the North Pacific. Limnology and Oceanography, 62(3):884–897, May 2017. CODEN LIOCAH. ISSN 0024-3590.

Stachelek:2019:DFC

[SS19] Jemma Stachelek and Patricia A. Soranno. Does freshwater connectivity influence phosphorus retention in lakes? Limnology and Oceanography, 64(4):1586–1599, July 2019. CODEN LIOCAH. ISSN 0024-3590.

Steinle:2016:LSW

[SSB^{+}16] Lea Steinle, Mark Schmidt, Lee Bryant, Matthias Haeckel, Peter Linke, Stefan Sommer, Jakob Zopfi, Moritz F. Lehmann, Tina Treude, and Helge Niemannn. Linked sediment and watercolumn methanotrophy at a man-made gas blowout in the North Sea: Implications for methane budgeting in seasonally stratified shallow seas. Limnology and Oceanography, 61(S1):S367–S386, November 2016. CODEN LIOCAH. ISSN 0024-3590.

Schwefel:2018:USS

[SSB^{+}18] Robert Schwefel, Thomas Steinsberger, Damien Bouffard, Lee D. Bryant, Beat Müller, and Alfred Wüest. Using smallscale measurements to estimate hypolimnetic oxygen depletion in a deep lake. Limnology and Oceanography, 63(S1):S54–S67, March 2018. CODEN LIOCAH. ISSN 0024-3590.

Stubbins:2010:IDM

[SSC^{+}10] Aron Stubbins, Robert G. M. Spencer, Hongmei Chen, Patrick G. Hatcher, Kenneth Mopper, Peter J. Hernes, Vincent L. Mwamba, Arthur M. Mangangu, Jose N. Wabakanghanzi, and Johan Six. Illuminated darkness: Molecular signatures of Congo River dissolved organic matter and its photochemical alteration as revealed by ultrahigh precision mass spectrometry. Limnology and Oceanography, 55(4):1467–1477, July 2010. CODEN LIOCAH. ISSN 0024-3590.

Suter:2017:NBS

[SSC^{+}17] Elizabeth A. Suter, Mary I. Scranton, Stephanie Chow, Dallyce Stinton, Luis Medina Faull, and Gordon T. Taylor. Niskin bottle sample collection aliases microbial community composition and biogeochemical interpretation. Limnology and Oceanography, 62 (2):606–617, March 2017. CODEN LIOCAH. ISSN 0024-3590.

Sampei:2012:SEF

[SSFF12] Makoto Sampei, Hiroshi Sasaki, Alexandre Forest, and Louis Fortier. A substantial export flux of particulate organic carbon linked to sinking dead copepods during winter 2007–2008 in the Amundsen Gulf (southeastern Beaufort Sea, Arctic Ocean). Limnology and Oceanography, 57(1):90–96, January 2012. CO-DEN LIOCAH. ISSN 0024-3590.

Strayer:2019:LTR

[SSFR19] David L. Strayer, Christopher T. Solomon, Stuart E. G. Findlay, and Emma J. Rosi. Long-term research reveals multiple relationships between the abundance and impacts of a non-native species. Limnology and Oceanography, 64(S1):S105–S117, January 2019. CODEN LIOCAH. ISSN 0024-3590.

Santoro:2017:TED

[SSG^{+}17] Alyson E. Santoro, Mak A. Saito, Tyler J. Goepfert, Carl H. Lamborg, Chris L. Dupont, and Giacomo R. DiTullio. Thaumarchaeal ecotype distributions across the equatorial Pacific Ocean and their potential roles in nitrification and sinking flux attenuation. Limnology and Oceanography, 65(9):1984–2003, September 2017. CODEN LIOCAH. ISSN 0024-3590.

Seitaj:2017:SOD

[SSGB^{+}17] Dorina Seitaj, Fatimah Sulu-Gambari, Laurine D. W. Burdorf, Alicia Romero-Ramirez, Olivier Maire, Sairah Y. Malkin, Caroline P. Slomp, and Filip J. R. Meysman. Sedimentary oxygen dynamics in a seasonally hypoxic basin. Limnology and Oceanography, 62(2):452–473, March 2017. CODEN LIOCAH. ISSN 0024-3590.

Swarbrick:2019:DSS

[SSGL19] Vanessa J. Swarbrick, Gavin L. Simpson, Patricia M. Glibert, and Peter R. Leavitt. Differential stimulation and suppression of phytoplankton growth by ammonium enrichment in eutrophic hardwater lakes over 16 years. Limnology and Oceanography, 64 (S1):S130–S149, January 2019. CODEN LIOCAH. ISSN 0024-3590.

Stukel:2018:RSG

[SSGM18] Michael R. Stukel, Hajoon Song, Ralf Goericke, and Arthur J. Miller. The role of subduction and gravitational sinking in particle export, carbon sequestration, and the remineralization length scale in the California Current ecosystem. Limnology and Oceanography, 63(4):363–383, January 2018. CODEN LIOCAH. ISSN 0024-3590.

Smith:2014:LSB

[SSH^{+}14] K. L. Smith, Jr., A. D. Sherman, C. L. Huffard, P. R. McGill, R. Henthorn, S. Von Thun, H. A. Ruhl, M. Kahru, and M. D. Ohman. Large salp bloom export from the upper ocean and benthic community response in the abyssal northeast Pacific: Day to week resolution. Limnology and Oceanography, 59(3): 745–757, May 2014. CODEN LIOCAH. ISSN 0024-3590.

Schoen:2016:RFM

[SSH^{+}16] Lee S. Schoen, James J. Student, Joel C. Hoffman, Michael E. Sierszen, and Donald G. Uzarski. Reconstructing fish movements between coastal wetland and nearshore habitats of the Great Lakes. Limnology and Oceanography, 62(3):1800–1813, September 2016. CODEN LIOCAH. ISSN 0024-3590.

Staehr:2010:DMN

[SSJR^{+}10] Peter A. Staehr, Kaj Sand-Jensen, Ane L. Raun, Bertel Nilsson, and Jacob Kidmose. Drivers of metabolism and net heterotrophy in contrasting lakes. Limnology and Oceanography, 55(4):817– 830, March 2010. CODEN LIOCAH. ISSN 0024-3590.

Schreiber:2014:NOT

[SSKdB14] Frank Schreiber, Peter Stief, Marcel M. M. Kuypers, and Dirk de Beer. Nitric oxide turnover in permeable river sediment. Limnology and Oceanography, 59(4):1310–1320, July 2014. CODEN LIOCAH. ISSN 0024-3590.

Shapiro:2012:ATG

[SSL^{+}12] Karen Shapiro, Mary W. Silver, John L. Largier, Patricia A. Conrad, and Jonna A. K. Mazet. Association of Toxoplasma gondii oocysts with fresh, estuarine, and marine macroaggregates. Limnology and Oceanography, 57(2):449–456, March 2012. CODEN LIOCAH. ISSN 0024-3590.

Strock:2019:RBL

[SSM^{+}19] Kristin E. Strock, Jasmine E. Saros, Suzanne McGowan, Mark B. Edlund, and Daniel R. Engstrom. Response of boreal lakes to changing wind strength: Coherent physical changes across two large lakes but varying effects on primary producers over the 20^{th} century. Limnology and Oceanography, 66(4):2237– 2251, September 2019. CODEN LIOCAH. ISSN 0024-3590.

Stevens:2012:IWD

[SSN12] Craig L. Stevens, Philip J. H. Sutton, and Cliff S. Law N. Internal waves downstream of Norfolk Ridge, western Pacific, and their biophysical implications. Limnology and Oceanography, 57 (4):897–911, July 2012. CODEN LIOCAH. ISSN 0024-3590.

Sulcius:2017:PPS

[ŠSP17] Sigitas Šulčius, Kristina Slavuckytė, and Ričardas Paškauskas. The predation paradox: Synergistic and antagonistic interactions between grazing by crustacean predator and infection by cyanophages promotes bloom formation in filamentous cyanobacteria. Limnology and Oceanography, 65(9):2189–2199, September 2017. CODEN LIOCAH. ISSN 0024-3590.

Silinski:2018:CWP

[SSP^{+}18] Alexandra Silinski, Ken Schoutens, Sara Puijalon, Jonas Schoelynck, Deirdre Luyckx, Peter Troch, Patrick Meire, and Stijn Temmerman. Coping with waves: Plasticity in tidal marsh plants as self-adapting coastal ecosystem engineers. Limnology and Oceanography, 63(2):799–815, March 2018. CODEN LIOCAH. ISSN 0024-3590.

Smyth:2012:PPR

[SSPK^{+}12] Robyn L. Smyth, Cristina Sobrino, Jesse Phillips-Kress, Hae-Cheol Kim, and Patrick J. Neale. Phytoplankton photosynthetic response to solar ultraviolet irradiance in the Ross Sea polynya: Development and evaluation of a time-dependent model with limited repair. Limnology and Oceanography, 57(6):1602–1618, November 2012. CODEN LIOCAH. ISSN 0024-3590.

Saxton:2016:BRG

[SSS^{+}16] Matthew A. Saxton, Vladimir A. Samarkin, Charles A. Schutte, Marshall W. Bowles, Michael T. Madigan, Sarah B. Cadieux, Lisa M. Pratt, and Samantha B. Joye. Biogeochemical and ^{16}S rRNA gene sequence evidence supports a novel mode of anaerobic methanotrophy in permanently ice-covered Lake Fryxell, Antarctica. Limnology and Oceanography, 61(S1):S119–S130, November 2016. CODEN LIOCAH. ISSN 0024-3590.

Sweetman:2019:KRB

[SSS^{+}19] Andrew K. Sweetman, Craig R. Smith, Christine N. Shulse, Brianne Maillot, Markus Lindh, Matthew J. Church, Kirstin S. Meyer, Dick van Oevelen, Tanja Stratmann, and Andrew J. Gooday. Key role of bacteria in the short-term cycling of carbon at the abyssal seafloor in a low particulate organic carbon flux region of the eastern Pacific Ocean. Limnology and Oceanography, 64(2):694–713, March 2019. CODEN LIOCAH. ISSN 0024-3590.

Schelker:2016:CES

[SSU^{+}16] Jakob Schelker, Gabriel A. Singer, Amber J. Ulseth, Sabrina Hengsberger, and Tom J. Battin. CO_{2} evasion from a steep, high gradient stream network: importance of seasonal and diurnal variation in aquatic pCO_{2} and gas transfer. Limnology and Oceanography, 62(3):1826–1838, September 2016. CODEN LIOCAH. ISSN 0024-3590.

Sommer:2019:RDD

[SSW19] Tobias Sommer, Martin Schmid, and Alfred Wüest. The role of double diffusion for the heat and salt balance in Lake Kivu. Limnology and Oceanography, 64(2):650–660, March 2019. CODEN LIOCAH. ISSN 0024-3590.

Sugimoto:2014:USN

[SSYT14] Ryo Sugimoto, Takahisa Sato, Takehito Yoshida, and Osamu Tominaga. Using stable nitrogen isotopes to evaluate the relative importance of external and internal nitrogen loadings on phytoplankton production in a shallow eutrophic lake (Lake Mikata, Japan). Limnology and Oceanography, 59(1):37–47, January 2014. CODEN LIOCAH. ISSN 0024-3590.

Semeniuk:2016:ICI

[STB^{+}16] David M. Semeniuk, Rebecca L. Taylor, Randelle M. Bundy, W. Keith Johnson, Jay T. Cullen, Marie Robert, Katherine A. Barbeau, and Maria T. Maldonado. Iron–copper interactions in iron-limited phytoplankton in the northeast subarctic Pacific Ocean. Limnology and Oceanography, 61(1):279–297, January 2016. CODEN LIOCAH. ISSN 0024-3590.

Shadwick:2011:SVI

[STC^{+}11] E. H. Shadwick, H. Thomas, M. Chierici, B. Else, A. Fransson, C. Michel, L. A. Miller, A. Mucci, A. Niemi, T. N. Papakyriakou, and J.-É. Tremblay. Seasonal variability of the inorganic carbon system in the Amundsen Gulf region of the southeastern Beaufort Sea. Limnology and Oceanography, 56(1):303–322, January 2011. CODEN LIOCAH. ISSN 0024-3590.

Sartoris:2010:BDA

[STCS10] Franz Josef Sartoris, David N. Thomas, Astrid Cornils, and Sigrid B. Schnack Schiela. Buoyancy and diapause in Antarctic copepods: The role of ammonium accumulation. Limnology and Oceanography, 55(5):1860–1864, September 2010. CODEN LIOCAH. ISSN 0024-3590.

Stratmann:2018:RHP

[SVG^{+}18] Tanja Stratmann, Ilja Voorsmit, Andrey Gebruk, Alastair Brown, Autun Purser, Yann Marcon, Andrew K. Sweetman, Daniel O. B. Jones, and Dick van Oevelen. Recovery of Holothuroidea population density, community composition, and respiration activity after a deep-sea disturbance experiment. Limnology and Oceanography, 64(4):2140–2153, September 2018. CODEN LIOCAH. ISSN 0024-3590.

Soissons:2018:SLV

[SvKP^{+}18] L. M. Soissons, M. M. van Katwijk, G. Peralta, F. G. Brun, P. G. Cardoso, T. F. Grilo, B. Ondiviela, M. Recio, M. Valle, J. M. Garmendia, F. Ganthy, I. Auby, L. Rigouin, L. Godet, J. Fournier, N. Desroy, L. Barillé, P. Kadel, R. Asmus, P. M. J. Herman, and T. J. Bouma. Seasonal and latitudinal variation in seagrass mechanical traits across Europe: The influence of local nutrient status and morphometric plasticity. Limnology and Oceanography, 63(1):37–46, January 2018. CODEN LIOCAH. ISSN 0024-3590.

Samper-Villarreal:2016:OCS

[SVLS^{+}16] Jimena Samper-Villarreal, Catherine E. Lovelock, Megan I. Saunders, Chris Roelfsema, and Peter J. Mumby. Organic carbon in seagrass sediments is influenced by seagrass canopy complexity, turbidity, wave height, and water depth. Limnology and Oceanography, 61(3):938–952, May 2016. CODEN LIOCAH. ISSN 0024-3590.

Stark:2015:OTS

[SVMT15] J. Stark, T. Van Oyen, P. Meire, and S. Temmerman. Observations of tidal and storm surge attenuation in a large tidal marsh. Limnology and Oceanography, 60(4):1371–1381, July 2015. CO-DEN LIOCAH. ISSN 0024-3590.

Strand:2019:SMB

[SVS^{+}19] Kjersti Opstad Strand, Frode Vikebø, Svein Sundby, Ann Kristin Sperrevik, and øyvind Breivik. Subsurface maxima in buoyant fish eggs indicate vertical velocity shear and spatially limited spawning grounds. Limnology and Oceanography, 64(3):1239– 1251, May 2019. CODEN LIOCAH. ISSN 0024-3590.

Sperfeld:2011:TCI

[SW11] Erik Sperfeld and Alexander Wacker. Temperature- and cholesterol-induced changes in eicosapentaenoic acid limitation of Daphnia magna determined by a promising method to estimate growth saturation thresholds. Limnology and Oceanography, 56(4):1273–1284, July 2011. CODEN LIOCAH. ISSN 0024-3590.

Shi:2014:ORS

[SW14] Wei Shi and Menghua Wang. Ocean reflectance spectra at the red, near-infrared, and shortwave infrared from highly turbid waters: a study in the Bohai Sea, Yellow Sea, and East China Sea. Limnology and Oceanography, 59(2):427–444, March 2014. CODEN LIOCAH. ISSN 0024-3590.

Schmidt:2012:RSE

[SWCL12] Allison L. Schmidt, Jessica K. C. Wysmyk, Susanne E. Craig, and Heike K. Lotze. Regional-scale effects of eutrophication on ecosystem structure and services of seagrass beds. Limnology and Oceanography, 58(1):1389–1402, September 2012. CODEN LIOCAH. ISSN 0024-3590.

Smith:2011:RSP

[SWD11] Lydia Smith, Mary C. Watzin, and Gregory Druschel. Relating sediment phosphorus mobility to seasonal and diel redox fluctuations at the sediment–water interface in a eutrophic freshwater lake. Limnology and Oceanography, 56(6):2251–2264, November 2011. CODEN LIOCAH. ISSN 0024-3590.

Schloss:2014:PBC

[SWD^{+}14] I. R. Schloss, A. Wasilowska, D. Dumont, G. O. Almandoz, M. P. Hernando, C.-A. Michaud-Tremblay, L. Saravia, M. Rzepecki, P. Monien, D. Monien, E. E. Kopczyńska, A. V. Bers, and G. A. Ferreyra. On the phytoplankton bloom in coastal waters of southern King George Island (Antarctica) in January 2010: an exceptional feature? Limnology and Oceanography, 59 (1):195–210, January 2014. CODEN LIOCAH. ISSN 0024-3590.

Schutte:2018:DOP

[SWE^{+}18] Charles A. Schutte, Alicia M. Wilson, Tyler Evans, Willard S. Moore, and Samantha B. Joye. Deep oxygen penetration drives nitrification in intertidal beach sands. Limnology and Oceanography, 63(S1):S193–S208, March 2018. CODEN LIOCAH. ISSN 0024-3590.

Simpson:2011:ISM

[SWL11] J. H. Simpson, P. J. Wiles, and B. J. Lincoln. Internal seiche modes and bottom boundary-layer dissipation in a temperate lake from acoustic measurements. Limnology and Oceanography, 56(5):1893–1906, September 2011. CODEN LIOCAH. ISSN 0024-3590.

Schlappy:2010:HOR

[SWM^{+}10] M.-L. Schläppy, M. Weber, D. Mendola, F. Hoffmann, and D. de Beer. Heterogeneous oxygenation resulting from active and passive flow in two Mediterranean sponges, Dysida avara and Chondrosia reniformis. Limnology and Oceanography, 55 (3):1289–1300, May 2010. CODEN LIOCAH. ISSN 0024-3590.

Schmale:2018:CZM

[SWM^{+}18] Oliver Schmale, Janine Wäge, Volker Mohrholz, Norbert Wasmund, Ulf Gräwe, Gregor Rehder, Matthias Labrenz, and Natalie Loick-Wilde. The contribution of zooplankton to methane supersaturation in the oxygenated upper waters of the central Baltic Sea. Limnology and Oceanography, 63(4):412–430, January 2018. CODEN LIOCAH. ISSN 0024-3590.

Small:2011:DPD

[SWP11] Gaston E. Small, John P. Wares, and Catherine M. Pringle. Differences in phosphorus demand among detritivorous chironomid larvae reflect intraspecific adaptations to differences in food resource stoichiometry across lowland tropical streams. Limnology and Oceanography, 56(1):268–278, January 2011. CODEN LIOCAH. ISSN 0024-3590.

Scicluna:2015:DDP

[SWZ^{+}15] Todd R. Scicluna, Ryan J. Woodland, Yafei Zhu, Michael R. Grace, and Perran L. M. Cook. Deep dynamic pools of phosphorus in the sediment of a temperate lagoon with recurring blooms of diazotrophic cyanobacteria. Limnology and Oceanography, 60 (6):2185–2196, November 2015. CODEN LIOCAH. ISSN 0024-3590.

Schwaderer:2011:EED

[SYdTP^{+}11] Anne S. Schwaderer, Kohei Yoshiyama, Paula de Tezanos Pinto, Nathan G. Swenson, Christopher A. Klausmeier, and Elena Litchman. Eco-evolutionary differences in light utilization traits and distributions of freshwater phytoplankton. Limnology and Oceanography, 56(2):589–598, March 2011. CODEN LIOCAH. ISSN 0024-3590.

Sugie:2018:ICE

[SYW18] Koji Sugie, Takeshi Yoshimura, and Masahide Wakita. Impact of CO_{2} on the elemental composition of the particulate and dissolved organic matter of marine diatoms emerged after nitrate depletion. Limnology and Oceanography, 63(5):1924–1943, September 2018. CODEN LIOCAH. ISSN 0024-3590.

Schwerin:2010:HCD

[SZH^{+}10] Susanne Schwerin, Bettina Zeis, Wolfgang Horn, Heidemarie Horn, and Rüdiger J. Paul. Hemoglobin concentration in Daphnia (D. galeata-hyalina) from the epilimnion is related to the state of nutrition and the degree of protein homeostasis. Limnology and Oceanography, 55(2):639–652, March 2010. CODEN LIOCAH. ISSN 0024-3590.

Titze:2014:WTS

[TA14] Daniel J. Titze and Jay A. Austin. Winter thermal structure of Lake Superior. Limnology and Oceanography, 59(4):1336–1348, July 2014. CODEN LIOCAH. ISSN 0024-3590.

Tadonleke:2010:EWE

[Tad10] Rémy D. Tadonléke. Evidence of warming effects on phytoplankton productivity rates and their dependence on eutrophication status. Limnology and Oceanography, 55(3):973–982, May 2010. CODEN LIOCAH. ISSN 0024-3590.

Tartarotti:2018:SPP

[TAE^{+}18] B. Tartarotti, A. Alfreider, M. Egg, N. Saul, T. Schneider, R. Sommaruga, A. Tischler, and J. Vetter. Seasonal plasticity in photoprotection modulates UV-induced hsp gene expression in copepods from a clear lake. Limnology and Oceanography, 63 (4):1579–1592, July 2018. CODEN LIOCAH. ISSN 0024-3590.

Tilstone:2010:HCM

[TAV^{+}10] Gavin H. Tilstone, Ruth L. Airs, Victor Martinez Vicente, Claire Widdicombe, and Carole Llewellyn. High concentrations of mycosporine-like amino acids and colored dissolved organic matter in the sea surface microlayer off the Iberian Peninsula. Limnology and Oceanography, 55(5):1835–1850, September 2010. CODEN LIOCAH. ISSN 0024-3590.

Thorlacius:2018:ILS

[TB18] Magnus Thorlacius and Tomas Brodin. Investigating large-scale invasion patterns using-small scale invasion successions — phenotypic differentiation of the invasive round goby (Neogobius melanostomus) at invasion fronts. Limnology and Oceanography, 63(2):702–713, March 2018. CODEN LIOCAH. ISSN 0024-3590.

Timoner:2014:DBB

[TBAS14] Xisca Timoner, Carles M. Borrego, Vicenç Acuña, and Sergi Sabater. The dynamics of biofilm bacterial communities is driven by flow wax and wane in a temporary stream. Limnology and Oceanography, 59(6):2057–2067, November 2014. CODEN LIOCAH. ISSN 0024-3590.

Tsui:2013:PMS

[TBF^{+}13] Martin Tsz Ki Tsui, Joel D. Blum, Jacques C. Finlay, Steven J. Balogh, Sae Yun Kwon, and Yabing H. Nollet. Photodegradation of methylmercury in stream ecosystems. Limnology and Oceanography, 58(1):13–22, January 2013. CODEN LIOCAH. ISSN 0024-3590.

Talmy:2013:OMP

[TBHM^{+}13] David Talmy, Jerry Blackford, Nick J. Hardman-Mountford, Alex J. Dumbrell, and Richard J. Geider. An optimality model of photoadaptation in contrasting aquatic light regimes. Limnology and Oceanography, 58(5):1802–1818, September 2013. CO-DEN LIOCAH. ISSN 0024-3590.

Tomonaga:2015:ADN

[TBK15] Yama Tomonaga, Matthias S. Brennwald, and Rolf Kipfer. Attenuation of diffusive noble-gas transport in laminated sediments of the Stockholm Archipelago. Limnology and Oceanography, 60(2):497–511, March 2015. CODEN LIOCAH. ISSN 0024-3590.

Terseleer:2014:TBR

[TBLG14] N. Terseleer, J. Bruggeman, C. Lancelot, and N. Gypens. Traitbased representation of diatom functional diversity in a plankton functional type model of the eutrophied southern North Sea. Limnology and Oceanography, 59(6):1958–1972, November 2014. CODEN LIOCAH. ISSN 0024-3590.

Tiwari:2017:ISD

[TBSL17] Tejshree Tiwari, Ishi Buffam, Ryan A. Sponseller, and Hjalmar Laudon. Inferring scale-dependent processes influencing stream water biogeochemistry from headwater to sea. Limnology and Oceanography, 62(S1):S58–S70, November 2017. CODEN LI-OCAH. ISSN 0024-3590.

Trimborn:2013:SAP

[TBSR13] Scarlett Trimborn, Tina Brenneis, Elizabeth Sweet, and Björn Rost. Sensitivity of Antarctic phytoplankton species to ocean acidification: Growth, carbon acquisition, and species interaction. Limnology and Oceanography, 58(3):997–1007, May 2013. CODEN LIOCAH. ISSN 0024-3590.

Tamburello:2019:DER

[TCFP19] N. Tamburello, B. M. Connors, D. Fullerton, and C. C. Phillis. Durability of environment–recruitment relationships in aquatic ecosystems: insights from long-term monitoring in a highly modified estuary and implications for management. Limnology and Oceanography, 64(S1):S223–S239, January 2019. CODEN LI-OCAH. ISSN 0024-3590.

Tang:2017:SDB

[TCG^{+}17] Xiangming Tang, Jianying Chao, Yi Gong, Yongping Wang, Steven W. Wilhelm, and Guang Gao. Spatiotemporal dynamics of bacterial community composition in large shallow eutrophic Lake Taihu: High overlap between free-living and particle-attached assemblages. Limnology and Oceanography, 62 (4):1366–1382, July 2017. CODEN LIOCAH. ISSN 0024-3590.

Tournois:2017:LNM

[TDF^{+}17] Jennifer Tournois, Audrey M. Darnaude, Franck Ferraton, Catherine Aliaume, Lény Mercier, and David J. McKenzie. Lagoon nurseries make a major contribution to adult populations of a highly prized coastal fish. Limnology and Oceanography, 63 (3):1219–1233, May 2017. CODEN LIOCAH. ISSN 0024-3590.

Teneva:2013:HRC

[TDM^{+}13] L. Teneva, R. B. Dunbar, D. A. Mucciarone, J. F. Dunckley, and J. R. Koseff. High-resolution carbon budgets on a Palau backreef modulated by interactions between hydrodynamics and reef metabolism. Limnology and Oceanography, 58(5):1851–1870, September 2013. CODEN LIOCAH. ISSN 0024-3590.

Taylor:2010:SRM

[TDS^{+}10] David Taylor, Sebastien Delaux, Craig Stevens, Roger Nokes, and David Schiel. Settlement rates of macroalgal algal propagules: Cross-species comparisons in a turbulent environment. Limnology and Oceanography, 55(1):66–76, January 2010. CO-DEN LIOCAH. ISSN 0024-3590.

Teece:2011:HAA

[TEGL11] Mark A. Teece, Benjamin Estes, Eric Gelsleichter, and Diego Lirman. Heterotrophic and autotrophic assimilation of fatty acids by two scleractinian corals, Montastraea faveolata and Porites astreoides. Limnology and Oceanography, 56(4):1285– 1296, July 2011. CODEN LIOCAH. ISSN 0024-3590.

Tao:2018:TVC

[TEZ^{+}18] Shuqin Tao, Timothy I. Eglinton, Liang Zhang, Zhiwei Yi, Daniel B. Montluçon, Cameron McIntyre, Meng Yu, and Meixun Zhao. Temporal variability in composition and fluxes of Yellow River particulate organic matter. Limnology and Oceanography, 63(S1):S119–S141, March 2018. CODEN LIOCAH. ISSN 0024-3590.

Taylor:2011:STC

[TF11] John R. Taylor and Raffaele Ferrari. Shutdown of turbulent convection as a new criterion for the onset of spring phytoplankton blooms. Limnology and Oceanography, 56(6):2293–2307, November 2011. CODEN LIOCAH. ISSN 0024-3590.

Tansik:2017:ICS

[TFH17] Anna L. Tansik, William K. Fitt, and Brian M. Hopkinson. Inorganic carbon is scarce for symbionts in scleractinian corals. Limnology and Oceanography, 65(9):2045–2055, September 2017. CODEN LIOCAH. ISSN 0024-3590.

Treibergs:2014:NIR

[TFLS14] Lija A. Treibergs, Sarah E. Fawcett, Michael W. Lomas, and Daniel M. Sigman. Nitrogen isotopic response of prokaryotic and eukaryotic phytoplankton to nitrate availability in Sargasso Sea surface waters. Limnology and Oceanography, 59(3):972– 985, May 2014. CODEN LIOCAH. ISSN 0024-3590.

Treibergs:2017:ELI

[TG17] Lija A. Treibergs and Julie Granger. Enzyme level N and O isotope effects of assimilatory and dissimilatory nitrate reduction. Limnology and Oceanography, 62(1):272–288, January 2017. CODEN LIOCAH. ISSN 0024-3590.

Tang:2010:DOA

[TGC^{+}10] Xiangming Tang, Guang Gao, Jianying Chao, Xiaodong Wang, Guangwei Zhu, and Boqiang Qin. Dynamics of organicaggregate-associated bacterial communities and related environmental factors in Lake Taihu, a large eutrophic shallow lake in China. Limnology and Oceanography, 55(2):469–480, March 2010. CODEN LIOCAH. ISSN 0024-3590.

Tang:2011:CGB

[TGG^{+}11] Kam W. Tang, Ronnie N. Glud, Anni Glud, Søren Rysgaard, and Torkel Gissel Nielsen. Copepod guts as biogeochemical hotspots in the sea: Evidence from microelectrode profiling of Calanus spp. Limnology and Oceanography, 56(2):666–672, March 2011. CODEN LIOCAH. ISSN 0024-3590.

Tremblay:2010:ROS

[TGGZS^{+}10] Nelly Tremblay, Jaime Gómez-Gutiérrez, Tania Zenteno-Savín, Carlos J. Robinson, and Laura Sánchez-Velasco. Role of oxidative stress in seasonal and daily vertical migration of three krill species in the Gulf of California. Limnology and Oceanography, 55(6):2570–2584, November 2010. CODEN LIOCAH. ISSN 0024-3590.

Thrane:2017:PAS

[THA17] Jan-Erik Thrane, Dag O. Hessen, and Tom Andersen. Plasticity in algal stoichiometry: Experimental evidence of a temperatureinduced shift in optimal supply N : P ratio. Limnology and Oceanography, 62(4):1346–1354, July 2017. CODEN LIOCAH. ISSN 0024-3590.

Tong:2016:EVG

[THFG16] Shanying Tong, David A. Hutchins, Feixue Fu, and Kunshan Gao. Effects of varying growth irradiance and nitrogen sources on calcification and physiological performance of the coccolithophore Gephyrocapsa oceanica grown under nitrogen limitation. Limnology and Oceanography, 61(6):2234–2242, November 2016. CODEN LIOCAH. ISSN 0024-3590.

Thottathil:2013:BCF

[THH^{+}13] Shoji D. Thottathil, Kazuhide Hayakawa, Yoshikuni Hodoki, Chikage Yoshimizu, Yuki Kobayashi, and Shin ichi Nakano. Biogeochemical control on fluorescent dissolved organic matter dynamics in a large freshwater lake (Lake Biwa, Japan). Limnology and Oceanography, 58(6):2262–2278, November 2013. CODEN LIOCAH. ISSN 0024-3590.

Thorson:2019:MIO

[Tho19] James T. Thorson. Measuring the impact of oceanographic indices on species distribution shifts: The spatially varying effect of cold-pool extent in the eastern Bering Sea. Limnology and Oceanography, 64(6):2632–2645, November 2019. CODEN LI-OCAH. ISSN 0024-3590.

Takahashi:2015:SOD

[TIF^{+}15] Kazutaka Takahashi, Tadafumi Ichikawa, Chika Fukugama, Misaki Yamane, Shigeho Kakehi, Yuji Okazaki, Hiroshi Kubota, and Ken Furuya. In situ observations of a doliolid bloom in a warm water filament using a video plankton recorder: Bloom development, fate, and effect on biogeochemical cycles and planktonic food webs. Limnology and Oceanography, 60(5):1763–1780, September 2015. CODEN LIOCAH. ISSN 0024-3590.

Tanaka:2014:NAA

[TIN^{+}14] Yasuaki Tanaka, Akira Iguchi, Kozue Nishida, Mayuri Inoue, Takashi Nakamura, Atsushi Suzuki, and Kazuhiko Sakai. Nutrient availability affects the response of juvenile corals and the endosymbionts to ocean acidification. Limnology and Oceanography, 59(5):1468–1476, September 2014. CODEN LIOCAH. ISSN 0024-3590.

Takahashi:2013:SCP

[TIS^{+}13] Kazutaka Takahashi, Tadafumi Ichikawa, Hiroaki Saito, Shigeho Kakehi, Yasunori Sugimoto, Kiyotaka Hidaka, and Koji Hamasaki. Sapphirinid copepods as predators of doliolids: Their role in doliolid mortality and sinking flux. Limnology and Oceanography, 58(6):1972–1984, November 2013. CODEN LIOCAH. ISSN 0024-3590.

Taucher:2015:CEC

[TJJ^{+}15] J. Taucher, J. Jones, A. James, M. A. Brzezinski, C. A. Carlson, U. Riebesell, and U. Passow. Combined effects of CO_{2} and temperature on carbon uptake and partitioning by the marine diatoms Thalassiosira weissflogii and Dactyliosolen fragilissimus. Limnology and Oceanography, 60(3):901–919, May 2015. CO-DEN LIOCAH. ISSN 0024-3590.

Testa:2012:HIS

[TK12] Jeremy Mark Testa and W. Michael Kemp. Hypoxia-induced shifts in nitrogen and phosphorus cycling in Chesapeake Bay.

Limnology and Oceanography, 57(3):835–850, May 2012. CO-DEN LIOCAH. ISSN 0024-3590.

Testa:2018:SST

[TKB18] Jeremy M. Testa, W. Michael Kemp, and Walter R. Boynton. Season-specific trends and linkages of nitrogen and oxygen cycles in Chesapeake Bay. Limnology and Oceanography, 64(4):2045– 2064, September 2018. CODEN LIOCAH. ISSN 0024-3590.

Tyssebotn:2017:CBU

[TKK^{+}17] Inger Marie B. Tyssebotn, Joanna D. Kinsey, David J. Kieber, Ronald P. Kiene, Alison N. Rellinger, and Jessie Motard-Côté. Concentrations, biological uptake, and respiration of dissolved acrylate and dimethylsulfoxide in the northern Gulf of Mexico. Limnology and Oceanography, 63(3):1198–1218, May 2017. CO-DEN LIOCAH. ISSN 0024-3590.

Tanzil:2016:LDB

[TLB^{+}16] J. T. I. Tanzil, J. N. Lee, B. E. Brown, Rick Quax, J. A. Kaandorp, J. M. Lough, and P. A. Todd. Luminescence and density banding patterns in massive Porites corals around the Thai-Malay Peninsula, southeast Asia. Limnology and Oceanography, 61(6):2003–2026, November 2016. CODEN LIOCAH. ISSN 0024-3590.

Tank:2011:MTD

[TLG^{+}11] Suzanne E. Tank, Lance F. W. Lesack, Jolie A. L. Gareis, Christopher L. Osburn, and Ray H. Hesslein. Multiple tracers demonstrate distinct sources of dissolved organic matter to lakes of the Mackenzie Delta, western Canadian Arctic. Limnology and Oceanography, 56(4):1297–1309, July 2011. CODEN LIOCAH. ISSN 0024-3590.

Tonolla:2011:CSH

[TLH^{+}11] Diego Tonolla, Mark S. Lorang, Kurt Heutschi, Chris C. Gotschalk, and Klement Tockner. Characterization of spatial heterogeneity in underwater soundscapes at the river segment scale. Limnology and Oceanography, 56(6):2319–2333, November 2011. CODEN LIOCAH. ISSN 0024-3590.

Thurber:2013:MMM

[TLR^{+}13] Andrew R. Thurber, Lisa A. Levin, Ashley A. Rowden, Stefan Sommer, Peter Linke, and Kerstin Kröger. Microbes, macrofauna, and methane: a novel seep community fueled by aerobic methanotrophy. Limnology and Oceanography, 58(5):1640–1656, September 2013. CODEN LIOCAH. ISSN 0024-3590.

Tang:2014:PRM

[TMF^{+}14] Kam W. Tang, Daniel F. McGinnis, Katharina Frindte, Volker Brüchert, and Hans-Peter Grossart. Paradox reconsidered: Methane oversaturation in well-oxygenated lake waters. Limnology and Oceanography, 59(1):275–284, January 2014. CODEN LIOCAH. ISSN 0024-3590.

Trimmer:2010:PCF

[TMH^{+}10] Mark Trimmer, Susanna Maanoja, Alan G. Hildrew, James L. Pretty, and Jonathan Grey. Potential carbon fixation via methane oxidation in well-oxygenated river bed gravels. Limnology and Oceanography, 55(2):560–568, March 2010. CODEN LIOCAH. ISSN 0024-3590.

Tong:2018:CPW

[TMH^{+}18] Chuan Tong, James T. Morris, Jiafang Huang, Hui Xu, and Siang Wan. Changes in pore-water chemistry and methane emission following the invasion of Spartina alterniflora into an oliogohaline marsh. Limnology and Oceanography, 63(4):384–396, January 2018. CODEN LIOCAH. ISSN 0024-3590.

Taylor:2013:MEP

[TMK^{+}13] Joe D. Taylor, Boyd A. McKew, Alison Kuhl, Terry J. McGenity, and Graham J. C. Underwood. Microphytobenthic extracellular polymeric substances (EPS) in intertidal sediments fuel both generalist and specialist EPS-degrading bacteria. Limnology and Oceanography, 58(4):1463–1480, July 2013. CODEN LIOCAH. ISSN 0024-3590.

Tsunogai:2018:QND

[TMO^{+}18] Urumu Tsunogai, Takanori Miyauchi, Takuya Ohyama, Daisuke D. Komatsu, Masanori Ito, and Fumiko Nakagawa. Quantifying nitrate dynamics in a mesotrophic lake using triple oxygen isotopes as tracers. Limnology and Oceanography, 63(S1):S458– S476, March 2018. CODEN LIOCAH. ISSN 0024-3590.

Tachibana:2019:ILT

[TNI19] Aiko Tachibana, Hideaki Nomura, and Takashi Ishimaru. Impacts of long-term environmental variability on diapause phenology of coastal copepods in Tokyo Bay, Japan. Limnology and Oceanography, 64(S1):S273–S283, January 2019. CODEN LIOCAH. ISSN 0024-3590.

Twining:2014:DRM

[TNK^{+}14] Benjamin S. Twining, Scott D. Nodder, Andrew L. King, David A. Hutchins, Gary R. LeCleir, Jennifer M. DeBruyn, Elizabeth W. Maas, Stefan Vogt, Steven W. Wilhelm, and Philip W. Boyd. Differential remineralization of major and trace elements in sinking diatoms. Limnology and Oceanography, 59(3):689– 704, May 2014. CODEN LIOCAH. ISSN 0024-3590.

Twining:2010:VSC

[TNMV^{+}10] Benjamin S. Twining, Daliangelis Nuñez-Milland, Stefan Vogt, Rodney S. Johnson, and Peter N. Sedwick. Variations in Synechococcus cell quotas of phosphorus, sulfur, manganese, iron, nickel, and zinc within mesoscale eddies in the Sargasso Sea. Limnology and Oceanography, 55(2):492–506, March 2010. CO-DEN LIOCAH. ISSN 0024-3590.

Toffolon:2014:PST

[TPM^{+}14] Marco Toffolon, Sebastiano Piccolroaz, Bruno Majone, Anna-Maria Soja, Frank Peeters, Martin Schmid, and Alfred Wüest. Prediction of surface temperature in lakes with different morphology using air temperature. Limnology and Oceanography, 61 (1):2185–2202, November 2014. CODEN LIOCAH. ISSN 0024-3590.

Tuttle-Raycraft:2019:LHT

[TRA19] Shaylah Tuttle-Raycraft and Josef Daniel Ackerman. Living the high turbidity life: The effects of total suspended solids, flow, and gill morphology on mussel feeding. Limnology and Oceanography, 64(6):2526–2537, November 2019. CODEN LI-OCAH. ISSN 0024-3590.

Thibodeau:2014:LNM

[TRD^{+}14] Patricia S. Thibodeau, Collin S. Roesler, Susan L. Drapeau, S. G. Prabhu Matondkar, Joaquim I. Goes, and P. Jeremy Werdell. Locating Noctiluca miliaris in the Arabian Sea: an optical proxy approach. Limnology and Oceanography, 59(6): 2042–2056, November 2014. CODEN LIOCAH. ISSN 0024-

Thamdrup:2019:AMO

[TSB^{+}19] Bo Thamdrup, Herdís G. R. Steinsdóttir, Anthony D. Bertagnolli, Cory C. Padilla, Nastassia V. Patin, Emilio Garcia-Robledo, Laura A. Bristow, and Frank J. Stewart. Anaerobic methane oxidation is an important sink for methane in the ocean’s largest oxygen minimum zone. Limnology and Oceanography, 64(6):2569–2585, November 2019. CODEN LIOCAH. ISSN 0024-3590.

Till:2019:ILM

[TSC^{+}19] C. P. Till, J. R. Solomon, N. R. Cohen, R. H. Lampe, A. Marchetti, T. H. Coale, and K. W. Bruland. The iron limitation mosaic in the California Current System: Factors governing Fe availability in the shelf/near-shelf region. Limnology and Oceanography, 64(1):109–123, January 2019. CODEN LIOCAH. ISSN 0024-3590.

Townsend-Small:2016:QEM

[TSDF^{+}16] Amy Townsend-Small, Doug Disbennett, Julianne M. Fernandez, Rebecca W. Ransohoff, Ross Mackay, and Rick A. Bourbonniere. Quantifying emissions of methane derived from anaerobic organic matter respiration and natural gas extraction in Lake Erie. Limnology and Oceanography, 61(S1):S356–S366, November 2016. CODEN LIOCAH. ISSN 0024-3590.

Tiselius:2013:SCF

[TSK13] Peter Tiselius, Enric Saiz, and Thomas Kiørboe. Sensory capabilities and food capture of two small copepods, Paracalanus parvus and Pseudocalanus sp. Limnology and Oceanography, 58 (5):1657–1666, September 2013. CODEN LIOCAH. ISSN 0024-3590. See reply [PJ16].

Thibodeau:2019:ECP

[TSSH19] P. S. Thibodeau, D. K. Steinberg, S. E. Stammerjohn, and C. Hauri. Environmental controls on pteropod biogeography along the Western Antarctic Peninsula. Limnology and Oceanography, 64(S1):S240–S256, January 2019. CODEN LIOCAH. ISSN 0024-3590.

Tsuchiya:2019:SVR

[TST^{+}19] Kenji Tsuchiya, Tomoharu Sano, Noriko Tomioka, Ayato Kohzu, Kazuhiro Komatsu, Ryuichiro Shinohara, Noriko Takamura, Megumi Nakagawa, Youta Sugai, Victor S. Kuwahara, Tatsuki Toda, Hideki Fukuda, and Akio Imai. Seasonal variability and regulation of bacterial production in a shallow eutrophic lake. Limnology and Oceanography, 64(6):2441–2454, November 2019. CODEN LIOCAH. ISSN 0024-3590.

Tweel:2012:WLU

[TT12] Andrew W. Tweel and R. Eugene Turner. Watershed land use and river engineering drive wetland formation and loss in the Mississippi River birdfoot delta. Limnology and Oceanography, 57(1):18–28, January 2012. CODEN LIOCAH. ISSN 0024-3590.

Tarling:2014:IMK

[TT14] Geraint A. Tarling and Sally E. Thorpe. Instantaneous movement of krill swarms in the Antarctic Circumpolar Current. Limnology and Oceanography, 59(3):872–886, May 2014. CODEN LIOCAH. ISSN 0024-3590.

Thomson:2019:BSL

[TTTM^{+}19] A. C. G. Thomson, S. M. Trevathan-Tackett, D. T. Maher, P. J. Ralph, and P. I. Macreadie. Bioturbator-stimulated loss of seagrass sediment carbon stocks. Limnology and Oceanography, 64 (1):342–356, January 2019. CODEN LIOCAH. ISSN 0024-3590.

Toming:2013:CAA

[TTV^{+}13] Kaire Toming, Lea Tuvikene, Sirje Vilbaste, Helen Agasild, Malle Viik, Anu Kisand, Tõnu Feldmann, Tõnu Martma, Roger I. Jones, and Tiina Nõges. Contributions of autochthonous and allochthonous sources to dissolved organic matter in a large, shallow, eutrophic lake with a highly calcareous catchment. Limnology and Oceanography, 58(4):1259–1270, July 2013. CODEN LIOCAH. ISSN 0024-3590.

Tamborski:2019:TVL

[TvBR^{+}19] Joseph Tamborski, Pieter van Beek, Valentí Rodellas, Christophe Monnin, Erwin Bergsma, Thomas Stieglitz, Christina Heilbrun, J. Kirk Cochran, Céline Charbonnier, Pierre Anschutz, Simon Bejannin, and Aaron Beck. Temporal variability of lagoon–sea water exchange and seawater circulation through a Mediterranean barrier beach. Limnology and Oceanography, 66(4):2059– 2080, September 2019. CODEN LIOCAH. ISSN 0024-3590.

Tan:2010:IDC

[TW10a] Qiao-Guo Tan and Wen-Xiong Wang. Interspecies differences in calcium content and requirement in four freshwater cladocerans explained by biokinetic parameters. Limnology and Oceanography, 55(3):1426–1434, May 2010. CODEN LIOCAH. ISSN 0024-3590.

Thor:2010:FRC

[TW10b] Peter Thor and Ida Wendt. Functional response of carbon absorption efficiency in the pelagic calanoid copepod Acartia tonsa. Limnology and Oceanography, 55(4):1779–1789, July 2010. CODEN LIOCAH. ISSN 0024-3590.

Tan:2011:CPC

[TW11] Qiao-Guo Tan and Wen-Xiong Wang. Contrasting patterns of cadmium bioaccumulation in freshwater cladocerans. Limnology and Oceanography, 56(1):257–267, January 2011. CODEN LIOCAH. ISSN 0024-3590.

Twining:2013:HCN

[TWP13] Cornelia W. Twining, Derek C. West, and David M. Post. Historical changes in nutrient inputs from humans and anadromous fishes in New England’s coastal watersheds. Limnology and Oceanography, 58(4):1286–1300, July 2013. CODEN LIOCAH. ISSN 0024-3590.

True:2018:CAT

[TWWY18] A. C. True, D. R. Webster, M. J. Weissburg, and J. Yen. Copepod avoidance of thin chemical layers of harmful algal compounds. Limnology and Oceanography, 63(3):1041–1055, May 2018. CODEN LIOCAH. ISSN 0024-3590.

Tang:2019:HMU

[TYX^{+}19] Yali Tang, Xiaoqin Yang, Ruohua Xu, Xiufeng Zhang, Zhengwen Liu, Yongdong Zhang, and Henri J. Dumont. Heterotrophic microbes upgrade food value of a terrestrial carbon resource for Daphnia magna. Limnology and Oceanography, 64(2):474–482, March 2019. CODEN LIOCAH. ISSN 0024-3590.

Tzortziou:2015:CDO

[TZD^{+}15] Maria Tzortziou, Christina Zeri, Elias Dimitriou, Yan Ding, Rudolf Jaffé, Emmanouil Anagnostou, Elli Pitta, and Angeliki Mentzafou. Colored dissolved organic matter dynamics and anthropogenic influences in a major transboundary river and its coastal wetland. Limnology and Oceanography, 60(4):1222–1240, July 2015. CODEN LIOCAH. ISSN 0024-3590.

Uthicke:2010:WCN

[UA10] Sven Uthicke and Christine Altenrath. Water column nutrients control growth and C : N ratios of symbiont-bearing benthic Foraminifera on the Great Barrier Reef, Australia. Limnology and Oceanography, 55(4):1681–1696, July 2010. CODEN LI-OCAH. ISSN 0024-3590.

Ullrich:2016:PEM

[UCOG16] Nina Ullrich, Peter Casper, Andreas Otto, and Mark O. Gessner. Proteomic evidence of methanotrophy in methane-enriched hypolimnetic lake water. Limnology and Oceanography, 61(S1): S91–S100, November 2016. CODEN LIOCAH. ISSN 0024-3590.

Uchimiya:2018:BOC

[UFW^{+}18] Mario Uchimiya, Hideki Fukuda, Masahide Wakita, Minoru Kitamura, Hajime Kawakami, Makio C. Honda, Hiroshi Ogawa, and Toshi Nagata. Balancing organic carbon supply and consumption in the ocean’s interior: Evidence from repeated biogeochemical observations conducted in the subarctic and subtropical western North Pacific. Limnology and Oceanography, 64(4):2015–2027, September 2018. CODEN LIOCAH. ISSN 0024-3590.

usler:2011:KRH

[uGH^{+}11] Eva Rothä usler, Ivàn Gómez, Ivàn A. Hinojosa, Ulf Karsten, Leonardo Miranda, Fadia Tala, and Martin Thiel. Kelp rafts in the Humboldt Current: Interplay of abiotic and biotic factors limit their floating persistence and dispersal potential. Limnology and Oceanography, 56(5):1751–1763, September 2011. CO-DEN LIOCAH. ISSN 0024-3590.

Urabe:2011:WLW

[UIY^{+}11] Jotaro Urabe, Tomoya Iwata, Yosuke Yagami, Eriko Kato, Takao Suzuki, Shuji Hino, and Syuhei Ban. Within-lake and watershed determinants of carbon dioxide in surface water: a comparative analysis of a variety of lakes in the Japanese Islands. Limnology and Oceanography, 56(1):49–60, January 2011. CODEN LIOCAH. ISSN 0024-3590.

Urakawa:2014:AAS

[UMHH^{+}14] Hidetoshi Urakawa, Willm Martens-Habbena, Carme Huguet, Jose R. de la Torre, Anitra E. Ingalls, Allan H. Devol, and David A. Stahl. Ammonia availability shapes the seasonal distribution and activity of archaeal and bacterial ammonia oxidizers in the Puget Sound Estuary. Limnology and Oceanography, 59 (4):1321–1335, July 2014. CODEN LIOCAH. ISSN 0024-3590.

Uitz:2010:VOP

[USB^{+}10] Julia Uitz, Dariusz Stramski, Anne-Claire Baudoux, Rick A. Reynolds, Vanessa M. Wright, Jean Dubranna, and Farooq Azam. Variations in the optical properties of a particle suspension associated with viral infection of marine bacteria. Limnology and Oceanography, 55(6):2317–2330, November 2010. CO-DEN LIOCAH. ISSN 0024-3590.

Uusikivi:2010:CML

[UVGS10] Jari Uusikivi, Anssi V. Vähätalo, Mats A. Granskog, and Ruben Sommaruga. Contribution of mycosporine-like amino acids and colored dissolved and particulate matter to sea ice optical properties and ultraviolet attenuation. Limnology and Oceanography, 55(2):703–713, March 2010. CODEN LIOCAH. ISSN 0024-3590.

Villar-Argaiz:2012:DFQ

[VABMS^{+}12] Manuel Villar-Argaiz, Francisco José Bullejos, Juan Manuel Medina-Sánchez, Eloísa Ramos-Rodríguez, José Antonio Delgado-Molina, and Presentación Carrillo. Disentangling food quantity and quality effects in zooplankton response to P-enrichment and UV radiation. Limnology and Oceanography, 57(1):235– 250, January 2012. CODEN LIOCAH. ISSN 0024-3590.

Verburg:2011:DCD

[VAH11] Piet Verburg, Jason P. Antenucci, and Robert E. Hecky. Differential cooling drives large-scale convective circulation in Lake Tanganyika. Limnology and Oceanography, 56(3):910–926, May 2011. CODEN LIOCAH. ISSN 0024-3590.

Voss:2017:EMR

[VB17] Kristofor A. Voss and Emily S. Bernhardt. Effects of mountaintop removal coal mining on the diversity and secondary productivity of Appalachian rivers. Limnology and Oceanography, 62 (5):1754–1770, July 2017. CODEN LIOCAH. ISSN 0024-3590.

vanBergen:2019:SDV

[vBBM^{+}19] Tamara J. H. M. van Bergen, Nathan Barros, Raquel Mendonça, Ralf C. H. Aben, Inge H. J. Althuizen, Vera Huszar, Leon P. M.

Lamers, Miquel Lürling, Fábio Roland, and Sarian Kosten. Seasonal and diel variation in greenhouse gas emissions from an urban pond and its major drivers. Limnology and Oceanography, 66(4):2129–2139, September 2019. CODEN LIOCAH. ISSN 0024-3590.

Valipour:2015:NIW

[VBBR15] Reza Valipour, Damien Bouffard, Leon Boegman, and Yerubandi R. Rao. Near-inertial waves in Lake Erie. Limnology and Oceanography, 60(5):1522–1535, September 2015. CODEN LIOCAH. ISSN 0024-3590.

Valipour:2017:SRM

[VBBR17] Reza Valipour, Leon Boegman, Damien Bouffard, and Yerubandi R. Rao. Sediment resuspension mechanisms and their contributions to high-turbidity events in a large lake. Limnology and Oceanography, 63(3):1045–1065, May 2017. CODEN LIOCAH. ISSN 0024-3590.

VandeBogert:2012:SHS

[VBC^{+}12] Matthew C. Van de Bogert, Darren L. Bade, Stephen R. Carpenter, Jonathan J. Cole, Michael L. Pace, Paul C. Hanson, and Owen C. Langman. Spatial heterogeneity strongly affects estimates of ecosystem metabolism in two north temperate lakes. Limnology and Oceanography, 57(6):1689–1700, November 2012. CODEN LIOCAH. ISSN 0024-3590.

Villafane:2013:ISV

[VBGG^{+}13] Virginia E. Villafañe, Anastazia T. Banaszak, Sergio D. Guendulain-García, Sebastian M. Strauch, Silvana R. Halac, and E. Walter Helbling. Influence of seasonal variables associated with climate change on photochemical diurnal cycles of marine phytoplankton from Patagonia (Argentina). Limnology and Oceanography, 58(2):203–214, January 2013. CODEN LI-OCAH. ISSN 0024-3590.

Venables:2013:WCS

[VCM13] Hugh J. Venables, Andrew Clarke, and Michael P. Meredith. Wintertime controls on summer stratification and productivity at the western Antarctic Peninsula. Limnology and Oceanography, 58(3):1035–1047, May 2013. CODEN LIOCAH. ISSN 0024-3590.

Vila-Costa:2016:MLM

[VCPC^{+}16] Maria Vila-Costa, Cristina Pulido, Eglantine Chappuis, Adelina Calviño, Emilio O. Casamayor, and Esperança Gacia. Macrophyte landscape modulates lake ecosystem-level nitrogen losses through tightly coupled plant-microbe interactions. Limnology and Oceanography, 61(1):78–88, January 2016. CODEN LI-OCAH. ISSN 0024-3590.

vandenHeuvel:2019:IND

[vdHHC^{+}19] Michael R. van den Heuvel, Jesse K. Hitchcock, Michael R. S. Coffin, Christina C. Pater, and Simon C. Courtenay. Inorganic nitrogen has a dominant impact on estuarine eelgrass distribution in the Southern Gulf of St. Lawrence, Canada. Limnology and Oceanography, 64(6):2313–2327, November 2019. CODEN LIOCAH. ISSN 0024-3590.

vanderJagt:2018:BES

[vdJFS^{+}18] Helga van der Jagt, Carmen Friese, Jan-Berend W. Stuut, Gerhard Fischer, and Morten H. Iversen. The ballasting effect of Saharan dust deposition on aggregate dynamics and carbon export: Aggregation, settling, and scavenging potential of marine snow. Limnology and Oceanography, 63(3):1386–1394, May 2018. CODEN LIOCAH. ISSN 0024-3590.

Valiadi:2019:MBB

[VdRA^{+}19] Martha Valiadi, Tristan de Rond, Ana Amorim, John R. Gittins, Chrysoula Gubili, Bradley S. Moore, Maria Debora Iglesias-Rodriguez, and Michael I. Latz. Molecular and biochemical basis for the loss of bioluminescence in the dinoflagellate Noctiluca scintillans along the west coast of the U.S.A. Limnology and Oceanography, 64(6):2709–2724, November 2019. CODEN LIOCAH. ISSN 0024-3590.

Vandeperre:2016:EPH

[VdSLC^{+}16] Frederic Vandeperre, Alexandre Aires da Silva, Cleridy Lennert-Cody, Ricardo Serrão Santos, and Pedro Afonso. Essential pelagic habitat of juvenile blue shark (Prionace glauca) inferred from telemetry data. Limnology and Oceanography, 61(5):1605– 1625, September 2016. CODEN LIOCAH. ISSN 0024-3590.

vonEinem:2010:EFD

[vEG10] Jessica von Einem and Wilhelm Granéli. Effects of fetch and dissolved organic carbon on epilimnion depth and light climate in small forest lakes in southern Sweden. Limnology and Oceanography, 55(4):920–930, March 2010. CODEN LIOCAH. ISSN 0024-3590.

Vogel:2010:MAH

[VF10] Catherine Vogel and Nicholas S. Fisher. Metal accumulation by heterotrophic marine bacterioplankton. Limnology and Oceanography, 55(2):519–528, March 2010. CODEN LIOCAH. ISSN 0024-3590.

Valdes:2018:NEC

[VFME18] Valentina Valdés, Camila Fernandez, Veronica Molina, and Rubén Escribano. Nitrogen excretion by copepods and its effect on ammonia-oxidizing communities from a coastal upwelling zone. Limnology and Oceanography, 63(1):278–294, January 2018. CODEN LIOCAH. ISSN 0024-3590.

Vogel:2015:CGA

[VFS^{+}15] Nikolas Vogel, Katharina Elisabeth Fabricius, Julia Strahl, Sam Hamilton Croft Noonan, Christian Wild, and Sven Uthicke. Calcareous green alga Halimeda tolerates ocean acidification conditions at tropical carbon dioxide seeps. Limnology and Oceanography, 60(1):263–275, January 2015. CODEN LIOCAH. ISSN 0024-3590.

Visser:2017:CHL

[VGJ17] Andre W. Visser, Josephine Grønning, and Sigrún Huld Jónasdóttir. Calanus hyperboreus and the lipid pump. Limnology and Oceanography, 63(3):1155–1165, May 2017. CODEN LIOCAH. ISSN 0024-3590.

Vedamati:2014:ISE

[VGM14] Jagruti Vedamati, Tyler Goepfert, and James W. Moffett. Iron speciation in the eastern tropical South Pacific oxygen minimum zone off Peru. Limnology and Oceanography, 59(6):1945–1957, November 2014. CODEN LIOCAH. ISSN 0024-3590.

vanHaren:2019:TCH

[vH19] Hans van Haren. Turbulent convection and high-frequency internal wave details in 1-m shallow waters. Limnology and Oceanography, 64(3):1323–1332, May 2019. CODEN LIOCAH. ISSN 0024-3590.

Vagle:2010:MBC

[VHM^{+}10] S. Vagle, J. Hume, F. McLaughlin, E. MacIsaac, and K. Shortreed. A methane bubble curtain in meromictic Sakinaw Lake, British Columbia. Limnology and Oceanography, 55(3):1313– 1326, May 2010. CODEN LIOCAH. ISSN 0024-3590.

vanHees:2019:CAP

[vHOM^{+}19] Daniel H. van Hees, Ylva S. Olsen, Lydiane Mattio, Leonardo Ruiz-Montoya, Thomas Wernberg, and Gary A. Kendrick. Cast adrift: Physiology and dispersal of benthic Sargassum spinuligerum in surface rafts. Limnology and Oceanography, 64(2): 526–540, March 2019. CODEN LIOCAH. ISSN 0024-3590.

Verbruggen:2011:SOI

[VHR^{+}11] F. Verbruggen, O. Heiri, G. J. Reichart, C. Blaga, and A. F. Lotter. Stable oxygen isotopes in chironomid and cladoceran remains as indicators for lake-water δ ^{18}O. Limnology and Oceanography, 56(6):2071–2079, November 2011. CODEN LI-OCAH. ISSN 0024-3590.

Vilhena:2010:RCC

[VHrI10] Leticia C. Vilhena, Ingrid Hillmer, and Jö rg Imberger. The role of climate change in the occurrence of algal blooms: Lake Burragorang, Australia. Limnology and Oceanography, 55(3): 1188–1200, May 2010. CODEN LIOCAH. ISSN 0024-3590.

Vermilyea:2010:DPH

[VHV10] Andrew W. Vermilyea, S. Paul Hansard, and Bettina M. Voelker. Dark production of hydrogen peroxide in the Gulf of Alaska. Limnology and Oceanography, 55(2):580–588, March 2010. CODEN LIOCAH. ISSN 0024-3590.

Vogt:2013:FCL

[VIS^{+}13] Robert A. Vogt, Toni R. Ignoffo, Lindsay J. Sullivan, Julian Herndon, Jonathon H. Stillman, and Wim J. Kimmerer. Feeding capabilities and limitations in the nauplii of two pelagic estuarine copepods, Pseudodiaptomus marinus and Oithona davisae. Limnology and Oceanography, 58(6):2145–2157, November 2013. CODEN LIOCAH. ISSN 0024-3590.

Vehmaa:2018:LTC

[VKC18] Anu Vehmaa, Tarja Katajisto, and Ulrika Candolin. Longterm changes in a zooplankton community revealed by the sediment archive. Limnology and Oceanography, 64(4):2126–2139, September 2018. CODEN LIOCAH. ISSN 0024-3590.

Vachon:2019:IWC

[VLDM19] Dominic Vachon, Timon Langenegger, Daphne Donis, and Daniel F. McGinnis. Influence of water column stratification and mixing patterns on the fate of methane produced in deep sediments of a small eutrophic lake. Limnology and Oceanography, 66(4):2114–2128, September 2019. CODEN LIOCAH. ISSN 0024-3590.

Varin:2010:MPA

[VLJ^{+}10] Thibault Varin, Connie Lovejoy, Anne D. Jungblut, Warwick F. Vincent, and Jacques Corbeil. Metagenomic profiling of Arctic microbial mat communities as nutrient scavenging and recycling systems. Limnology and Oceanography, 55(5):1901–1911, September 2010. CODEN LIOCAH. ISSN 0024-3590.

Valle-Levinson:2011:TVS

[VLMTEW11] Arnoldo Valle-Levinson, Ismael Mariño-Tapia, Cecilia Enriquez, and Amy F. Waterhouse. Tidal variability of salinity and velocity fields related to intense point-source submarine groundwater discharges into the coastal ocean. Limnology and Oceanography, 56(4):1213–1224, July 2011. CODEN LIOCAH. ISSN 0024-3590.

VanCleave:2014:RSL

[VLWV14] Katherine Van Cleave, John D. Lenters, Jia Wang, and Edward M. Verhamme. A regime shift in Lake Superior ice cover, evaporation, and water temperature following the warm El Niñ winter of 1997–1998. Limnology and Oceanography, 59(6):1889– 1898, November 2014. CODEN LIOCAH. ISSN 0024-3590.

Vick-Majors:2016:BMD

[VMAS^{+}16] Trista J. Vick-Majors, Amanda Achberger, Pamela Santibáñez, John E. Dore, Timothy Hodson, Alexander B. Michaud, Brent C. Christner, Jill Mikucki, Mark L. Skidmore, Ross Powell, W. Peyton Adkins, Carlo Barbante, Andrew Mitchell, Reed Scherer, and John C. Priscu. Biogeochemistry and microbial diversity in the marine cavity beneath the McMurdo Ice Shelf, Antarctica. Limnology and Oceanography, 61(2):572–586, March 2016. CODEN LIOCAH. ISSN 0024-3590.

Vogt:2013:FWC

[VMC^{+}13] R. J. Vogt, B. Matthews, T. P. Cobb, M. D. Graham, and P. R. Leavitt. Food web consequences of size-based predation and vertical migration of an invertebrate predator (Leptodora kindtii). Limnology and Oceanography, 58(5):1790–1801, September 2013. CODEN LIOCAH. ISSN 0024-3590.

Villamana:2017:RIW

[VMCM^{+}17] Marina Villamaña, Beatriz Mouriño-Carballido, Emilio Marañón, Pedro Cermeño, Paloma Chouciño, José C. B. da Silva, Patricio A. Díaz, Bieito Fernández-Castro, Miguel Gilcoto, Rocío Graña, Mikel Latasa, Jorge M. Magalhaes, José Luis Otero-Ferrer, Beatriz Reguera, and Renate Scharek. Role of internal waves on mixing, nutrient supply and phytoplankton community structure during spring and neap tides in the upwelling ecosystem of Ría de Vigo (NW Iberian Peninsula). Limnology and Oceanography, 63(3):1014–1030, May 2017. CODEN LIOCAH. ISSN 0024-3590.

Vidussi:2011:EEW

[VMF^{+}11] Francesca Vidussi, Behzad Mostajir, Eric Fouilland, Emilie Le Floc’h, Jean Nouguier, Cècile Roques, Patrice Got, Delphine Thibault-Botha, Thierry Bouvier, and Marc Troussellier. Effects of experimental warming and increased ultraviolet B radiation on the Mediterranean plankton food web. Limnology and Oceanography, 56(1):206–218, January 2011. CODEN LIOCAH. ISSN 0024-3590.

Vilhena:2013:INI

[VMI13] Leticia C. Vilhena, Clelia L. Marti, and Jörg Imberger. The importance of nonlinear internal waves in a deep subalpine lake: Lake Iseo, Italy. Limnology and Oceanography, 58(5):1871–1891, September 2013. CODEN LIOCAH. ISSN 0024-3590.

Venkatachalam:2019:IOV

[VML^{+}19] Siddarthan Venkatachalam, Gwynneth F. Matcher, Tarron Lamont, Marcel van den Berg, Isabelle J. Ansorge, and Rosemary A. Dorrington. Influence of oceanographic variability on near-shore microbial communities of the sub-Antarctic Prince Edward Islands. Limnology and Oceanography, 64(1):258–271, January 2019. CODEN LIOCAH. ISSN 0024-3590.

Vidal:2013:TSV

[VMMS^{+}13] Javier Vidal, Sally MacIntyre, Erika E. McPhee-Shaw, William J. Shaw, and Stephen G. Monismith. Temporal and spatial variability of the internal wave field in a lake with complex morphometry. Limnology and Oceanography, 58(5):1557–1580, September 2013. CODEN LIOCAH. ISSN 0024-3590.

vanOevelen:2012:CFB

[vOSH12] Dick van Oevelen, Karline Soetaert, and Carlo Heip. Carbon flows in the benthic food web of the porcupine abyssal plain: The (un)importance of labile detritus in supporting microbial and faunal carbon demands. Limnology and Oceanography, 57 (3):645–664, March 2012. CODEN LIOCAH. ISSN 0024-3590.

Valois:2015:GDP

[VP15a] Amanda E. Valois and Robert Poulin. Global drivers of parasitism in freshwater plankton communities. Limnology and Oceanography, 60(5):1707–1718, September 2015. CODEN LI-OCAH. ISSN 0024-3590.

VanZuidam:2015:WFL

[VP15b] Bastiaan G. Van Zuidam and Edwin T. H. M. Peeters. Wave forces limit the establishment of submerged macrophytes in large shallow lakes. Limnology and Oceanography, 60(5):1536–1549, September 2015. CODEN LIOCAH. ISSN 0024-3590.

Vachon:2010:RBN

[VPC10] Dominic Vachon, Yves T. Prairie, and Jonathan J. Cole. The relationship between near-surface turbulence and gas transfer velocity in freshwater systems and its implications for floating chamber measurements of gas exchange. Limnology and Oceanography, 55(4):1723–1732, July 2010. CODEN LIOCAH. ISSN 0024-3590.

Voynova:2019:IRC

[VPG^{+}19] Yoana G. Voynova, Wilhelm Petersen, Martina Gehrung, Steffen Aßmann, and Andrew L. King. Intertidal regions changing coastal alkalinity: The Wadden Sea–North Sea tidally coupled bioreactor. Limnology and Oceanography, 64(3):1135–1149, May 2019. CODEN LIOCAH. ISSN 0024-3590.

Valerio:2012:SBS

[VPMrI12] Giulia Valerio, Marco Pilotti, Clelia Luisa Marti, and Jö rg Imberger. The structure of basin-scale internal waves in a stratified lake in response to lake bathymetry and wind spatial and temporal distribution: Lake Iseo, Italy. Limnology and Oceanography, 57(3):772–786, May 2012. CODEN LIOCAH. ISSN 0024-3590.

Volkenborn:2010:OPB

[VPWW10] N. Volkenborn, L. Polerecky, D. S. Wethey, and S. A. Woodin. Oscillatory porewater bioadvection in marine sediments induced by hydraulic activities of Arenicola marina. Limnology and Oceanography, 55(3):1231–1247, May 2010. CODEN LIOCAH. ISSN 0024-3590.

Vaquer-Sunyer:2010:SEA

[VSD10] Raquel Vaquer-Sunyer and Carlos M. Duarte. Sulfide exposure accelerates hypoxia-driven mortalit. Limnology and Oceanography, 55(3):1075–1082, May 2010. CODEN LIOCAH. ISSN 0024-3590.

Vachon:2017:RSD

[VSdG17] Dominic Vachon, Christopher T. Solomon, and Paul A. del Giorgio. Reconstructing the seasonal dynamics and relative contribution of the major processes sustaining CO_{2} emissions in northern lakes. Limnology and Oceanography, 62(2):706–722, March 2017. CODEN LIOCAH. ISSN 0024-3590.

Greve:2017:MBP

[vSGAK17] Hans van Someren Gréve, Rodrigo Almeda, and Thomas Kiørboe. Motile behavior and predation risk in planktonic copepods. Limnology and Oceanography, 62(5):1810–1824, September 2017. CODEN LIOCAH. ISSN 0024-3590.

VandenWyngaert:2011:QDS

[VSP^{+}11] Silke Van den Wyngaert, Michaela M. Salcher, Jakob Pernthaler, Michael Zeder, and Thomas Posch. Quantitative dominance of seasonally persistent filamentous cyanobacteria (Planktothrix rubescens) in the microbial assemblages of a temperate lake. Limnology and Oceanography, 56(1):97–109, January 2011. CODEN LIOCAH. ISSN 0024-3590.

VanDam:2018:CLC

[VTH^{+}18] Bryce R. Van Dam, Craig Tobias, Andreas Holbach, Hans W. Paerl, and Guangwei Zhu. CO_{2} limited conditions favor cyanobacteria in a hypereutrophic lake: an empirical and theoretical stable isotope study. Limnology and Oceanography, 63 (4):1643–1659, July 2018. CODEN LIOCAH. ISSN 0024-3590.

Veuger:2011:LTP

[VvO11] Bart Veuger and Dick van Oevelen. Long-term pigment dynamics and diatom survival in dark sediment. Limnology and Oceanography, 57(4):1065–1074, May 2011. CODEN LIOCAH. ISSN 0024-3590.

Vlahos:2017:OCP

[VW17] Penny Vlahos and Michael M. Whitney. Organic carbon patterns and budgets in the Long Island Sound estuary. Limnology and Oceanography, 62(S1):S46–S57, November 2017. CODEN LIOCAH. ISSN 0024-3590.

Vizza:2017:LPS

[VZJ^{+}17] Carmella Vizza, Jacob A. Zwart, Stuart E. Jones, Scott D. Tiegs, and Gary A. Lamberti. Landscape patterns shape wetland pond ecosystem function from glacial headwaters to ocean. Limnology and Oceanography, 62(S1):S207–S221, November 2017. CODEN LIOCAH. ISSN 0024-3590.

Walsh:2014:CEC

[WA14] Michael J. Walsh and Beth A. Ahner. Copper export contributes to low copper levels and copper tolerance in Emiliania huxleyi. Limnology and Oceanography, 59(3):827–839, May 2014. CO-DEN LIOCAH. ISSN 0024-3590.

Wheeler:2017:RPP

[WAB^{+}17] Sarah G. Wheeler, Todd W. Anderson, Tom W. Bell, Steven G. Morgan, and James A. Hobbs. Regional productivity predicts individual growth and recruitment of rockfishes in a northern California upwelling system. Limnology and Oceanography, 62 (2):754–767, March 2017. CODEN LIOCAH. ISSN 0024-3590.

Weidberg:2019:LTV

[WB19] Nicolas Weidberg and Sünnje L. Basedow. Long-term variability in overwintering copepod populations in the Lofoten Basin: The role of the North Atlantic oscillation and trophic effects. Limnology and Oceanography, 66(4):2044–2058, September 2019. CO-DEN LIOCAH. ISSN 0024-3590.

Winder:2017:IAZ

[WBB^{+}17] Monika Winder, Jean-Marie Bouquet, J. Rafael Bermúdez, Stella A. Berger, Thomas Hansen, Jay Brandes, Andrey F. Sazhin, Jens C. Nejstgaard, Ulf Båmstedt, Hans H. Jakobsen, Jörg Dutz, Marc E. Frischer, Christofer Troedsson, and Eric M. Thompson. Increased appendicularian zooplankton alter carbon cycling under warmer more acidified ocean conditions. Limnology and Oceanography, 62(4):1541–1551, July 2017. CODEN LIOCAH. ISSN 0024-3590.

Waite:2016:CST

[WBG^{+}16] Anya M. Waite, Lynnath E. Beckley, Lionel Guidi, Jason P. Landrum, David Holliday, Joseph Montoya, Harriet Paterson, Ming Feng, Peter A. Thompson, and Eric J. Raes. Cross-shelf transport, oxygen depletion, and nitrate release within a forming mesoscale eddy in the eastern Indian Ocean. Limnology and Oceanography, 61(1):103–121, January 2016. CODEN LIOCAH. ISSN 0024-3590.

Wessels:2010:DMF

[WBS^{+}10] Martin Wessels, Ingeborg Bussmann, Stefan Schloemer, Michael Schlü ter, and Volker Bö dere. Distribution, morphology, and formation of pockmarks in Lake Constance, Germany. Limnology and Oceanography, 55(6):2623–2633, November 2010. CODEN LIOCAH. ISSN 0024-3590.

Wenk:2013:AAO

[WBZ^{+}13] Christine B. Wenk, Jan Blees, Jakob Zopfi, Mauro Veronesi, Annie Bourbonnais, Carsten J. Schubert, Helge Niemann, and Moritz F. Lehmann. Anaerobic ammonium oxidation (anammox) bacteria and sulfide-dependent denitrifiers coexist in the water column of a meromictic south-alpine lake. Limnology and Oceanography, 58(1):1–12, January 2013. CODEN LIOCAH. ISSN 0024-3590.

Williamson:2014:LSL

[WBZ^{+}14] Craig E. Williamson, Jennifer A. Brentrup, Jing Zhang, William H. Renwick, Bruce R. Hargreaves, Lesley B. Knoll, Erin P. Overholt, and Kevin C. Rose. Lakes as sensors in the landscape: Optical metrics as scalable sentinel responses to climate change. Limnology and Oceanography, 59(3):840–850, May 2014. CODEN LIOCAH. ISSN 0024-3590.

Weyhenmeyer:2017:LDB

[WC17] Gesa A. Weyhenmeyer and Daniel J. Conley. Large differences between carbon and nutrient loss rates along the land to ocean aquatic continuum — implications for energy : nutrient ratios at downstream sites. Limnology and Oceanography, 62(S1):S183– S193, November 2017. CODEN LIOCAH. ISSN 0024-3590.

Wallace:2010:CZV

[WCB^{+}10] Margaret I. Wallace, Finlo R. Cottier, Jørgen Berge, Geraint A. Tarling, Colin Griffiths, and Andrew S. Brierley. Comparison of zooplankton vertical migration in an ice-free and a seasonally ice-covered Arctic fjord: an insight into the influence of sea ice cover on zooplankton behavior. Limnology and Oceanography, 55(4):831–845, March 2010. CODEN LIOCAH. ISSN 0024-3590.

Weber:2017:ARI

[WCC^{+}17] Sarah C. Weber, Edward J. Carpenter, Victoria J. Coles, Patricia L. Yager, Joaquim Goes, and Joseph P. Montoya. Amazon River influence on nitrogen fixation and export production in the western tropical North Atlantic. Limnology and Oceanography, 62(2):618–631, March 2017. CODEN LIOCAH. ISSN 0024-3590.

Wilkinson:2014:UDA

[WCCP14] Grace M. Wilkinson, Stephen R. Carpenter, Jonathan J. Cole, and Michael L. Pace. Use of deep autochthonous resources by zooplankton: Results of a metalimnetic addition of ^{13}C to a small lake. Limnology and Oceanography, 59(3):986–996, May 2014. CODEN LIOCAH. ISSN 0024-3590.

Winder:2017:LSI

[WCG^{+}17] Monika Winder, Jacob Carstensen, Aaron W. E. Galloway, Hans H. Jakobsen, and James E. Cloern. The land–sea interface: a source of high-quality phytoplankton to support secondary production. Limnology and Oceanography, 62(S1):S258–S271, November 2017. CODEN LIOCAH. ISSN 0024-3590.

Wu:2014:OAE

[WCI^{+}14] Yaping Wu, Douglas A. Campbell, Andrew J. Irwin, David J. Suggett, and Zoe V. Finkel. Ocean acidification enhances the growth rate of larger diatoms. Limnology and Oceanography, 59 (3):1027–1034, May 2014. CODEN LIOCAH. ISSN 0024-3590.

Wear:2015:SSD

[WCJ^{+}15] Emma K. Wear, Craig A. Carlson, Anna K. James, Mark A. Brzezinski, Laura A. Windecker, and Craig E. Nelson. Synchronous shifts in dissolved organic carbon bioavailability and bacterial community responses over the course of an upwellingdriven phytoplankton bloom. Limnology and Oceanography, 60 (2):657–677, March 2015. CODEN LIOCAH. ISSN 0024-3590.

West:2016:PDR

[WCJ16] William E. West, Kevin P. Creamer, and Stuart E. Jones. Productivity and depth regulate lake contributions to atmospheric methane. Limnology and Oceanography, 61(S1):S51–S61, November 2016. CODEN LIOCAH. ISSN 0024-3590.

Wang:2017:DBD

[WCJ^{+}17] Bin Wang, Jianfang Chen, Haiyan Jin, Hongliang Li, Daji Huang, and Wei-Jun Cai. Diatom bloom-derived bottom water hypoxia off the Changjiang Estuary, with and without typhoon influence. Limnology and Oceanography, 62(4):1552–1569, July 2017. CODEN LIOCAH. ISSN 0024-3590.

Walsh:2019:CLT

[WCM19] Jake R. Walsh, Jessica R. Corman, and Samuel E. Munoz. Coupled long-term limnological data and sedimentary records reveal new control on water quality in a eutrophic lake. Limnology and Oceanography, 64(S1):S34–S48, January 2019. CODEN LI-OCAH. ISSN 0024-3590.

Wilkinson:2015:PBC

[WCP^{+}15] Grace M. Wilkinson, Jonathan J. Cole, Michael L. Pace, Robert A. Johnson, and Maxwell J. Kleinhans. Physical and biological contributions to metalimnetic oxygen maxima in lakes. Limnology and Oceanography, 60(1):242–251, January 2015. CODEN LIOCAH. ISSN 0024-3590.

Wahl:2018:MMM

[WCS^{+}18] M. Wahl, S. Schneider Covachã, V. Saderne, C. Hiebenthal, J. D. Müller, C. Pansch, and Y. Sawall. Macroalgae may mitigate ocean acidification effects on mussel calcification by increasing pH and its fluctuations. Limnology and Oceanography, 63(1): 3–21, January 2018. CODEN LIOCAH. ISSN 0024-3590.

Wormer:2012:CHG

[WCV^{+}12] Lars Wörmer, Samuel Cirés, David Velázquez, Antonio Quesada, and Kai-Uwe Hinrichs. Cyanobacterial heterocyst glycolipids in cultures and environmental samples: Diversity and biomarker potential. Limnology and Oceanography, 57(6):1775– 1788, November 2012. CODEN LIOCAH. ISSN 0024-3590.

Wollrab:2015:BRL

[WD15] Sabine Wollrab and Sebastian Diehl. Bottom-up responses of the lower oceanic food web are sensitive to copepod mortality and feeding behavior. Limnology and Oceanography, 60(2):641– 656, March 2015. CODEN LIOCAH. ISSN 0024-3590.

Wernberg:2016:PRH

[WdBJF16] Thomas Wernberg, Thibaut de Bettignies, Bijo Arackal Joy, and Patrick Michael Finnegan. Physiological responses of habitatforming seaweeds to increasing temperatures. Limnology and Oceanography, 61(6):2180–2190, November 2016. CODEN LI-OCAH. ISSN 0024-3590.

Wang:2018:ICO

[WDCH18] Shiyu Rachel Wang, Daniela Di Iorio, Wei-Jun Cai, and Charles S. Hopkinson. Inorganic carbon and oxygen dynamics in a marsh-dominated estuary. Limnology and Oceanography, 63 (1):47–71, January 2018. CODEN LIOCAH. ISSN 0024-3590.

Wilson:2017:CCN

[WDH^{+}17] Shaun K. Wilson, Martial Depczynski, Thomas H. Holmes, Mae M. Noble, Ben T. Radford, Paul Tinkler, and Christopher J. Fulton. Climatic conditions and nursery habitat quality provide indicators of reef fish recruitment strength. Limnology and Oceanography, 62(5):1868–1880, September 2017. CODEN LIOCAH. ISSN 0024-3590.

Ward:2012:SSF

[WDJF12] B. A. Ward, S. Dutkiewicz, O. Jahn, and M. J. Follows. A size-structured food-web model for the global ocean. Limnology and Oceanography, 57(6):1877–1891, November 2012. CODEN LIOCAH. ISSN 0024-3590.

Wu:2017:DPD

[WDL^{+}17] Kai Wu, Minhan Dai, Xiaolin Li, Feifei Meng, Junhui Chen, and Jianrong Lin. Dynamics and production of dissolved organic carbon in a large continental shelf system under the influence of both river plume and coastal upwelling. Limnology and Oceanography, 63(3):973–988, May 2017. CODEN LIOCAH. ISSN 0024-3590.

Ward:2013:IPN

[WDMF13] Ben A. Ward, Stephanie Dutkiewicz, C. Mark Moore, and Michael J. Follows. Iron, phosphorus, and nitrogen supply ratios define the biogeography of nitrogen fixation. Limnology and Oceanography, 58(6):2059–2075, November 2013. CODEN LI-OCAH. ISSN 0024-3590.

Wang:2011:MCD

[WDX^{+}11] Da-Zhi Wang, Hong-Po Dong, Zhang-Xian Xie, Min-Han Dai, and Hua-Sheng Hong. Metaproteomic characterization of dissolved organic matter in the water column of the South China Sea. Limnology and Oceanography, 56(5):1641–1652, September 2011. CODEN LIOCAH. ISSN 0024-3590.

Wells:2019:PTS

[WE19] Naomi S. Wells and Bradley D. Eyre. δ ^{15}N patterns in three subtropical estuaries show switch from nitrogen “reactors” to “pipes” with increasing degradation. Limnology and Oceanography, 64(3):860–876, May 2019. CODEN LIOCAH. ISSN 0024-3590.

Williamson:2011:TMC

[WFB^{+}11] Craig E. Williamson, Janet M. Fischer, Stephen M. Bollens, Erin P. Overholt, and Joanne K. Breckenridge. Toward a more comprehensive theory of zooplankton diel vertical migration: Integrating ultraviolet radiation and water transparency into the biotic paradigm. Limnology and Oceanography, 56(5):1603– 1623, September 2011. CODEN LIOCAH. ISSN 0024-3590.

Wenk:2016:DDT

[WFK^{+}16] Christine B. Wenk, Caitlin H. Frame, Keisuke Koba, Karen L. Casciotti, Mauro Veronesi, Helge Niemann, Carsten J. Schubert, Naohiro Yoshida, Sakae Toyoda, Akiko Makabe, Jakob Zopfi, and Moritz F. Lehmann. Differential N_{2}O dynamics in two oxygen-deficient lake basins revealed by stable isotope and isotopomer distributions. Limnology and Oceanography, 62(3): 1735–1749, September 2016. CODEN LIOCAH. ISSN 0024-

Wyatt:2012:OFN

[WFL^{+}12] Alex S. J. Wyatt, James L. Falter, Ryan J. Lowe, Stuart Humphries, and Anya M. Waite. Oceanographic forcing of nutrient uptake and release over a fringing coral reef. Limnology and Oceanography, 57(2):401–419, March 2012. CODEN LIOCAH. ISSN 0024-3590.

Wright:2010:IRP

[WFR10] Derek D. Wright, Thomas K. Frazer, and John R. Reinfelder. The influence of river plume dynamics on trace metal accumulation in calanoid copepods. Limnology and Oceanography, 55 (6):2487–2502, November 2010. CODEN LIOCAH. ISSN 0024-3590.

Wong:2013:DGD

[WGC^{+}13] Wei Wen Wong, Michael R. Grace, Ian Cartwright, M. Bayani Cardenas, Peter B. Zamora, and Perran L. M. Cook. Dynamics of groundwater-derived nitrate and nitrous oxide in a tidal estuary from radon mass balance modeling. Limnology and Oceanography, 58(5):1689–1706, September 2013. CODEN LIOCAH. ISSN 0024-3590.

Wong:2014:SFN

[WGCC14] Wei Wen Wong, Michael R. Grace, Ian Cartwright, and Perran L. M. Cook. Sources and fate of nitrate in a groundwater-fed estuary elucidated using stable isotope ratios of nitrogen and oxygen. Limnology and Oceanography, 59(5):1493–1509, September 2014. CODEN LIOCAH. ISSN 0024-3590.

Weber:2019:CEU

[WGDAA19] Laura Weber, Patricia Gonzalez-Díaz, Maickel Armenteros, and Amy Apprill. The coral ecosphere: a unique coral reef habitat that fosters coral–microbial interactions. Limnology and Oceanography, 64(6):2373–2388, November 2019. CODEN LI-OCAH. ISSN 0024-3590.

Wheatcroft:2010:RED

[WGH^{+}10] Robert A. Wheatcroft, Miguel A. Goñi, Jeff A. Hatten, Gregory B. Pasternack, and Jonathan A. Warrick. The role of effective discharge in the ocean delivery of particulate organic carbon by small, mountainous river systems. Limnology and Oceanography, 55(1):161–171, January 2010. CODEN LIOCAH. ISSN 0024-3590.

Waldbusser:2016:SSB

[WGH^{+}16] George G. Waldbusser, Matthew W. Gray, Burke Hales, Chris J. Langdon, Brian A. Haley, Iria Gimenez, Stephanie R. Smith, Elizabeth L. Brunner, and Greg Hutchinson. Slow shell building, a possible trait for resistance to the effects of acute ocean acidification. Limnology and Oceanography, 61(6):1969–1983, November 2016. CODEN LIOCAH. ISSN 0024-3590.

Weidberg:2019:FSD

[WGJ^{+}19] Nicolas Weidberg, Wayne Goschen, Jennifer M. Jackson, Paula Pattrick, Christopher D. McQuaid, and Francesca Porri. Fine scale depth regulation of invertebrate larvae around coastal fronts. Limnology and Oceanography, 64(2):785–802, March 2019. CODEN LIOCAH. ISSN 0024-3590.

Werner:2016:EMN

[WGM16] Franziska Julie Werner, Angelika Graiff, and Birte Matthiessen. Even moderate nutrient enrichment negatively adds up to global climate change effects on a habitat-forming seaweed system. Limnology and Oceanography, 62(3):1891–1899, September 2016. CODEN LIOCAH. ISSN 0024-3590.

Wing:2017:SGR

[WGRS^{+}17] S. R. Wing, M. Gault-Ringold, C. H. Stirling, L. C. Wing, O. A. Shatova, and R. D. Frew. δ ^{56}Fe in seabird guano reveals extensive recycling of iron in the Southern Ocean ecosystem. Limnology and Oceanography, 62(4):1671–1681, July 2017. CODEN LIOCAH. ISSN 0024-3590.

Wheeler:2015:IHT

[WHAM15] Jeanette D. Wheeler, Karl R. Helfrich, Erik J. Anderson, and Lauren S. Mullineaux. Isolating the hydrodynamic triggers of the dive response in eastern oyster larvae. Limnology and Oceanography, 60(4):1332–1343, July 2015. CODEN LIOCAH. ISSN 0024-3590.

Wicks:2010:IIT

[WHD10] Laura C. Wicks, Ross Hill, and Simon K. Davy. The influence of irradiance on tolerance to high and low temperature stress exhibited by Symbiodinium in the coral, Pocillopora damicornis, from the high-latitude reef of Lord Howe Island. Limnology and Oceanography, 55(6):2476–2486, November 2010. CODEN LIOCAH. ISSN 0024-3590.

Wilken:2011:DFS

[WHH^{+}11] Susanne Wilken, Bernd Hoffmann, Nils Hersch, Norbert Kirchgessner, Sabine Dieluweit, Wolfgang Rubner, Linn J. Hoffmann, Rudolf Merkel, and Ilka Peeken. Diatom frustules show increased mechanical strength and altered valve morphology under iron limitation. Limnology and Oceanography, 56(4):1399– 1410, July 2011. CODEN LIOCAH. ISSN 0024-3590.

Wetz:2011:SDR

[WHL^{+}11] Michael S. Wetz, Emily A. Hutchinson, Ross S. Lunetta, Hans W. Paerl, and J. Christopher Taylor. Severe droughts reduce estuarine primary productivity with cascading effects on higher trophic levels. Limnology and Oceanography, 56(2):627– 638, March 2011. CODEN LIOCAH. ISSN 0024-3590.

Wolf:2018:RDL

[WHR18] Klara K. E. Wolf, Clara J. M. Hoppe, and Björn Rost. Resilience by diversity: Large intraspecific differences in climate change responses of an Arctic diatom. Limnology and Oceanography, 63(4):397–411, January 2018. CODEN LIOCAH. ISSN 0024-3590.

Williams:2018:FSV

[WJHS18] Joel Williams, Gregory P. Jenkins, Jeremy S. Hindell, and Steve E. Swearer. Fine-scale variability in elemental composition of estuarine water and otoliths: Developing environmental markers for determining larval fish dispersal histories within estuaries. Limnology and Oceanography, 63(1):262–277, January 2018. CODEN LIOCAH. ISSN 0024-3590.

Wlodarska-Kowalczuk:2019:TDM

[WKAM^{+}19] Maria Włodarska-Kowalczuk, Magnus Aune, Loïc N. Michel, Agata Zaborska, and Joanna Legeżyńska. Is the trophic diversity of marine benthic consumers decoupled from taxonomic and functional trait diversity? Isotopic niches of Arctic communities. Limnology and Oceanography, 66(4):2140–2151, September 2019. CODEN LIOCAH. ISSN 0024-3590.

White:2010:PMT

[WKB^{+}10] A. E. White, D. M. Karl, K. M. Björkman, L. J. Beversdorf, and R. M. Letelier. Phosphonate metabolism by Trichodesmium IMS101 and the production of greenhouse gases. Limnology and Oceanography, 55(4):1755–1767, July 2010. CODEN LIOCAH. ISSN 0024-3590.

Wang:2016:ISM

[WKG^{+}16] Zhaohui Aleck Wang, Kevin D. Kroeger, Neil K. Ganju, Meagan Eagle Gonneea, and Sophie N. Chu. Intertidal salt marshes as an important source of inorganic carbon to the coastal ocean. Limnology and Oceanography, 62(3):1916–1931, September 2016. CODEN LIOCAH. ISSN 0024-3590.

Wolff:2014:MSL

[WKJS^{+}14] Christian Wolff, Iris Kristen-Jenny, Georg Schettler, Birgit Plessen, Hanno Meyer, Peter Dulski, Rudolf Naumann, Achim Brauer, Dirk Verschuren, and Gerald H. Haug. Modern seasonality in Lake Challa (Kenya/Tanzania) and its sedimentary documentation in recent lake sediments. Limnology and Oceanography, 59(5):1621–1636, September 2014. CODEN LIOCAH. ISSN 0024-3590.

White:2011:LVV

[WKK^{+}11] Jeffrey D. White, RajReni B. Kaul, Lesley B. Knoll, Alan E. Wilson, and Orlando Sarnelle. Large variation in vulnerability to grazing within a population of the colonial phytoplankter, Microcystis aeruginosa. Limnology and Oceanography, 56(5): 1714–1724, September 2011. CODEN LIOCAH. ISSN 0024-3590.

Wain:2013:IWD

[WKSR13] Danielle J. Wain, Michael S. Kohn, Joshua A. Scanlon, and Chris R. Rehmann. Internal wave-driven transport of fluid away from the boundary of a lake. Limnology and Oceanography, 58 (2):429–442, March 2013. CODEN LIOCAH. ISSN 0024-3590.

Wang:2016:CSC

[WLG^{+}16] Xuchen Wang, Chunle Luo, Tiantian Ge, Caili Xu, and Yuejun Xue. Controls on the sources and cycling of dissolved inorganic carbon in the Changjiang and Huanghe River estuaries, China: 14C and ^{13}C studies. Limnology and Oceanography, 61(4):1358– 1374, July 2016. CODEN LIOCAH. ISSN 0024-3590.

Wyatt:2013:PNF

[WLHW13] Alex S. J. Wyatt, Ryan J. Lowe, Stuart Humphries, and Anya M. Waite. Particulate nutrient fluxes over a fringing coral reef: Source-sink dynamics inferred from carbon to nitrogen ratios and stable isotopes. Limnology and Oceanography, 58(3): 409–427, January 2013. CODEN LIOCAH. ISSN 0024-3590.

West:2011:MTC

[WLL^{+}11] A. J. West, C.-W. Lin, T.-C. Lin, R. G. Hilton, S.-H. Liu, C.- T. Chang, K.-C. Lin, A. Galy, R. B. Sparkes, and N. Hovius. Mobilization and transport of coarse woody debris to the oceans triggered by an extreme tropical storm. Limnology and Oceanography, 56(1):77–85, January 2011. CODEN LIOCAH. ISSN 0024-3590.

Wang:2019:PTI

[WLO^{+}19] Qing Wang, Zhao Lyu, Said Omar, Stephen Cornell, Zhou Yang, and David J. S. Montagnes. Predicting temperature impacts on aquatic productivity: Questioning the metabolic theory of ecology’s “canonical” activation energies. Limnology and Oceanography, 64(3):1172–1185, May 2019. CODEN LIOCAH. ISSN 0024-3590.

Wu:2017:DKB

[WLR17] Bin Wu, Su Mei Liu, and Jing Ling Ren. Dissolution kinetics of biogenic silica and tentative silicon balance in the Yellow Sea. Limnology and Oceanography, 62(4):1512–1525, July 2017. CODEN LIOCAH. ISSN 0024-3590.

Wai:2011:SSO

[WLS^{+}11] Tak-Cheung Wai, Kenneth M. Y. Leung, Shadow Y. T. Sin, Andy Cornish, David Dudgeon, and Gray A. Williams. Spatial, seasonal, and ontogenetic variations in the significance of detrital pathways and terrestrial carbon for a benthic shark, Chiloscyllium plagiosum (Hemiscylliidae), in a tropical estuary. Limnology and Oceanography, 57(4):1035–1053, May 2011. CO-DEN LIOCAH. ISSN 0024-3590.

Walsh:2017:IIP

[WLV17] Jake R. Walsh, Richard C. Lathrop, and M. Jake Vander Zanden. Invasive invertebrate predator, Bythotrephes longimanus, reverses trophic cascade in a north-temperate lake. Limnology and Oceanography, 62(6):2498–2509, November 2017. CODEN LIOCAH. ISSN 0024-3590.

Walsh:2018:UIW

[WLV18] Jake R. Walsh, Richard C. Lathrop, and M. Jake Vander Zanden. Uncoupling indicators of water quality due to the invasive zooplankter, Bythotrephes longimanus. Limnology and Oceanography, 63(3):1313–1327, May 2018. CODEN LIOCAH. ISSN 0024-3590.

Walker:2012:EIC

[WM12] B. D. Walker and M. D. McCarthy. Elemental and isotopic characterization of dissolved and particulate organic matter in a unique California upwelling system: Importance of size and composition in the export of labile material. Limnology and Oceanography, 57(6):1757–1774, November 2012. CODEN LI-OCAH. ISSN 0024-3590.

Widner:2017:CDU

[WM17] Brittany Widner and Margaret R. Mulholland. Cyanate distribution and uptake in North Atlantic coastal waters. Limnology and Oceanography, 62(6):2538–2549, November 2017. CODEN LIOCAH. ISSN 0024-3590.

Waeles:2013:CWA

[WMBR13] Matthieu Waeles, Jean-François Maguer, François Baurand, and Ricardo D. Riso. Off Congo waters (Angola Basin, Atlantic Ocean): a hot spot for cadmium-phosphate fractionation. Limnology and Oceanography, 58(4):1481–1490, July 2013. CODEN LIOCAH. ISSN 0024-3590.

Watras:2015:DCF

[WMC^{+}15] C. J. Watras, K. A. Morrison, J. T. Crawford, C. P. McDonald, S. K. Oliver, and P. C. Hanson. Diel cycles in the fluorescence of dissolved organic matter in dystrophic Wisconsin seepage lakes: Implications for carbon turnover. Limnology and Oceanography, 60(2):482–496, March 2015. CODEN LIOCAH. ISSN 0024-3590.

Weinstock:2018:TSV

[WMC^{+}18] Jane B. Weinstock, Scott L. Morello, LeAnn M. Conlon, Huijie Xue, and Philip O. Yund. Tidal shifts in the vertical distribution of bivalve larvae: Vertical advection vs. active behavior. Limnology and Oceanography, 63(6):2681–2694, November 2018. CODEN LIOCAH. ISSN 0024-3590.

Woodward:2017:WBD

[WMI^{+}17] B. L. Woodward, C. L. Marti, J. Imberger, M. R. Hipsey, and C. E. Oldham. Wind and buoyancy driven horizontal exchange in shallow embayments of a tropical reservoir: Lake Argyle, Western Australia. Limnology and Oceanography, 62(4):1636– 1657, July 2017. CODEN LIOCAH. ISSN 0024-3590.

Widner:2018:CDU

[WMM18] Brittany Widner, Calvin W. Mordy, and Margaret R. Mulholland. Cyanate distribution and uptake above and within the Eastern tropical South Pacific oxygen deficient zone. Limnology and Oceanography, 63(S1):S177–S192, March 2018. CODEN LI-OCAH. ISSN 0024-3590.

Wooster:2019:FRG

[WMP^{+}19] Michael K. Wooster, Steven E. McMurray, Joseph R. Pawlik, Xosé A. G. Morán, and Michael L. Berumen. Feeding and respiration by giant barrel sponges across a gradient of food abundance in the Red Sea. Limnology and Oceanography, 64(4):1790– 1801, July 2019. CODEN LIOCAH. ISSN 0024-3590.

Woodson:2012:CFS

[WMT^{+}12] C. B. Woodson, M. A. McManus, J. A. Tyburczy, J. A. Barth, L. Washburn, J. E. Caselle, M. H. Carr, D. P. Malone, P. T. Raimondi, B. A. Menge, and S. R. Palumbi. Coastal fronts set recruitment and connectivity patterns across multiple taxa. Limnology and Oceanography, 57(3):582–596, March 2012. CO-DEN LIOCAH. ISSN 0024-3590.

Williams:2018:ECN

[WOC^{+}18] Richard A. J. Williams, Hannah L. Owens, John Clamp, A. Townsend Peterson, Alan Warren, and Mercedes Martín-Cereceda. Endemicity and climatic niche differentiation in three marine ciliated protists. Limnology and Oceanography, 63(6): 2727–2736, November 2018. CODEN LIOCAH. ISSN 0024-3590.

Wahl:2014:ECC

[WP14] Bernd Wahl and Frank Peeters. Effect of climatic changes on stratification and deep-water renewal in Lake Constance assessed by sensitivity studies with a 3D hydrodynamic model. Limnology and Oceanography, 59(3):1035–1052, May 2014. CO-DEN LIOCAH. ISSN 0024-3590.

Wangpraseurt:2014:SLM

[WPL^{+}14] Daniel Wangpraseurt, Lubos Polerecky, Anthony W. D. Larkum, Peter J. Ralph, Daniel A. Nielsen, Mathieu Pernice, and Michael Kühl. The in situ light microenvironment of corals. Limnology and Oceanography, 59(3):917–926, May 2014. CO-DEN LIOCAH. ISSN 0024-3590.

Waite:2019:PES

[WRB^{+}19] Anya M. Waite, Eric Raes, Lynnath E. Beckley, Peter A. Thompson, David Griffin, Megan Saunders, Christin Säwström, Richard O’Rorke, Miao Wang, Jason P. Landrum, and Andrew Jeffs. Production and ecosystem structure in cold-core vs. warmcore eddies: Implications for the zooplankton isoscape and rock lobster larvae. Limnology and Oceanography, 64(6):2405–2423, November 2019. CODEN LIOCAH. ISSN 0024-3590.

Winslow:2017:SCS

[WRH^{+}17] Luke A. Winslow, Jordan S. Read, Gretchen J. A. Hansen, Kevin C. Rose, and Dale M. Robertson. Seasonality of change: Summer warming rates do not fully represent effects of climate change on lake temperatures. Limnology and Oceanography, 65 (9):2168–2178, September 2017. CODEN LIOCAH. ISSN 0024-3590.

Wordenweber:2018:PNS

[WRH^{+}18] Robin Wördenweber, Sebastian D. Rokitta, Elena Heidenreich, Katrin Corona, Frank Kirschhöfer, Kirsten Fahl, Jessica L. Klocke, Tilman Kottke, Gerald Brenner-Weiß, Björn Rost, Jan H. Mussgnug, and Olaf Kruse. Phosphorus and nitrogen starvation reveal life-cycle specific responses in the metabolome of Emiliania huxleyi (Haptophyta). Limnology and Oceanography, 63(1):203–226, January 2018. CODEN LIOCAH. ISSN 0024-3590.

Weissbach:2011:PAI

[WRO^{+}11] Astrid Weissbach, Maria Rudström, Martin Olofsson, Christian Béchemin, John Icely, Alice Newton, Urban Tillmann, and Catherine Legrand. Phytoplankton allelochemical interactions change microbial food web dynamics. Limnology and Oceanography, 56(3):899–909, May 2011. CODEN LIOCAH. ISSN 0024-

Wilson:2013:ZFP

[WRS13] S. E. Wilson, H. A. Ruhl, and K. L. Smith, Jr. Zooplankton fecal pellet flux in the abyssal northeast Pacific: a 15 year timeseries study. Limnology and Oceanography, 58(3):881–892, May 2013. CODEN LIOCAH. ISSN 0024-3590.

Wall:2019:SVB

[WRWPG19] Christopher B. Wall, Raphael Ritson-Williams, Brian N. Popp, and Ruth D. Gates. Spatial variation in the biochemical and isotopic composition of corals during bleaching and recovery. Limnology and Oceanography, 66(8):2011–2028, September 2019. CODEN LIOCAH. ISSN 0024-3590. See corrigendum [Ano21c].

Woodland:2013:BPC

[WS13] Ryan J. Woodland and David H. Secor. Benthic-pelagic coupling in a temperate inner continental shelf fish assemblage. Limnology and Oceanography, 58(3):966–976, May 2013. CODEN LIOCAH. ISSN 0024-3590.

Wagner:2018:CNP

[WS18] Tyler Wagner and Erin M. Schliep. Combining nutrient, productivity, and landscape-based regressions improves predictions of lake nutrients and provides insight into nutrient coupling at macroscales. Limnology and Oceanography, 63(6):2372–2383, November 2018. CODEN LIOCAH. ISSN 0024-3590.

Wurgaft:2013:MPD

[WSB^{+}13] Eyal Wurgaft, Ofer Shamir, Eugeni Barkan, Nathan Paldor, and Boaz Luz. Mixing processes in the deep water of the Gulf of Elat (Aqaba): Evidence from measurements and modeling of the triple isotopic composition of dissolved oxygen. Limnology and Oceanography, 58(4):1373–1386, July 2013. CODEN LIOCAH. ISSN 0024-3590.

Webb:2019:GSD

[WSM^{+}19] Jackie R. Webb, Isaac R. Santos, Damien T. Maher, Douglas R. Tait, Tyler Cyronak, Mahmood Sadat-Noori, Paul Macklin, and Luke C. Jeffrey. Groundwater as a source of dissolved organic matter to coastal waters: Insights from radon and CDOM observations in 12 shallow coastal systems. Limnology and Oceanography, 64(1):182–196, January 2019. CODEN LIOCAH. ISSN 0024-3590.

Wynne:2010:CCB

[WSTD10] Timothy T. Wynne, Richard P. Stumpf, Michelle C. Tomlinson, and Julianne Dyble. Characterizing a cyanobacterial bloom in Western Lake Erie using satellite imagery and meteorological data. Limnology and Oceanography, 55(5):2025–2036, September 2010. CODEN LIOCAH. ISSN 0024-3590.

Wunsch:2018:USD

[WSTG18] Urban J. Wünsch, Colin A. Stedmon, Lars J. Tranvik, and François Guillemette. Unraveling the size-dependent optical properties of dissolved organic matter. Limnology and Oceanography, 63(2):588–601, March 2018. CODEN LIOCAH. ISSN 0024-3590.

Wilken:2018:PPC

[WSUC^{+}18] Susanne Wilken, Margarida Soares, Pablo Urrutia-Cordero, Jens Ratcovich, Mattias K. Ekvall, Ellen Van Donk, and Lars-Anders Hansson. Primary producers or consumers? Increasing phytoplankton bacterivory along a gradient of lake warming and browning. Limnology and Oceanography, 63(S1):S142–S155, March 2018. CODEN LIOCAH. ISSN 0024-3590.

Williams:2017:CCS

[WTC^{+}17] Branwen Williams, Benoit Thibodeau, Yoshito Chikaraishi, Naohiko Ohkouchi, Andrew Walnum, Andréa G. Grottoli, and Patrick L. Colin. Consistency in coral skeletal amino acid composition offshore of Palau in the western Pacific warm pool indicates no impact of decadal variability in nitricline depth on primary productivity. Limnology and Oceanography, 62(2):399– 407, March 2017. CODEN LIOCAH. ISSN 0024-3590.

Woodland:2015:NLE

[WTN^{+}15] Ryan J. Woodland, James R. Thomson, Ralph Mac Nally, Paul Reich, Victor Evrard, Fiona Y. Wary, Jeffrey P. Walker, and Perran L. M. Cook. Nitrogen loads explain primary productivity in estuaries at the ecosystem scale. Limnology and Oceanography, 60(5):1751–1762, September 2015. CODEN LIOCAH. ISSN 0024-3590.

Waterkeyn:2010:ESS

[WVGB10] Aline Waterkeyn, Bram Vanschoenwinkel, Patrick Grillas, and Luc Brendoncka. Effect of salinity on seasonal community patterns of Mediterranean temporary wetland crustaceans: a mesocosm study. Limnology and Oceanography, 55(4):1712–1722, July 2010. CODEN LIOCAH. ISSN 0024-3590.

Woolway:2018:GTV

[WVL^{+}18] R. Iestyn Woolway, Piet Verburg, John D. Lenters, Christopher J. Merchant, David P. Hamilton, Justin Brookes, Elvira de Eyto, Sean Kelly, Nathan C. Healey, Simon Hook, Alo Laas, Don Pierson, James A. Rusak, Jonna Kuha, Juha Karjalainen, Kari Kallio, Ahti Lepistö, and Ian D. Jones. Geographic and temporal variations in turbulent heat loss from lakes: a global analysis across 45 lakes. Limnology and Oceanography, 63(6): 2436–2449, November 2018. CODEN LIOCAH. ISSN 0024-3590.

Waterkeyn:2011:LTE

[WVV^{+}11] Aline Waterkeyn, Bram Vanschoenwinkel, Hanne Vercampt, Patrick Grillas, and Luc Brendonck. Long-term effects of salinity and disturbance regime on active and dormant crustacean communities. Limnology and Oceanography, 57(4):1008–1022, May 2011. CODEN LIOCAH. ISSN 0024-3590.

Wang:2013:MIC

[WWC^{+}13] Zhaohui Aleck Wang, Rik Wanninkhof, Wei-Jun Cai, Robert H. Byrne, Xinping Hu, Tsung-Hung Peng, and Wei-Jen Huang. The marine inorganic carbon system along the Gulf of Mexico and Atlantic coasts of the United States: Insights from a transregional coastal carbon study. Limnology and Oceanography, 58 (2):325–342, January 2013. CODEN LIOCAH. ISSN 0024-3590.

Wang:2018:TSP

[WWC^{+}18] Mei Wang, Jianghua Wu, Huai Chen, Zicheng Yu, Qiu’an Zhu, Changhui Peng, Nicholas John Anderson, and Junwei Luan. Temporal-spatial pattern of organic carbon sequestration by Chinese lakes since 1850. Limnology and Oceanography, 63(3): 1283–1297, May 2018. CODEN LIOCAH. ISSN 0024-3590.

Whiting:2011:PMP

[WWS11] Daniel P. Whiting, Matt R. Whiles, and Mandy L. Stone. Patterns of macroinvertebrate production, trophic structure, and energy flow along a tallgrass prairie stream continuum. Limnology and Oceanography, 56(3):887–898, May 2011. CODEN LIOCAH. ISSN 0024-3590.

Wang:2015:WMW

[WXF^{+}15] Zongling Wang, Jie Xiao, Shiliang Fan, Yan Li, Xiangqing Liu, and Dongyan Liu. Who made the world’s largest green tide in China? — An integrated study on the initiation and early development of the green tide in Yellow Sea. Limnology and Oceanography, 60(4):1105–1117, July 2015. CODEN LIOCAH. ISSN 0024-3590.

White:2010:RNW

[WXMS10] Michael S. White, Marguerite A. Xenopoulos, Robert A. Metcalfe, and Keith M. Somers. On the role of natural water level fluctuation in structuring littoral benthic macroinvertebrate community composition in lakes. Limnology and Oceanography, 55(6):2275–2284, November 2010. CODEN LIOCAH. ISSN 0024-3590.

Wang:2016:CIF

[WYL16] Shilu Wang, Kevin M. Yeager, and Weiqi Lu. Carbon isotope fractionation in phytoplankton as a potential proxy for pH rather than for [CO_{2} (aq)]: Observations from a carbonate lake. Limnology and Oceanography, 61(4):1259–1270, July 2016. CO-DEN LIOCAH. ISSN 0024-3590.

Williams:2010:URL

[WYW^{+}10] Clayton J. Williams, Youhei Yamashita, Henry F. Wilson, Rudolf Jaffé, and Marguerite A. Xenopoulos. Unraveling the role of land use and microbial activity in shaping dissolved organic matter characteristics in stream ecosystems. Limnology and Oceanography, 55(3):1159–1171, May 2010. CODEN LI-OCAH. ISSN 0024-3590.

Wohlers-Zollner:2011:TNS

[WZBW^{+}11] Julia Wohlers-Zöllner, Petra Breithaupt, Katja Walther, Klaus Jürgens, and Ulf Riebesell. Temperature and nutrient stoichiometry interactively modulate organic matter cycling in a pelagic algal–bacterial community. Limnology and Oceanography, 56(2): 599–610, March 2011. CODEN LIOCAH. ISSN 0024-3590.

Wurl:2013:APB

[WZC13] Oliver Wurl, Louise Zimmer, and Gregory A. Cutter. Arsenic and phosphorus biogeochemistry in the ocean: Arsenic species as proxies for P-limitation. Limnology and Oceanography, 58(2): 729–740, March 2013. CODEN LIOCAH. ISSN 0024-3590.

Wenk:2014:PBB

[WZG^{+}14] Christine B. Wenk, Jakob Zopfi, Wayne S. Gardner, Mark J. McCarthy, Helge Niemann, Mauro Veronesi, and Moritz F. Lehmann. Partitioning between benthic and pelagic nitrate reduction in the Lake Lugano south basin. Limnology and Oceanography, 59(4):1421–1433, July 2014. CODEN LIOCAH. ISSN 0024-3590.

Wilson:2019:SAC

[WZR19] Will Wilson, Qiong Zhang, and Rosalind E. M. Rickaby. Susceptibility of algae to Cr toxicity reveals contrasting metal management strategies. Limnology and Oceanography, 66(4):2271–2282, September 2019. CODEN LIOCAH. ISSN 0024-3590.

Weitzman:2015:ACW

[WZTK15] Joel S. Weitzman, Robert B. Zeller, Florence I. M. Thomas, and Jeffrey R. Koseff. The attenuation of current- and wavedriven flow within submerged multispecific vegetative canopies. Limnology and Oceanography, 60(6):1855–1874, November 2015. CODEN LIOCAH. ISSN 0024-3590.

Xiao:2018:MMF

[XBR^{+}18] Ke-Qing Xiao, Felix Beulig, Hans Røy, Bo Barker Jørgensen, and Nils Risgaard-Petersen. Methylotrophic methanogenesis fuels cryptic methane cycling in marine surface sediment. Limnology and Oceanography, 63(4):1519–1527, July 2018. CODEN LIOCAH. ISSN 0024-3590.

Xiu:2019:CWI

[XDC^{+}19] Peng Xiu, Minhan Dai, Fei Chai, Kuanbo Zhou, Lili Zeng, and Chuanjun Du. On contributions by wind-induced mixing and eddy pumping to interannual chlorophyll variability during different ENSO phases in the northern South China Sea. Limnology and Oceanography, 64(2):503–514, March 2019. CODEN LIOCAH. ISSN 0024-3590.

Xenopoulos:2017:HOE

[XDK^{+}17] Marguerite A. Xenopoulos, John A. Downing, M. Dileep Kumar, Susanne Menden-Deuer, and Maren Voss. Headwaters to oceans: Ecological and biogeochemical contrasts across the aquatic continuum. Limnology and Oceanography, 62(S1):S3–S14, November 2017. CODEN LIOCAH. ISSN 0024-3590.

Xenopoulos:2019:ELT

[Xen19] Marguerite A. Xenopoulos. Editorial: Long-term studies in limnology and oceanography. Limnology and Oceanography, 64(S1): S1, January 2019. CODEN LIOCAH. ISSN 0024-3590.

Xu:2014:CRT

[XFH14] Kai Xu, Fei-Xue Fu, and David A. Hutchins. Comparative responses of two dominant Antarctic phytoplankton taxa to interactions between ocean acidification, warming, irradiance, and iron availability. Limnology and Oceanography, 59(6):1919–1931, November 2014. CODEN LIOCAH. ISSN 0024-3590.

Xing:2014:TSV

[XFLM14] Zikun Xing, Derek A. Fong, Edmond Yat-Man Lo, and Stephen G. Monismith. Thermal structure and variability of a shallow tropical reservoir. Limnology and Oceanography, 59(1): 115–128, January 2014. CODEN LIOCAH. ISSN 0024-3590.

Xu:2019:CES

[XLS^{+}19] Min Nina Xu, Xiaolin Li, Dalin Shi, Yao Zhang, Minhan Dai, Tao Huang, Patricia M. Glibert, and Shuh-Ji Kao. Coupled effect of substrate and light on assimilation and oxidation of regenerated nitrogen in the euphotic ocean. Limnology and Oceanography, 64(3):1270–1283, May 2019. CODEN LIOCAH. ISSN 0024-3590.

Xu:2018:FRA

[XNK18] Jiayi Xu, Lasse Tor Nielsen, and Thomas Kiørboe. Foraging response and acclimation of ambush feeding and feedingcurrent feeding copepods to toxic dinoflagellates. Limnology and Oceanography, 63(4):1449–1461, July 2018. CODEN LIOCAH. ISSN 0024-3590.

Xu:2010:NPI

[XPQ^{+}10] Hai Xu, Hans W. Paerl, Boqiang Qin, Guangwei Zhu, and Guang Gaoa. Nitrogen and phosphorus inputs control phytoplankton growth in eutrophic Lake Taihu, China. Limnology and Oceanography, 55(1):420–432, January 2010. CODEN LI-OCAH. ISSN 0024-3590.

Xiao:2013:CFI

[XSAHV13] Yi-Hua Xiao, Timo Sara-Aho, Helinä Hartikainen, and Anssi V. Vähätalo. Contribution of ferric iron to light absorption by chromophoric dissolved organic matter. Limnology and Oceanography, 58(2):653–662, March 2013. CODEN LIOCAH. ISSN 0024-3590.

Xu:2012:EPU

[XSAM12] Yan Xu, Dalin Shi, Ludmilla Aristilde, and François M. M. Morel. The effect of pH on the uptake of zinc and cadmium in marine phytoplankton: Possible role of weak complexes. Limnology and Oceanography, 57(1):293–304, January 2012. CODEN LIOCAH. ISSN 0024-3590.

Xiao:2019:CNO

[XXZ^{+}19] Qitao Xiao, Xiaofeng Xu, Mi Zhang, Hongtao Duan, Zhenghua Hu, Wei Wang, Wei Xiao, and Xuhui Lee. Coregulation of nitrous oxide emissions by nitrogen and temperature in China’s third largest freshwater lake (Lake Taihu). Limnology and Oceanography, 64(3):1070–1086, May 2019. CODEN LIOCAH. ISSN 0024-3590.

Xu:2016:ORH

[XZC^{+}16] Jun Xu, Huan Zhang, Yongjiu Cai, Jorge García Molinos, and Min Zhang. Optimal response to habitat linkage of local fish diversity and mean trophic level. Limnology and Oceanography, 61(4):1438–1448, July 2016. CODEN LIOCAH. ISSN 0024-3590.

Xue:2017:MEM

[XZGW17] Yuejun Xue, Li Zou, Tiantian Ge, and Xuchen Wang. Mobilization and export of millennial-aged organic carbon by the Yellow River. Limnology and Oceanography, 62(S1):S95–S111, November 2017. CODEN LIOCAH. ISSN 0024-3590.

Ye:2019:IWM

[YAC^{+}19] Xinyu Ye, Eric J. Anderson, Philip Y. Chu, Chenfu Huang, and Pengfei Xue. Impact of water mixing and ice formation on the warming of Lake Superior: a model-guided mechanism study. Limnology and Oceanography, 64(2):558–574, March 2019. CO-DEN LIOCAH. ISSN 0024-3590.

Yao:2017:RCS

[YH17] Hongming Yao and Xinping Hu. Responses of carbonate system and CO_{2} flux to extended drought and intense flooding in a semiarid subtropical estuary. Limnology and Oceanography, 62 (S1):S112–S130, November 2017. CODEN LIOCAH. ISSN 0024-

Yamashita:2017:FCG

[YHS^{+}17] Youhei Yamashita, Fuminori Hashihama, Hiroaki Saito, Hideki Fukuda, and Hiroshi Ogawa. Factors controlling the geographical distribution of fluorescent dissolved organic matter in the surface waters of the Pacific Ocean. Limnology and Oceanography, 62(6):2360–2374, November 2017. CODEN LIOCAH. ISSN 0024-3590.

Yates:2019:VDO

[YJO^{+}19] Christopher A. Yates, Penny J. Johnes, Alun T. Owen, Francesca L. Brailsford, Helen C. Glanville, Christopher D. Evans, Miles R. Marshall, David L. Jones, Charlotte E. M. Lloyd, Tim Jickells, and Richard P. Evershed. Variation in dissolved organic matter (DOM) stoichiometry in U.K. freshwaters: Assessing the influence of land cover and soil C : N ratio on DOM composition. Limnology and Oceanography, 64 (6):2328–2340, November 2019. CODEN LIOCAH. ISSN 0024-3590.

Young:2012:QDA

[YKBJL12] A. M. Young, L. Karp-Boss, P. A. Jumars, and E. N. Landis. Quantifying diatom aspirations: Mechanical properties of chainforming species. Limnology and Oceanography, 57(6):1789–1801, November 2012. CODEN LIOCAH. ISSN 0024-3590.

Yamamoto:2015:TAF

[YKT^{+}15] Shoji Yamamoto, Hajime Kayanne, Tatsuki Tokoro, Tomohiro Kuwae, and Atsushi Watanabe. Total alkalinity flux in coral reefs estimated from eddy covariance and sediment pore-water profiles. Limnology and Oceanography, 60(1):229–241, January 2015. CODEN LIOCAH. ISSN 0024-3590.

Yang:2016:PPN

[YLH^{+}16] Shan-Hua Yang, Sonny T. M. Lee, Chang-Rung Huang, Ching-Hung Tseng, Pei-Wen Chiang, Chung-Pin Chen, Hsing-Ju Chen, and Sen-Lin Tang. Prevalence of potential nitrogen-fixing, green sulfur bacteria in the skeleton of reef-building coral Isopora palifera. Limnology and Oceanography, 61(3):1078–1086, May 2016. CODEN LIOCAH. ISSN 0024-3590.

Ye:2011:SPA

[YLJ11] Zhi-Wei Ye, Guan-Nan Liu, and Jian-Guo Jiang. Structural and phylogenetic analysis of a novel ζ-carotene desaturase from Dunaliella bardawil, a unicellular alga that accumulates large amounts of β-carotene. Limnology and Oceanography, 56(1): 133–138, January 2011. CODEN LIOCAH. ISSN 0024-3590.

Young:2018:EST

[YMB^{+}18] Caitlin Young, Jonathan B. Martin, Jackie Branyon, Andrea Pain, Arnoldo Valle-Levinson, Ismael Mariño-Tapia, and Mario Rebolledo Vieyra. Effects of short-term variations in sea level on dissolved oxygen in a coastal karst aquifer, quintana roo, Mexico. Limnology and Oceanography, 63(4):352–362, January 2018. CODEN LIOCAH. ISSN 0024-3590.

Yuan:2018:CRB

[YP18] Lester L. Yuan and Amina I. Pollard. Changes in the relationship between zooplankton and phytoplankton biomasses across a eutrophication gradient. Limnology and Oceanography, 63(6): 2493–2507, November 2018. CODEN LIOCAH. ISSN 0024-3590.

Yang:2017:NAU

[YWL^{+}17] Jiachuan Yang, Zhi-Hua Wang, Qi Li, Nikki Vercauteren, Elie Bou-Zeid, and Marc B. Parlange. A novel approach for unraveling the energy balance of water surfaces with a single depth temperature measurement. Limnology and Oceanography, 62(1): 89–103, January 2017. CODEN LIOCAH. ISSN 0024-3590.

Yu:2015:GEB

[YWY^{+}15] Tsai-Luen Yu, Bo-Shian Wang, Chen-Feng You, George S. Burr, Chuan-Hsiung Chung, and Yue-Gau Chen. Geochemical effects of biomass burning and land degradation on Lanyu Islet, Taiwan. Limnology and Oceanography, 60(2):411–418, March 2015. CODEN LIOCAH. ISSN 0024-3590.

Yokokawa:2013:LSG

[YYMN13] Taichi Yokokawa, Yanhui Yang, Chiaki Motegi, and Toshi Nagata. Large-scale geographical variation in prokaryotic abundance and production in meso- and bathypelagic zones of the central Pacific and Southern Ocean. Limnology and Oceanography, 58(2):61–73, January 2013. CODEN LIOCAH. ISSN 0024-3590.

Zhang:2015:MEL

[ZBSR15] Yong Zhang, Lennart T. Bach, Kai G. Schulz, and Ulf Riebesell. The modulating effect of light intensity on the response of the coccolithophore Gephyrocapsa oceanica to ocean acidification. Limnology and Oceanography, 60(6):2145–2157, November 2015. CODEN LIOCAH. ISSN 0024-3590.

Zwart:2016:MPR

[ZCK^{+}16] Jacob A. Zwart, Nicola Craig, Patrick T. Kelly, Stephen D. Sebestyen, Christopher T. Solomon, Brian C. Weidel, and Stuart E. Jones. Metabolic and physiochemical responses to a whole-lake experimental increase in dissolved organic carbon in a north-temperate lake. Limnology and Oceanography, 61(2): 723–734, March 2016. CODEN LIOCAH. ISSN 0024-3590.

Zhang:2019:NEH

[ZCL^{+}19] You Zhang, Long Cheng, Kuanyi Li, Lu Zhang, Yongjiu Cai, Xiaolong Wang, and Jani Heino. Nutrient enrichment homogenizes taxonomic and functional diversity of benthic macroinvertebrate assemblages in shallow lakes. Limnology and Oceanography, 64 (3):1047–1058, May 2019. CODEN LIOCAH. ISSN 0024-3590.

Zhang:2015:PBC

[ZCY^{+}15] Run Zhang, Min Chen, Qing Yang, Yuanshao Lin, Huabin Mao, Yusheng Qiu, Jinlu Tong, E. Lv, Zhi Yang, Weifeng Yang, and Jianping Cao. Physical-biological coupling of N_{2} fixation in the northwestern South China Sea coastal upwelling during summer. Limnology and Oceanography, 60(4):1411–1425, July 2015. CODEN LIOCAH. ISSN 0024-3590.

Zhan:2018:CUA

[ZCZ^{+}18] L. Zhan, L. Chen, J. Zhang, Y. Li, M. Wu, and J. Liu. Contribution of upwelling to air-sea N_{2}O flux at the tip of the Antarctica Peninsula. Limnology and Oceanography, 63(6):2737–2750, November 2018. CODEN LIOCAH. ISSN 0024-3590.

Zheng:2018:DPD

[ZD18] Guangming Zheng and Paul M. DiGiacomo. Detecting phytoplankton diatom fraction based on the spectral shape of satellite-derived algal light absorption coefficient. Limnology and Oceanography, 63(S1):S85–S98, March 2018. CODEN LIOCAH. ISSN 0024-3590.

Zhang:2015:CER

[ZEXH15] Huan Zhang, Mattias K. Ekvall, Jun Xu, and Lars-Anders Hansson. Counteracting effects of recruitment and predation shape establishment of rotifer communities under climate change. Limnology and Oceanography, 60(5):1577–1587, September 2015. CODEN LIOCAH. ISSN 0024-3590.

Zakem:2017:TBN

[ZF17] E. J. Zakem and M. J. Follows. A theoretical basis for a nanomolar critical oxygen concentration. Limnology and Oceanography, 62(2):795–805, March 2017. CODEN LIOCAH. ISSN 0024-3590.

Zimmer:2016:UCF

[ZHD^{+}16] Kyle D. Zimmer, William O. Hobbs, Leah M. Domine, Brian R. Herwig, Mark A. Hanson, and James B. Cotner. Uniform carbon fluxes in shallow lakes in alternative stable states. Limnology and Oceanography, 61(1):330–340, January 2016. CODEN LIOCAH. ISSN 0024-3590.

Zimmerman:2015:PEO

[ZHG15] Richard C. Zimmerman, Victoria J. Hill, and Charles L. Gallegos. Predicting effects of ocean warming, acidification, and water quality on Chesapeake region eelgrass. Limnology and Oceanography, 60(5):1781–1804, September 2015. CODEN LI-OCAH. ISSN 0024-3590.

Zhang:2010:LTP

[ZHN^{+}10] Jan Zhang, Jeff Hudson, Richard Neal, Jeff Sereda, Thomas Clair, Michael Turner, Dean Jeffries, Peter Dillon, Lewis Molot, Keith Somers, and Ray Hesslein. Long-term patterns of dissolved organic carbon in lakes across eastern Canada: Evidence of a pronounced climate effect. Limnology and Oceanography, 55 (1):30–42, January 2010. CODEN LIOCAH. ISSN 0024-3590.

Zhang:2014:BWP

[ZKL^{+}14] Yong Zhang, Regina Klapper, Kai T. Lohbeck, Lennart T. Bach, Kai G. Schulz, Thorsten B. H. Reusch, and Ulf Riebesell. Between- and within-population variations in thermal reaction norms of the coccolithophore Emiliania huxleyi. Limnology and Oceanography, 59(5):1570–1580, September 2014. CODEN LI-OCAH. ISSN 0024-3590.

Zell:2013:DOB

[ZKMT^{+}13] Claudia Zell, Jung-Hyun Kim, Patricia Moreira-Turcq, Gwenaël Abril, Ellen C. Hopmans, Marie-Paule Bonnet, Rodrigo Lima Sobrinho, and Jaap S. Sinninghe Damsté. Disentangling the origins of branched tetraether lipids and crenarchaeol in the lower Amazon River: Implications for GDGT-based proxies. Limnology and Oceanography, 58(2):343–353, January 2013. CODEN LIOCAH. ISSN 0024-3590.

Zhang:2010:SCD

[ZLLM10] Leiyan Zhang, Kuanyi Li, Zhengwen Liu, and Jack J. Middelburg. Sedimented cyanobacterial detritus as a source of nutrient for submerged macrophytes (Vallisneria spiralis and Elodea nuttallii): an isotope labeling experiment using ^{15}N. Limnology and Oceanography, 55(5):1912–1917, September 2010. CODEN LIOCAH. ISSN 0024-3590.

Zhuang:2018:EPM

[ZMS^{+}18] Guang-Chao Zhuang, Andy Montgomery, Ryan J. Sibert, Mary-Katherine Rogener, Vladimir A. Samarkin, and Samantha B. Joye. Effects of pressure, methane concentration, sulfate reduction activity, and temperature on methane production in surface sediments of the Gulf of Mexico. Limnology and Oceanography, 64(4):2080–2092, September 2018. CODEN LIOCAH. ISSN 0024-3590.

Zigah:2011:RSC

[ZMWM11] Prosper K. Zigah, Elizabeth C. Minor, Josef P. Werne, and S. Leigh McCallister. Radiocarbon and stable carbon isotopic insights into provenance and cycling of carbon in Lake Superior. Limnology and Oceanography, 56(3):867–886, May 2011. CODEN LIOCAH. ISSN 0024-3590.

Zhong:2016:RAW

[ZNVF16] Yafang Zhong, Michael Notaro, Stephen J. Vavrus, and Michael J. Foster. Recent accelerated warming of the Laurentian Great Lakes: Physical drivers. Limnology and Oceanography, 62 (3):1762–1786, September 2016. CODEN LIOCAH. ISSN 0024-3590.

Zigah:2015:MOP

[ZOB^{+}15] Prosper K. Zigah, Kirsten Oswald, Andreas Brand, Christian Dinkel, Bernhard Wehrli, and Carsten J. Schubert. Methane oxidation pathways and associated methanotrophic communities in the water column of a tropical lake. Limnology and Oceanography, 60(2):553–572, March 2015. CODEN LIOCAH. ISSN 0024-3590.

Zingel:2012:CCF

[ZPK^{+}12] Priit Zingel, Tiit Paaver, Katrit Karus, Helen Agasild, and Tiina Nõges. Ciliates as the crucial food source of larval fish in a shallow eutrophic lake. Limnology and Oceanography, 57(4): 1049–1056, July 2012. CODEN LIOCAH. ISSN 0024-3590.

Zacher:2018:FED

[ZS18] Leah Sloan Zacher and Richard R. Strathmann. A field experiment demonstrating risk on the seafloor for planktonic embryos. Limnology and Oceanography, 63(6):2708–2716, November 2018. CODEN LIOCAH. ISSN 0024-3590.

Zhou:2014:NLM

[ZSM14] Yuntao Zhou, Donald Scavia, and Anna M. Michalak. Nutrient loading and meteorological conditions explain interannual variability of hypoxia in Chesapeake Bay. Limnology and Oceanography, 59(2):373–384, March 2014. CODEN LIOCAH. ISSN 0024-3590.

Zimmer:2012:FFM

[ZSZ12] Cheryl Ann Zimmer, Victoria R. Starczak, and Richard K. Zimmer. Flocs, flows, and mechanisms decoupling larval supply from settlement. Limnology and Oceanography, 57(4):936–944, July 2012. CODEN LIOCAH. ISSN 0024-3590.

Zamora-Terol:2013:EFC

[ZTS13] Sara Zamora-Terol and Enric Saiz. Effects of food concentration on egg production and feeding rates of the cyclopoid copepod Oithona davisae. Limnology and Oceanography, 58(3):376–387, January 2013. CODEN LIOCAH. ISSN 0024-3590.

Zhu:2011:DHA

[ZTW^{+}11] Chun Zhu, Helen M. Talbot, Thomas Wagner, Jian-Ming Pan, and Richard D. Pancost. Distribution of hopanoids along a land to sea transect: Implications for microbial ecology and the use of hopanoids in environmental studies. Limnology and Oceanography, 56(5):1850–1865, September 2011. CODEN LIOCAH. ISSN 0024-3590.

Zeller:2014:IPS

[ZWA^{+}14] Robert B. Zeller, Joel S. Weitzman, Morgan E. Abbett, Francisco J. Zarama, Oliver B. Fringer, and Jeffrey R. Koseff. Improved parameterization of seagrass blade dynamics and wave attenuation based on numerical and laboratory experiments. Limnology and Oceanography, 59(1):251–266, January 2014. CODEN LIOCAH. ISSN 0024-3590.

Zhang:2014:TSR

[ZWL^{+}14] Yunlin Zhang, Zhixu Wu, Mingliang Liu, Jianbo He, Kun Shi, Mingzhu Wang, and Zuoming Yu. Thermal structure and response to long-term climatic changes in Lake Qiandaohu, a deep subtropical reservoir in China. Limnology and Oceanography, 59 (4):1193–1202, July 2014. CODEN LIOCAH. ISSN 0024-3590.

Zhuang:2019:BMA

[ZXL^{+}19] Guang-Chao Zhuang, Lei Xu, Qianyong Liang, Xibei Fan, Zhen Xia, Samantha B. Joye, and Fengping Wang. Biogeochemistry, microbial activity, and diversity in surface and subsurface deepsea sediments of South China Sea. Limnology and Oceanography, 66(4):2252–2270, September 2019. CODEN LIOCAH. ISSN 0024-3590.

Zhang:2011:SVF

[ZXM^{+}11] Xiaowen Zhang, Dong Xu, Yuze Mao, Youxun Li, Suyan Xue, Jian Zou, Wei Lian, Chengwei Liang, Zhimeng Zhuang, Qingyin Wang, and Naihao Ye. Settlement of vegetative fragments of Ulva prolifera confirmed as an important seed source for succession of a large-scale green tide bloom. Limnology and Oceanography, 56(1):233–242, January 2011. CODEN LIOCAH. ISSN 0024-3590.

Zafiriou:2012:EDC

[ZXN^{+}12] Oliver C. Zafiriou, Huixiang Xie, Norman B. Nelson, Raymond G. Najjar, and Wei Wang. Erratum: Diel carbon monoxide cycling in the upper Sargasso Sea near Bermuda at the onset of spring and in midsummer. Limnology and Oceanography, 57 (3):895, May 2012. CODEN LIOCAH. ISSN 0024-3590.

Zou:2017:III

[ZXZ17a] Youjia Zou, Xiangying Xi, and Chaoyang Zhang. Issue information — instr to contrib. Limnology and Oceanography, 62 (4):2339–2341, September 2017. CODEN LIOCAH. ISSN 0024-3590.

Zou:2017:RSM

[ZXZ17b] Youjia Zou, Xiangying Xi, and Chaoyang Zhang. Retracted: Southward migrations of the Atlantic Equatorial Currents during the Younger Dryas. Limnology and Oceanography, 62(4): 1732–1741, July 2017. CODEN LIOCAH. ISSN 0024-3590. See retraction notice [Ano17l].

Zhu:2019:TSD

[ZYZ19] Rong Zhu, Gui-Peng Yang, and Hong-Hai Zhang. Temporal and spatial distributions of carbonyl sulfide, dimethyl sulfide, and carbon disulfide in seawater and marine atmosphere of the Changjiang Estuary and its adjacent East China Sea. Limnology and Oceanography, 64(2):632–649, March 2019. CODEN LIOCAH. ISSN 0024-3590.

Ziegelgruber:2013:SCS

[ZZAC13] Kate L. Ziegelgruber, Teng Zeng, William A. Arnold, and Yu-Ping Chin. Sources and composition of sediment pore-water dissolved organic matter in prairie pothole lakes. Limnology and Oceanography, 58(3):1136–1146, May 2013. CODEN LIOCAH. ISSN 0024-3590.

Zhonghua:2016:PAH

[ZZW16] Zhao Zhonghua, Lu Zhang, and Jinglu Wu. Polycyclic aromatic hydrocarbons (PAHs) and organochlorine pesticides (OCPs) in sediments from lakes along the middle-lower reaches of the Yangtze River and the Huaihe River of China. Limnology and Oceanography, 61(1):47–60, January 2016. CODEN LIOCAH. ISSN 0024-3590.

Zhang:2010:CSC

[ZZY^{+}10] Yunlin Zhang, Enlou Zhang, Yan Yin, Mark A. van Dijk, Longqing Feng, Zhiqiang Shi, Mingliang Liu, and Boqiang Qina. Characteristics and sources of chromophoric dissolved organic matter in lakes of the Yungui Plateau, China, differing in trophic state and altitude. Limnology and Oceanography, 55(6):2645– 2659, November 2010. CODEN LIOCAH. ISSN 0024-3590.